A Complete Bibliography of Publications in Annals of Mathematical Statistics

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

08 November 2023 Version 2.10

Title word cross-reference

(1001) → (0110) [Whi61a]. (α, ξ) [Kun67]. (d) [Dav65]. (n ≥ 2) [Sen55]. +∞ [Bau63a]. 0 [QW71]. 0 < p < 2 [Cho71]. 1 [Gov62, Her67b, Her70, Jag63, QW71]. 1024 [DM70a]. 11 ≤ n ≤ 15 [SG58]. 1 ≤ r ≤ 2 [vBE65]. 2 [Ala68, Har69b, Her67b, Her70, Sea61]. 256 [DM68b]. 2_{R}^{k−p} [DM67]. 2^{m} [Pat62, SC71a]. 2^{n} [BC57, RF70a, PF72]. 2_{n−r} [Joh66]. 2 × 2 [HK64b, Nic58]. 3 [Her67b, Her70, Mah63, Yos49]. 4 × 4, 5 × 5 [Fre66d]. 50 [Gna67]. 512 [DM68b, DM70a]. 6 × 6 [Fre66d]. 7 × 7 [Sad51]. [0, 1] [Pit68].

∗ [HK65a]. α [RS67, Sob67]. B [Kul48, Sug67a]. b = v [Rao65]. β [Gut70b]. β_{1} [Mer33]. β_{2} [Mer33]. c [Pur64, Tam71]. C[0, ∞) [Whi70]. χ^{2} [CL54, Coc52, Gol63a, GL46, HC58, JB61, Rob33, Wis56]. D [Egy59, Kie62, Wyn70, GC64]. D[0, 1] [Set65]. D_{n}^{+} [BP58, Dem59]. d_{2} [Ros63b]. E(|X_{1}|) = ∞ [Wil70]. E^{2} [Hsu45d]. E_{n} [DeH70, GWA72]. E_{p} (S_{t}/t) [CS68]. η [Fel32]. exp(−e^{−x}) [BD72]. F [Aro41b, Bra52, Dav67, Fox56, Gau59, GG58, Gut70a, HS70a, Her59, Lau60, Oga63, Rag65a, Sug67a].

\frac{1}{y} \frac{dy}{dx} = \frac{a_{0}+a_{1}x}{b_{0}+b_{1}x+b_{2}x^{2}} ≡ \frac{N}{D} [Bea37]. Γ [RH71, Gut70b, Mon72b]. ≥ 5 [DM68b, DM70a]. ≥ 6 [DM68b, DM70a]. i [Lau63, MM69]. ∞ [Bau63b, Bro70c, Far68b]. K [Bas67b, Kie59, Qua66, Sim67a, Bah50, BG46, Con65a, Con65b, Con67b, Dav66, Goo52, Gut70b, Her67c, Ken71, Lor72, MT65, Mos48a, MT50, Pau52b, Pau63, Pau64b, RSS68, Sch69d, Stu67, Tra68, Tru53, ZD59, uD54]. K_{10} [uD54]. k_{11} [ZD59]. k_{9} [uD54]. k ≥ 3 [SW72a]. L [HF70, Maz67, Rak67, cS67, Wol71c, Wol71a, Wol71d]. L^{1} [Gun68]. L^{p} [Cha69]. L_{1} [Bro39, Moo39, OS70b]. L_{2} [Rag60b]. L_{p} [Cho71]. λ_{2} = 3 [Agg72]. M [Goo59, Chu41, Fri40, Goo58a, Goo58b, KS39, Mar48, New64, Sen68a, Wat54]. D_{n}^{+} [Dwa59]. L_{2} [Shr59b]. m × n [Van66d, Van66c]. N [Cam48, Cra31a, O’T34, Ski68, Sta56, BJK56, Bue65, Chu41, Fer36, GD55, Hsu48, Jee54, Joh67d, Mau40, Pie40, Rao66, Roy53b, Sen55, WF57a]. n^{−1/p}S_{n} [Cho71]. N^{th} [LL39]. n_{1} = 12 [Sei66]. n_{2} = 15 [Sei66]. nω_{n}^{2} [Mar58]. O(n, n − 1) [HF70]. ω_{n} [She57]. P [Bic64, Bic69a, Con67d, Far66a, Far66b, Sen72a]. p_{11}^{2} = 4 [Sei66]. p ≥ 1 [Sen72a]. ϕ [Sen71c]. Pr{Y < X} [BM58a]. Q [HPW64a]. R [HS72b, O’T34, Ban60, Dwy40a, PR68, Tai66, Wal50a, vBE65]. R^{2} [GK64b, Gir65a, KS72b]. r × c [Fie68]. s [Gna67]. s = 1 [Gna67]. S_{1}S_{2}^{−1} [Cha70a, Kha67b]. S_{n} [Ser70a]. S_{n}/n [MS70a, SSF68, TBO71, TO72]. S_{n}/n^{\frac{1}{2}} [OS69, RM71]. Σ [Boe69, Coh70, Day64, Day65, BF64c, Bru65, Dan40, Dyk70a, LR72b, Nev72, Rob66, Rob67, Vil64b]. T [DK58, Hsu38, Bra52, BC64, Bro67, BF63b, Che65a, Cra41b, Dal46, DK56, Dav67, Dic67a, Dic67b, Dun61, Dun65, GL46, Gup63b, HL56, HL68, HPW64a, HPW64b, Kap71, Kru54, Lau60, MC65, Mil68d, Ord68, Par64, Pil51, PW63, Res62, Rob48c, Sac65, Sch43b, Sid67a, Sid67b, Sob67, Sto63, Wal47d, Wal47e, Wal48c, Wal49e, Wel64a, Whi57b, Whi58a]. t, τ, [Chu56]. T^{2} [Bic65a, Bic65b, GKS63, Gir65a, Hsu45d, JB61, Ste56, KS72b]. T^{2}−, R^{2}− [KS65]. T_{0}^{2} [Ito56]. T_{0}^{2} [Con66a, Dav68a, Ito60, Mui72a, Sio71]. T_{1} [Ajg67]. U [Arv69, Arv70, MS72b, Sho66, Whi51]. U^{2} [Bur64b]. v = 21, b = 35, r = 10, k = 6, λ_{1} = 2 [Agg72]. v = 28 [Sei66]. V^{(s)} [Pil68]. V^{(}s)_{1} [Gna67]. V_{NM} [MS68a]. ε [Pit68, Ros71c, Tay70, Wol51]. W [Con67d, AM67]. W^{2} [Bur64b]. W^{∗} [MO70]. X [BM58a]. X^{(n)} [Jos39]. xy [Cra36b]. Y [BM58a]. Z [BK70, Coc40b, McC39, Aro41b, Wis57]. Z^{∗} [BK70]. Z^{γ} [BK70]. |S_{n}| ≥ cn^{\frac{1}{2}} [Woo70c].

-Algebra [Coh70]. -Algebras [Nev72, Vil64b]. -bounded [Gun68]. -Complete [Wol51, Boe69]. -Content [Gut70b]. -Convergence [Cho71, Cha69]. -Decision [Lor72]. -Dependent [New64, Sen68a, Tai66, Wat54]. -Dimensional [Ken71, HS72b, Sta56]. -Dimensions [Jee54]. -Distribution [PW63, Res62, Sch43b, Wal48c, Wis57]. -distributions [Mil68d]. -Entropy [Pit68]. -Expansions [Gut70a].

-Factorial [PF72]. -Field [BF64c]. -Fields [LR72b]. -Lattice [Bru65, Dyk70a, Rob67]. -Lattices [Rob66]. -Mean [PR68]. -Minimax [RH71]. -Mixing [HK65a, Sen71c]. -Optimal [Ros71c, Tay70]. -Optimality [Kie62]. -Optimum [Wyn70]. -Point [Far66a, Far66b]. -Quantile [RS67]. -Quantiles [Sob67]. -Ratio [Kap71]. -Restrictional [HF70, Maz67, Rak67]. -Run [DM68b, DM70a]. -Sample [Con67b, Dav66, MT65, Pau52b, Sch69d, MT50, Bas67b, Kie59, Qua66, Con65a, Con65b, Mos48a, Pur64, Tam71]. -Sequences [AM67]. -Sphere [Yos49]. -Square [HS70a]. -Squares [HS70a]. -Statistic [Egy59, Oga63, Sho66, Sio71, ZD59]. -Statistics [Arv70, Tra68, uD54, MS72b, Arv69]. -Subprocess [Kun67]. -Test [Dal46, DK56, DK58, Gau59, HL56, Wal47d, Wal47e, Wal49e, Whi57b, Whi58a, Bra52, GK64b, Her59, Hsu45d, McC39, Rag65a, Ste56, Fox56, HL68, Sac65]. -Tests [Gir65a, JB61]. -th [Goo58a, Goo58b, MM69]. -Transformations [GWA72]. -Variate [Bic69a, Con67d, Gut70b, Mau40, Bic64]. -Variation [Mon72b]. -Way [Ban60, Har69b, Sea61, Mah63].

1 [Hea61b, Igl72, Pak71, Sch70b]. 1949 [Fis50a]. 1950 [Arn51a]. 1951 [Fis52a]. 1953 [Ano53e]. 1954 [Ano54g]. 1971 [Kes72]. 1977 [Efr79].

2 [Che66].

A. [Cra62, Get67a, Hot44a, MR64, Sch64]. A.P.O. [BY69]. Above [Mus68b]. Abraham [BB55, Ano52j, Gir48, Men52, Tin52, Wol52a]. Absence [Ylv68a]. Absolute [Bel47, Dun70, Eri71b, Fri72, KZ71, Kat60, Kni65a, PR70, Sam41, Sin40, Wen61, vBE65]. Absolutely [Bla57a, Far64b, Koo61, Mor52, Sto66b]. Absorbing [Kac45a, Wat61, Wee61]. Absorption [MD71]. Abstract [AJ65, Ano62h, Geb68, Gup62, KA65, Kum65, Sch61a, Sin66, Sri65a, Sri69, Tuk58a]. Abstracts [Ano39a, Ano40a, Ano40b, Ano41a, Ano41b, Ano42a, Ano42b, Ano43a, Ano44a, Ano45a, Ano46a, Ano47a, Ano47b, Ano47c, Ano48a, Ano48b, Ano48c, Ano49a, Ano49b, Ano49c, Ano49d, Ano50a, Ano50b, Ano50c, Ano50d, Ano51a, Ano51b, Ano51c, Ano51d, Ano52a, Ano52b, Ano52c, Ano52d, Ano53a, Ano53b, Ano53c, Ano54a, Ano54b, Ano54c, Ano54d, Ano55a, Ano55b, Ano55c, Ano56a, Ano56b, Ano56c, Ano56d, Ano57a, Ano57b, Ano57c, Ano58a, Ano58b, Ano58c, Ano58d, Ano59a, Ano59b, Ano59c, Ano59d, Ano60a, Ano60b, Ano60c, Ano60d, Ano61a, Ano61b, Ano61c, Ano61d, Ano62a, Ano62b, Ano62c, Ano62d, Ano63a, Ano63b, Ano63c, Ano63d, Ano64a, Ano64b, Ano64c, Ano64d, Ano65a, Ano65b, Ano65c, Ano65d, Ano65e, Ano65f, Ano66a, Ano66b, Ano66c, Ano66d]. Abstracts [Ano66e, Ano66f, Ano67d, Ano67e, Ano67f, Ano67a, Ano67b, Ano67c, Ano68a, Ano68b, Ano68c, Ano68d, Ano68e, Ano68f, Ano69a, Ano69b, Ano69c, Ano69d, Ano69e, Ano69f, Ano70a, Ano70b, Ano70c, Ano70d, Ano70e, Ano70f, Ano71a, Ano71b, Ano71c, Ano71d, Ano71e, Ano71f, Gup62, Sch61a]. Accelerated [Kes58]. Acceptance [GJ59]. Accidents [Bat55]. According [Jor32, Jor59]. Accumulation [Ble40]. Accuracy [Blo71, BS70, Coc46, Hoe43a, WS34]. Accurate [Alb56]. Achieve [Ahl71]. Acknowledgement [Ber73, Ste60]. Acknowledgment [Bau63a, BD65a, Bir65, God61a, Kin61b, KS72c, KA64a, MM63, Mik60a, PL71, Pra66b, Rao59, Rob47, Sta59, SN70, Whi58a]. Action [Fur72, JV71, JV72, Kak71, Kar57b, Lip71]. Adaptive [Zac70b]. Added [Pit63]. Addenda [BK58b, CW59]. Addendum [AG70, Ano54e, Bre69, Mas72]. Adding [Pit64]. Additional [KS66b]. Additions [TW79]. Additive [Get65, Get67b, HS72b, KS68, Luc67, Ols68, Pat63, PJ70, Sud69a, Wol42]. Additively [Cro58]. Adequacy [TW34]. Adequate [Ski67]. ADF [Kou70a]. Adjacent [CMPT60]. Adjusted [Ksh59b, SA57]. Adjusting [Ste42]. Adjustment [Bar51, DS40, Kav41, SM50, Thi31l, Thi31m, Thi31n]. Adjustments [Ano36a, Pie40]. Admissibility [Bly51, Bro66, BW63, Coh68, Far68b, Far68a, Far69a, FR64, Fox71, GJ65, God69, Han68, Jos65a, Jos65b, Jos66a, Jos66b, Jos68a, Jos69a, Jos70, Jos71b, Jos71c, Jos71d, Kar57b, Kar58a, LS53, MR66, Ste55, Ste56, Ste59b, Sto67, SC71b, Tsa68, Zid70, Gho64a]. Admissible [Bir54, Bro71d, Cac66, Coh66, Cop32, Eri70b, Joh71, Jos69b, Jos71f, Kat61b, KS65, KS72b, Leh47a, Mee72b, Pra58b, Rob58, Sch67c, SV69b, Wal47a]. Admitting [Pit57]. Adoption [Dwy48a]. Advanced [Bar47b, Kie64, Wat71b, Cup69a]. Advances [Cam42, Cra42, Dal62]. AEP [Tho60]. After [Hoy68b]. Against [BB69, Cha63a, CLA62, Gna59, Gna60, Høy68a, Kri68, Mik62, SS66b, Sho67, Ter52, Wei53b]. Age [Ath69, Dur71, Kur70, Wei65a, Wei65b, Wei66, Wei72a]. Age-Dependent [Dur71, Wei72a]. Aggregates [Lot39a]. Ahead [Ros71b]. Alan [Kie64]. Albert [Fab69a]. Alexander [Wel64a]. Alfréd [RV72, RS72a]. Algebra [Coh70, Jam57, Man60, OI65]. Algebraic [Dre40, Hog51, Hol68, Kal65, Sat41]. Algebras [Boe69, BM59b, Fis72b, Nev72, Rob70b, Rob71b, Vil64b]. Algorithm [BW64b, Gut70a, Rob71a, Wah71b]. Aligned [Sen68b, Sen71a, Sen72b]. Allocation [Dem60, Dem61, Elf52, SF39]. Almost [BB68b, Ber70c, CT71, FK51, Gil71b, Gil72, GHRM71, Her63, KsW69, Smi68a, Sto68c, Var68, Wic70]. Alms [Hal44]. Alphabets [BK60]. Alternative [Bic64, Ell62, Fis72a, Fri40, Klo65, Kui59, Leh63a, Pan67, Rao54, Sch68a, Sch56a, Woi72]. Alternatives [Arn65, Ber70a, BJ70, Cha63a, CLA62, Dix53, Dok69a, DH69, Eat70a, Gna59, Gna60, Háj68, HG66, Hol67, Kri68, Mik62, Nüe66, PS68a, Ram66b, RG62, SS66a, SS67, SS66b, Set70b, Sho66, Sho67, Str72d, Ter52, SE43]. Ambiguity [Con70]. Amendments [Ano44b]. Amiel [Blu67a]. Among [Bro45, Chu41, Dav62, Dav63b, EP72, Goo64b, Goo76, Mar48, Nor35, Sea56a, Set50, SF39, Var60, Con52a]. Analog [Bha63, Mar60]. Analogous [Dia63, Gre32]. Analogue [Dal46, Ger71, Pil51, RSS67, Wal47d, vE72]. Analogues [Gir65a, Kie59, MP67, RP58]. Analyses [FM58, Ney41]. Analysis [Alb47, Alb48, And63a, And54, Bah58, BF68a, BW57a, Bar51, BPSW70, Bha62, Bha63, Bis39, Bla57b, Bol34, Box54a, Box54b, BBH56, BK58b, Bra65, Coc40a, Coh68, Con63b, Cur43, DJ52, DA59, Dem64, Dia63, Dix50, Ehr70, Far69b, Fle62a, Fle62b, Fra57c, Gab62, Gav59, Gei42b, GG58, Gei65, Gir46a, Gir46b, GW70, Goo63c, GR53, Gru36, HWG65, HWG95, Han66a, Haq70, Har47, Har44, Har45, Hen35a, Her59, HL62, Hoe37, HK44, Ifr65b, Ifr65a, Jam54b, Joh64b, Joh53, Jon62, Kab68, Kem64, Ken53, Kha65c, KP68b, Kim51, Koo64b, KB57, Kul35b, Kul52, Kul56a, Kul56b, LB63, Leh63c, Lig71, Lin64, Lin60, Lot39b, LR53, Mad37, Mad48a, MO70, Mij61, Mud65, Mud66, MD68, Nic54, OI65, Par61, Pau50, Pau42a, Pau49, Per69a, Per71]. Analysis [Pha63, Pil55, Pil56, Pil67, PS69b, PP69, Qua66, Ram56a, Rao61b, Rao68c, Rei40, Roy53a, Roy58, RR59, RC60b, RC60c, Rub68, Sat41, Sat42b, Sch56a, Sch56b, Sch65, Sch34, Sch67c, Sea56b, Seb64b, Sha70c, Sho67, Spj68b, Sta39, Sug65, Sug66, Sug67a, Tro71, Tuk47b, Tuk62a, Vaj51, Wah68, Wal40b, Wal41b, Wal42b, Wal46a, Wal47b, Wij57, Wij59c, WK56, WF57b, Wol67a, AT63, Goo52, Tuk62b, Gir48]. Analytic [CL70, CL71, Hen34, Lah58]. anatomy [Tuk57a]. Ancillary [San68, Spr61]. Anderson [Lew61, Wol61]. Angular [Ber67b, Tha63, Tha65, Win47, FT50]. Anhang [Sch64]. Animal [CJ56]. Annals [And51b, And52, Ano35a, Cam38a, Kin30, Wil50]. Anniversary [Cra60]. Announcement [Ano39b, Ano41c]. Annual [Ano46b, Ano46c, Ano46d, Arn51a, Coc47, Cra37, Cra38b, Cra39a, Cra43c, Cra44a, Dwy45b, Fis50a, Fis52a, Kim42, Old42a, Old43b, Old44, Rid40, She45, Wil47]. Anova [Bha60, LaM72, RG59a, RG59b]. Ante [Gab62]. Ante-dependence [Gab62]. Any [Dwy37a, Hen34, Sha65a, Wil36, Bra65, HS49b, Mor52]. Appendix [Kem58, Sch64]. Applicable [Kni65c, Ney39]. Application [Ail70, AJ58, BS70, Bre57c, Bro59, Bro70d, Bro71f, CH45, Cle71, Cra41b, Der56a, Dwy38a, EM70, Fer58, Gho71a, HS49a, Har52, Har70a, Har69b, Hen31, Hen32b, Hen35a, Hey66, Jog67, Joh53, Kul34, Kul52, Kul56a, Kul56b, Kus69, Law33, Lot48, MW42, McC39, MG69, Mos41, Mus67, Nor40, Pha63, Pie40, Por71, Pra60a, Rak67, RC62, Rei40, RP49, Rog64, RO72, See70b, Sto59, Suk72, Tak64, Var55, Wal58b, Wil36, Won35]. Applications [Agr66b, Bar65, Bas68, BW71, Bir55, BG46, Bru65, BM53, CGJ67, Dic68b, Dwy38c, Eps48, EPW67, Fis62, Gho54, GMS46, GS55, HWG65, HWG95, Har59a, Hoo56a, Hsu44, Ifr65b, Iye58, Iye61a, Jen70a, Jen70b, Kab65, Kha67a, KW56b, Kot65, Kou68, MR65, Mil61b, Mor53, Noe70, OS70b, PB70, Rao62, Rob37a, Rub56, Rus57, Sau63b, Sen71b, She55, Sho72a, Sri62, SC71a, SA71, Str66b, Tho38, Wic69, Wij59a, Wol43, ZF64, da 45, Ben63, Gra67, Spi67, Tuk47b, Ylv68b]. Applied [Box54a, Box54b, BK70, Had63, Kim40, Kim42, Rid61]. Approach [Ban60, BJ68a, Bur32, Dan39, Don52, Doo49, Eat70b, GR52, Gre50a, Gut46a, GT64b, Ham72, Kem68b, Kem48, Kru68, Leh63a, Par61, PS68c, Rob64a, Sam63b, Wil67, Wij68b]. Approachability [fH69, fH71, Sac72]. Appropriate [Bro71e]. Approximate [Cam64, Chu46a, Coc40b, Dem61, Ekm59a, Ela62, GL46, Hsu45b, Hsu45c, Koz56, McC47, MS71, Obe47, Pau42a, Que49a, Ros62b, Ros62c, Ros63a, Tuk48a, WS34, Whi57a, Wis57]. Approximately [Mee71, SC72]. Approximating [DG57b]. Approximation [AG70, Aro43, Bla55a, Blo57, Blu54a, Blu54b, Bur56, Bur59, Cam51, Chu54, Com64, Cra48, Dav33, Der56a, Der56b, DS59, Dud69a, Dup65, Ekm59b, Eze30, Fab67, Fab68a, Fab68b, Fab69a, Fab69b, Fel45b, Geb68, Gho64b, Gil69, Gre32, Ham72, HH63, Hen36, HL60, HPW64a, Jor32, Jor59, Kal54, Kes58, Kim49, Kur70, Kus66, Nic56, O’T33b, Oat72, Ric66, RM51, Sac58, Sar49, Sch69g, Smi39, TW46, Ven66, Ven67b, Wil46b, Wol52b, Wol56, Dup70]. Approximations [Bac63, Bah60a, Bak40, Bir62, BS70, Chu56, Dan54b, Fri63, Jag71, Kem70, KP66a, KP68a, PM59, Ram72, RP63, Rei63b, Sid62a, Sid65, Tsa57, Wal58a, Wal59a, Wal60, War60, Yar72]. April [Kul46]. Arbitrarily [Ahl70, Gei38, Gei39, Kim44]. Arbitrary [Chu41, CH45, Enn70, Enn73, Gei42a, Jos71b, KS68, Lip71, MM69, Ros68a, Sin64]. Arc [Ber62a, Hor71, Imh68]. Arc-Sine [Hor71]. Arcsine [Bax62]. Arcwise [Den70]. Area [Gor41b]. Areas [PW63]. Argument [Dem63a, Gor41b, Mui70b, Sha70a, vS50]. Arises [GS57]. Arising [Bac48, Bar51, DV67, Dod37, DS69b, GR53, Hoe69, Lev49b, MG69, Ney41, Nor66, Sch43a, Wal44b]. Arithmetic [Agg66, Gur55b, Gur59, Kir35, Lar39, Pol34, Wer37, Zoc35a]. Armed [FV70, Fel62, Sam68b, SP65, Vog60a, Vog60b, Ber72f]. Arrangements [KZ62, Rag62, Thi31e, Zel58, ZF64, Olm46]. Array [Hoo56b, Sei55a, Sei55b, Sei56]. Arrays [AK61, BB52, Bus52b, Cha61, Sei54, SZ66]. Arrival [Fre62c]. Arrivals [Ben57, Pra60a]. Arrow [Had63]. Arthur [Dav63a, Fab69a]. Article [Gre50a]. Ascent [JF64]. Ash [Wol67b]. Ashton [Wan68]. Aspects [BM65, Rag60c, Rub62b, Wat62, Wat63, WK56]. Assay [KvE64a, Tsu67]. Assays [Bli46]. Assets [Mor32]. Assigned [AJ51, Zac68]. Associate [BK58b, Kag72, KB57, Rag60b, Rao61b, Sei61]. Associated [BS68, Blu63, BK60, BBFH69, Dwy37b, Dwy49b, Fos53, Gei67, IW71, Joh67d, Joh70c, KP57, Kei72, MH62, Mud66, Pau41, Ros52a, RP58, RB58, SR68, Sid62b, Tho38, War60, WD39, Wol49b]. Association [Aig71, Ano60i, BM59b, Con58, EPW67, Hof60, Joh67b, Joh70a, Joh70b, Mes65, Rag60b, Shr59a, Shr59b, Sil64, SA71, Sur68a, Sur71, OT47]. Associative [BM59b]. Assuming [Loh69, Sta38]. Assumption [HP36]. Assumptions [CR51, FG52, Le 70a]. Asymmetric [Jae71a, Mij72, Mil72a, PS69a, Por70, Tor52]. Asymmetrical [Zel58]. Asymptotes [Eli70]. Asymptotic [AGW71, Adi67a, Agn59, Ail70, And67, AR50, AD52, And59, And63a, AW64, And65a, And70, And54, AJ65, Bah64, BP67b, BK63, BB65, Beh71, Beh72, Ber69a, Ber72c, Ber73, Ber71d, Bha61d, Bha71b, Bha67b, BH67, Bil56, Bin72, Blu63, Bro71e, Cap61, CLA62, Cha63b, Cha70a, Cha70b, Che49, Che52, CS58, CGJ67, CR65b, Dan71, DT71, DH69, DKW56, Dwa53, Dwa55b, Eic63a, Fab67, Fab68a, Fab64, Far64a, Far72, Fel51, Fra68a, Fre63b, Fre65b, GW68, Gho54, GK64a, Gle65, Gle66b, Goo61a, Goo58a, Goo58b, Gui71, GG51, Háj64, Háj68, Hal60, HR55a, HM67, Hil69b, HD68, Hoa71, Hoe61, Hög72, Hsu46, Hsu48, Ifr65b, Ito56, Ito60, Jen69, Jog68a, Joh67d, Joh67c, Joh70c, Jur69, KM71, KS62, Kap45, Kem59, KW59a, Kim44, Kim50, Koo60, Kou69, Kou70b]. Asymptotic [Kou71, Kou72, KP70, Lad55, Le 70a, sL71, LPC70, Lig71, Mal72, Mar72a, McC69, McC64, MS67, Moo54, Moo68b, Mui70a, Mui72a, Nea70, New64, Noe50, Oka63, Oka68, Pei43, Pei49, Per69b, Pfa70, Pur64, Qua65, QW71, QW72, Rea72, Ros72a, Ros72b, Rou65, RSvZ72, Sac58, SA66, Sam63a, Sch69b, Sch72, Sch62, Sch68b, Sen66, Sen68a, SP70, Sen71a, Sev57, Sho69, Sie68, Sin70, Sin71, Sio71, SHB64, Sta66a, Str72d, Sug69, SF69, Tsu67, Tsu68, Van66a, Vil66, Vog60a, Vor72, Wal48a, Wal58a, Wei65a, Wei57b, WK69, Wol44a, Wyn69, Yar72, Zei50, Van63, vE72, vM47]. Asymptotically [Bic65a, Bic65b, BY68, Boz71, Chi56, Dok66, Dok67a, Dok68, Fer58, Gas65, Gir68b, Gup67, Háj62, Har67a, Har65b, Her67a, Hoe48a, Hoe65a, Hol66, Hol71a, JR69, JR70, KW72, KS63, Leh63a, Leh64, Lor67, Mac59, Moo69, Rob63, Rob64b, Sen67, Str71b, Str72c, Tay53, Wal41a, Wal41c, Wal42a, Won68]. Atlantic [Dwy47a]. Atomic [Bre58b]. Attachment [Sri68d]. Attaining [GG39]. Attainment [Fen59]. Attracted [Bel70a, Bel70b, Hey67, Tuc68]. Attraction [BD72, Roo72]. Attributes [Jon52]. Authorship [Wat66]. Autocorrelated [Gre54b]. Autocorrelation [Ban65, LZ57, Jag63]. Autocorrelations [AW64]. Automorphism [HF70, NP66]. Autoregressive [And71a, And71b, And72a, Fra63, Haq70, Hil69b, Pie70, Sid58c]. Auxiliary [Gol63a, Sch70a, Sen72a, Sri69, Thi31o]. Average [Car34a, CT56, Der66a, Gru33, Hoe53a, Hoe60, Jen32, JM63, Lip71, Pie70, Ros71c, SR68, Sim67b, Wil38c]. Average-Reward [Lip71]. Averages [Bru71, Cra33a, Dok67a, Dok68, Nor35, She64a, Tom71]. Averaging [KS70, KS72c]. Axiomatizations [Fis72a]. Axioms [Lee64].

B [Get66, Gre64, Gre50a, MR64, Pea64, Pyk63, Sal67, Wel64a, Aro37, Kul47]. B. [Cup69a, Gre64]. Bacteria [Sri68d]. Bacterial [Fer58]. Bacteriology [Ney39]. Bahadur [Abr67, Gho71a, Gle64, JK62, Kie67, KH72, Leh66a, LF70, Meh72, Rag70a, Rao72, Woo70d]. Bailey [Str66b]. Balakrishnan [Ylv61]. Balance [And72b, HST53]. Balanced [Agr66a, AJ56, Ban48, Bha60, Bos49, BC55, BM59b, Car68, Cha61, Con52b, Con53, CC54, Cor60, Cox40, DN62, Far68b, Gau59, GW59b, Han61, Han65, Har69b, HF70, HSF72, Her59, Hin64, Hof63, Joh64a, Joh67b, Kag71, KP69, KDS72, Law71, Maj62b, Man69a, Maz67, Mes67, Mik60a, Mik60b, Mot58, Mur61b, Nai69, Oga59, Oga63, OI65, Rag66, Rak67, Rao58b, Rao60, RL57, Sei61, Sei66, Ses66b, Sha59b, Sha60a, Sha60b, Sha66b, Sha67, Shr50, Shr65, Spr55, SA70, SC71a, SA71, SM66, Sur68b, Sur71, Tho56, Yal68, Zel54, Zys63, Tuk57e, Tuk56b]. Balancing [Sha58a, Sha59a]. Ball [DL66]. Ballot [Tak64]. BAN [Wij59b, Kat50, Wij59d]. Banach [PS72]. Banach-Valued [PS72]. Bandit [Ber72f, FV70, Fel62, Sam68b, SP65, Vog60a, Vog60b]. Bands [Hoe54]. Bark [Owe67b]. Barlow [McC65]. Barrier [Leh63d, Wee61]. Barriers [Kac45a]. Bartlett [SN70, Ksh59a, SN69]. Barton [Rio62b]. Based [Adi67b, BP66b, BP67a, BP69, Ber70d, BD69, BM58a, BH60a, Blo51, BKR61, Bur72a, Che52, Col32, Dem66, Ell62, ET53, FP37, Gee70, Ges70, Goo63c, GS58b, HL63, Hoe52, KKW55, KS62, KB71, Kha65c, Kha66, Kim49, Kou71, KS72d, Lev52, Mac59, Mad48b, Man45c, Mat55, PS69b, PP69, PR70, RK57, Rao54, RW70, Ros63b, Sal66, SS66a, SS67, Sch43b, Sch67b, Sen68c, SG71, Set70b, Sta72b, Tam69, TK70, WW44, Wal46e, Wal50a, Weg56, Wei57b, Wei61, Wij67, AD52, Fis67b, Tam71, WW43]. Basic [Bel64b, Kha65a, Kha65b, Kim47, McM53, SA70]. Basis [Kul63, Ric66]. batch [FN61, FP64]. Baye [Ber30, Mol31, Spr61, vM38, vM42]. Bayes [KS72b, Agg59, Agg66, Ams63, And64b, Bha64, BB67, BY68, Cha70b, CR65a, DR63, Dun61, Fab64, Far68c, Far69b, Fra72a, Fre63b, Fre65b, FP69b, Gil69, GR52, GJ65, God69, GJ59, Joh57b, JV71, JV72, Jos65a, Jos65b, Jos66a, Ker63, KS65, Lor67, Mee72a, Mee72b, Oat72, PB71, Por71, Ray65, Rob64a, Sac63, Sam63b, Sca59, Sch69f, Sch62, Ski60, Sto67, Str71a, WW50, Wij70, Zac68, Zac70a, Zid70]. Bayes-Fiducial [PB71]. Bayesian [Che66, Bha68, BW67b, BT65, Dic68b, Fel72, Fer72b, Gei65, GT64b, KW70, KvE64a, Lin64, Rol68, Shu69, Vil72]. Be [Pol34, Bar69, Bri63a, Dha63b, Min68, Par58, Pit63, Sam41]. Bearing [Bri62]. Bechhofer [Gup72, Hal59]. Become [Dod41b]. Been [Fer36, Len39]. Behavior [And54, Ber66a, Ber66b, BF68c, Boa67, BE70, Cha70b, Dan71, Fab64, Far64a, Fin62, Fre63b, Fre65b, Har67a, Ifr65b, KM71, Kou69, Kou71, Lad55, McC69, Pit65, QW71, Sev57, Str72d, Tak61c, Hur48]. Behaviour [BN63, Dal69b, Fin61]. Behrens [Ban67, EO72, Fra53c, MB66, Pot63, Ram66a, RSS67, Sch43b, Sch44, Wal49e]. Bell [Tor52, Jog66]. Bell-Shaped [Tor52]. Bellman [Ros66a]. Below [Eic70]. Beltrami [Jam68]. Benes [Rei64]. Bergström [Dar65b]. Berkeley [Ano48j, Ano48k, Hod49, Hod50]. Berkson [Tuk58b]. Bernard [Fre65a]. Bernoulli [Ber72f, Eat70c, FS71, Lar36]. Bernoullian [Ste45b]. Bernstein [BKR71, Cra33b]. Berry [Ger71, Kat63a]. Best [And62b, Bar49, Bur59, Chi56, Col32, Cra43a, Dav51a, Dem64, Far72, Fer58, FP69a, GT64b, Hub63b, Lan44, Mee70, Mee71, O’T33c, O’T34, Pau67, PS71, SW72a, Ste69b, Ste73, Stu67, Sug72, Tay53, TD63, Wal49e, Zid69, Zys67]. Beta [Aro59, Ban45, Ban49a, Dav70, Dav72, Hod55b, KP65, Ksh61, Mau59, OR64, OR66, Par64, Ski69, Sri68b, Sri68c, TW79, dW70, dW71, TW46, Aro41a, Jam54a]. Beta-Distribution [Hod55b, Mau59]. Beta-Distributions [Sri68b]. Bets [Mv67]. Better [GS58a, HP72, Sha64a]. Betting [Tru70]. Between [Abr70, Bak31, Blo50, Box54b, Col51, Cra33a, DJ68, Dri52, Dur68, FW69, Fré47b, HWG65, HWG95, Hil60, Hil61, Hoe47, Hsu46, JS72, KW70, LF70, Mas51a, Mas52b, Mer33, Mud66, Mui72b, Ney51, Noe51, Pie70, Rob37b, Sco50, Sha65a, Sin64, Sri62, SHB64, SPZ56, Sta38, Ste71a, Ste59c, Ste69c, Sur68b, Tat54, Tsa54a, Tsu68, Vaj51, Wee61, Wol46a, Van63, Agr66a, AG64b, Bru62, Fre66c, KS62, KR66, Rao62, Sri67, Thi31j, vB31]. Bhattacharyya [Fen59, FG52, Mur56]. Bias [AGW71, Alt39, BH57a, Kim56, Mad48a, NR67, NS60, Que56, Tuk58a, van61b]. Biased [FK51, HS70b, Sub66]. Biases [Ban44]. BIB [Sei63]. BIBD [Law71, Nem71]. BIBDS [Sta57]. Bibliography [Bri62, Gup63a, Moo64]. Big [BF68b]. Bigraph [Moo63a]. Bilinear [Cra47a, Mad40b]. Billingsley [Haj67]. Bimodal [Bak30, O’T33a]. Binary [Ber63a, Bir61a, BK60, Mar71a, Wal49d]. Binomial [Ahu71, Bah60a, Bli62, BN61, Byr35b, Cam51, Cla37, Coc40a, Cra48, Dav60, Dav61, DeG59, Fel45b, Geb68, GMS46, Gou63, HL60, Hoe47, Jog67, JS68, Joh71, Kir35, Kni65c, Mer33, MS64a, MS64b, Mur56, Nad60, Pau67, Pla48, Rio37, Sav47, Sch65, SJ71, Ski69, SW72a, Twe67, Wol46b, Wol47a]. Binomial-Type [Twe67]. Binomially [Eva40]. Binomials [HH63]. Bio [KvE64a, Tsu67]. Bio-Assay [KvE64a, Tsu67]. Bioassay [FP71]. Biological [Cha54, Eis39a, Tho38]. Biometrika [Owe67a]. Biorthogonal [Ber69b]. Biostatistics [Bro71a]. Bipartite [Hof64]. Bipolykays [Day66a, Day66b, Hoo56a]. Birkhoff [Aus67]. Birnbaum [Kui59]. Birth [Joh57a, Ken48, McN70, Mil68a, Sin70]. Birth-and-Death [Ken48]. Biserial [Lev49a, Tat54]. Bispectral [Van66a]. Bispectrum [RV65]. Bivariate [BH69, Ber72b, BJ70, Bhu64, Cha66, Eag64, EP72, Fer62b, Gri69, Hil60, Hil61, Hod55a, Hsu40, Hsu41, JK62, Jog68b, Jos69a, Jos71c, KH72, Klo59, Klo64, Kon56, Kon58, Kot64, Lan58, Lan63b, Lan64, Mus68b, Mus71, Obe47, Owe56, PL70, PL71, Pau42b, Rot71, Rub61b, Sha66a, Sid67b, Sim51, SHB64, VA43, Whi51, ZS60, Sid67a]. Blackwell [Bar50, LS08, OQ72]. Block [Agr66a, AJ56, Bha60, Bos49, BC52, BSB53, BK58b, Cau68, Con52b, Con53, Cor60, Cox40, DN62, Fre57, GW59b, Han61, Han65, Har45, HST53, Har69b, Hin64, Hof63, Joh67b, Joh70b, Kag71, Kap66, Kem56b, KB57, Ksh58, Ksh59b, KDS72, Maj62b, Maj53, Man69a, Mes67, Mik60a, Mik60b, Mot58, Mur61b, Nai66, Nai69, Nig70, Nig76, Oga63, OI65, PS67b, Rag60b, Rag62, Rag66, Rao60, Rob70b, Rob71b, Rob72b, RL57, Sei66, Sen71a, Sha59b, Sha60c, Sha64a, Sha66b, Sha67, Sha71a, Shr50, Shr65, Spr55, Spr56, Spr57, Spr59, SM66, Sur68b, Tho55a, Tho56, Tsu67, Yal68, Zys63, Rao65]. Blocks [Agr66a, Ban48, Con52a, Fra51b, Ges70, GW59b, Gre66, Hol67, Kem53, Ksh59b, Maj62b, McC39, Oga59, Pat62, PS68a, Ses66b, Sha64b, Sha65a, Sur68b, Var60, Wal59c, Tuk47a, Tuk48b]. Blumen [Klo64]. Blumenthal [Dym70]. Bolshev [Owe66, Owe67b]. Bonferroni [Mey69, SU72]. Bonferroni-Type [SU72]. Bookies [FP69b]. Boolean [Boe69, Fis72b]. Bootstrap [Efr79]. Borel [Bla68, DF65, Nas54, SB57b]. Borge [Geb68]. Boston [Dwy47b, Fis52a, Lit52]. Both [DK72, FP37, Rub58, Wal40a]. Boulder [Gua49]. Bound [Bah64, BM58a, BG47, BS64, HW71, Hoe53a, KT60b, KT61, Koo61, KQ72, Kul69, Mar70, Mar48, Mor52, Sha64b, Sto72]. Boundaries [Fed68, Moo71, Pla48]. Boundary [Abr70, Ahm61, Bro71d, Kem63, Lor70, RS70, Sta72a]. Bounded [Dub72, Far62, Far66a, Far66b, Fis67a, GRS55, SG71, She58, Sin63, Wal49f, Wal51, Wei56a, Wij71, Gun68]. Bounding [Gor41b]. Bounds [Agr66a, AC65, And65c, BM64a, BM64b, Bar65, Beh72, Bha60, BK57a, BF72, Boy58, Cha62a, DFJ68, DJ69, DT71, Dub72, Eli70, Fen59, FG52, Fra52a, GW68, Gle65, Gle66b, Ham57, Har59a, Har62, HD54, Her48, HS55, Hoe60, Ker63, Kha67a, Kot61, Kou68, KP70, Mad59, Maj62a, Mar72a, Mic59, MB66, Mud66, Mur56, Ray65, Ros70a, RP58, RB58, Roy53b, Sha65a, Sim67b, Sri71, Swi64, Tay72a, Wal59a, Wal60, Wij68a, Ylv64]. Box [GG58, McC47]. Bradley [Bro59]. Branching [AK68b, Ath68, Ath69, Ath70, AK71b, AK71a, Büh67a, Dur71, Fos71, Har48, Jag71, Jof67a, KM67, Kur70, tCM67, SS69, Smi68a, SW69, ST64, Wei65a, Wei65b, Wei66, Wei72a, Wel65]. Breathing [Alb70]. Breiman [Kal69]. Broad [Dav50]. Brown [Pyk63, Pfa71, Whe35]. Brownian [Cla71, Doo68, Ber67b, Bha71a, Cla70, Eis69, GHRM71, Huf69, Kin63b, Kni65a, Mon72a, Rod71, Roo69, She58, Yos49]. Brunswick [Old43c]. Budgets [Bol34]. Buffon [Man53]. Building [KB71, LB63]. Bulk [Fin62, Neu67]. Burgie [LeV66a]. Business [Ano53e]. By-Laws [Ano37a, Ano39c, Ano40c, Ano43b, Ano44b, Ano44c, Ano45b, Ano46e, Ano47d, Ano48e, Ano48d, Ano49e, Dwy49a]. Byars [Wel64a].

C [Kem64, Kri71a, McC65, Mul66]. C. [Gra67, Kri71a, Man45a, Wol67a]. calculable [Car67]. Calculating [Roy33, Too34]. Calculation [Bur55, Bur67, Hot43a, Hot43b, Jam68, Jen32, NV68, Noé72, Sat44]. Calculus [Blu67a, KZ62, ZF64]. Calendars [Rob36]. Can [Pit63, Pol34]. Canonical [Bar47a, CJ58, Gri69, Lan63b, sL71, Zys67]. Cantelli [Tuc59a, DeH70, Deh71, DF65, Dud69c, Gän70, Kli67, Nas54, Shu72, Tuc59b, Wol54b]. Capacities [BBT60]. Capacity [Ahl70, Ahl71, BBT59, Con70, Fin62]. Card [And43, Car34b]. Carefoot [AH71]. Carlo [Mul66, Mos62, Mul56, Wen57]. Case [Bic69a, Bri38, Bro64, Bur61, Che56, Che65b, Cra31c, DJ52, Eps54b, ES55, Far64b, Fre63b, Fre65b, Gau59, GKS63, GK64b, Gre54b, Har52, Her59, Hoa71, JV71, JV72, KM67, Kha64a, Kim47, KJB67a, KJB67b, LS50, Mee72a, Mik63, NV68, PJ71, Pol70b, RSS60, Ros62c, Ros63a, RC60b, RC60c, Sac70, Sav56, Sav57a, Sav59, SSW66, Sch43c, Sch65, SW72b, Vil47, Wal41b, Whi58b, Whi59, Win47, WK62, AT63, Dar55, HW58, Jam70, Tuk47a, Tuk48b, WW43]. Cases [Far67]. Cataloging [Har59a]. Cataloguing [Kno67]. Categorical [Bha61c, Dia63]. Categories [Pau52a, Pau62, Rob63, Rob64b]. Cauchy [Den70, Fer62b, GK50b, Lah59, Men62, Men66, PW67, PS69a, Wil69, Win47]. Cauchy-Distributed [PW67, Wil69]. Causality [Thi31c]. CDF [Hoa67]. Cell [Dic68a, Dok67b, DL66, Leh64, Moo71, Sch60c, Smi47, Swe70]. Cell-Frequencies [Sch60c]. Censored [Bas67b, DV55, Dix60, Gas65, Hal61, HM67, JM72, Pla58, RSS60, Sal66, SG56, SG58, SG69, Sär65, Wat59]. Censoring [Nad60]. Cent [RW64]. Center [Mur71a]. Centered [Mad53, Sha70d]. Centering [Pyk60]. Central [Abe69b, AJ51, And46, And64a, Car50, Con63b, Cra40, Cra41b, Eav69, HPW64a, HPW64b, Kem70, KP65, KP66a, KP68b, KJB67b, Ksh61, LLN60, Law65, Mad40a, McN62, Pie69a, Pre66, Res62, Ros70a, Ser68a, SK61, Sha63, Sri68b, Sri68c, Wal48c, WAC62, Gil64b, Har71b, Par64, PG68, SK63, Sha68, Wel64a, Abe69a, Agn57, Agn59, Bha71b, Bro69, Bro70a, Bro71b, CT58, Coc72, Eic63b, Eic66, Ger71, GZ61, GS67, Her69, Hoe51a, IW71, Ken71, KJB67a, Noe70, Par57a, Phi69b, Pie69b, vB65]. Certain [Agr66a, Ahu70, Aig71, And46, And64a, AJ65, Ano35b, BS56, Bai31, Bak35b, Bea41, Beh71, Beh72, Ber69c, Ber71c, BBT60, BW57b, BBH56, BK56, Bro40b, Cap61, CS58, CT71, Coc46, Cra33a, Cra38a, Cra43b, CD62, Dav67, DJ69, Dod40, Doo35, DF66a, DWW51, Dwa53, Dwa57b, Eis69, Eis39a, Far64a, Far67, Fel68, Fis33a, Fis33b, Fos53, Fox70, Fox71, Gei57, Ghu54, Ghu55, GMS46, Gna59, Gna60, Goo63c, Har62, Har44, Har45, Hat49, Her72, HB70, Hog51, HC58, Hog63, Høy68a, Hsu46, HK44, Jog68b, Joh51, Joh54, Joh71, Kag72, KZ71, KS62, Kem70, Kha59, Kha61, Kha65c, Kha65d, Kha66, Kha67a, Kim56, Kim68, Kni65a, Kon53, Kon54, Kon56, Kon58, Kri68, Kul37, Kul54, Lam71b, Lar35, Law71, sL71, Leh51a, LS53, Loy69, Mal54, McC69, McC47]. Certain [MH62, MS67, MD71, Mer33, Moo54, Mus67, Mus71, Nan50, NS51, NS52, O’T33b, OP58b, Pan67, Pat63, Pau41, Pau49, Pei43, Pei49, Pit68, Pol34, PP69, Rag60b, Rid31, Rie31, Rie36, Rie39, Roo69, Ros66b, Rue70, Rus57, Sam63b, Sam64, Sat41, Sch68a, Sea56a, Sha64b, Sha65a, Sha66b, Sha67, SH72, Sha55, Shr50, Sob53, Str61, Suk57, Sur68b, Tam63, Var59, Wal47c, Wei35, Wei56b, Wei56c, Wei57b, Wei57c, Wij59a, Won68, Wyn69, Zoc34, van61a, AD52]. Chain [Abr69, BF64c, BR58, Dha63b, DL70a, DL70b, DL72, DS69b, Eps49a, Eps51, Eri72, Fos52, Fre67d, Gav59, Gol63a, Goo61c, KT60b, KS61, Ken53, Lam60, Mac67, McC69, Mil61a]. Chains [AG57, BP66c, BPSW70, Bha61d, Bha61e, Bil61a, Bil61c, Bir62, BK57b, Bou68, Boz71, Bre57c, Bre58b, Bre60b, Car67, CC46, Coh70, Coh72, Dha63a, Dha64, Dha65, DN70, Dym66, Eri70a, Gil59b, Goo58a, Goo58b, Goo59, Hac63, Hel65, Ios72, KS58, Kim69, Kur69, Ley67, Ley70, Mon47, Pet69, Por64b, Por66, RR64b, RR66, RM68, Sin71, Spi72, Tsa68, Wat61, Bha68, Ore68b]. Chakrabarti [Kem64]. Chance [BG46, Cur41, Dod39, GS58b, KW59a, Wei56a, HW58, Rio62b]. Chances [DL70a, DL70b, DL72]. Change [Ano61e, Ano62e, Ano62f, Ano62g, Lig70, Lor71, NR71, Rod71, SM50, Yac68]. Changes [Bur60, Cam46, Cau72, CZ64, Gar69, KZ66, Mus68a]. Changing [Bae45]. Channel [Ahl70, Ahl71, Bil61b, BBT60, JN63, Kot61]. Channels [Ahl70, Ahl71, BBT58, BBT59, Bla59, KW72, Tho63]. Chapel [Rob50]. Chapman [Hac63, Kar61]. Chapter [Han43a, Han43b, Kul46]. Character [Dal40, DKW56, Far69b, GKS63, GK64b, Hal59, KW59a, KS65, KS72b, Nar50, RM61, WW48, Wol66]. Characterisation [Dha65, Lah54]. Characteristic [AK68a, BS69, Boa67, Bro70a, Cha70a, Cup69b, Cup69a, CL70, CL71, Hec60, Hor35, Kac48, Kha64b, Kha67b, Kin52, Kul34, Lah58, Luk57, Men66, Mij72, Moo51a, PB70, Pit56, Ram62, Sam42, Ses66a, SC71a, Ste69b, Ste73, WK62, Zyg47]. Characteristics [Alb54, Boh68, FGW46, KW56b, KR50, LD66, NR71, SD68, van61a]. Characterization [AGW71, Aig71, And71c, Bah55a, Bal59, Bar61, Beh72, Bel64a, BEM66, Cra66, Dav69c, Des71, Dyk70a, Fer64, Fis70, GO62, Gov66, Kap43, KM70, Kha62a, Kim68, KG70, Lah56a, Lah57a, Lah57b, Luk42, Luk55, Men62, OR62, PR70, RF70a, Rao58a, Rao59, Rao68b, Rao70b, Rue70, Shr59a, Spi72, Sta63, Sta71, Sud72, Tam67, Tei61b, Twe67, WK70, Wil69]. Characterizations [Bir55, BH65, Cac67, Dos69, GM72, Jog68b, Mes68, Pfa67a]. Characterizing [Kri71b]. Characters [Gho63a]. Charles [Tak61a]. Charlier [Kul48, Aro37, Bak35a, Boa49, DG57b, Hil31, Kul47, Sam43, Tru47]. Chart [Cra36a, How49, How50, Kin52]. Charts [FM58, Fox56, Hec60, Mos41]. Cheating [Dic69]. Chebyshev [BM61, Hoe66, MR69a, MO60a, MO60b, RP63, Sam66]. Chebyshev-Type [Sam66]. Check [Byr35a]. Cherenkov [Owe67b]. Chernoff [PS68a, PS68b, PS68c, PS69d]. Chi [Ala70b, Aro43, Dia63, Gil64b, Goo61d, Gov62, Gum43b, Har69a, Har71b, Hoe38, Jof67b, KA65, MW42, McN62, Mit58, Moo71, MG70, Pre66, Sat42c, Wal59a, Wal60, WAC62]. Chi-Distribution [Gov62]. Chi-Square [Ala70b, Aro43, Dia63, Gil64b, Gum43b, Har69a, Hoe38, KA65, McN62, Mit58, Moo71, MG70, Pre66, Sat42c, Wal59a, Wal60, WAC62]. Chi-Squared [Goo61d, Jof67b]. Chi-Squares [Har71b]. Chiang [Bro71a]. Chicago [Ano48l, Arn50, Arn51a, Arn51b, Old41, Rid41]. Chin [Bro71a]. Choice [And62a, And64b, Bro45, Che62, Fab68b, MW42]. Choosing [Mar72b, Par57b, Pau63, Sim67a, SW49, Woo70b]. Chosen [Kim44, Kul63]. Chow [Nád69]. Chung [LeP70, Der56a, Gil71b, VJ66]. Circle [Rao72]. Circles [Obe47]. Circular [AA50, Ber69a, OP69, Sch69c]. Circularly [Lei47]. City [Dwy47a]. Clarification [Fel68]. Class [Adi67a, Ajg67, Alb54, All53, Bea41, Ber69a, Bhu64, BBT59, BHP63, BB60, BK58b, Bre60b, Bro40b, Bur56, Bur60, Cac66, Cha63b, Cha62b, Coc46, Dan39, Dar53b, Dav50, DeG61, Dem72, Des70a, Dha65, Eat70a, Ehr66, ET72, Eri72, Eri70b, Far68b, Fel43, Fer70a, GS71, GW59a, Gre50a, GT64a, GP72, Hal51, HK64a, HB70, Hoe48a, Ifr66, Joh51, Jos71c, Kab62, Kar58b, Kem53, Kie53a, Kie53b, Kni70, Kou70a, Kou71, KB57, Kud66, Kup58a, Kup58b, Lam63, Lea66, MW42, McC47, MS69b, Mud65, Mus71, Nel70, Neu67, Ney39, Noa50, PS69c, Rao61b, Ros66b, RR59, Sch69c, Sch71a, Sch61b, Sch60c, Sea55, Sen66, Sen68b, Sen69, Sen71a, Sen72b, SH72, Shr65, Sim67a, Ski60, Sob53, Sri68a, Ste68, cS67, Tre65, Wal40b, Wal41b, Wal47a, Wei53a, Wei56b, Wei56c]. Class [Whi65, Wij68a, Wol42, Zac63, Zid71, Dar62, Mor62, Pur64, Sin70]. Classes [BP66b, BP67a, Bir55, BBT60, BC55, Dav69c, Dem60, Fel68, Goo49, Kad66, Kag72, Kon53, Kon54, Rag60b, Rao61b, Sei61, Sur71, Wol51]. Classical [Ail70, Ber45, Bha71b, DKW56, Gum43b, Kha65c, KS65, KS72b, LPC70, Par72a, PN71, Rei49, Wij57, Wij59c, Wol49b]. Classification [AB62, Ban60, Box54a, Box54b, Ell62, Ell65, Far69a, Gup62, Gup65, Gup70, Har51, Har67b, Har69b, HP49, Kab63, Kab64, Kos45, Mah63, MC65, MO70, Mos52, Nai69, Sam63c, Sax67, Sea61, Sit52, Wal44b, Wes59a, WK69, vM45, Tuk57d, Tuk57e]. Classifications [Wal41b]. Cleveland [Dwy46c, Van49]. Closed [Cro58, Dem68, KS72a, Pau64a]. Closeness [Zac66b, Zac67]. Closest [Set50]. Cluster [Gho63b, Jos68b]. Co [Mac71]. Co-Spectral [Mac71]. Cochran [Dal64a, Smi64]. Codes [Ahl71, Bla59, BS64]. Coding [Bil61b, BBT60, KW72]. Coefficient [And42b, AA50, Bac38, Ber72b, Col32, Eri71a, For37, Gou63, Hea61a, HR36, Hor32, Kem70, Kir37, Koo64a, Lei47, Lev49a, Mad45, Rob37b, Rub45, Sek51, Sid58b, Sta39, Str35, Whe31, Whi57a, Whi57b, Whi58a, Whi58b, Whi59, Wil32b, Van63]. Coefficients [And51a, And80, Cha50a, Con66b, Dun31, Dwy37b, Dwy49b, Fis33b, Fra53c, Gir42a, Gir42b, Gol63b, Gre54b, Gri31, Gri69, Ham56, Hor35, Jae72, Jam68, Jon70, Jos39, Jur71, Koo61, Lar35, Law63, Loh69, Lon42, Oga51, OP58b, Pak71, Que49b, Ros56a, Sam42, Sid60, Wri34]. Cohen [Tak71b]. Coherence [Jon69, Koo64a]. Coin [BDF64, BF64b, Dwa72, HS70b]. Coin-tossing [BDF64]. Coincidence [Fin61, Ric48, Tak61b]. Coincidences [Tak59]. Collapsed [Hac63]. Collected [Coc44, CS39, Tau62, Mal90]. Collection [GW70]. Collective [Pra61]. Collector [BB65, Ros70b]. Collinearity [Kab68]. Collineation [HF70]. Collineations [Maz67, Rak67]. Columbia [Tuk47b]. Combinability [MZ66]. Combination [HL62, Spr61, vZO67]. Combinations [Bak31, Bak46, BF63a, BP66a, CGJ67, Coh65a, HK65b, HK65a, Har69a, Har71b, Koo66, Pre66, Pro65, RO72, Sho69, Sto68c, Wei71]. Combinatorial [Bax61, BC52, Cac61, CH45, Dwy40a, DT64, Gra63, Har66a, HP63a, HP63b, HP64, Hoe51a, Hsu44, Hsu45a, Rob72a, Tra68, Rio62b]. Combinatorics [Maj53]. Combined [Dwy38a, Dwy38c, Ses66b, Spr56, Spr57]. Combining [Bat34, Fra57b, GW59b, Rak69, Rak70]. Coming [Wei54b]. Comment [Gna60, San70]. Comments [BL55, Hot40a, KP68c]. Committee [Ano40g, BCD^{+}45, CDDK45]. Common [Agr66a, Dun61, FGW46, Fis33a, Fis42, Ken39, Len39, Ric60, Sha65a, Sur68b, Zac70a]. Commutativity [Koc58]. Compact [Fur72, Sta63, Teh63, WD45, HL68]. Comparable [Bir48]. Comparative [Cha58, Cha65, Mad37, MS67, HMTW47, Tuk57a]. Compared [KM71]. Comparing [Pau62, Pau70, Rob63, Rob64b, Sri67, Wal48c]. Comparison [Abr70, Agr66a, Bah60b, Bak40, BP64, BK56, BH63, Cha50a, Coc43, Dav68b, Dav69a, Fri40, Gol63b, HK64b, Hol66, Hub63a, Hub63b, Joh67a, Kur71, MM62, MM63, McF55, Meh64, Pau52a, Per69b, RSW68, Sch68a, SB57a, Ses66b, SO56, Spj72a, Spj72b, Sto59, Tak71a, Tam69, Tor72, TD63, Mad53]. Comparisons [Bla53, Dun61, Gab69, Gho71b, GZ66, Gut46a, HL54, Kad66, O’B72, RG62, SD68, SP69, ST68, Sto61a, TR64, WG64, Gil64a]. Competing [Ber63b]. Competition [Igl64]. Competitor [Lit72]. Competitors [Bic65a, Bic65b, HL56]. Complements [RF70b]. Complete [AR49, AR50, ACH64, BF63a, Bar39, Bir55, Boe69, Bur60, Car68, Dic67a, Eat70a, Efr65b, Ehr66, Hof64, Imh60, Ios65, Ios69, Kar58b, Kie53a, Kie53b, Lan63a, Moo63a, Nev72, Sob53, Sto68c, Wal47a, Wei53a, Wol72, Wol51, Zac63].

Completely [HR55a, HV65, Nar50, Por70, Ste69e, WK56]. Completeness [BBB60, DeG61, LS50, See71, See72b]. Completion [Har66b, Tei71]. Complex [Bax61, Ber72b, BC59, BK70, Gir65a, Gir42a, Gir42b, Goo63b, Goo63c, Kha64b, Kha65c, Kha65d, Kot65, Kri70, LPC70, MS71, PN71, PJ71, Sri65b]. Component [And63a, Ber67b, Gaf59, HH72b, Har69a, Har69b, Hea61a, Hoe37, Joh54, MS71, Sta56, Sug65, Sug66]. Components [AFPR59, Bak37, BW57a, Bir58, BEM66, Con63a, Dar65c, FK72, Gra54, Gre54a, GD55, Har67b, Hen72, Hoo56a, KP69, Mah63, MD60, Obe72b, OK68, Sea56b, Sea58, Sea61, Spj67, Sto66b, Tho55b, Tho62, WGL66, Tuk56b, Tuk57d, Tuk57e]. Composed [Bak40]. Composite [All53, BP67b, Bro40b, DeG60, Leh47b, LS48, MT67a, MT67b, Rei61, Ros62c, Ros63a, Sch42b, Ste51b, Wij71, Wil38b, Woi72]. Compound [Bro33, Con69, Gil68, GH69, HR55a, HV65, HH72b, Hua72, Mac48, Oat72, Sam63a, Sam64, Sam65b, San70, Van66d, Van66c, Vot48, Wal49d]. Compounding [ZJ59]. Compression [PR71b]. Computation [Alb54, Ano33, Ano39b, Ano41c, Bis39, Bow46, Coh40, Dwy37b, Dwy38b, Kim53, Rei40, Res62, Sav60, WS34, WD45]. Computations [Har71a, LD66, Tuk56a]. Computer [Har71a]. Computing [And34, Car32b, Gut46b, Nic54, Owe56, Rob71a, Ste58a, Ste59a]. Concavity [Kei72]. Concentration [Law68]. Concept [Leh51b, Pra59, da 45]. Concepts [EP72, Leh66b]. Concerning [AB61, Ano39b, Ano41c, Bre58a, Dem63c, Der61, Dur71, Fel68, Fis60, Gar32, GH61, Gup69, Kha64b, wK68, Luk57, MW69, Pau63, Pie70, RW70, RR64b, Sen68e, She64a, Sim67a, SW49, Tak59, Tak61b, VN67, Wal47c, Wal49d, Wau45, vN42]. Concise [Sat41]. Condition [BD72, DeH70, DM68a, Eps49a, Eps51, Far68a, Gil71a, Isa63, LS60, Nan50, Oga59, Rub58, Rue71, Ste70, Ste55, Sto70a]. Conditional [BB68a, Bat65, Bla47, BRN63, Boe69, BH65, Bru65, Bur62b, Cog62, Con63a, Dyk70a, Ekm59a, GP71, HH63, Hoe51b, LR72a, Lau63, Luc68, MR69a, Pfa67a, Rob65b, Rob66, Rod66b, Str65b, Wal59b, Bro67, BF63b, Fra68b, Wen57]. Conditionally [MS69b, Teh63]. Conditioned [Bel70a, Bel70b, DK63, Lig70]. Conditions [Bar69, BE72, Bri63a, Cau72, Dha63b, Eis69, Gil71b, Ken40, Kes72, Kha62b, Kha73, Koz51, Luc67, MS70b, Min72, MR69b, Par58, Pie69a, Pie69b, Sam41, Sei63, Smi68a, Wal49g, Zid70, van61a]. Cones [Rag70b]. Confidence [Alb66, And65c, Ban67, BP66b, BP67a, Bha60, BT51, BM58a, BF72, CR65b, Dwa55a, Far62, Far66a, Far66b, Fra52a, Gee70, Gle65, Gle66b, Goo64b, Goo76, Gra58, Gre54a, Gut48a, Gut70b, Hal61, Har70c, Hea61a, Hoe54, HS61, Jos66b, Jos67a, Jos67b, Jos69a, Jos70, Jos71c, Jos71d, Kha69, Kha67a, Kol41, Kou69, Kou71, KF40, Lan71, Leh63b, Mal54, Man58, Mas50b, Mud66, MS68b, MS71, Nád69, Ney35, Noe48, Noe51, Pra63, Ros63b, RB53, Roy54, Roy56, RP58, RB58, Sch70c, Sco67, Sco68, SG71, Sen71b, Sim68, Sin63, Sri71, SW47, Ste59c, Tho36, Vil64a, WW39, WW41, Wal42a, Wal58b, Wal59b, Wei55b, Wei55a, Wel39, Wer39, Wil38c, WD39, Wol46a, Tuk58a]. Configurations [Ber69b, Cox40, Hed70, RW70, Spr55]. Confluence [Rei40].

Confounded [Zel58]. Confounding [Ban70, WH65]. Congestion [Ney66]. Conjugacy [Bro63]. Conjugate [Coh40]. Connected [Ber42, Bha60, Cox58, Den70, Hil31, Kha67b, Lah57b, Mal47, dW70, dW71]. Connection [Hoe69, Noe51]. Connections [Ios65, Ios69]. Conover [Dav66]. Consecutive [Kap45, Wol44b]. Consensus [EG59, Nor67]. Consequence [Fis71]. Consequences [Zys63]. Conservative [Dav60, Dav61, Sco67, Sco68]. Considered [Hen32b, Hen35a]. Consistency [Ber69c, Ber70b, Ber71c, Ber72c, Ber73, BW57b, Bow44, Eic63a, KW56a, Lan72a, Leh51a, Loy69, Mal70, MH69, MH70, OQ72, Rao61a, Van69, Wal49a, Wol47a, Wol49a]. Consistent [AKKW69, Ber51a, BLS64, Dav50, Far67, Gaf59, Ges70, LS60, Ney51, Par72b, Par57c, Pfa69a, Sco50, Yak69]. Constant [Bar43, Cac67, Eri71a, Fel32, Rao70b]. Constants [Kim68, O’T33a]. Constituent [Vaj45]. Constitution [Ano37b, Ano39d, Ano40d, Ano41d, Ano43c, Ano44b, Ano44d, Ano45c, Ano46f, Ano47e, Ano48e, Ano49f, Dwy48a]. Constrained [Mal68b]. Constraint [Leh47b, Sch42b]. Constraints [Sei55a, Sei55b, Sei56]. Constructed [NR67]. Constructing [AJ56, Fre57, God61b, HF70, Pat62, RC62, Sha60b]. Construction [AD70, Agg72, AJ58, BF66, Ban70, BSB53, BC59, Byl70, Cha61, Cox40, DN62, DK68a, DM67, DM68b, DM70a, Gul72, Gut70b, Han61, Her67c, Hoe71, Kem68a, Lam71b, Man43, Maz67, Rak67, Roy53a, Sei54, Sei61, Sei66, Shi71, SA71, SM68, WH65, Wic70, Eli72, Man42]. Contact [Alt39]. Contagion [CC47, Yan68]. Contagious [Fel43, Ney39, Rut54]. Contain [AJ51]. Containing [Bla68, RS67, Sta56, Stu67]. Contains [GS58a]. Contamination [Jae71a]. Content [Gut70b, Rub60a, Rub60b, Rub61a, Rub61b, Rub62a]. Contents [Thi31a]. Contiguity [Beh71]. Contingency [Cau72, Fie70, Fie71, Goo63a, Goo64a, HK64b, Kul71, Lin64, Moo49, RK56, Roy33, Wei34, Wil35, Fie68]. Continued [Aro41a, Aro59, Dwy38c, TW79]. Continues [Dol69]. Continuity [Dun70, Has63, Huf71, KZ71, Kus69, Ore68a, Pre72, Wij70, Wol71c]. Continuous [AE69, Alb62, AK68b, Ath68, Bat30, Bir53a, Bir53b, BM61, Bla57a, CL56, Dal69b, DS53, Den64, DLS57, DJL59, Dod43, Dol69, DKW53b, DKW53a, DKW59, Eav67, Ekm59a, Far64b, Fra63, Fre63a, Gav59, Get67b, Goo61c, Han59, Isa71, KW72, Kie57, Kur69, Law33, Lie53a, LS55, Mar70, Mar72a, Mon72a, Mor52, Mul56, OT61, OT65, Rao68a, Sak64, Sak66, Sto66b, Sto69b, Tat54, Tuc64, WW39, WW41, WW45, Wal62, Whi65, Tuk47a, JV72, Tau62]. Continuous-time [KW72]. Continuously [Kak71]. Continuum [Kei63]. Contour [McN62]. Contour-Integral [McN62]. Contours [BT51, Mal54]. Contrast [Lan72a]. Contrasts [BP65a, Goo64b, Goo76, SP58]. Contribution [Lot39a]. Contributions [AN53, Blu66a, Blu66b, Dix44, Fel62, Gir46a, Gir46b, Mad37, RSS60, RG59a, RG59b, Sav56, Sav57a, Sav59, Sav60, Sch69d, Ser68a, Tin52, Wal39, MR64].

Control [Ano43j, BH57a, Com64, DJL59, Der64b, DS66, GR52, Hol66, How49, How50, Kin52, Mos41, Pau52a, Pau62, RH71, RSW68, Rob63, Rob64b, Sat44, Ton69, Wei54b, Wol43, Zac70b]. Controlling [DS66, Wei52]. Controls [See72a]. Converge [Blu54a]. Convergence [Bah67, Bar61, Bau63a, Bau63b, Ber45, Bha71b, BW71, Blo71, Blu55, BW64b, Bro70b, Bro71c, Cha64, Cho65, Cho66b, Cho68, Cho71, Dav68d, Dav68c, Dav68b, Dav69c, DeH70, Den66, Dud69c, Far72, Fer70b, Fer71, Gän70, GP71, Gil72, GHRM71, HV65, HK65b, HK65a, Han67, Her72, Her69, Hey69b, IT68, JK70, Jiř72, JN63, Jog67, JS68, JK66, JV71, JV72, Kaw59, Ken72, Koo66, Kou70c, KS72d, KsW69, Koz47, Koz51, LR72a, Le 70b, Lig70, Mar69, Mas70a, MS72b, Neu71, Nor67, OS70b, Pfa69b, Pic68, PR68, PS68b, PS68c, Rag70a, Rao62, Rei49, Rob48a, RO72, Rub56, Sam64, Saw72, Sch71a, Sch47, Sch70e, Sen70b, Sen71c, Sen72c, Ser70a, Sho72a, SJ71, Sto65c, Sto70a, Sto68c, Top67, Ull44, Var66, Var68, Ven67b, Wag69, Wal70a, Wei66, Wei55c, Wei57c, Wel71]. Convergence [Wen61, Whi70, Wic69, Wic71, Wol72, WZ65, Wri72, vB65, Cha69]. Convergent [Koo61, Wic70]. Converse [Rob72a]. Convex [BL69, DR63, Dem72, Far68a, Fis69c, Kin63b, Mad59, Pro65, Rag70b, She55, Yar72]. Convexity [Bec42, Hil62, Mil61a, Nor37]. Convolution [BH59, Gaf59, KS72a, Mas70a, Old52, Tei54a]. Convolutions [KP60a, Mas70b, Ney63, Tuc68, Dar65b]. Convolutive [GW59a]. Coordinate [Eat70b, Kru68, Rot71]. Coordinate-Free [Kru68]. Coordinates [Kot64]. corner [OT47]. Cornish [HD68]. Correct [Cha35, vM42]. Correcting [BS64]. Correction [Abe69b, AG60, Alb48, Alt39, AG65, And80, AJ65, Ano47f, Ano49g, Ano62h, Ano69g, Aro59, Arv70, BR51, Ban69, BP67a, Bat31, Bel70a, Ber66a, Ber71c, Ber73, Ber58a, Bic65a, Bla58a, Bre60a, Bre70, Bro71f, Bur67, Cha65, Cla71, Con53, Csö65a, CL71, Dar53a, Dar62, DK58, Day65, Day66a, DL70a, DL72, Dok68, DKW59, Dwy60, Enn73, Eps51, Fer63, Fie71, Fre68, Geb68, Gir64a, Gir68a, Gir42a, Gle66b, Gle68, Gna60, Goo76, Gup62, Gup70, Gur59, Gur62, HWG95, Har60b, Har64a, Her70, Hil61, HP64, Ios69, Iye61b, Iye61a, Jen70a, Joh61b, Jor59, Jos71d, Kad73, Kha65a, Kha61, Kie53a, KS72b, KHS74, Kin61b, Kon54, Kon58, KA65, KP68c, KP60b, Kul48, Kul56b, Kul68a, Kum65, Kup58a, Lan64, Lin52, Man46, Mar63, Mas52a, MT67a, MH70, MS64a]. Correction [NS52, NS60, Nig76, Obe72a, Oka68, Old49, OT65, OR66, Pei49, Per71, Pie69a, RR72a, Rid59a, Rob71b, Ros63a, Rub61a, Sak66, San62, SG69, SS67, Sch61a, Sco68, Sei56, SP68, Set70a, Sha59a, SK63, Sha67, Sha68, Sid67b, Sin66, Spr57, Sri65a, Sri68b, Sri69, Ste59a, Ste73, Str72c, TC80, Tei61c, Tuc59a, Tuk62a, Wal60, Wal53, Wat63, Wij59b, Wij59c, Wij75, dW71]. Corrections [AG60, And64a, AG64a, Aro59, Ber71c, Bic65a, Bil61a, Bra52, Cra36c, Cra41a, Dav61, DKW59, Goo57a, Goo61b, Gur59, HW48, Hil61, Iye61b, Iye61a, Jor59, Kab64, Ken41, Kha73, KP68c, Kul35a, Lew35, Rid59a, Rub61a, Sha59a, Ste59a, TW79, Tuc59a, Wij59b]. Correlated [Cle71, Cra31c, Cra47a, Gur55b, Gur59, Kaw50, KA64a, KA64b, KA65, Lei47, Mat49, RP58, SY66, SY68, SY70a, Sid58a, Sid60]. Correlates [Thi31m]. Correlation [And42b, AA50, Bac38, Bac48, Bak37, Bar47a, Bel62, Ber72b, Ber37, Box54b, Col32, CJ58, Cra33a, Dal68, Dix44, Dun31, Dwy37b, Dwy44b, Dwy49b, Fis33a, Fis33b, Fis42, Gri31, Gri69, Gut38, Hal51, Hoe40, Hor32, HP36, KS62, Kem70, Kha64a, Kir37, Koo42, KP67, KP68c, Lan63a, Law63, Lei47, Lev49a, Mad45, McF55, MG69, Mor62, Oga51, OP58b, OT61, OT65, Pak71, Pre69b, Que49b, Rob37b, Rub45, Sek51, Sid58b, Sri65c, Sri68b, Sri68c, Sta39, Str35, Tat54, Wal47c, WD51, Whe31, Whi57a, Whi57b, Whi58a, Whi58b, Whi59, Wil32b, Woo40, Yua33, Van63, Que49a, WW43]. Correlations [Cha62a, De 38, Lan63b, sL71, Šid68, Šid71]. Correspondence [KW70]. Corresponding [BM59b, SM50, Sug66]. Cost [Der66a, Ros71c, Sim68]. Costs [Sch67b]. Cotterman [Man45a]. Council [Dwy49a]. Count [Goo61c]. Countable [Bla55b, Bla58b, DV67, Kak71, RS72b]. Countably [tCM67, Wei72b]. Counter [AN53, Dav67, Pyk58]. Counter-Example [Dav67]. counterbalancing [Hou66]. Counterexample [Bre57a, Hen72, Roo71, Tor71, Pfa71]. Counterexamples [LR72b, RW70]. Counters [San61, San62, Tak58, Tak61b]. Counting [GRS55, IW71, Mal47, Moo68a]. Counts [Kin63a]. Coupon [BB65, Ros70b]. Course [Jai70, LeP70]. Courses [Ano43j, Hed70, Hed70]. Covariance [AB62, And63b, AG64c, And65a, AG65, CB48, Con69, Dem63c, Dem64, Dyk70b, Fed51, Fra63, Gir68a, Gir68c, pH69, Jam60, Jam61, Kha65d, KP69, KP67, KP68c, Kri68, LaM72, Lay72, Mad48a, PAAJ69, PN71, PS69b, Sea56b, Sid58c, SN68, Sug69, Sug67b, Sug70, Vil69a, Ylv64, Zys67, LW44]. Covariances [Bal59, Ell65, HM67, Lie53b, Sid58a, Wil46a]. Covariants [Zoc34]. Covariate [MO70]. Coverage [GT64a, Hal60, MM69, Sau63a]. Cox [Der64a, Han66a, Pra66a]. Craig [Car50, Hot44a]. Cramér [Fel47, Ylv68b, And62c, Bur63, Bur64b, Dar55, Dar57, Fen59, Kie59, Kul54, RW72, Gar59, Ylv66]. Cramér-Rao [Gar59]. Criteria [AJ65, BMT66, Bha61d, BW71, Bil56, Bue59, Cha51, Con66b, Con67c, Con69, Goo58a, Gru50, Kab62, KS62, KP68b, Kie62, Kus69, Mil71a, Pil55, Sha60c, SN68, Sug69, SF69, Tro71, Vei69, Wil46a, AD52]. Criterion [And62c, Ber45, Bur63, Con67d, Der66a, Dix40, Hol69, Jaq72, MR71, Pil68, PAAJ69, PN71, Tho63, Tho35, TW34, Bic69b]. Critical [Ham57, Hey70, KM67, Wei72a, Wei57a]. Cross [Har69b, Koe70, Mor62, Mur63, Rao67, Rao68c]. Cross-Spectrum [Mur63]. Crossed [Ban60]. Crossing [NR62, RS70, Sta72a]. Crossings [Ber71d, CL65, LC65, Lea66, Qua68]. Crystals [Gil62]. Cubic [Joh70a, RR64a]. Cumulants [Dos69, Lor59]. Cumulative [Agg55, Bur42, Dri52, Dwy38b, Dwy40b, Eri66, Hat49, Joh59, Joh61b, Kem59, Mas51a, Rob54, Ser70b, Ste46, Tsa54a, Wal44c, Wal45b, Wis57]. Cumulatives [Kou70c, KS72d, Mas52b]. Current [CZ64]. Curve [Ail70, Car32a, Dwy40a, Enl34, Gre32, Hil62, LC65, Que36, Ros71a, Tor52].

Curves [And34, Bai31, Cas50, Cra36a, Hoe54, Lea66, Mol46, Mye34, O’T33c, Pin47, Wil30, Win47, Zoc34, Zoc35b, Thi31f]. Curvilinear [Eze30]. Cusp [Rao68a]. Cutting [DK68b]. Cycle [New60]. Cycles [Dod39, Pow30]. Cyclic [DW65, Dod41a, Dwa62, Maz67, Pre69a, Rak67]. Cylindrically [Her66, Her67b, Her70].

D [Gre64, Mul66, Pfa71, Pra66a, Pyk63, Rio62b, Wan68, Wel64a]. D. [Der64a, Han66a, Pra66a, Rub68]. d.f [Gov62]. Dallas [Old42b]. Dam [Gho60, Pha63, Pra58a]. Dams [Gan61]. Daniels [Hil61, Hil60]. Dantzig [Gre50a, Hem60]. Darling [Lew61]. Darmois [Gup72, Dug61, Pau70]. Data [AN53, BP69, Bas67b, Bha61c, Bha67a, CS39, Dem66, Dia63, Dod30, EC44, Fer72b, Fra67a, Gas65, Hun55, JM72, Kal59b, MP71, Pie40, Pla58, PR71b, Sen67, Swi67, TB64, Tuk47b, Tuk62a, vM45, Tuk62b]. David [Gil64a, Ore64, Rio62b, Sav71, Wat66, Hem60, Hub63b]. Death [Joh57a, Ken48, McN70, Mil68a, Sin70]. Decapitated [Ahu71]. Decaying [Hai61]. December [Arn51a, Fis50a, Fis52a]. Decimal [HW48]. Decision [Alb54, And62a, ACH64, BG52, Bah54, BL55, Bec54, BS54, BEM59, Bha64, Bha68, Bly51, Bly70, Cog67, DK56, DK58, Der66a, Der66b, DV67, DWW51, DKW53b, DKW53a, DKW59, Efr65b, Enn70, Fab62, FR68a, Fis68, Fox71, Fur72, GH69, GKR57, Hal58, Hal59, HR55a, HV65, HH72b, Hua72, Jaq72, Kak71, KR56, Kon53, Kon54, Kud66, Lad55, Le 55, Lec62, Leh57a, Leh57b, Lip71, Lor72, Mat55, Mei69, Nel66, Oat72, Pau49, Pau63, Pau64a, PP69, Rad62, RK57, Rob64a, RS72b, Ros68a, Ros68b, Ros71c, Sam63a, Sam64, San70, SY70b, Sch69a, Sea55, Shu69, SW49, Sob53, Ste70, Sto67, Van66d, Van66c, Wal47a, Wal49c, WW50, Whi65, Wij68b, WF57a, Wol51]. Decision-Making [Shu69]. Decisions [HL52, Wei54b]. Deck [Gre44b]. Decomposition [HC58, Kni70, Ksh59a, Lut65, PS72, Rao69, Sim70, Gun68, Luk65]. Decreasing [Bre68, Bre69, Bre70, Smi39]. Defective [Dor43, Ste57b]. Deficiency [HL70]. Defined [Coh72, Mil61a, Vaj51]. Definite [Gan66, Gur55a, Gur62, Kha66, Pac55, Rob48b, SK61, SK63, Sha63, Sha68, Sha70a, Suk72]. Definiteness [Dyk70b]. Definition [AA63, Bri62, CL70, CL71, Fré47a, Gei45b, Ham71, Har70b, Ols67b]. Definitions [Fis67b, Pyk61a]. Degeneracy [BJ68a]. Degree [And62a, Aro48, Day64, Day65, Kha62b, Kha73, KA65, Lah56b, Moo63b, O’T33b, O’T33a, SU72]. Delays [Bur64a]. Delicate [Bre68, Bre69, Bre70]. Delta [Woo70b]. Delta-Sequence [Woo70b]. Demand [Bur32]. Demand-Supply-Price [Bur32]. Dennis [Che66]. Denominator [Gor41a]. Densities [Dic67a, Fer58, GO69, Ifr65b, Koe70, Kul68a, Kul68b, Kul69, MS69a, Sha69b, Ste69e, Ste71c, Str59, WV67]. Density [Bar69, BG68, Bue65, Cra65, Dem69, Far67, Far72, Fri72, Ges70, Joh67d, Kei63, Klo70, Kru54, LQ65, MH69, MH70, Mur65b, Par57b, Par62, PB70, Pet54, Pic69, Pol70a, Rao68a, Rob67, Ros56b, Sch69e, Sch70e, Sch67e, SQ71, Spr53, TK70, Van69, Wah71b, WL63, Wat69, Weg69, Weg70a, Weg70b, Woo67].

Denumerable [Bre58b, Der66a, FR68a, Fis68, Ros68b, Ore68b]. Denumerably [Mur72, Por66]. Denumerably-Valued [Mur72]. Departure [HB70]. departures [FN61, FP64]. Dependence [Bel62, Ber72a, Bha72, Blo50, Bur64a, EP72, Høy68a, Lan63a, Leh66b, Moo43, Šid68, Šid71, Tak62, Gab62]. Dependent [Ath69, Ber62c, Bha64, Blo70, Blo71, CR63, Coc72, Dun58, Dun59, Dur71, Fos71, GM62, GZ61, KK60, KK62, Kim51, Kur70, New64, SR68, Sen68a, Ser68a, Tai66, Tho72, Wat54, Wei65a, Wei65b, Wei66, Wei72a]. Depending [Fra52b]. Depreciation [Mor32]. Derivable [BB60]. Derivation [Cac61, Dav50, Fos52, God49a, Gut38, Kir37, McN62, Kri71a]. Derivative [Byr35a, Var67]. Derivatives [Cap64, DM48, Isa69b, KZ71, Kai71, Pit56, Sch69e, She66, Sal30a]. Derived [Hel65, Jam64, Kul37]. Derman [Sch67a]. Design [Agg72, Alb61, AJ58, Ban49b, BH57a, Bli46, Boh68, BK56, Che59, Dra60b, DB71, Fab68b, Far69a, GP58, Hag66, Han65, Har44, Har45, Hed70, HS70a, Jam57, Jos68b, Kem56b, Kem64, KS63, Kis45, Mes67, Mot58, MD68, Oga63, OI65, Owe70, Rao60, Rob63, Sha70c, Sto59, Tho58, Vog60b, Wah71a, Wal43b, WK56, Zac68, Zys63]. Designing [Gru49]. Designs [Agr66a, Agr66b, And72b, AJ56, Atw69, Ban48, Ban65, Ban66a, Ban66b, Ban69, Ban70, Bha60, Bos49, BC52, BSB53, BC55, BC59, BM59b, BD59, BS64, BH57b, BB60, BK58b, BJK56, Cau68, Che53, Con52b, Con52a, Con53, CC54, Cor60, DN62, DW65, Dem60, Dem61, Dra60a, Dra60c, Dra61, DM67, DM68b, DH68, DM70a, Ehr55, Far68b, Fre57, Fre58, Fre61, GGH59, Gre66, Gul72, Han61, HST53, Har69b, HF70, HSF72, Her64, Her66, Her67b, Her67c, Her70, Hin64, Hoe65b, Hoe65c, Hof63, Joh64a, Joh66, Joh67b, Joh67a, Joh70b, Jos71f, Kag71, Kag72, Kap66, KS66a, Kem53, Kie58, KW59b, Kie61, Kie62, KW65, KB57, Ksh58, Ksh59b, KDS72, Maj62b, Maj53, Man69a, Maz67, Mes67, Mik60a, Mik60b, Mur61b, Mur71a, Mur71b, Nai66, Nai69, Nig70, Nig76, Oga59]. Designs [Pat62, PS67b, Rag59, Rag60b, Rag60a, Rag60c, RR64a, Rag64, Rag66, Rak67, Rao58b, Rao61b, RC62, Rob70b, Rob71b, Rob72b, RL57, SY66, SY68, SY70a, Sch60c, Sei61, Sei63, Sei66, Sen71a, Sha59c, Sha59b, Sha60b, Sha60c, Sha64b, Sha65a, Sha66b, Sha67, Shr50, Shr51, Shr65, Spr55, Spr56, Spr57, Spr59, Sri68a, SA70, SC71a, SA71, SM66, Stu68, SV69b, Stu71a, Stu71b, Sur68a, Sur68b, Sur71, Tha63, Tha65, Tho56, Var55, Var59, Var60, WH65, Wyn70, Yal68, Yal71, Zac64, Zac66a, Zel54, Zel58, Rao65, Rao66, Tuk56b]. Detect [Dod30]. Detecting [Gar69, NR71]. Detection [Bat67, Dor43, IT68, Ste57b]. Determinant [Goo63b, Rei40, Suk72]. Determinantal [And65c, Gir39, Mic59, Nan48a, Nan48b, Nan50]. Determinants [Hoe40, Nic58]. Determination [BZ49, Dun31, Fra72b, HH49, Hot41, Thi31r, Wil41a]. Determined [DK68b, Fra51b, Mac67, Wei61]. Determining [Gra58, Har59a, Har62, Hor35, Kha69, O’T33a, Que36, Rob30, Sam42, Sim67c]. Deterministic [Kni72]. Detroit [Ano56n]. Deutsch [Mid66]. Developed [ZD38]. Developing [Kim47]. Development [Tuk47b, BCD^{+}45]. Developments [For37]. Deviates [Dwy64, Fin46, BM58b]. Deviation [Bak35b, Bak46, Boo72, Cra32, Dix57, Fré47a, God49a, GG39, Gru44, Har59b, Har60b, Hey67, Jen32, Kat60, Kni63, Mas51a, Mas52b, PR54, Sar54, Sar55a, Sar55b, Spr53, Tho35, Tsa54a, Wal58c, Woo67, WV67, Zei50]. Deviations [BR60, Bak32, Bak35a, Dav68c, Dun31, ET68, Hoa67, Kac48, KHS72, KHS74, Set64, Set65, Set70a, Sie69c, Sin40, TW34, Wei56a, Woo70d, Tuk46]. Diagonalization [HA72]. Dice [Fer70a]. Dichotomization [Blo51]. Dichotomous [AE69, BK56]. Differ [Bak35b, Suk58]. Difference [And59, Bru62, Dri52, FW69, Gei56a, Har42, HvN42, Hil31, Hsu48, Kat60, Kul36, Lom52, McG68, Mil68c, PR70, Rao61a, SPZ56, Sta38, Wil41b, Wol54a, vN41, vNKBH41, vN42]. Differences [Bak52, Bar35, Cho68, Dod41a, Goo61a, Hsu44, KS62, Kim50, KOvE72, LL39, Man45c, MG47, Old38, Old49, Sto68c, Wei55c, Wil30]. Different [AB62, Coc43, Sav57b, Tei57]. Differentiability [Has63, LR72c]. Differentiable [Gan66, MS72b, vM47]. Differential [Bea37, Dav68a, Eri71a, Hil31, Huf69, Mui70b, RT59, Sal67, ZD38]. Differentiation [Har47, Wal46a]. Diffusion [Ber64c, Hel70, Jag71, Wat62, Wat63, Spi66]. Diffusions [Bro71d]. Digital [Har71a]. Digits [Pin61, HS49b, Mar71a]. Dilution [FP71]. Dimension [Den69, Hen72]. Dimensional [Ath69, BN45, Cau68, Dem58, Gut70a, HS72b, Hoe65b, Ken71, KP66b, Leh63d, MD71, Rai72, Sch34, Sim71b, Sta56, Sto65a, Wic69, HP63b, Mil68a, Sol53]. Dimensions [BRZ47, BG46, BD59, Cam48, Dra60a, Dra60c, Dra60b, Dra61, Her64, Her67c, Jee54, PS71, RG59a, RG59b]. Direct [Gre65b, HS49b]. Direction [Kab68]. Directional [Ney51]. Directions [Sch70c]. Directory [Ano35a, Ano37c, Ano39e, Ano40e, Ano43d]. Dirichlet [Mul56]. Discontinuous [Fré47b, Isa71, Sch58, Tuk48b]. Discount [Swe70, Vei69]. Discounted [Bla65, DT67, Ger71, Kak71, Ros68b, Swe70]. Discounting [DM68a, Vei66]. Discrepancy [Ste59c, Tsu68, Weg56]. Discrete [Ahl71, Bat30, Bla51b, Bla57a, Bla62, Bro65, Bru69, Bur55, Bur67, Cha63c, Che65b, Cra36c, DM68a, Den72, Dwa72, Fis42, Fos52, Fre63b, Fre65b, Gil72, HH72a, Har72, Hin65, Kat60, KT61, Kin62, Kot61, Mee72a, MV69, Mur72, Noa50, OT61, OT65, Por65, Sac70, SY70b, Tat54, Tsa56, Tsa68, Vei66, Vei69, Vor72, Wal48a, WK71, Ord68, JV71]. Discrete-Time [SY70b]. Discriminant [Bar51, Bro47a, CB48, pH69, Kab68, MO70, Oka63, Oka68]. Discriminants [Gei67]. Discriminating [Hoe47]. Discrimination [Bro71g, Cam70, Fre67b, Fre68, Ham72, Joh61a, Koo60, KK66, QG68, Rao54, Yar71]. Discriminator [Dem64]. Discussion [Dem40, NCC65, vMD41]. Disjoint [Sha64b]. Dispersion [Gna59, Gna60, Gup69, Hen35b, Hoe43b, Jae72, Mik62, MG47, Mos63, Nor38, Ros53, RG62]. Dispersions [AB60]. Displacement [KOvE72]. Disputed [Wat66]. Dissected [Pol34]. Distance [AJ65, Gei45b, Ham50, KKW55, Lor54, Mat55, Sri67, Tay53, Wol54a, Wol57]. Distances [Dud68, SHB64, Sri67]. Distinct [Agr66b, BW49, HH72c, KS70, KS72c, Rak69, Sch72, vM45]. Distinguishability [FV69, FV71, HW58]. Distinguishable [Lan72b]. Distinguishing [Kan69, She65]. Distriburtions [Bak30]. Distributed [Aro47, Bak37, Bha72, CR63, CT71, Eva40, Goo63b, Gor41a, Hoa71, HS55, Koo68, KA64a, KA64b, Kul36, Lah56b, PW67, Sim71b, Wil69, HW58, Jag63]. Distribution [Abr64, Agg55, Ahu71, Ail70, Ala70b, Ala70a, Ala72, Alb56, And42b, AG44, And46, AA50, And62c, AW64, And64a, AG64a, And65a, And70, And71c, Ano31a, Ano35b, Aro41b, Aro43, Aro48, AFLP57, ABE^{+}55, Bag65, Bah60a, BW64a, Bak32, Bak46, Bal67, Bar70, BGP69, Bar47a, Bas64, Bat31, BK63, Bel60, Bel64a, Bel64b, BD65a, BD65b, BMT66, BD66, BD67, Ber69a, Ber71b, Ber71a, Ber69b, Ber62c, Ber64c, Ber58a, Ber58b, Bha61d, Bic69a, Bin72, BT51, Bir53a, Bir54, BR54, BM58a, Bla55a, Bla56, Bla58a, BW67b, Blu55, BKR61, BB69, Bow46, BC64, Bre63, BK70, Bro70b, Bru62, Bur72a, Bur59, Bur55, Bur63, Bur67, But69, Cac61, Cad53, Cam38b, Cam46, Car32a, Car46, Cha63b, Cha70a, Cha56, Che54, CZ64, Che65a, Che65b, CGJ67, Chu55, Chu56, Chu46a, Cla37, Coh65a]. Distribution [Con63a, Con67a, CJ58, Con63b, Con66a, Con67d, Cra31a, Cra32, Cra41b, Csö65a, Csö65b, Cur41, Dan54a, Dan39, DD64, Dar64a, Dav68a, Dav70, DK68a, Des70a, Des71, Dic67b, Dok66, Dri52, Dun31, Dur68, DKW56, Dwa57b, Dwa58, Dwa59, Egy59, Eic70, Ekm59a, Ell64, Eps49a, Eps51, ES54, Eps54a, Far66a, Far66b, Fed68, Fel45b, Fel51, Fer62b, Fer64, FP37, Fin46, Fis70, Fox70, FG60, Gal72, Geb68, Gei56b, Gei57, GG58, Gho54, GO62, Gil72, Gir42a, Gir42b, Gla59, God49a, Goo56, Goo57a, Goo57c, Goo61b, Goo61d, Goo63b, Goo63c, GM72, Gov62, GS55, GG50, GD55, Gre41, Gre50a, GH68, Gru44, Gui71, Gum47, GvS50, GH51, GG51, GS58b, Gup61, GS65, Gup71, Gur48, Gur53, Gur55a, Gur55b, Gur59, Gur62, Gut48a, Gut70b, Ham50, pH69, HH63]. Distribution [Har65a, Har52, Har70b, Har51, Har65b, Har71b, Hec60, Hel70, Hen36, HR36, Her48, Hil69b, Hod55b, HL60, Hoe48a, Hoe53b, HS55, Hoe56, Hoe38, Hoe69, HPW64a, Hog56, Hol69, Hol66, HF38, Hsu45c, Hun39, Ito56, Ito60, Jam60, Jam61, Jen70a, Jen70b, Jiř72, JK62, Jon48, Kab62, Kab63, Kab64, Kab65, Kap43, Kap45, Kap71, Kat61a, Kat52, Kat63b, KM70, Kel71, KM59, Kha62a, Kha64a, Kha64b, Kha65c, KP65, Kha65d, Kha66, KP66a, Kha67b, KP68a, Kie52b, KW59a, Kim42, Kim44, Kim46, Kim47, Kim49, Kim50, Kim56, KG70, KP66b, Kir35, Klo59, Kni65a, Kot64, KA64a, KA64b, KvE64b, KA65, Kri71b, KP51, KP60b, Ksh59a, Ksh59b, Ksh61, Kul34, Kul36, Kul37, Kul63, Lad39, Lah54, Lah57a, Lar36, Lar39, Lau63, Law33, sL71, Lei47]. Distribution [Lew61, Lie53b, Lie55, Lin50, Lin52, Lor54, Lor71, LL39, Luk42, LK54, Luk55, Mac71, Mad40a, Mad45, Mal68a, Mal54, Mar70, Mar58, Mas50b, Mas51a, Mas52b, MS69a, MR71, Mau40, Mau59, McN62, Mee70, Mee72a, MS69b, Men62, Men66, Mes68, Mic59, MB66, Moo40, Moo41, Moo51a, Mor52, MD60, Mor62, Mos62, MS64a, MS64b, Mui72a, Mur56, Mus68b, Nad60, Nan48a, Nan48b, Nan50, New68, Nic56, Nic58, OQ72, Obe47, Oga63, Oka63, Oka68, OR54, OR62, OR64, OR66, Olm46, Ols38, Ore68a, OS62, Pac55, PL70, PL71, Pau41, Pau42a, Pau42b, Pau52b, Pau63, Pil51, PR54, PM59, Pil67, PG68, PN71, PJ71, Pin61, PW63, Por70, Pra58b, Pra63, Qua68, Que49b, Que36, Rao58a, Rao59, Ras48, Rea72, Res62, Rie31, Rie36, Rie39, Rob44c, Rob48b, Rob48c]. Distribution [Rob54, Rog64, Rou69, Roy53b, Rub60b, Rub61a, Rub62a, Rub63b, Rub45, Rut54, Sac58, SA66, Sal66, SG58, SGO63, SG69, Sar57, Sar60, Sat42a, Sch69b, Sch43b, Sch72, SMR72, Sen66, Ses66a, Set50, SO56, Sev57, SK61, SK63, Sha63, Sha66a, Sha68, Sha70a, She58, Sho66, Sid58b, Sid65, Sid67b, Sim51, Sin66, Sio71, Sit52, SW49, Sol53, Spr53, SHB64, Sri65b, Sri68b, Sri68c, SPZ56, SP58, Sta62, Sta63, Sta39, Ste58b, Ste60, Ste71a, SW47, Sto65a, Sto68b, Sto69b, SC71b, Sug65, Sug66, Sug67b, Sug67a, Sug70, cS67, Tak59, Tak71a, Tak69b, Tan64, Tat53, TG58, Tei56, Tei61c, The64, Tho35, Tho36, Tho38, Tsa54a, Tsa56, Tsa68, Tuc64, Vil69a, Vot46, Vot48, Wah68, WW39, WB41, WW41, WW46, Wal48c, Wal62, WK70, Wei35, Wei54a, Wei71]. Distribution [Wei55b, Wei59, Wel64a, Wel47b, WAC62, Whe56, Whi58b, Whi59, Whi61b, Wij57, Wij59c, Wij68a, Wil32b, Wil38b, Wil46a, Wil69, WB64, Win47, Wis56, Wis57, Wol71b, Wol72, Wol49b, Wol50, Woo49, Woo72, Wyn69, Yua33, Zac66b, Zac67, Zei50, van61a, vM39, vM47, vN41, vS51, Dun65, Kol41, Ord68, Sch58, Sid67a, Sto63, TW46, Wol44a, vN42, Joh69]. Distribution-Free [Bel60, Bel64a, Bel64b, BD65a, BD65b, BMT66, BD66, BD67, Ber71b, Ber71a, BR54, BM58a, Dan54a, Dri52, Gut48a, MS69b, Wal62, Bir53a, DD64, Des70a, Dok66, Hol66, KvE64b, Rob44c, Sen66]. Distributions [AD70, AB67, And42a, And51a, And59, AB62, And80, Bak34, Bak35a, Bak40, Bak46, BMP63, BM64a, BM64b, BP66b, BP67a, Bar68, BG69a, Bar39, Bat55, Bat65, BB65, BM59a, Ber66a, Ber66b, BP68, Bil56, BP58, BH60b, BRN63, Bli62, BM63, Blu66a, Boa49, Boe66, BH65, Bon72, BK70, BBFH69, Bur64b, Cam37a, Cas50, Cha50a, CL56, CW58, CW59, Coh51, Coh57, Con66b, Con67d, Con67c, Con69, Cra66, Cra70, Cro58, CS72, Dan39, Dav72, DV55, Dem66, DK68b, Doo35, DG57b, Eag64, Ell62, Fel32, Fel43, Fel50, Fer62a, Fer63, Fer72b, Fin71b, Fis62, Fos52, FD65, Gle69, Goo58a, Gov66, Gre50a, Gri69, GK50b, Gup69, Gut57, Ham56, HR55a, HV65, HK64a, Har60a, Har64b, Her67a, HW58, Hoe65a, Hoe47, Hsu45b, Hsu46, Huf71, Isa71, Iye50, Iye61b]. Distributions [Jac69, Jam64, Jen32, Jog68b, JR51, Joh67c, Joh70c, Joh72, Kab68, Kac48, KZ66, KR56, Kar57a, Kar57b, Kar58b, KP60a, Kat60, KP57, KS72a, Kem68a, Kha67a, KP68b, Kni65c, KJB67a, KJB67b, Kul37, Lan58, Lan63b, Lan64, Lau60, Law65, sL71, LS48, Leh52a, Leh55, LPC70, Loy62, Luk39, Lut65, Mac48, Mad40b, Mad48b, Mar62, Mar63, MM69, Mar65, MP65, Mas70b, MH62, MZ65, McN70, MW69, Mij64, Mil67b, MvZ66, Moo43, Mor51, Mor53, Mou30, Mui70a, MG70, Mus71, Ney39, Noa50, NV68, Noé72, Nor66, Old38, Old52, OR64, OR66, Olm40, Par64, Pei43, Pei49, Pfa64, Pfa69b, Pil50, PAAJ69, Pil71, Pol34, PRR69b, Pro65, Rao54, Rao61a, Rao70a, Rei61, Rei63b, Rid61, Rie31, Rio37, Rob48a, Rob48d, Rol68, Rub60a, Rub60b, Rub61a, Rub61b, Rub62a, Rue71]. Distributions [RP65, San61, San62, Sch69c, Sch47, Sch34, Sch67b, Sco71, See72b, Sen70a, Sha65b, SH72, Shi71, Šid71, Sid63, SJ71, Smi48, Sri68b, Sri68c, Sta56, Ste69d, Sto65c, Sto70a, ST64, Sug69, SF69, Sug72, Szá72, Tan70, Tei54a, Tei54b, Tei60, Tei61b, Tei57, Tho72, Tho36, Tro71, Tru53, Tuc68, Twe57a, Twe57b, Twe67, Van66b, VA43, Vor72, Wal58a, Wal59a, Wal60, Wal46d, WK71, Wei58b, Wei58a, Wei58d, Wei60, Wij72, Wil32a, Wil38a, Zac70a, dW70, dW71, den55, Ala68, BT46, Fel48a, Iye51, Jam54a, Mil68d, TD67, Tho67, Tuk49a, Luk65, Owe67b, Rub66]. Disturbance [Gre54b]. Disturbances [Hil69b]. Divergence [Dav69b]. Divergent [Joh57a]. Divided [Bak32]. Divisibility [Kat67, Kel71, Rue71, Ste67, WK71]. Divisible [Ber64a, BS70, BC52, BSB53, BK58b, Con52a, DT57, Fis62, Fre57, Her72, Hin64, Huf71, Joh67a, Kap66, KS72a, KB57, Ore68a, Pol72, Pyk60, Rag60a, Rue70, Sha69b, Smi68b, Spr59, Ste68, Ste69d, Ste71c, Tuc64, Wol71b, Zel58]. Division [Dar53b, Dar62, SC72]. do [Ahl71]. Dobrushin [Sto68a]. Dodge [Lie53a]. Doeblin [Isa63, Lev71]. Doksum [Jog66]. Domain [BD72]. Domains [CL70, CL71, Hil69a, Roo72]. Dominated [Tei71]. Dominating [SR68]. Donald [Pea64]. Doob [Don52, GR70]. Doolittle [Dwy41b, Gri31]. Double [Cam37b, Owe53, Raj64, Tat53]. Doubly [Coh50, Kim49, SG56, SG58, SG69, Sin64, Spr53]. Down [Alb70, Lev52, Roy58, Tsu67, Der57, LW44, Olm46, Wol44a]. Downward [Mad48a]. Drawn [Cra31a, Dun31, Fel35, Fer36, GG50, Lad39, Len39]. Drift [Sim67a]. Driven [Eri71a, Eri72]. Droite [Dol69]. Dual [Ber69b, Kha64a, Rao60]. Duality [Pie70]. Duals [Hof63, Rag66]. Dubins [Sud72]. Dudley [Gän70]. Due [Abe33, Alt39, Ban44, Fis42, BZ44, Bus52a]. Duration [DL66, Was51]. During [vS51]. Dvoretzky [DS59, Ven66]. Dynamic [Bla62, Bla64, Bla65, Bro65, DM68a, Dup65, Har72, Mai66, Mai69, MV69, Sti63, Str66a, Vei66, Vei69]. Dynamical [NP66, Doo68]. Dynkin [Get66, Pyk63, Den69, Kun67].

E. [Coc44, Get66, Owe67a, Pyk63, Sch60a]. Earlier [Wei61]. Earliest [Ore64]. Ecology [Hoe43a]. Econometrics [Tin52]. Economic [Hur48, Wol67a]. Economical [Hal58, Hal59, Wei52]. Edges [Abe69a, Abe69b]. Editor [And51b, And52, Ano31b, Ano46b, Thi31b, Wil47, Wil48, Wil49, Wil50]. Editorial [Ano30, Ano31a, Ano33, Ano36a, Car30, Car32b, Car34a, Car34b, O’T34]. Editorship [Ano61e]. Edward [Doo68]. Effect [AK61, Bas64, Bha65, Bir50, Bra65, Bur60, Cam46, DJ51, Geb66, HW48, Hel70, PF70, RF70a, RF71, Wal47c, Wil65, Box54a]. Effects [BF64a, Dod41a, FM58, Fle62b, HA72, Imh60, Imh62, Law71, Por71, RC60b, RC60c, Vaj51, Box54b]. Efficiencies [Abr67, KH72, Klo65, Rao72, Yu71, Kat50]. Efficiency [Adi67a, And67, AJ65, Beh72, Bha67b, BJ68a, Bly58, Cap61, CLA62, Cha63c, Che52, CS58, Dix53, Ehr55, Gle64, HL56, HR55b, Hoe46, Hoe51b, Hoe58, Hoe61, JK62, Kem56b, Klo63, LF70, Lit72, MS67, Meh72, Mik63, Moo54, Pau47, Pfa70, Rag65a, Ram66b, Sen66, Sen71a, Sin67, Sta66a, Str58, Tsu68, Vil66, Wal47e, Wit60, Wol47b, Woo70d, Van63, Van70, Pur64]. Efficiency-Robust [Van70]. Efficient [AT63, Atw69, Bha67a, Mac59, McG68, Moo69, Nor38, Pic69, Str61, Vat46, Wal43b, Eli72]. Ehrenfest [Igl68]. Eigenvalues [Bin72, OK68, Pru64]. einem [Sch64]. Eisen [Pru71]. Eisenhart [GH61]. Either [Eav67]. Election [Dwy49a, Rio64]. Element [SM50]. Elementary [Bre58a, BE70, Doo44, GW68, KP68a, KHS72, KHS74, Mij61, Mij64, Moo68b, Pil64, PG68, Sho30b, Sho30c, Tho60, dW72]. Elements [AFLP57, Ben63, Dod39, Fis33a, Fis42, FG60, Gla59, GM69, Kap45, Ken39, Lan72b, Rak69, Rak70, Str66b, Thi31n, Wol44b]. Elfving [Stu71a]. Elimination [Abe33, DWW51, Rob36, Shr51, ZF64]. Ellipsoids [Alb66]. Elliptical [Tal63]. Emanuel [Smi65]. Embedding [AK68b, HF71, Mon72a]. Emigration [Rub62b, Rub63a]. Emphasis [Dre40]. Empiric [Bla55a, Blu55, Fox70]. Empirical [AF62, ABE^{+}55, Bru62, Csö65a, Csö65b, Dri52, Dun31, Dwa58, Fel50, Fin71b, Hoa67, Joh57b, JV71, JV72, Kou70c, KS72d, Mee72b, PS68c, Rob64a, Sam63b, Sen71c, Smi48, Ste71a, Tak71a, Tei55b, Fel48a]. Emptiness [Gan61, Gho60]. Emptive [Wel64b]. End [Cha72, Pin47]. Ends [FP37, Sid58b]. Energy [KQ72]. Engineering [Ros66a]. Enjoyed [Ste58b, Ste60]. Enlargement [Gut46b]. Entire [Lot42, Ram62]. Entitled [Hub63b]. Entomology [Ney39]. Entropy [Bir62, Bro63, Dym66, FS71, Goo63a, Ios65, Ios69, MP71, Mur72, Pit68, PRR67, PRR69a, PRR69b, PR71b]. Entry [Tay70]. Enumeration [Cox40, HF71, Rio62a, Rio64, WF57a]. Environment [Smi68a]. Environments [AK71b, AK71a, SW69]. Epidemics [Yan68]. Epsilon [PRR67, PRR69a, PR71b]. Equal [Ber72b, Bur63, Dod41b, Gil71b, Joh66, Sta38, SC72, Wal51, Woo40]. Equal-Tail [Wal51]. Equalities [tCM60]. Equality [And63b, AG64c, AG65, Bro39, Dem64, Gir68a, Gir68c, Gna59, Gna60, KP67, KP68c, Kri68, Law63, Lay72, Loh69, Mik62, tCM60, OS70a, Ram56b, SN68, Wil46a]. Equation [AR49, AR50, Bea37, Bla46, Den70, Fel41, Hac63, Hor35, JS68, KL56, Lot39b, McG68, Min31, Nan48a, Nan48b, Nan50, Ras48, Sam42, Sud71a, Wei65b, Wol47b, Kar61, Rei58, Rei59]. Equations [AR49, AR50, And59, And65c, Ber45, Bri72, Dav68a, DV67, Dwy37b, Eri71a, Gir39, Ham56, Hil31, Hoe41, Hot43a, Kin48b, Lon42, Lon44, MG55, Mic59, Mil68c, Mui70b, Rao61a, Rei49, Str72a, Tuc41, Wol54a]. Equiconvergence [Boy71]. Equicorrelated [OS62]. Equilibrium [BDF64, Bro70d, Bro71f]. Equimodal [Mou30]. Equivalence [Cam70, Cho69a, Day64, Day65, Eis69, Fel68, Fre66a, Fre66c, HF70, IW71, Kad66, Sac67]. Equivalent [Bla53, Fra51b, Ges70, Gor72, Kai67, Kie62, Sto61a, Var67, Tuk47a, Tuk48b]. Equivariant [Ber67a, HH72b, Zac70a]. Era [Ore64]. Ergodic [Aus67, Bax65, Ber68, Bre57b, Bre58a, Bre60a, Bur65, Chu61, Isa64b, Kei66, wK68, Kli69, Kre71, Met69, Par58, Paz70, Tei71, Mad71, Haj67]. Ergodicity [DBTL68, Ios72, Joh65]. Ergodocity [Sac60]. Erlang [Tak69a]. Erlangian [Sha65b]. Errata [Ano35b, Ano35c, Fel50, Gei39, How50, Iye51]. Error [Ahl70, Ano35b, Bak52, Blo57, Blo70, BS64, Dod36, DK72, Fra68a, GRS55, GW70, Gre50b, Hen34, Hen35a, Her72, Kin48a, Kot61, Lom52, Lon42, LR53, Min31, Nea71, Pal30, RC60b, RC60c, Rub58, Sat44, Sch61a, Sha55, Sny72, Thi31s, Twe67, Wal40a, Wei35, Wei54a, Wer32, WK69, WF57b, Won36, Zid69]. Errors [Abe33, Ban65, Box54b, Chu56, Cle71, Coc40a, DB71, Har70c, HW48, HS72b, Hoe40, Joh61a, KS68, Leh44, Lon44, Lor59, LL39, Nor40, SY66, SY68, SY70a, Sid60, Thi31f, Thi31d, Thi31h, Thi31i, Thi31k, Thi31j, Zid70]. Escape [Por64a]. Especially [Goo63a]. Esseen [Ger71, Kat63a]. Essential [DeG61]. Essentially [Ehr66, Sob53, Wal47a]. Establishing [Dyk70b]. Estimability [Har67b, Mil71a]. Estimate [BH67, BG68, Cha67, Cra43a, Fen59, Høy68a, Jos68a, Jur71, LQ65, Par57b, PR54, Wal48a, Wal49a, Wol49a]. Estimated [And65a, Car32a, Kim49, Sch72]. Estimates [Adi67b, Agn57, AR50, And59, And62b, AKKW69, Bah67, BF63a, BF64a, Bar49, Bic64, Bic65c, BB67, BH60a, BG47, Blo57, BLS64, Bru55, CL54, Chi56, Cla57, Coh65a, Coh66, Cra38a, Dav50, Dav51a, Dic68a, Dix57, Dow54, Dre40, Dre41, Fab64, Fer58, FR64, Fra57b, Fre63b, Fre65b, Gho64b, GMS46, GRS55, Gra54, Gre65b, Her72, HL63, Hoe51b, Høy65, Jae71a, Jae71b, Joh40b, Jos69b, Jos71b, Kac48, Kat61b, Ken72, Kim56, KvE72, Lan44, Lan72a, LS60, Le 70a, Mad48b, MM62, MM63, Mah63, MH69, MH70, MR66, MS64a, MS64b, Nea70, Nea71, Ney51, Par57c, Pet54, Pfa70, Pie70, Ram66a, Rod66a, Ros56b, Ros71a, SGO63, Sch70e, Sco50, Sea58, Set49, Sha55, Sid58a, SC63, Ste50, Str58, TK70, Tay53, Tho62, Van66a, Van69, Wal50a, Wij59b]. Estimates [Wij59d, Wil32a, Wol47a, Wol47b, Wol50, Kat50]. Estimating [Agg55, Agg66, Alb62, And51a, And80, Bha66, Bly59, Bos66, Car46, Cha56, Che53, CZ64, Coh50, DK68a, Fox70, GW70, GC64, Har59b, Har60b, Har61, Har64a, Jae72, Jos70, Kat63c, Mos62, NS51, NS52, Pol70a, Ric60, Rob68, Rob67, RW68, RC60a, RT59, Sam65a, San72, Sen66, Smi48, Sta66a, Wal57, Weg69]. Estimation [AS58, AJ51, AR49, Ban44, BF66, Bas55b, Ber67a, Bha67a, BP65a, BY68, BY69, BL69, Bir61b, Bla47, Blo70, BR63, BR69, BC68a, BC68b, Boe66, Bro47b, Bro70c, Bru58, Cam70, Cha72, CR51, Cha54, CJ56, CGJ67, Coh51, CS70, Cra65, CS72, Dal41, Dav51b, DeG59, DR63, DV55, Dem60, Dix60, DS69b, Dun58, Dun59, DKW53a, EM70, Eat70b, Far72, FT66, Fer58, Fie70, Fie71, Gei65, Gei67, GS66, GO69, Gil68, God60, GJ65, God69, Goo49, Gor41a, GS60, Gre66, Gre54b, Hal46, Hal52, HB65, Has42, Hea56, Hin65, HS72b, HL50, Hoe58, Hoe61, Hoo56a, HB66, HB48, Hoy68b, Hub64, Jee54, Jof67b, Jog67, Jon69, Jon70, Jos65a, Jos65b, Jos66a, KS68, Kar58a, KW52, Kie52b, Kie57, Kim46, KW70, KP69, Kni65c, Koo71, KOvE72, Kul39b]. Estimation [Kul63, Leh63c, Len69, LZ57, Man62, Man45b, Man46, MP65, Mas50b, Mat55, McC72, Mee72a, MG47, Mur61a, Mur63, Mur65a, Mur65b, OP58b, OP69, Par72b, Par60, Par62, Pat63, PJ70, Pau42b, Pau64a, Pau69, Pfa69a, Pic69, Pic70, Pla58, Por71, PS67b, Que56, Rad58, Raj64, Rao68a, Rao70a, Rol68, Ros56a, RV65, Rou69, RB53, Roy54, Roy56, Rub63a, Rub58, Sac70, Sac63, SY72, SA66, Sal66, Sam68a, San68, Sar54, Sar55a, Sar55b, SG56, SG58, SG69, SV69a, Sav47, Sch70c, Sch60b, Sch69e, Sch67e, See70a, See70b, SZ71, SQ71, Sim51, Sim71b, Ski68, Smi47, Spj68b, Spr56, Spr57, Sri68d, Sta66b, SW72b, Ste57a, Str61, Tan70, TG58, Tat59, TB64, Try58, Val68, Ven67c, Vil47, Vil61, Vil69b, Wah71b, Wal39, Wal48b, Wal58c, WL63, Wat69, Wat59]. Estimation [Weg70a, Weg70b, Weg71, WGL66, Wol46b, Wol54a, Woo72, Yen64, Ylv62, Zac66b, Zac68, den55, Van70, FW51, Roy70, ST45, Tuk47a, Tuk48b, Mid66, Sch70c]. Estimator [Bha62, DKW56, Fin66, Gaf59, Ges70, Hoe69, KW56a, RT69, Sen68c, Sha71a, Sri69, Ste59b, Sub66, Yak69, Zac67, Zid69]. Estimators [AD70, Ajg67, Bar70, Ber69c, Ber71c, Bli62, Bro66, Bro68, Bro71d, Cha70b, Eic63a, Eri70b, Far64b, Far67, Far68a, GR69, Geb66, Goo53b, Han68, HM67, Har69a, Har69b, Hil69b, Hoa71, Jen69, Joh51, Joh71, Jos71e, Jos72, Kie52a, Kru68, Loy69, Man69b, McG68, Moo69, OQ72, Pat64a, Rao61a, Rob58, Sam65b, SMR72, Ses66b, Sha64a, Sid63, SC71b, Str71a, Str71b, Str72c, Tor71, Vil66, Win72, Zac63, Zac64, Zac70a, Zeh66, Zid70, Zid71, Zys67, Tuk58b]. Evaluated [Ber37, Hen34]. Evaluating [Hoe40, Lar36, Str35]. Evaluation [AF62, Ber64a, Cam37b, CD62, Lie53b, Whi65]. Even [DM68b, DM70a, Ros70a]. Event [AT72, BEM59]. Events [CC46, Chu41, Chu42, Chu43a, Chu43c, CH45, Dvo49, Gei38, Gei39, Gei42a, Kin63a, Ric48, Sch61b, vS51, Han66a]. Everett [tCM67]. Every [Ste70]. Everywhere [BRN63]. Exact [Abr67, Bak40, Bon72, Bur55, Bur67, Con66b, Con67d, Con67c, Con69, DG57a, Ela62, Goo58b, Gov62, GH51, GS65, Har70c, HG66, Hea61a, Jon48, Kab62, Lie53b, Mal68a, MR71, Mil67b, Nea71, Pra58a, Sie69c, Wal59c, WD51, Whi61b, WK62, WW43]. Example [Dav67, FR68a, FW69, Spr61, Ste59c, Tsu68, Wat61]. Examples [Bri62, Cau68, Dem63a, Wal41c, Tuk57b]. Exceedances [Eps54a, GvS50, Har52, Sar57, Sar60]. Exceeds [Sim43]. Except [NR67]. Excess [Lor70]. Excessive [Get65, Get71]. Exchangeable [Ber70d, DJ68, Dha64, Dha69, HK65a]. Excursion [Ber72e]. Excursions [KS59, Kus66]. Exhibiting [Fis42]. Existence [Ahl70, BRN63, Bur59, Dan40, Eri71b, Fin71a, Han61, Lan72a, LS60, Led71, Leh52a, Luc67, Met69, Mil72b, Ney51, Oga59, PS66, Roo69, Sei63, SSF68, Sin63, Str65a, TBO71, Var59, Wij63, Zac67]. Exit [Por70]. Exogamous [NR67]. Expansion [And65a, And70, Bin72, Hsu48, Joh67c, Jos39, Oka63, Oka68, PW63, RP55, Sio71]. Expansions [Agn59, Bha71b, Cha63b, Dic67a, Dwy38a, Dwy38c, Eag64, Fri72, Gut70a, Hil31, HD68, Joh67d, Joh70c, Kem59, sL71, LPC70, Sug69, SF69, TK70]. Expectation [BB68a, Bla47, Bru65, Dyk70a, Fel35, Har47, Hoe51b, Hsu45a, Lah56a, Mad59, Rob65b, Thi31s, Tho36, Wal46a]. Expectations [Bur62b, Fri69, GP71, Jos71b, LR72a, MR69a, NV60, Pfa67a, Rob66, Str65b].

Expected [AC65, Cho67b, DS40, Dre41, DF66b, Far64a, Fer72a, Fis67a, Fis69a, Fis71, Fis72a, Fis72b, Hai61, Har62, Hoe53b, Hoe55a, Hoe60, KR66, MM69, Mar48, Per69b, Rob44a, Ste42, Ste45b, SC69, Tei56, Tei61c, Wal46b, Ylv65, dW72, Van67, vM39]. Experience [HL52, Shu69]. Experiment [BF63a, Hub63b, KF40, Lin56, Ols67b, Owe70, Rob68, Rob33, TD63]. Experimental [AJ58, Bak52, Bli46, BH57b, Coc40a, Coc43, DB71, DK72, Ehr55, Hag66, Hed70, Hot41, Hot44b, Jam57, KS66a, Kem53, Oga59, Pau52a, Pau62, Rob63, Rob64b, Wyn70]. Experimentation [Der57]. Experiments [Alb61, Atw69, Bir61a, Bla53, BK56, Che59, Coh58, DeG62, EZ61, EZ63, Ehr66, EZ67, Fel72, Fre58, Gra68, HSF72, HL62, Kis45, Len69, Meh64, MS67, MS69b, MD68, Nig70, Nig76, Ric60, SB57a, Sen71a, Sha58a, Sha59a, Sha60a, Sri68a, Sto59, Sto61a, Tor72, WK56, WG64, Zac63, Zac68, Kem64]. Explicit [KS47, She69]. Explicitly [Sam42, Sim71b]. Exploring [BH57b, GGH59]. Explosive [Rao61a, Whi58b, Whi59]. Exponent [Mij72]. Exponential [AD70, Alb56, Aro48, BW64a, Bas64, Ber72c, Ber73, BP68, BP65b, Bou68, Coh65b, Cra66, DV55, Den72, Des71, Dos69, ET68, ET53, ES54, Eps54b, ES55, Fer62a, Fer63, Fer64, Fer72b, Gre49, Gru36, Har61, Har64a, HG66, Hoe69, Jos69b, KT60b, Kee51, Kim49, Kni65c, Koo61, KQ72, Kot61, Kri71b, Kul63, Kup58a, Kup58b, Lau63, MT67a, MT67b, McG68, MR66, MS71, Nad60, O’T33a, Pau41, Rid59a, Rid59b, Rid61, SA66, Sal66, SGO63, Sch69f, Sid63, Spj68a, SW72b, Ste67, Sto67, Tak69b, Tan64, Vil47, Wat71a, WK62, Mar61b]. Exponential-Class [Kup58a, Kup58b]. Exponential-type [BP68]. Exponentially [Hai61, KA64a, KA64b, Wij71]. Exponentials [Bar69]. Exposition [Cra36a]. Expository [Mol31]. Expressed [Thi31i]. Expressing [Mar61a]. Expression [Har69a, uD54, ZD59]. Expressions [Ekm59a]. Extended [GH69, Har65a, Hey69b, Hin64, Law33]. Extension [Bar50, Ber62a, Bla56, Bla58a, Car50, Con67b, DJ52, Day66a, Day66b, Don52, GG58, Gir58, Hey70, HM68, Koo68, Le 55, Man53, Nád69, OS70a, Tsa54a, Wal43a, Wal47d, Whi51, da 45, Gar59, HS70c, Nas54, Ven67a]. Extensions [AG44, AG64a, BD66, BM61, Cho69b, GWA72, Háj61, Han59, tCM67, van61b]. Exterior [Dod38]. External [Dod37]. Extinction [AK71b, Eri71b, Goo68, Smi68a]. Extrapolation [DM70b, Hoe65c, Hoe66, KW65, Lev66b, Stu71b, Tre65]. Extrema [Hoe55a]. Extremal [Dwa64, GK50a, GH51, GP67, Mor51, Sen72c]. Extreme [BN63, Ber63b, BH59, Dar52, Dar53a, Dav63b, Dix50, Dix51, Eps49a, Eps49b, Eps51, GG51, Har70a, HD54, Igl72, Kim56, Koz56, Lam64, Lie53b, Loy65, McC64, Mor51, Sug70, Wat54, Wel72, Zid69, Mos48a]. Extreme-Value [Kim56, Lie53b]. Extremes [Gum46, New64, Pic68].

F. [Ore64, Rio62b]. Fabins [Get66]. Facilitate [Bai31, Res62]. Factor [BH57b, DB71, Kem56b, NP66, Owe70, Rub68]. Factorial [And72b, BF63a, Ban70, BS64, EZ61, EZ63, EZ67, Gla59, Gul72, Iye58, Iye61a, Kem48, KZ62, Len69, Mar65, MS69b, PF72, Pre69a, RF70a, Sch60c, Sha58a, Sha59a, Sha60a, SC71a, SA71, WH65, WK56, Zac63, Zac68, Zel58, ZF64].

Factorials [BK58b, KB57, Rak69, CT56]. Factorization [Lah58, Tei54b]. Factors [Bow46, Bra65, HH72c, Imh60, Kab68, Wel64a]. Failure [BP66b, BP67a, BP69, BD69, Dok67a, Dok68, Fer72b, MP65, Rao70a, Bic69c]. Failure-time [Fer72b]. Fair [Fel45a, Mv67]. Falling [AM67, Bak49, Rio65, Wei61]. Families [BP66a, Bar68, BG69a, BG69b, BL69, BP65b, BBFH69, Cro58, Dem69, Den72, ET68, Eic63a, Eic63b, Fer62a, Fer63, Fre58, GW59a, HSF72, Jog68b, KS72a, Leh47a, Leh55, Mar69, Pfa64, Pfa67b, Pfa69b, Sch69f, Set64, Set70a, Spj68a]. Family [Bar70, Coh65b, Dos69, Har66a, Has63, MT67a, MT67b, Mes67, Nem71, PR71a, Sav57b, See72b, Sto67, Wat71a]. Fatou [Sud71a]. Favorable [Chu42, Leh52a, Rei61, RW70, Tru70]. Features [Zoc35b]. Federalist [Wat66]. Feinstein [Blu67a]. Feller [Spi67, Str72a, Wal70b]. Ferguson [Wij68b]. Fertility [Bak52]. Few [HH72c]. Fiducial [Bar39, Bri62, Dem63a, Fis39, Fra61, HB66, HB67, PB71, Ste59c, Wil38a, Zac70a, Tuk57b]. Field [Abr69, Bak52, BF64c]. Fields [Bru69, Bru71, Eav67, Eav69, Kni72, LR72b, Rak69, Rak70, SB57b, Ghu68, Won69]. Figures [Tuc41]. Filling [Ney62]. Filtered [New68]. Filtering [Cle71, Kal59b, Sny72]. Finding [DR67, Mee71, Pow30, Vei66]. Fine [Nii72, SSW66]. Finetti [Fre62a, Fre63a, HS72a]. Finite [Agg59, Alb62, ACH64, Bar35, BP66c, Bax62, Beh61, BK57b, BBT58, Bla64, Blo71, Bou68, Boz71, Bro65, Car67, Car68, Chu43a, Chu43c, Coh70, Coh72, Cov69, Cup69b, DR67, Dha63a, Dha63b, Dha65, Dha69, DN70, Efr65b, Eri70b, Fin62, Fis69b, Fis71, Gho60, Gil59b, GJ65, God69, Háj64, HH72b, HH43, Han68, Hed70, HC70, HY64, Hol72, Jos65a, Jos65b, Jos66a, Jos67a, Jos68a, Jos71b, Jos71e, Jos71f, Jos72, Kal59b, Kin62, KPS59, Lah56a, Ley67, Mad48b, Mil61a, Mor52, Mor72, O’T34, Pet69, Pie40, Pra58a, Rag70b, Rak69, Rak70, Rao68b, Rio62a, RC60a, Roy70, Sam68b, Sch72, Sco71, Smi64, SP65, SA57, Spi72, Sri69, SB57b, Tei63, Tho63, Tsa68, Van66d, Van66b, Van66c, YS68, Yak69, vS51, Gil69]. Finite-Stage [Bla64]. Finite-State [BBT58]. Finitely [Luc67, Ols68, Pyk61b, Sud69a]. First [Ano35b, Bak35b, BB60, DS53, DL66, Eze30, Gan61, Gol71, Han43a, Haq70, Hey66, KS62, Kei63, Ken72, Kin61b, Kin61c, Lam60, Mil67a, OS69, Pin61, Roo72, Roy53b, RM71, San47, She67, She71, Ski68, Sle61, Teu68, Wei35, Woo70c]. First-Passage [Lam60]. Fisher [Bir65, Aro41b, Bah64, Ban67, Bir64, Bos49, EO72, Fra53c, HD68, MB66, Pot63, Ram66a, RSS67, Sch43b, Sch44, Wal49e]. Fit [Bha61d, Bil56, Bir53b, Blu66b, Cha58, Cha65, CL54, Col32, Goo58a, KKW55, Kim47, Kim49, Mat55, Smi48, Tsa55, VN67, Wei56c, Wei57b, Wei57d, Wei58d, Wei61, AD52, Bir53a, Coc52, Ros62a]. Fitted [AA50, Ant62]. Fitting [Ail70, Bai31, Gue51, Kim40, O’T33c, Pin47, Sam43, Too34, Wal40a, Wil30]. Fixed [Alb66, Bak49, BR69, CR65b, FM58, Gle65, Gle66b, HH63, Imh62, Joh71, Kha69, Mah67, Mv67, New60, Rob64b, Sim68, Sri71, Sta66b, Str72d, Zac67]. Fixed-Cycle [New60]. Fixed-Sample [Mah67]. Fixed-Size [Gle65, Gle66b].

Fixed-Width [CR65b, Kha69, Sim68, Sri71, Sta66b]. Flett [AH71]. Flexible [Jae71b]. Flood [Gum41, Kim42]. Flow [Wan68]. Flows [FS71, Gum41, Kim42]. Fluctuation [Dwa62, Hey69a, HP63a, HP63b, HP64, Imh71, KvE64b]. Fluctuations [Rao68c, Wil70]. Folded [Joh70a]. Follow [Coc40a]. Following [Mus71]. Follows [Kot64]. Force [Dod36]. Forces [Bor67]. Forecasters [HB71]. Form [Alt37, Bal67, CMPT60, DF65, Fra56b, Mas71b, Mas72, Mey69, Pac55, Rob48b, SK61, SK63, Sha70a, SC72, Tho36]. Formal [Por71, Thi31q, Zid70]. Formative [Men52]. Formed [Wal47e]. Forms [Box54a, Box54b, Cra43b, Cra47a, Eic66, Gei57, GS55, GM69, Gur53, Gur55a, Gur62, HW71, Hog51, Kac45b, Kad68, Koo42, KJB67a, KJB67b, Lan63b, Lut65, Mad40a, Mad40b, Mat49, RP49, Rub63b, Sat41, Sha63, Sha68, Sid65, Tan64, Tho38, Var66, Var68, Zys67, GM57, Kaw50]. Formula [Bac38, Chu43b, CH45, Coc40b, Goo61a, Gre44a, Gre47, NR67, Nea70, Nic54, Tak69a, Whe31, Whe35, Wis57, vS49, vS50, ST44, vM42]. Formulae [Ban49a, DT64, Gur48, Gut38, Ito56, Ito60, Kem70, Sny72]. Formulas [Ano35c, Dwy40a, Gri31, Hsu44, Hsu45a, Jon52, O’T33b, Obe47, Pei43, Pei49, Rio49, Rob35, Rob37a, Sar49, Sug71, GL46]. Formulation [Bur32, Goo63a, Str72d]. Forsäkringstidskrift [Ano31c]. Found [Gou63]. Foundation [Bol34]. Foundations [Bir61a, Blu67a, vM41, RS72a]. Four [Che69, Dra60a, Dra60b, Har44, Har45, Her64, Wal46d]. Four-Group [Har44, Har45]. Fourier [AA50, RP55]. Fourth [Aro48, O’T33a, vS50]. Fraction [Obe47, Wol46a]. Fractional [BF64a, BF66, Ban66a, BF68a, BC57, Len69, PF70, PF72, Pat62, Pre69a, RF71, Rie31, Zac64, Zac66a, Zac68]. Fractions [Aro59, BF63a, TW79, Aro41a]. Fragmentary [Wil32a]. Frank [McC65, Dar65a, Joh69]. Franklin [Sal67]. Frederick [Wat66]. Fredholm [Suk72]. Free [Bel60, Bel64a, Bel64b, BD65a, BD65b, BMT66, BD66, BD67, Ber71b, Ber71a, Bic69a, BR54, BM58a, BKR61, Bur72a, Dan54a, Dri52, Eat70b, Gut48a, Jon69, Kru68, MS69b, Wal62, Bir53a, DD64, Des70a, Dok66, Hol66, Kei63, KvE64b, Rob44c, Sen66, Thi31g]. Freedom [Cas50, KA65]. French [Dol69]. Frequencies [BW57a, GG50, Hoe45, Sch60c, Wal40b, Wal41b]. Frequency [Ano31a, Bak34, Bat31, Bur42, Coh51, Cra36b, Cra36a, DS40, Dod37, Efr65a, Fin46, Goo61c, Gre41, Hun39, Ker63, Kim49, Kim50, Kir35, Lar39, Mit58, Mou30, O’T32, Pol34, Rie31, Rie36, Rio37, Ste42, Tor52, Yua33, Zoc34, Zoc35b]. Friedman [Fre65a, Fre66a]. Fuchs [Gil71b]. Full [RP67]. Fully [KS65, KS72b]. Function [Agg55, Aro47, Aro48, Aro59, Aus66, ABE^{+}55, Bah60a, Ban45, Ban49a, Bar69, Ber70e, Bha66, BS72, Bla55a, BG68, BKR61, Bro40a, Bru62, BR58, Bur59, Bur62b, But69, Cam46, Cra36b, Dan39, Dha63b, Dur68, DKW56, Dwy37a, Efr65a, Eic70, Ela62, Far62, Far66a, Far66b, Far67, Far72, FM58, Fox70, Gor72, Gre50a, Gre50b, GW65, pH69, Hat49, Hen34, HS55, Hoe55a, Hot41, Kat55, Kei69, Kha65a, Kha65b, KW52, KW59a, Kim47, Kin48a, KP66b, Lev52, LQ65, Mad59, Mas50b, Mei72, Mit58, MH69, MH70, Mur71b, OQ72, O’T33a, Oka63, Oka68, Par57b, Par62, PB70, Pic69, Pie40, Pil68, PG68, Pit56, Rea72, Ros56b, Rou69, Roy53b, Sal30a, Sal30b, Sch69e, Sch72, Sim51, Ste71a, Sto72, Swe70, Thi31k, TW79, Wal42b, Wal46c, Wal49e, Weg70a]. Function [Wei55c, Wen68, Wer32, Whe56, Whi70, WD39, Wil36, Ylv68b, Zim72, vM39, Aro41a, Jag63, Kol41]. Functional [Ber64a, IW71, Ken71, Ras48, Sto69b, Thi31j, Vil61, Wyn69, Thi31h]. Functionals [Cla70, Cla71, Fre67d, Get65, Get67b, McN70, NV60, Por64b, Saw72]. Functions [Ahl70, AK68a, AG64b, AG64c, AG65, Ano35b, Bah54, BL55, BC54, Bat34, BP66c, BPSW70, Bec42, Ber58a, Ber58b, Bha70, BS69, Bir62, BT51, Bir53a, Bir53b, BG46, BK57b, Blu55, Blu67b, Boa67, Bou68, Bri38, Bro47a, Bro70a, Bur42, CG66, Car67, Cha63b, CGJ67, CB48, Coh40, Con67a, Cra70, Csö65a, Csö65b, Cup69b, CL70, CL71, Cur42, Dan40, Dha63a, Dha64, Dha65, DN70, Dix53, Dod30, Dod37, Dos69, Dri52, Dwa58, Eav67, Eri70a, Fab69b, Fel65, FR68b, FR69, FR70, Gan66, Get65, Ghu54, Ghu55, GO69, Gil59b, Gre32, Gre49, GAM64, Gup69, HH72a, Han67, HK64b, Har62, Har70b, Has63, Her48, Hoa67, Hoo56b, Hsu45d, HK44, Ifr65a, Isa72a, Kac48, KT60b, Kei72, Kha66, KP68a, Kim46, Koz47, Koz51, KR66, Kru54, Kud66]. Functions [Kul34, Lah58, Lam71b, Lam71a, Lan71, Le 55, Ley67, Luk57, Mac71, Mal54, Man69b, Mas51a, Mic59, MB66, Mij61, Mij64, MS72b, Moo68b, Mud65, Mui70b, Neu64, Nor37, O’T31, O’T32, Ore68a, Par72b, Pet69, Pil64, Pit63, PW67, Ram62, Rod66a, Rog59, RP55, RM61, Rub62a, Rub56, RM68, Sam53, Sch69b, Sea56a, SQ71, Sho69, Sho72b, Smi39, Smi47, Sob53, Spj72a, Spr53, Sta71, Ste71a, Ste51a, Sti69, Sto65a, Sto65b, Sto67, cS67, Tak71a, Tat59, Tay53, Tei55b, Thi31g, Thi31i, Tsa54a, Tsa68, Tuc64, WW39, WW41, Wal47a, Wal49c, Wei35, Wei57c, Wei58b, Wil69, Wol71c, Wol71a, Wol71b, Wol71d, Wol72, Wol42, Wol51, Zyg47, dW72, vM47, Get71, Sch58, Tru47, Cup69a]. Fundamental [Alb47, Alb48, Ano36a, Bah58, Bla57b, Blo49, CS52, Chu43c, DW51, Gri31, Har47, Loè50, Mol46, Rub63a, Wal46a, Mad37]. Fundamentals [Ano30, Car30, She36, Moo48]. Further [Chu43a, Dem63a, Dix44, DH68, Fre57, Fre61, Gei45a, Hot43a, PJ70, Pfa69b, Roy54, Roy56, Sei55a, Sen68c, SW72b, vN42]. Future [DT67, Lig70, Tuk62a, Wal57, Tuk62b].

G [Bar47b, Bro70c, Dal64a, Get66, Hea61b, Igl72, JS72, Kie64, MR64, Ore68b, Sch70b, Wat71b, dH70]. G. [Der66b, Pla71]. Galton [Ath70, BH67, Eri71b, Goo68, GH68, Hey70, Jof67a, KS66b, KS66c, Sen68e, Sti66]. Gambling [Ore64, Sud69b, Sud71b, Sud71c]. Game [BDF64, BF64b, Cov66, DL66, HS72a, fH69, fH71, Sac72, TD67, Tho67]. Games [Bañ68, DWW51, Fel45a, Fer70a, Gil69, Hur48, Ore64, Rob61, Sid62b, Sob71, Tru70, Van66b]. Gamma [Ber58a, Ber58b, Cha56, Coh40, Gri69, GW65, HM67, Kha65a, Kha65b, KA64a, KA64b, KP51, KP60b, Lah54, Luk55, Mal68a, Rog64, Sta62, Tak61c, WGL66, Jam54a].

Gamma-Type [KP51, KP60b]. Gamma-Variables [KA64a, KA64b]. gap [BB61]. Gardner [Fab69a]. Gatlinburg [Ano58m]. Gauss [Cam70, Eat70b, Kru68, Lah59, Won69]. Gaussian [Dwy60, Rub66, Ber72b, Ber62b, Ber70e, Ber71d, Ber72e, Bha71a, BS72, BE72, Cap64, Che69, Cho69a, Doo44, Dud65, Dwy58, DM70b, Eav67, Eis70, Eis72, Fel68, FK72, Fra51a, Fra53a, Gem72, Goo63c, Hin65, HS72b, Jam70, Kac45b, KS47, KS59, Kha62a, Kha65c, Kha65d, KQ72, Kul69, Lam65, LPC70, Mar70, Mar72a, NR62, Par70, PJ71, Pit63, PRR69a, PRR69b, Pre72, Qua68, QW71, QW72, RC72, Rao67, Rao68c, SR68, She66, She71, Sle61, Twe57a, Twe57b, Var67, WK70, Ylv65, Ylv68a, Zim72]. Gaussianating [Hoy68b]. Gautschi [Blu60]. Geiringer [Lin66]. Gene [Gei45b]. General [Ahl71, Alb54, Bar47a, Ben62, Bha66, Bra65, CJ58, DJ51, Dem60, Dem63b, Dod40, DT64, Fel43, Fis69a, Fle62a, Fré47b, Gre38, Gre41, Ham71, Hoe55b, Hot40a, Isa63, Jai69, Joh65, Joh72, KK62, Kem70, Kem63, Kem68b, Kha69, KW56b, KvE72, Leh51b, Neu67, Pra59, PS69b, PS69c, Rao54, Ray57, Sri68a, Tak62, Tan70, Wal49g, Wij67, Wil36, da 45, van61a, GM57, Sam43, AT63, Rei64]. Generalisation [Sha59c, Sha59b]. Generalised [Con66a, Kha64a]. Generalization [Bax62, BRZ47, Blo49, Bus52a, Cam48, CS52, Chu43b, Ell72, Gei40, Gre44a, Gre47, HK68, HS70a, Hot31, Hun55, JS68, Kab65, wK68, Mau59, Mil61b, Rag60a, Shu72, Sta62, Tuc59a, Wal38, Wol54b, Wor60, Ylv61, Zys63, Tuc59b]. Generalizations [Agr66b, Bax65, Dav67, Deh71, Dic67b, Dod41b, Dwy58, Dwy60, Wal45b]. Generalize [Fre62a, Fre63a]. Generalized [Bag65, Ban66b, Ban69, Bas67a, Bas68, Bat55, Bec42, BM63, BBFH69, Car68, Cha35, CL56, Che67a, DR72, Dav68a, Day66a, Day66b, DeG61, Dem59, Dur71, Dwa59, Eat67a, Eav69, Far68c, Fra51a, Fra53a, Gov66, HD68, Hsu38, Ito56, Ito60, Kem56a, Ken48, KW57, Kin61b, Kin61c, Kot65, Len69, Lev71, Mac48, Mal68a, Mui72a, Nor37, Por65, Pot63, Rak70, RT69, RP67, Rog64, Sac63, Sat42b, Sat42a, Sho66, Sio71, Sri69, Ste51a, Sto67, Sug67a, Tha65, Tra68, Wer32, Wit60, Zac64, Der66b]. Generated [BP65b, Dem68, Wei58c]. Generating [Cur42, Fer58, Han67, Kei72, Koz47, Koz51, Mar61b, Neu64, Pil68, Sto65b]. Generation [Sho66, Wyn70, BM58b]. Generations [CS39]. Generator [Alb62]. Generators [RF71]. Genesis [Cas50]. Genetic [Wat61]. Genetics [Kim57, Lev49b, Wat62, Wat63]. Geometric [Cam38b, Cra66, GK50b, Kem68b, Klo70, Nor40]. Geometrical [CE67, Sei61]. Geometry [BK60, Men52, RC62, Tau62, Fie68]. George [Ano72, Gre50a, Mos48b]. Georges [Dug61]. Germ [Kni72]. German [Sch64]. Getoor [Dym70]. GI [Igl72, JS72, Pak71, Sch70b]. GI/G/1 [Igl72, Sch70b]. GI/G/k [JS72]. GI/M/1 [Pak71]. Gil [Wen61]. Gini [Lom52]. Girshick [BB55]. Given [Bru65, Bur59, Chu46b, GG39, Kim49, Kul68a, Kul68b, Rie31, Rob66, Roy53b, SM50, Str65a, Wei55b, Wic70, Wol46a, vM39]. Glivenko [Tuc59a, DeH70, Deh71, Dud69c, Gän70, Kli67, Shu72, Tuc59b, Wol54b]. Global [Agn57, Agn59, Koe70]. Gnedenko [Gre64]. Goal [Sud69b]. Gods [Ore64]. Good [Per70]. Goodman [Leh66a]. Goodness [BMT66, Bha61d, Bil56, Blu66b, Cha58, Cha65, CL54, Goo58a, KKW55, Ros62a, Smi48, AD52, Coc52]. Goodness-of-Fit [Cha58, Cha65, Ros62a]. Graduate [Jai70]. Graduated [Dod41a]. Graduation [Dod39, Jor32, Jor59, Key38]. Graduations [Dod41a]. Gram [Aro37, Bak35a, DG57b, Sam43]. Granger [Wol67a]. Graph [Bla68, Hof64, Moo63a, NR67]. Graphic [Eze30]. Graphical [BZ49, WG64]. Graphs [Abe69a, Abe69b, AFPR59, DG57a, Gil59a, Rio62a, ZS60]. Greater [BR50, BR51, Kim44]. Green [Wen68]. Greenberg [Get66, MR64]. Greenwood [Dav63a]. Grenander [Kal65]. Gretchen [LeV66a]. Griffins [Hed70]. Group [Ahl71, Ber67a, BC52, BSB53, BK60, BK58b, Bri63a, Con52a, Cra65, Fre57, Har44, Har45, HF70, Hin64, Jam54b, Joh67a, Jon52, Kap66, Kat52, KB57, Kum65, Rag60a, Spr59, Sta63, Tuk47b, Zel58, Tuk49a]. Group-Testing [Kum65]. Grouped [Bha67a, Pie40, Sen67]. Grouping [Abe33, Alt39, Dwy42, Gue51, SE43]. Groups [BE72, BP65b, Bru69, Dwy37b, Eis72, Pfa72, Pit65, WB41, Wal44b, Wen57, vM45]. Growth [DS69b, Mil68a, Sin70, Sin71, Sto66a, Sto69a, Wil36, dHH72]. Gruel [Luk65]. Grundlagen [Pyk63]. Guide [Dav63a].

H [Dav63a, Man45a, Owe67a, Pra66a, Spi66]. H. [BZ44, Gil64a, Sav71, dH70]. Haar [Sto65c]. Haight [Joh69]. Hajek [Hoe68, Sen71b]. Half [Por64a, Thi31j]. Half-Invariants [Thi31j]. Hammersley [Mul66]. Handbook [Joh69, Mos65, Pea64]. Handscomb [Mul66]. Hanover [Ano40f]. Harald [Dar65b, Fel47, Ylv68b]. Harmonic [Cra70, Lam71a, Nor40]. Harris [Wel65, tCM67]. Hartley [Dav63a, Owe67a, Ram56b]. Hartman [Sto70b]. Hatanaka [Wol67a]. Hausdorff [Hen72]. Haven [Cra47b]. Having [Dan40, Joh66, Ken39, PN71, Rag60b, Zac66b, Zac67, vS50]. Hayes [Kri71a]. Hazard [BMP63, BM64a, BM64b, HK64a, Har70b, Mur65a]. Hemmerle [Har71a]. Henry [Cra44b]. Herbach [Gau59]. Herbert [Had63]. Heredity [Ber42, Gei44, Gei45a]. Hermite [Gre32]. Heterogeneity [DJ51, Shr51, ZF64]. Heterogeneous [Sen70a]. Heuristic [Don52, Doo49, God61b, Roy53a]. Hewitt [HS70c]. Hidden [Dod30]. Hide [MP71]. High [Alt39, Ber71d, Ber72e, Bra52, Dan71, Dem58, Qua68]. Higher [Wei53a]. Highest [BEM59]. Hilda [Lin66]. Hill [Rob50]. Histogram [Alt39]. History [Ore64]. Hit [Fra51a, Fra53a]. Hitting [Cra70, Joh72, Por67, Por68, Shi71]. Hodge [Hil61]. Hodges [Fin66, Hil60, Høy65, JK62, Klo59, Klo64]. Hoeffding [Fra56b, NCC65, Ros67]. Hoel [KW65]. Homogeneity [Bak41, Cha63a, CS71, Dar52, Dar53a, Hoe45, Kul71, SN69, SN70, vB31]. Homogeneous [Bak32, Bru69, Bru71, Eav67, Ios65, Ios69, Joh57a, Kei63, Loh69, Luk52, RR64b, Rub62a, Sha69a, Sha70d, Wei58b, Won69, RR66].

Hopf [Kem63]. Hornich [BZ44]. Horvitz [Ajg67]. Hotelling [Ban69, Bic65a, Ban65, Ban66b, Bic65b, Con66a, Dav68a, Dem40, GKS63, Hsu38, Ito56, Ito60, Moo46, Mui72a, Sio71, Ste56, Wau45]. Household [Bol34]. Houses [Sud69b]. Howard [Jai70]. Hsu [Erd50, Erd49]. Huber [SY72]. Hull [Kin63b]. Hulls [Fis69c]. Hultquist [Ban70]. Hunter [McC65]. Hybrid [Coh65b]. Hydrodynamic [Bor67]. Hyper [Han68]. Hyper-Admissibility [Han68]. Hyperadmissibility [Jos71e, Jos72]. Hypergeometric [Har65a, Ifr65a, Kni65c, Lar39, Mui70b, Nic56, Rio37, vS49, vS50]. Hyperplanes [DK68b]. Hypersphere [CE67, Ham50, Lor54]. Hypotheses [AG59, AG60, All53, BP67b, BW49, Bir55, Bro45, Bro40b, Cha51, Che49, Con66b, Cou44, DeG60, Dem69, DL70a, DL70b, Dic71, DL72, DKW53b, Dwa57a, EZ67, FP69a, Gho54, Isa51, Kat61a, Kup58a, Kup58b, Leh47b, LS48, LS49, Leh50, Leh52b, Lig71, MT67a, MT67b, MG55, Mee70, Mee71, Ney38, Pau63, PAAJ69, Rei61, Ros62b, Ros62c, Ros63a, Sam63b, Sav57b, Sch42b, Sim67a, SW49, Ste51b, Tan70, Tsa54b, Wal39, Wal41a, Wal45a, Wij71, Wil38b, Wil65, Wol42, vM43, Gna67]. Hypothesis [Che52, Con67c, Cov69, Dan40, DS63, DJ51, Dix40, DKW59, Fed68, Fer36, Goo63a, GAM64, Hoe55b, Imh62, Kha64b, KvE72, Len39, Man60, Mat43, PS69c, Ray57, RP67, RK56, Seb64a, Sim43, Sim67c, Ste45a, SF69, Suk58, Tai66, Wei53b, GM57]. Hypothetical [Bru62].

I. [Moo64]. I.I.D. [PR70]. Idea [van61b]. Ideas [Ore64]. Idempotent [Ban64, GM57, GM69, Loy66]. Identical [Kru68]. Identically [CR63, CT71, Hoa71, HS55, KA64a, KA64b, HW58, Koo68]. Identifiability [BK57b, Gil59b, Tei61a, Tei63, Tei67, YS68]. Identification [Car67, Gup72, Jee54, KR50, Rub58, Yak69]. Identifying [Sim71a]. Identities [Dic68b, Har66a]. Identity [Alb47, Alb48, Bah58, Ban45, Blo49, BC57, Har47, Joh66, Mil61b, Sen72b, Wal46a]. if [Sta72b, Wal40a, vM39]. IFRA [Dok69a]. Ignorance [ACH64, Efr65b, Tho38]. II [Gau59, Kul56b, Rob47, Rub61a, AK71a, BM64b, Bic69c, Blu66b, Box54b, BC64, Car32a, Cra42, Dem61, DH69, DG57b, Far66a, Far66c, Far69b, FP64, FR69, Fra53a, Fre65b, FP67, Get66, Get71, Ghu55, Gir65b, Gir46b, GH61, HKcY68, Her59, Hey69a, Ito60, JV72, Jos65a, Kar57a, Kie61, KJB67b, Kul56a, Leh57b, Mad49, Pyk58, Rei59, Rob45, Ros72b, RG59b, RC60c, Rub60b, Sar55a, SG58, Set70b, Shi71, Sin71, Thi31d, Tho67, Tuk47a, Tuk57d, Twe57b, WK71, Weg70b, Whi59, ZF64, Der66b, Bar47b]. II-Type [Her59]. III [Boy58, Bro40a, Che65a, Far66b, Far66d, FR70, Gir46b, Jos65b, Kar57b, Kul36, Mad53, Ols38, Raj53, Rub61b, SY70a, Sal30a, Sal30b, Sar55b, Thi31e, Tuk48b, Tuk57e, den55]. Image [Wic70]. Imaginary [Coh40]. Imbedded [Gav59, Ken53]. Imbedding [Kus69]. Immigration [Dur71, Fos71, Rub62b, Rub63a, Wei72a]. Impartial [BG52]. Implications [BP64]. Impossibility [Shr50]. Improper [Sch70c]. Improved [Dwy44b, Fab67, Kag71, Maj62a, Nea70]. Improvements [Hot44b, Wal46b].

Improving [Goo53b, Mye34]. Inadmissibility [Bro68, FP69a, Jos67b, Per70, PS71, Rea72, SC63, Zid69, Zid71]. Incidental [KW56a, Pfa69a]. Include [Obe47]. Included [Bak52]. Including [NV68]. Income [Swe70]. Incomplete [Agr66a, AJ56, Aro59, ABE^{+}55, Ban45, Ban49a, Ban48, BP69, Bha60, Bos49, BC52, BSB53, Cau68, Con52b, Con53, Cor60, Cox40, DN62, Fre57, GW59b, Gre66, GW65, Han61, Han65, HST53, Har69b, Hin64, Hof63, Joh67b, Kag71, Kem56b, Kha65a, Kha65b, Ksh58, Ksh59b, Maj62b, Maj53, Man69a, Mes67, Mik60a, Mik60b, Mot58, Mur61b, Nai69, Oga59, Oga63, OI65, PS67b, Rag66, Rao60, Rob70b, Rob71b, Rob72b, SY70b, Sei66, Sen71a, Ses66b, Sha59b, Sha60c, Sha66b, Sha67, Shr50, Shr65, Spr55, Spr56, Spr57, Spr59, SM66, Sur68b, Tho55a, Tho56, TB64, TW79, Wij58, Yal68, Zys63, Aro41a, Rao65]. Incompletely [BBH56]. Inconsistencies [Dem63a]. Inconsistency [Bas55a]. Incorrect [Ber66a, Ber66b]. Increases [Wal48b]. Increasing [BD62, Dok67a, Dok68, Efr65a, HK64a, Har70b]. Increment [FK72]. Increments [Ber70e, Ber72e, Hey69a, Mil69, Mon72b]. Indecomposability [Kat55]. Indecomposable [BBT58, KS66b, Tho63]. Indefinite [Gur55a, Gur62, Sha63, Sha68]. Indefinitely [Wal48b]. Independence [Ahm61, AG64b, BD67, Ber71d, Bha59, Bhu64, BKR61, Con67c, Cra38a, Cra43b, Dal40, GM62, Gum46, HK64b, Hoe48b, Hog63, Hsu46, Jog68b, Kac45b, Kaw50, Kha59, Kha61, Kha62b, Kha73, Kon56, Kon58, Lah56b, LLN60, Lan65, sL71, Luk52, Mat49, Moo49, Mud66, Mui72b, Nar50, OR64, OR66, PS67a, Rog59, Rot71, RB58, RSvZ72, SF69, WB41, Wil35]. Independent [Ahu71, AH71, Bel61, BM58a, BG46, BS70, Bro64, Bur68b, Cho66b, CT71, Dan40, Dar64a, Fel51, FK72, Fer70b, Fri69, Gle69, Gre50b, HH63, HK65b, HW71, Har70c, Har69a, Hey69a, HS72b, Hoa71, HL62, HS55, Hoe55a, Hoe56, Hol69, Hsu45b, Jof67b, Kac45b, Kes72, Kin48a, Lin50, Lin52, Mal68a, Mil69, Mil67a, MS70b, Min72, Mon72b, PR54, Pit64, Ros70a, Sam65c, Sam66, Sam69, Sek51, Sha57, Sha58b, Sha60d, SA57, Ste45a, Sto68c, Tom71, Var66, Var68, ZJ59, vZO67, HW58, Mar71a]. Independently [Kul36]. Indeterminate [Dod41b]. Index [Bus52b, Goo56, Goo57a, Nor40, Red51]. Indexed [Ber70a]. Indices [Bak37, Car32b, Gui71, Hoe43b]. Individual [Bre57b, Bre60a, Kre71, Wal44b, Wei54b]. Individuals [Old38]. Induced [Dem67]. Inductive [SM66]. Industrial [Eis39a, Lot39a]. Inefficient [Mos46]. Inequalities [BP66a, BM61, Bru58, Cam48, Dav69a, GW65, Kei72, Kha67a, Kul54, Kur72, Law65, MR69a, MO60a, Mey69, MR67, Nor35, Pan67, Sam53, SS66b, Ser70b, SU72, Tom72, Wic69, Wol71a, vBE65]. Inequality [Bel47, Bir42, BRZ47, Bos49, Box54a, Box54b, CS69, Cra33b, Fab72, Gor72, Gut48b, Jog70, Kad68, Kag71, KsW69, Kul54, Mal68b, Mar60, MO60b, Mik60a, Mik60b, Mor53, Mur61b, OP58a, Ros67, Sam66, Sam69, Sen71b, Sha71b, Ste69b, Ste73, Tat53, Tei55a, VN67, Wal38, Van67, BZ44, Gar59, Tuk46]. Inference [AG57, BP66c, Ber71b, Ber71a, Bir61a, BR67, Bly70, Cha62b, Cox58, Dem63c, Fis39, Fra61, Gra67, Har68, KS63, Kie64, Kni65b, Leh63a, Leh64, Mus68a, Pea62, Rou65, Sch43c, Vil72, Wil38a, Wat66]. Inferences [Bue59, Dem69, Ker63]. Infimum [Pyk59]. Infinite [Agg59, Bla59, Der39, Fel35, FS71, Fer70a, FP69a, Fri69, GG39, Kat67, Kel71, Kli67, Mil67a, Paz70, Pin47, Rue71, Sim71b, Spr53, Ste67, Teu68, TBO71, WK71]. Infinite-Dimensional [Sim71b]. Infinitely [Alb61, Ber64a, BS70, DT57, Fis62, Gan66, Her72, Huf71, KS72a, KW56a, Ore68a, Pol72, Pyk60, Rue70, Sha69b, Smi68b, Ste68, Ste69d, Ste71c, Tuc64, Wol71b]. Infinitesimal [Alb62, Kni70, Ros71b]. Inflation [Sin66]. Influence [BK63]. Information [ABE^{+}55, Bel62, BS68, BD62, BR67, Bly59, BK60, Bre57b, Bre60a, Cam70, Chu61, Chu46b, DeG62, Eis39b, Ell62, Fra65, Ghu68, GW59b, GS60, Har45, Kem69, KL51, Kul52, Kul54, Kul56a, Kul56b, KK66, Kup58a, Kup58b, Lee64, Lin56, MZ66, McM53, Owe70, Rao54, Sch64, Sha64a, Sha71a, SC63, Sri69, SC71a, Sto59, Tho60, Wol58, dlG54, AT63, Wol67b, Haj67]. Information-Statistics [Kup58a, Kup58b]. Informationstheorie [Sch64]. Informed [GS71]. Inheritance [CS39]. Inhomogeneous [Coh70, Coh72]. Initial [Mos58]. Inner [Cle71]. Input [Neu67, Tak61c, Tak62]. Insoluble [Bro71d]. Inspection [CR65a, Cur46, Dod43, Gru49, Lie53a, Moo43, PL46, WW45, Whi65, vdW60]. Inspections [Ney41]. Instantaneous [Bla58b, She58]. Institute [Ano35d, Ano36b, Ano37c, Ano39d, Ano39e, Ano40d, Ano40e, Ano40f, Ano40g, Ano41d, Ano43c, Ano43d, Ano43i, Ano44b, Ano44d, Ano45c, Ano46c, Ano46d, Ano46f, Ano46g, Ano47e, Ano47f, Ano48e, Ano48j, Ano48k, Ano48l, Ano49f, Ano49e, Ano56n, Ano58m, Arn48, Arn50, Arn51b, Arn51c, Bir49, BCD^{+}45, CDDK45, Coc47, Cra37, Cra38b, Cra39a, Cra43c, Cra44a, Cra47b, Doo51, Dwy44a, Dwy45b, Dwy45a, Dwy46c, Dwy46b, Dwy46a, Dwy47a, Dwy47b, Dwy47c, Dwy48c, Dwy48b, Dwy49c, Dwy52, Eis52, Fel48b, Fis50b, Fis51, Fis52b, Gua49, Han43a, Han43b, Hod49, Hod50, Kul46, Lit48, Lit49, Lit52, Mad44, Moo51b, Mos51, Ney50, Old41, Old42a, Old42b, Old42c, Old42d, Old43b, Old43c, Old43a, Old44, Rid40, Rid41, Rob50, She45, Van49, Wal49b]. Institutes [Ano50i]. Institutional [Ano52e, Ano53d, Ano54f, Ano55d, Ano56e, Ano57d]. Insure [WW45]. Integer [Kat67, McC72, Min72, Rob58]. Integer-Valued [Kat67, Min72, Rob58]. Integrability [Boa67, Has63, Isa72b]. Integrable [Bur62b, Isa72b]. Integral [AK68a, Cad54, Dwy37a, Fel65, Fel41, Gor41b, GW65, Lot39b, McF55, McN70, McN62, Rub61b, Rus57, Wei65b, Wen61, Wil46b, Dic68b]. Integrals [Bar65, Bat30, Bro70b, Cla70, Cla71, CD62, Dic67a, Gup63a, Gup63b, Har59a, Isa69b, Par72b, Pra60b, Pra66b, Sho30a, WZ65]. Integrated [Gol71, Lor67]. Integration [Gre49, Pow30]. integrodifferential [Rei58, Rei59]. Intensity [Igl65]. Inter [GW59b, Har45, Sha64a, Sha71a]. Inter-Block [GW59b, Har45, Sha64a, Sha71a]. Interaction [Har67b, Hol72, RK56, Rub63a]. Interactions [Goo64a, Hol72, MS69b, Vaj51]. Interblock [Fra57b, GS60, Spr56, Spr57]. Intercepts [Sen72b]. Interchangeable [CT58]. Interchanging [Pra60b, Pra66b]. Interdependence [Ano31a]. Interest [Den71, Jaq72, MV69]. Interesting [Hen72, Zoc35b]. Interior [Dod38]. Intermediate [RG62]. Internal [Dod37, WG64, van65]. Interpenetrating [Koo67]. Interpolation [Ano35c, Bru69, DM70b, Gre50b, Kin48a, Kin48b, Rob30, Rob35, Sny72]. Interpretation [Eis39a]. Interruption [Kei62]. Intersection [Abe69a, Abe69b, Maj62b, Wal58b]. Intersections [SU72]. Interval [Bak49, Ban67, Dar53b, Dar62, Dem68, Fel68, Geb64a, Gra58, Gre54a, Gut48a, Ney62, Pau69, Por70, RB53, Roy54, Roy56, Sim68, Sta66b, Val68]. Intervals [AJ51, Bat55, CR65b, CMPT60, Dix48, Dwa55a, Ekm59a, Ell64, Far62, Far66a, Far66b, Gee70, Goo64b, Goo76, Hal61, Har70c, Jos66b, Jos67a, Kha69, KF40, Lan71, Leh63b, Man58, MW42, Nád69, Ney35, Noe51, PB71, Pra63, Ros63b, Sco68, SG71, Sin63, SW47, Ste59c, Wal42a, Wal62, Wei55b, Wer39, Wil38c, Wil67]. Intra [BK58b, GW59b, Hal51, KB57, Mor62]. Intra-Block [BK58b, GW59b, KB57]. Intra-Class [Hal51, Mor62]. Intrablock [Fra57b, Spr56, Spr57]. Intraclass [KP67, KP68c, Sri65c, Wal47c]. Intrinsically [Mac67]. Introduced [Der39, Has63, NS60]. Introduction [Bri65, Che66, Goo63c, Spi67, Bro71a, Mil64a, Pru71]. Intuitive [KPS59]. Invariance [BB68b, Ber70c, Ber72d, Dro72, EM70, Fra57c, Fra67b, Fre67d, HWG65, HWG95, Hey69b, Kie57, Kue68, Lam62, Wes59b, Zeh66]. Invariant [Agg55, BG69b, Bat65, Bri63a, Bro66, DF66a, Fel65, FP69a, Fox71, Har64b, HB66, HB67, Imh68, Jai66, Jos70, Jos71d, KS65, KS72b, Kli67, KS69, Lam60, LS53, Leh59, MT65, Met69, Obe71, Obe72a, Ols68, PR71a, Per70, PS71, Sch69f, Spj67, Sta71, Sto70a, Sug72, Tor71, Val68, Wij67, Wij68a, Yan71, Zid69, dO66]. Invariants [Cra31b, Cra40, Dre40, Fre62a, Fre63a, Koe70, Thi31j, Zoc34]. Inventories [Mor66]. Inventory [Sca59]. Inverse [Ban66b, Ban69, Bar51, Gut46b, Kha62a, Leh44, Mus67, Nad60, She36, SM50, SW72a, Twe57a, Twe57b, WK70, WD45, Wer39]. Inverses [Joh64b, RP67]. Inversion [Gur48, Gut70a, LR53, Rio49, Sid58c, Wen61]. Inverted [Dic67b]. invertibility [Tuk58c]. Inverting [Ull44, Wau45]. Investigating [DJ51, DJ52]. Investigations [KP57, Wal43b]. Involved [Hoe40]. Involving [AB66, BC54, Dix51, Hen35b, HP36, Kac45b, LS53, MR69a, Meh64, Sto61a, Tod70, Wes59a, WF57a]. Iosifescu [Nor70]. Irreducible [ASS72]. Irregular [Eav67]. Irregularity [Pol70a]. Isbell [SP65]. Isobaric [Dwy37a]. Isolates [Kat52]. Isomorphism [MS66]. Itô [wK68]. Items [Cur46, Lad39, Vaj45]. Iterated [Bre68, Bre69, Bre70, Dav68d, Fin71b, Fre67c, JP72, Mij72, PS70, Phi69a, Rob70a, Sto70b, Str72a, Tom71]. Iteration [Nel66]. Iterations [Fra68a]. Iterative [Ber45, Bro65, DR72, Fie70, Fie71, Rei49, Ste42, Ull44]. ith [PR54]. Ithaca [Dwy46b]. Ito [Spi66]. Itself [She65]. IV [Che65b, FW51, Jos66a, Kar58b, Rub62a, Tau62, Thi31f]. IX [Thi31g].

J [Bha68, Dav66, ER72, Get66, Gup72, Had63, Har71a, Luk65, Ore68b, Str66b, Wel64a, Wol67a, Zac70b]. J. [Bha68, Blu67a, Dav63a, LeV66a, Mul66, Ore68b, Sch60a, Tak71b, Wen68]. Jackknife [Efr79, GWA72, Mil64b, Mil74]. Jackknifing [AGW71, Arv69, Arv70, Mil68e]. James [Win72]. Jaroslav [Hoe68]. Jensen [CS69]. Jirina [Sau63b, Sau63a]. John [Coc44, Hur48, Ken64, Mal90, Mos65, Tau62]. Joint [Bat55, Imh71, Jen70a, Jen70b, Kem68a, Lor59, Mal72, Meh72, Moo41, Que49b, Rob54, RP65, Sch72, Spr53, Sug70, Wei59]. Joint-Ranking [Meh72]. Jointly [GM62, Wah68]. Jordan [Tak61a]. Joseph [Pyk63]. Jr. [Ano72]. Jumps [Mur65a]. Jurecková [vE72]. Justification [Don52, Dwy44b].

K. [Dud69b, Spi66]. Kac [AB67, God49a]. Kai [LeP70]. Karlin [Had63, Mar71b, Jos69b, MR66]. Kean [Spi66]. Keeping [Tuk56a]. Keilson [Wen68]. Kelley [Moo48]. Kemeny [Ore68b]. Kendall [Bar47b, Kie64, Wat71b]. Kenneth [Had63, Rub66]. Kernel [Par72b, Suk72]. Kernels [Fel65, OS70b]. Kesten [Ath70]. Khinchin [Cra62]. Kiefer [Gup72, Sak66, Der56a, Sak64, Ven67b]. Kincaid [Gre50b]. Kind [Leh44]. Kingman [VJ66]. Kingston [Ano53e]. Klaus [Han66b]. Knapp [Ore68b]. Knockout [Har66b, Har68]. Knot [Mur71a]. Knowing [Sim68]. Known [DS40, Ell65, Fri69, Jam61, Kot64, Mur72, Pra63, SO56, Sev57, Sim67c, Ski68, SW47, Ste42, Wol50, Bec54]. Kolmogoroff [Ste73, Chu48, Ste69b]. Kolmogoroff-Type [Ste73, Ste69b]. Kolmogorov [Bla58a, Fel50, Abr67, And67, Bic69b, Bla56, Bre57a, Con65b, Dar57, Dav58, Don52, Doo49, Egy59, Fab72, Fel48a, Hac63, Kar61, Kie59, Lit72, Mar60, NV68, Noé72, Qua65, Ram66b, Ros62b, Sch58, Sta72a, Ste69c, VN67, Whi61b, Yu71]. Kolmogorov-Type [Fab72]. Koopman [Gup72, Pau70]. Koshal [Mye34]. Krickeberg [Han66b]. Kronecker [Sha59c, Sur68a, Var55, Var60]. Kubilius [LeV66a]. Kuiper [Abr67, Lit72]. Kurtosis [Gut48b, Pic51, Wil44]. Kuznetsov [Owe67b].

L [Ben63, Luk65, Ore68b, Owe67b, Pla71, Wat66]. L. [Has63, Man45a, Owe66, Owe67b, Pfa71]. labels [Roy70]. Lack [Alt39, Far67]. Lag [Sim71b]. Lagrangian [Sil59]. Lai [LeP70]. Lajos [Mil64a]. Laplace [Jam68]. Large [Bah66, Bas67a, Bau63a, Bau63b, Ber62b, Boo72, Bow46, BC64, Bre60b, Che56, Cho65, Cho67a, Dor43, Dwa56, ET68, Fel45a, Gei40, Gor41b, GJ59, Gut70b, HK65b, HK65a, Han67, Hey67, Hin65, Hoa67, Hoe52, KS59, KT61, Kat68, KHS72, KHS74, Koo61, Koo66, Kur72, Lay72, McM49, PS67a, RR64b, Sco71, Seb64a, Set64, Set65, Set70a, Sie69c, Ste57b, Str71b, Str72c, Tay72b, Tom72, Tsu68, Wag69, Wal42c, Wal50b, Wal53, Wei56a, Whi65, Wil38b, Wil38c, Woo70d, Tuk58a]. Large-Sample [Che56, Dwa56, Hin65, Hoe52, Sco71, Str71b, Str72c, Wil38b]. Larger [BC68a, MW47]. Largest [Gum54, Hec60, How49, How50, Kha64b, Mic59, Pau64b, Pil67, RS67, RSS68, Sob67, SPZ56, Sug66, Sug67b, Sug67a, Wal50b, Wal53, Wei59]. Larry [McC65]. Last [BDF64, CS70, vS51]. Latent [And65a, Dav72, Jam60, Jam64, Sug65, Sug66, Sug67b, Sug67a, van61a]. Latin [Fre66d, HS70a, HF71, Joh70b, Man43, Mes67, Hou66, Man42]. Lattice [Bru65, Dyk70a, Har44, Har45, HF70, Iye50, Iye61b, Kue68, Mas71b, Mas72, Maz67, MS70b, Rak67, Rob67, Yar72, Iye51]. Lattices [Rob66]. Law [Ano35b, Bau63a, Bau63b, Bax62, Bel70a, Bel70b, Ber62a, Ber62b, Bre60b, Bre68, Bre69, Bre70, Cho65, Cho67a, Dav68d, DT68, DF65, Fel45a, Fin71b, Fre67c, Gei40, HK65b, HK65a, Han67, Hey67, Hor71, Imh68, JP72, KT61, Kat68, Koo61, Koo66, Kur72, Lah56a, Mij72, PS70, Phi69a, Rob70a, RR64b, She64a, Sto70b, Thi31c, Tom72, Wag69, Wei35, Zid69, HS70c, BP72]. Lawley [Rub68]. Laws [Ano37a, Ano39c, Ano40c, Ano43b, Ano44b, Ano44c, Ano45b, Ano46e, Ano47d, Ano48e, Ano48d, Ano49e, BKS71, BS70, Coc40a, Dwy49a, Her72, Imh68, Imh71, Kul36, Lah59, Mas70a, Mas70b, Pil71, Rue70, Sch70d, Smi68b, Str72a, Tay72b, Thi31h, Thi31j, Tuc68, Wel72, Wic70, Thi31d, Thi31i, Thi31k]. Layer [BJ70]. Layout [Bra65, Imh60, KP69]. Layouts [Sen68b]. Lead [Rio64]. Leadbetter [Ylv68b, Ylv66]. Leading [Kri70]. Learning [BW64b, BM53, HC70, Koc58, MH62, Nor66, Shu69, Nor70]. Least [AT72, And62b, Beu65, Bha62, Ble40, Chu41, Col32, Dav33, DS40, Dow54, Dwy44b, Dwy49b, Eic63a, Gol63b, HB65, Hen31, Hen32b, Hen34, Jen69, Joh40b, Jor32, Jor59, Kim40, Kru68, Leh52a, Len69, MM62, MM63, MR69b, Pie70, Rei61, RW70, Sid58a, Vaj45, Vil66, Vil69b, War60, Wat67, Wis37, Won35, Zac64, Zys67]. Least-Squares [Beu65, Dow54, Len69, Sid58a]. lecture [Hub72, Efr79, Kes72]. Lee [Moo48]. Left [Dal69b, MM69]. Left-Continuous [Dal69b]. Lehmann [SS67, SN70, Fin66, Høy65, SS66a, Set70b, Sho66, SN69, Sun54]. Leland [Fab69a]. Lemma [Bax61, CS52, Chu48, DW51, Der56a, DF65, HK68, Kab65, Kni65b, Pfa67b, Wal44a, Nas54, Pfa71]. Length [Ban67, Dod39, Far62, Far66a, Far66b, GD55, Joh66, SG71, Sin63, Wei55b]. Lengths [Dix48]. Leo [Kal69]. Less [Bak35b, Jof67a, Sam41, Tei56, Tei61c]. Leve [Der66b]. Level [Ber71d, Ber72e, CL65, DLS57, Gut70b, Leh58, LS55, Qua68, Wal59b, Wal51, Bro67, Gul72]. Levels [Aro43, Bak52, Bra52, Coc40b, DB71, Har42, Lea66, Old49, Pei43, Pei49, Wal49f, MT50]. Levine [KW65]. Levy [FW69, Dev59, Fre67e]. Lévy-Raikov [Dev59]. Lewis [Cra44b, Han66a]. Lexis [Hen35b]. Liapounoff [Bel47]. Liar [Fer70a]. Lie [Bri63a]. Lies [Dur68, Eic70, Ste71a, Wol46a, Yar72]. Life [Bas68, CS72, ES54, Eps54b, ES55, Har52, Ley70, MD60, Wal57, WK62]. Life-Distributions [CS72]. Life-Testing [Har52]. Lifetime [Mei72]. lifting [Tul65]. Light [New60, Dar64b]. Like [PW63, Tuk56a]. Likelihood [Abr70, AS58, Bar68, Bas55a, Bir55, BB69, Bro71e, Bru55, Car32a, Cha67, Che54, CL54, Cla57, Con66b, Con67c, CS72, Dal40, DL70a, DL70b, Dic71, DL72, Efr67, Fed68, Fra72a, Fra72b, GW59a, Gir64a, Gir64b, Gir65b, GW70, God60, Hal52, HM67, Hil69b, Hoa71, Hog56, Jog67, KR56, Kat61a, KW56a, KP69, KP70, KL56, LD66, Le 70a, Lor72, MP65, Mye34, Pau41, Pfa67b, Pit65, Sam68a, Sav57b, SP70, Str72d, Sug69, SF69, Tei61b, TB64, Vil61, Wal48a, Wal49a, Weg70a, Weg70b, Wil35, Wil38b, WD39, Wol49a, Woo72, Yar71, Zeh66]. Likelihood-Ratio [Dal40, Pau41, Sav57b]. Limit [Abe69a, Abe69b, Agn57, Agn59, AK68b, AK71a, BN63, Bel70a, Bel70b, Ben62, Ber64b, Bha71b, Bil62, Blo70, Blu66a, Blu67b, BK58a, Bro69, Bro70a, Bro71b, Bue65, CT58, Coc72, Coh72, Csö65a, Csö65b, DK63, Dym66, Eav69, Eic63b, Eic66, Far64c, Far66c, Far66d, Fel50, Fos71, Fri69, Geb64a, Ger71, GZ61, Gre65a, GS67, HH72a, Her69, Hoe51a, Hol71b, Igl65, Igl68, IW71, Isa72a, Jai69, Kei66, Ken71, Ken72, KS66b, KS66c, Kes70, Kim69, Lam65, Lev71, Lin72, Loè50, Mac71, MW43, Mas70b, MW69, MS70b, Min72, tCM67, Ney63, Ney38, Noe70, Par57a, Phi69b, Pie69a, Pie69b, Pil71, PS64, Qua68, Rao61a, Rob72a, RR66, Ros52a, Ros66b, San61, San62, Ser68a, SH72, Sha55, She64a, She64b, Sta65, Sto65a, ST64, Szá72, WW45, Wel72]. Limit [vB65, Fel48a, Sch58, Dar65b, Gol67]. Limited [Cas50, Kim42]. Limiting [Ber66a, Ber66b, Ber62c, Ber64c, Dia63, Doo35, Fis69c, Gle69, Har52, Mad40b, Mad48b, MH62, MZ65, McN70, Mei72, Mit58, MG70, Nan48b, Nor66, She58, Sug72, Tak59, Tho72, Wei58b, Wei58a, Wei59, Whi58b, Whi59]. Limits [BP66b, BP67a, BZ49, Bow46, DD64, FD65, Gar32, HK64a, Hea61a, HS32, Kem56a, Mas50b, Mil72b, Mur48, MS68b, MS71, Noe48, Pau43, Pra60b, Pra66b, Rob44c, Sim71a, Som58, WW39, WW41, Wal42c, Wal43a, Wal46b, WW46, Wel39, Wel64a, Wil41a, Wil42, Wol46a, vM39, Kol41, ST44]. Lindley [Che66]. Line [Cla56, Col32, Eic70, Gav59, Hof64, Moo63a, Por64a, PS69a, Rub58, Too34, Won36]. Line-Graph [Moo63a]. Linear [Ajg67, Alb66, And51a, AW64, AKKW69, And80, And34, AFPR59, BF63a, BP66a, Ber70a, Ber45, Beu65, Bha59, BP65a, Bir50, BM59b, Cac67, Cha72, CGJ67, Cle71, Coh65a, Coh66, Col32, Con66b, Cra43a, DR72, DJ51, Dav68a, DG57a, Dem64, Dic71, Dod30, Dod41a, Dok67b, Dos69, DH69, Dwy41a, Eat70b, Eic63a, Eic66, Elf52, Eri71a, Eze30, Fra67a, GR69, Gei67, Ghu54, Ghu55, Gle65, Gle66b, Gre49, GAM64, Háj68, HK65b, HK65a, Har69a, Har71b, Hil69b, Hoe55b, HB48, Jee54, Jen69, Kac45b, Kat61a, KG70, Koo61, Koo66, Kou69, Kou70a, KvE72, Lah57a, Lah57b, Lan71, Law33, Leh63a, Leh64, Lon42, Lon44, Luk52, Man60, Man69b, MZ66, MG55, Mil68a, Mil71a, Moo68b, MR66, Nag51, NS51, Ney51, NS52, Obe71, Obe72a, Oka63, Oka68, Pla58, Pre66]. Linear [PS69c, Rad58, Rao69, Ray57, RP67, Rei49, Rie31, Rio62a, Rob65a, RO72, Sat42c, SS66b, Sch69b, Sco50, Sea56a, Seb64a, See70a, See70b, SZ71, Sen68d, SP70, Sho69, Smi47, Sny72, SP58, Ste51a, Ste45a, Sti69, Sto68c, SF69, Tan64, Tay72b, Tuc41, Vil61, Vil64a, Vil69b, Vil72, Vot46, Wat67, Wat59, Wei71, Woo70d, Ylv62, Zac63, Zys67, GM57, Rob71a, Gra67]. Linearity [Hil62, Jur69, Ols67a, vE72]. Linearized [Cha67, KvE72]. Linearly [dO66]. Lines [Dur68, Fin46, Sen69, Sen72b, Wal40a]. Link [Rob30]. Linkage [Gei44, Gei45a, Joh40a]. Linked [Gei38, Gei39, Joh70b, Nai66, RL57]. Linnik [Luk65]. Lionel [Le 54]. Lipschitz [Boa67]. List [Sad51]. Lists [Byr35a, Goo52]. Literature [GT64a]. Lives [MS71]. Local [And67, And71a, Ber70e, BF68c, Bro71e, Cho65, GK64a, Lin70, Mas71b, Mas72, MS70b, Min72, Ros66b, She64b, Sto65a]. Locally [Ahm61, AFLP57, Bar49, Cap61, Che53, FG60, Gir68b, Gla59, JM72, Kra71, Sch67d, Sug72, Uza60]. Location [AE69, Bic65c, BH67, Bro66, Bro68, Cha66, Dic71, Far64b, Fer62a, Fer63, FR64, Fox71, GS66, HL63, Høy68a, Hub64, Jae71a, Jae71b, Jos70, Jos71d, KH72, Klo70, Man69b, Moo69, NR71, Nüe66, Pfa69a, PB71, Pol70a, Rao68a, Ros54, SY72, SG56, SG58, SG69, SP67, SP68, Ste59b, Sto61a, Suk58, Tak69b, Tat59, Val68, Wal58b, Wei57d, Wel39, Win72, Van70]. Locations [Dix48, Rag65b, See72a, Ton69]. Locks [Wel64a]. Log [DR72, Fed68, Kei72, Zac66b, Zac67]. Log-Concavity [Kei72]. Log-Linear [DR72]. Log-Normal [Zac66b, Zac67]. Logarithm [Bre68, Bre69, Bre70, Dav68d, Fin71b, Fre67c, JP72, Mij72, PS70, Phi69a, Rob70a, Sto70b, Str72a, Tom71]. Logarithmic [Dod41b, Jen32, Que36, WB64, Yua33]. Logarithmico [SO56, Sev57]. Logarithmico-Normal [SO56, Sev57]. Logistic [Bai31, BD63, GS65, Sha66a, Sha70b, TC65, TC80]. Lois [Luk65]. Long [Bro71a, Pow30]. Look [Efr79, GR70, Ros71b]. Look-Ahead [Ros71b]. Loose [Huf69]. Loss [Bha66, BY69, DR63, Far68a, JN63, JM63, Kar58a, San68, Sta71, SC71b, Stu71a, Van66d, Van66c, Zid69, Zid70, vdW60]. Losses [Sam64]. Lots [GJ59]. Louis [Ano36b]. Low [Ber71d, God49b, HMTW47]. Lower [Ber71b, Ber71a, BG47, Dav69c, Dem67, Dem68, Dem69, GW68, Gov62, Ham57, Hoe53a, Hoe60, Jon48, Kle71, Mar48, Mor52, Sim67b, Ylv64]. Lusin [DZ67]. Lyapunov [Bla51a].

M [Bur68a, Dym70, Gre50b, Gup72, Luk65, Mul66, Pak71, Wel64a, Wol67a, Ylv68b, Bro70c, Bur68a, Hea61b]. M. [AB67, Dal69a, Dup70, Fis60, Kem64, Nor70, Wat71b, Wel64a]. M/G/ [Bro70c]. M/G/1 [Hea61b]. M/M/s [Bur68a]. Machine [Too34]. MacNeish [Bus52a]. Macroscopic [Bro70d, Bro71f]. Madison [Arn48]. Magnitude [KOvE72]. Main [AK61, Bra65, Imh62, PF70, RF70a, RF71]. Main-Effect [AK61, Bra65]. Main-Effects [Imh62]. Maintenance [Fle62a]. Maitra [Get66]. Majone [Get66]. Making [Kis45, Shu69, Sim68]. Man [Alb70]. Management [Had63]. Manifolds [Jee54]. Mann [Bas67a, Zar65, BK57a, Cha63c, HG66, HF70, Het68, Sho66]. Manova [Bha60, Gup71, Gho64a]. Many [Alb61, KW56a, Moo39, tCM67, Por66, Pyk61b, SY68, Wei72b]. Mapping [Dem67, Kat55, Rio49]. Mappings [Har60a, Rio62a, Top67]. Marcinkiewicz [Cho69b, Gun67]. Marginal [Dav72, DS40, Fle62b, Isa71, Kul71, Que36, Ste42]. Marginally [Tai66]. Marginals [Kul68a, Kul68b, Str65a]. Markoff [Eps51, Eps49a, Fel65, Mon47, Wal38]. Markoffschen [Pyk63]. Markov [Bha68, DL70a, DL72, Ore68b, Pyk63, Abr69, ASS72, Alb62, AG57, AK68b, Ath68, BP66c, BPSW70, Ber63a, Ber63b, Beu63, Bha61d, Bha61e, Bha64, Bil61c, Bir62, Bla55b, BK57b, Bla58b, BF64c, BF68c, BKR71, Bou68, Boz71, Bre57c, Bre58b, Bre60b, Bro65, BR58, Car67, Cla56, Coh70, Coh72, Cra70, DS53, DG57a, DBTL68, Den71, Der61, Dha63a, Dha63b, Dha64, Dha65, DN70, DL70b, DF66a, DS69b, Dym66, Dym70, Eat70b, Eri70a, Eri72, Fos52, Fre62b, Fre67d, Gav59, Get66, Gil59b, Gol63a, Goo61c, Goo58a, Goo58b, Goo59, Hac63, Hel65, Hor71, Ios72, Isa64a, Isa64b, Isa67, Isa72a, Jai69, Jaq72, JR69, JR70, Kak71, KT60b, KT61, Kei69, KS58, KS61, Ken53, Kim69, Kni70, Kni72, Koo60, Koo66, Kra59, Kru68, Kun67, Kur69, Kur71]. Markov [Lam60, Ley67, Ley70, Lin72, Lip71, MR65, Mac67, McC69, Mil61a, Neu64, O’B72, OS70b, PS70, Pet69, Por64b, Por66, Pyk61a, Pyk61b, PS64, PS66, RR64b, RR66, Ros71c, Rou65, Rou69, RM68, Sam69, Sch71a, Sch70a, Ser71, Shi71, Spi72, Sto68b, Sto69b, Tay68, Tay70, Tsa68, Wat61, Won69, Yac68, Yan71]. Markovian [Der66b, BR58, Der66a, DV67, Dha68, Fel59, Fle62b, FR68b, FR69, FR70, Fur72, Hol72, Ros68a, Ros68b, SY70b, Swi65, Tay65, Whi65]. Markow [Bil61a]. Martin [Bha68, Sal67, Abr70, Pru71]. Martingale [Bro71b, Bur66, Cho68, Cho69b, Dav69a, Dav69b, DF66b, Gor72, Gun67, Sim70, Sto68c, Wal70a]. Martingales [Aus66, Boy71, Cha64, Cho65, Cho67a, Dav68b, DFJ68, Dol69, Dro72, Gil72, HB70, Isa71, Isa72b, Lam71a, Mar69, Mil69, Mon72a, Nel70, PS72, Ste69b, Ste73, Sto70b, Sud69a, Gun68, Kri71a]. Marvin [Mad63]. Mass [Ney41]. Massey [CLA62, Tsa54a]. Match [BF68b]. Matching [And43, Bat42, Gre38, Gre41, Gre44b, HL54, KR45, Kul39a, Lev49b]. Matchings [Cha35]. Matchmaking [DFG71]. Mathematical [Alb63, Ano35a, Ano35d, Ano36b, Ano37c, Ano39b, Ano39d, Ano39e, Ano40d, Ano40e, Ano40g, Ano41c, Ano41d, Ano43c, Ano43d, Ano44d, Ano45c, Ano46f, Ano46g, Ano47e, Ano47f, Ano48e, Ano49f, Ano49e, Ano56n, Ber42, Bol34, Bur32, Cam38a, Cam42, Cra37, Cra38b, Cra42, Dav63a, Dep30, Fel35, Hsu45a, Kin30, Lar35, Pru71, Ros66a, Sho30a, Sho30b, Sho30c, Wij68b, Cra62, Mal90, Blu67a, Cox62, Fel47, Hoe62, Lin66, McC65, Owe66, Thi31s]. Mathematically [Tuc41]. Mathematicians [Wol58]. Mathematics [Cra39b, Hed70, Kem64, Zac70b]. matric [Hot44a]. Matrices [AB62, And63b, AG64c, AG65, Ban64, Dem64, Fed51, Fos53, FK60, Gir68a, Gir68c, Gna59, Gna60, GM69, Gup69, pH69, HA72, Jen70a, Jen70b, Kab63, Kab64, KP65, KP68a, Kra59, KP67, KP68c, Kri68, Lay72, LPC70, Loy66, Lut65, Mik62, Moo51a, Pil64, PN71, Pru64, RG62, Sin64, Sit52, Smi64, SN68, Sur68a, Tho58, Var55, Ylv64, Zys67, GM57]. Matricvariate [Dic67b]. Matrix [And65a, Bac48, Bar51, Bin72, Bow47, Cha51, Con69, Dav72, Dem63c, Dwy44b, DM48, Dyk70b, ESM72, Fra63, Goo63b, Gut46b, Hot43a, Hot43b, Jam60, Jam64, Kei69, Kha64a, Kha65d, KP66a, LaM72, LR53, MG55, Mij61, Mij64, Mui70b, Mus67, Nag51, OK68, PM59, Pil67, PG68, PAAJ69, Sat44, Sea56b, Sha60b, Sha70a, SM50, Sid58c, Sri68b, Sri68c, SC71a, Sug69, Sug67b, Sug70, Ull44, Van66d, Van66c, Vil69a, WD45, Wau45, Wij59a, dW72, van61a, LW44].

Maurice [Bar47b, Kie64]. Max [Usi64]. Max-Min [Usi64]. Maxima [Gum54, Mar72a, Pic67, dHH72]. Maximal [AFLP57, Bax65, Bel61, Bel62, Che67a, FG60, Gla59, Koe70, Lip71]. Maximising [Enn73]. Maximization [BPSW70, Gil64b, Mur72]. Maximize [Goo68]. Maximizing [BJK56, CMPT60, Enn70, Rus57]. Maximum [AS58, BGP69, Bas55a, Ber62c, Ber62b, Ber64c, Ber64b, Ber72e, Bru55, Car32a, Cha67, CL54, Cla57, CS72, Gal72, GW70, God60, Goo63a, GG50, Gur55b, Gur59, Hal52, HM67, Hil69b, Hoa71, Hot41, Jac69, Jog67, KW52, KW56a, Kim42, Kim46, Kim49, KP69, Kni65a, KL56, KA65, Kus66, Le 70a, Lin70, Mar70, MP65, Mas51a, Mye34, Pic51, Ray65, Sam68a, Sei55a, Sei55b, Sei56, Ser70b, Sie69a, SA57, Ste71b, Sto68b, Tei61b, TB64, Tsa54a, Vil61, Wal48a, Wal49a, Wal58b, Weg70a, Weg70b, WK69, Wol49a, Woo67, WV67, Woo72, Zeh66]. Maximum-Likelihood [AS58, HM67]. Maxwell [Rub68]. Mc [Spi66]. Mean [Agg66, Alb56, BR50, BR51, BR60, Bak35b, Bar70, Bec42, Bha66, BS72, Bir54, BC64, Cam38b, Cha72, CZ64, Che65a, Che65b, CR65b, Coh50, Coh65a, Coh66, CS70, Cra32, Cra40, Dav62, Dav63b, Dix57, Dod40, Dod41b, Dud69c, Dwy40a, DL66, Ekm59a, Fré47b, Gän70, Gar69, Geb66, Gei56a, GW70, God49a, Gum54, Gur55b, Gur59, Gut48a, Har42, HvN42, HD54, Her48, Hoe46, Hsu45b, Jof67a, Jos67a, Jos67b, Jos68a, Jos69a, Kir35, KOvE72, LLN60, Lan71, Lar39, LC65, LR72c, Lom52, McC72, Mei72, Mos52, Mur72, Nád69, Nor37, Pau50, Pau42b, Pau63, Pau64b, Pol34, Pra58b, Pra63, PR68, Rao68b, Rei63b, Ric60, RC60a, Roy33, Sar54, Sar55a, Sar55b, SMR72, Sco67, Sco68, Set65, SO56, Sev57, Sim43, Sim67c, Sim68, SW49, SA57, Spr53, Sri69, Sta66a, Sta66b]. Mean [SW47, SC71b, Str71a, Thi31s, Twe67, Wei52, Wei55b, Wel47b, Wer37, Wil41b, Wol50, Zac66b, Zac67, Zac70a, Zei50, Zoc35a, CT56, vN41, vNKBH41, vN42]. Mean-Square-Error [Twe67]. Means [And63b, AF62, Bak31, Bak32, Bak40, BM59a, Bec54, Beh61, BBFH69, Car32a, Cha50a, Cra31a, Dan71, Dod37, Dod38, Dod41b, Dun31, Dun58, Fab62, Fin46, Fra52a, Gof53, Gre32, Jam61, Joh40b, Kin52, Kle71, Loh69, McC47, Mus68b, Nor40, OS70a, PL70, PL71, Pau69, Ram56b, Rob48c, RSS68, Rod71, RP58, Sea55, Set64, Set70a, Sta38, TW34, Wal46c, Wil46a, WF57a, BT46]. Measurability [Dar71, Han59]. Measurable [DZ67, Led71, Rob67, Sud69b, Sud71c]. Measure [Ail70, Blo50, BN45, Che52, Doo41, Gar32, Gle64, HS32, Kai67, KZ71, KM67, Lam60, Lin56, Lin60, Luc67, Maj62a, Met69, Rob44b, Rob45, Rob47, San47, Sch43a, Sil64, Sol53, Sto59, Sto65c, Var58, Vot46, Weg56]. Measure-Spaces [Var58]. Measurement [Coc43, Lor59]. Measurements [Kis45]. Measures [Bar61, BG69b, Bel62, Ber51b, Cap64, Dud68, Dud69b, Dun70, Eri72, Fel65, FV69, Fis69c, Fre66c, HH72c, HY64, Isa64a, Isa64b, Jai66, Jiř72, Kai67, KZ71, Kai71, Kan69, Kli67, Le 70b, Led71, Nor38, PR71a, Pit57, Pit68, PS66, RC72, Rao62, RR72a, RR72b, Rod66b, Str65a, Tei67, Tsu68, Var67, Wei34, Whi70, Yan71, da 45, She66]. Mechanical [McP41]. Mechanics [Kos45].

Median [Bha61b, Chu55, CH55, CRWC72, Hil62, Pfa70, Pol34, Sar60, Sid62a, TC65, Tho36, Wal49g, Zei50]. Medians [Dun59, HS61, Moo41, RR65, Tuk46]. Meeting [Ano36b, Ano40f, Ano43i, Ano48j, Ano48k, Ano48l, Ano50i, Ano53e, Ano54g, Ano56n, Ano58m, Arn48, Arn50, Arn51a, Arn51b, Arn51c, Bir49, Cra37, Cra38b, Cra39a, Cra47b, Dwy44a, Dwy45a, Dwy46c, Dwy46b, Dwy46a, Dwy47a, Dwy47b, Dwy48c, Eis52, Fis50a, Fis52a, Gua49, Han43a, Han43b, Hod49, Hod50, Kul46, Lit48, Lit49, Lit52, Mad44, Moo51b, Mos51, Old41, Old42b, Old42c, Old42d, Old43c, Old43a, Rid40, Rid41, Rob50, Van49]. Mellin [Eps48, Kot65]. Members [Ano46g, Ano47f, Ano52e, Ano53d, Ano54f, Ano55d, Ano56e, Ano57d, CW58, CW59, Dor43, Eps49a, Eps51, Ste57b]. Membership [Ano54g, Arn51a, Fis50a, Fis52a, CDDK45]. Memoriam [Cra44b]. Memory [HC70, HC71, Sam68b, SP65]. Memoryless [Bla59, Bla64, DS66, KW72, Kot61]. Mendelian [CS39, Gei44, Gei45a]. Merging [BD62]. Messages [Hai61]. Method [And34, AJ56, Bas55a, Ber37, Ber45, Bol34, Bro65, Car32a, Chu54, Col32, Cra39b, DJ51, DJ52, Dav33, Die42, Dod38, Dre41, Dup65, DT64, EG59, Eze30, Fer58, Fer36, Fra57c, FP69b, GW68, Gil64a, God61b, Goo61d, Goo53b, GS60, Gri31, Gue51, HF70, Hen31, Hen32b, Hen34, Her67c, Hoe71, Hor32, Hor35, Jen32, JF64, Kal54, KR45, Kem63, Ken53, Kha69, Kni65c, Lar36, Mat43, Mos58, Mye34, O’T33a, O’T33c, Old40, Pow30, Que36, Rei49, Rei40, Rid61, RP49, RM51, Roy53a, Roy33, Sam42, Sei61, Sen66, Sha60b, Sin40, Ste42, Tra68, Tsu67, Vaj45, Wal43a, Wau45, Wil30, Wol52b, Wol54a, Wol57, Wri34, Hor48, HS49b, Mul66, Rub68]. Methods [Bil61a, Bil61c, Blu54a, Blu54b, Cam37a, Cha61, Dal62, Dem60, Dem63b, Dem66, Dwy42, Dwy44b, Efr79, Eis39a, Fab62, Fab68b, Fel47, Fre57, GZ66, Goo57c, Goo61b, Gut46b, HL62, Hoe41, Hoe43a, Hot43b, Jog68a, Joh53, KKW55, LeV66a, Mos62, Mul56, NS51, NS52, Pha63, Rob70a, Rob30, Sea56b, SM66, Thi31o, Wen68, WG64, Wol56, Mul66, Beu65, Dal69a, Mos48b]. Metric [BE72, Dud69b, Eis72, Led71, Nev72, Ros63b, VR58]. Metrics [FW69, Tho57]. Meyer [Kni66, BB55]. Midrange [GCM65, Pil50, Wal49f]. Midranges [Gum44]. Mijares [Gna67]. Mill [BC54, Bir42, Raj53, Sam53]. Miller [Pra66a, Rub66]. Mills [Boy58, Gor41b, PW63, RP63]. Mills’-Ratio-Like [PW63]. Min [Usi64]. Minima [Fab67, Gre65a, Hög72, McM49]. Minimal [Bur60, Ehr66, LR72b, Vat46]. Minimax [Agg55, Agg59, Agg66, All53, Ant62, Bar70, BY68, Bly51, DeG60, Dok66, Dok69a, DT71, DKW56, FK51, Gho64b, GKS63, GK64a, GK64b, Gir68b, HL50, Hoe65b, Hol71a, Jos68b, Kat61b, Kie52b, Kie57, KW59a, Lev66b, Mot58, Nel66, Rad58, RH71, Rob58, Sch69a, Sch67d, Ski68, Ste70, Str71a, Teh63, Try58, Vog60a, Wes59b, Wol50]. Minimization [OK68]. Minimize [Goo68]. Minimizing [Jae72, Rus57, Sin40]. Minimum [AD70, BGP69, Cam70, CR51, Dav51b, Gol63b, GG50, Jof67b, Kie52a, KK66, Lan72a, MM62, MM63, Pat63, PJ70, Pla48, Pol70a, Pol70b, SV69a, Sea56a, SZ71, Sri67, Ste51a, Ste50, TG58, Wol54a, Wol57, Ylv64, vdW60, Tuk58b]. Minneapolis [Arn51c]. Minors [Rei40]. Minutes [Ano53e, Ano54g, Arn51a, Fis50a, Fis52a]. Mises [Fis60, Lin66, And62c, BD72, BKR71, Bur63, Bur64b, Cra53, Dar57, Kie59, MS72b, Ros52a, RW72]. Misleading [Ker63]. Missed [Jon62, Sch65]. Mixed [Ban70, Coh58, DJ52, HA72, Imh60, Imh62, Jos71a, OT61, OT65, Rak69, RC60b, RC60c, Sch56b, See70b, Tho55a, Wei58c, WH65]. Mixing [AB61, BM59a, Boe66, DK68a, Fre62a, Fre63a, HK65a, KS72a, Loy65, Mee72a, Phi69a, Phi69b, Rol68, Sen71c, Ser68b, Wel71, Wel72]. Mixture [Bar69, Bli62, Fis72b, MS72a, Min68, Nig70, Nig76, Rid61, Rob48d, Tei60, Tei57]. Mixtures [Fre62b, Kel71, MvZ66, MD68, RP49, Ste67, Ste68, Tei61a, Tei63, Tei67, YS68, Yak69]. MLE [Ber73, Ber72c]. Mode [Fré47b, Gre65b, Par62, Ven67c, Weg71]. Model [AE69, Ber66a, Ber66b, BM53, Dav66, DS69b, Fle62a, Fle62b, Fra67a, Fra72b, Gau59, GR52, GH61, Hil69b, HA72, Igl68, Imh60, Imh62, Kak71, KB71, KP67, KP68c, LaM72, LB63, Lau63, Oga63, OP69, Por71, RT69, Rob72b, RC60b, RC60c, Sch56b, See70b, Shu69, Sin70, Sin71, Sri65c, Sri68d, Tho55a, Tho55b, Con65a, Far69b, Her59, Leh61, Sho67]. Models [Ber70d, Ber72c, Ber73, Bha59, BP65a, Bly70, BBH56, DR72, DJ52, Dok67a, Dok67b, Dok68, Eat70b, Eri72, Fra67b, FP71, Haq70, Kno67, Koc58, Leh63a, Leh64, MZ66, McF55, MH62, Mil71a, Nor66, OT61, OT65, Pyk58, Ros68b, Ros71c, Sch56a, Sen68d, SP70, Spj67, Whi65, Yan68, Zys67]. Moderate [Dav62, Dav68c]. Modes [Gup61, Wol71a]. Modification [And60, Ano33, Mor53, Rob30, Tam63, vM38]. Modifications [Blo57]. Modified [Ams63, Bru62, Dwa57a, Eps49b, Gei56a, Lan72a, Ros62b, Wes59b]. modulo [Jag63]. Moment [Alt39, Bli62, Cra40, Cur42, Dwy37a, Dwy40a, FP69a, Han67, JR51, Kad71, Kad73, Kei72, Kem68b, Kes72, Kha54, Koz47, Koz51, LLN60, Mil67a, Moo51a, Mor52, Nag51, Pic68, Pie40, Pil68, Rio37, Ser70a, Ser70b, Sto65b, Teu68, TBO71, Vat46, Wei56b, da 45, vBE65, Tuk56a]. Moment-Function [Pie40]. moment-like [Tuk56a]. Moments [Ano33, Bak35a, Bat31, BK63, Beh61, Bel47, BH65, Bro69, BE70, Bro70a, Bro71c, CRT65, Cho66a, CT66, CH55, Chu46b, Col51, CL65, DJ68, DFJ68, DJ69, Dod38, Dwy37a, Dwy38b, Eri71b, Eri66, Eva40, Fel35, Gla59, God49b, Gov62, Gov63, GS65, HPW64b, Hoo56a, Iye58, Iye61a, Jon48, Jos71a, Kat60, Kem70, Ken40, Ken41, KP65, KP66a, KP68a, Kir35, Lam60, Lar36, Lar39, Lie55, McC64, Mij61, O’T33c, OS62, PM59, Pil64, Rid61, Ros70a, Roy53b, RP65, San47, Sch61a, Sha66a, Sha70b, Sid62a, Sie69a, Ski68, Ski69, Str72b, Tuk49a, Wei66, Whi57a, Wil32a, Wil41b, Wol71b, uD54, vB65, HMTW47, Tuk57e, ZD59]. Monica [Moo51b]. Monographs [Ano56o, Hed70]. Monotone [BMP63, BM64a, BM64b, BP69, BD69, Bru55, CS70, CRWC72, Dem69, GW59a, Jog67, JS68, KR56, MR65, MP65, Mud65, Pfa67b, Rao70a, RW68, Sac70, Ste69e, Wei66, Bic69c]. Monotonic [RM61]. Monotonicity [Ala70b, Ala70a, AG64b, AG64c, AG65, CS69, GAM64, Hoe70, Kru54, Wij60, Wij63, Wij75, CS68]. Monro [Blo57, DK72, Kal54, Ven67a, Wol52b]. Monte [Mul66, Mos62, Mul56, Mul66, Wen57]. Mood [CLA62, Cro64]. Morgenstern [Hur48]. Most [Cap61, Fra57a, Gas65, Háj62, Hal58, Hal59, JR69, JM72, Kra71, LS48, SS66b, Sch68a, Sen67, Spj68a, Uza60, Wal41a, Wal41c, Wei52, Ahm61, Ter52]. Most-Economical [Hal59]. Mosteller [Wat66]. Motion [Ber67b, Bha71a, Cla70, Cla71, Eis69, GHRM71, Hel70, Huf69, Kin63b, Mon72a, Rod71, Roo69, She58, Yos49, Doo68]. Motions [Kni65a]. Moving [Bru71, Hoe46, KMP59, MR57, Pie70, She64a]. Mr. [Zoc37]. Mullin [Byl70]. Multi [BH60b, BH57b, Cau68, DLS57, HP63b, Igl68, Kar57b, LS55, MP67, Mil68a, Par70, RK56, Sri69, Sto65a, Wic69]. Multi-Action [Kar57b]. Multi-Auxiliary [Sri69]. Multi-Dimensional [Cau68, Sto65a, Wic69, HP63b, Mil68a]. Multi-Factor [BH57b]. Multi-Level [DLS57, LS55]. Multi-Parameter [Par70]. Multi-Sample [BH60b, MP67]. Multi-urn [Igl68]. Multi-way [RK56]. Multicollinearity [Tin45]. Multidimensional [Blu54b, Dic68b, Eat70a, Goo63a, Goo64a, KS66b, KS66c, Kul71, Luk39, Neu71, SA70, SA71, Rub66]. Multidimensional-integral [Dic68b]. Multihypothesis [Sim67b]. Multilinear [Par57a]. Multinomial [Ala70a, AT72, BEM59, BW67b, Cas50, Dic68a, DKW56, Goo57c, Goo61b, Goo61d, Goo64b, Goo76, GG50, Hoe65a, KM59, Koz56, Kul37, San72]. Multinomials [Wes59a]. Multinormal [Bir50, BC50, Coh57]. Multiparameter [BW71, BT65, Leh52b]. Multiparametric [Bir55]. Multiple [And71b, And72a, And62a, Bac38, Bak35b, Bar43, Bat42, Bec54, BS54, BEM59, Bha59, Che62, DBTL68, Dun61, Dwy37b, Fab62, Fis33b, Fox71, Gab69, Gre44b, Gut38, Hal58, Hol66, Hor32, JN63, KR45, Kha64a, Kir37, Kni65b, Kou69, Kou70a, Leh57a, Leh57b, Mei69, Min31, Mur71a, Par72b, Pau49, PP69, RB58, Sch69a, Sch70c, Sil48, Spj72a, Spj72b, Sri68b, Sri68c, Sta39, Ste70, Tin45, Tra68, Tra69, Wal41b, Whe31, Wil32b]. Multiple-Decision [BEM59, Hal58]. Multiplex [Eva63]. Multiplication [Car70]. Multiplicative [Goo68, Ott49]. Multiplicity [ET72]. Multiplier [Sil59]. Multiresponse [Sen71a, Sri68a, WG64]. Multisample [Bel64a, Des70a]. Multistage [BR63, GKR57, HSF72, Sch67b]. Multitype [Ath68, Ath70, Kur70, tCM67]. Multivalued [Dem67]. Multivariate [AE69, And46, And51a, AB62, And64a, And71c, And80, AF62, Bac63, Bar70, Beh61, Beu63, Bha59, Bha66, Bha67b, BJ68a, BP68, BM61, Cha62b, Coh65a, Con63b, CD62, Dav67, Dav70, Dav72, Dem63b, Dic67b, Eat70b, Eic66, Eli70, Ell65, ET72, Fin46, Fis70, FD65, GG58, Gei65, Ges70, GO62, GO69, Gir64a, GK64a, Gir64b, Gir65b, Gir68b, Gna67, Goo56, Goo57a, Goo61d, Goo63c, GM72, Gup63a, Gup63b, GAM64, Gup65, Gup69, Gup70, Har70b, Her67a, HS61, Igl64, Jam54b, Jog68b, Jog70, Jos67b, Kab62, Kab65, Kap71, Kat61a, Kha64b, Kha65c, KP65, KP68b, KS65, KS72b, KP69, Koo64b, KOvE72, KP51, KP60b, Ksh61, Kul35c, Kul52, Kul56a, Kul56b, LQ65, Mad59, Mar62, Mar63, Mar64, MO60a, MT67a, MT67b, McF55, Mey69, Mij61, Mij64].

Multivariate [Moo41, Moo69, Mud65, Mud66, MR67, Mui70a, Nar50, Nüe66, Obe71, Obe72a, OR54, OP58a, OT61, OR64, OT65, OR66, OS62, PB70, Per69a, Per71, Pil55, Pil56, Pil67, PJ71, Ros52b, Ros62a, Roy53a, Roy58, RR59, RC60c, RM61, Sch71a, Sch67c, SMR72, Sco67, Sco68, See72b, SP67, SP68, SP70, Ses66a, SP69, Šid68, Šid71, Sri65a, Sri67, Sri68b, Sri68c, SC71b, Str71a, Stu71b, SF69, Sug67a, Tam66, Tam69, Tan70, TK70, TB64, Tro71, Vot48, Wat59, Wei58d, Wei60, Wij57, Wij59c, Wij67, Wil46a, Win72, dW70, dW71, Dun65, Mil68d, Tuk48b]. Multivariate-Normal [Ell65]. Murray [Gne64]. Murthy [Dal69a]. Must [Sta72b]. Mutual [Bel62, Fre66b, Kha59, Kha61, Won36]. Mutually [Chu42, HSF72]. Mutuel [EG59]. my [Erd50].

N [Dal69a, Ore64, Owe66, Owe67b, Rio62b, Rub68]. N. [Dwa67a, MW69, Mor66, Sal67]. Natural [Str66b]. Nature [Alb61, Ano36a, Gau59, Her59]. Near [Fed68, Lin70, Pre69a, Sid58b]. Near-Cyclic [Pre69a]. Nearest [Enn70, Enn73]. Nearly [MS72a, Moo69]. Necessarily [Blo71]. Necessary [Bri63a, DeH70, Far68a, Koz51, LS60, Oga59, Rue71, Sei63, Smi68a, Ste70, Ste55, Sto70a]. Necessity [Pit57]. Neck [Alb70]. Need [Min68]. Needle [Man53]. Negative [Ahu71, Har61, Har64a, Mat49, Mur56, Pru64, SA66, Sid63, Smi68b, Ste45b, Str66a, Tho62, Zys67]. Nelson [Doo68]. Nested [LaM72, Mah63, Sea61]. Neumann [Hur48, Tau62]. Neveu [Blu67a]. News [Ano43e, Ano43f, Ano43g, Ano43h, Ano44e, Ano44f, Ano44g, Ano45d, Ano45e, Ano45f, Ano46h, Ano46i, Ano46j, Ano46k, Ano47g, Ano47h, Ano47i, Ano47j, Ano48f, Ano48g, Ano48h, Ano48i, Ano49g, Ano49h, Ano49i, Ano49j, Ano49k, Ano50e, Ano50f, Ano50g, Ano50h, Ano51e, Ano51f, Ano51g, Ano51h, Ano52f, Ano52g, Ano52h, Ano52i, Ano53f, Ano53g, Ano53h, Ano53i, Ano54h, Ano54i, Ano54j, Ano54k, Ano55e, Ano55f, Ano55g, Ano55h, Ano56f, Ano56g, Ano56h, Ano56i, Ano57e, Ano57f, Ano57g, Ano57h, Ano58e, Ano58f, Ano58g, Ano58h, Ano59e, Ano59f, Ano59g, Ano59h, Ano60e, Ano60f, Ano60g, Ano60h, Ano61f, Ano61g, Ano61h, Ano61i, Ano62i, Ano62j, Ano62k, Ano62l, Ano63e, Ano63f, Ano63g, Ano63h, Ano64e, Ano64f, Ano64g, Ano64h, Ano65g]. News [Ano65h, Ano65i, Ano65j, Ano65k, Ano65l, Ano66g, Ano66h, Ano66i, Ano66j, Ano66k, Ano66l, Ano67g, Ano67h, Ano67i, Ano67j, Ano67k, Ano67l, Ano68g, Ano68h, Ano68i, Ano68j, Ano68k, Ano68l, Ano69g, Ano69h, Ano69i, Ano69j, Ano69k, Ano69l, Ano69m, Ano70g, Ano70h, Ano70i, Ano70j, Ano70k, Ano70l, Ano71g, Ano71h, Ano71i, Ano71j, Ano71k, Ano71l]. Newtonian [Ore64]. Neyman [CS52, Cou44, DW51, HK68, Kul39b, Oga63, Sch60a]. Nicholson [Ano72]. Nikodym [Cap64, HS49a, KZ71, Kai71, Var67, She66]. No [HP36, RK56, Vei66, DM68a, Imh62, Tay70]. Noether [Háj61]. Noise [Eri71a, Hin65, KS68]. Noiseless [Bil61b]. Non [Adi67a, And46, And64a, And34, Arn65, Bak32, Bak34, Ban66a, Bha65, Bic65a, BRN63, Bra65, Bre68, Bre69, Bre70, Bro71e, Car50, CR63, Con63b, Cra41b, Dan40, DBTL68, Der57, DN70, Dwa55b, Ela62, FT66, Fel59, Fre66d, Gil64b, Gil69, Gut59, Har71b, Hil60, Hil61, Hoe48b, HPW64a, HPW64b, Isa64a, Jen69, Joh57b, KP65, KP66a, KP68b, Kim47, KG70, Koo68, Ksh61, Lad39, Law65, LC65, LS49, Mad40a, Mas50a, Mat49, McN62, MS70b, Mud66, Mur48, Par64, PG68, Pol70b, Pre66, Pru64, Rei63b, Res62, Rid31, Rie39, Rob65a, Rob71a, RR64b, RR66, Ros68b, RC60b, RC60c, Rub62a, Sch69b, Sch43a, Sch43c, ST45, SMR72, Sen66, SK61, SK63, Sha63, Sha68, Sio71, Smi68b, Som58, Spj68a, Sri68b, Sri68c, Sto61a, SF69, Tuk47a, Var59]. Non [Vil69b, Wal48c, Wel64a, WAC62, Wen61, Yen64, Zac67, Zys67, AT63, WW43, KJB67b]. Non-Absolute [Wen61]. Non-Central [And46, And64a, Car50, Con63b, Cra41b, HPW64a, HPW64b, KP65, KP66a, KP68b, Ksh61, Law65, Mad40a, McN62, Pre66, Res62, SK61, Sha63, Sri68b, Sri68c, Wal48c, WAC62, Gil64b, Har71b, Par64, PG68, SK63, Sha68, Wel64a, KJB67b]. Non-Decreasing [Bre68, Bre69, Bre70]. Non-Discounted [Ros68b]. Non-Equivalent [Sto61a]. Non-Existence [BRN63, Dan40, Var59, Zac67]. Non-Exponential [Spj68a]. Non-finite [Gil69]. Non-Homogeneous [Bak32, RR64b, Rub62a, RR66]. Non-Identically [CR63, Koo68]. Non-Independence [Mud66]. Non-Lattice [MS70b]. Non-Linear [And34, Jen69, KG70, Rob65a, Vil69b, Rob71a]. Non-Local [Bro71e]. Non-Markovian [Fel59]. Non-Negative [Mat49, Pru64, Smi68b, Zys67]. Non-Normal [Arn65, Bak34, Gut59, Lad39, Rei63b, Rid31, Rie39, RC60b, RC60c]. Non-Normality [Bha65]. Non-Null [Sio71, SF69]. non-numerical [AT63]. Non-Optimality [SMR72]. Non-Orthogonal [Bra65, Fre66d]. Non-Parametric [Adi67a, Bic65a, Der57, Dwa55b, Ela62, Hil60, Hil61, Hoe48b, Joh57b, Kim47, LS49, Mas50a, Mur48, Sch43a, Sch43c, Yen64, ST45, Sen66, Som58, Tuk47a, WW43]. Non-randomized [Ban66a]. Non-Regular [Pol70b, DN70]. Non-Singular [Isa64a]. Non-Stationary [LC65, DBTL68]. Non-Unique [FT66]. Nonatomic [RR72a, RR72b]. Noncentral [And70, Gup71, Har69a, Hod55b, Jam61, Kru54, Lau60, Mui72a, Str72d, dW72]. Noncircular [WD51]. Noncooperative [Sob71]. Nonexistence [BS56, Gil71a, Joh70b, Rag66, Ros71c]. Nonhomogeneous [Ios72]. Nonindependent [Gil71a]. Nonlattice [Sto65a]. Nonleavable [Sud69b]. Nonlinear [AG70, Fab69a, HB65, Mal70, Rod66a]. nonnormal [Fra68b]. Nonnormality [Bra52, DJ51]. Nonoptimality [Kie58]. Nonparametric [BS56, BW67a, BH69, Bha61a, Bha61b, BJ68b, Bhu64, Bic65b, CS58, Dix54, Dwa57a, FW51, Fra53b, Ges70, Gho54, Gup67, HL56, Hol71a, Jog68a, Jur71, Klo62, Kru52, Leh51a, Leh63a, Leh63b, Leh64, LQ65, Mas52a, Mik63, Moo54, Moo64, Mos65, Noe70, Put55, QG68, Ram66a, RS67, RSW68, Ros53, Ros54, Ros56b, Rou69, RSvZ72, Sch69c, Sen68d, Sho67, Sob67, Suk57, Tam66, Tei55b, Tuk48b, Wal50b, Wal53, Wal58c, Wal59c, Whi61a, Wit60, WK69, Woi72, Wol54a]. Nonrandomized [Kie58]. Nonregular [Dav51b]. Nonuniqueness [Ano60i]. Nordisk [Ano31c]. Norm [Bow47]. Normal [Abr64, AJ51, And51a, AB62, And71c, And80, Arn65, Bac63, Bak31, Bak34, Bak40, Bak41, Bak46, Bak49, Bal67, Bar70, Bec54, BS54, Ber69b, Bha59, Bha66, Bha67b, BA49, Bow46, BC64, BK70, Cad53, CZ64, Che65a, Che65b, Chi56, Chu56, Coh50, Con67d, CL65, CD62, Dal40, Dan39, Dav63b, Dia63, Dic71, Dix48, Dix53, Dix57, Dix60, Dwy40a, Ell64, Ell65, Enl34, Fel32, Fel45b, Fer58, Fin46, Fis70, Gar69, Gei56b, GO62, Gir64a, Gir64b, Gir65b, Goo56, Goo57a, Gor41b, GM72, Gov63, Gov66, Gre50a, Gup61, Gup63a, Gup63b, Gup65, GW65, Gup69, Gup70, Gut57, Gut59, Gut70b, Hal51, HH63, Har67a, Her67a, Hey67, Hoe48a, Hoe41, Hsu40, Hsu41, Hsu46, Jam54b, Jam64, Jog70, Jon48, Jos67b, Jos69a, Kac48, Kad68, Kap43, KM70, Kel71, Key38]. Normal [Kha62b, Kha64b, Kha66, Kha67a, Kha73, KS65, KS72b, KG70, Klo63, Koo42, KJB67a, KJB67b, KOvE72, Kul35c, Lad39, Lah57a, Lan71, LC65, Len39, Lin70, Loh69, Luk42, LK54, Mad40a, Mar64, MS72a, Mau40, MG55, McF55, Mes68, Mic59, Mik62, Mil67b, Moo51a, Mor62, Mus68b, Nar50, NP66, Nic56, Nic58, Obe47, Owe56, OS62, Pac61, PB70, PL70, PL71, Pau52b, Pau63, Pau64b, Pau69, Pil50, PR54, PN71, Pol34, Pra63, Que36, Rag65a, Ram66b, Rao58a, Rao59, Rei63b, Rid31, Rie39, Rio49, RP49, RP58, RR59, RC60b, RC60c, Rub60a, Rub60b, Rub61a, Rub61b, Rub62a, SG58, SG69, Sch42a, SMR72, Sco67, Sco68, Sea55, See72a, See72b, Ses66a, SO56, Sev57, SK61, SK63, Sha63, Sha68, Sho67, Šid68, Sim43, Sim67c, SW49, SA57, Sta56, Sta38, Ste58a]. Normal [Ste58b, Ste59a, Ste60, SW47, SC71b, Str71a, Str71b, Str72c, Tal63, Tat53, Tay53, Tei56, Tei57, Tei61c, Tho38, Ton69, Tru53, VA43, Vot48, WW46, Wal59a, Wal60, Wal48c, Wat59, Wei55b, Wij67, Wil46a, Wil41b, Wil65, Wol46a, Wol49b, Wol50, Zac66b, Zac67, Zac70a, ZS60, Zid69, Zid71, BM58b, Fra68b, LS48]. Normality [Ber72c, Ber73, Bha65, Blu63, Cac67, CS58, Dix54, DH69, Dwa53, Dwa55b, Eic63a, Fab68a, Háj68, HB70, Hög72, HP36, Jog68a, KKW55, Kou70b, Le 70a, Mal72, Moo68b, RSvZ72, Sen68a, Sho69, Sie68, Tam67, Van66a]. Normalization [Pau42a, GL46]. Normalized [BD69, Kes70, PR68]. Normalizing [Lau60]. Normally [Aro47, Bak37, Bha72, Cra47a, Gor41a, Kaw50, Lah56b, Mat49]. Norman [Str66b]. Normed [Fer71, Sie69a, Ste71b, Tay72b]. Norming [Kim68]. Norms [Tho57]. Norton [Sad51]. not-quite [Tuk58a]. Note [Agg72, Ahl70, Aig71, Alb47, Alb48, AJ65, Ano31b, Ano33, Ano60i, Aus67, Bac38, Bah58, Bah66, BW64a, Bak35a, Ban45, Ban49b, Ban64, Ban69, BP67a, BP69, Bas55b, BW67a, Bel47, Ber71a, Ber71c, Ber72d, Ber63b, Bha59, BB67, BS69, Bis39, BP72, Bly58, Bly59, Bos49, Bow44, BN61, Bri38, Bri63b, Bro59, Bro40a, Bro71c, BF63b, Bur72b, Bur67, Cam37b, Car50, Cha64, Che67a, Chu61, Cla37, Coc40b, Coh68, Cox56, Cra31a, Cra41b, Cra41a, Cra43a, Cra43b, Cro64, CL71, Cur42, Cur46, Dav49, Dav66, Dav51a, De 38, Den69, DT68, DS66, Dha69, Die42, Dub66, Dwa55a, Dym66, Eat70c, Efr65b, Eis72, Eis39b, Eva40, Fab72, Fel45a, Fin71a, Fis39, Fis70, Fre66b, FG60, Gän70, Gei42a, Gei56b, Ghu58, Gil64b, Gir68a, Gir42a, Gir42b, Gla59, Gle66b, Gle66a, Gle68]. Note [God49a, Gof53, Gou63, Gre47, Gre50b, GCM65, Gut38, Har47, HST53, Hem52, Hol69, Hor71, Hot44a, Hsu48, Hua72, Ios69, Isa69a, Jai66, JK70, Jen70a, Jos68b, Jos71d, Jos71f, Jos72, Kad73, Kal54, Kal59a, Kap71, KMP59, Kat63a, Kat68, Kav41, Kha73, Kim44, Kim53, Kin48a, Kir37, KS69, KR66, Kri71b, Ksh58, Kul35b, Kul35a, Kul39a, Kul39b, KK66, Kun67, Kur69, Kus66, Lam71a, Lan44, Law71, Le 54, Lie53a, Mac41, Mad45, Mai66, Mai69, Mal72, Man58, Man45a, Man69a, Mas50b, Mas50a, Mas51b, Mas52a, MG55, Mei69, Mes65, Mey69, Mil67a, Mil68c, Moo39, Mos41, MS64a, Mur56, Mye34, Nev72, Nii72, O’B72, Oga51, Old52, OT65, OR66, Olm40, Pac55, Par72a, Pau42b, Pau43, Pau47, Pau70, PL46, Pic51, PB71, Pie69a, Pol72, Pra62, Rao58b]. Note [RP67, Red51, Rei63a, Rob65b, Rob71b, Rog59, Ros62a, Roy56, RR59, RM71, Sac72, SY72, Sac67, Sam63c, Sam65a, SG69, SS67, Sax67, Sch70e, Sco50, Sco67, Sco68, See72a, Sei66, Sen70a, Sen71c, Sha67, Sha68, Sha70b, Sha65b, Sid67b, Sim51, Sin66, Smi57, Spj68b, Spr56, Spr57, Sta57, Ste46, Ste67, Ste69e, Str72c, Sub66, Sud69a, Sug71, cS67, Tak69b, Teu72, The64, Thi31b, Tho56, Tho58, Tin45, VJ66, Wal40b, WW41, Wal44a, Wal47b, Wal49a, WK70, Wat71a, Wau45, Weg69, Weg71, Wei55a, Wer37, Wer39, Wic71, Wil44, Wol72, Wol44b, Woo70a, Ylv68a, Zar65, Zei50, Zel54, dH70, dW71, van65, BM58b, Mad71, PS69d, Rao65, Sac65, Sch44, Tru47]. Notes [Abe69b, AG60, AG65, Ano69g, Aro59, Arv70, Bel70a, Ber66a, Ber73, Bic65a, Bre60a, Bre70, Bro71f, Cam38b, Cha65, Cla71, Csö65a, Dar62, Day65, Day66a, DL70a, DL72, Dok68, DKW59, Dwy60, Enn73, Fer63, Fie71, Fre68, Geb68, Gir64a, Gle68, Gna60, Goo61b, Goo76, Gru36, Gup62, Gup70, Gur59, Gur62, Har60b, Her70, Hil61, Hsu38, Iye61b, Iye61a, Joh53, Jor59, Kha65a, KS72b, KHS74, KA65, KP68c, KP60b, Kul68a, Kum65, LaM72, Lan64, Mar63, MH70, Nig76, Obe72a, Oka68, Per71, Que56, RR72a, Rid59a, Ros63a, Rub61a, Sak66, San62, Sco68, SP68, Set70a, Sha59a, SK63, Sri65a, Sri68b, Sri69, Ste59a, Ste73, Tuc59a, Tuk62a, Wal60, Wij59b, Wij75]. Notice [Ano35d, Cam38a]. Notices [Ano43e, Ano43f, Ano43g, Ano43h, Ano44e, Ano44f, Ano44g, Ano45d, Ano45e, Ano45f, Ano46h, Ano46i, Ano46j, Ano46k, Ano47g, Ano47h, Ano47i, Ano47j, Ano48f, Ano48g, Ano48h, Ano48i, Ano49g, Ano49h, Ano49i, Ano49j, Ano49k, Ano50e, Ano50f, Ano50g, Ano50h, Ano51e, Ano51f, Ano51g, Ano51h, Ano52f, Ano52g, Ano52h, Ano52i, Ano53f, Ano53g, Ano53h, Ano53i, Ano54h, Ano54i, Ano54j, Ano54k, Ano55e, Ano55f, Ano55g, Ano55h, Ano56f, Ano56g, Ano56h, Ano56i, Ano57e, Ano57f, Ano57g, Ano57h, Ano58e, Ano58f, Ano58g, Ano58h, Ano59e, Ano59f, Ano59g, Ano59h, Ano60e, Ano60f, Ano60g, Ano60h, Ano61f, Ano61g, Ano61h, Ano61i, Ano62i, Ano62j, Ano62k, Ano62l, Ano63e, Ano63f, Ano63g, Ano63h, Ano64e, Ano64f, Ano64g, Ano64h, Ano65g]. Notices [Ano65h, Ano65i, Ano65j, Ano65k, Ano65l, Ano66g, Ano66h, Ano66i, Ano66j, Ano66k, Ano66l, Ano67g, Ano67h, Ano67i, Ano67j, Ano67k, Ano67l, Ano68g, Ano68h, Ano68i, Ano68j, Ano68k, Ano68l, Ano69h, Ano69i, Ano69j, Ano69k, Ano69l, Ano69m, Ano70g, Ano70h, Ano70i, Ano70j, Ano70k, Ano70l, Ano71g, Ano71h, Ano71i, Ano71j, Ano71k, Ano71l]. Nuisance [Fra56a, Hot40a, Wei57d]. Null [Con67c, Dav70, JK62, Kha64b, Klo59, Oga63, Sio71, SPZ56, SF69].

Null-Distribution [Oga63]. Number [Abe69a, Abe69b, Agr66a, AB66, AFLP57, AFPR59, Bla55b, Bra65, Bre58b, Bri72, Bur72a, Chu43a, Chu43c, CL65, Cur46, Dar64a, Enn70, Enn73, Eps54a, Fer36, Fis33a, Gem72, Gle69, Goo49, GvS50, Hed70, HSF72, Hoe53a, Hoe56, Kat52, LC65, MW42, Mv67, Pat63, PJ70, RF71, Rio64, Sam65c, Sar57, SS66b, Sch72, Sei55a, Sei55b, Sei56, Sha64b, Sha65a, Sim67b, SA57, Szá72, Tak59, Tho72, Tod70, Wal46b, Wal48b, Wei61, Woo66, Ylv65, vS51, HS49b]. Numbers [Bau63a, Bau63b, Bax61, Ber62b, Ber63a, Ber71d, Bre60b, Cho65, Cho67a, Cop32, Dwy40b, FM58, Fel45a, Gei40, Gum43a, HK65b, HK65a, Han67, Joh64a, KT61, Kim44, Koo61, Koo66, Kur72, Nor40, Old38, RR64b, Tay72b, Tom72, Wag69, WS34, Cam52, Hor48, Kat68, LeV66a]. Numerical [Cam37b, Coh40, CD62, Geb68, Gre49, Luc68, AT63].

O [Dav63a, Fre61, Owe67a, Wel64a]. Oak [Mos51]. Objects [Bor67]. Observation [Dok67b, HD54, HW48, KS68, Leh64, PR54, SY70b, vM45]. Observational [Wil30]. Observations [And64b, Ant62, BF63a, Ban70, Bha64, BJK56, Che52, Cra31c, Dow54, ET53, FM58, Geb64b, Gei42b, Gru50, Gue58, Hen32a, Hoe52, Hsu46, Jon62, KS62, Kav41, Kim40, Mac59, Mus68b, Sau63b, Sch65, Set50, Thi31k, Tho35, WW44, Wal46b, Wal48b, Wal50b, Wal53, Wei61, Wij67]. Observed [Fer36, HS72b, Olm40]. Obstructive [Wij72]. Obtain [Sha64a]. Obtainable [Far72, Rod71]. Obtained [Dod38, Rie31]. Obtaining [Cam37a, Dre41, Enn70, Enn73, GW68, Joh67b, Maz67, NR67, Som58, Str71b, Str72c, Hor48]. OC [Hoe70]. Occasions [Raj65]. Occupancy [Ail70, ESM72, FP56, Hol71b, Par72a, Wei58a]. Occupation [Ber67b, Wei72b]. Occurrence [Bat55, Chu41]. Occurrences [Olm40]. Occurring [BPSW70, KZ66, Ken53, Man45b, Man46]. Odd [Chu46b, Rao66]. Oddsmaking [HS72a]. Office [Tuk47b]. Officers [Dwy49a]. Omission [Sad51]. One [Abr67, All53, Arn65, Ath69, BD66, Bic64, BT51, Bir53b, BP72, BF72, BP65b, Box54a, BLS64, Bro45, Bro66, Bur59, Cha58, Cha65, Cho66a, Con67b, CS39, DS63, DK56, DK58, Den66, DT71, Eat70a, Far64a, Far69a, Fis71, Fis72a, FP69a, Gre44b, Gut70a, Har61, Har64a, Igl65, Jof67a, Klo63, KP69, KA65, Kul63, Leh47b, Leh64, Leh63d, Mar60, MO60b, MP67, Mij72, Mik62, NT70, O’T32, Pau63, Pau67, Per69a, Per71, Pra58b, Qua65, Rai72, Ram66b, RG59a, RG59b, Sav59, Sch42b, Sch34, SP67, SP68, Sen70b, SG71, SM50, Sid63, Sie67a, Sie68, Sie69b, Sim43, Sim67a, SW49, SN68, Tan70, Wal44b, Wel64a, Whi61b, Wij67, Blu54a, Dok67b, Hag67, HS70c, MW47, Wei53b]. One- [Abr67]. One-Dimensional [Gut70a, Sch34]. One-Parameter [BP65b]. One-Sample [BD66, MP67, Qua65, Sav59, Sen70b, SG71, DT71]. One-Sided [BT51, Bir53b, BF72, Cha58, Cha65, Eat70a, Mar60, MO60b, Per69a, Per71, Pra58b, Sie67a, Sie69b, Wel64a, Whi61b, All53, Cho66a, Con67b, Mik62, Ram66b, Sie68]. One-Way [Box54a, Fis72a, KP69, Fis71]. Only [Bla58b, Suk58]. Ontario [Ano53e]. Operating [Boh68, FGW46, Kin52, LD66, WK62]. Operational [Ban60, Ols67b]. Operator [Jam68, Kli69, OS70b, Sta65, ZD38]. Operators [Cog62, Koc58, MR65]. Opinion [Man45b, Man46, Sto61b]. Opinions [BD62]. Optimal [Atw69, BP67b, BD66, BY68, Bla54, Boz71, Che53, DLT72, Dar65c, DA59, DS69a, Dok67a, Dok68, DT67, EZ63, Fre67f, FP67, Gup62, Hag66, Hag67, Her67a, Hoe65a, HL64, Hoe66, JR70, KM71, KS66a, MT65, Mik63, Mur71a, Mur71b, RF70a, Rob63, Rob64b, Ros71c, RM68, Sch69g, She69, Sie67b, SSF68, Sny72, Sta72b, Stu68, Stu71a, Stu71b, Sud71b, Sud72, Tay68, Tay70, TBO71, TO72, Vei66, Yah66]. Optimality [Beh71, Bro71e, DM68a, Gho71b, Hil69a, Jaq72, Kie58, Kie62, Lig71, Mat63, MR69b, Obe72b, OK68, SMR72, Sha60c, Spj72b, Vei69]. Optimally [Alb70]. Optimals [Rag70b]. Optimization [Ric66]. Optimum [And65c, BC50, BW63, Eat67b, Elf52, Enn73, Gau59, Gho63a, Gir58, God60, Gup65, Gup70, Gut57, Gut59, HH49, Han68, Her59, HP49, Hoe65c, Hub63a, Kal59b, Kem69, KW59b, Kie61, KS63, Lec62, Leh47b, Leh59, Man69b, Owe70, Pau52b, Rag59, SA66, Sal66, Sid63, Spj67, Tru53, WW48, WD39, Wol66, Won68, Wyn70]. Optional [JO69]. Order [Ala72, AC65, BP66a, BGP69, BN63, Beh71, Beh72, BBB60, Bha72, Bha70, BD63, BD59, BB60, Bri38, Bur62a, CGJ67, CMPT60, Col51, Dan71, Dav62, Dav63b, Dav65, Dav66, DJ68, Des71, DL70a, DL70b, DL72, Dra60a, Dra60c, Dra60b, Dra61, DH68, Dwa53, Dwa56, Dwa57b, Dwa67b, GGH59, GS71, Gho71b, God49b, Goo58a, Goo58b, Goo59, Gov62, Gov63, Gov66, Gri31, Gup61, GS65, Gut46a, Háj62, Haq70, Har61, Har64a, Her64, Her67c, Hoe53b, JS72, Jon48, Jos71a, Kad71, Kad73, KR66, Kul63, Lam64, Lau63, Lie53b, Lie55, MW43, Mar64, MM69, MR64, Moo68b, OS62, Pol70a, Pol70b, PS69c, Ram62, RSS60, Rog59, Ros70a, RO72, Sal66, Sar54, Sar55a, Sar55b, SG56, SG58, SG69, Sav56, Sav57a, Sav59, Sav60, SS66a, SS67, Sch69g, Sea56a]. Order [Sen67, SP67, Sen68b, SP68, Sen69, SP70, Sen70a, Sen70b, Sen71a, SG71, Sen72b, Set70b, Sha66a, Sha70b, SP69, Sha71b, Sho69, Sho72b, Sid63, Sri62, SHB64, SP58, Ste71a, Sti69, Str61, Tak64, Tan64, TC65, TC80, Tei56, Tei61c, Ter52, Wal46d, Wal46e, Wei71, Wei53a, Wel71, Wel72, Wen60, WGL66, AT63, BS68, Con65a, HMTW47, ST45, Sav71]. Ordered [BJ70, BC68b, Car70, Cha63a, CW58, CW59, Dow54, ET53, Gab62, GS58b, Kat63c, Leh55, Pfa64, Pfa69b, PR54, PS68a, Sam65a, Sho67, You41]. Ordering [Hol67, SSW66, Whi61a]. Orderings [Fel72, Kon66, Swi64]. orderly [Tuk58b]. Orders [Mil67b, Sär65, Sav60, SSW66]. Ordinal [vS51]. Ordinary [WZ65]. Ordinate [Gor41b]. Orey [BF64c]. Organic [Wil36]. Organization [Ano35d]. Oriented [DG57a]. Origin [Pit56]. Original [Tho38]. Origins [Ore64]. Ornstein [Met69, Str58]. Orthant [Abr64, Bac63, Che69]. Orthogonal [AK61, BF63a, Bea41, BB52, Bra65, Bus52b, Fre66d, HS70a, HF71, Jam54b, Jor32, Jor59, Kim40, Kim53, Man43, Nel70, Sch67e, Sei54, Sei55a, Sei55b, Sei56, SZ66, TK70, Wat69, Wij57, Wij59c, Man42]. Orthogonality [Seb64b]. Orthogonalization [Won35]. Oscillating [Fre67a]. Oscillations [Str72b].

Osculatory [Rob37a]. Oskar [Hur48, Wol61]. Other [Hot44b, KKW55, KS65, KS72b, Kis45, MW47, RP63, RC60b, RC60c, Ste45b, TC65, Tau62, Tho36]. O’Toole [ZD38]. Our [Cra60]. Outcomes [Rob68]. Outgoing [WW45]. Outliers [Ste71b]. Outlying [Geb64b, Gru50]. Output [Bur68a, Dal68]. Overlapping [CMPT60, da 45]. Owen [Pea64, Wel64a].

P [Han66a, Owe67b, Spi66]. P. [Beu65]. P.B.I.B. [Agg72, Tho58]. Page [NR71]. Paired [Gho71b, Gil64a, Gut46a, HL54, Joh67a, Kad66, Meh64, SD68, SP69, ST68, TR64, TD63, Hub63b]. Paired-Comparison [Meh64, TD63, Hub63b]. Pairing [Wal47e]. Pairs [Rai72, Sav57b]. Pairwise [BH63, GM62, Hub63a, Lan65]. Paper [Byr35b, CL56, Dem40, Erd50, Gau59, Hub63b, Ken41, Man45a, Man46, NCC65, Wer37, Zoc37, dH70]. Papers [Ano39a, Ano40a, Ano40b, Ano41a, Ano41b, Ano42a, Ano42b, Ano43a, Ano44a, Ano45a, Ano46a, Ano47a, Ano47b, Ano47c, Ano48a, Ano48b, Ano48c, Ano49a, Ano49b, Ano49c, Ano49d, Ano50a, Ano50b, Ano50c, Ano50d, Ano51a, Ano51b, Ano51c, Ano51d, Ano52a, Ano52b, Ano52c, Ano52d, Ano53a, Ano53b, Ano53c, Ano54a, Ano54b, Ano54c, Ano54d, Ano55a, Ano55b, Ano55c, Ano56a, Ano56b, Ano56c, Ano56d, Ano57a, Ano57b, Ano57c, Ano58b, Ano58c, Ano58d, Ano59a, Ano59b, Ano59c, Ano59d, Ano60a, Ano60b, Ano60c, Ano60d, Ano61a, Ano61b, Ano61c, Ano61d, Ano62a, Ano62b, Ano62c, Ano62d, Ano63a, Ano63b, Ano63c, Ano63d, Ano64a, Ano64b, Ano64c, Ano64d, Ano65a, Ano65b, Ano65c, Ano65d, Ano65e, Ano65f, Ano66b, Ano66c, Ano66d, Ano66e, Ano66f]. Papers [Ano67d, Ano67e, Ano67f, Ano68a, Ano68b, Ano68c, Ano68d, Ano68e, Ano68f, Ano69a, Ano69b, Ano69c, Ano69d, Ano69e, Ano69f, Ano70b, Ano70c, Ano70d, Ano70e, Ano70f, Ano71b, Ano71c, Ano71d, Ano71e, Ano71f, Gup62, Ios72, Sch61a, vMD41, Coc44]. Parabolic [Joh40a]. Paradox [Dem63c, Usi64, Rob61]. Parallel [Dur68, Gan61, Kin61a]. Parallelism [Sen69]. Parameter [Bas64, BG69b, Ber70a, Bha67a, BM61, Bla51b, BC68a, BP65b, Dos69, DKW53b, DKW53a, DKW59, Far64b, Fed68, FR64, Fox70, Fra52b, Fre63a, HS72b, Hub64, Jos70, Jos71d, Jur69, Klo70, Kur69, Kus69, Lec62, LS53, Leh63b, Mal68a, Mee70, Mee71, NR71, Neu71, Ney51, Par70, PR71a, Pfa69a, Pie70, Rob58, Rub63a, Sac70, SY72, Sim67a, Sim71b, Sri68d, Ste59b, Str61, Tak69b, Val68, Wal48a, Wal48b, WGL66, Win72, Woo72, den55, Zim72]. Parameter-Testing [DKW59]. Parameters [Adi67a, Adi67b, Agg72, AS58, Alb66, AR49, AR50, And59, AKKW69, Bli62, BC68b, Bro66, Bru55, Bru58, Car46, Cha56, Che53, Coh51, Coh65a, Dan54a, DV55, Dic71, DS69b, Dud65, Fer62a, Fer63, Fra56a, GS66, Gle65, Gle66b, Har61, Har64a, HM67, Hen34, Hot40a, Isa51, Joh40b, KZ66, Kat61b, KW56a, Kim46, Kim56, Kul63, Len69, MM62, MM63, Maj62b, Man69b, Mes65, Moo69, Mus68a, Mye34, Rao61a, RW68, RT59, SY68, SA66, Sal66, Sam65a, SG56, SG58, SG69, Sch68b, Sei66, Sid63, Sri71, Sto61a, Tat59, Wal48b, Wei57d, Wel39, Wic69, Wil32a, Wil36, Zac68, Bro68]. Parametric [Adi67a, BH69, Bic65a, Blu66b, CLA62, Che56, Der57, Dwa55b, Ela62, Hil60, Hil61, Hoe48b, Joh57b, Kim47, LS49, Mas50a, Mur48, Sam63b, Sch43a, Sch43c, Str72a, Ter52, Yen64, Dar55, ST45, Sen66, Som58, Tuk47a, WW43]. Pareto [Mar63, Cas50, Mar62, Mar64]. Pari [EG59]. Pari-Mutuel [EG59]. Part [SG69, RC60b, RC60c, Sar55a, Sar55b, SG58]. Parthasarathy [Dud69b]. Partial [And72b, Ber37, Bil62, BR67, DLT72, Fis33b, Gra63, Gut38, IW71, Mas70b, Mui70b, PR68, Sie69a, SA57, Str35, Tay72a, Wen60, vS49, vS50]. Partially [Agr66a, AJ56, Bha60, BC55, BM59b, Cha61, CC54, Cor60, GS71, Hin64, Hof63, Joh67b, Leh63d, Mes67, Nai66, Nai69, Oga59, OI65, Rag66, RL57, Sei61, Sei66, Sha59b, Sha60b, Sha66b, Sha67, Shr65, SA70, SA71, Sur68b, Sur71, Zel54]. Partially-Balanced [Hof63]. Particle [Bro70d, Bro71f, GRS55, Tak61b]. Particles [Mal47]. Particular [Dre40, Law33, NV68, Pol34, She71, Sle61, Weg56, Wel39]. Particularly [Var67]. Partition [Wol42]. Partitioned [Wau45]. Partitioning [LR53, See72a, Ton69]. Partitions [Cac66, Car70, Fre66d, Tuk57c]. Partly [GZ61]. Parts [Der66b, Rob47]. Parzen [Smi65]. Passage [DS53, DL66, Gol71, Hey66, Kei63, Kei66, Ken72, Kin61b, Kin61c, Lam60, OS69, Roo72, She67, She71, Sle61]. Passages [RM71]. Past [Wal57]. Path [Sha69a, Wri34]. Paths [Rai72, Sch70a, Sha70d, Ylv68a, Spi66]. Patrick [Haj67]. Pattern [Alb63]. Patterns [Hed70, Swi67, WF57a]. Pauc [Kri71a]. Paul [Kni66]. Paulson [HM68]. Paz [Mad71]. PBIB [Kag72, Mes65, Rao61b, RC62, Sha64b, Sha65a, Var59, Rag60b, Sei63]. PDF [Sri62]. Peakedness [Bir48, Pro65]. Pearson [Bir65, Coc44, Owe67a, Bir64, Bro40a, Car32a, Cas50, CS52, Coh51, Cou44, Cra36a, DW51, HK68, Hil31, Kul36, Ols38, Sal30b, den55]. Pearsonian [Bak40, Dwy49b]. peculiar [Fre67e]. Pelaez [Wen61]. Percentage [Gup61, GS65, Hec60, Koz56, Res62, Wal48c, Wal58c, GL46, Gna67]. Percentiles [She57]. Percolation [Ham57]. Perfect [DZ67, Dar71, Kal59a, Rod66b]. Performance [Eli70, SO68, Sta66b]. Period [Gum41, Pow30]. Periodic [Her63]. Periodically [Ric48]. Periodogram [Alt37, BK58a, Rao67, Sta39]. Periods [Dod30]. Permutation [Wor59]. Permutations [AM67, Hoe52, Rio65, WW44]. Perpetual [Rob36]. Persisting [Dod41a]. Person [DWW51, fH69, fH71, Sac72]. Petersburg [Rob61]. Petrovskii [Bic69b]. Phage [Sri68d]. Phase [Jon69, Sin67]. Physical [Mor32]. Physics [Ros66a]. Pillai [Pil68]. Pitman [Noe55, Ram66b, Ste59b, Wit60, Yu71, Van63]. Pittsburgh [Han43a, Han43b]. Place [Kni63]. Places [Por67, Por68]. Plan [CR65a, Dod43, Lie53a]. Plane [Ber67b, BY65, HS72b, Kin63b, Que49c, Swi65]. Plans [AK61, Ban70, BN61, Cur46, DLS57, Gru49, LS55, Moo43, Nad60, Pre69a, RF70a, RF70b, WW45, WH65, Whi65, Wel64a]. Play [Fre67f, FP67, SW72a, Van66b]. Play-the-Winner [SW72a]. Plots [Kem53]. Poincaré [Chu43b]. Point [BJ68b, BT65, Byr35b, Cam51, Cop32, Far66a, Far66b, Far72, Gho64b, HL50, Lev49a, Mer33, Mil71b, Pet54, RT69, Wal58c, Gol67, Goo57c, Goo61b, Mar72b, Tat54]. Point-Biserial [Tat54]. Points [CE67, Gna67, Gup61, GS65, Hec60, Hot43a, Iye50, Iye61b, KZ66, Kes70, Koz56, Mar48, NV68, Pin47, Res62, Sch72, Stu68, Wal48c, GL46, Iye51]. Pointwise [wK68]. Poisson [Bro71f, Joh69, Sin66, Ahm61, Ano35b, Ben57, Bos66, Bre63, Bro70d, Bro71g, Cac61, Coc40a, Cup69b, Fre62c, FD65, Her48, HL60, Hoe45, Joh51, Kin63a, Kni65c, Lec62, Mac48, McC72, Neu67, Pra60a, Pyk59, Rio37, Sat42a, SJ71, Tak62, TG58, Tei55a, Teu72, The64, Tru47, Wei35]. Poisson-Subordination [Teu72]. Poisson-Type [Cup69b]. Policies [BP64, Fle62a, Lip71, Ros71c, Vei66]. Policy [RS72b]. Polls [Man45b, Man46]. Pólya [Efr65a, Kar57a, Kar57b, Kar58b, KP60a, Ala68, Bat55]. Pólya-type [Kar57b]. Polychotomy [Gho66]. Polykays [Car68, Car70, Day64, Day65, Day66a, Day66b, DT64, Dwy64, NT70, Tra69]. Polynomial [And62a, Ant62, Bar35, Dav33, Eag64, Gir42a, Gir42b, Gue58, Hoe54, Hoe58, Hoe61, HL64, Hoe65c, Hsu44, Jam68, Lah56b, Lev66b, Mur71a, RP63, Sam41, SV69b, Stu71b, Wah71b, War60, dlG54]. Polynomial-Product [Hsu44]. Polynomials [Bea37, Bea41, Dwy40b, Gre32, Gue51, Hil31, Jor32, Jor59, Kha62b, Kha73, Kim40, McP41, Sug71, Woo70a]. Polyspectra [Bri65]. Polytopes [Dem72]. Pool [Sto61b]. Pooling [Coh68, Hun55, Pau50]. Population [Agg66, AJ51, Bae45, Bak35a, Bak49, Bha66, Bir50, Boy58, Bur55, Bur67, Cad53, Cad54, Cam46, CJ56, Cra31a, DR67, Dix40, Dix57, Eri70b, Fab62, Fel32, Fer36, Fin46, Goo49, Goo53a, Goo68, GT64b, Háj64, Har59b, Har60b, Har64a, Hsu40, Hsu46, Jos67a, Jos67b, Jos68a, Jos69a, Jos71b, Jos71f, Kac48, Kul35c, Len39, Loh69, Lot39b, Lot42, Mat43, Moo41, Moo43, O’T34, Pac61, Pau64b, Pil50, PR54, Raj53, Rao54, Rid59a, Rid59b, RS67, Rob48c, RC60a, Sam68a, San72, Sim43, Sim67c, Sin70, Sin71, Sob67, Sri69, Tsa56, WW40, Wat62, Wat63, Wei52, Wil32a, Wol46a, Mos48a, Roy70, ST44]. Populations [Agg59, Bah50, Bak31, Bak32, Bak40, Bak41, Bec54, BS54, Beh61, Bhu64, BC50, Bro33, Car68, Cha63c, Cha54, Coc46, Coh50, CMPT60, DeG59, Dor43, Dun31, ET53, Fel35, GJ65, God69, Goo64b, Goo76, Gov63, GS58a, Gup65, Gup70, Gup72, HH43, Han68, Har61, HM67, Hsu41, Joh51, Jos65a, Jos65b, Jos66a, Jos69b, Jos71c, Jos71e, Jos72, Kha64b, Kon56, Kon58, KOvE72, Kup58a, Kup58b, LPC70, Mar64, Mor72, MR66, Nüe66, Pau64b, Pau67, Pau69, Pau70, Pie40, PN71, Rei63b, Rie39, Sch42a, Sco71, Sea55, See72a, SW72a, SHB64, Sri67, Sta38, Ste57b, Stu67, Suk58, Tal63, Tho36, Ton69, Wal47d, Wal48c, Wal62, Wat59, Wil65]. Populations-IV [Jos66a]. Positions [Rio64]. Positive [BF72, Bur72a, Dyk70b, Gan66, Kha66, Mai69, Rob65b, Sha70a, Sin64, Ste45b, Suk72, Woo66, Sin71]. Positivity [Mar71b]. Possible [KZ66, Ste69b, Ste73]. Possibly [RC60b, RC60c]. Post [BCD^{+}45, Gho63b]. Post-War [BCD^{+}45]. Posterior [Ber66a, Ber66b, Dem66, Joh67c, Joh70c, Kle71, Sco71, Sto65c, Sto70a, Sto63]. Posteriori [Ber70b, Thi31r]. Postulate [Hen32a, Wer37, Zoc35a]. Potential [Cur68, Dym70, Sch67a]. Potentials [KS61, Kni66]. Poughkeepsie [Old42d]. Power [And67, AG64b, AG64c, AG65, Arn65, Ber58a, Ber58b, Bir53b, BBH56, Bro39, Bue65, Dan40, Dia63, Dix53, Dix54, Dok69b, DT71, Dwa56, Dwy38a, Dwy38c, Efr67, Ela62, FM58, GW68, GAM64, Gup69, Gut57, Gut59, HK64b, HG66, HSF72, Hoe52, Hsu45d, Joh67d, KS62, Klo63, Klo64, Kon56, Kon58, KP70, Leh53, Leh58, Lev52, LF70, Mac41, Mas50a, Mas52a, Mit58, Mud65, Nic54, Qua65, RM61, Ste45a, Tei55b, Tsa57, Wal42b, Wal46c, Wal47e, Wal49e, Wei57b, Wol49b, uD54, Fox56, HL68, ZD59]. Powered [HF70, Maz67, Rak67]. Powerful [Ahm61, Cap61, Fra57a, Gas65, Háj62, JR69, JM72, Kra71, LS48, SS66b, Sch68a, Sen67, Spj68a, Ter52, Uza60, Wal41a, Wal41c]. Powers [Goo61a, Kim50, Sha58b, Sha60d, Ste45b]. PP [Fre61]. PPS [Raj64]. Prabhu [Dwa67a, Mor66]. Practical [Ano35c, Rob35]. Practice [Dem65]. Pre [Wel64b, Zac68]. Pre-Assigned [Zac68]. Pre-Emptive [Wel64b]. Preassigned [Ban67]. Precision [BR69, Fel32, Gru33, Lin60]. Predicting [Whe31]. Prediction [Beu63, HL64, HB67, Hot40a, KB71, LB63, Rob65a, Sny72, Wil42, Wol48, dO66, Beu65, Sal67]. Predictive [Lin60]. Predictor [Bha62, Rob71a]. Predictors [GS71, Kim53]. Preference [Fis67b]. Preference-based [Fis67b]. Preliminary [Ban44, Hun55, Pau50, Pyk61a, SMR72, Cam52]. Preliminary-Test [SMR72]. Preparedness [Ano40g]. Prescribed [BH60a, MG55, RP65, WW45, WK69, Zac66b, Zac67]. Presence [Hal51, Kac45a, KW56a, Mor62, Pfa69a, Rao68c, SC63, Wei57d]. Present [Dal64b, Hem52]. Presentation [Dwy44b, Mol31]. Preservation [Top67]. President [Ano46c, Coc47, Cra43c, Cra44a, Doo51, Dwy52, Fel48b, Ney50, She45, Wal49b]. Preventive [Fle62a]. Previous [HL52]. Price [Bur32]. Prices [Car34a]. Primary [Kim42, Sen55]. Prime [HF70, HSF72, Maz67, Rak67]. Prime-Powered [HF70, Maz67]. Primes [HSF72]. Princeton [Dwy46a]. Principal [And63a, Dar65c, Obe72b, OK68, Rei40, Sug65, Sug66]. Principle [Cam70, Cou44, Dro72, Hey69b, Jor32, Jor59, Kue68, Lam62, Wes59b]. Principles [Dem65, Fre67d, Kar58b, Leh50, Dar65a]. Prior [Fer72b, Har64b, Har65b, Owe70, Sch67b, SC63]. Priori [BR67, Eis39b, Hol67, Thi31r, Vil69a]. Priorities [Joh65]. Priority [Bau63a, BD65a, Ber73, Bir65, God61a, Kin61b, KS72c, KA64a, MM63, Mik60a, Mil60, PL71, Pra66b, Rao59, Rob47, Sta59, Ste60, SN70, Wel64b, Whi58a]. Priors [Fre66b]. Probabilistic [BP66c, BPSW70, Har66b, Kri70, Paz70, Pet69, Por65, Tsa68, LeV66a]. Probabilities [Abr64, AK71b, Bac63, Ber71b, Ber71a, Bla55b, BF68c, Boo72, Cam70, Cha72, Che69, Chu43a, Chu43c, Dav68c, DG57a, Dem67, Dem68, Dic68a, Dix48, DF66a, EG59, Fis69a, FH55, Fra51a, Fra53a, Goo58b, Goo68, Háj64, HH49, HW71, HvN42, HR62, JN63, Jog67, JS68, Jog70, Kat63c, Kle71, Kup58a, Kup58b, Leh44, MZ65, MD71, Mil67b, NR62, Nii72, Nor67, Ols68, Owe56, RS70, Rod66b, Ros72a, Ros72b, Sav60, SSW66, Šid68, Šid71, Sta72a, Ste58a, Ste59a, Ste71a, Tei55a, Thi31r, Usi64, Wei54b, ZS60, Kal65, Luk65]. Probability [Ahl70, And60, AA63, Aro47, Bak35b, BMP63, Bar68, BG69a, Bar69, Bar43, Bat30, Bat34, BEM59, Ber51b, BS72, BT51, Bir53b, BH59, Blu54a, Boa49, Boe69, Bri62, Byr35b, Cad54, Cam37a, CC47, CL56, Che66, Chu41, Chu43b, CH45, Cog62, CMPT60, Cop32, Cov66, CE67, Cra47c, Cra53, Csö65c, Dar71, DK56, DK58, DeG61, Dem69, Den66, Doo41, Dud68, Dur68, Eic70, Enn70, Enn73, Fer72a, Fin71a, Fis67b, Fis69b, FV69, Gei38, Gei39, Gei42a, Gei44, GO69, Gil72, Gil64b, Gir58, Gor41b, GG39, GP67, Gup63b, Had63, Ham57, Har60a, Hat49, Hel70, HB71, Hoa67, HY64, Hub65, Ifr66, Iye50, Iye61b, Joh51, Kal59a, Kat55, KP57, Kat63b, Kei72, Kem68a, KW57, KHS72, KHS74, Kim42, Kim44, Kim47, KQ72, Kot61, Kou68, KPS59, Kul68a]. Probability [Kul68b, Law33, Le 70b, LeP70, Led71, Lin66, Loè50, Luc67, Luc68, Luk65, Mat63, Mor52, Mur65b, NR67, Nev72, New68, Ols67b, Ore64, Par62, PR71a, Pic69, Por64a, Pra59, Pru71, Raj65, Ray57, Ric48, Rob68, Rod66b, Rub60a, Rub60b, Rub61a, Rub61b, Rub62a, Rue71, SS66a, SS67, Saw72, Sch47, Sch64, Sch69e, Sch67e, Set64, Set65, Set70a, Set70b, Sha55, Sie69c, Sim71a, Ski68, SU72, Spi67, Spr53, Ste71a, Sto65c, Sto70a, Str65a, Sub66, Swi64, Tak58, Thi31q, Thi31p, Tom72, Ull44, Vot46, Wal46b, WW48, WL63, Wei56a, Wen68, Wer39, Whe56, Whi70, Wij60, Wij63, Wij67, Wij68a, Wij71, Wij75, Wil46b, WK69, Wol66, Woo49, Yar72, vMD41, vM41, Cra62, Iye51, Vil64b, Dud69b, Han66b, Kni66, Blu67a, Gre64, Jai70, Kal69, RS72a]. Probable [AT72, Fré47a, Hsu44]. Problem [Ail70, And62a, And64b, BR50, Bah50, BR51, Ban67, Bat67, Bat42, BB65, Ber42, Bha61a, Bha64, Bic64, BG46, Bla51b, BK57b, Blu63, Bow44, Bre57c, Bri63a, Bro70c, Bur60, Cau72, Cha35, Coh65b, Dal41, DS53, Des70b, Dix44, Dun61, Dur71, Dwa56, DL66, Eat67a, Eav69, Ehr70, ESM72, Ell62, Eps49b, Far69a, Fel62, Fin66, Fra53c, Fri40, Gas65, Gil59b, GH69, Gir64a, Gir64b, Gir65b, Gir68b, God61b, Gre38, Gre41, Hal60, Ham72, HR55a, HV65, HH72b, Hen35b, Hey66, HP49, Hot40a, Isb59, JR51, JM72, Kad71, Kad73, Kal59b, KMP59, Kem68b, Kem48, KS39, Ken40, Ken41, Kha54, Kru52, Kul34, Kul39a, Kum65, Lev49b, Lev66b, Mac61, Mal47, Man45b, Man46, Man53, MT65, MR57, Moo46, Moo68a, Mor72, Mul56, Ney62, Ney35, Ney41, Nüe66]. Problem [Par72a, Pau52b, Per69b, Pot63, Pra61, Qua66, Rag65b, RSS67, Rob65a, Sam63a, Sam63c, Sam64, Sca59, Sch43a, Sch43b, Sch60a, Sch69d, Sei54, SP67, SP68, She67, SP65, Sri65a, Sri67, SO68, Ste57a, Tak58, Tak61b, Tay70, Tho62, TBO71, TO72, Van66d, Van66c, Var58, Vog60a, Wah71a, Wal44b, Wal49e, Wei56b, Wes59a, Wil42, Wis37, vdW60, vM43, EO72, Sch44, Tuk58b, vM38]. Problems [Alb54, And46, And64a, ACH64, BS56, Bar65, Bha68, Blu66b, Box54a, Box54b, BT65, Bro68, Bro71d, Cha62b, Che67a, Cog67, Con63b, Cox58, Cra47c, Dar53b, Dar62, Dud69a, Dun61, DKW53b, DKW53a, DKW59, Efr65b, FP71, Gho54, GMS46, Goo53a, GT64a, GS57, GT64b, Har59a, Har70c, Hey67, HL50, Hoe58, Hol71b, IT68, JV71, JV72, Joh54, Joh71, Kar57b, KT60a, Kha66, Kha67b, KW59b, Kie61, KS65, KS72b, Kim57, Kno67, Lah57b, Leh57a, Leh57b, Leh61, Lor72, Mah67, Man58, Mat55, McC47, Mol46, Mud65, Mus68a, Nel66, Oat72, OS69, Pat63, Pau49, Per69a, Per71, PB71, PJ71, PP69, Que49c, Rad62, RH71, Rob64a, RS72b, RR59, Sac63, SY66, SY68, SY70a, San68, Sch68a, Sch69a, She69, Sie67b, Sob53, Ste70, Sud71c, Tod70, Try58, Uza60]. Problems [WW50, Wei58a, Wij57, Wij59c, Ylv62, Ylv64]. Procedure [AF62, Ban70, Bec54, BS54, BEM59, Che67b, Der56a, DK72, Fie70, Fie71, Hoe60, HM68, Hol66, Hua72, Hub63a, Hun55, Kar58b, Meh72, Mos58, NR71, Owe53, PF70, PF72, Pau49, Pau62, Pau63, Pau64b, Pau70, RK57, Ray65, RSS68, Rob64b, Roy58, Sak64, Sak66, Sau63b, Sau63a, SW49, Sta66a, Sta66b, Ste70, Tsa54b, Ven67b, Wor60, Ven67a]. Procedures [Agg55, Agg59, Agg66, And64b, ACH64, BS56, BG69a, Bha61b, BY68, Bly51, BBH56, Cha62b, Com64, DJL59, Dok66, Dok67b, DWW51, Eat67b, Efr65b, Fox71, FK51, Gab69, Ghu58, GKR57, GP72, Gup72, GJ59, Hag67, HH72b, Hea56, Hoe71, Jog66, Joh57b, Joh54, Jos70, Jos71d, Kab63, Kab64, KM71, KZ66, KR56, Kon53, Kon54, Lad55, Lec62, Lor71, Mee72b, Mei69, Mik63, Obe71, Obe72a, Pau64a, Pau67, Per69b, Per70, PP69, Ran70, RH71, RS67, RSW68, RW70, Sac58, Sch67b, Sea55, Sen68d, Sim67a, Sit52, Sob67, Spj72b, SO68, Tam69, Tho64, Wal58b, Wei54b]. Proceedings [Ney66]. Process [Alb62, AW64, Ath70, Ber63a, Ber64c, Ber72e, BS72, Bla58b, Blo57, BR64, Bos66, Bro65, Bur68a, Cla56, Com64, CL65, Dal68, Dal64b, DS53, Dha63b, Dha68, Enn70, Eri71b, Fos71, Gav59, Geb64a, Gem72, Gol71, Goo68, Hea61b, Hel70, Hen72, Hey70, Jof67a, KS47, KS68, KM67, Ken48, KW56b, KQ72, Kun67, Lec62, Lin70, Mil68a, tCM59, tCM67, NV60, Ott49, PS70, Par70, Par60, Par58, Pit63, Por68, Pyk59, PS68c, Qua68, Rao67, Rao68c, Rao70b, RS70, Rou69, Rub62b, Rub63a, RT59, SS69, Sen68e, She67, She71, Sid58c, Sid62b, Sim67a, Sle61, Smi68a, Sto68b, ST64, Str58, Swi65, Tak61c, Tay68, Ull44, Wal48a, Wei65a, Wei72a, Won35, Wyn69, Ylv65, Zim72, Sti66]. Processes [AK68b, Ath68, Ath69, AK71b, AK71a, Bel61, Ben62, Ber72a, Ber64a, Ber63b, Ber70e, Ber71d, Beu63, Bha71a, BW71, BM61, Bla55b, BE72, BKR71, Bou68, Bre57c, Bre58a, Bre68, Bre69, Bre70, Bro71a, Bro70d, Bro71f, Bro71g, Büh67a, Bur56, Cha72, Cho69a, Cra70, DBTL68, Der61, Der64b, Der66a, DS66, Der66b, DV67, Dha64, Dha69, Doo44, DF66a, Dud65, Dur71, DKW53b, DKW53a, DKW59, Dwa64, DM70b, Dym70, Eis69, Eis70, Eis72, Ell72, ET72, Fel68, Fel59, FK72, FR68a, Fis68, FK58, Fos53, FR68b, Fre62b, Fre62c, Fur72, Goo68, GR53, GHRM71, Ham57, HK65a, Har48, Hel65, Hey69a, Hol68, Hol72, Huf69, Igl64, IW71, Isa64a, Isa64b, Isa67, Jag71, Jai69, JK70, Jaq72, Joh57a, JR69, JR70, Joh72, Jon63, KS59, KT61, Kaw59]. Processes [KK60, Kei66, Kei69, Ken53, KS66b, KS66c, Kie57, KW70, Kim57, Kni70, Kni72, Koo60, Koo61, Koo64a, Koo64b, Koo66, Kur70, Kur71, Kus66, Kus69, Lam63, Lam65, LC65, Lea66, Lin72, Lip71, Loy65, Mac71, MR65, Mar70, Mar72a, MS72a, McN70, Mil68b, Mil72a, Mil72b, Mil71b, Mon72b, tCM60, tCM67, Neu71, Neu64, NR62, Nor70, O’B72, Par57a, Phi69a, Phi69b, Pie70, Pit63, Pit64, Pit65, Por67, PS67a, PS69a, Por70, PRR67, PRR69a, Pre72, Pyk58, Pyk60, Pyk61a, Pyk61b, PS64, PS66, QW71, QW72, RC72, Rao68b, Rei64, Rod71, Ros66a, Ros68a, Rou65, Sak64, Sak66, SY70b, Sch71a, Sch70a, Sen68a, Sen71c, Sen72c, Ser71, Sha69a, Sha70d, Shi71, Sho72a, SW69, Spi66, Sti63, Sto69b, Str59, Str61, Str66b, Tai66, Tay65, Wat54, Wei65b, Wei66, Wel71]. Processes [Wel72, Wic69, Wic71, Wol47b, Yac68, Yan71, Ylv68b, AD52, FR69, FR70, Gol67, Jam70, Ser68b, Sin70, Get66, Pyk63, Dwa67a, Get67a, Gne64, Mas71a, Pra66a, Smi65, Wel65, Zac70b]. Processes-Average [Der66a]. Processes/Sampling [Ylv68b]. producing [HS49b]. Product [Aro47, Cle71, Coh40, Cra40, Fel35, Fis69c, HH72c, Hsu44, Hun39, Mal68a, MS69a, Nag51, PRR69b, Sha59c, Sur68a, Tei67, Var55, Var60, WAC62, TW46]. Product-Moment [Nag51]. Production [Dod43, Ney41, WW45]. Products [Cup69b, Dwy38a, Dwy38c, DT64, FK60, Mor62, NT70, Tei56, Tei61c, Tra68, Tra69]. Professional [Dem65]. Professor [Dem40]. Progeny [Lot42]. Programming [Bla62, Bla64, Bla65, Bro65, DM68a, Har72, Mai66, Mai69, MV69, Str66a, Vei66, Vei69, Vot60]. Programs [Den71]. Progress [BCD^{+}45]. Projection [Cle71]. Prokhorov [FW69]. Proof [Ano36a, Ath70, Bir65, BBT58, Dub68b, Fri69, Gho71a, Gil71b, Jog70, Joh59, Joh61b, Kie53a, Kui59, Moh71, Moo68b, Por65, Ray57, Sch67a, Tho60, Wij67, Wol49a, Bir64]. Proofs [LR72a, Mil71b, Pan67]. Proper [BRN63, Cor60, HB71, Sch69f, Str71a]. Properties [Ala70a, And71a, AR50, BMP63, Bas67a, Bec42, Ber70e, Bha61e, Bha62, BM63, BC52, BW64b, BW63, CG66, DJ52, Dav69b, Den66, Dwy64, Eat67b, Efr65a, Fel72, GK64a, GM72, Gup69, Har65a, Her59, Hoa71, Hol68, Hub63a, Jen69, KM67, Kha65a, Kha65b, KW57, KP66b, KP70, Lah57a, Lin70, Luk39, MG69, Mei72, Mes67, Mor51, Noe50, Nor37, OR64, OR66, Ols40, Ore68a, Pre72, Pyk61a, QW72, Rie31, Ros62b, Sen70b, Ser70a, Ski60, Ski69, SA70, TC65, Tsu67, Twe57a, Twe57b, Vor72, Wal48a, Wal70a, Wal59b, Wal51, Weg56, Wei65a, Wol71c, Won68, Woo66, Ylv68b, Zim72, Jam54a]. Property [AH71, Ala70b, AG64c, AG65, Aus66, Bro70d, Bro71f, Bur68b, Cra31b, Cro58, Dar65c, Dav60, Dav61, Fel59, Gir58, God60, Jai69, JF64, KM59, Kri71b, Kud66, LK54, Men66, Mik62, Mil61a, Mot58, PB70, Sen68c, Ses66a, Ste58b, Ste60, Ste69b, Ste73, Ste51b, Swi65, Tho64, Wal70b, Wij60, Wij75, WD39, BF63b, Che51, Dun65, Tul65]. Prophecy [Whe35]. Proportion [Mus68b]. Proportional [BW57a, pH69, Sub66, Zac66b, Zac67]. Proportionality [Fed51]. Proportionate [Raj65]. Proportions [AJ51, Bak49, Smi47]. Proposed [Bow44, Gle64]. Proschan [McC65]. Prove [Le 70a]. Provided [Lin56]. Providing [Pat64a]. Prozesse [Pyk63]. Pseudo [Bañ68, Joh64b, Led71, Cam52]. Pseudo-Games [Bañ68]. Pseudo-Inverses [Joh64b]. Pseudo-Metric [Led71]. pseudo-random [Cam52]. PSI [Bha61d, Goo58a]. PSI-Squared [Bha61d]. Public [Man45b, Man46]. Publications [Ano51i, Ano51j, Ano51k, Ano51l, Ano52j, Ano52k, Ano52l, Ano52m, Ano52n, Ano53j, Ano53k, Ano53l, Ano53m, Ano54l, Ano54m, Ano54n, Ano54o, Ano55i, Ano55j, Ano55k, Ano55l, Ano56j, Ano56k, Ano56l, Ano56m, Ano57i, Ano57j, Ano57k, Ano57l, Ano58i, Ano58j, Ano58k, Ano58l, Ano59i, Ano59j, Ano59k, Ano59l, Ano60j, Ano60k, Ano60l, Ano60m, Ano61j, Ano61k, Ano61l, Ano61m, Ano62m, Ano63i, Ano63j, Ano63k, Ano64i, Ano64j, Ano65n, Ano65o, Ano65p, Ano65q, Ano65r, Ano65s, Ano65m, Ano66m, Ano66n, Ano66o, Ano66p, Ano66q, Ano66r, Ano67m, Ano67n, Ano67o, Ano67p, Ano67q, Ano67r, Ano68m, Ano68n, Ano68o, Ano68p, Ano68q, Ano68r, Ano69n, Ano69o, Ano69p, Ano69q, Ano69r, Ano69s, Ano70m, Ano70n, Ano70o, Ano70p, Ano70q, Ano70r]. Publications [Ano71m, Ano71n, Ano71o, Ano71p, Ano71q, Ano71r]. Punched [Car34b]. Pyke [Bru62, Kui59].

Quadratic [Bal67, Bha66, BY69, Box54a, Box54b, Cra43b, Gei57, GS55, Gra54, GM69, Gur53, Gur55a, Gur62, Ham56, HH72a, HW71, Kac45b, Kad68, Kar58a, Kaw50, Kha66, Koo42, KJB67a, KJB67b, LR72c, Luk52, Lut65, Mad40a, Mad40b, Mat49, Pac55, Rob48b, RP49, Rub62a, Rub63b, See71, SK61, SK63, Sha63, Sha68, Sha70a, Sid65, SC71b, Stu71a, Var66, Var68, Wal58b, GM57, Dol69]. Quadratique [Dol69]. Quadrature [Enl34, Mac71, O’T33b]. Quadrics [RC62]. Quadrivariate [Abr64]. Qualitative [Fis69b, Ham71, Luc68, Nii72, Vil64b]. Quality [Ano43j, GR52, Mos41, WW45, Wol43]. Quantifying [Gut46a]. Quantile [FR64, RS67, Sho72a, Zid71]. Quantiles [Bah66, FT66, Gho71a, Kie67, Noe48, RR65, SA66, Sen68a, Sob67, Zid69, Van67]. Quantitative [Fin71a, AT63]. Quantizers [Eli70]. Quantum [Sal67]. Quasi [Cad53, Chu57, Dal69b, Gov63, Har59b, Har60b, MS72a, PS72, Rid59a, Rid59b, Sto65c, dO66]. Quasi-Linearly [dO66]. Quasi-Martingales [PS72]. Quasi-mixture [MS72a]. Quasi-Ranges [Cad53, Chu57, Har59b, Har60b, Rid59a, Rid59b, Gov63]. Quasi-Stationary [Dal69b]. Questionnaires [Che62]. Queue [Cla57, Dar64b, Igl72, KK62, Kin62, Pak71, Tak62, Tak71b]. Queueing [Ben63, Bur68a, Dal68, Fin62, Hea61b, KW56b, Sha65b, Wis56]. Queues [Ben57, Bur64a, FN61, FP64, Hai61, Igl65, Joh65, Kei62, Ken53, Kin61a, Loy62, MS63, Mil60, Neu67, New60, Pra60a, Rei57, Rei63a, Sac60, Wel64b, Mor66, Mil64a, Rei64]. Queuing [Fos53, KK60, Sch70b, Tak61c]. Quickest [Bat67, IT68]. quite [Tuk58a]. Quotient [Cur41, Gor41a, GK50a, GH51, GP67, Hun39, Kot64, Kul36]. Quotients [Law68, EO72].

R [Der64a, Dud69b, Dym70, Geb68, Gle64, Gra67, Han66a, Nor70, Pra66a, Ylv68b]. R. [Aro41b, BD72, Dym70, Gle64, Gup72]. Radial [Gru44, Sch61a, Tal63, Wei54a]. Radii [Obe47]. Radioactive [Mal47]. Radon [She66, Cap64, HS49a, KZ71, Kai71, Var67]. Radon-Nikodým [She66]. Raikov [Dev59]. Raked [Cau72]. Raktoe [HF70]. Ralph [Mid66]. Ramanchandran [Cup69a]. Ramanujan [JS68]. Random [Abe69a, Abe69b, Ail70, AB66, AK71b, AK71a, AFLP57, AFPR59, Bal67, BM65, BK63, Bel70a, Bel70b, Ber62c, Bir48, BA49, BG46, BBT60, Blo70, Blo71, Blo50, BHP63, BR64, Bor67, Bri38, Bri72, BN45, Bro64, Bru69, Bru71, Bur72b, CT58, CR63, Cho66b, CT71, Coc46, Coh72, Cov66, CE67, Dal69b, Dar53b, Dar62, Dar65a, Dav49, DJ52, Dem60, Dem61, Dem68, Dem72, DJ69, Dud68, DG57b, DT57, Dwa62, Dwa67b, Dwa72, Eat70c, EO72, Eav67, Eav69, Eic63b, Eic66, Enn70, Enn73, ET72, EPW67, Far64c, Fel51, Fer70b, Fer71, Fis33a, Fre62c, FG60, Fri69, FK60, Gal72, Geb64a, Get67a, Ghu54, Ghu55, Gil71a, Gil71b, Gil59a, Gil62, Gir42a, Gir42b, Gla59, Gle69, GM69, GD55, Gui71, Gur55b, Gur59, Hal44, Ham56, Han59, HK65b, HW71]. Random [Har60a, Hey67, Hil69a, HS55, Hoe55a, Hög72, Hol69, Hor71, HK44, Imh60, Ios65, Ios69, JP72, Jiř72, Kab63, Kab64, Kac45a, Kat67, Kat55, KP57, Kem63, Ken72, Ken39, Kes70, Kes72, KW56b, Kim44, Kin63a, Kin61b, Kin61c, KP66b, Koo68, Kot64, Kot65, Kou70b, Kou72, Kue68, LaM72, Lan63a, Law68, Leh63d, Ley70, Lig70, Mad59, Mad40a, Mad53, Mal68b, MW47, Mar48, Mar71a, MS72a, Mas71a, Mas71b, Mas72, Mee70, Mee71, MD71, Mil61a, Mil61b, Mil67a, Min68, MS70b, Min72, Moo71, Mur72, New64, New68, Ney62, Nic58, Noa50, Nor66, Pic70, Pol72, Por71, PR70, PS68b, Rai72, Rob44c, Rob44b, Rob45, Rob47, Rob65b, Rod71, RP55, Ros70a, Ros66b, RC60b, RC60c, Rub56, Sak64, Sak66, Sam66, Sam69, San47, Sen71c, Sen72c, Ser71, Set65, Sha57]. Random [Sha58b, Sha60d, She65, She50, Sit52, Ski68, Smi68b, Smi68a, SW69, Sol53, Sto66a, Sto69a, Sto68c, Str65b, SC72, Swi65, Szá72, Tei65, Tho72, Tho57, Tod70, Tom71, Tsa56, Var66, Var68, Vot46, Wal44c, Wal45b, Was51, Wee61, Wei55a, Wei72b, Wic69, Wic70, Wij57, Wij59c, Woo70a, Wri72, Yac68, Yar72, van61a, vBE65, BB61, BM58b, Cam52, Eli72, Far66c, Far66d, Goo57b, Hor48, HS49b, Mar61a, Mar61b, Tuk49a, Tuk57c, Won69, Gne64, Nor70]. Random-Effects [Imh60]. Randomization [Cox56, DWW51, Dwa57a, EZ61, HC71, Rob72b, Wal49d, Zys63]. Randomized [Ahm61, EZ67, Far68b, Far69a, FP71, Hol67, Jog66, Kie58, Len69, McC39, Oga63, PF70, PF72, PS68a, Ros71c, Sam68b, Wal59c, WK56, Zac63, Zac64, Zac66a, Ban66a]. Randomly [Bil62, CRT65, CS39, Hin65, NR67, Sch65, Sud71a]. Randomness [Lev52, Man45c, Moo63b, Noe50, Cam52, SE43, WW43, You41]. Range [Bak46, Bru62, Cad54, Dal46, Fel51, Gum47, GK50b, Gum54, GS65, Har60c, HD54, Hoe46, Hol72, JP72, Kem59, Kie52b, Kni63, Mos52, Pil50, Rub60b, Rub61a, Sid58b, Spr53, Tho38, Wal47d, Wal49f]. Range-Midrange [Wal49f]. Ranges [Bur55, Bur67, Cad53, Chu57, Gum44, Har59b, Har60b, Lin50, Lin52, Rid59a, Rid59b, Tho36, Gov63]. Rank [Adi67b, And54, AB60, Bas67b, Beh71, Beh72, Ber70a, Ber72a, BS68, BJ70, Cap61, Dha69, Dok69b, DT71, DH69, Dwa53, Dwa56, Dwa57b, Dwa67b, Fra57a, Gas65, GW68, Gee70, GS71, Gho71b, Gut46a, Háj62, Háj68, HL62, HL63, Hoe68, Hol67, HP36, Jog66, JM72, Jur69, Klo65, Kou70b, Kou72, KvE72, Kra71, Leh53, MT65, Meh64, MS67, Mil67b, Mos63, Old38, Old49, PS69b, PS69c, RSS60, RP67, Sär65, Sav56, Sav57a, Sav59, Sav60, SSW66, SS66a, SS67, Sch69b, Sen67, SP67, Sen68b, SP68, Sen68c, Sen69, SP70, Sen70b, Sen71a, SG71, Sen72b, Set70b, SP69, Ste69a, Swi64, Tam63, Ter52, Tin45, Tsa55, Tsa57, Uza60, Vor72, Woo40, Woo70d, vE72]. Rank-Order [Beh71, Beh72, Gho71b, Háj62, SS66a, SS67, Set70b, BS68]. Rank-Sum [AB60]. Rank-type [Fra57a]. Ranking [Bec54, BS54, BH63, Eat67a, Eat67b, Fab62, GS57, Gup72, Hub63a, Mah67, Meh72, Per69b, RSW68, Sea55, SO68]. Rankings [Fri40, KS39, TR64]. Ranks [Dwa57b, Kou70c, Kou71, KS72d, PP69, RW70, Tam69, Tuk57c]. Rao [Gra67, Fen59, Kul54, OQ72, Gar59]. Rare [CC46]. Rate [BMP63, BM64a, BM64b, BP66b, BP67a, BP69, BD69, Dok67a, Dok68, Fre62c, HV65, HK65b, Har70b, Hey69b, Kem69, Koo60, MP65, MV69, Mur65a, Rag70a, Rao70a, Wag69, dHH72, Bic69c]. Rates [Bah67, Bha71b, Dav68d, Dav68c, Dav69c, Den71, Far72, HK64a, HK65a, Han67, Her69, Jaq72, JM63, JV71, JV72, Ken72, Koo66, RO72, Saw72, Wri72]. Ratio [And60, Bak46, Ban49a, Bar68, BC54, BS72, Bir42, Bir55, Bli46, BB69, Boy58, Bro71e, Che54, Con66b, Con67c, Dal40, DeG61, DL70a, DL70b, Dic71, DL72, Efr67, Fed68, GW59a, Gir64a, Gir64b, Gir65b, Gir58, Gor41b, GG39, Har42, HvN42, Hog56, Hot31, Hub65, Ifr66, Jai69, Kad68, Kap71, KR56, Kat61a, Kha65a, Kha65b, KW57, KP70, Kru54, Lad39, LD66, Lev71, Lin72, Lin50, Lin52, Lor72, Mat63, Mos62, Ney63, Pat64a, PL71, Pau41, Pfa67b, PR54, PW63, Raj53, Ray57, RP63, Rie39, RP58, Sam53, Sav57b, SS66a, SS67, Sch42a, SP70, Set70b, Sha71a, Str72d, Sug69, SF69, Tho35, Tho55a, Tho55b, Wal46b, WW48, Wij60, Wij63, Wij67, Wij68a, Wij71, Wij75, Wil38b, Wil41b, Wol66, vN41, vN42]. Rational [Dwy37a]. Ratios [Abr70, BGP69, Bir58, Blu66a, Dem69, Dix51, Gei57, Gur48, Gur53, Hog51, Hsu45c, PL70, PAAJ69, Pit65, Rie36, Sch70c, Yar71]. Rayleigh [Owe67b, BM63]. Re [Bri63b]. Re-Use [Bri63b]. Reaches [DK56, DK58]. Reaching [HL52]. Reacting [Lor71]. Real [Enn70, Enn73]. Reality [Kha65d]. Realization [Arb67]. Reasoning [Dem66]. Rebranching [SS69]. Received [Ano51i, Ano51j, Ano51k, Ano51l, Ano52k, Ano52l, Ano52m, Ano52n, Ano53j, Ano53k, Ano53l, Ano53m, Ano54l, Ano54m, Ano54n, Ano54o, Ano55i, Ano55j, Ano55k, Ano55l, Ano56j, Ano56k, Ano56l, Ano56m, Ano57i, Ano57j, Ano57k, Ano57l, Ano58i, Ano58j, Ano58k, Ano58l, Ano59i, Ano59j, Ano59k, Ano59l, Ano60j, Ano60k, Ano60l, Ano60m, Ano61j, Ano61k, Ano61l, Ano61m, Ano62m, Ano63i, Ano63j, Ano63k, Ano64i, Ano64j, Ano65n, Ano65o, Ano65p, Ano65q, Ano65r, Ano65s, Ano66m, Ano66n, Ano66o, Ano66p, Ano66q, Ano66r, Ano67m, Ano67n, Ano67o, Ano67p, Ano67q, Ano67r, Ano68m, Ano68n, Ano68o, Ano68p, Ano68q, Ano68r, Ano69n, Ano69o, Ano69p, Ano69q, Ano69r, Ano69s, Ano70m, Ano70n, Ano70o, Ano70p, Ano70q, Ano70r, Ano71m, Ano71n]. Received [Ano71o, Ano71p, Ano71q, Ano71r]. Reciprocal [Ber69b, BG68, Jam70, Rob65b, Ste45b]. Reciprocity [Cou44]. Recognition [Alb63]. Reconsideration [Lew35]. Reconstructing [Swi67]. Recovery [Har45, Sha71a]. Rectangle [Ste71a]. Rectangles [Jog70, Šid68, Šid71]. Rectangular [Bir54, Car46, Dun31, HG66, Kie52b, Pra58b]. Recurrence [Ahu70, Ban49a, DJ68, Dos69, Jos71a, KR66, MS70b, Min72, tCM59, Rio37, Rio65, Sri62]. Recurrent [Bel70a, Bel70b, Ber68, Bro71d, Buc65, Fis68, Isa64a, Lot48, Min68, Rei40, Rob61, Sch61b, Sin71, Sto66a, Tak61c, Wei72b]. Recurring [Ric48]. Recursive [Jon70, Sho66]. Reduce [And60]. Reduced [Joh67a, Kha54, Lau63]. Reduction [AGW71, Bro40b, Kag72, Vaj45]. Reference [Ano62e, Ano62f, Ano62g, DM68b, Gup72, Lot39a, Pol34, Wel39, Wil42]. Reflecting [Leh63d, Wee61]. Regarding [Jos69b, MR66, Rob33]. Regenerative [Kei66, Mil72b]. Region [She58, Vil64a, Wal58b]. Regions [BF72, Fed68, Fra53b, FG56, Gut57, Gut59, Gut70b, Hoe48c, Kou69, Kou71, QG68, Rub60a, Rub60b, Rub61a, Rub61b, Rub62a, Sau60, Sav57b, Sch62, Sch69g, Sco67, Sen71b, Wei57a, WD39, Che51, Tuk47a, Tuk48b]. Regression [Adi67a, Adi67b, Alb66, AG70, And51a, And62a, And80, And34, Ant62, Ber37, Bha62, Bha63, Bri38, Cac67, Cha50a, Cha51, Cha62b, Cle71, Con66b, CRWC72, Dan54a, DA59, Dwy37b, Dwy41a, Eis39a, Elf52, Far62, Fin46, Fra57c, Fra72b, Fra53c, GS71, Gle65, Gle66b, Gol63b, GM72, Gre54b, Gue58, Hal51, Har70c, Hil62, Hoe65b, Hoe66, Jae72, Jon70, Jur69, Jur71, Kee51, KB71, Ken39, KW52, KW59b, Kie61, Kou69, Kou70a, Kou71, LB63, Lin60, MM62, MM63, Man58, McG68, Min31, Mur71a, Mur71b, Obe72b, Que36, RT69, Rao70b, Ros56a, Rub58, SY66, SY68, SY70a, Sch72, Sen69, Sen72b, Sid60, Spj72a, Sri71, Sto59, Str61, SV69b, Tin45, Vil47, Wah71a, Wal47b, Wal58b, Wat67, Wic30, Wil67, WF57b, Won36, Ylv62, Ylv64, dlG54, Fab69a]. Regressions [Eic63a, Eze30, Mal70, Rad58]. Regular [Bha61e, Che67a, Chi56, DN70, Fre57, God60, Oga59, Pol70b, Rod66b, Ste70, Tay53]. Regularity [CR51, FG52, Get71]. Regularly [Jon62]. Regulation [Beu65]. Rejecting [Ste71b]. Rejection [Tho35]. Rejective [Háj64]. Related [Aro41b, AK68b, Bea37, BH67, BP58, Blu67b, Bro69, Cha50a, Cor60, Dar53b, Dar62, Dic67a, Dun61, DG57b, Dwa58, Fer70a, Gei56a, GO69, Goo68, Gup63a, Har60a, HPW64a, HPW64b, Imh68, Joh51, Kai71, McN70, Nel70, Pit65, Pyk58, Rob70a, Sam53, She50, SSF68, Tom72, Ylv68b, FT50]. Relating [Dav67, Han67, Maj53, Yac68]. Relation [Ahl70, Ahu70, Bak31, Dem64, Dre40, Eis69, Fré47b, GWA72, LF70, NS51, Ney51, NS52, Sco50, Ste69c, Vat46, Vil64a, Van63]. Relations [Col51, Con52a, DJ68, Dos69, Fre66c, Ghu54, Ghu55, Jos71a, KR66, Lah57b, Rao62, Rio37, Sri62, Vaj51, Var60, Vil69b, Vil72, vB31, Thi31j]. Relationship [BC57, Car68, HWG65, HWG95, Hil60, Hil61, Jam57, Kie64, OI65, Rob70b, Rob71b, Sin64, Vil61]. Relationships [Bla57b, Bur32, EP72, Goo58b, JS72, Joh66, MW43, Mer33]. Relative [Beh72, Bly58, CLA62, Coc46, GRS55, Hen31, Hen32b, KZ71, Lon44, Nor38, Pol34, Rag65a, Rob30, Sin67, Van63, Kat50]. relevance [Tuk57b]. Relevant [ES54]. Reliability [Bar65, CG66, DS69b, Eze30, Fle62a, Fle62b, Gum43b, Har70a, MD60, Mur65a, MS68b, SV69a, Mad63, McC65]. Remainder [Phi69b, Vaj45]. Remark [Ber51b, Ber70c, BF64b, Cha67, Der61, Erd50, Fre67c, Fre67b, Fre68, Gre47, Gre50a, Gre50b, Hub63b, Kac45b, KL56, LR72c, Por68, RR72a, RR72b, Rob54, Sei55a, Wol66, Zyg47, Bic69b, vN42]. Remarks [FV70, Gau57, Gau59, Gei45a, Get67b, LS08, Luk57, Mil72a, Pfa69b, RK58, Ros52b, Ros56b, SW72b, Wal70b, Wic30]. Renewal [BP64, Ben62, BY65, Bro40a, CR63, CS72, Dal64b, Den71, Fel41, Hey66, Hey68, JM63, Kei69, Ken71, Lam62, Lot48, Neu64, Por65, PS67a, Pyk58, Pyk61a, Pyk61b, PS64, PS66, Roo71, Sch70a, Sid62b, Smi68b, Ste69d, Sto65b, Sto66b, Sto72, Teu68, Der64a]. Renewing [Lot39a]. Rényi [Csö65a, RS72a, Sch64, AB61, Csö65b, RV72, Sen71b]. Reparametrization [Wat71a]. Repeated [KB71]. Repeating [Hin65]. Repetitive [Olm40, Van66b]. Replacement [BP64, HR62, KS70, KS72c, Lan72b, Lot39a, RK58, Ros72a, Ros72b, Sub66, Tay65]. Replicated [Koo67]. Replicates [BF64a, BF68a, BC57, Joh64a, PF70, PF72, Pat62, RF71]. Replication [BF66, Zac64]. Replications [KDS72]. Reply [Zoc37]. Report [And51b, And52, Ano36b, Ano40f, Ano40g, Ano43i, Ano46b, Ano46c, Ano46d, Ano48j, Ano48k, Ano48l, Ano50i, Ano56n, Ano58m, Arn48, Arn50, Arn51b, Arn51c, Bir49, BCD^{+}45, CDDK45, Coc47, Cra37, Cra38b, Cra39a, Cra43c, Cra44a, Cra47b, Doo51, Dwy44a, Dwy45b, Dwy45a, Dwy46c, Dwy46b, Dwy46a, Dwy47a, Dwy47b, Dwy47c, Dwy48c, Dwy48b, Dwy49c, Dwy52, Eis52, Fel48b, Fis50b, Fis51, Fis52b, Gua49, Han43a, Han43b, Hod49, Hod50, Kul46, Law55, Lit48, Lit49, Lit52, Mad44, Moo51b, Mos51, Ney50, Old41, Old42a, Old42b, Old42c, Old42d, Old43b, Old43c, Old43a, Old44, Rid40, Rid41, Rob50, She45, Van49, Wal49b, Wil47, Wil48, Wil49, Wil50, Cam52]. Representation [Boa49, BC59, Cha70a, Cla70, Cla71, Fer62b, FK72, Gho71a, Hol68, KS47, Kie67, Luc68, Rob66, RO72, Wij59a]. Representations [KJB67a, KJB67b, Pre69a]. Representative [Cra39b]. Representatives [Agr66b]. Representing [Ols68]. Reprint [Ano31c]. Reproduction [Kon66]. Required [GC64, Hag67, Wal46b]. Requiring [Cur46]. Research [Ano56o, Eis39a, Tuk47b]. Residual [Hen31, Law71, Mei72, Ros56a, Ste71b]. Residuals [AA50, Dwy41a, Hen32b, Jae72, Sid58a]. Resolution [DM68b, DM70a, Gul72, Pre69a, SC71a]. Resolutions [Ano40h]. Respect [Dyk70a, Kai71, Rob67, See72a, Ton69]. Response [BH57b, GGH59, MZ65]. Restraints [AS58]. Restricted [BP66a, Bar68, BG69a, Bru58, FP56, Jac69, Nüe66, SS66b, Sch68a]. Restriction [Coh57]. Restrictional [HF70, Maz67, Rak67]. Restrictions [And51a, And80, Sat42c, Ser70a]. Result [Ath70, Dod39, Fis60, Gän70, Hey70, Moh71, Rub63b, Shu69]. Results [Ath68, BK70, Cam52, Chu43a, Coc43, Dok69a, DT71, Eze30, Fis62, FP56, Geb68, GG58, Hal60, Han67, Hey69b, HP63a, HP63b, HP64, Kab63, Kab64, KP65, Kou72, Mar64, MW69, PR71a, PJ70, Pil56, Pra58a, Roy54, Roy56, Sch71a, Sin70, Sin71, Ste69c, Sto68c, Whi61b, Ala68]. Resume [Wel64b]. Return [BDF64, Gum41, She58]. Returns [Rio64]. Reverse [Bha70].

Reversed [And72a]. Review [Bar47b, Ben63, Beu65, Bha68, Blu67a, Bro71a, Che66, Coc44, Cox62, Cup69a, Dal64a, Dal69a, Dar65a, Dar65b, Dav63a, Der64a, Der66b, Doo68, Dud69b, Dup70, Dwa67a, Dym70, Fab69a, Fel47, Get66, Get67a, Gil64a, Gir48, Gna67, Gne64, Gra67, Gre64, GT64a, Gup72, Had63, Haj67, Han66a, Han66b, Har71a, Hed70, Hoe62, Hoe68, Hur48, Jai70, Joh69, Kal65, Kal69, Kem64, Ken64, Kie64, Kni66, Kri71a, LeP70, LeV66a, Lin66, Luk65, Mad63, Mar71b, Mas71a, McC65, Mid66, Mil64a, Mil74, MR64, Moo48, Moo64, Mor66, Mos65, Mos48b, Mul66, Ney66, Nor70, Ore64, Ore68b, Owe66, Owe67b, Owe67a, Pea64, Pra66a, Pru71, Pyk63, Rei64, Rio62b, RS72a, Ros66a, Rub66, Rub68, Sal67, Sav71, Sch64, Smi65, Spi66, Spi67, Str66b, Tak71b, Tuk47b, Wan68, Wat66]. Review [Wat71b, Wel64a, Wel65, Wen68, Wij68b, Wol67a, Wol67b, Ylv68b, Zac70b, Hub72]. Revision [Dwy49a]. Reward [Lip71]. Rewards [Har72, SSF68]. Rice [Owe67b]. Richard [Moo64, Cra53, Lin66, McC65, Ros66a]. Ridge [Mos51]. Riesz [PS72]. Rietz [Efr79, Kes72, Cra44b]. Right [Dol69, Kni72, Mon72a, Sto65c, Tha63, Tha65]. Right-Continuous [Dol69]. Right-Deterministic [Kni72]. Riordan [Ken64, Gou63]. Rising [AM67, Rio65]. Risk [Che67a, Geb64b, Geb66, Gil69, Hoe60, Lor67, Oat72, Pra61, Sch69a, Ste70, Wij70]. Risks [Ber63b]. Road [Wan68]. Robbins [Erd50, Blo57, DK72, Erd49, Hua72, Isb59, Kal54, Nád69, SP65, Ven67a, Wol52b]. Robert [Wol67b]. Robust [Bic65c, Dok67b, GR69, Gre66, Hub64, Hub65, Jae71a, Leh63c, PS67b, Ran70, RT69, SY72, Spj68b, Van70, Hub72]. Robustness [Ham71, Høy65, Høy68a, Mes68, Sen68c, Sen68d]. Role [Man62]. Room [Byl70, AJ58, Sha70c, SM68]. Root [Hec60, Kha64b, Nan48a, Nan48b, PL70, PL71, Pil67, Sug66, Sug67b, Sug67a, FT50]. Roots [And65a, And65c, Cha70a, Dav70, Dav72, Gir39, Gir42a, Gir42b, Gna67, Ham56, Jam60, Jam64, Kha67b, KP68a, KL56, LPC70, Mic59, Mij61, Mij64, Moo51a, Nan50, Pil64, PG68, PAAJ69, Sam41, SC71a, Sug70, van61a]. Rosen [Koo68]. Rosenblatt [Gne64, Fis60]. Rota [Sta65]. Rotatable [BC59, BD59, BB60, DN62, Dra60a, Dra60c, Dra60b, Dra61, DH68, GGH59, Her64, Her66, Her67b, Her67c, Her70]. Rotation [Eck55]. Roulette [Dub68b]. Routine [Ney41]. Row [BE70, HST53, Hub63a]. Rozanov [Mas71a]. Ruin [Pra61]. Rule [Ams63, BY69, Bro69, Car32b, GS67, SW72a, TBO71, TO72, RW64]. Ruled [BKS71, Str72b]. Rules [BG52, Boh68, Cho66a, CT66, DS66, Dun61, Gup62, Gup65, Gup70, Hal58, Hal59, Mat55, Ros71b, Sam64, Sam68b, Sav60, Sie67a, Sie67b, Sie68, SSF68, Sie69b, Tra68]. Run [DM68b, DM70a, FD65]. Runs [Kap45, KR45, Lev52, Moo40, Mos41, Sta72b, Wol43, Wol44b, LW44, Olm46, Wol44a]. Rutgers [Dwy45a].

’s [Day65, Ano65m, Coc44, Cox62, Hed70, Hoe62, Luk65, Owe67b, Owe67a, Rub66, Wij68b, Bur68a]. S. [Ano65m, Hed70]. S.P.R.T [Ray57]. Saaty [Ben63]. Sacks [Wah71a]. Saddle [Goo57c, Goo61b]. Saddle-point [Goo57c, Goo61b]. Saddlepoint [Dan54b, Goo61d]. Salaries [Rob37b]. Sale [Alb70]. Same [Dix40, Fer36, Mal68a, Mat43, Sav57b, SU72, WW40]. Sample [Abr67, AE69, Alt39, And60, And62c, AW64, Arn65, AFLP57, Aus66, BR60, Bak35b, Bak46, Bas67a, Bas67b, Bec54, BS54, BEM59, BMT66, BD66, Ber72b, Ber70e, Bha65, Bic64, Bic69a, BZ49, BH60b, BW57b, BKR61, Blu63, Blu66a, Blu66b, Blu67b, Bow46, Bra52, Bro47b, Bur63, Bur64b, But69, CLA62, Cha72, Cha50b, Che56, Chu55, CH55, CW58, CW59, Coh58, Con67b, Dal62, Dal46, Dav62, Dav63b, Dav66, Dem58, Dem66, Dud69a, Dur68, DKW56, Dwa56, Dwy37a, Dwy40a, Dyk70b, Eav67, Eic70, Far64a, Fer36, FP37, Fin66, FG60, Gas65, Gla59, God61b, Gra58, GC64, GG50, Gru50, Gum46, Har59b, Har60b, Hea56, Hem52, Hin65, Hoe52, Hoe53a, Hoe60, Hsu45b, Ifr66, Joh71, Jos68a, KP57, KW59a, Kie59, Kie67, Kin52, Klo63, Klo64, Koo71, Kru52]. Sample [Kus66, LLN60, Lay72, Lit72, MS68a, Mah67, Mar58, Mas51a, Mas51b, Mas52b, MT65, MP67, Mik63, Moh71, Moo54, Mos58, Mus68b, Owe53, PL70, PL71, Pau52b, Per69b, Pic67, Pic68, PR54, Pit63, PS67a, PS68b, Qua65, Qua66, Rag65b, Ram72, RSS60, Ray65, Rea72, RS70, Rob64b, Rot71, RG62, Sau63a, Sav56, Sav59, SSW66, Sch69b, Sch69c, Sch69d, Sco71, Seb64a, SP67, Sen68a, SP68, Sen70b, SG71, Sen72c, Set64, Set70a, Sha69a, Sha70d, She50, Sid58c, Sid62a, Sim67b, Spi66, SPZ56, Ste69a, Ste45a, Ste42, Str71b, Str72c, Sun54, Swi67, Tam66, Tei55b, Tho35, TW34, Tsa54a, Uza60, Wal42c, Wal47d, Wal47e, Wal50a, Wal58c, Weg56, Wei52, Wei55c, Wei57b, Wei57c, Wei58b, Wei59, Wei60, Wei61, Wij68a, Wil38b, Wil41a, Wil46a, Woo67, WV67, Wor60, Ylv68a, Zac67]. Sample [Zei50, Zim72, dHH72, uD54, Van67, BT46, Cha63c, Con65a, Con65b, Dav58, DT71, JM72, Mos48a, MT50, Olm46, Pur64, PS68c, ST44, Ser68b, Suk57, Tam71, ZD59]. Sampled [DS40, Nad60]. Samples [Bah66, Bak31, Bak32, Bak35a, Bak40, Bak49, Bha61a, BM58a, Bow44, Bri63b, Bur63, Cad53, Cad54, Car32a, Coc46, Coh50, CMPT60, Cra31a, Cra32, Cra33a, Cra40, Dix40, Dix60, Ell62, Eps49a, Eps51, Fel32, Fel35, Fin46, Fis69c, Geb66, Gov62, Gut70b, Hal52, Hal61, HM67, Hoe38, Hsu40, Hsu41, Jam64, Jon48, Kac48, Koo67, Kul35c, Lad39, Len39, Lie53b, Lin50, Lin52, Mad48b, Mat43, Mat55, McM49, Moo39, Moo41, O’T34, Pil50, Rao54, Rei63b, Rid31, Rid59a, Rid59b, Rie39, Rub60b, Rub61a, Sal66, SG56, SG58, SG69, Sär65, Sek51, Spr61, Sta38, Tan70, Tei55b, Tei56, Tei61c, Tsa56, WW40, Wal46d, Wat54, Wat59, Wei58c, Wil32a, Wil38c, Wil41b, Goo52, HMTW47, Tam71, Tuk58a]. Sampling [Agg59, Agg66, Alb54, And42a, Ano30, Ano31a, ABE^{+}55, Bae45, Bak40, Bar43, Beh61, BH60a, BR64, BN61, Bro33, Bur55, Bur67, Car30, CR65a, Cra31c, Cra39b, Cur46, DS69a, DLS57, DJL59, Die42, Dod43, Dwy38a, Dwy38c, Eck55, Eva63, Eze30, Gau57, Geb64a, Gho63b, Gho66, Gir39, GMS46, GJ65, God69, GZ66, Goo53a, Goo61e, Gru49, GJ59, Gut57, Gut59, Háj64, HH43, HH49, Han68, HR62, Has42, HR36, Hil69a, Hoe43a, Hoe69, Ios65, Ios69, Jon52, Jos65a, Jos65b, Jos66a, Jos68b, Jos71f, Kim49, KS70, KS72c, Koz56, KF40, Kup58a, Kup58b, Lan72b, LS55, Lin50, Lin52, MM44, Mad49, Mad53, Mah63, Mol46, Moo43, Mor72, Nad60, Ney41, Old40, Pat64a, Pat64b, PL46, Pla48, Que49c, RK58, Raj64, Raj65, Rei63b, Rob44c, Ros72a, Ros72b, Sch67b]. Sampling [Sch69g, Sea58, See53, Sen55, Sen72a, She36, Sil48, Sin67, SW72a, Spr53, Sub66, WW45, Wel64a, Whi65, Wil32b, Ylv68b, vdW60, Roy70, Tuk56a, Dal64a, Dal69a]. Samplings [SF39]. Samuel [Had63, And65b, Cox62, Hoe62, Mar71b]. Santa [Moo51b]. Sarhan [MR64]. Satisfying [Ghu54, Ghu55]. Saturated [DM67, PF72, DM68b, Gul72, PF70, RF70a, RF70b, RF71]. Savage [Moo64, Bas68, HS70c, PS68a, PS68b, PS68c, PS69d, Sud72]. Saved [HC71]. Scale [Bas64, BW67a, BM59a, Bro68, Dok69a, Fer62a, Fer63, GS66, Klo62, Man69b, Pfa69a, PB71, Rag65b, SG56, SG58, SG69, Tat59, Wei57d, WGL66, Woi72]. Scales [Coc43]. Scaling [DR72, Dod37, LR53]. Scarf [Had63]. Scatter [van65]. Scheduling [Ehr70]. Scheffe [Imh62]. Scheme [Con58, Hof60, Joh70a, See53, Shr59a, Shr59b, Bat55]. Schemes [Aig71, Ano60i, AK68b, BM59b, CC47, Joh67b, Joh70b, Mes65, Pat64a, Rag60b, Sin67, SA71]. Scholarship [Rob37b]. Schumann [Bro59]. Schuur [LeV66a]. Schwarz [Mor53]. Science [Had63]. Sciences [Str66b]. Scientific [Tuk47b]. Score [Bha67b, Jog66]. Scores [HB71, Klo63, PS69b, Rag65a]. Scoring [Che62]. Scott [Sch60a]. Search [Che67b, Hel70, Swe70]. Searched [Swe70]. Seasonal [Car32b, Rob30]. Seasonals [Dep30, Pal30]. Seattle [Bir49]. Seckler [Gre64]. Second [Bak35a, Bak35b, BD59, BB60, Bri38, CT66, Day64, Day65, Dra60a, DH68, GG51, Han43b, Her67c, KS62, Kha62b, Kha73, Lah56b, Leh44, Moo51a, PG68, Sch69g, Str61, Teu68, TBO71, vS50]. Second-Degree [Lah56b]. Second-Moment [Moo51a]. Second-Order [Sch69g]. Secretary [Ano46d, Dwy45b, Dwy47c, Dwy48b, Dwy49c, Fis50b, Fis51, Fis52b, Old42a, Old43b, Old44]. Secretary-Treasurer [Ano46d, Dwy45b, Dwy47c, Dwy48b, Dwy49c, Fis50b, Fis51, Fis52b, Old42a, Old43b, Old44]. Sections [Koe70]. Secured [Eze30]. Security [Car34a]. Seek [MP71]. Segment [Eic70]. Selected [Bil62]. Selecting [AT72, BEM59, GS58a, Kar58b, Mos58, Pau64b, Pau67, RS67, RSS68, Sob67, SW72a, Stu67]. Selection [BG69a, BH57a, Coh57, Des70b, Dud69a, Fra52b, Gra68, GS57, GP72, HM68, Hoe71, Hot40a, Hub63b, Leh61, Mah67, Ram72, Ran70, RH71, RW70, Sen55, TD63, Tuk49b]. Selections [BC50]. Self [Lot39a]. Self-Renewing [Lot39a]. Semi [Ber63b, Cra31b, Cra40, Der61, Kur71, Lip71, Pil50, Pil71, Ser71, Sto68b, Sto69b, Yac68, Yua33]. Semi-Invariants [Cra31b, Cra40]. Semi-Logarithmic [Yua33]. Semi-Markov [Ber63b, Der61, Kur71, Lip71, Ser71, Sto68b, Sto69b, Yac68]. Semi-Range [Pil50]. Semigroup [Fel59, Fre67a, Fre67e]. Semimartingale [JK66]. Semimartingales [Dub66, Dub72, KS58]. Seminvariants [Dre40]. Seneta [Hey70]. sense [Beu63]. Sensitive [Vei69]. Sensitivities [SB57a]. Sensitivity [Cau72, EC44]. Separable [Ber51b, Ell72]. Separated [Dod30, Mil67b]. September [Ano53e, Ano54g]. Sequence [BM65, Ber62b, CS70, Dvo49, Far67, Fis33b, Fis42, IT68, Koz51, She65, SC69, SC72, Woo70b, Eli72, Jag63, SE43]. Sequences [AB61, AM67, AH71, Ber62c, Ber64b, Bha61d, BG46, BK60, Bur68b, DK68a, Eic63b, Fox70, Gil69, Gui71, HY64, Kim68, Kin63a, Koz47, Kur69, Pic67, Sac70, Sen71c, BB61, Goo57b, You41]. Sequential [Alb47, Alb48, Alb54, Alb56, Alb61, And60, AF62, Bah58, Ber69c, Ber71c, Bha64, BS72, BY68, BY69, Bir58, Bla47, BG47, Bla57b, Boh68, BJK56, BC64, BW63, Cha62b, Che59, CR65a, Che65a, Che65b, Che67b, CR65b, DK56, DK58, DeG59, DeG60, DeG61, DeG62, Der64b, DS66, DKW53b, DKW53a, DKW59, Ehr70, ES55, Fre67b, Fre68, Gee70, Gil68, Gir46a, Gir46b, Gir58, Gle65, Gle66b, Goo53a, Gra68, Hag67, HWG65, HWG95, Har47, Hoe53a, Hoe60, Hoe55b, Hoe71, Hol71b, Hou66, Ifr65b, Ifr65a, Ifr66, JB61, Joh53, Joh54, Jon52, Kie52b, Kie57, KW57, KS63, Kni65c, LB63, LS50, Lor67, Lor72, Loy69, Mat63, McC72, Mei69, Nád69, Pau47, Pau62, Pau63, Pau64a, Pau64b, Pau67, Pau69, Pau70, Per69b, Pha63, Ray65, RSS67, RSS68]. Sequential [Rob63, RS72b, Sam63a, Sam63c, Sam64, Sam65b, Sam68a, Sau60, Sau63b, Sau63a, Sav47, SS66a, SS67, Sch62, Sch68b, Sch71b, SG71, Set70b, Sim67a, Sim67b, Sim67c, SW49, Sob53, SO68, Sta66a, Sta66b, SW72b, SW47, Swe70, Tai66, Tay65, Tho64, Tsa54b, Van66c, Vog60b, Wal45a, Wal46a, Wal46b, WW48, WW50, Wei54b, Wei56d, Wij60, Wij63, Wij67, Wij68a, Wij71, Wij75, WK69, Wol46b, Wol47a, Wol47b, Wol66, Won68, WK62, Wyn70, Yar71, Zac66b, Zac68, Sac65, Gir48, Gup72, Ray57, Tuk47b]. Sequentially [Fra51b, Her48]. Serial [And42b, AA50, Ber72a, Cha62a, Dix44, Gho54, Gum43a, KS62, Kem70, Koo42, Lei47, Mad45, Oga51, Pak71, Pre69b, Que49b, Rub45, Sid58b, WD51, Whi57a, Whi57b, Whi58a, Whi58b, Whi59, Goo57b, WW43]. Series [And71a, And71b, And72a, AA50, Aro37, Bha72, BR69, Boa49, BC57, Cam37b, Cle71, Der39, DG57b, Fra63, Gei57, Gil71a, GG39, GR53, Han66a, HH72a, Isa69a, Joh65, KJB67a, KJB67b, Kul48, Lig71, LZ57, Lot48, MW69, Moo63b, Nel70, Par57c, Par57b, Par61, RP55, Ros56a, Sac60, Sch67e, Sny72, Spr59, Tin39, Tre65, Wah68, Wat69, Wil30, WB64, Wol48, Yal68, Yal71, Ylv62, Ylv64, vS49, vS50, Kul47, Que49a, Sam43, Wol67a]. Server [Dal68, KMP59, KK62, Loy62, MR57, Tak61c, Tak62, Tak71b]. Service [Fin62, Kei62, Ken64, Sha65b, Tak61c, Tak62, Wis56]. Service-Time [Wis56]. Servicing [Hai61]. Set [Bla68, BN45, Bur62a, DM68b, DM70a, Fel35, Gab62, HF70, Hen34, Kim44, Law63, MS69a, Maz67, Nev72, Rob44b, Rob45, Rob47, RS72b, San47, See72a, Set50, Sol53, Swi65, Ton69, Vot46, Wal50b, Wei72b, Yar72, Fra52a, Ste51a, Wal53]. Sets [AB61, Ail70, AG64b, BG69b, BLS64, Buc65, Con63a, Con67c, Dod30, Fis69b, Fis72b, Fis69c, HW58, HS61, Hor71, Jos67b, Jos69a, Jos71c, KPS59, KS69, Man43, Mud66, Mui72b, PR71a, Pit57, Rio62a, Sch70c, She55, SB57b, Teh63, Wei55a]. Setting [Sud69a, Wal42c, Wal43a, Wal46b, Wil41a]. Seven [RW64]. Several [Bha61a, Bro45, Bro39, Cha58, Cha65, CMPT60, Con65b, Cup69b, Dix54, Dud65, Dwa58, Hoa67, Kri70, Kru52, Meh64, Pau52a, Pau62, Pau67, Rob63, Rob64b, Sen69, Sen72b, SF39, Tam66, Tan70, WB41, Wel47a]. Several-Sample [Tam66]. Shannon [Ahl70, Ahl71, BBT58]. Shape [Mal68a, Win47]. Shaped [Tor52]. Shapes [Sch62, Sch68b]. Sharp [BF72, DL70a, DL70b, DL72, Dub72, MR67]. Sharper [DF65]. Sheppard [Ano36a, Cra36c, Cra41a, Kul35a, Lew35]. Shift [Arn65, Bha67a, BJ68b, Bic64, Fin66, Høy65, KS69, Leh63b, Ram66a]. Shift-Invariant [KS69]. Shifts [ASS72]. Short [Fri69, Hor32, Jen32, LR72a, Moh71]. Shorter [Pra63]. Shortest [Blu67b, Wal42a, Wei61, Wil38c]. Shreider [Mul66]. Shrinkage [Whe31, Whe35]. sic [Sac60]. Sidak [Hoe68]. Sided [BT51, Bir53b, BF72, Cha58, Cha65, Eat70a, Ell64, Far64a, Kem63, Mar60, MO60b, Noé72, Per69a, Per71, Pra58b, Sie67a, Sie69b, Wel64a, Whi61b, All53, Cho66a, Con67b, Mik62, NR71, Ram66b, Sie68]. Sidney [Zac70b]. Siegel [Sal67]. Sievers [Pla71]. Sign [Cha66, Dix53, Har47, Hil60, Hil61, Hod55a, JK62, Klo59, Klo64, Lig70, Mac41, Ols67a, Wal46a, Wal46c, Wal51]. Signal [New68]. Signed [Klo65, KvE72, Sen68c, vE72]. Signed-Rank [KvE72]. Significance [Aro43, Ban44, Bra52, Coc40b, Dem58, FGW46, FH55, Fri40, Gna67, Har42, Hoe37, HP36, Kee51, KB71, Kim51, Leh58, Mau40, MT50, NV68, Old49, Pei43, Pei49, Sta38, Swi64, Tho38, Tin39, VA43, Vil47, Wal46e, Wal47c, Wal49f, Wal49g, Wal51, ZJ59]. Significant [Pin61, Tuc41]. Signs [Man45c]. Silver [Cra60]. Similar [Ahm61, Chu56, Dod41b, Hoe48c, Kin61a, tCM60, SB57a, Wij58, HKcY68]. Similarity [da 45]. Simple [Alt37, Ath70, BH67, BG68, BHP63, BN61, Cla57, CS39, Day66a, Day66b, Dem61, DH69, Dwa67b, For37, Gil71b, Goo53b, Háj68, Hen32b, Hen35a, Hil69a, Hoe66, Isa51, Jog70, KM67, KF40, MZ66, Mil71b, MR69b, Por65, Roy33, Tsa54b, Var67, Wei53b, Zys67, AT63, Tuk56a]. Simpler [Dub68b]. Simplest [GKS63, GK64b]. Simplex [BB60, DK68b]. Simplex-Sum [BB60]. Simplification [Ifr66]. Simplified [Dix60, SGO63]. Simplify [HF38]. Simulated [Con63a]. Simultaneous [Bar39, Bat30, Ber45, Cra32, Dwa55a, Dwy37b, Gab69, Goo64b, Goo76, Hea56, Hot43a, HA72, Kha67a, Kri68, Ram56a, Rei49, RB53, Roy54, Roy56, Sen71b, Try58, Tuc41, Wil67, Zei50, vS51]. Sine [Ber62a, Hor71, Imh68]. Single [AR49, AR50, Bec54, BS54, BEM59, Coh58, Cur46, Dal68, Dha68, FR64, Gru49, Kar58b, KMP59, KK62, Loy62, MR57, Mur71a, Ste59b, Tak61c, Tak62, Tak71b, TW46, Tuk57d, Tuk57e]. Single-Sample [Bec54, BS54, BEM59]. Single-Server [Loy62, Tak62]. Singly [Coh50, SG56, SG58, SG69]. Singular [BF68a, Isa64a, Kap66, Rag64, Tuc64]. Singularity [Ban66b, Ban69, Fre66a, Fre66b]. Situation [Kra71, Ram66a]. Six [Har44, Har45]. Six-By-Six [Har44, Har45]. Size [Alb66, And60, BZ49, CJ56, DR67, Dud69a, Far64a, Fer36, Gle65, Gle66b, Goo53a, Gra58, GC64, Hoe60, Ifr66, Mos58, Pla48, Raj65, Ram72, Ray65, Rob64b, Sam68a, San72, SG58, SG69, Sau63a, Sub66, Tei56, Tei61c, Wal46d, Wei52, Wij68a, Tuk49a]. Sizes [Koo67, KDS72, Per69b, PS68b, Rag62, Sen72c, Wil41a, ST44].

Skewness [Dwy41a, Gar32, HS32, Maj62a, Wil44]. Skip [Kei63]. Skip-free [Kei63]. Skorohod [Dub68a]. Skorokhod [Get67a, RO72]. Slippage [HK68, KT60a, Pau52b, Sri65a, Tru53, Wal46c, Woi72, HKcY68, Mos48a, MT50]. Slope [Bli46]. Slope-Ratio [Bli46]. Slopes [Sie69c]. Slowly [Büh67a]. Small [Arn65, BH60b, Bro47b, Bur63, Bur64b, Cam46, Car32a, Cra32, Cra33a, Cur46, Den71, Dod30, Ekm59a, Geb66, Hoe38, Jaq72, Jon48, Klo63, Klo64, Mar58, MV69, Old38, Rid31, Rie39, Sha60d, Tei55b, TW34, Wal50b, Wal53, HMTW47, Che65a]. Small-Sample [Bro47b, Bur64b]. Smallest [Ala70b, How49, How50, Kha64b, Mic59, Sug67a, Wal50a, Wei59]. Smirnov [Fel50, Owe66, Abr67, And67, Bic69a, BH60b, Con65b, Con67a, Con67b, Csö65c, Dar55, Dar57, Dav58, Don52, Doo49, Egy59, Fel48a, Fis60, KM71, Kem59, Kie59, Lit72, MW69, Moh71, NV68, Noé72, Qua65, Ram66b, Ros62b, Sch58, Sta72a, Ste69a, Ste69c, VN67, Whi61b, Yu71]. Smith [Ney66, Dub68b, Hey68]. Smooth [Fab69b, Tuk58c]. Smoothed [Dic68a]. Smoothest [Sar49]. Smoothing [Dic69, KW70]. Snedecor [Mos48b]. Snell [Ore68b]. Snowball [Goo61e]. Snyder [Man45a]. Sobel [Gup72, Hal59]. Sobolev [Kus69]. Social [Dod36, Kat52]. Sociometric [KP57]. Solution [Ber42, Bha64, DV67, Dwy44b, HP49, Hoe66, Kin48b, Mol46, Nel66, Pau52b, Tuc41, Wal49e, Wil36]. Solutions [Bri72, Cog67, HR55a, McC47, Sac63, Sam63a, Sca59, Sch43b, She69, WW50, Wei56b]. Solving [Ber45, Hoe41, Hor32, Rei49]. Some [AK61, Agg55, Agr66b, Ala68, AC65, AK68a, AG44, AG64a, AG64b, AG64c, AG65, And65c, And54, Ano35c, Ath68, Bah60a, Ban49a, Ban70, BC54, BM65, Bel64b, BD65a, BD65b, BS68, Bha61b, Bha61c, Bha61e, Bha62, BJ68a, Bha70, Bic64, Bic65a, Bic65b, Bic65c, BW71, Bir55, BP58, BM61, BG47, Blo51, Boh68, BH65, BC55, Box54a, Box54b, BT46, Bue59, Cam42, CG66, Cau68, Cha61, CLA62, Cha50b, Cho66a, Cho66b, Chu57, Com64, Con52a, CC54, Con63b, Cox58, Csö65a, Csö65b, Csö65c, Cur46, Dal68, Dav69b, DeG60, Der39, DN70, Dod41b, Dok67a, Dok67b, Dok68, Dra61, Dri52, Dwa55b, DT64, Eat67b, Eps48, ET53, ES54, EP72, FV70, Fel72, For37, Fre67d, Fre57, Fre61, Fre66d, FP56, Gab69, Gau57, Gau59, Geb64b, Geb66, Geb68]. Some [Get67b, GO69, Gil71b, God49b, GZ66, Goo58b, Goo59, Goo68, GS55, Gre65b, Gru36, GAM64, Gup69, Gut59, GT64b, Háj61, Hal59, Hal60, Han67, Hec60, HSF72, Her48, Hey66, HY64, HL50, HL56, Hol68, Hol71b, Hoo56a, Hot40a, Hot43b, Hot44b, Hsu44, Hsu45a, Imh68, Imh71, Jae71b, Jai69, Jam54a, Jen70a, Jen70b, Joh53, Kab63, Kab64, Kad66, Kaw59, Kha65a, Kha65b, KP65, Kha67b, KW57, Kim47, Kim57, Kot64, Kou70c, Kou72, Lah57b, Law65, LS49, Leh50, Leh57b, Leh61, Leh66b, Lin70, Luk39, MR65, Mah67, Mar64, Mee72b, MP67, MG69, MW69, Mes68, Mes67, Mic59, Mil72a, Mil68d, Mol46, MH69, MH70, Mos46, MR67, Mui70a, Mul56, MG70, NR71, Nor66, Nor38, OS69, PR71a, PB70, Pau42b, Pea62, Pfa64, Pfa69b, Pil55, Pil56, PJ71, Por64b]. Some [Pra58a, Pre69a, Pre72, Put55, Rag59, Rag60c, RK58, Ram66a, Ran70, Rao72, Rob35, RR64b, RR66, Ros56b, Ros62b, Roy54, Roy56, RP58, RG59a, RG59b, Rub62b, Sam53, Sau63b, Saw72, Sch69b, Sch70d, Sei63, Sen68d, SP70, Sen70b, She69, Sie67a, Sie67b, Sin70, Sin71, Sin67, Ski60, SC63, Ski69, Spj68a, Spj72b, Sri62, Sri65c, SO68, SA70, Ste51b, Sto68c, SN68, Sug71, Tam69, Ter52, Tho64, Tod70, Tom71, Tom72, Tra68, Tra69, Tro71, Try58, Tsa68, Tuk57b, VA43, Wah68, Wal41c, Wal45b, Wal46b, Wal46d, Wal46e, Wal49f, Wal49g, Wal50a, Wal50b, Wal51, Wal53, Wal70b, Wat62, Wat63, Weg56, Wei58a, Whi61b, Wij57, Wij59c, WK56, Wor59, Yan71, Zim72, Zoc35b, Zys63, van61b, vS49, Cam52, Che51, Gho64a, Goo53b, Leh57a, Leh64, Mad71]. some [Maj53, Mil71b, Rag66, Sen71a, Wol43, Kar58b]. Somewhere [SS66b, Sch68a]. Source [Mad48a]. Space [Bro65, CC47, DZ67, Fer71, Gil62, Kak71, KP57, Led71, Lip71, Nev72, Ols67b, Sal67, Sim71b, Whi70]. Spaces [Ber51b, Cor60, Den70, Fis71, Fur72, Hed70, Kat61b, See70a, See70b, SZ71, Sto69b, Tay72b, Tho57, Var58, VR58, Dud69b]. Spacing [Ant62, DA59, Gue58, HL64, She50, dlG54]. Spacings [BD69, Blu66a, Blu66b, Blu67b, Sho72a, Wei57b, Wei57c, Wei58b, Wei58c, Wei59, Wei61]. Spare [MD60]. Spearman [Whe35]. Special [Ano43j, Ber67a, DM68b, Gup72, Lot39a, Sch61b, Suk72, Tre65, Wil42, Thi31o]. Specific [Dra61, Ter52]. Specification [WF57b]. Specifications [Byr35a]. Specificity [RG62]. Specified [BBH56, Fer36, Gra58, HR55a, HV65, Obe47]. Specifying [FM58, Isa51]. Spectral [Ber72a, GR53, Jon62, Lig71, Mac71, Par57b, Pic70, Rao68c, Rod66a, Sch65, Smi64, WV67, Wol67a]. Spectrum [Mur61a, Mur63, Nea70, Nea71, Par60, Par57c]. Speed [Dud69c, Fab67, Gän70]. Sphere [Jon63, Yos49, Mar72b]. Spherical [Rub60a, Rub60b, Rub61a, Rub61b, Rub62a]. Sphericity [Con67d, Gle66a, Gle68, Mau40, PN71, Sug72]. Spitzer [Dar65a]. Spline [Mur71a, Mur71b, SV69b]. Splines [KW70]. Splitting [Dha68, Koo71]. Spread [McM49]. SPRT [BS68, Hoe70]. SPRTS [Ber70d]. Square [Ala70b, Aro43, Dia63, Fel32, Gei56a, Gil64b, Gor72, Gum43b, Har42, HvN42, Har69a, HS70a, Hoe38, Isa72b, Joh70b, Kim40, KP66b, KA65, MW42, McN62, Mes67, Mit58, Moo71, MG70, PL70, PL71, Pre66, Roy33, Sat42c, Sha70c, Twe67, Wal59a, Wal60, WAC62, Wil41b, FT50, vN41, vNKBH41, vN42]. Squared [Bak31, Bha61d, Goo61d, Goo58a, Jof67b, Zid69, Zid70]. Squares [And62b, AJ58, Beu65, Bha62, Ble40, Byl70, Cho68, Col32, Dav33, DS40, Dow54, Dwy44b, Dwy49b, Eic63a, Fre66d, GW70, Gol63b, HB65, Har71b, HS70a, HF71, Hen31, Hen32b, Hen34, Jen69, Joh40b, Jor32, Jor59, Kru68, Ksh59b, Len69, MM62, MM63, Man43, MR69b, Mor62, Old38, Old49, Pau50, Pie70, Sad51, Sid58a, Vaj45, Vil66, Vil69b, War60, Wat67, Wis37, Won35, Zac64, Zys67, CT56, Hou66, Man42, SM68]. St [Ano36b]. Stability [Dvo49, For37, HH72c, Pol34, vB31]. Stable [Bel70a, Bel70b, BS70, Bre68, Bre69, Bre70, Fri72, Hen72, Kim69, Lah56a, Mas70a, Mas70b, Mij72, Pil71, Por67, Por68, Por70, Rao68b, Sch70d, Tuc68, Wol72]. Stage [BH60a, Bla64, DS69a, Ghu58, GZ66, Ric60, Ski60, Tho64, Yen64, Agg66, See53]. Staircase [GP58]. Standard [Bak32, Bak35a, Bak35b, Bak46, Cra32, Dix57, Dod36, Dun31, Dwy40a, Far69b, GG39, Gru44, GS58a, Har59b, Har60b, Hen34, Jen32, KM71, Kni63, Kun67, Lom52, Min31, Nor40, Pau62, PR54, Sar54, Sar55a, Sar55b, SC63, Sob53, Spr53, Tho35, Wal58c, Won36]. Standardized [Sal30a]. Stands [Has42]. Stanley [And65b]. Stanton [Byl70]. Starshaped [Dok69b]. Start [Mad53]. State [Alb62, BP66c, BBT58, Car67, Der61, Der66a, Dha68, Fos71, HH72b, HY64, Kak71, Lip71, Pet69, Ros68a, SY70b, Sto69b, Tsa68]. State-calculable [Car67]. State-Dependent [Fos71]. States [Alb61, Bla55b, Bla58b, Bre58b, Dha68, Dha69, FR68b, Pyk61b, Wat61, FR69, FR70]. Stationary [And72a, Ber62b, Ber64b, Ber70e, Ber71d, Ber72e, Bha72, Bla55b, BE72, Bre58a, Bur68a, Cap64, Cho69a, CL65, Dal69b, Dha63b, Dha64, Dha68, DM70b, Eis69, Eis72, Eri72, Gem72, GR53, HY64, Isa64a, Isa64b, Isa67, KS47, KW72, KM67, Kaw59, Koo64a, Koo64b, LC65, Loy62, Loy65, Mon72b, Moo63b, tCM59, tCM60, NR62, OP69, Pak71, Par60, Par57c, Par57b, PS66, Qua68, Rao67, Ros56a, Ros71c, Sid58c, Sny72, Tai66, Tre65, Wah68, Wat54, Wol48, Ylv65, DBTL68, Jam70, Mas71a, Ylv68b]. Statistic [Ala72, Bas67a, Bas68, BK57a, BP58, Cac61, Cac67, Chu46a, Con67b, Dwa59, Egy59, Fed68, GWA72, GH68, GS57, GS58b, Gup71, Har51, Ito56, Ito60, Jur69, Lau63, MM69, MG69, MO70, MB66, Moo71, MG70, Oga63, Pot63, RK57, Ros52a, RW72, Sch69d, Sen68c, Ser68b, Sho66, Sio71, Spr61, Str72d, Sug67a, WB41, Wal46d, Whi51, dW70, dW71, AB67, Het68, Lew61, She57, Zar65, ZD59]. Statistical [Agr66b, AN53, AB66, AG57, Ano43j, Ano56o, Bah54, BL55, BS56, BP67b, BP66c, BPSW70, Ber71b, Ber71a, Bil61a, Bil61c, Bir61a, BR67, Bly51, Bol34, Bri63a, Bro45, Bro40b, Bue59, Bur32, CG66, Cog67, Cou44, Cox58, Dem65, Dre40, DWW51, FGW46, FK58, Fra67b, GKR57, Goo59, Goo63c, Gra67, GR53, Han66a, Hed70, Hem60, HL52, HK44, Isa51, KZ66, Kha65c, Kim46, Kon53, Kon54, Kon66, Kud66, Kul39b, Lan65, Le 55, LS53, Mad37, Mal47, MS72b, Mor72, MS66, Mud65, Mul66, Nel66, Ney38, Ney41, Ore64, Owe66, Pea62, Pea64, Per70, Ric66, Rob64a, Rob70a, Rub68, Rus57, Sac67, Sca59, Sch43c, Sch34, Tsa54b, Tuk47b, Twe57a, Twe57b, Van66b, Var55, Vot60, Wal39, Wal41a, Wal43b, Wal44b, WW44, Wal45a, Wal49c, Wil42, Wol42, Woo66]. Statistical [vM47, Har71a, Mad63, Mos48b, Tuk47b]. Statistically [Fra51b, Ges70, Tuk47a, Tuk48b]. Statisticians [Owe67a]. Statistics [Abr67, AC65, And46, And64a, Ano40h, Ano48m, Arv70, BG52, Bah55b, Bah67, BP66a, BGP69, BN63, Bar47b, BBB60, Bel60, Bel64a, Bel64b, BD65a, BD65b, Ber70a, Bha70, BR54, BH60b, BD63, Blu63, Bly70, Bro64, Cac67, Cam42, Car34b, CS58, Che66, CGJ67, CMPT60, Col51, Con67a, Cov69, Cra42, Cra53, Dan54b, Dav62, Dav63b, Dav66, DJ68, Dem59, Den64, Den69, Den70, Den72, Des71, Dok67a, Dok68, Doo35, DH69, Dwa53, Dwa55b, Dwa57b, Dwa58, Dwa67b, EC44, Eps48, Fra52b, Fra56a, Fra65, Gei56a, Gho54, Ghu58, God49b, Goo58b, Gov62, Gov63, Gov66, Gup61, GS65, Háj68, HS49a, Har61, Har64a, Hoe48a, Hoe53b, HF38, Hot40b, Hsu46, Jof67b, Jon48, Jos71a, Kad71, Kad73, KM70, Kha59, Kha61, Kie64, Kou70b, Kou72, KR66, Kul34, Kul63]. Statistics [Kup58a, Kup58b, Lah56b, Lah57a, Lam64, LR72b, Lar35, Lau63, sL71, Lie53b, Lie55, Luk52, Mar64, McC64, MS72b, Moh71, Moo68b, Mos46, NV68, Noé72, Ols40, OS62, Pfa72, Pit57, Rag70a, RSS60, Rob44a, Rog59, RO72, Rub56, Sal66, San68, Sar54, Sar55a, Sar55b, SG56, SG58, SG69, Sav56, Sav57a, Sav59, Sav60, Sea56a, Sen70a, Sen70b, SG71, Sha66a, Sha70b, Sha71b, Sho30a, Sho30b, Sho30c, Sho69, Sho72b, Sid63, Smi57, Sri62, SHB64, SP58, Sta72a, Ste69c, Sti69, Tak64, Tan64, TC65, TC80, Tei56, Tei61c, Tra68, Wah68, Wal46e, WD51, Wei71, Wel71, Wel72, Wen60, Wij58, Wij68b, WGL66, Woo70d, uD54, Van63, vE72, vZO67, vM41, Arv69, Con65a, Dav65, HMTW47, Hub72, ST45, Ste71a, Cox62, Dav63a, Fel47, Hoe62, Lin66, MR64, Moo48]. Statistics [Moo64, Mos65, Sav71, Wat71b, Ano35a, Ano35d, Ano36b, Ano37c, Ano39d, Ano39e, Ano40d, Ano40e, Ano40g, Ano41d, Ano43c, Ano43d, Ano44d, Ano45c, Ano46f, Ano46g, Ano47e, Ano47f, Ano48e, Ano49f, Ano49e, Ano56n, Cam38a, Cra37, Cra38b, Kin30]. Statistics-the [Sav56]. Steck [Moh71]. Steepest [JF64]. Stein [Bha65, Bur68b, Cha62b, Kni63, Mos58, See53, Win72, Wor60]. Stencil [Too34]. Step [Mas51a, Roy58, Tsa54a]. Step-Down [Roy58]. Stepwise [Dem63b]. Stieltjes [Bat30, Sho30a]. Stigum [Ath70]. Stirling [Har67a]. Stochastic [AG70, AR49, AR50, And59, AW64, Arb67, Bah60b, Bel61, Ben62, Ber64a, Bha62, BW71, BEM66, Blo57, Blu54b, BR64, Bos66, Bro71a, Bro70b, Bur56, BM53, Cas50, Cha72, Chu54, Cla70, Cla71, Com64, Der56a, Der56b, DS59, Dup65, Ell72, Fab67, Fab68a, Fab68b, Fab69b, FK58, Fos53, Fri63, Geb64a, GS71, Gol67, GR53, Gre65a, Ham72, HK65a, Haq70, Hel65, Hol68, Isa69b, JS72, JK70, Joh72, Jon63, Kal54, KS68, Kaw59, Ken53, Kes58, KW52, KW70, Kim57, Koc58, Koo61, Koo64a, Koo64b, Kus66, Lah56b, Lam63, Lea66, Loy65, Luk52, MW43, MH62, Mil68b, Mil68c, Mus71, NV60, Neu71, Par60, Par57a, Par58, Phi69a, Phi69b, Pit64, Pit65, PRR67, Pra66a, Rao61a, Rei64, RM51, Rog59, Sac58, Sen71b, Sin64, Sob71, Sri68d, Sti63]. Stochastic [SC69, ST64, Str59, Str66b, Ven66, Wal48a, Wat54, Wei55c, Wic71, Wol52b, Wol54a, Wol56, WZ65, Yan68, Ylv68b, AD52, Dup70, Dwa67a, Fab69a, Ken64, Ros66a, Smi65]. Stochastically [MW47]. Stock [Alb70]. Stop [Hag67]. Stopped [CRT65, Cho67b, DF66b, Far64c, SC69, Sud71a, Tay72a, Far66c, Far66d]. Stopping [Ams63, BS68, Ber70d, BG64, Bro69, Che67a, Cho66a, CT66, DLT72, DT67, GS67, Hag66, JO69, Roo69, Ros71b, RM68, SS66a, SS67, Set70b, She69, Sie67a, Sie67b, Sie68, SSF68, Sie69b, Sta72b, Sud71b, Tay68, Tay70, TBO71, TO72, Wij71, Yah66]. Storage [Eri72, Kin62, MP71, Pra60a]. Straggler [Mos52]. Stragglers [Geb66]. Straight [Dur68, Too34, Wal40a]. Strata [SF39]. Strategies [Bla64, EZ63, Geb64b, Sud69a, Sud72]. Strategy [Enn73]. Stratification [Ekm59b, Gho63a]. Stratified [And42a, CJ56, Coc46, DS69a]. Stratum [Sen55]. Strength [AK61, BB52]. Strengthened [Mas71b, Mas72]. Strictly [Far68a, Fri72]. Striking [Ski69]. Stringent [Cog67, SS66b, Sch68a]. Strip [Has42]. Strong [BKS71, Ber69c, Ber71c, Bre60b, Cho67a, DF65, Dvo49, Jai69, Koo61, Lin72, OQ72, RR64b, Van69, Wel71, Wel72]. Strong-Mixing [Wel71, Wel72]. Strongly [Ser68b]. Structural [Bon72, Fra72a, Haq70, KR50, Lah57b, MS69a, NS51, Ney51, NS52, Sco50]. Structure [BF68a, Bel60, Ber67a, Bur62a, Con52b, Con53, Dal68, HY64, Kai71, Kap66, KP57, Lan58, Lan64, Mor72, Mus71, Pfa64, Rag60b, Rao69, SSW66, Sch70d, Sen55, Sha66b, Sha67, TO72, Dav65]. Structures [Kal65]. Stuart [Kie64, Wat71b]. Student [Rie39, Wal60, Bro67, BF63b, Chu46a, Dal46, Dun61, Hot31, MC65, Ord68, Pil51, Rob48c, Sch71b, Šid71, Wal59a, Wal47d]. Student- [Dun61]. Studentization [Che49, Wel47a]. Studentized [Ala70b, Har60c]. Student’s [Coc44, Dan40, Hen36, Lad39, TW34]. Studies [Get67a, Had63]. Study [Aro41b, BW67b, Box54a, Box54b, Car32a, Cha58, Cha65, Pie40, HMTW47]. Style [Ano62e, Ano62f, Ano62g]. Sub [Cha64, Ghu68, Nev72]. Sub-fields [Ghu68]. Sub-Martingales [Cha64]. Subclasses [Dod40]. Subdivisions [Gil62]. Subfamilies [Kon66]. Subfields [Bah55b, Bur62a, Pit57, Ski67]. Subgroup [Zac68]. Subhypothesis [Kou70a]. Subject [AS58, DK72, Eps49a, Eps51, Kei62, Lon42, Rub58, Wal40a]. Subjected [CZ64, Mus68a]. Subjective [AA63, EG59, Fer72a, Fis67b, Fis69a, Fis72b, Nor67]. Submartingale [Bha70, Cho67b]. Submartingales [Che67a]. Subminimax [FK51]. Subordination [Huf69, Teu72]. Subplots [Bak52]. Subprocess [Kun67]. Subscribers [Ano35a, Cam38a]. Subsequences [Kre71]. Subset [GS58a, GP72, Jon70, RS67, Stu67]. Subspaces [See71]. Substitution [BB68a, Sha60b]. Substitutive [Dod40]. Subvector [Len69]. Success [Sta72b]. Successes [Dar64a, Hoe56, Sam65c]. Successions [Rio65]. Successive [Bat55, Bur62b, Eze30, Fre58, Gei56a, Har42, HvN42, Hsu46, KS62, tCM59, Pow30, Ros72a, Ros72b, Sen72a, Wei55c, Wil41b, vN41, vNKBH41, vN42]. Sufficiency [Bah54, Bah55a, BL55, Ber72d, Bur61, Cam64, HWG65, HWG95, KL51, Man62, Pat64b, Ski67, Tuk49b, Wel39]. Sufficient [BG52, Bar69, Bri63a, Bro64, Bur62a, Cac61, Den64, Den69, Den70, Den72, Dha63b, Far68a, Fra52b, Fra56a, Ghu58, Ghu68, HS49a, KM70, Kim46, Koz51, LR72b, LS60, Luc67, Ols40, Pfa72, Pit57, Smi57, Ste70, Ste55, Sto70a, Wij58, Zid70]. Suggestions [Ail70]. Suitable [Lah57a]. Suitably [Kul63]. Sum [Ahu71, AB60, Bla57a, BB60, BJK56, Dav70, DWW51, Dwy40a, Fri69, Gna67, GD55, HH63, HS55, Hub63a, Kim44, Kim50, KA64a, KA64b, Ksh59b, Nan50, PL70, PL71, Ros70a, Ser70b, Sin40, Tsa55, Tsa56, Yar72, vBE65, vS50, Che52]. Summability [CT71]. Summation [Cra48]. Sums [Bak40, BK63, BKS71, Ber37, Bil62, BA49, Blo71, Blu66a, Boo72, BS70, BE70, Bur72a, Car68, CT58, CRT65, Cho68, Coc72, Coh72, DLT72, Dwy38a, Dwy38c, Eri66, Fel51, Fer70b, Fer71, Fis33a, Fis33b, Gle69, Gou63, Gra63, Gre49, Gre65a, Hey67, Hög72, IW71, Joh59, Joh61b, Kem59, Ken39, Kes72, Lig70, Mac59, Mas70a, Mas70b, Mil67a, MS70b, Min72, Mor62, Old38, Old49, PR68, Rob54, RS70, Sam66, Sam69, Sha57, Sha58b, Sha60d, Sie69a, SA57, Ste46, Str72b, SC72, Sug71, Szá72, Tay72a, Tei65, Tuk57c, Wal44c, Wal45b, Wei56a, Wen60, Woo66, Wri72, uD54, vS49, ZD59]. Super [Hey70].

Super-Critical [Hey70]. Supercritical [Ath69, Ath70, Sen68e]. Supermartingales [Cur68]. Superpositions [Huf71]. Supply [Bur32]. Supremum [Pyk59, Sto69b, MS70a]. Sure [Gil72, GHRM71, KsW69, Smi68a, Sto68c, Var68]. Surely [Gil71b]. Surface [Yos49, Yua33, Mar72b]. Surfaces [Bak37, BH57b, Fis33a, GGH59]. Surprise [Goo56, Goo57a, Red51]. Survey [Dal62]. Survival [Mac61]. Survive [Mv67]. Susan [LeV66a]. Sverdrup [Kab65]. Symbolic [DM48, KR45]. Symmetric [And72a, BM59a, Bur72b, Dre41, Dwy38a, Dwy38c, Eat70c, Fer62b, Har70c, Hof63, Hoo56b, HA72, KP68a, Luk52, Lut65, Mij61, Mij64, O’T31, O’T32, Pil64, PG68, Rao68b, Sch70d, Sch69d, Sim43, Tho64, dW72, van61a]. Symmetrical [Cad54, Con52a, Far69a, Gul72, Kie58, Oga59, Rag62, Sei63, Shr50, Spr59, Thi31i, Wal62]. Symmetrically [BA49]. Symmetrizable [Kra59]. Symmetrization [HH72c]. Symmetry [BH69, Bur72a, But69, Cha51, Con69, Gup67, Jae71a, RW72, Vor72, Vot48, Wor59]. Symposium [Ney66]. Synopsis [Sho30b, Sho30c]. System [AR49, AR50, AKKW69, Bur68a, Con63a, Cra36a, Dav68a, DV67, Fin62, Fle62b, GW65, KS68, Ley70, MD60, NP66, O’T33c, Por66, Sch70b, Wis56, HS49b, Wal49d]. Systematic [Abe33, Coc46, Gau57, Koo71, MM44, Mad49, Mad53, Too34]. Systems [Arb67, BEM66, Bla54, Bri72, BE70, Bro70d, Bro71f, Car34b, Chu43c, Dal68, Dal40, Fer72b, Fle62a, Hil31, Ios65, Ios69, JS72, KS68, Ken39, Lon42, Lon44, Mui70b, MS68b, MS71, Nar50, Sal67, Sha65b, Tru70, Ken64].

T [Dup70, Hot44a, Sal67, Str66b]. Table [Bai31, Bro59, Cau72, DS40, Dix48, Gna67, GG51, Mas52b, NV68, RK56, Smi48, Ste58a, Ste59a, Fie68, HL68, Sal30a]. Tables [Ano39b, Ano41c, DD64, Dav63a, Eps54a, Fie70, Fie71, Goo63a, Goo64a, GP67, HK64b, Har60c, Jen32, KA65, Kul71, Leh44, Lin64, Moo49, Owe56, Owe67a, Pac61, Paz70, Res62, Ros53, Ros54, Sal30b, Som58, Ste42, SE43, Tei56, Tei61c, Wei57a, Wel64a, Wil35, WB64, Owe67b, SG58, Owe66, Pea64]. Tabular [LL39, Thi31e]. Tabulation [HvN42, McP41]. Tactical [Spr55]. Tagging [Goo53a]. Tail [Abr69, BH59, BF64c, Cha72, Coh70, FP37, HW71, Kat63b, PW63, Wal51]. Takacs [Mil64a, Has63, Rei58, Rei59]. Tandem [Bur64a, MS63, Rei57, Rei63a]. Tangencies [Ylv68a]. Target [Fra51a, Fra53a, Hel70]. Tax [Alb70]. Taylor [Luk65]. Tchebychef [Cra33b]. Tchebycheff [Cam48, MR67, OP58a, Smi39, Stu68]. Tchebycheffian [Mur71b]. Teaching [Ano40h, Ano48m, Hot40b]. Team [Rad62]. Technical [Pfa67b]. Technique [BPSW70, Dwy41b]. Techniques [Hot44b, Kri70, Dal64a]. Telegraph [New68]. Telephone [Fin61, Tak59]. Tendency [Bre63, Ela62, The64]. Tennessee [Ano58m]. Term [Ber62c, Ber64b]. Terminates [Ray57]. Termination [Wij67]. Terms [Dav69c, uD54, Mar61a, ZD59]. Test [Ahm61, AB66, And60, And63b, Arn65, Bah67, Bak41, Bas64, Bha61a, Bha65, BS72, Bha67b, BJ70, BH67, Bic69a, Bir53b, Blu63, Blu67b, Bow44, BBH56, Bra52, Bro39, Büh67b, But69, Cha51, Cha63c, Cha66, CS58, Che65a, Con67b, Cro64, Dal46, Dan54a, Dar52, Dar53a, DK56, DK58, Dav69a, Dem58, Dix53, Efr67, Ela62, Far69b, Fed68, FP37, Fis60, FH55, FP69a, Gab69, Gau59, GKS63, Gir64b, GK64b, Gir65b, Gir58, Gle64, Gle66a, Gle68, Gum43b, Gup67, HK64b, HG66, Hem52, Her59, Hil60, Hil61, Hil62, Hod55a, HL56, Hoe48b, Hoe53a, Hoe37, Hoe55b, Hsu45d, Hub65, JK62, Joh40a, Kab62, KM71, KS62, KZ66, Kee51, Kem59, Kha65d, KP68b, Klo59, Kni63, KvE64b, Kra71, sL71, Loh69, MS68a, Mac41, MW42, Man45c, MW47, Mas50a, Mas51b, Mas52a]. Test [Mau40, McC39, Mik62, Mit58, Moo39, Mos62, Mui72b, Nic54, Noe50, Oga51, OS70a, Owe53, Pau50, Pau47, Pil55, PN71, PS67a, PS71, Rag65a, Rag70a, Ram56a, Ram56b, Ray57, Rob33, Ros53, Ros54, Roy53a, Sav57b, SS66a, SS67, Sch68a, Sch71b, SMR72, Sen68c, Set70b, Sil59, Sim67c, Sin66, Sta38, Ste45a, Ste56, SN68, SN69, SN70, Sug72, Sun54, Tak69b, Tro71, Tru53, Vil47, WW40, Wal42b, Wal46b, WW48, Wal46c, Wal47d, Wal47e, Wal49e, Wal49f, Wal51, Wei58d, Whi57b, Whi58a, Wij60, Wij75, Wil35, Woi72, Wol66, Van63, vZO67, BB61, Bro67, Coc52, Dar55, Dav58, Fox56, Goo57b, HL68, Kru52, Mos48a, MT50, OT47, Sac65, Tam71, WW43]. Testing [AF62, Bas68, Bos66, BB69, Bur60, Cha63a, Che49, Con66b, Con67d, Con67c, Con69, Cou44, Cov69, CS72, DS63, Dix40, DKW59, Eat67a, EZ67, ES54, Fed51, Fer36, Fle62b, Gir64a, Gir64b, Gir65b, Gru50, Har52, Her48, Hoe45, HP72, Hsu46, Hun55, Imh62, Kat61a, KOvE72, Kum65, Law63, Leh50, Leh52b, Len39, Lig71, Mat43, Mee70, Mee71, MG69, Mos52, Mul66, Nüe66, OP69, Per69a, Per71, Rei61, Ros62c, Ros63a, RW72, Sam63b, Sch42b, Sch70c, Sch62, Sch68b, Sho67, Suk58, Tan70, Tsa54b, Vot48, Wal39, WB41, Wal57, Wei53b, Wil38b, Wil46a, vM43, SE43, DKW53b]. Testing/Monte [Mul66]. Tests [Adi67a, Adi67b, AE69, AG59, AG60, Alb56, All53, And67, AG64b, AG64c, AG65, And54, AB60, Bah60b, Ban44, Bar68, BP69, BP67b, BW67a, Bas67b, Beh71, Beh72, BMT66, BD66, BD67, BH69, Ber69a, Ber72a, Ber51a, Bha61c, BJ68a, BJ68b, Bhu64, BD69, Bic69c, Bir53a, Bir54, Bir55, Bir58, Blo51, BW57b, BKR61, Blu66b, Bly58, Boz71, BC64, Bro71e, Bur72a, Bur60, BW63, Cap61, CLA62, Cha50b, Cha58, Cha65, Che52, CL54, Che65b, CS71, Con63a, CS39, Dal40, Dan40, DJ51, DJ52, Dav60, Dav61, Dav68b, DeG60, DeG61, Dem64, Des70a, Dia63, Dix54, Dok69b, DT71, Dri52, Dwa55b, Dwa56, Dwa57a, ET53, Eps54b, ES55, Far64a, Far68c, FGW46, Fra57a, Fri40, Gas65, GW68, Gee70, Gho54, GS71, Gho71b, GK64a, Gir65a, Gir68a, Gir68b]. Tests [Gir68c, Gna67, God61b, Gol63a, Goo59, GAM64, Gup69, Háj62, HK68, Her59, Her67a, Hil69a, HL63, Hoe52, HR55b, Hoe65a, Hol67, Hol71a, HP36, Ifr66, Isa51, JR69, JR70, JM72, KKW55, KB71, KW57, Kie59, KS65, KS72b, KH72, Kim47, Kim51, Klo62, Klo63, Klo64, Klo65, Kon56, Kon58, Kou70a, KP70, KP67, KP68c, Kri68, Lay72, Leh47a, Leh47b, LS48, Leh51a, LS53, Leh53, Leh59, Lev52, Lit72, Lor67, Lor72, Mac59, Mat63, MT65, MT67a, MT67b, Mee70, Mee71, Meh64, MS67, MP67, MS69b, Moo49, Moo54, Mos63, Mud65, Mui70a, Nar50, Ney38, Ols67a, Pac61, Pau41, PAAJ69, Pra58b, Pra62, PS68a, PS69c, Put55, Qua65, Ram66b, Rao72, Ros62a, Ros62b, Ros63b, Rot71, RP58, RB58, RM61, RSvZ72, SS66b, Sch69c, Sch43a, Sch67c]. Tests [Sch67d, Sch69f, Sen66, Sen67, SP67, Sen68b, SP68, Sen69, SP70, Sen71a, Sen72b, SO56, Sev57, SP69, Sid60, Sim43, Sim67b, Ski60, Spj67, Spj68a, Sri65c, Ste69a, Ste51b, Str71b, Str72c, Suk57, Tai66, Tam63, Tam66, Tei55b, Ter52, Tho38, Tin39, Tsa55, Tsa57, Tsu68, Uza60, VN67, Vat46, VA43, Vor72, Wal41a, Wal41c, WW44, Wal45a, Wal46e, Wal47c, Wal49f, Wal49g, Wal50a, Wal50b, Wal53, Wal59c, WD51, Weg56, Wei56c, Wei56d, Wei57b, Wei57d, Wei60, Wei61, Whi61b, Wij58, Wij63, Wij67, Wij68a, Wij71, Wil65, Wit60, Wol49b, Won68, WK62, Wor59, Yu71, ZJ59, Van63, Cam52, Con65b, Dar57, Gho64a, HKcY68, JB61, Pur64, Que49a, Hoe68]. th [Bue65, Dwy40a, Goo58a, Goo58b, Goo59, Hsu48, Joh67d, Lau63, MM69, vBE65]. Their [BD66, Dre40, Dwy40b, Eis39a, Jos71b, KH72, Kot64, KP70, Nor40, Pie70, RF70b, See72a, Šid68, Šid71, Sto61a, Ton69, Wij59c, Ylv68b, Van63, Bly51, Hoo56a, HF38, Ken53, Kha67a, Ros62b, Spi66, Wij57]. Theodore [Wel65]. Theodorescu [Nor70]. Theorem [AB61, Abe69b, Aus67, Bar50, Bax65, Bel70a, Bel70b, Ben62, Ber67b, Bil61b, Bir65, Bla51a, BBT58, BF64c, Ble40, Blo70, BG64, BK58a, BKR71, Bre57a, Bre57b, Bre58a, Bre60a, Bur65, Bus52a, Car50, CR63, Cho69b, Chu61, Den69, DS59, Dub68b, Dub68a, Dwa62, Dwy58, Dwy60, Eat70a, Erd49, Erd50, Fer70b, Fer71, Fos71, Fra56b, Fre62a, Fre63a, Fri69, Geb64a, GR70, Gil71b, Gra63, Gre65a, Gun67, Háj61, HS49a, HS72a, Hey68, Hey69a, Isa64b, Isa69a, Iye58, Iye61a, JO69, JK66, Joh59, Joh61b, Jos69b, KW72, Kei66, KS66c, KW65, KHS72, KHS74, wK68, Kim69, Kli67, Kli69, Koo68, KvE64b, Kre71, Kri70, Kui59, Lah58, Leh66a, LS08, Lev71, Met69, MS70b, MR66, tCM67, Nád69, Noe49, Noe55, Nor67, OS70b, Por65, PS68b, Rag70a, RF70b, Rao69]. Theorem [Rob72a, Ros66b, RO72, Sär65, Sav47, Sch67a, Sch47, Sek51, She64b, She55, Shu72, Sim70, Smi64, Smi68b, Sta65, Sto65a, Sto68a, Stu71a, Sud71b, Tak64, Tei71, Tom71, Tuc59a, VR58, VJ66, Vog60a, Wei53a, Wel71, Wij72, Wol54b, Ylv61, vE72, Bir64, Cha69, Deh71, Dev59, Fre67e, Hot44a, Kat63a, Le 54, Pla71, Sti66, Tuc59b, Tuk58c, Ylv66, Abe69a, Ber30, Bro69, Bro70a, CT58, Eic66, Ger71, GZ61, GS67, Her69, Hoe51a, Ken71, Mol31, Noe70, Par57a, Phi69b, Spr61, vB65]. Theorems [AK68b, AK71a, Bel64b, Ber64b, Bil62, Blu67b, Box54a, Box54b, Cho66b, Coh72, CC54, Csö65a, Csö65b, Csö65c, Dar65b, Don52, Doo49, DK63, Dym66, ES54, Far64c, Fel50, GH61, HH72a, Hey66, HY64, Hol71b, Igl65, Igl68, Isa72a, JN63, Kaw59, Ken72, KS66b, Kie53a, Kie53b, Kim47, Kou70c, Kus69, Lam65, LR72a, Lin72, Loè50, Mad37, Maj53, Mas71b, Mas72, McM53, Mil71b, Min72, Paz70, Pie69a, Por64b, PS64, PS68c, Rag66, RR64b, RR66, Ros52a, Sch70d, Sen68e, Sha55, Szá72, Teu68, Teh63, Ven66, Far66c, Far66d, Fel48a, Gol67, Mad71, PS69d, Sch58, Agn57, Agn59, Bha71b, Bro71b, Coc72, Eic63b, IW71, Pie69b]. Theoretic [Sch67a, Wij68b]. Theoretical [Olm40, Wat62, Wat63]. Theorie [Pyk63]. Theories [Doo68]. Theory [Abr70, AN53, Alb63, And63a, Ano30, Ano31a, Bac48, Bah50, BP64, Bar47b, Bas55b, Ben63, Ber69a, Ber68, Ber42, Beu63, BY65, BH67, Bir61b, Blu66a, Blu66b, BK60, BS64, Bre57b, Bre60a, Bro40a, Byr35a, Car30, Che56, CR65b, Chu43b, CH45, Chu61, Cop32, Cou44, Cra47c, Csö65c, Cup69a, Cur68, Cur42, Dal62, Dal69a, Dem61, Dem63b, Dep30, Dwy41a, Dwy44b, Dwy49b, ET68, Elf52, Far68c, Fel41, Fis69a, FP56, Gab69, Gei38, Gei39, Gei42b, Gei44, Gei45b, Gei45a, Get67a, Gir39, Gir46a, Gir46b, Gle65, Gle66b, GP58, Gre64, Háj64, Haj67, Hal46, HS49a, Hal51, HH43, Har70a, HS72a, Hen35a, Hen35b, HP63a, HP63b, HB66, HB67, Hur48, Imh71, Isa51, Iye50, Iye61b, Kab65, KR56, Ken53, Ken71, Kie64, Kul39b, Kul52, Kul54, Kul56a]. Theory [Kul56b, Lah58, Lam62, Law33, Le 55, LeV66a, Lee64, LS49, Leh50, Leh57a, Leh57b, Lin66, Loè50, Lot39a, Lot48, Mad63, Mad37, MM44, Mad49, Mad53, McC65, McM53, Mil61a, Mil64a, MR69b, Mon47, Moo40, Mor32, Mos62, OR54, Ols67b, Pat64b, Pat63, PJ70, Pha63, Pra60a, Pra61, Pra66a, Rao54, RSS60, Rei64, Rob65a, Roo71, Ros72a, Ros72b, San70, San61, San62, Sav56, Sav57a, Sav59, Sav60, Sch42b, Sch43a, Sch64, SB57a, Seb64a, SP67, SP68, SP70, Ser68a, Sha70a, She36, Sho67, Sie67b, Spi67, Sto65b, Sto66b, Sud71b, Tak58, Thi31q, Thi31p, Tho60, Wal39, Wal45b, Wat71b, Wat62, Wat63, Wel65, Wes59b, Wij59b, Wij59d, Wil36, Wol43, Wol58, Won35, vMD41, AT63, AD52, Iye51, Hoe68, Pyk63, Wan68, Der64a, Dym70, Han66b, LeP70, Mid66, Ney66]. Theory [Pru71, Wen68, Wol67b]. There [Woo40]. Thermodynamics [Man62]. Thiele [Cra31b]. Third [Dra60c, Dra60b, Dra61, GGH59, Her64, Tuk57e]. Thomas [Ben63, Wij68b]. Thompson [Ajg67]. Those [BB60]. Thoughts [Pea62]. Three [BD59, Bri38, Dic68b, Dra60c, Dra61, Gil71a, Imh60, Isa69a, PAAJ69, PS71, Sek51, Set50, Sim67c, SW49, BB52, Dav58]. three-sample [Dav58]. Three-Way [Imh60]. Threshold [Kei72]. Thrifty [Sud69a]. Tied [Kra71]. Ties [Büh67b, Hem52, Put55, ST68]. Tight [Nii72]. Tightened [DLS57]. Tightness [Nev72]. Timber [Has42]. Time [Alb62, AK68b, Ath68, Bat55, BS68, Ber67b, BJ68b, BR69, Bro65, CC47, CZ64, Cle71, DKW53a, DKW59, Eri71b, Fra63, Fre63a, Gav59, Gei57, Goo61c, GR53, HH72a, Joh57a, Joh72, KZ66, KK60, KK62, Kei63, Kie57, Kin62, Lig71, Loy62, Moo63b, Mus68a, Neu71, Par57c, Par57b, Par61, Rod71, Ros70b, Ros56a, SS66a, SS67, SY70b, Sch70b, Ser71, Set70b, Sha65b, She71, Sle61, Sny72, Tak61c, Tak62, Tay70, Tin39, Tre65, Wah68, Wei72b, Wic69, Wij71, Wis56, Wol48, Wol67a, Woo70c, Yac68, Ylv62, Ylv64, vS51, DKW53b, Fer72b, KW72, LZ57, Que49a]. Time-Series [Fra63, Que49a]. Times [Ber70d, Ber70e, BG64, Kei66, Ken72, Kin61b, Kin61c, Ley70, tCM59, Ore64, Pak71, Rei57, Roo69, Roo72, Tak62]. Timid [Fre67f, FP67]. Timing [Alb70]. Tito [Gna67]. Tolerance [BP66b, BP67a, BZ49, Bow46, DD64, Ell64, Fra53b, FG56, Gut57, Gut59, Gut70b, HK64a, Kem56a, Mur48, Noe51, Pau43, QG68, Rob44c, Sau60, Som58, Wal42c, Wal43a, WW46, Wal62, Wel64a, Wil41a, Wil42, Wil67, ST44, Tuk47a, Tuk48b]. too [Wal50b, Wal53]. Topics [Gup63a, Tau62]. Topological [Cra65, Pfa64, Sta63]. Topology [FV71, Pfa69b]. Torus [Rot71]. Tossing [BF64b, Dwa72, HS70b, BDF64]. Total [Dub72, Eri70b, Rob68, Mar71b]. Totals [Cau72, DS40, Dwy38b, Ste42]. Tournament [Har66b, Har68]. Tournamental [Yal71]. Tournaments [BH63]. Trace [KP66a, PM59]. Traces [Jen70a, Jen70b, KP65]. Traffic [Bre63, Dar64b, Fin61, Igl65, New60, Tak59, The64, Wan68]. Traffic-light [Dar64b]. Transducer [Con70]. Transform [Eps48, Kot65]. Transformation [Bak34, Bla57b, BP65b, HF38, Hoy68b, NS60, Pfa72, Rie31, Ros52b]. Transformations [Bak30, Bha71a, Cur43, Dok69b, Fra67a, FT50, GWA72, Nag51, Obe71, Obe72a, Ols38, Pit65, Ser71, Wij57, Wij59c, Tuk57a]. Transformed [Fra72b]. Transforms [AK68a, Bur66, Dav69b, Han59, Mil68b, Wal70a]. Transient [Bla55b, Bre57c, Bre58b, DS69b, Fin61, Fin62, Kin62, Por66, Por68, Tak61c]. Transition [Bla55b, BF68c, Fel65, Goo61c, Isa72a, Lam71b, Rou69]. Transitivity [BE72, Bon72, Eis72]. Translate [She65]. Translates [Kan69]. Translating [Eis70]. Translation [Ano31c, BG69b, Bla51b, BC68a, BC68b, Kon66, LS53, PR71a, Per70, Tor71, Tor72, Woo72]. Translation-Parameter [BG69b]. Transmission [BBT58]. Transmitting [Kem69]. Transport [GHRM71]. Trapezoidal [Car32b]. Travel [Mar48]. Trawinski [Hub63b]. Treasurer [Ano46d, Dwy45b, Dwy47c, Dwy48b, Dwy49c, Fis50b, Fis51, Fis52b, Law55, Old42a, Old43b, Old44]. Treatment [Bat30, Büh67b, FK58, Hub63b, Put55, ST68, TD63]. Treatments [Agr66a, HSF72, Hol66, Ksh59b, Meh64, Nai69, RH71, Sha65a, Sur68b]. Tree [Moo68a, NR67]. Trend [Bos66, BB69, Sau63b, Sav57a, Str58]. Trends [HB48, Too34]. Trials [Bak52, Dar64a, GG39, Hoe56, Sam65c]. Triangular [Agg72, Ano60i, Con58, Cor60, Hof60, Oga59, Rag60b, Sei63, Shr59a]. trimmed [Het68, Tam71]. Trivariate [Ste58a, Ste59a]. True [Fed68, GC64, Lor59]. Truman [Moo48]. Truncated [BW64a, BA49, Cac61, Cha56, Coh50, Coh51, Con67a, DV55, Eps54b, Gov66, Hal52, Hoe69, Kat61b, Key38, Kle71, Sha71a, Sha57, TG58, WK62, Woo72, den55]. Truncation [AG59, AG60, Bas64, Bir50, CS69, Pic70, Sch68b, Smi57, Tal63, Wil65, Tuk49b]. Trustworthy [Mil64b]. Tsao [Con67a]. Tschuprow [Der39]. Tshebyshev [BRZ47]. Tucker [Jai70, dH70]. Tukey [Mal90, Ram56b, RK57]. Twenty [Tei56, Tei61c, RW64]. Twenty-Seven [RW64]. Twin [HSF72]. Two [Abe69a, Abe69b, Abr67, AK61, AE69, Agg66, Ahm61, And62c, AG64b, AG64c, AG65, Aro47, BL55, Bak37, Bas67b, Bat31, Bat34, BMT66, BD66, Ber72f, Bha65, Bha72, Bic69a, Bil56, BRZ47, BH60a, Bli62, Blo57, BW57b, Blu63, Blu66b, Blu67b, BC68b, Bow44, Box54b, Bra52, BK58b, BN45, Bur63, Bur64b, CLA62, Cha50a, Cha63c, Con67b, CS39, Cur41, DS69a, Dem58, Dem64, Der61, Dix40, Dod41b, Dur68, DWW51, Dwa56, DT64, Ell64, FV70, Fel62, Fin66, Fre66c, Fre58, Gan61, Gas65, Gho63a, Ghu58, Gil71b, Gir68a, Gir68c, Gna59, Gna60, God61b, GZ66, Gou63, Gre50b, Gul72, Gup65, Gup70, HR55a, HH63, HV65, Har61, Har64a, Har67b, Hea56, Hem52, Her64, Hoe65b, Hoo56b, fH69, fH71, Hsu41, Imh60, Ios72, Isa51, JV71, JV72, JM72]. Two [Kab63, Kab64, Kad68, Kem63, Kem53, Ken39, KP65, KP68a, KP68b, Kie62, Kin48a, Kin61a, KOvE72, KB57, Kru54, Kul63, Lad39, Lah56b, LR72a, Lay72, LS08, Len39, LPC70, Lin50, Lin52, Lit72, Loh69, MS68a, MW47, Mas51b, Mat43, Mat55, MC65, MD71, Mik62, Mik63, Mil67b, Moh71, Moo54, Mos52, Mos58, MS64a, MS64b, Mud65, Mud66, Mui72b, NR71, NT70, Nai69, Ney51, Noé72, Ols40, Owe70, Pau70, Per69b, Pil64, PS67a, PR70, PS68c, Rag65b, Rag60b, Raj65, RSS60, Rao61b, Ric48, Ric60, Rid61, Rob44a, Rob37a, RP58, RM61, RG62, Sac72, Sam68b, San47, Sär65, Sav56, SSW66, Sch69b, Sch69c, Sch42a, Sco50, See72a, See53, Sei61, Sen68b, Ser68b, Set50, Sha65a, Shr51, Sid63, Sin67, Ski60, SP65, Sol53, Spr61, SPZ56, Sta38, Ste69a]. Two [Ste71a, Ste45a, Ste69c, SC72, SN68, Suk57, Suk58, Sun54, Sur71, Tak71a, Tan70, Tei55b, Tho64, Tsa54a, Uza60, Vog60a, Vog60b, WW40, Wal44b, Wal47d, Weg56, Wei60, WAC62, WK69, Wol46a, Wor60, Yen64, Zac70a, ZJ59, ZF64, Zim72, Van63, vM39, AB62, Blo50, BB52, Cha50b, Dri52, ET53, Mas51a, Mas52b, Rag62, Sit52, LR72b]. Two-Action [JV71, JV72]. Two-Armed [FV70, Fel62, Ber72f]. Two-Armed-Bandit [Sam68b, SP65]. Two-Associate [BK58b, KB57]. Two-Dimensional [BN45, MD71, Sol53]. Two-Factor [Owe70]. Two-level [Gul72]. Two-parameter [Zim72]. Two-Person [DWW51, fH69, fH71, Sac72]. Two-Phase [Sin67]. Two-Sample [Abr67, And62c, BMT66, BD66, BW57b, Blu63, Blu67b, Bra52, Bur63, Con67b, Dwa56, Gas65, Hea56, Hem52, Lit72, MS68a, Moh71, Moo54, Rag65b, RSS60, RG62, Sav56, Sch69b, Sch69c, Ste45a, Sun54, Tei55b, Tsa54a, Uza60, Weg56, Wei60, Wor60, Blu66b, Bur64b, Cha63c, JM72, PS68c, Ser68b, Suk57]. Two-Sided [Ell64, Kem63, Noé72, NR71]. Two-Stage [BH60a, DS69a, Ghu58, Ric60, Ski60, Tho64, Yen64, Agg66, See53]. Two-State [Der61]. Two-Way [Bha72, Box54b, Har67b, MC65, Mos52, Nai69, Shr51, WK69, Hoo56b, Sen68b]. Type [And67, BH60b, Bos66, Bou68, Bur59, Cam48, Car34a, Cha62b, Cla56, Csö65a, Csö65b, Csö65c, Cup69b, Fab72, GK50b, Her59, HD68, Kar57a, Kar58b, KP60a, Kem63, Kim42, KQ72, Kou69, KP51, KP60b, Kul48, MR69a, Mar64, MO60b, NV68, Noé72, Oga59, Que36, Ram62, RW72, Sam66, Sch69a, Sen71b, SU72, Ste69b, Ste73, Ste69c, Suk72, Twe67, VN67, Wei57a, Whi61b, Ala68, BP68, Che51, Fra57a, Kar57b, Kri70, Aro37, Boy58, Bro40a, Car32a, Cra31a, Fre61, Kul36, Kul47, Ols38, Pyk58, Raj53, Sal30a, Sal30b, Ste70, den55]. Types [Chu56, Her70, HK44, Kis45, Law71, tCM67, Her67b].

U [Dwa67a, Mor66]. Uhlenbeck [Str58]. Ulam [tCM67]. Ulf [Kal65]. Unbalanced [Har69b, Mah63, Sea61, Spj67, Tuk57d]. Unbiased [AD70, And62b, Bar49, Bas55b, BL69, Bla47, BG47, Dal40, Dav51a, DeG59, Dre41, Dwa72, EM70, GS66, GO69, GMS46, Hal46, Har65b, HS70b, Isa51, Jos71b, Ney38, OP58b, Pac61, Pat64a, Pat63, PJ70, Pet54, Pfa70, Pra62, SV69a, Sav47, Sch60b, See70a, See70b, SZ71, SQ71, Ste70, Ste50, TG58, Tat59, Zac63, Eli72]. Unbiasedness [CS71, GS60, Leh51a, Leh51b, Sch69a, SN68]. Unbiassed [Nar50]. Unbounded [Har72]. Uncertainty [DeG62]. Unconditional [Moo49]. Uncorrelated [MZ66]. Undiscounted [Mai66]. Undominated [Bur61]. Unequal [And63b, HR62, Joh64a, Koo67, KDS72, MC65, Rag62, Rob48c, Sch60c, Wal40b, Wal41b, WGL66]. Unfair [Cov66]. Unified [Bir61b, Gna67]. Uniform [Bur65, Gho64b, GHRM71, GS58b, Hol69, Isa72b, JK70, Kli69, Mar69, MH69, MH70, OQ72, Old52, Rao62, Rub56, Sch70e, Sta63, Ste71a, Wei71, Mar61a]. Uniformity [Bak52, Ber69a, Rao72]. Uniformly [AD70, Ber51a, Dav51a, Dub72, Eri70b, Far67, Loy65, Sha64a, Wic70, Jag63]. Unimodal [Cad54, Jac69, JR51, Weg69, Weg70a, Weg70b]. Unimodality [Ala72, cS67, Wol71d]. Union [Kou68]. Unions [SU72]. Unique [DM68b, FT66]. Uniqueness [Ano60i, Ant62, Bre57c, Con58, Hoe48c, Hof60, Isa64b, Ken40, Ken41, PS66, Sav47, Shr59b, Ste58b, Ste60, Wei56d, Wij63]. Unit [DG57b, GD55, Kim50]. Units [GC64, KS70, KS72c, Sen55]. Unity [Bus52b, Ray57, Sam41, vS50]. Univariate [AJ51, Ekm59b, Gna67, Jos69a, RC60b]. Universal [Dar71, HD54]. Universe [Ano35b, Lei47, Pie40, Wei35, Wil41b]. Universes [Gut59, Lad39, Mad48b, Rid31]. University [Tuk47b]. Unknown [BJ68b, Bro68, Hin65, KZ66, Kie52b, Klo70, Mee70, Pau63, PN71, Rag65b, Sim67a, SW49, Tho36, Van66b, Wal48a, Wal48b, Wei55b, Kol41]. Unobserved [Rob68]. Unreliability [Pal30]. Unrestricted [Car68]. Unsymmetrical [Sei63]. Unusual [Rub58]. Up-and-Down [Tsu67, Der57]. Upcrossing [Pan67]. Upcrossings [Dub66]. Upon [Hel70, KB71, Moo43]. Upper [Ber71b, Ber71a, BM58a, Dav69c, Dem67, Dem68, Dem69, Hec60, Kle71, Koz56, Mar72a, Sha64b, Sto72]. Urn [AK68b, Fre65a, Igl68, McF55]. Urns [Fre66a]. Use [Ban44, Bat30, Bra52, Bri63b, Bro40a, CL54, CL56, Com64, Dal46, Dod30, Dwy38b, Eis39a, For37, GG58, GW70, Har59b, Har60b, Hen31, Hot40a, Jam68, Kni63, Mos62, Mye34, Pot63, Rei61, Sha64a, Sha71a, Sto61a, Wal48c, Wij59c, Zel58, vM42, HL52, Roy53a, Wij57]. Used [Cur43, Dwa55b, Far68a, Her48, HP72, Hsu46, Kab63, Lar35, Le 70a, Kab64, Sit52]. Useful [BS64, Ekm59b, Mos46, Pil56, Sch47, Wah68, Wor60]. Usefulness [DG57b]. Uses [Chu57]. Using [But69, Car70, Dow54, Fra57c, GM72, Gov66, GS60, Gri31, Har45, Hoy68b, Jof67b, Kem53, Mos52, Owe70, QG68, Rod66a, Sri69, Wal47d, WD51]. Usual [Bro68, Jos67b, Jos69a, Jos71c]. Utilities [Fer72a, Fis69a]. Utility [Fis67a, Fis71, Fis72a, Fis72b, Hai61, Led71]. utilization [AT63].

V [Che66, Get66, Get67a, God69, Gre64, Luk65, MW69, Owe66, SC71a, Thi31h, Tul65]. Vaclav [Rei64]. Vajda [Hed70]. Valid [Wal49g]. Validation [Byr35a, ST45]. Validity [Bue59, MR69b]. Valiron [Kri70]. Valiron-type [Kri70]. Value [Bec42, BS72, Bro71d, Cho67b, Dal64b, DF66b, Eps49b, Fri72, GC64, Gum54, Har70a, Hoe55a, Kim44, Kim49, Kim56, Lie53b, Mor51, Nor37, O’T34, Pic51, Sam41, Ste45b, SC69, TD67, Tho67]. Valued [Kat67, Min72, Mur72, PS72, Rob58, Ros56a]. Values [AC65, Ant62, Ber63b, Dar52, Dar53a, Dix50, Dix51, Dod41a, Dre41, Eps49a, Eps51, Gor41b, Har62, Hoe53b, How49, How50, Hsu44, Igl72, Isa51, Kim46, Kim49, Koz56, KR66, Lor59, Loy65, Pau42b, Rob44a, She50, Sin40, SPZ56, Tei56, Tei61c, Wal47e, Wal50a, Wat54, dW72, Van67, vM39, CT56]. Variability [Eze30, da 45]. Variable [Cra36c, Enn70, Enn73, Gre44b, Kot65, Law33, Mad59, O’T32, Rio49, Rob65b, She50, Tat54, Whe56, Mar61a, TW46]. Variables [AE69, Ahm61, Ahu71, AB66, Aro47, Bal67, Bat31, BK63, Ber62c, Bha62, Bir48, BA49, Bla51b, Bla57a, Blo70, Blo71, Blo50, BS70, Bri38, Bro64, Bur72b, CT58, CR63, Cho66b, CT71, Coc72, Coh72, Cra47a, Cup69b, Cur41, DJ69, Dud68, Dun58, Dun59, DK72, Dwa62, Dwa72, Eat70c, Eic63b, EPW67, Fel51, Fer70b, Fer36, Fis42, Fri69, Gab62, Gal72, GM62, Gil71a, Gil71b, GZ61, Gre65a, Gre50b, Gui71, GS58b, Gur55b, Gur59, HK65b, HW71, Hey67, HS55, Hoe55a, HC58, Hog63, Hög72, Hol69, Hsu45b, Kac45b, Kat67, KP57, Kaw50, Kes72, KW59a, Kim44, Kin48a, Koo42, Koo68, Kot64, KA64a, KA64b, KJB67a, KJB67b, Kri70, Kue68, Kul36, Lan63a, Lig70, Mad40a, Mal68a, Mal68b, MW47, Mas71b, Mas72, Mat49, Mil61a, Mil67a, MS70b, Min72, Mur72]. Variables [New64, Ney51, NS60, Noa50, OT61, OT65, PR70, Rao69, Ros70a, Rub62a, Rub58, Sam66, Sam69, Sco50, Sek51, Sen71c, Sen72a, Ser68a, Set65, Sha57, Sha58b, Sha60d, She65, Sil48, Smi68b, Sto68c, Str65b, SC72, Sud71a, Szá72, Tho72, Tho57, Tod70, Tom71, Var66, Var68, Wal40a, Wal44c, Wal45b, Wei56a, Wei55a, Wel64a, Wic70, Wri72, vBE65, HW58, Mar61a, Mar61b, Mar71a, TW46]. Variance [AD70, And54, Ant62, Bag65, Bak49, BW57a, Ber72b, BK57a, Bir58, BH60a, Bis39, BG47, Box54a, Box54b, BBH56, BN45, CR51, CS69, Coc40a, Coh50, Coh68, Cra38a, Cur43, DJ51, DJ52, Dav51b, Dia63, Dub72, Eat67a, Ekm59a, Far69b, FM58, Fen59, FP37, Gei42b, Gem72, GW70, Gol63b, Gra54, GC64, Gre54a, Gre38, Har42, HvN42, Har67b, Har69a, Har69b, Hea61a, Hen35a, Her59, HL62, Hoo56a, Hsu45b, Jac69, Joh64b, Joh53, Joh54, Kel71, KP68b, Kie52a, Kim49, Kim51, Koo71, Kul35b, Lan71, Lan44, Leh63c, Lev66b, Mad48a, MM62, MM63, Mah63, Mor52, Mud66, Nea70, Nic54, OI65, Pac61, Pat63, PJ70, Pau50, Pau42a, Pau49, Pol70a, Pol70b, Pra63, PP69, Qua66, Ram56a, Rao68c, RR59, RC60b, RC60c, SV69a, Sat42b, Sch56a, Sch56b, Sch67c, Sea56a]. Variance [Sea56b, Sea58, Sea61, Seb64b, SZ71, Set49, SO56, Sev57, Sha71a, Sho67, Sie69b, Sim67c, Sim68, SA57, Spj67, Spj68b, Ste51a, Ste45a, SW47, Ste50, Ste45b, TG58, Tho55b, Tho62, Vaj51, Vat46, Wal40b, Wal41b, Wal42b, Wei55b, Wol50, Tuk56b, Tuk57d, Tuk57e, Tuk58b, vN41, vN42]. Variance-Component [Har69a, Har69b]. Variances [Bah64, Bak31, Bak40, Bec54, BS54, Blo71, Bro39, CS71, Gna59, Gna60, HM67, Har69b, Her63, Hoy68b, Lie53b, Mah63, MC65, McC47, Mil68e, PL70, PL71, Ram56b, RP58, Sch42a, Sea58, Sta38, SN69, SN70, Suk57, Tei57, Tho55a, Tho55b, Tru53, Tuk56b, Tuk57d, Tuk57e, Wel47a, Wil46a]. Variants [Ros52a]. Variate [Bic64, Bic69a, Con67d, Eva40, Fin61, Fré47b, Gut70b, HPW64a, HPW64b, Mau40, Ste45b]. Variates [AG64b, Car50, Che69, Con67c, DJ68, Dod41b, Gar69, Har69a, Hot40a, Jam64, Kad68, KA65, Lah56b, Mor62, Mud66, Mui72b, PW67, Pre66, Rie31, RP49, RP58, SK61, SK63, SA57, TC65, TC80, WB41, WAC62, Woo40]. Variation [Dol69, Gof53, HR36, Kul69, Loh69, Mon72b, Rob30]. Variational [MW69, Spi72]. Variations [CS69, Hey68, MG70, Par64, Sha69a, Sha70d]. Various [Dix48, Fle62a, Gna67, Kac48, Per70, Wei54b]. Variously [Vaj51]. Varying [Ahl70, Háj64, Ros72a, Ros72b]. Vector [Ber70a, Coc72, Coh65a, Coh66, Eic66, Fra63, Fra56b, Fri72, Kha62b, Kha73, KW59a, Kus69, Rao67, Rao68c, Rao69, Ros56a, RP58, Sch69b, SC71b, Sug66]. Vector-Valued [Ros56a]. Vectors [BG46, DG57b, DT57, Gle69, GD55, Kha66, KOvE72, Pol72, SR68, Sen71b, Sug65, Tei65, Yar72]. Veinott [Sch67a]. Version [Bic69a, Gun67, Hub65, Isa63, Shu72]. Versus [Hil62, RH71, Sch70c]. Very [Eav67, Wal49g]. VI [Gul72, Mal90, Pre69a, Thi31i]. View [BT65, Cop32]. Viewpoint [Che66]. VII [Thi31j]. VIII [Thi31k]. Vol [Bar47b]. Vols. [Che66]. Volume [Has42, Kie64, Mal90, Owe67a, Tau62]. Volumes [Get66]. Votaw [Con69]. Voting [Usi64]. vs [Hol66].

W [Han66a, Kud66, Mal90, Man45a, Mos48b, Ney66, Ore68b, Sal67, Tak71b, Wol67a]. W. [Dav66, ER72, Gre50b, Ney66]. WAGR [DK58, DK56]. Wahrscheinlichkeitsrechnung [Sch64]. Waiting [Cla56, Gav59, Loy62, Pak71, Rei57, Ros70b, Sch70b]. Waiting-Line [Gav59]. Waiting-Time [Loy62]. Wald [Gir48, Joh61b, Ano52j, Bla46, Blo49, Fra56b, Gho71b, Háj61, Har51, Hub72, Joh59, Kie53a, Kie53b, Le 55, Men52, Mil61b, tCM60, Noe49, Noe50, Pau47, Str72d, Tin52, Wei56d, Wol47b, Wol49a, Wol52a]. Wald-Optimality [Gho71b]. Walk [Bel70a, Bel70b, BG46, Dal69b, Dav49, Dwa67b, JP72, Kac45a, Kem63, Kes70, KW56b, Kin61b, Kin61c, Leh63d, Rai72, Sto66a, Sto69a, Wee61, Dar65a]. Walks [BHP63, Far64c, Kei63, Ken72, MD71, Mil61b, Min68, Nor66, Ros66b, Was51, Wei72b, Far66c, Far66d]. Wallace [Wat66]. Walsh [Mos65]. War [BCD^{+}45, Sta72b, Ano40g]. Waring [Gre47, Gre44a]. Wasan [Dup70]. Washington [Eis52, Kul46, Mad44, Old43a]. Watson [Ath70, Bur64b, Eri71b, Goo68, Hey70, Jof67a, KS66b, KS66c, Sen68e, Sti66]. Wave [HS72b]. Waveform [Hin65]. Way [Ban60, Bha72, Box54a, Box54b, Far69a, Fis72a, Har67b, Har69b, Imh60, KP69, MC65, Mos52, Nai69, Sea61, Shr51, WK69, ZF64, Fis71, Hoo56b, Mah63, RK56, Sen68b]. Weak [Ahl70, Bar61, Bha71b, DT68, Fer70b, Fer71, Fis69b, Her72, fH69, IT68, KS72d, Le 70b, MS72b, Neu71, PS72, PS68b, PS68c, Rao62, Sch71a, Sen71c, Sen72c, Shu72, Tay72b, Top67, Wel71, Whi70, Wic69, Wic71, Kat68]. Weakly [Koo64a, Koo64b, Wic70]. Wear [BEM66]. Wear-out [BEM66]. Weibull [HM67, Lie55]. Weighing [Ban48, Ban49b, Ban65, Ban66a, Ban66b, Ban69, Hot44b, Kem48, Kis45, Moo46, Rag59, Rag60c, Rag64, Rao66, Sur71, Zac66a]. Weight [NT70]. Weighted [Boo72, Cox56, DL70a, DL70b, Dic71, DL72, Gol63b, Gru33, Gue51, Hen32b, Kim40, Kou70c, KS72d, Sug71, Tom71, Wri72, ZJ59]. Weighting [HL64, MZ66]. Weights [Ber68, Tuk48a]. Weiss [Le 54]. Wellesley [Dwy44a]. Werner [Blu60]. Wertheimer [Zoc37]. Weyl [Hol69]. Where [Kot64]. Whether [HP72, MW47, WW40, Wal50b, Wal53]. Which [BEM59, Dod41b, Fre63a, Gil71b, GS58a, Hin65, Ney38, O’T33c, Per70, Pol34, Tay70, AJ51, Blu54a, CZ64, CR63, Dha64, Dod39, Ell62, Fre62a, GS57, Hey67, Kim44, Mus68a, Obe47, Pit63, Rob67, Ter52, WW45, Wal49g, Wol46a, vS51]. White [Eri71a, KS68, Ban70]. Whitney [Sho66, Bas67a, BK57a, Cha63c, HG66, Het68, Zar65]. Whittle [Beu65]. Whose [Hel70, Ski68, Ste45a, Bla57a, Eps49a, Eps51]. Wide [Beu63, Lea66, Ste59c]. Wide-sense [Beu63]. Widely [Mil67b]. Width [CR65b, Gra58, Kha69, Sim68, Sri71, Sta66b]. Wiener [Sal67, Gol71, Kai67, KZ71, Kai71, Kem63, Par72b, Rao70b, RS70, She67, Sim67a, Wyn69]. Wilcoxon [Arn65, Bas67a, Büh67b, FH55, Hem52, Høy68a, Klo63, Kou69, Kra71, Mos62, Ols67a, Pot63, Sen68c, Ser68b, Sho66, Zar65]. Wilkinson [Ney66]. Wilks [Cox62, Hoe62, And65b, Ano65m, BZ49, Gup71, MR71, OS70a, PAAJ69, WB41, Wal43a, van65]. Will [Bak35b]. William [Dal64a, Har71a, Spi67]. Win [Sta72b]. Winifred [Wan68]. Winner [SW72a]. Wintner [Sto70b]. Wishart [And64a, AG64a, Coc44, Jen70a, OR66, AG44, And46, And70, Bin72, Fra68b, Goo63b, Hog63, Jen70b, Ksh59a, OR62, OR64, Ras48, Sri65b, Wij59a, dW72]. Wishartness [Kha73, Kha62b]. Within [Far68b, GC64, Bak49]. Without [Bon72, CR51, DJL59, Kes72, Ros72a, Ros72b, SS69, Sta38, Str65b, AM67, FK72, FG52, HR62, Rio65]. Wloka [Pyk63]. Wolfowitz [Fra56b, Háj61, Sak66, Der56a, Noe49, Noe50, Sak64, Ven67b]. Words [Joh66]. Work [Cra53, Hem60, Men52, Cra62]. Works [Tau62, Mal90]. wrong [Cle71].

X [Thi31l]. XI [Thi31m]. XII [Thi31n]. XIII [Thi31o]. XIV [Thi31p]. XV [Thi31q]. XVI [Thi31r]. XVII [Thi31s].

Yakowitz [Zac70b]. Yates [GS60]. Years [Men52]. Ylvisaker [Wah71a]. York [Ano43i, Ano50i, Dwy48c, Fis50a, Lit48, Lit49, Old42c]. Youden [ER72, HSF72]. Yu [Luk65, Mul66]. Yu. [Luk65, Mas71a].

Zbynek [Hoe68]. Zelen [Mad63]. Zero [Ahl70, BM59a, BLS64, DWW51, Far62, Goo61a, Gri31, Hsu44, Hsu48, Joh51, NR62, OS70b, Pin47, HS70c, BP72]. Zero-One [BLS64, HS70c]. Zero-Order [Gri31]. Zero-Sum [DWW51]. Zeroes [Sha69b]. Zeros [Gem72, Ste71c, Ylv65]. Zoch [Wer37]. Zonal [Jam68, Sug71]. Zygmund [Cho69b, Gun67].

References

Anderson:1950:DCS

[AA50] R. L. Anderson and T. W. Anderson. Distribution of the circular serial correlation coefficient for residuals from a fitted Fourier series. Annals of Mathematical Statistics, 21(1):59–81, March 1950. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177729886.

Anscombe:1963:DSP

[AA63] F. J. Anscombe and R. J. Aumann. A definition of subjective probability. Annals of Mathematical Statistics, 34(1):199–205, March 1963. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177704255.

Ansari:1960:RST

[AB60] A. R. Ansari and R. A. Bradley. Rank-sum tests for dispersions. Annals of Mathematical Statistics, 31(4):1174–1189, December 1960. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177705688.

Abbott:1961:TRC

[AB61] J. H. Abbott and J. R. Blum. On a theorem of Rényi concerning mixing sequences of sets. Annals of Mathematical Statistics, 32(1): 257–260, March 1961. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177705157.

Anderson:1962:CTM

[AB62] T. W. Anderson and R. R. Bahadur. Classification into two multivariate normal distributions with different covariance matrices. Annals of Mathematical Statistics, 33(2):420–431, June 1962. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177704568.

Allen:1966:STI

[AB66] J. L. Allen and J. A. Beekman. A statistical test involving a random number of random variables. Annals of Mathematical Statistics, 37 (5):1305–1311, October 1966. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177699274.

Allen:1967:DMK

[AB67] J. L. Allen and J. A. Beekman. Distributions of a M. Kac statistic. Annals of Mathematical Statistics, 38(6):1919–1923, December 1967. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177698629.

Abernethy:1933:ESE

[Abe33] John R. Abernethy. On the elimination of systematic errors due to grouping. Annals of Mathematical Statistics, 4(4):263–277, November 1933. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177732755.

Ayer:1955:EDF

[ABE^{+}55] Miriam Ayer, H. D. Brunk, G. M. Ewing, W. T. Reid, and Edward Silverman. An empirical distribution function for sampling with incomplete information. Annals of Mathematical Statistics, 26(4):641– 647, December 1955. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177728423.

Abe:1969:CLT

[Abe69a] O. Abe. A Central Limit Theorem for the number of edges in the random intersection of two graphs. Annals of Mathematical Statistics, 40(1):144–151, February 1969. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/ euclid.aoms/1177697811. See correction [Abe69b].

Abe:1969:CNC

[Abe69b] O. Abe. Correction notes: Correction to “A Central Limit Theorem for the Number of Edges in the Random Intersection of Two Graphs”. Annals of Mathematical Statistics, 40(4):1510, August 1969. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177697527. See

[Abe69a].

Abrahamson:1964:OPQ

[Abr64] I. G. Abrahamson. Orthant probabilities for the quadrivariate normal distribution. Annals of Mathematical Statistics, 35(4):1685–1703, December 1964. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177700391.

Abrahamson:1967:EBE

[Abr67] Innis G. Abrahamson. Exact Bahadur efficiencies for the Kolmogorov– Smirnov and Kuiper one- and two-sample statistics. Annals of Mathematical Statistics, 38(5):1475–1490, October 1967. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http:/

/projecteuclid.org/euclid.aoms/1177698702.

Abrahamse:1969:TFM

[Abr69] Allan F. Abrahamse. The tail field of a Markov chain. Annals of Mathematical Statistics, 40(1):127–136, February 1969. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177697809.

Abrahamse:1970:CBM

[Abr70] Allan F. Abrahamse. A comparison between the Martin boundary theory and the theory of likelihood ratios. Annals of Mathematical Statistics, 41(3):1064–1067, June 1970. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177696982.

Ali:1965:SBE

[AC65] Mir M. Ali and Lai K. Chan. Some bounds for expected values of order statistics. Annals of Mathematical Statistics, 36(3):1055– 1057, June 1965. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177700081.

Atkinson:1964:DPF

[ACH64] F. V. Atkinson, J. D. Church, and B. Harris. Decision procedures for finite decision problems under complete ignorance. Annals of Mathematical Statistics, 35(4):1644–1655, December 1964. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177700387.

Anderson:1952:ATC

[AD52] T. W. Anderson and D. A. Darling. Asymptotic theory of certain ‘goodness of fit’ criteria based on stochastic processes. Annals of Mathematical Statistics, 23(2):193–212, June 1952. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http:/

/projecteuclid.org/euclid.aoms/1177729437;   http://www.
jstor.org/stable/2236446.

Abbey:1970:CUM

[AD70] J. L. Abbey and H. T. David. The construction of uniformly minimum variance unbiased estimators for exponential distributions. Annals of Mathematical Statistics, 41(4):1217–1222, August 1970. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177696896.

Adichie:1967:AEC

[Adi67a] J. N. Adichie. Asymptotic efficiency of a class of non-parametric tests for regression parameters. Annals of Mathematical Statistics, 38(3): 884–893, June 1967. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177698882.

Adichie:1967:ERP

[Adi67b] J. N. Adichie. Estimates of regression parameters based on rank tests. Annals of Mathematical Statistics, 38(3):894–904, June 1967. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177698883.

Afifi:1969:MTS

[AE69] A. A. Afifi and R. M. Elashoff. Multivariate two sample tests with dichotomous and continuous variables. I. The location model. Annals of Mathematical Statistics, 40(1):290–298, February 1969. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177697824.

Appleby:1962:EEM

[AF62] R. H. Appleby and R. J. Freund. An empirical evaluation of multivariate sequential procedure for testing means. Annals of Mathematical Statistics, 33(4):1413–1420, December 1962. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177704373.

Austin:1957:DNL

[AFLP57] T. Austin, R. Fagen, T. Lehrer, and W. Penney. The distribution of the number of locally maximal elements in a random sample. Annals of Mathematical Statistics, 28(3):786–790, September 1957. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177706893.

Austin:1959:NCR

[AFPR59] T. L. Austin, R. E. Fagen, W. F. Penney, and John Riordan. The number of components in random linear graphs. Annals of Mathematical Statistics, 30(3):747–754, September 1959. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177706204.

Anderson:1944:SEW

[AG44] T. W. Anderson and M. A. Girshick. Some extensions of the Wishart distribution. Annals of Mathematical Statistics, 15(4):345–357, December 1944. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/

1177731206. See corrections [AG64a].

Anderson:1957:SIA

[AG57] T. W. Anderson and Leo A. Goodman. Statistical inference about Markov chains. Annals of Mathematical Statistics, 28(1):89–110, March 1957. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177707039.

Aggarwal:1959:TTH

[AG59] Om P. Aggarwal and Irwin Guttman. Truncation and tests of hypotheses. Annals of Mathematical Statistics, 30(1):230–238, March 1959. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/

1177706378. See correction [AG60].

Aggarwal:1960:CNC

[AG60] Om P. Aggarwal and Irwin Guttman. Correction notes: Corrections to “Truncation and Tests of Hypotheses”. Annals of Mathematical Statistics, 31(4):1213, December 1960. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177705695. See [AG59].

Anderson:1964:CSE

[AG64a] T. W. Anderson and M. A. Girshick. Corrections to “Some Extensions of the Wishart Distribution”. Annals of Mathematical Statistics, 35(2):923, June 1964. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177703606. See [AG44].

Anderson:1964:MPF

[AG64b] T. W. Anderson and S. Das Gupta. Monotonicity of the power functions of some tests of independence between two sets of variates. Annals of Mathematical Statistics, 35(1):206–208, March 1964. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177703743.

Anderson:1964:MPP

[AG64c] T. W. Anderson and S. Das Gupta. A monotonicity property of the power functions of some tests of the equality of two covariance matrices. Annals of Mathematical Statistics, 35(3):1059–1063, September 1964. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/

1177703264. See correction [AG65].

Anderson:1965:CNC

[AG65] T. W. Anderson and S. Das Gupta. Correction notes: Correction to “A Monotonicity Property of the Power Functions of Some Tests of the Equality of Two Covariance Matrices”. Annals of Mathematical Statistics, 36(4):1318, August 1965. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177700008. See [AG64c].

Albert:1970:ASA

[AG70] A. E. Albert and L. A. Gardner, Jr. An addendum to “Stochastic Approximation and Nonlinear Regression”. Annals of Mathematical Statistics, 41(2):646–647, April 1970. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177697105.

Aggarwal:1955:SMI

[Agg55] Om P. Aggarwal. Some minimax invariant procedures for estimating a cumulative distribution function. Annals of Mathematical Statistics, 26(3):450–463, September 1955. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177728490.

Aggarwal:1959:BMP

[Agg59] Om P. Aggarwal. Bayes and minimax procedures in sampling from finite and infinite populations — I. Annals of Mathematical Statistics, 30(1):206–218, March 1959. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177706376.

Aggarwal:1966:BMP

[Agg66] Om P. Aggarwal. Bayes and minimax procedures for estimating the arithmetic mean of a population with two-stage sampling. Annals of Mathematical Statistics, 37(5):1186–1195, October 1966. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177699264.

Aggarwal:1972:NCT

[Agg72] K. R. Aggarwal. A note on construction of triangular P.B.I.B. design with parameters v = 21, b = 35, r = 10, k = 6, λ_{1} = 2 and λ_{2} = 3. Annals of Mathematical Statistics, 43(1):371, February 1972. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177692736.

Agnew:1957:EGC

[Agn57] Ralph Palmer Agnew. Estimates for global Central Limit Theorems. Annals of Mathematical Statistics, 28(1):26–42, March 1957. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177707035.

Agnew:1959:AEG

[Agn59] Ralph Palmer Agnew. Asymptotic expansions in global Central Limit Theorems. Annals of Mathematical Statistics, 30(3):721–737, September 1959. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177706201.

Agrawal:1966:CBN

[Agr66a] Hira Lal Agrawal. Comparison of the bounds of the number of common treatments between blocks of certain partially balanced incomplete block designs. Annals of Mathematical Statistics, 37(3):739– 740, June 1966. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177699473.

Agrawal:1966:SGD

[Agr66b] Hiralal Agrawal. Some generalizations of distinct representatives with applications to statistical designs. Annals of Mathematical Statistics, 37(2):525–528, April 1966. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177699539.

Adams:1971:ACB

[AGW71] J. E. Adams, H. L. Gray, and T. A. Watkins. An asymptotic characterization of bias reduction by jackknifing. Annals of Mathematical Statistics, 42(5):1606–1612, October 1971. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177693159.

Al-Hussaini:1971:ISC

[AH71] A. Al-Hussaini. Independent sequences with the Carefoot–Flett property. Annals of Mathematical Statistics, 42(2):789–793, April 1971. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177693433.

Ahlswede:1970:NEW

[Ahl70] R. Ahlswede. A note on the existence of the weak capacity for channels with arbitrarily varying channel probability functions and its relation to Shannon’s zero error capacity. Annals of Mathematical Statistics, 41(3):1027–1033, June 1970. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177696979.

Ahlswede:1971:GCD

[Ahl71] R. Ahlswede. Group codes do not achieve Shannon’s channel capacity for general discrete channels. Annals of Mathematical Statistics, 42 (1):224–240, February 1971. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177693508.

Ahmed:1961:LMP

[Ahm61] Mohamad Salahuddin Ahmed. On a locally most powerful boundary randomized similar test for the independence of two Poisson variables. Annals of Mathematical Statistics, 32(3):809–827, September 1961. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177704975.

Ahuja:1970:CRR

[Ahu70] J. C. Ahuja. On a certain recurrence relation. Annals of Mathematical Statistics, 41(2):665–666, April 1970. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177697110.

Ahuja:1971:DSI

[Ahu71] J. C. Ahuja. Distribution of the sum of independent decapitated negative binomial variables. Annals of Mathematical Statistics, 42 (1):383–384, February 1971. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177693527.

Aigner:1971:NCC

[Aig71] Martin Aigner. Note on the characterization of certain association schemes. Annals of Mathematical Statistics, 42(1):363–367, February 1971. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177693522.

Ailam:1970:ADM

[Ail70] Gedalia Ailam. The asymptotic distribution of the measure of random sets with application to the classical occupancy problem and suggestions for curve fitting. Annals of Mathematical Statistics, 41 (2):427–439, April 1970. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177697082.

Albert:1951:ECI

[AJ51] G. E. Albert and Ralph B. Johnson. On the estimation of central intervals which contain assigned proportions of a normal univariate population. Annals of Mathematical Statistics, 22(4):596–599, December 1951. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177729551.

Archbold:1956:MCP

[AJ56] J. W. Archbold and N. L. Johnson. A method of constructing partially balanced incomplete block designs. Annals of Mathematical Statistics, 27(3):624–632, September 1956. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177728172.

Archbold:1958:CRS

[AJ58] J. W. Archbold and N. L. Johnson. A construction for Room’s squares and an application in experimental design. Annals of Mathematical Statistics, 29(1):219–225, March 1958. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177706719.

Andrews:1965:CNC

[AJ65] Fred Andrews and Kumar Jogdeo. Correction note: Correction to abstract “Asymptotic Efficiency of Certain Distance Criteria”. Annals of Mathematical Statistics, 36(2):715, April 1965. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177700184. See [Ano64d].

Ajgaonkar:1967:HTC

[Ajg67] S. G. Prabhu Ajgaonkar. On Horvitz and Thompson’s T_{1} class of linear estimators. Annals of Mathematical Statistics, 38(6):1882–1886, December 1967. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177698621.

Addelman:1961:SME

[AK61] Sidney Addelman and Oscar Kempthorne. Some main-effect plans and orthogonal arrays of strength two. Annals of Mathematical Statistics, 32(4):1167–1176, December 1961. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177704855.

Andersen:1968:SIT

[AK68a] G. Andersen and T. Kawata. Some integral transforms of characteristic functions. Annals of Mathematical Statistics, 39(6):1923–1931, December 1968. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177698022.

Athreya:1968:EUS

[AK68b] Krishna B. Athreya and Samuel Karlin. Embedding of urn schemes into continuous time Markov branching processes and related limit theorems. Annals of Mathematical Statistics, 39(6):1801–1817, December 1968. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177698013.

Athreya:1971:BPRb

[AK71a] Krishna B. Athreya and Samuel Karlin. Branching processes with random environments, II: Limit theorems. Annals of Mathematical Statistics, 42(6):1843–1858, December 1971. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177693051.

Athreya:1971:BPRa

[AK71b] Krishna B. Athreya and Samuel Karlin. On branching processes with random environments: I: Extinction probabilities. Annals of Mathematical Statistics, 42(5):1499–1520, October 1971. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177693150.

Anderson:1969:CEP

[AKKW69] William N. Anderson, Jr., George B. Kleindorfer, Paul R. Kleindorfer, and Michael B. Woodroofe. Consistent estimates of the parameters of a linear system. Annals of Mathematical Statistics, 40 (6):2064–2075, December 1969. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177697286.

Alam:1968:SRP

[Ala68] Khursheed Alam. Some results on Pólya type 2 distributions. Annals of Mathematical Statistics, 39(5):1759–1761, October 1968. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177698163.

Alam:1970:MPM

[Ala70a] Khursheed Alam. Monotonicity properties of the multinomial distribution. Annals of Mathematical Statistics, 41(1):315–317, February 1970. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177697211.

Alam:1970:MPD

[Ala70b] Khursheed Alam. A monotonicity property of the distribution of the Studentized smallest chi-square. Annals of Mathematical Statistics, 41(1):318–320, February 1970. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177697212.

Alam:1972:UDO

[Ala72] Khursheed Alam. Unimodality of the distribution of an order statistic. Annals of Mathematical Statistics, 43(6):2041–2044, December 1972. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177690881.

Albert:1947:NFI

[Alb47] G. E. Albert. A note on the fundamental identity of sequential analysis. Annals of Mathematical Statistics, 18(4):593–596, December 1947. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/

1177730352. See correction [Alb48].

Albert:1948:CNF

[Alb48] G. E. Albert. Correction to “A Note on the Fundamental Identity of Sequential Analysis”. Annals of Mathematical Statistics, 19(3):426– 427, September 1948. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177730213. See [Alb47].

Albert:1954:CSC

[Alb54] G. E. Albert. On the computation of the sampling characteristics of a general class of sequential decision problems. Annals of Mathematical Statistics, 25(2):340–356, June 1954. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177728790.

Albert:1956:AST

[Alb56] G. E. Albert. Accurate sequential tests on the mean of an exponential distribution. Annals of Mathematical Statistics, 27(2):460– 470, June 1956. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177728269.

Albert:1961:SDE

[Alb61] Arthur E. Albert. The sequential design of experiments for infinitely many states of nature. Annals of Mathematical Statistics, 32(3):774– 799, September 1961. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177704973.

Albert:1962:EIG

[Alb62] Arthur Albert. Estimating the infinitesimal generator of a continuous time, finite state Markov process. Annals of Mathematical Statistics, 33(2):727–753, June 1962. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177704594.

Albert:1963:MTP

[Alb63] Arthur Albert. A mathematical theory of pattern recognition. Annals of Mathematical Statistics, 34(1):284–299, March 1963. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177704265.

Albert:1966:FSC

[Alb66] Arthur Albert. Fixed size confidence ellipsoids for linear regression parameters. Annals of Mathematical Statistics, 37(6):1602–1630, December 1966. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177699151.

Albert:1970:OTS

[Alb70] Arthur Albert. Optimally timing the sale of stock when the tax man is breathing down your neck. Annals of Mathematical Statistics, 41 (2):626–641, April 1970. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177697103.

Allen:1953:CMT

[All53] S. G. Allen. A class of minimax tests for one-sided composite hypotheses. Annals of Mathematical Statistics, 24(2):295–298, June 1953. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177729036.

Alter:1937:SFP

[Alt37] Dinsmore Alter. A simple form of periodogram. Annals of Mathematical Statistics, 8(2):121–126, June 1937. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177732434.

Alter:1939:CSM

[Alt39] Dinsmore Alter. Correction of sample moment bias due to lack of high contact and to histogram grouping. Annals of Mathematical Statistics, 10(2):192–195, June 1939. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177732217.

Abramson:1967:PRF

[AM67] Morton Abramson and W. O. J. Moser. Permutations without rising or falling w-sequences. Annals of Mathematical Statistics, 38(4):1245– 1254, August 1967. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177698793.

Amster:1963:MBS

[Ams63] Sigmund J. Amster. A modified Bayes stopping rule. Annals of Mathematical Statistics, 34(4):1404–1413, December 1963. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177703872.

Albert:1953:CST

[AN53] G. E. Albert and Lewis Nelson. Contributions to the statistical theory of counter data. Annals of Mathematical Statistics, 24(1):9–22, March 1953. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177729079.

Andersson:1934:NMC

[And34] Walter Andersson. On a new method of computing non-linear regression curves. Annals of Mathematical Statistics, 5(2):81–106, June 1934. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177732708.

Anderson:1942:DSS

[And42a] Paul H. Anderson. Distributions in stratified sampling. Annals of Mathematical Statistics, 13(1):42–52, March 1942. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177731641.

Anderson:1942:DSC

[And42b] R. L. Anderson. Distribution of the serial correlation coefficient. Annals of Mathematical Statistics, 13(1):1–13, March 1942. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177731638.

Anderson:1943:CM

[And43] T. W. Anderson. On card matching. Annals of Mathematical Statistics, 14(4):426–435, December 1943. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177731362.

Anderson:1946:NCW

[And46] T. W. Anderson. The non-central Wishart distribution and certain problems of multivariate statistics. Annals of Mathematical Statistics, 17(4):409–431, December 1946. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/ euclid.aoms/1177730882. See corrections [And64a].

Anderson:1951:ELR

[And51a] T. W. Anderson. Estimating linear restrictions on regression coefficients for multivariate normal distributions. Annals of Mathematical Statistics, 22(3):327–351, September 1951. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177729580. See [And80].

Anderson:1951:REB

[And51b] T. W. Anderson. Report of the Editor of the Annals for 1950. Annals of Mathematical Statistics, 22(1):163–164, March 1951. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177729714.

Anderson:1952:REB

[And52] T. W. Anderson. Report of the Editor of the Annals for 1951. Annals of Mathematical Statistics, 23(1):166–167, March 1952. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177729500.

Andrews:1954:ABS

[And54] Fred C. Andrews. Asymptotic behavior of some rank tests for analysis of variance. Annals of Mathematical Statistics, 25(4):724–736, December 1954. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177728658.

Anderson:1959:ADE

[And59] T. W. Anderson. On asymptotic distributions of estimates of parameters of stochastic difference equations. Annals of Mathematical Statistics, 30(3):676–687, September 1959. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177706198.

Anderson:1960:MSP

[And60] T. W. Anderson. A modification of the sequential probability ratio test to reduce the sample size. Annals of Mathematical Statistics, 31(1):165–197, March 1960. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177705996.

Anderson:1962:CDP

[And62a] T. W. Anderson. The choice of the degree of a polynomial regression as a multiple decision problem. Annals of Mathematical Statistics, 33(1):255–265, March 1962. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177704729.

Anderson:1962:LSB

[And62b] T. W. Anderson. Least squares and best unbiased estimates. Annals of Mathematical Statistics, 33(1):266–272, March 1962. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177704730.

Anderson:1962:DTS

[And62c] T. W. Anderson. On the distribution of the two-sample Cramér–von Mises criterion. Annals of Mathematical Statistics, 33(3):1148–1159, September 1962. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177704477.

Anderson:1963:ATP

[And63a] T. W. Anderson. Asymptotic theory for principal component analysis. Annals of Mathematical Statistics, 34(1):122–148, March 1963. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177704248.

Anderson:1963:TEM

[And63b] T. W. Anderson. A test for equality of means when covariance matrices are unequal. Annals of Mathematical Statistics, 34(2):671– 672, June 1963. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177704181.

Anderson:1964:CNC

[And64a] T. W. Anderson. Corrections to “The Non-Central Wishart Distribution and Certain Problems of Multivariate Statistics”. Annals of Mathematical Statistics, 35(2):923–924, June 1964. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177703607. See [And46].

Anderson:1964:BPP

[And64b] T. W. Anderson. On Bayes procedures for a problem with choice of observations. Annals of Mathematical Statistics, 35(3):1128–1135, September 1964. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177703270.

Anderson:1965:AED

[And65a] George A. Anderson. An asymptotic expansion for the distribution of the latent roots of the estimated covariance matrix. Annals of Mathematical Statistics, 36(4):1153–1173, August 1965. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177699989.

Anderson:1965:SSW

[And65b] T. W. Anderson. Samuel Stanley Wilks, 1906–1964. Annals of Mathematical Statistics, 36(1):1–23, February 1965. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177700267.

Anderson:1965:SOC

[And65c] T. W. Anderson. Some optimum confidence bounds for roots of determinantal equations. Annals of Mathematical Statistics, 36(2): 468–488, April 1965. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177700158.

Andel:1967:LAP

[And67] Jiří Anděl. Local asymptotic power and efficiency of tests of Kolmogorov–Smirnov type. Annals of Mathematical Statistics, 38 (6):1705–1725, December 1967. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177698605.

Anderson:1970:AEN

[And70] George A. Anderson. An asymptotic expansion for the noncentral Wishart distribution. Annals of Mathematical Statistics, 41(5):1700– 1707, October 1970. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177696814.

Andel:1971:LPA

[And71a] Jiří Anděl. Local properties of the autoregressive series. Annals of Mathematical Statistics, 42(1):67–74, February 1971. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177693495.

Andel:1971:MAS

[And71b] Jiří Anděl. On the multiple autoregressive series. Annals of Mathematical Statistics, 42(2):755–759, April 1971. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177693425.

Anderson:1971:CMN

[And71c] Mary R. Anderson. A characterization of the multivariate normal distribution. Annals of Mathematical Statistics, 42(2):824–827, April 1971. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177693441.

Andel:1972:SRM

[And72a] Jiří Anděl. Symmetric and reversed multiple stationary autoregressive series. Annals of Mathematical Statistics, 43(4):1197–1203, August 1972. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177692471.

Anderson:1972:DPF

[And72b] Donald A. Anderson. Designs with partial factorial balance. Annals of Mathematical Statistics, 43(4):1333–1341, August 1972. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177692486.

Anderson:1980:CEL

[And80] T. W. Anderson. Correction to “Estimating Linear Restrictions on Regression Coefficients for Multivariate Normal Distributions”. Annals of Statistics, 8(6):1400, November 1980. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176345214. See correc-

tion [And51a].

Anonymous:1930:EFT

[Ano30] Anonymous. Editorial: Fundamentals of the theory of sampling. Annals of Mathematical Statistics, 1(1):101–121, February 1930. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177733263.

Anonymous:1931:EIS

[Ano31a] Anonymous. Editorial: The interdependence of sampling and frequency distribution theory. Annals of Mathematical Statistics, 2(1):

82–98, February 1931. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177733044.

Anonymous:1931:EN

[Ano31b] Anonymous. Editor’s note. Annals of Mathematical Statistics, 2(4): 479, November 1931. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177732953.

Anonymous:1931:RTB

[Ano31c] Anonymous. Reprint and translation from Nordisk Forsäkringstidskrift 1930. Annals of Mathematical Statistics, 2(4):485–492, November 1931. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177732954.

Anonymous:1933:ENC

[Ano33] Anonymous. Editorial: Note on the computation and modification of moments. Annals of Mathematical Statistics, 4(3):229–239, August 1933. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177732781.

Anonymous:1935:DSB

[Ano35a] Anonymous. 1935 directory of subscribers to the Annals of Mathematical Statistics. Annals of Mathematical Statistics, 6(4):228–232, December 1935. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177732570.

Anonymous:1935:ECD

[Ano35b] Anonymous. Errata: On certain distribution functions when the law of the universe is Poisson’s first law of error. Annals of Mathematical Statistics, 6(2):110, June 1935. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177732616. See [Wei35].

Anonymous:1935:ESP

[Ano35c] Anonymous. Errata: Some practical interpolation formulas. Annals of Mathematical Statistics, 6(3):142, September 1935. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177732590. See [Rob35].

Anonymous:1935:NOI

[Ano35d] Anonymous. Notice of the organization of the Institute of Mathematical Statistics. Annals of Mathematical Statistics, 6(4):227, December 1935. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177732569.

Anonymous:1936:EFN

[Ano36a] Anonymous. Editorial: The fundamental nature and proof of Sheppard’s adjustments. Annals of Mathematical Statistics, 7(3):154–163, September 1936. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177732506.

Anonymous:1936:RMI

[Ano36b] Anonymous. Report of the meeting of the Institute of Mathematical Statistics at st. Louis. Annals of Mathematical Statistics, 7(1): 54, March 1936. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177732547.

Anonymous:1937:L

[Ano37a] Anonymous. By-laws. Annals of Mathematical Statistics, 8(4):234– 235, December 1937. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177732393.

Anonymous:1937:C

[Ano37b] Anonymous. Constitution. Annals of Mathematical Statistics, 8 (4):232–234, December 1937. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177732392.

Anonymous:1937:DIM

[Ano37c] Anonymous. Directory of Institute of Mathematical Statistics. Annals of Mathematical Statistics, 8(4):236–238, December 1937. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177732394.

Anonymous:1939:AP

[Ano39a] Anonymous. Abstracts of papers. Annals of Mathematical Statistics, 10(1):84–90, March 1939. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177732251.

Anonymous:1939:ACC

[Ano39b] Anonymous. Announcement concerning computation of mathematical tables. Annals of Mathematical Statistics, 10(4):399–401, December 1939. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177732155.

Anonymous:1939:L

[Ano39c] Anonymous. By-laws. Annals of Mathematical Statistics, 10(1):93– 94, March 1939. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177732253.

Anonymous:1939:CIM

[Ano39d] Anonymous. Constitution of the Institute of Mathematical Statistics. Annals of Mathematical Statistics, 10(1):91–92, March 1939. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177732252.

Anonymous:1939:DIM

[Ano39e] Anonymous. Directory of Institute of Mathematical Statistics. Annals of Mathematical Statistics, 10(1):95–98, March 1939. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177732254.

Anonymous:1940:APa

[Ano40a] Anonymous. Abstracts of papers. Annals of Mathematical Statistics, 11(1):110–115, March 1940. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177731951.

Anonymous:1940:APb

[Ano40b] Anonymous. Abstracts of papers. Annals of Mathematical Statistics, 11(4):475–478, December 1940. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177731837.

Anonymous:1940:L

[Ano40c] Anonymous. By-laws. Annals of Mathematical Statistics, 11(1):117– 119, March 1940. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177731953.

Anonymous:1940:CIM

[Ano40d] Anonymous. Constitution of the Institute of Mathematical Statistics. Annals of Mathematical Statistics, 11(1):116–117, March 1940. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177731952.

Anonymous:1940:DIM

[Ano40e] Anonymous. Directory of the Institute of Mathematical Statistics. Annals of Mathematical Statistics, 11(1):120–123, March 1940. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177731954.

Anonymous:1940:RHM

[Ano40f] Anonymous. Report of the Hanover meeting of the Institute. Annals of Mathematical Statistics, 11(4):473–474, December 1940. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177731836.

Anonymous:1940:RWP

[Ano40g] Anonymous. Report of the War Preparedness Committee of the Institute of Mathematical Statistics. Annals of Mathematical Statistics, 11(4):479–484, December 1940. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177731838.

Anonymous:1940:RTS

[Ano40h] Anonymous. Resolutions on the teaching of statistics. Annals of Mathematical Statistics, 11(4):472, December 1940. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177731835.

Anonymous:1941:APa

[Ano41a] Anonymous. Abstracts of papers. Annals of Mathematical Statistics, 12(1):123–125, March 1941. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177731795.

Anonymous:1941:APb

[Ano41b] Anonymous. Abstracts of papers. Annals of Mathematical Statistics, 12(4):470–473, December 1941. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177731688.

Anonymous:1941:ACC

[Ano41c] Anonymous. Announcement concerning computation of mathematical tables. Annals of Mathematical Statistics, 12(4):465–467, December 1941. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177731686.

Anonymous:1941:CIM

[Ano41d] Anonymous. Constitution of the Institute of Mathematical Statistics. Annals of Mathematical Statistics, 12(4):474, December 1941. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177731689.

Anonymous:1942:APa

[Ano42a] Anonymous. Abstracts of papers. Annals of Mathematical Statistics, 13(1):110–113, March 1942. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177731653.

Anonymous:1942:APb

[Ano42b] Anonymous. Abstracts of papers. Annals of Mathematical Statistics, 13(4):450–453, December 1942. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177731547.

Anonymous:1943:AP

[Ano43a] Anonymous. Abstracts of papers. Annals of Mathematical Statistics, 14(4):445–447, December 1943. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177731368.

Anonymous:1943:L

[Ano43b] Anonymous. By-laws. Annals of Mathematical Statistics, 14(1):104– 105, March 1943. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177731502.

Anonymous:1943:CIM

[Ano43c] Anonymous. Constitution of the Institute of Mathematical Statistics. Annals of Mathematical Statistics, 14(1):102–104, March 1943.

CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177731501.

Anonymous:1943:DIM

[Ano43d] Anonymous. Directory of the Institute of Mathematical Statistics. Annals of Mathematical Statistics, 14(4):448–468, December 1943. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177731369.

Anonymous:1943:NNa

[Ano43e] Anonymous. News and notices. Annals of Mathematical Statistics, 14 (1):96–98, March 1943. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177731498.

Anonymous:1943:NNb

[Ano43f] Anonymous. News and notices. Annals of Mathematical Statistics, 14 (2):199–201, June 1943. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177731463.

Anonymous:1943:NNc

[Ano43g] Anonymous. News and notices. Annals of Mathematical Statistics, 14(3):301–302, September 1943. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177731423.

Anonymous:1943:NNd

[Ano43h] Anonymous. News and notices. Annals of Mathematical Statistics, 14(4):442–443, December 1943. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177731365.

Anonymous:1943:RNY

[Ano43i] Anonymous. Report of the New York meeting of the Institute. Annals of Mathematical Statistics, 14(2):204, June 1943. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177731465.

Anonymous:1943:SCS

[Ano43j] Anonymous. Special courses in statistical quality control. Annals of Mathematical Statistics, 14(2):202–203, June 1943. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177731464.

Anonymous:1944:AP

[Ano44a] Anonymous. Abstracts of papers. Annals of Mathematical Statistics, 15(3):342–343, September 1944. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177731247.

Anonymous:1944:ACL

[Ano44b] Anonymous. Amendments to the constitution and by-laws of the Institute. Annals of Mathematical Statistics, 15(3):340–341, September 1944. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177731246.

Anonymous:1944:L

[Ano44c] Anonymous. By-laws. Annals of Mathematical Statistics, 15(1):116– 118, March 1944. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177731326.

Anonymous:1944:CIM

[Ano44d] Anonymous. Constitution of the Institute of Mathematical Statistics. Annals of Mathematical Statistics, 15(1):114–116, March 1944. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177731325.

Anonymous:1944:NNa

[Ano44e] Anonymous. News and notices. Annals of Mathematical Statistics, 15(1):109–110, March 1944. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177731322.

Anonymous:1944:NNb

[Ano44f] Anonymous. News and notices. Annals of Mathematical Statistics, 15 (2):222–225, June 1944. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177731290.

Anonymous:1944:NNc

[Ano44g] Anonymous. News and notices. Annals of Mathematical Statistics, 15(3):336–337, September 1944. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177731244.

Anonymous:1945:AP

[Ano45a] Anonymous. Abstracts of papers. Annals of Mathematical Statistics, 16(4):402–404, December 1945. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177731070.

Anonymous:1945:L

[Ano45b] Anonymous. By-laws. Annals of Mathematical Statistics, 16(1):114– 116, March 1945. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177731184.

Anonymous:1945:CIM

[Ano45c] Anonymous. Constitution of the Institute of Mathematical Statistics. Annals of Mathematical Statistics, 16(1):112–113, March 1945. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177731183.

Anonymous:1945:NNa

[Ano45d] Anonymous. News and notices. Annals of Mathematical Statistics, 16(1):100–102, March 1945. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177731178.

Anonymous:1945:NNb

[Ano45e] Anonymous. News and notices. Annals of Mathematical Statistics, 16 (2):218–220, June 1945. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177731125.

Anonymous:1945:NNc

[Ano45f] Anonymous. News and notices. Annals of Mathematical Statistics, 16(3):313–316, September 1945. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177731098.

Anonymous:1946:AP

[Ano46a] Anonymous. Abstracts of papers. Annals of Mathematical Statistics, 17(4):500–504, December 1946. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177730891.

Anonymous:1946:ARE

[Ano46b] Anonymous. Annual report of the Editor. Annals of Mathematical Statistics, 17(1):116, March 1946. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177731034.

Anonymous:1946:ARP

[Ano46c] Anonymous. Annual report of the President of the Institute. Annals of Mathematical Statistics, 17(1):95–113, March 1946. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177731032.

Anonymous:1946:ARS

[Ano46d] Anonymous. Annual report of the Secretary-Treasurer of the Institute. Annals of Mathematical Statistics, 17(1):114–115, March 1946. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177731033.

Anonymous:1946:L

[Ano46e] Anonymous. By-laws. Annals of Mathematical Statistics, 17(1):119– 121, March 1946. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177731036.

Anonymous:1946:CIM

[Ano46f] Anonymous. Constitution of the Institute of Mathematical Statistics. Annals of Mathematical Statistics, 17(1):117–118, March 1946. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177731035.

Anonymous:1946:MIM

[Ano46g] Anonymous. Members of the Institute of Mathematical Statistics. Annals of Mathematical Statistics, 17(4):518–556, December 1946. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177730895. See

correction [Ano47f].

Anonymous:1946:NNa

[Ano46h] Anonymous. News and notices. Annals of Mathematical Statistics, 17 (1):89–91, March 1946. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177731030.

Anonymous:1946:NNb

[Ano46i] Anonymous. News and notices. Annals of Mathematical Statistics, 17 (2):249–253, June 1946. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177730989.

Anonymous:1946:NNc

[Ano46j] Anonymous. News and notices. Annals of Mathematical Statistics, 17(3):370–376, September 1946. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177730953.

Anonymous:1946:NNd

[Ano46k] Anonymous. News and notices. Annals of Mathematical Statistics, 17(4):505–513, December 1946. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177730892.

Anonymous:1947:APa

[Ano47a] Anonymous. Abstracts of papers. Annals of Mathematical Statistics, 18(2):298–300, June 1947. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177730452.

Anonymous:1947:APb

[Ano47b] Anonymous. Abstracts of papers. Annals of Mathematical Statistics, 18(3):455–462, September 1947. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177730395.

Anonymous:1947:APc

[Ano47c] Anonymous. Abstracts of papers. Annals of Mathematical Statistics, 18(4):607–611, December 1947. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177730357.

Anonymous:1947:L

[Ano47d] Anonymous. By-laws. Annals of Mathematical Statistics, 18(1):162– 164, March 1947. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177730512.

Anonymous:1947:CIM

[Ano47e] Anonymous. Constitution of the Institute of Mathematical Statistics. Annals of Mathematical Statistics, 18(1):160–161, March 1947. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177730511.

Anonymous:1947:CMI

[Ano47f] Anonymous. Correction: Members of the Institute of Mathematical Statistics. Annals of Mathematical Statistics, 18(3):465, September 1947. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/

1177730397. See [Ano46g].

Anonymous:1947:NNa

[Ano47g] Anonymous. News and notices. Annals of Mathematical Statistics, 18(1):145–148, March 1947. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177730506.

Anonymous:1947:NNb

[Ano47h] Anonymous. News and notices. Annals of Mathematical Statistics, 18 (2):301–305, June 1947. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177730453.

Anonymous:1947:NNc

[Ano47i] Anonymous. News and notices. Annals of Mathematical Statistics, 18(3):463–464, September 1947. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177730396.

Anonymous:1947:NNd

[Ano47j] Anonymous. News and notices. Annals of Mathematical Statistics, 18(4):612–615, December 1947. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177730358.

Anonymous:1948:APa

[Ano48a] Anonymous. Abstracts of papers. Annals of Mathematical Statistics, 19(1):116–121, March 1948. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177730300.

Anonymous:1948:APb

[Ano48b] Anonymous. Abstracts of papers. Annals of Mathematical Statistics, 19(3):428–435, September 1948. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177730215.

Anonymous:1948:APc

[Ano48c] Anonymous. Abstracts of papers. Annals of Mathematical Statistics, 19(4):595–602, December 1948. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177730157.

Anonymous:1948:L

[Ano48d] Anonymous. By-laws. Annals of Mathematical Statistics, 19(1):147– 149, March 1948. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177730311.

Anonymous:1948:CLI

[Ano48e] Anonymous. Constitution and by-laws of the Institute of Mathematical Statistics. Annals of Mathematical Statistics, 19(1):145–146, March 1948. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177730310.

Anonymous:1948:NNa

[Ano48f] Anonymous. News and notices. Annals of Mathematical Statistics, 19(1):127–131, March 1948. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177730303.

Anonymous:1948:NNb

[Ano48g] Anonymous. News and notices. Annals of Mathematical Statistics, 19 (2):285–289, June 1948. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177730258.

Anonymous:1948:NNc

[Ano48h] Anonymous. News and notices. Annals of Mathematical Statistics, 19(3):438–443, September 1948. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177730217.

Anonymous:1948:NNd

[Ano48i] Anonymous. News and notices. Annals of Mathematical Statistics, 19(4):603–606, December 1948. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177730158.

Anonymous:1948:RBMa

[Ano48j] Anonymous. Report on the Berkeley meeting of the Institute. Annals of Mathematical Statistics, 19(1):132–133, March 1948. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177730304.

Anonymous:1948:RBMb

[Ano48k] Anonymous. Report on the Berkeley meeting of the Institute. Annals of Mathematical Statistics, 19(3):444–446, September 1948. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177730218.

Anonymous:1948:RCM

[Ano48l] Anonymous. Report on the Chicago meeting of the Institute. Annals of Mathematical Statistics, 19(1):136, March 1948. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177730306.

Anonymous:1948:TS

[Ano48m] Anonymous. The teaching of statistics. Annals of Mathematical Statistics, 19(1):95–115, March 1948. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177730299.

Anonymous:1949:APa

[Ano49a] Anonymous. Abstracts of papers. Annals of Mathematical Statistics, 20(1):130–143, March 1949. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177730106.

Anonymous:1949:APb

[Ano49b] Anonymous. Abstracts of papers. Annals of Mathematical Statistics, 20(2):317–320, June 1949. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177730045.

Anonymous:1949:APc

[Ano49c] Anonymous. Abstracts of papers. Annals of Mathematical Statistics, 20(3):464–469, September 1949. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177730003.

Anonymous:1949:APd

[Ano49d] Anonymous. Abstracts of papers. Annals of Mathematical Statistics, 20(4):620–624, December 1949. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177729959.

Anonymous:1949:LIM

[Ano49e] Anonymous. By-laws of the Institute of Mathematical Statistics. Annals of Mathematical Statistics, 20(2):330–331, June 1949. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177730050.

Anonymous:1949:CIM

[Ano49f] Anonymous. Constitution of the Institute of Mathematical Statistics. Annals of Mathematical Statistics, 20(2):327–330, June 1949. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177730049.

Anonymous:1949:CNN

[Ano49g] Anonymous. Correction: News and notices. Annals of Mathematical Statistics, 20(2):324, June 1949. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177730047.

Anonymous:1949:NNa

[Ano49h] Anonymous. News and notices. Annals of Mathematical Statistics, 20(1):144–150, March 1949. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177730107.

Anonymous:1949:NNb

[Ano49i] Anonymous. News and notices. Annals of Mathematical Statistics, 20 (2):321–324, June 1949. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177730046.

Anonymous:1949:NNc

[Ano49j] Anonymous. News and notices. Annals of Mathematical Statistics, 20(3):470–474, September 1949. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177730004.

Anonymous:1949:NNd

[Ano49k] Anonymous. News and notices. Annals of Mathematical Statistics, 20(4):624–626, December 1949. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177729960.

Anonymous:1950:APa

[Ano50a] Anonymous. Abstracts of papers. Annals of Mathematical Statistics, 21(1):139–147, March 1950. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177729898.

Anonymous:1950:APb

[Ano50b] Anonymous. Abstracts of papers. Annals of Mathematical Statistics, 21(2):302–315, June 1950. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177729851.

Anonymous:1950:APc

[Ano50c] Anonymous. Abstracts of papers. Annals of Mathematical Statistics, 21(3):461–462, September 1950. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177729809.

Anonymous:1950:APd

[Ano50d] Anonymous. Abstracts of papers. Annals of Mathematical Statistics, 21(4):616–620, December 1950. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177729760.

Anonymous:1950:NNa

[Ano50e] Anonymous. News and notices. Annals of Mathematical Statistics, 21(1):147–151, March 1950. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177729899.

Anonymous:1950:NNb

[Ano50f] Anonymous. News and notices. Annals of Mathematical Statistics, 21 (2):315–317, June 1950. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177729852.

Anonymous:1950:NNc

[Ano50g] Anonymous. News and notices. Annals of Mathematical Statistics, 21(3):462–467, September 1950. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177729810.

Anonymous:1950:NNd

[Ano50h] Anonymous. News and notices. Annals of Mathematical Statistics, 21(4):620–622, December 1950. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177729761.

Anonymous:1950:RNY

[Ano50i] Anonymous. Report of the New York meeting of the Institutes. Annals of Mathematical Statistics, 21(1):151–155, March 1950. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177729900.

Anonymous:1951:APa

[Ano51a] Anonymous. Abstracts of papers. Annals of Mathematical Statistics, 22(1):134–144, March 1951. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177729708.

Anonymous:1951:APb

[Ano51b] Anonymous. Abstracts of papers. Annals of Mathematical Statistics, 22(2):310–317, June 1951. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177729658.

Anonymous:1951:APc

[Ano51c] Anonymous. Abstracts of papers. Annals of Mathematical Statistics, 22(3):483–488, September 1951. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177729603.

Anonymous:1951:APd

[Ano51d] Anonymous. Abstracts of papers. Annals of Mathematical Statistics, 22(4):604–612, December 1951. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177729554.

Anonymous:1951:NNa

[Ano51e] Anonymous. News and notices. Annals of Mathematical Statistics, 22(1):144–149, March 1951. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177729709.

Anonymous:1951:NNb

[Ano51f] Anonymous. News and notices. Annals of Mathematical Statistics, 22 (2):317–322, June 1951. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177729659.

Anonymous:1951:NNc

[Ano51g] Anonymous. News and notices. Annals of Mathematical Statistics, 22(3):488–492, September 1951. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177729604.

Anonymous:1951:NNd

[Ano51h] Anonymous. News and notices. Annals of Mathematical Statistics, 22(4):612–617, December 1951. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177729555.

Anonymous:1951:PRa

[Ano51i] Anonymous. Publications received. Annals of Mathematical Statistics, 22(1):164, March 1951. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177729715.

Anonymous:1951:PRb

[Ano51j] Anonymous. Publications received. Annals of Mathematical Statistics, 22(2):325, June 1951. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177729661.

Anonymous:1951:PRc

[Ano51k] Anonymous. Publications received. Annals of Mathematical Statistics, 22(3):495, September 1951. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177729606.

Anonymous:1951:PRd

[Ano51l] Anonymous. Publications received. Annals of Mathematical Statistics, 22(4):621, December 1951. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177729557.

Anonymous:1952:APa

[Ano52a] Anonymous. Abstracts of papers. Annals of Mathematical Statistics, 23(1):135–148, March 1952. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177729493.

Anonymous:1952:APb

[Ano52b] Anonymous. Abstracts of papers. Annals of Mathematical Statistics, 23(2):299–303, June 1952. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177729452.

Anonymous:1952:APc

[Ano52c] Anonymous. Abstracts of papers. Annals of Mathematical Statistics, 23(3):472–483, September 1952. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177729395.

Anonymous:1952:APd

[Ano52d] Anonymous. Abstracts of papers. Annals of Mathematical Statistics, 23(4):638–647, December 1952. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177729348.

Anonymous:1952:IM

[Ano52e] Anonymous. Institutional members. Annals of Mathematical Statistics, 23(4):655, December 1952. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177729351.

Anonymous:1952:NNa

[Ano52f] Anonymous. News and notices. Annals of Mathematical Statistics, 23(1):148–152, March 1952. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177729494.

Anonymous:1952:NNb

[Ano52g] Anonymous. News and notices. Annals of Mathematical Statistics, 23 (2):304–312, June 1952. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177729453.

Anonymous:1952:NNc

[Ano52h] Anonymous. News and notices. Annals of Mathematical Statistics, 23(3):483–491, September 1952. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177729396.

Anonymous:1952:NNd

[Ano52i] Anonymous. News and notices. Annals of Mathematical Statistics, 23(4):647–655, December 1952. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177729349.

Anonymous:1952:PAW

[Ano52j] Anonymous. The publications of Abraham Wald. Annals of Mathematical Statistics, 23(1):29–33, March 1952. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177729483.

Anonymous:1952:PRa

[Ano52k] Anonymous. Publications received. Annals of Mathematical Statistics, 23(1):167–168, March 1952. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177729501.

Anonymous:1952:PRb

[Ano52l] Anonymous. Publications received. Annals of Mathematical Statistics, 23(2):313, June 1952. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177729454.

Anonymous:1952:PRc

[Ano52m] Anonymous. Publications received. Annals of Mathematical Statistics, 23(3):491, September 1952. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177729397.

Anonymous:1952:PRd

[Ano52n] Anonymous. Publications received. Annals of Mathematical Statistics, 23(4):655, December 1952. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177729350.

Anonymous:1953:APa

[Ano53a] Anonymous. Abstracts of papers. Annals of Mathematical Statistics, 24(1):136–147, March 1953. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177729098.

Anonymous:1953:APb

[Ano53b] Anonymous. Abstracts of papers. Annals of Mathematical Statistics, 24(3):490–499, September 1953. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177728990.

Anonymous:1953:APc

[Ano53c] Anonymous. Abstracts of papers. Annals of Mathematical Statistics, 24(4):680–693, December 1953. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177728927.

Anonymous:1953:IM

[Ano53d] Anonymous. Institutional members. Annals of Mathematical Statistics, 24(4):708, December 1953. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177728931.

Anonymous:1953:MBM

[Ano53e] Anonymous. Minutes of the business meeting, Kingston, Ontario, September 3, 1953. Annals of Mathematical Statistics, 24(4):705– 707, December 1953. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177728929.

Anonymous:1953:NNa

[Ano53f] Anonymous. News and notices. Annals of Mathematical Statistics, 24(1):147–165, March 1953. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177729099.

Anonymous:1953:NNb

[Ano53g] Anonymous. News and notices. Annals of Mathematical Statistics, 24 (2):313–318, June 1953. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177729041.

Anonymous:1953:NNc

[Ano53h] Anonymous. News and notices. Annals of Mathematical Statistics, 24(3):500–510, September 1953. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177728991.

Anonymous:1953:NNd

[Ano53i] Anonymous. News and notices. Annals of Mathematical Statistics, 24(4):694–705, December 1953. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177728928.

Anonymous:1953:PRa

[Ano53j] Anonymous. Publications received. Annals of Mathematical Statistics, 24(1):165, March 1953. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177729100.

Anonymous:1953:PRb

[Ano53k] Anonymous. Publications received. Annals of Mathematical Statistics, 24(2):318, June 1953. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177729042.

Anonymous:1953:PRc

[Ano53l] Anonymous. Publications received. Annals of Mathematical Statistics, 24(3):511, September 1953. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177728992.

Anonymous:1953:PRd

[Ano53m] Anonymous. Publications received. Annals of Mathematical Statistics, 24(4):708, December 1953. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177728930.

Anonymous:1954:APa

[Ano54a] Anonymous. Abstracts of papers. Annals of Mathematical Statistics, 25(1):171–178, March 1954. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177728861.

Anonymous:1954:APb

[Ano54b] Anonymous. Abstracts of papers. Annals of Mathematical Statistics, 25(2):408–414, June 1954. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177728802.

Anonymous:1954:APc

[Ano54c] Anonymous. Abstracts of papers. Annals of Mathematical Statistics, 25(3):616–620, September 1954. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177728734.

Anonymous:1954:APd

[Ano54d] Anonymous. Abstracts of papers. Annals of Mathematical Statistics, 25(4):806–814, December 1954. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177728673.

Anonymous:1954:A

[Ano54e] Anonymous. Addendum. Annals of Mathematical Statistics, 25(2): 394, June 1954. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177728797.

Anonymous:1954:IM

[Ano54f] Anonymous. Institutional members. Annals of Mathematical Statistics, 25(4):826, December 1954. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177728677.

Anonymous:1954:MMM

[Ano54g] Anonymous. Minutes of membership meeting, September 12, 1954. Annals of Mathematical Statistics, 25(4):824–826, December 1954. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177728675.

Anonymous:1954:NNa

[Ano54h] Anonymous. News and notices. Annals of Mathematical Statistics, 25(1):178–201, March 1954. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177728862.

Anonymous:1954:NNb

[Ano54i] Anonymous. News and notices. Annals of Mathematical Statistics, 25 (2):414–421, June 1954. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177728803.

Anonymous:1954:NNc

[Ano54j] Anonymous. News and notices. Annals of Mathematical Statistics, 25(3):620–630, September 1954. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177728735.

Anonymous:1954:NNd

[Ano54k] Anonymous. News and notices. Annals of Mathematical Statistics, 25(4):814–824, December 1954. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177728674.

Anonymous:1954:PRa

[Ano54l] Anonymous. Publications received. Annals of Mathematical Statistics, 25(1):201, March 1954. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177728863.

Anonymous:1954:PRb

[Ano54m] Anonymous. Publications received. Annals of Mathematical Statistics, 25(2):421, June 1954. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177728804.

Anonymous:1954:PRc

[Ano54n] Anonymous. Publications received. Annals of Mathematical Statistics, 25(3):630, September 1954. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177728736.

Anonymous:1954:PRd

[Ano54o] Anonymous. Publications received. Annals of Mathematical Statistics, 25(4):826, December 1954. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177728676.

Anonymous:1955:APa

[Ano55a] Anonymous. Abstracts of papers. Annals of Mathematical Statistics, 26(1):150–162, March 1955. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177728610.

Anonymous:1955:APb

[Ano55b] Anonymous. Abstracts of papers. Annals of Mathematical Statistics, 26(3):536–542, September 1955. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177728502.

Anonymous:1955:APc

[Ano55c] Anonymous. Abstracts of papers. Annals of Mathematical Statistics, 26(4):767–775, December 1955. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177728436.

Anonymous:1955:IM

[Ano55d] Anonymous. Institutional members. Annals of Mathematical Statistics, 26(4):785, December 1955. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177728439.

Anonymous:1955:NNa

[Ano55e] Anonymous. News and notices. Annals of Mathematical Statistics, 26(1):163–188, March 1955. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177728611.

Anonymous:1955:NNb

[Ano55f] Anonymous. News and notices. Annals of Mathematical Statistics, 26 (2):353–360, June 1955. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177728556.

Anonymous:1955:NNc

[Ano55g] Anonymous. News and notices. Annals of Mathematical Statistics, 26(3):543–549, September 1955. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177728503.

Anonymous:1955:NNd

[Ano55h] Anonymous. News and notices. Annals of Mathematical Statistics, 26(4):776–785, December 1955. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177728437.

Anonymous:1955:PRa

[Ano55i] Anonymous. Publications received. Annals of Mathematical Statistics, 26(1):188, March 1955. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177728612.

Anonymous:1955:PRb

[Ano55j] Anonymous. Publications received. Annals of Mathematical Statistics, 26(2):364, June 1955. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177728558.

Anonymous:1955:PRc

[Ano55k] Anonymous. Publications received. Annals of Mathematical Statistics, 26(3):549, September 1955. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177728504.

Anonymous:1955:PRd

[Ano55l] Anonymous. Publications received. Annals of Mathematical Statistics, 26(4):785, December 1955. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177728438.

Anonymous:1956:APa

[Ano56a] Anonymous. Abstracts of papers. Annals of Mathematical Statistics, 27(1):204–223, March 1956. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177728361.

Anonymous:1956:APb

[Ano56b] Anonymous. Abstracts of papers. Annals of Mathematical Statistics, 27(2):543–552, June 1956. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177728280.

Anonymous:1956:APc

[Ano56c] Anonymous. Abstracts of papers. Annals of Mathematical Statistics, 27(3):860–874, September 1956. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177728200.

Anonymous:1956:APd

[Ano56d] Anonymous. Abstracts of papers. Annals of Mathematical Statistics, 27(4):1184–1186, December 1956. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177728090.

Anonymous:1956:IM

[Ano56e] Anonymous. Institutional members. Annals of Mathematical Statistics, 27(4):1212, December 1956. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177728095.

Anonymous:1956:NNa

[Ano56f] Anonymous. News and notices. Annals of Mathematical Statistics, 27(1):224–249, March 1956. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177728362.

Anonymous:1956:NNb

[Ano56g] Anonymous. News and notices. Annals of Mathematical Statistics, 27 (2):552–566, June 1956. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177728281.

Anonymous:1956:NNc

[Ano56h] Anonymous. News and notices. Annals of Mathematical Statistics, 27(3):874–878, September 1956. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177728201.

Anonymous:1956:NNd

[Ano56i] Anonymous. News and notices. Annals of Mathematical Statistics, 27 (4):1187–1197, December 1956. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177728092.

Anonymous:1956:PRa

[Ano56j] Anonymous. Publications received. Annals of Mathematical Statistics, 27(1):249, March 1956. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177728363.

Anonymous:1956:PRb

[Ano56k] Anonymous. Publications received. Annals of Mathematical Statistics, 27(2):567, June 1956. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177728282.

Anonymous:1956:PRc

[Ano56l] Anonymous. Publications received. Annals of Mathematical Statistics, 27(3):878, September 1956. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177728202.

Anonymous:1956:PRd

[Ano56m] Anonymous. Publications received. Annals of Mathematical Statistics, 27(4):1212, December 1956. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177728094.

Anonymous:1956:RDM

[Ano56n] Anonymous. Report of the Detroit meeting of the Institute of Mathematical Statistics. Annals of Mathematical Statistics, 27(4):1198– 1211, December 1956. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177728093.

Anonymous:1956:SRM

[Ano56o] Anonymous. Statistical research monographs. Annals of Mathematical Statistics, 27(4):1186, December 1956. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177728091.

Anonymous:1957:APa

[Ano57a] Anonymous. Abstracts of papers. Annals of Mathematical Statistics, 28(2):523–533, June 1957. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177706987.

Anonymous:1957:APb

[Ano57b] Anonymous. Abstracts of papers. Annals of Mathematical Statistics, 28(3):816–818, September 1957. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177706902.

Anonymous:1957:APc

[Ano57c] Anonymous. Abstracts of papers. Annals of Mathematical Statistics, 28(4):1056–1075, December 1957. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177706816.

Anonymous:1957:IM

[Ano57d] Anonymous. Institutional members. Annals of Mathematical Statistics, 28(4):1098, December 1957. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177706819.

Anonymous:1957:NNa

[Ano57e] Anonymous. News and notices. Annals of Mathematical Statistics, 28(1):269–279, March 1957. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177707057.

Anonymous:1957:NNb

[Ano57f] Anonymous. News and notices. Annals of Mathematical Statistics, 28 (2):533–545, June 1957. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177706988.

Anonymous:1957:NNc

[Ano57g] Anonymous. News and notices. Annals of Mathematical Statistics, 28(3):818–822, September 1957. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177706903.

Anonymous:1957:NNd

[Ano57h] Anonymous. News and notices. Annals of Mathematical Statistics, 28 (4):1076–1097, December 1957. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177706817.

Anonymous:1957:PRa

[Ano57i] Anonymous. Publications received. Annals of Mathematical Statistics, 28(1):279, March 1957. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177707058.

Anonymous:1957:PRb

[Ano57j] Anonymous. Publications received. Annals of Mathematical Statistics, 28(2):545, June 1957. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177706989.

Anonymous:1957:PRc

[Ano57k] Anonymous. Publications received. Annals of Mathematical Statistics, 28(3):822, September 1957. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177706904.

Anonymous:1957:PRd

[Ano57l] Anonymous. Publications received. Annals of Mathematical Statistics, 28(4):1097–1098, December 1957. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177706818.

Anonymous:1958:A

[Ano58a] Anonymous. Abstracts. Annals of Mathematical Statistics, 29(1): 325–328, March 1958. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177706739.

Anonymous:1958:APa

[Ano58b] Anonymous. Abstracts of papers. Annals of Mathematical Statistics, 29(2):614–623, June 1958. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177706647.

Anonymous:1958:APb

[Ano58c] Anonymous. Abstracts of papers. Annals of Mathematical Statistics, 29(3):936–939, September 1958. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177706560.

Anonymous:1958:APc

[Ano58d] Anonymous. Abstracts of papers. Annals of Mathematical Statistics, 29(4):1279–1291, December 1958. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177706464.

Anonymous:1958:NNa

[Ano58e] Anonymous. News and notices. Annals of Mathematical Statistics, 29(1):328–350, March 1958. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177706740.

Anonymous:1958:NNb

[Ano58f] Anonymous. News and notices. Annals of Mathematical Statistics, 29 (2):624–631, June 1958. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177706648.

Anonymous:1958:NNc

[Ano58g] Anonymous. News and notices. Annals of Mathematical Statistics, 29(3):939–946, September 1958. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177706561.

Anonymous:1958:NNd

[Ano58h] Anonymous. News and notices. Annals of Mathematical Statistics, 29 (4):1292–1308, December 1958. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177706465.

Anonymous:1958:PRa

[Ano58i] Anonymous. Publications received. Annals of Mathematical Statistics, 29(1):350, March 1958. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177706741.

Anonymous:1958:PRb

[Ano58j] Anonymous. Publications received. Annals of Mathematical Statistics, 29(2):634, June 1958. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177706650.

Anonymous:1958:PRc

[Ano58k] Anonymous. Publications received. Annals of Mathematical Statistics, 29(3):946, September 1958. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177706562.

Anonymous:1958:PRd

[Ano58l] Anonymous. Publications received. Annals of Mathematical Statistics, 29(4):1308, December 1958. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177706466.

Anonymous:1958:RGT

[Ano58m] Anonymous. Report of the Gatlinburg, Tennessee meeting of the Institute. Annals of Mathematical Statistics, 29(2):631–634, June 1958. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177706649.

Anonymous:1959:APa

[Ano59a] Anonymous. Abstracts of papers. Annals of Mathematical Statistics, 30(1):255–257, March 1959. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177706388.

Anonymous:1959:APb

[Ano59b] Anonymous. Abstracts of papers. Annals of Mathematical Statistics, 30(2):611–625, June 1959. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177706281.

Anonymous:1959:APc

[Ano59c] Anonymous. Abstracts of papers. Annals of Mathematical Statistics, 30(3):835–837, September 1959. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177706214.

Anonymous:1959:APd

[Ano59d] Anonymous. Abstracts of papers. Annals of Mathematical Statistics, 30(4):1271–1278, December 1959. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177706123.

Anonymous:1959:NNa

[Ano59e] Anonymous. News and notices. Annals of Mathematical Statistics, 30(1):257–269, March 1959. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177706389.

Anonymous:1959:NNb

[Ano59f] Anonymous. News and notices. Annals of Mathematical Statistics, 30 (2):626–639, June 1959. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177706282.

Anonymous:1959:NNc

[Ano59g] Anonymous. News and notices. Annals of Mathematical Statistics, 30(3):838–844, September 1959. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177706215.

Anonymous:1959:NNd

[Ano59h] Anonymous. News and notices. Annals of Mathematical Statistics, 30 (4):1279–1286, December 1959. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177706124.

Anonymous:1959:PRa

[Ano59i] Anonymous. Publications received. Annals of Mathematical Statistics, 30(1):269, March 1959. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177706390.

Anonymous:1959:PRb

[Ano59j] Anonymous. Publications received. Annals of Mathematical Statistics, 30(2):639–640, June 1959. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177706283.

Anonymous:1959:PRc

[Ano59k] Anonymous. Publications received. Annals of Mathematical Statistics, 30(3):844, September 1959. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177706216.

Anonymous:1959:PRd

[Ano59l] Anonymous. Publications received. Annals of Mathematical Statistics, 30(4):1286, December 1959. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177706125.

Anonymous:1960:APa

[Ano60a] Anonymous. Abstracts of papers. Annals of Mathematical Statistics, 31(1):230–246, March 1960. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177706005.

Anonymous:1960:APb

[Ano60b] Anonymous. Abstracts of papers. Annals of Mathematical Statistics, 31(2):524–539, June 1960. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177705923.

Anonymous:1960:APc

[Ano60c] Anonymous. Abstracts of papers. Annals of Mathematical Statistics, 31(3):811–823, September 1960. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177705814.

Anonymous:1960:APd

[Ano60d] Anonymous. Abstracts of papers. Annals of Mathematical Statistics, 31(4):1215–1217, December 1960. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177705696.

Anonymous:1960:NNa

[Ano60e] Anonymous. News and notices. Annals of Mathematical Statistics, 31(1):247–267, March 1960. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177706006.

Anonymous:1960:NNb

[Ano60f] Anonymous. News and notices. Annals of Mathematical Statistics, 31 (2):540–556, June 1960. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177705924.

Anonymous:1960:NNc

[Ano60g] Anonymous. News and notices. Annals of Mathematical Statistics, 31(3):824–833, September 1960. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177705815.

Anonymous:1960:NNd

[Ano60h] Anonymous. News and notices. Annals of Mathematical Statistics, 31 (4):1219–1254, December 1960. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177705697.

Anonymous:1960:NUN

[Ano60i] Anonymous. Note [on the uniqueness and nonuniqueness of the triangular association schemes]. Annals of Mathematical Statistics, 31 (2):497, June 1960. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177705915.

Anonymous:1960:PRa

[Ano60j] Anonymous. Publications received. Annals of Mathematical Statistics, 31(1):268, March 1960. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177706007.

Anonymous:1960:PRb

[Ano60k] Anonymous. Publications received. Annals of Mathematical Statistics, 31(2):556, June 1960. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177705925.

Anonymous:1960:PRc

[Ano60l] Anonymous. Publications received. Annals of Mathematical Statistics, 31(3):833–834, September 1960. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177705816.

Anonymous:1960:PRd

[Ano60m] Anonymous. Publications received. Annals of Mathematical Statistics, 31(4):1254, December 1960. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177705698.

Anonymous:1961:APa

[Ano61a] Anonymous. Abstracts of papers. Annals of Mathematical Statistics, 32(1):340–342, March 1961. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177705165.

Anonymous:1961:APb

[Ano61b] Anonymous. Abstracts of papers. Annals of Mathematical Statistics, 32(2):621–639, June 1961. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/ euclid.aoms/1177705077. See correction [Sch61a].

Anonymous:1961:APc

[Ano61c] Anonymous. Abstracts of papers. Annals of Mathematical Statistics, 32(3):914–926, September 1961. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177704990.

Anonymous:1961:APd

[Ano61d] Anonymous. Abstracts of papers. Annals of Mathematical Statistics, 32(4):1346–1352, December 1961. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177704881.

Anonymous:1961:CE

[Ano61e] Anonymous. Change of editorship. Annals of Mathematical Statistics, 32(1):343, March 1961. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177705166.

Anonymous:1961:NNa

[Ano61f] Anonymous. News and notices. Annals of Mathematical Statistics, 32(1):344–355, March 1961. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177705167.

Anonymous:1961:NNb

[Ano61g] Anonymous. News and notices. Annals of Mathematical Statistics, 32 (2):640–650, June 1961. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177705078.

Anonymous:1961:NNc

[Ano61h] Anonymous. News and notices. Annals of Mathematical Statistics, 32(3):927–950, September 1961. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177704992.

Anonymous:1961:NNd

[Ano61i] Anonymous. News and notices. Annals of Mathematical Statistics, 32 (4):1353–1365, December 1961. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177704882.

Anonymous:1961:PRa

[Ano61j] Anonymous. Publications received. Annals of Mathematical Statistics, 32(1):355, March 1961. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177705168.

Anonymous:1961:PRb

[Ano61k] Anonymous. Publications received. Annals of Mathematical Statistics, 32(2):650, June 1961. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177705079.

Anonymous:1961:PRc

[Ano61l] Anonymous. Publications received. Annals of Mathematical Statistics, 32(3):950, September 1961. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177704993.

Anonymous:1961:PRd

[Ano61m] Anonymous. Publications received. Annals of Mathematical Statistics, 32(4):1365, December 1961. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177704883.

Anonymous:1962:APa

[Ano62a] Anonymous. Abstracts of papers. Annals of Mathematical Statistics, 33(1):295–312, March 1962. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177704734.

Anonymous:1962:APb

[Ano62b] Anonymous. Abstracts of papers. Annals of Mathematical Statistics, 33(2):814–833, June 1962. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177704606. See correction [Gup62].

Anonymous:1962:APc

[Ano62c] Anonymous. Abstracts of papers. Annals of Mathematical Statistics, 33(3):1206–1212, September 1962. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177704486.

Anonymous:1962:APd

[Ano62d] Anonymous. Abstracts of papers. Annals of Mathematical Statistics, 33(4):1480–1506, December 1962. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177704386.

Anonymous:1962:CRSa

[Ano62e] Anonymous. Change in reference style. Annals of Mathematical Statistics, 33(1):313–314, March 1962. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177704735.

Anonymous:1962:CRSb

[Ano62f] Anonymous. Change in reference style. Annals of Mathematical Statistics, 33(2):834–835, June 1962. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177704608.

Anonymous:1962:CRSc

[Ano62g] Anonymous. Change in reference style. Annals of Mathematical Statistics, 33(3):1213–1214, September 1962. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177704487.

Anonymous:1962:CA

[Ano62h] Anonymous. Correction to abstract. Annals of Mathematical Statistics, 33(2):833, June 1962. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177704607.

Anonymous:1962:NNa

[Ano62i] Anonymous. News and notices. Annals of Mathematical Statistics, 33(1):315–324, March 1962. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177704736.

Anonymous:1962:NNb

[Ano62j] Anonymous. News and notices. Annals of Mathematical Statistics, 33 (2):836–846, June 1962. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177704609.

Anonymous:1962:NNc

[Ano62k] Anonymous. News and notices. Annals of Mathematical Statistics, 33 (3):1215–1226, September 1962. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177704488.

Anonymous:1962:NNd

[Ano62l] Anonymous. News and notices. Annals of Mathematical Statistics, 33 (4):1507–1527, December 1962. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177704387.

Anonymous:1962:PR

[Ano62m] Anonymous. Publications received. Annals of Mathematical Statistics, 33(1):324, March 1962. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177704737.

Anonymous:1963:APa

[Ano63a] Anonymous. Abstracts of papers. Annals of Mathematical Statistics, 34(1):355–358, March 1963. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177704283.

Anonymous:1963:APb

[Ano63b] Anonymous. Abstracts of papers. Annals of Mathematical Statistics, 34(2):678–695, June 1963. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177704185.

Anonymous:1963:APc

[Ano63c] Anonymous. Abstracts of papers. Annals of Mathematical Statistics, 34(3):1112–1132, September 1963. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177704040.

Anonymous:1963:APd

[Ano63d] Anonymous. Abstracts of papers. Annals of Mathematical Statistics, 34(4):1615–1629, December 1963. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177703901.

Anonymous:1963:NNa

[Ano63e] Anonymous. News and notices. Annals of Mathematical Statistics, 34(1):359–374, March 1963. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177704284.

Anonymous:1963:NNb

[Ano63f] Anonymous. News and notices. Annals of Mathematical Statistics, 34 (2):696–704, June 1963. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177704186.

Anonymous:1963:NNc

[Ano63g] Anonymous. News and notices. Annals of Mathematical Statistics, 34 (3):1133–1146, September 1963. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177704041.

Anonymous:1963:NNd

[Ano63h] Anonymous. News and notices. Annals of Mathematical Statistics, 34 (4):1630–1659, December 1963. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177703902.

Anonymous:1963:PRa

[Ano63i] Anonymous. Publications received. Annals of Mathematical Statistics, 34(1):374, March 1963. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177704285.

Anonymous:1963:PRb

[Ano63j] Anonymous. Publications received. Annals of Mathematical Statistics, 34(3):1146, September 1963. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177704042.

Anonymous:1963:PRc

[Ano63k] Anonymous. Publications received. Annals of Mathematical Statistics, 34(4):1659, December 1963. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177703903.

Anonymous:1964:APa

[Ano64a] Anonymous. Abstracts of papers. Annals of Mathematical Statistics, 35(1):454–462, March 1964. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177703778.

Anonymous:1964:APb

[Ano64b] Anonymous. Abstracts of papers. Annals of Mathematical Statistics, 35(2):926–940, June 1964. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177703612.

Anonymous:1964:APc

[Ano64c] Anonymous. Abstracts of papers. Annals of Mathematical Statistics, 35(3):1389–1408, September 1964. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177703306.

Anonymous:1964:APd

[Ano64d] Anonymous. Abstracts of papers. Annals of Mathematical Statistics, 35(4):1834–1852, December 1964. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid. org/euclid.aoms/1177700412. See corrections [AJ65, Sri65a].

Anonymous:1964:NNa

[Ano64e] Anonymous. News and notices. Annals of Mathematical Statistics, 35(1):463–474, March 1964. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177703779.

Anonymous:1964:NNb

[Ano64f] Anonymous. News and notices. Annals of Mathematical Statistics, 35 (2):941–948, June 1964. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177703613.

Anonymous:1964:NNc

[Ano64g] Anonymous. News and notices. Annals of Mathematical Statistics, 35 (3):1409–1417, September 1964. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177703307.

Anonymous:1964:NNd

[Ano64h] Anonymous. News and notices. Annals of Mathematical Statistics, 35 (4):1853–1880, December 1964. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177700413.

Anonymous:1964:PRa

[Ano64i] Anonymous. Publications received. Annals of Mathematical Statistics, 35(1):474, March 1964. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177703780.

Anonymous:1964:PRb

[Ano64j] Anonymous. Publications received. Annals of Mathematical Statistics, 35(3):1417, September 1964. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177703308.

Anonymous:1965:APa

[Ano65a] Anonymous. Abstracts of papers. Annals of Mathematical Statistics, 36(1):351–365, February 1965. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177700304. See correction [KA65].

Anonymous:1965:APb

[Ano65b] Anonymous. Abstracts of papers. Annals of Mathematical Statistics, 36(2):726–736, April 1965. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177700188. See correction [Kum65].

Anonymous:1965:APc

[Ano65c] Anonymous. Abstracts of papers. Annals of Mathematical Statistics, 36(3):1073–1087, June 1965. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177700089.

Anonymous:1965:APd

[Ano65d] Anonymous. Abstracts of papers. Annals of Mathematical Statistics, 36(4):1321–1329, August 1965. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177700011.

Anonymous:1965:APe

[Ano65e] Anonymous. Abstracts of papers. Annals of Mathematical Statistics, 36(5):1592–1615, October 1965. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177699926.

Anonymous:1965:APf

[Ano65f] Anonymous. Abstracts of papers. Annals of Mathematical Statistics, 36(6):1905–1907, December 1965. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177699827.

Anonymous:1965:NNa

[Ano65g] Anonymous. News and notices. Annals of Mathematical Statistics, 36(1):366–368, February 1965. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177700305.

Anonymous:1965:NNb

[Ano65h] Anonymous. News and notices. Annals of Mathematical Statistics, 36 (2):737–744, April 1965. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177700189.

Anonymous:1965:NNc

[Ano65i] Anonymous. News and notices. Annals of Mathematical Statistics, 36(3):1088–1094, June 1965. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177700090.

Anonymous:1965:NNd

[Ano65j] Anonymous. News and notices. Annals of Mathematical Statistics, 36(4):1330–1337, August 1965. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177700012.

Anonymous:1965:NNe

[Ano65k] Anonymous. News and notices. Annals of Mathematical Statistics, 36(5):1616–1624, October 1965. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177699927.

Anonymous:1965:NNf

[Ano65l] Anonymous. News and notices. Annals of Mathematical Statistics, 36 (6):1908–1926, December 1965. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177699828.

Anonymous:1965:PW

[Ano65m] Anonymous. The publications of S. S. Wilks. Annals of Mathematical Statistics, 36(1):24–27, February 1965. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177700268.

Anonymous:1965:PRa

[Ano65n] Anonymous. Publications received. Annals of Mathematical Statistics, 36(1):368, February 1965. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177700306.

Anonymous:1965:PRb

[Ano65o] Anonymous. Publications received. Annals of Mathematical Statistics, 36(2):744–745, April 1965. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177700190.

Anonymous:1965:PRc

[Ano65p] Anonymous. Publications received. Annals of Mathematical Statistics, 36(3):1095, June 1965. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177700091.

Anonymous:1965:PRd

[Ano65q] Anonymous. Publications received. Annals of Mathematical Statistics, 36(4):1338, August 1965. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177700013.

Anonymous:1965:PRe

[Ano65r] Anonymous. Publications received. Annals of Mathematical Statistics, 36(5):1624–1625, October 1965. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177699928.

Anonymous:1965:PRf

[Ano65s] Anonymous. Publications received. Annals of Mathematical Statistics, 36(6):1926–1927, December 1965. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177699829.

Anonymous:1966:A

[Ano66a] Anonymous. Abstracts. Annals of Mathematical Statistics, 37(6): 1858–1866, December 1966. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177699186.

Anonymous:1966:APa

[Ano66b] Anonymous. Abstracts of papers. Annals of Mathematical Statistics, 37(1):313–315, February 1966. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177699629.

Anonymous:1966:APb

[Ano66c] Anonymous. Abstracts of papers. Annals of Mathematical Statistics, 37(2):539–555, April 1966. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177699543.

Anonymous:1966:APc

[Ano66d] Anonymous. Abstracts of papers. Annals of Mathematical Statistics, 37(3):759–775, June 1966. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177699482. See correction [Sin66].

Anonymous:1966:APd

[Ano66e] Anonymous. Abstracts of papers. Annals of Mathematical Statistics, 37(4):1066–1078, August 1966. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177699392.

Anonymous:1966:APe

[Ano66f] Anonymous. Abstracts of papers. Annals of Mathematical Statistics, 37(5):1414–1433, October 1966. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177699290.

Anonymous:1966:NNa

[Ano66g] Anonymous. News and notices. Annals of Mathematical Statistics, 37(1):316–318, February 1966. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177699630.

Anonymous:1966:NNb

[Ano66h] Anonymous. News and notices. Annals of Mathematical Statistics, 37 (2):556–564, April 1966. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177699544.

Anonymous:1966:NNc

[Ano66i] Anonymous. News and notices. Annals of Mathematical Statistics, 37 (3):776–781, June 1966. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177699483.

Anonymous:1966:NNd

[Ano66j] Anonymous. News and notices. Annals of Mathematical Statistics, 37(4):1079–1085, August 1966. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177699393.

Anonymous:1966:NNe

[Ano66k] Anonymous. News and notices. Annals of Mathematical Statistics, 37(5):1434–1436, October 1966. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177699291.

Anonymous:1966:NNf

[Ano66l] Anonymous. News and notices. Annals of Mathematical Statistics, 37 (6):1867–1887, December 1966. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177699187.

Anonymous:1966:PRa

[Ano66m] Anonymous. Publications received. Annals of Mathematical Statistics, 37(1):318–319, February 1966. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177699631.

Anonymous:1966:PRb

[Ano66n] Anonymous. Publications received. Annals of Mathematical Statistics, 37(2):564–565, April 1966. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177699545.

Anonymous:1966:PRc

[Ano66o] Anonymous. Publications received. Annals of Mathematical Statistics, 37(3):781, June 1966. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177699484.

Anonymous:1966:PRd

[Ano66p] Anonymous. Publications received. Annals of Mathematical Statistics, 37(4):1085–1086, August 1966. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177699394.

Anonymous:1966:PRe

[Ano66q] Anonymous. Publications received. Annals of Mathematical Statistics, 37(5):1436–1437, October 1966. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177699292.

Anonymous:1966:PRf

[Ano66r] Anonymous. Publications received. Annals of Mathematical Statistics, 37(6):1887–1888, December 1966. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177699188.

Anonymous:1967:Aa

[Ano67a] Anonymous. Abstracts. Annals of Mathematical Statistics, 38(2): 634–643, April 1967. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177698993.

Anonymous:1967:Ab

[Ano67b] Anonymous. Abstracts. Annals of Mathematical Statistics, 38(3): 953–971, June 1967. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177698897.

Anonymous:1967:Ac

[Ano67c] Anonymous. Abstracts. Annals of Mathematical Statistics, 38(4): 1310–1316, August 1967. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177698809.

Anonymous:1967:APa

[Ano67d] Anonymous. Abstracts of papers. Annals of Mathematical Statistics, 38(1):297–299, February 1967. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177699090.

Anonymous:1967:APb

[Ano67e] Anonymous. Abstracts of papers. Annals of Mathematical Statistics, 38(5):1595–1598, October 1967. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177698718.

Anonymous:1967:APc

[Ano67f] Anonymous. Abstracts of papers. Annals of Mathematical Statistics, 38(6):1932–1940, December 1967. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177698633.

Anonymous:1967:NNa

[Ano67g] Anonymous. News and notices. Annals of Mathematical Statistics, 38(1):300–301, February 1967. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177699091.

Anonymous:1967:NNb

[Ano67h] Anonymous. News and notices. Annals of Mathematical Statistics, 38 (2):644–657, April 1967. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177698994.

Anonymous:1967:NNc

[Ano67i] Anonymous. News and notices. Annals of Mathematical Statistics, 38 (3):972–976, June 1967. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177698898.

Anonymous:1967:NNd

[Ano67j] Anonymous. News and notices. Annals of Mathematical Statistics, 38(4):1317–1322, August 1967. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177698810.

Anonymous:1967:NNe

[Ano67k] Anonymous. News and notices. Annals of Mathematical Statistics, 38(5):1599–1604, October 1967. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177698719.

Anonymous:1967:NNf

[Ano67l] Anonymous. News and notices. Annals of Mathematical Statistics, 38 (6):1941–1947, December 1967. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177698634.

Anonymous:1967:PRa

[Ano67m] Anonymous. Publications received. Annals of Mathematical Statistics, 38(1):301–302, February 1967. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177699092.

Anonymous:1967:PRb

[Ano67n] Anonymous. Publications received. Annals of Mathematical Statistics, 38(2):658, April 1967. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177698995.

Anonymous:1967:PRc

[Ano67o] Anonymous. Publications received. Annals of Mathematical Statistics, 38(3):976, June 1967. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177698899.

Anonymous:1967:PRd

[Ano67p] Anonymous. Publications received. Annals of Mathematical Statistics, 38(4):1322, August 1967. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177698811.

Anonymous:1967:PRe

[Ano67q] Anonymous. Publications received. Annals of Mathematical Statistics, 38(5):1604, October 1967. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177698720.

Anonymous:1967:PRf

[Ano67r] Anonymous. Publications received. Annals of Mathematical Statistics, 38(6):1947–1948, December 1967. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177698635.

Anonymous:1968:APa

[Ano68a] Anonymous. Abstracts of papers. Annals of Mathematical Statistics, 39(1):294–312, February 1968. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177698541.

Anonymous:1968:APb

[Ano68b] Anonymous. Abstracts of papers. Annals of Mathematical Statistics, 39(2):687–708, April 1968. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177698433.

Anonymous:1968:APc

[Ano68c] Anonymous. Abstracts of papers. Annals of Mathematical Statistics, 39(3):1086–1097, June 1968. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177698343.

Anonymous:1968:APd

[Ano68d] Anonymous. Abstracts of papers. Annals of Mathematical Statistics, 39(4):1360–1372, August 1968. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177698267.

Anonymous:1968:APe

[Ano68e] Anonymous. Abstracts of papers. Annals of Mathematical Statistics, 39(5):1770–1795, October 1968. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177698168.

Anonymous:1968:APf

[Ano68f] Anonymous. Abstracts of papers. Annals of Mathematical Statistics, 39(6):2171–2173, December 1968. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177698058. See correction [Sri69].

Anonymous:1968:NNa

[Ano68g] Anonymous. News and notices. Annals of Mathematical Statistics, 39(1):313–322, February 1968. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177698542.

Anonymous:1968:NNb

[Ano68h] Anonymous. News and notices. Annals of Mathematical Statistics, 39 (2):709–724, April 1968. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177698434.

Anonymous:1968:NNc

[Ano68i] Anonymous. News and notices. Annals of Mathematical Statistics, 39(3):1098–1104, June 1968. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177698344.

Anonymous:1968:NNd

[Ano68j] Anonymous. News and notices. Annals of Mathematical Statistics, 39(4):1373–1378, August 1968. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177698268.

Anonymous:1968:NNe

[Ano68k] Anonymous. News and notices. Annals of Mathematical Statistics, 39(5):1796–1798, October 1968. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177698169.

Anonymous:1968:NNf

[Ano68l] Anonymous. News and notices. Annals of Mathematical Statistics, 39 (6):2174–2187, December 1968. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177698059.

Anonymous:1968:PRa

[Ano68m] Anonymous. Publications received. Annals of Mathematical Statistics, 39(1):322–323, February 1968. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177698543.

Anonymous:1968:PRb

[Ano68n] Anonymous. Publications received. Annals of Mathematical Statistics, 39(2):724–725, April 1968. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177698435.

Anonymous:1968:PRc

[Ano68o] Anonymous. Publications received. Annals of Mathematical Statistics, 39(3):1104–1105, June 1968. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177698345.

Anonymous:1968:PRd

[Ano68p] Anonymous. Publications received. Annals of Mathematical Statistics, 39(4):1378–1380, August 1968. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177698269.

Anonymous:1968:PRe

[Ano68q] Anonymous. Publications received. Annals of Mathematical Statistics, 39(5):1798–1799, October 1968. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177698170.

Anonymous:1968:PRf

[Ano68r] Anonymous. Publications received. Annals of Mathematical Statistics, 39(6):2187–2188, December 1968. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177698060.

Anonymous:1969:APa

[Ano69a] Anonymous. Abstracts of papers. Annals of Mathematical Statistics, 40(1):329–332, February 1969. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177697835.

Anonymous:1969:APb

[Ano69b] Anonymous. Abstracts of papers. Annals of Mathematical Statistics, 40(2):720–727, April 1969. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177697751.

Anonymous:1969:APc

[Ano69c] Anonymous. Abstracts of papers. Annals of Mathematical Statistics, 40(3):1142–1161, June 1969. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177697631.

Anonymous:1969:APd

[Ano69d] Anonymous. Abstracts of papers. Annals of Mathematical Statistics, 40(4):1511–1518, August 1969. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177697529.

Anonymous:1969:APe

[Ano69e] Anonymous. Abstracts of papers. Annals of Mathematical Statistics, 40(5):1856–1883, October 1969. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177697405.

Anonymous:1969:APf

[Ano69f] Anonymous. Abstracts of papers. Annals of Mathematical Statistics, 40(6):2216–2226, December 1969. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177697305.

Anonymous:1969:CNN

[Ano69g] Anonymous. Correction: News and notes. Annals of Mathematical Statistics, 40(3):1165, June 1969. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177697633.

Anonymous:1969:NNa

[Ano69h] Anonymous. News and notices. Annals of Mathematical Statistics, 40(1):333–337, February 1969. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177697836.

Anonymous:1969:NNb

[Ano69i] Anonymous. News and notices. Annals of Mathematical Statistics, 40 (2):728–740, April 1969. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177697752.

Anonymous:1969:NNc

[Ano69j] Anonymous. News and notices. Annals of Mathematical Statistics, 40(3):1162–1164, June 1969. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177697632.

Anonymous:1969:NNd

[Ano69k] Anonymous. News and notices. Annals of Mathematical Statistics, 40(4):1519–1520, August 1969. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177697530.

Anonymous:1969:NNe

[Ano69l] Anonymous. News and notices. Annals of Mathematical Statistics, 40(5):1884–1887, October 1969. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177697406.

Anonymous:1969:NNf

[Ano69m] Anonymous. News and notices. Annals of Mathematical Statistics, 40 (6):2227–2236, December 1969. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177697306.

Anonymous:1969:PRa

[Ano69n] Anonymous. Publications received. Annals of Mathematical Statistics, 40(1):337, February 1969. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177697837.

Anonymous:1969:PRb

[Ano69o] Anonymous. Publications received. Annals of Mathematical Statistics, 40(2):741, April 1969. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177697753.

Anonymous:1969:PRc

[Ano69p] Anonymous. Publications received. Annals of Mathematical Statistics, 40(3):1166, June 1969. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177697634.

Anonymous:1969:PRd

[Ano69q] Anonymous. Publications received. Annals of Mathematical Statistics, 40(4):1520–1521, August 1969. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177697531.

Anonymous:1969:PRe

[Ano69r] Anonymous. Publications received. Annals of Mathematical Statistics, 40(5):1887, October 1969. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177697407.

Anonymous:1969:PRf

[Ano69s] Anonymous. Publications received. Annals of Mathematical Statistics, 40(6):2237, December 1969. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177697307.

Anonymous:1970:A

[Ano70a] Anonymous. Abstracts. Annals of Mathematical Statistics, 41(5): 1783–1809, October 1970. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177696827.

Anonymous:1970:APa

[Ano70b] Anonymous. Abstracts of papers. Annals of Mathematical Statistics, 41(1):328–332, February 1970. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177697216.

Anonymous:1970:APb

[Ano70c] Anonymous. Abstracts of papers. Annals of Mathematical Statistics, 41(2):751–757, April 1970. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177697132.

Anonymous:1970:APc

[Ano70d] Anonymous. Abstracts of papers. Annals of Mathematical Statistics, 41(3):1133–1158, June 1970. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177697000.

Anonymous:1970:APd

[Ano70e] Anonymous. Abstracts of papers. Annals of Mathematical Statistics, 41(4):1384–1393, August 1970. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177696923.

Anonymous:1970:APe

[Ano70f] Anonymous. Abstracts of papers. Annals of Mathematical Statistics, 41(6):2187–2190, December 1970. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177696730.

Anonymous:1970:NNa

[Ano70g] Anonymous. News and notices. Annals of Mathematical Statistics, 41(1):333–335, February 1970. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177697217.

Anonymous:1970:NNb

[Ano70h] Anonymous. News and notices. Annals of Mathematical Statistics, 41 (2):758–763, April 1970. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177697133.

Anonymous:1970:NNc

[Ano70i] Anonymous. News and notices. Annals of Mathematical Statistics, 41(3):1159–1171, June 1970. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177697001.

Anonymous:1970:NNd

[Ano70j] Anonymous. News and notices. Annals of Mathematical Statistics, 41(4):1394–1395, August 1970. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177696924.

Anonymous:1970:NNe

[Ano70k] Anonymous. News and notices. Annals of Mathematical Statistics, 41(5):1810–1814, October 1970. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177696828.

Anonymous:1970:NNf

[Ano70l] Anonymous. News and notices. Annals of Mathematical Statistics, 41 (6):2191–2194, December 1970. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177696731.

Anonymous:1970:PRa

[Ano70m] Anonymous. Publications received. Annals of Mathematical Statistics, 41(1):335–336, February 1970. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177697218.

Anonymous:1970:PRb

[Ano70n] Anonymous. Publications received. Annals of Mathematical Statistics, 41(2):764, April 1970. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177697134.

Anonymous:1970:PRc

[Ano70o] Anonymous. Publications received. Annals of Mathematical Statistics, 41(3):1172, June 1970. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177697002.

Anonymous:1970:PRd

[Ano70p] Anonymous. Publications received. Annals of Mathematical Statistics, 41(4):1396, August 1970. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177696925.

Anonymous:1970:PRe

[Ano70q] Anonymous. Publications received. Annals of Mathematical Statistics, 41(5):1815, October 1970. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177696829.

Anonymous:1970:PRf

[Ano70r] Anonymous. Publications received. Annals of Mathematical Statistics, 41(6):2195, December 1970. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177696732.

Anonymous:1971:A

[Ano71a] Anonymous. Abstracts. Annals of Mathematical Statistics, 42(1): 416–421, February 1971. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177693539.

Anonymous:1971:APa

[Ano71b] Anonymous. Abstracts of papers. Annals of Mathematical Statistics, 42(2):843–846, April 1971. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177693447.

Anonymous:1971:APb

[Ano71c] Anonymous. Abstracts of papers. Annals of Mathematical Statistics, 42(3):1142–1153, June 1971. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177693352.

Anonymous:1971:APc

[Ano71d] Anonymous. Abstracts of papers. Annals of Mathematical Statistics, 42(4):1469–1492, August 1971. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177693267.

Anonymous:1971:APd

[Ano71e] Anonymous. Abstracts of papers. Annals of Mathematical Statistics, 42(5):1782–1799, October 1971. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177693189.

Anonymous:1971:APe

[Ano71f] Anonymous. Abstracts of papers. Annals of Mathematical Statistics, 42(6):2167–2202, December 1971. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177693089.

Anonymous:1971:NNa

[Ano71g] Anonymous. News and notices. Annals of Mathematical Statistics, 42(1):422–437, February 1971. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177693540.

Anonymous:1971:NNb

[Ano71h] Anonymous. News and notices. Annals of Mathematical Statistics, 42 (2):847–854, April 1971. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177693448.

Anonymous:1971:NNc

[Ano71i] Anonymous. News and notices. Annals of Mathematical Statistics, 42(3):1154–1155, June 1971. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177693353.

Anonymous:1971:NNd

[Ano71j] Anonymous. News and notices. Annals of Mathematical Statistics, 42(4):1493–1496, August 1971. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177693268.

Anonymous:1971:NNe

[Ano71k] Anonymous. News and notices. Annals of Mathematical Statistics, 42(5):1800–1813, October 1971. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177693190.

Anonymous:1971:NNf

[Ano71l] Anonymous. News and notices. Annals of Mathematical Statistics, 42 (6):2203–2211, December 1971. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177693090.

Anonymous:1971:PRa

[Ano71m] Anonymous. Publications received. Annals of Mathematical Statistics, 42(1):438, February 1971. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177693541.

Anonymous:1971:PRb

[Ano71n] Anonymous. Publications received. Annals of Mathematical Statistics, 42(2):854, April 1971. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177693449.

Anonymous:1971:PRc

[Ano71o] Anonymous. Publications received. Annals of Mathematical Statistics, 42(3):1156, June 1971. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177693354.

Anonymous:1971:PRd

[Ano71p] Anonymous. Publications received. Annals of Mathematical Statistics, 42(4):1497–1498, August 1971. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177693269.

Anonymous:1971:PRe

[Ano71q] Anonymous. Publications received. Annals of Mathematical Statistics, 42(5):1814, October 1971. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177693191.

Anonymous:1971:PRf

[Ano71r] Anonymous. Publications received. Annals of Mathematical Statistics, 42(6):2212–2213, December 1971. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177693091.

Anonymous:1972:GNJ

[Ano72] Anonymous. George E. Nicholson, Jr., 1918–1971. Annals of Mathematical Statistics, 43(5):1381–1382, October 1972. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177692371.

Antle:1962:USO

[Ant62] Charles E. Antle. The uniqueness of the spacing of observations in polynomial regression for minimax variance of the fitted values. Annals of Mathematical Statistics, 33(2):810–811, June 1962. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177704602.

Anderson:1949:EPS

[AR49] T. W. Anderson and Herman Rubin. Estimation of the parameters of a single equation in a complete system of stochastic equations. Annals of Mathematical Statistics, 20(1):46–63, March 1949. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177730090.

Anderson:1950:APE

[AR50] T. W. Anderson and Herman Rubin. The asymptotic properties of estimates of the parameters of a single equation in a complete system of stochastic equations. Annals of Mathematical Statistics, 21(4):570– 582, December 1950. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177729752.

Arbib:1967:RSS

[Arb67] Michael Arbib. Realization of stochastic systems. Annals of Mathematical Statistics, 38(3):927–933, June 1967. CODEN AASTAD.

ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177698889.

Arnold:1948:RMM

[Arn48] K. J. Arnold. Report on the Madison meeting of the Institute. Annals of Mathematical Statistics, 19(4):608–610, December 1948. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177730160.

Arnold:1950:RCM

[Arn50] J. K. Arnold. Report of the Chicago meeting of the Institute. Annals of Mathematical Statistics, 21(3):467–468, September 1950. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177729811.

Arnold:1951:MAM

[Arn51a] K. J. Arnold. Minutes of the annual membership meeting, Chicago, December 27, 1950. Annals of Mathematical Statistics, 22(1):157– 158, March 1951. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177729711.

Arnold:1951:RCM

[Arn51b] K. J. Arnold. Report of the Chicago meeting of the Institute. Annals of Mathematical Statistics, 22(1):149–157, March 1951. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177729710.

Arnold:1951:RMM

[Arn51c] K. J. Arnold. Report of the Minneapolis meeting of the Institute. Annals of Mathematical Statistics, 22(4):617–621, December 1951. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177729556.

Arnold:1965:SSP

[Arn65] Harvey J. Arnold. Small sample power of the one sample Wilcoxon test for non-normal shift alternatives. Annals of Mathematical Statistics, 36(6):1767–1778, December 1965. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177699805.

Aroian:1937:TBG

[Aro37] Leo A. Aroian. The Type B Gram–Charlier series. Annals of Mathematical Statistics, 8(4):183–192, December 1937. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177732388.

Aroian:1941:CFI

[Aro41a] Leo A. Aroian. Continued fractions for the incomplete Beta function. Annals of Mathematical Statistics, 12(2):218–223, June 1941. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177731751. See

additions and corrections [Aro59, TW79].

Aroian:1941:SRF

[Aro41b] Leo A. Aroian. A study of R. A. Fisher’s z distribution and the related F distribution. Annals of Mathematical Statistics, 12(4):429– 448, December 1941. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177731681.

Aroian:1943:NAL

[Aro43] Leo A. Aroian. A new approximation to the levels of significance of the chi-square distribution. Annals of Mathematical Statistics, 14 (1):93–95, March 1943. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177731497.

Aroian:1947:PFP

[Aro47] Leo A. Aroian. The probability function of the product of two normally distributed variables. Annals of Mathematical Statistics, 18 (2):265–271, June 1947. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177730442.

Aroian:1948:FDE

[Aro48] Leo A. Aroian. The fourth degree exponential distribution function. Annals of Mathematical Statistics, 19(4):589–592, December 1948. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177730155.

Aroian:1959:CNC

[Aro59] Leo A. Aroian. Correction notes: Corrections to “Continued Fractions for the Incomplete Beta Function”. Annals of Mathematical Statistics, 30(4):1265, December 1959. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177706114. See [Aro41a].

Arvesen:1969:JS

[Arv69] James N. Arvesen. Jackknifing U-statistics. Annals of Mathematical Statistics, 40(6):2076–2100, December 1969. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http:

//projecteuclid.org/euclid.aoms/1177697287. See correction

[Arv70].

Arvesen:1970:CNC

[Arv70] James N. Arvesen. Correction notes: Correction to “Jackknifing U-Statistics”. Annals of Mathematical Statistics, 41(4):1375, August 1970. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/

1177696920. See [Arv69].

Aitchison:1958:MLE

[AS58] J. Aitchison and S. D. Silvey. Maximum-likelihood estimation of parameters subject to restraints. Annals of Mathematical Statistics, 29(3):813–828, September 1958. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177706538.

Adler:1972:IMS

[ASS72] R. L. Adler, P. Shields, and M. Smorodinsky. Irreducible Markov shifts. Annals of Mathematical Statistics, 43(3):1027–1029, June 1972. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177692569.

Abelson:1963:EUN

[AT63] Robert P. Abelson and John W. Tukey. Efficient utilization of nonnumerical information in quantitative analysis: General theory and the case of simple order. Annals of Mathematical Statistics, 34(4): 1347–1369, December 1963. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177703869.

Alam:1972:SLP

[AT72] Khursheed Alam and James R. Thompson. On selecting the least probable multinomial event. Annals of Mathematical Statistics, 43 (6):1981–1990, December 1972. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177690870.

Athreya:1968:SRM

[Ath68] Krishna Balasundaram Athreya. Some results on multitype continuous time Markov branching processes. Annals of Mathematical Statistics, 39(2):347–357, April 1968. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177698395.

Athreya:1969:SOD

[Ath69] Krishna B. Athreya. On the supercritical one dimensional age dependent branching processes. Annals of Mathematical Statistics, 40 (3):743–763, June 1969. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177697585.

Athreya:1970:SPR

[Ath70] Krishna B. Athreya. A simple proof of a result of Kesten and Stigum on supercritical multitype Galton–Watson branching process. Annals of Mathematical Statistics, 41(1):195–202, February 1970. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177697200.

Atwood:1969:OED

[Atw69] Corwin L. Atwood. Optimal and efficient designs of experiments. Annals of Mathematical Statistics, 40(5):1570–1602, October 1969. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177697374.

Austin:1966:SFP

[Aus66] D. G. Austin. A sample function property of martingales. Annals of Mathematical Statistics, 37(5):1396–1397, October 1966. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177699283.

Austin:1967:NBE

[Aus67] D. G. Austin. A note on the Birkhoff ergodic theorem. Annals of Mathematical Statistics, 38(3):922–923, June 1967. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177698887.

Anderson:1964:ADA

[AW64] T. W. Anderson and A. M. Walker. On the asymptotic distribution of the autocorrelations of a sample from a linear stochastic process. Annals of Mathematical Statistics, 35(3):1296–1303, September 1964. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177703285.

Birnbaum:1949:SST

[BA49] Z. W. Birnbaum and F. C. Andrews. On sums of symmetrically truncated normal random variables. Annals of Mathematical Statistics, 20(3):458–461, September 1949. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177730001.

Bacon:1938:NFM

[Bac38] H. M. Bacon. Note on a formula for the multiple correlation coefficient. Annals of Mathematical Statistics, 9(3):227–229, September 1938. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177732313.

Bacon:1948:MAC

[Bac48] H. M. Bacon. A matrix arising in correlation theory. Annals of Mathematical Statistics, 19(3):422–424, September 1948. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177730211.

Bacon:1963:AMN

[Bac63] Ralph Hoyt Bacon. Approximations to multivariate normal orthant probabilities. Annals of Mathematical Statistics, 34(1):191– 198, March 1963. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177704254.

Baer:1945:SCP

[Bae45] Reinhold Baer. Sampling from a changing population. Annals of Mathematical Statistics, 16(4):348–361, December 1945. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177731061.

Bagai:1965:DGV

[Bag65] O. P. Bagai. The distribution of the generalized variance. Annals of Mathematical Statistics, 36(1):120–130, February 1965. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177700276.

Bahadur:1950:PTK

[Bah50] Raghu Raj Bahadur. On a problem in the theory of k populations. Annals of Mathematical Statistics, 21(3):362–375, September 1950. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177729795.

Bahadur:1954:SSD

[Bah54] R. R. Bahadur. Sufficiency and statistical decision functions. Annals of Mathematical Statistics, 25(3):423–462, September 1954. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177728715. See com-

ments [Wei53b].

Bahadur:1955:CS

[Bah55a] R. R. Bahadur. A characterization of sufficiency. Annals of Mathematical Statistics, 26(2):286–293, June 1955. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177728545.

Bahadur:1955:SS

[Bah55b] R. R. Bahadur. Statistics and subfields. Annals of Mathematical Statistics, 26(3):490–497, September 1955. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177728493.

Bahadur:1958:NFI

[Bah58] R. R. Bahadur. A note on the fundamental identity of sequential analysis. Annals of Mathematical Statistics, 29(2):534–543, June 1958. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177706628.

Bahadur:1960:SAB

[Bah60a] R. R. Bahadur. Some approximations to the binomial distribution function. Annals of Mathematical Statistics, 31(1):43–54, March 1960. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177705986.

Bahadur:1960:SCT

[Bah60b] R. R. Bahadur. Stochastic comparison of tests. Annals of Mathematical Statistics, 31(2):276–295, June 1960. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177705894.

Bahadur:1964:FBA

[Bah64] R. R. Bahadur. On Fisher’s bound for asymptotic variances. Annals of Mathematical Statistics, 35(4):1545–1552, December 1964. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177700378.

Bahadur:1966:NQL

[Bah66] R. R. Bahadur. A note on quantiles in large samples. Annals of Mathematical Statistics, 37(3):577–580, June 1966. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177699450.

Bahadur:1967:RCE

[Bah67] R. R. Bahadur. Rates of convergence of estimates and test statistics. Annals of Mathematical Statistics, 38(2):303–324, April 1967. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177698949.

Bailey:1931:TFF

[Bai31] Joshua L. Bailey. A table to facilitate the fitting of certain logistic curves. Annals of Mathematical Statistics, 2(3):355–359, August 1931. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177732978.

Baker:1930:TBD

[Bak30] G. A. Baker. Transformations of bimodal distriburtions. Annals of Mathematical Statistics, 1(4):334–344, November 1930. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177733063.

Baker:1931:RBM

[Bak31] G. A. Baker. The relation between the means and variances, means squared and variances in samples from combinations of normal populations. Annals of Mathematical Statistics, 2(3):333–354, August 1931. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177732977.

Baker:1932:DMD

[Bak32] G. A. Baker. Distribution of the means divided by the standard deviations of samples from non-homogeneous populations. Annals of Mathematical Statistics, 3(1):1–9, February 1932. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177732927.

Baker:1934:TNN

[Bak34] G. A. Baker. Transformation of non-normal frequency distributions into normal distributions. Annals of Mathematical Statistics, 5(2): 113–123, June 1934. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177732710.

Baker:1935:NDS

[Bak35a] G. A. Baker. Note on the distributions of the standard deviations and second moments of samples from a Gram–Charlier population. Annals of Mathematical Statistics, 6(3):127–130, September 1935. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177732587.

Baker:1935:PMS

[Bak35b] G. A. Baker. The probability that the mean of a second sample will differ from the mean of a first sample by less than a certain multiple of the standard deviation of the first sample. Annals of Mathematical Statistics, 6(4):197–201, December 1935. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177732565.

Baker:1937:CST

[Bak37] George A. Baker. Correlation surfaces of two or more indices when the components of the indices are normally distributed. Annals of Mathematical Statistics, 8(4):179–182, December 1937. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177732387.

Baker:1940:CPA

[Bak40] G. A. Baker. Comparison of Pearsonian approximations with exact sampling distributions of means and variances in samples from populations composed of the sums of normal populations. Annals of Mathematical Statistics, 11(2):219–224, June 1940. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177731918.

Baker:1941:THN

[Bak41] G. A. Baker. Test of homogeneity for normal populations. Annals of Mathematical Statistics, 12(2):233–236, June 1941. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177731754.

Baker:1946:DRS

[Bak46] G. A. Baker. Distribution of the ratio of sample range to sample standard deviation for normal and combinations of normal distributions. Annals of Mathematical Statistics, 17(3):366–369, September 1946. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177730952.

Baker:1949:VPS

[Bak49] G. A. Baker. The variance of the proportions of samples falling within a fixed interval for a normal population. Annals of Mathematical Statistics, 20(1):123–124, March 1949. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177730102.

Baker:1952:UFT

[Bak52] G. A. Baker. Uniformity field trials when differences in Fertility levels of subplots are not included in experimental error. Annals of Mathematical Statistics, 23(2):289–293, June 1952. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177729448.

Balakrishnan:1959:CC

[Bal59] A. V. Balakrishnan. On a characterization of covariances. Annals of Mathematical Statistics, 30(3):670–675, September 1959. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177706197.

Baldessari:1967:DQF

[Bal67] Bruno Baldessari. The distribution of a quadratic form of normal random variables. Annals of Mathematical Statistics, 38(6):1700– 1704, December 1967. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177698604.

Bancroft:1944:BED

[Ban44] T. A. Bancroft. On biases in estimation due to the use of preliminary tests of significance. Annals of Mathematical Statistics, 15(2):190– 204, June 1944. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/

1177731284. See note [Ban45].

Bancroft:1945:NII

[Ban45] T. A. Bancroft. Note on an identity in the incomplete beta function. Annals of Mathematical Statistics, 16(1):98–99, March 1945. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177731177. See

[Ban44].

Banerjee:1948:WDB

[Ban48] K. S. Banerjee. Weighing designs and balanced incomplete blocks. Annals of Mathematical Statistics, 19(3):394–399, September 1948. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177730204.

Bancroft:1949:SRF

[Ban49a] T. A. Bancroft. Some recurrence formulae in the incomplete beta function ratio. Annals of Mathematical Statistics, 20(3):451–455, September 1949. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177729999.

Banerjee:1949:NWD

[Ban49b] K. S. Banerjee. A note on weighing design. Annals of Mathematical Statistics, 20(2):300–304, June 1949. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177730040.

Bankier:1960:OAW

[Ban60] J. D. Bankier. An operational approach to the r-way crossed classification. Annals of Mathematical Statistics, 31(1):16–22, March 1960. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177705983.

Banerjee:1964:NIM

[Ban64] K. S. Banerjee. A note on idempotent matrices. Annals of Mathematical Statistics, 35(2):880–882, June 1964. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177703592.

Banerjee:1965:HWD

[Ban65] K. S. Banerjee. On Hotelling’s weighing designs under autocorrelation of errors. Annals of Mathematical Statistics, 36(6):1829–1834, December 1965. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177699812.

Banerjee:1966:NRF

[Ban66a] K. S. Banerjee. On non-randomized fractional weighing designs. Annals of Mathematical Statistics, 37(6):1836–1841, December 1966. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177699178.

Banerjee:1966:SHW

[Ban66b] K. S. Banerjee. Singularity in Hotelling’s weighing designs and a generalized inverse. Annals of Mathematical Statistics, 37(4):1021– 1032, August 1966. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.

aoms/1177699382. See correction [Ban69].

Banerjee:1967:CIP

[Ban67] Saibal Banerjee. Confidence interval of preassigned length for the Behrens–Fisher problem. Annals of Mathematical Statistics, 38(4): 1175–1179, August 1967. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177698787.

Banos:1968:PG

[Bañ68] Alfredo Baños. On pseudo-games. Annals of Mathematical Statistics, 39(6):1932–1945, December 1968. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177698023.

Banerjee:1969:CNC

[Ban69] K. S. Banerjee. Correction note: Correction to “Singularity in Hotelling’s Weighing Designs and a Generalized Inverse”. Annals of Mathematical Statistics, 40(2):719, April 1969. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177697750. See [Ban66b].

Banerjee:1970:SOW

[Ban70] K. S. Banerjee. Some observations on the White–Hultquist procedure for the construction of confounding plans for mixed factorial designs. Annals of Mathematical Statistics, 41(3):1108–1113, June 1970. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177696993.

Barkey:1935:FDP

[Bar35] I. H. Barkey. On the finite differences of a polynomial. Annals of Mathematical Statistics, 6(3):131–132, September 1935. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177732588.

Bartlett:1939:CSF

[Bar39] M. S. Bartlett. Complete simultaneous fiducial distributions. Annals of Mathematical Statistics, 10(2):129–138, June 1939. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177732211.

Bartky:1943:MSC

[Bar43] Walter Bartky. Multiple sampling with constant probability. Annals of Mathematical Statistics, 14(4):363–377, December 1943. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177731357.

Bartlett:1947:GCC

[Bar47a] M. S. Bartlett. The general canonical correlation distribution. Annals of Mathematical Statistics, 18(1):1–17, March 1947. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177730488.

Bartlett:1947:RBA

[Bar47b] M. S. Bartlett. Review: The Advanced Theory of Statistics. Vol. I; Maurice G. Kendall, The Advanced Theory of Statistics. Vol II. Annals of Mathematical Statistics, 18(1):139–142, March 1947. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177730504.

Barankin:1949:LBU

[Bar49] E. W. Barankin. Locally best unbiased estimates. Annals of Mathematical Statistics, 20(4):477–501, December 1949. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177729943.

Barankin:1950:ETB

[Bar50] E. W. Barankin. Extension of a theorem of Blackwell. Annals of Mathematical Statistics, 21(2):280–284, June 1950. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177729845.

Bartlett:1951:IMA

[Bar51] M. S. Bartlett. An inverse matrix adjustment arising in discriminant analysis. Annals of Mathematical Statistics, 22(1):107–111, March 1951. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177729698.

Bartoszynski:1961:CWC

[Bar61] Robert Bartoszyński. A characterization of the weak convergence of measures. Annals of Mathematical Statistics, 32(2):561–576, June 1961. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177705061.

Barlow:1965:BIA

[Bar65] Richard E. Barlow. Bounds on integrals with applications to reliability problems. Annals of Mathematical Statistics, 36(2):565–574, April 1965. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177700168.

Barlow:1968:LRT

[Bar68] Richard E. Barlow. Likelihood ratio tests for restricted families of probability distributions. Annals of Mathematical Statistics, 39(2): 547–560, April 1968. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177698416.

Bartholomew:1969:SCM

[Bar69] D. J. Bartholomew. Sufficient conditions for a mixture of exponentials to be a probability density function. Annals of Mathematical Statistics, 40(6):2183–2188, December 1969. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177697296.

Baranchik:1970:FME

[Bar70] A. J. Baranchik. A family of minimax estimators of the mean of a multivariate normal distribution. Annals of Mathematical Statistics, 41(2):642–645, April 1970. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177697104.

Basu:1955:IMM

[Bas55a] D. Basu. An inconsistency of the method of maximum likelihood. Annals of Mathematical Statistics, 26(1):144–145, March 1955. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177728606.

Basu:1955:NTU

[Bas55b] D. Basu. A note on the theory of unbiased estimation. Annals of Mathematical Statistics, 26(2):345–348, June 1955. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177728554.

Basu:1964:ETT

[Bas64] A. P. Basu. Effect of truncation on a test for the scale parameter of the exponential distribution. Annals of Mathematical Statistics, 35(1):209–213, March 1964. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177703744.

Basu:1967:LSP

[Bas67a] A. P. Basu. On the large sample properties of a generalized Wilcoxon– Mann–Whitney statistic. Annals of Mathematical Statistics, 38(3): 905–915, June 1967. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177698884.

Basu:1967:TSR

[Bas67b] A. P. Basu. On two K-sample rank tests for censored data. Annals of Mathematical Statistics, 38(5):1520–1535, October 1967. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177698706.

Basu:1968:GSS

[Bas68] A. P. Basu. On a generalized Savage statistic with applications to life testing. Annals of Mathematical Statistics, 39(5):1591–1604, October 1968. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177698140.

Baten:1930:STD

[Bat30] William Dowell Baten. Simultaneous treatment of discrete and continuous probability by use of Stieltjes integrals. Annals of Mathematical Statistics, 1(1):95–100, February 1930. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177733262.

Baten:1931:CMF

[Bat31] William Dowell Baten. Correction for the moments of a frequency distribution in two variables. Annals of Mathematical Statistics, 2(3): 309–319, August 1931. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177732974.

Baten:1934:CTP

[Bat34] William Dowell Baten. Combining two probability functions. Annals of Mathematical Statistics, 5(1):13–20, March 1934. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177732732.

Battin:1942:PMM

[Bat42] I. L. Battin. On the problem of multiple matching. Annals of Mathematical Statistics, 13(3):294–305, September 1942. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177731568.

Bates:1955:JDT

[Bat55] Grace E. Bates. Joint distributions of time intervals for the occurrence of successive accidents in a generalized Pólya Scheme. Annals of Mathematical Statistics, 26(4):705–720, December 1955. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177728429.

Bather:1965:ICD

[Bat65] J. A. Bather. Invariant conditional distributions. Annals of Mathematical Statistics, 36(3):829–846, June 1965. CODEN AASTAD.

ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177700057.

Bather:1967:QDP

[Bat67] J. A. Bather. On a quickest detection problem. Annals of Mathematical Statistics, 38(3):711–724, June 1967. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177698864.

Baum:1963:APC

[Bau63a] Leonard E. Baum. Acknowledgment of priority: “On Convergence to +∞ in the Law of Large Numbers”. Annals of Mathematical Statistics, 34(3):1111, September 1963. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177704039.

Baum:1963:CLL

[Bau63b] Leonard E. Baum. On convergence to + ∞ in the law of large numbers. Annals of Mathematical Statistics, 34(1):219–222, March 1963. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177704258.

Baxter:1961:CLC

[Bax61] Glen Baxter. A combinatorial lemma for complex numbers. Annals of Mathematical Statistics, 32(3):901–904, September 1961. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177704985.

Baxter:1962:GFA

[Bax62] Glen Baxter. On a generalization of the finite arcsine law. Annals of Mathematical Statistics, 33(3):909–915, September 1962. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177704459.

Baxter:1965:GME

[Bax65] Glen Baxter. Generalizations of the maximal ergodic theorem. Annals of Mathematical Statistics, 36(4):1292–1293, August 1965. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177700001.

Bose:1952:OAS

[BB52] R. C. Bose and K. A. Bush. Orthogonal arrays of strength two and three. Annals of Mathematical Statistics, 23(4):508–524, December 1952. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177729331.

Blackwell:1955:MAG

[BB55] David Blackwell and Albert H. Bowker. Meyer Abraham Girshick 1908–1955. Annals of Mathematical Statistics, 26(3):365–367, September 1955. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177728484.

Box:1960:SSD

[BB60] G. E. P. Box and D. W. Behnken. Simplex-sum designs: A class of second order rotatable designs derivable from those of first order. Annals of Mathematical Statistics, 31(4):838–864, December 1960. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177705661.

Bofinger:1961:GTR

[BB61] Eve Bofinger and V. J. Bofinger. The gap test for random sequences. Annals of Mathematical Statistics, 32(2):524–534, June 1961. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177705058;   http:
//www.jstor.org/stable/2237761.

Baum:1965:ADC

[BB65] Leonard E. Baum and Patrick Billingsley. Asymptotic distributions for the coupon Collector’s problem. Annals of Mathematical Statistics, 36(6):1835–1839, December 1965. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177699813.

Bickel:1967:NBE

[BB67] Peter J. Bickel and David Blackwell. A note on Bayes estimates. Annals of Mathematical Statistics, 38(6):1907–1911, December 1967. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177698625.

Bahadur:1968:SCE

[BB68a] R. R. Bahadur and P. J. Bickel. Substitution in conditional expectation. Annals of Mathematical Statistics, 39(2):377–378, April 1968. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177698399.

Berk:1968:IAI

[BB68b] Robert H. Berk and P. J. Bickel. On invariance and almost invariance. Annals of Mathematical Statistics, 39(5):1573–1576, October 1968. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177698138.

Boswell:1969:DLR

[BB69] M. T. Boswell and H. D. Brunk. Distribution of likelihood ratio in testing against trend. Annals of Mathematical Statistics, 40(2): 371–380, April 1969. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177697701.

Bell:1960:COS

[BBB60] C. B. Bell, David Blackwell, and Leo Breiman. On the completeness of order statistics. Annals of Mathematical Statistics, 31(3):794–797, September 1960. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177705808.

Brons:1969:GMA

[BBFH69] H. K. Brons, H. D. Brunk, W. E. Franck, and D. L. Hanson. Generalized means and associated families of distributions. Annals of Mathematical Statistics, 40(2):339–355, April 1969. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177697698.

Bozivich:1956:PAV

[BBH56] Helen Bozivich, T. A. Bancroft, and H. O. Hartley. Power of analysis of variance test procedures for certain incompletely specified models, I. Annals of Mathematical Statistics, 27(4):1017–1043, December 1956. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177728071.

Blackwell:1958:PST

[BBT58] David Blackwell, Leo Breiman, and A. J. Thomasian. Proof of Shannon’s transmission theorem for finite-state indecomposable channels. Annals of Mathematical Statistics, 29(4):1209–1220, December 1958. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177706452.

Blackwell:1959:CCC

[BBT59] David Blackwell, Leo Breiman, and A. J. Thomasian. The capacity of a class of channels. Annals of Mathematical Statistics, 30(4):1229– 1241, December 1959. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177706106.

Blackwell:1960:CCC

[BBT60] David Blackwell, Leo Breiman, and A. J. Thomasian. The capacities of certain channel classes under random coding. Annals of Mathematical Statistics, 31(3):558–567, September 1960. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177705783.

Birnbaum:1950:OSM

[BC50] Z. W. Birnbaum and D. G. Chapman. On optimum selections from multinormal populations. Annals of Mathematical Statistics, 21(3): 443–447, September 1950. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177729804.

Bose:1952:CPG

[BC52] R. C. Bose and W. S. Connor. Combinatorial properties of group divisible incomplete block designs. Annals of Mathematical Statistics, 23(3):367–383, September 1952. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177729382.

Barrow:1954:SFI

[BC54] D. F. Barrow and A. C. Cohen, Jr. On some functions involving Mill’s ratio. Annals of Mathematical Statistics, 25(2):405–408, June 1954. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177728801.

Bose:1955:SCP

[BC55] R. C. Bose and W. H. Clatworthy. Some classes of partially balanced designs. Annals of Mathematical Statistics, 26(2):212–232, June 1955. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177728539.

Burton:1957:IRF

[BC57] R. C. Burton and W. S. Connor. On the identity relationship for fractional replicates of the 2^{n} series. Annals of Mathematical Statistics, 28(3):762–767, September 1957. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177706888.

Bose:1959:CRC

[BC59] R. C. Bose and R. L. Carter. Complex representation in the construction of rotatable designs. Annals of Mathematical Statistics, 30 (3):771–780, September 1959. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177706206.

Breakwell:1964:STM

[BC64] John Breakwell and Herman Chernoff. Sequential tests for the mean of a normal distribution II (large t). Annals of Mathematical Statistics, 35(1):162–173, March 1964. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177703738.

Blumenthal:1968:ELT

[BC68a] Saul Blumenthal and Arthur Cohen. Estimation of the larger translation parameter. Annals of Mathematical Statistics, 39(2):502–516, April 1968. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177698413.

Blumenthal:1968:ETO

[BC68b] Saul Blumenthal and Arthur Cohen. Estimation of two ordered translation parameters. Annals of Mathematical Statistics, 39(2): 517–530, April 1968. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177698414.

Bliss:1945:PRC

[BCD^{+}45] C. I. Bliss, W. G. Cochran, W. E. Deming, P. S. Olmstead, and S. S. Wilks. Progress report of the Committee on Post-War Development of the Institute. Annals of Mathematical Statistics, 16(1): 109–111, March 1945. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177731182.

Bose:1959:SOR

[BD59] R. C. Bose and Norman R. Draper. Second order rotatable designs in three dimensions. Annals of Mathematical Statistics, 30(4):1097– 1112, December 1959. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177706093.

Blackwell:1962:MOI

[BD62] David Blackwell and Lester Dubins. Merging of opinions with increasing information. Annals of Mathematical Statistics, 33(3):882– 886, September 1962. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177704456.

Birnbaum:1963:LOS

[BD63] Allan Birnbaum and Jack Dudman. Logistic order statistics. Annals of Mathematical Statistics, 34(2):658–663, June 1963. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177704178.

Bell:1965:APS

[BD65a] C. B. Bell and K. A. Doksum. Acknowledgment of priority: Some new distribution-free statistics. Annals of Mathematical Statistics, 36(6): 1901, December 1965. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177699825. See [BD65b].

Bell:1965:SND

[BD65b] C. B. Bell and K. A. Doksum. Some new distribution-free statistics. Annals of Mathematical Statistics, 36(1):203–214, February 1965. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177700283. See

acknowledgment of priority [BD65a].

Bell:1966:OOS

[BD66] C. B. Bell and K. A. Doksum. ”optimal” one-sample distribution-free tests and their two-sample extensions. Annals of Mathematical Statistics, 37(1):120–132, February 1966. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177699603.

Bell:1967:DFT

[BD67] C. B. Bell and K. A. Doksum. Distribution-free tests of independence. Annals of Mathematical Statistics, 38(2):429–446, April 1967. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177698959.

Bickel:1969:TMFa

[BD69] Peter J. Bickel and Kjell A. Doksum. Tests for monotone failure rate based on normalized spacings. Annals of Mathematical Statistics, 40 (4):1216–1235, August 1969. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177697498.

Balkema:1972:RMC

[BD72] A. A. Balkema and L. De Haan. On R. von Mises’ condition for the domain of attraction of exp(−e^{−x}). Annals of Mathematical Statistics, 43(4):1352–1354, August 1972. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177692489.

Blackwell:1964:LRE

[BDF64] D. Blackwell, P. Deuel, and D. Freedman. The last return to equilibrium in a coin-tossing game. Annals of Mathematical Statistics, 35(3):1344, September 1964. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177703291.

Brown:1970:BMR

[BE70] B. M. Brown and G. K. Eagleson. Behavior of moments of row sums of elementary systems. Annals of Mathematical Statistics, 41(6):1853– 1860, December 1970. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177696686.

Blum:1972:CMT

[BE72] Julius R. Blum and Bennett Eisenberg. Conditions for metric transitivity for stationary Gaussian processes on groups. Annals of Mathematical Statistics, 43(5):1737–1741, October 1972. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177692412.

Beale:1937:PRD

[Bea37] Frank S. Beale. On the polynomials related to the differential equation \frac{1}{y} \frac{dy}{dx} = \frac{a_{0}+a_{1}x}{b_{0}+b_{1}x+b_{2}x^{2}} ≡ \frac{N}{D}. Annals of Mathematical Statistics, 8 (4):206–223, December 1937. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177732390.

Beale:1941:CCO

[Bea41] Frank S. Beale. On a certain class of orthogonal polynomials. Annals of Mathematical Statistics, 12(1):97–103, March 1941. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177731789.

Beckenbach:1942:CPG

[Bec42] E. F. Beckenbach. Convexity properties of generalized mean value functions. Annals of Mathematical Statistics, 13(1):88–90, March 1942. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177731646.

Bechhofer:1954:SSMa

[Bec54] Robert E. Bechhofer. A single-sample multiple decision procedure for ranking means of normal populations with known variances. Annals of Mathematical Statistics, 25(1):16–39, March 1954. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177728845.

Behnken:1961:SMM

[Beh61] D. W. Behnken. Sampling moments of means from finite multivariate populations. Annals of Mathematical Statistics, 32(2):406– 413, June 1961. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177705049.

Behnen:1971:AOC

[Beh71] Konrad Behnen. Asymptotic optimality and ARE of certain rankorder tests under contiguity. Annals of Mathematical Statistics, 42 (1):325–329, February 1971. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177693515.

Behnen:1972:CCR

[Beh72] Konrad Behnen. A characterization of certain rank-order tests with bounds for the asymptotic relative efficiency. Annals of Mathematical Statistics, 43(6):1839–1851, December 1972. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177690855.

Belz:1947:NLI

[Bel47] Maurice H. Belz. Note on the Liapounoff inequality for absolute moments. Annals of Mathematical Statistics, 18(4):604–605, December 1947. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177730355.

Bell:1960:SDF

[Bel60] C. B. Bell. On the structure of distribution-free statistics. Annals of Mathematical Statistics, 31(3):703–709, September 1960. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177705796.

Bell:1961:MIS

[Bel61] C. B. Bell. Maximal independent stochastic processes. Annals of Mathematical Statistics, 32(3):704–708, September 1961. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177704965.

Bell:1962:MIM

[Bel62] C. B. Bell. Mutual information and maximal correlation as measures of dependence. Annals of Mathematical Statistics, 33(2):587– 595, June 1962. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177704583.

Bell:1964:CMD

[Bel64a] C. B. Bell. A characterization of multisample distribution-free statistics. Annals of Mathematical Statistics, 35(2):735–738, June 1964. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177703571.

Bell:1964:SBT

[Bel64b] C. B. Bell. Some basic theorems of distribution-free statistics. Annals of Mathematical Statistics, 35(1):150–156, March 1964. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177703736.

Belkin:1970:CNC

[Bel70a] Barry Belkin. Correction notes: Correction to “A Limit Theorem for Conditioned Recurrent Random Walk Attracted to a Stable Law”. Annals of Mathematical Statistics, 41(4):1374, August 1970. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177696918. See

[Bel70b].

Belkin:1970:LTC

[Bel70b] Barry Belkin. A limit theorem for conditioned recurrent random walk attracted to a stable law. Annals of Mathematical Statistics, 41(1): 146–163, February 1970. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.

aoms/1177697195. See correction [Bel70a].

Bechhofer:1959:SSM

[BEM59] Robert E. Bechhofer, Salah Elmaghraby, and Norman Morse. A single-sample multiple-decision procedure for selecting the multinomial event which has the highest probability. Annals of Mathematical Statistics, 30(1):102–119, March 1959. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177706362.

Birnbaum:1966:SCW

[BEM66] Z. W. Birnbaum, J. D. Esary, and A. W. Marshall. A stochastic characterization of wear-out for components and systems. Annals of Mathematical Statistics, 37(4):816–825, August 1966. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177699362.

Benes:1957:QPA

[Ben57] V. E. Beneš. On queues with Poisson arrivals. Annals of Mathematical Statistics, 28(3):670–677, September 1957. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177706878.

Benes:1962:RLT

[Ben62] V. E. Beneš. A “renewal” limit theorem for general stochastic processes. Annals of Mathematical Statistics, 33(1):98–113, March 1962. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177704715.

Benes:1963:RTS

[Ben63] V. E. Benes. Review: Thomas L. Saaty, Elements of Queueing Theory, with Applications. Annals of Mathematical Statistics, 34(4):1610– 1612, December 1963. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177703899.

Berkson:1930:BT

[Ber30] Joseph Berkson. Bayes’ Theorem. Annals of Mathematical Statistics, 1(1):42–56, February 1930. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177733259.

Bernstein:1937:RCE

[Ber37] Felix Bernstein. Regression and correlation evaluated by a method of partial sums. Annals of Mathematical Statistics, 8(2):77–89, June 1937. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177732428.

Bernstein:1942:SMP

[Ber42] S. Bernstein. Solution of a mathematical problem connected with the theory of heredity. Annals of Mathematical Statistics, 13(1):53– 61, March 1942. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177731642.

Berry:1945:CCC

[Ber45] Clifford E. Berry. A criterion of convergence for the classical iterative method of solving linear simultaneous equations. Annals of Mathematical Statistics, 16(4):398–400, December 1945. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177731068.

Berger:1951:UCT

[Ber51a] Agnes Berger. On uniformly consistent tests. Annals of Mathematical Statistics, 22(2):289–293, June 1951. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177729649.

Berger:1951:RSS

[Ber51b] Agnes Berger. Remark on separable spaces of probability measures. Annals of Mathematical Statistics, 22(1):119–120, March 1951. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177729701.

Berndt:1958:CPF

[Ber58a] G. D. Berndt. Correction to “Power Functions of the Gamma Distribution”. Annals of Mathematical Statistics, 29(4):1278, December 1958. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/

1177706463. See [Ber58b].

Berndt:1958:PFG

[Ber58b] Gerald D. Berndt. Power functions of the gamma distribution. Annals of Mathematical Statistics, 29(1):302–306, March 1958. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177706732. See correc-

tion [Ber58a].

Berman:1962:EAS

[Ber62a] Simeon M. Berman. An extension of the arc sine law. Annals of Mathematical Statistics, 33(2):681–684, June 1962. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177704589.

Berman:1962:LLN

[Ber62b] Simeon M. Berman. A law of large numbers for the maximum in a stationary Gaussian sequence. Annals of Mathematical Statistics, 33 (1):93–97, March 1962. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177704714.

Berman:1962:LDM

[Ber62c] Simeon M. Berman. Limiting distribution of the maximum term in sequences of dependent random variables. Annals of Mathematical Statistics, 33(3):894–908, September 1962. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177704458.

Berman:1963:MPB

[Ber63a] Simeon M. Berman. A Markov process on binary numbers. Annals of Mathematical Statistics, 34(2):416–423, June 1963. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177704153.

Berman:1963:NEV

[Ber63b] Simeon M. Berman. Note on extreme values, competing risks and semi-Markov processes. Annals of Mathematical Statistics, 34(3): 1104–1106, September 1963. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177704036.

Berk:1964:EFI

[Ber64a] Kenneth Berk. An evaluation of a functional on infinitely divisible stochastic processes. Annals of Mathematical Statistics, 35(1):336– 340, March 1964. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177703756.

Berman:1964:LTM

[Ber64b] Simeon M. Berman. Limit theorems for the maximum term in stationary sequences. Annals of Mathematical Statistics, 35(2):502– 516, June 1964. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177703551.

Berman:1964:LDM

[Ber64c] Simeon M. Berman. Limiting distribution of the maximum of a diffusion process. Annals of Mathematical Statistics, 35(1):319–329, March 1964. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177703754.

Berk:1966:CNC

[Ber66a] Robert H. Berk. Correction notes: Correction to limiting behavior of posterior distributions when the model is incorrect. Annals of Mathematical Statistics, 37(3):745–746, June 1966. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177699477. See [Ber66b].

Berk:1966:LBP

[Ber66b] Robert H. Berk. Limiting behavior of posterior distributions when the model is incorrect. Annals of Mathematical Statistics, 37(1):51–58, February 1966. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/

1177699597. See correction [Ber66a].

Berk:1967:SGS

[Ber67a] Robert H. Berk. A special group structure and equivariant estimation. Annals of Mathematical Statistics, 38(5):1436–1445, October 1967. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177698698.

Berman:1967:OTT

[Ber67b] Simeon M. Berman. An occupation time theorem for the angular component of plane Brownian motion. Annals of Mathematical Statistics, 38(1):25–31, February 1967. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177699055.

Berk:1968:ETR

[Ber68] Kenneth N. Berk. Ergodic theory with recurrent weights. Annals of Mathematical Statistics, 39(4):1107–1114, August 1968. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177698235.

Beran:1969:ATC

[Ber69a] R. J. Beran. Asymptotic theory of a class of tests for uniformity of a circular distribution. Annals of Mathematical Statistics, 40(4): 1196–1206, August 1969. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177697496.

Berk:1969:BDC

[Ber69b] Robert H. Berk. Biorthogonal and dual configurations and the reciprocal normal distribution. Annals of Mathematical Statistics, 40(2): 393–398, April 1969. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177697703.

Berk:1969:SCC

[Ber69c] Robert H. Berk. Strong consistency of certain sequential estimators. Annals of Mathematical Statistics, 40(4):1492–1495, August 1969. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177697522. See

correction [Ber71c].

Beran:1970:LRS

[Ber70a] R. J. Beran. Linear rank statistics under alternatives indexed by a vector parameter. Annals of Mathematical Statistics, 41(6):1896– 1905, December 1970. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177696691.

Berk:1970:CP

[Ber70b] Robert H. Berk. Consistency a posteriori. Annals of Mathematical Statistics, 41(3):894–906, June 1970. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177696967.

Berk:1970:RAI

[Ber70c] Robert H. Berk. A remark on almost invariance. Annals of Mathematical Statistics, 41(2):733–735, April 1970. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177697125.

Berk:1970:STS

[Ber70d] Robert H. Berk. Stopping times of SPRTS based on exchangeable models. Annals of Mathematical Statistics, 41(3):979–990, June 1970. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177696974.

Berman:1970:GPS

[Ber70e] Simeon M. Berman. Gaussian processes with stationary increments: Local times and sample function properties. Annals of Mathematical Statistics, 41(4):1260–1272, August 1970. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177696901.

Beran:1971:NDF

[Ber71a] R. J. Beran. A note on distribution-free statistical inference with upper and lower probabilities. Annals of Mathematical Statistics, 42 (6):1943–1948, December 1971. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177693061.

Beran:1971:DFS

[Ber71b] R. J. Beran. On distribution-free statistical inference with upper and lower probabilities. Annals of Mathematical Statistics, 42(1):157– 168, February 1971. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177693503.

Berk:1971:CNC

[Ber71c] Robert H. Berk. Correction note: Corrections to “Strong Consistency of Certain Sequential Estimators”. Annals of Mathematical Statistics, 42(3):1135–1137, June 1971. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177693348. See [Ber69c].

Berman:1971:AIN

[Ber71d] Simeon M. Berman. Asymptotic independence of the numbers of high and low level crossings of stationary Gaussian processes. Annals of Mathematical Statistics, 42(3):927–945, June 1971. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177693322.

Beran:1972:RSP

[Ber72a] R. J. Beran. Rank spectral processes and tests for serial dependence. Annals of Mathematical Statistics, 43(6):1749–1766, December 1972. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177690850.

Berger:1972:CCB

[Ber72b] Toby Berger. On the correlation coefficient of a bivariate, equal variance, complex Gaussian sample. Annals of Mathematical Statistics, 43(6):2000–2003, December 1972. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177690873.

Berk:1972:CAN

[Ber72c] Robert H. Berk. Consistency and asymptotic normality of MLE’s for exponential models. Annals of Mathematical Statistics, 43(1):193– 204, February 1972. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.

aoms/1177692713. See acknowledgement of priority and correction [Ber73].

Berk:1972:NSI

[Ber72d] Robert H. Berk. A note on sufficiency and invariance. Annals of Mathematical Statistics, 43(2):647–650, April 1972. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177692645.

Berman:1972:MHL

[Ber72e] Simeon M. Berman. Maximum and high level excursion of a Gaussian process with stationary increments. Annals of Mathematical Statistics, 43(4):1247–1266, August 1972. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177692476.

Berry:1972:BTA

[Ber72f] Donald A. Berry. A Bernoulli two-armed bandit. Annals of Mathematical Statistics, 43(3):871–897, June 1972. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177692553.

Berk:1973:NAP

[Ber73] Robert H. Berk. Notes: Acknowledgement of priority and correction to “Consistency and Asymptotic Normality of MLE’s for Exponential Models”. Annals of Statistics, 1(3):593, May 1973. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1176342432. See [Ber72c].

Beutler:1963:MWS

[Beu63] Frederick J. Beutler. Multivariate wide-sense Markov processes and prediction theory. Annals of Mathematical Statistics, 34(2):424– 438, June 1963. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177704154.

Beutler:1965:RPW

[Beu65] Frederick J. Beutler. Review: P. Whittle, Prediction and Regulation by Linear Least-Squares Methods. Annals of Mathematical Statistics, 36(2):716–718, April 1965. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177700185.

Banerjee:1963:EFC

[BF63a] K. S. Banerjee and W. T. Federer. On estimates for fractions of a complete factorial experiment as orthogonal linear combinations of the observations. Annals of Mathematical Statistics, 34(3):1068– 1078, September 1963. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177704031.

Buehler:1963:NCP

[BF63b] R. J. Buehler and A. P. Feddersen. Note on a conditional property of Student’s t. Annals of Mathematical Statistics, 34(3):1098–1100, September 1963. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177704034.

Banerjee:1964:EEF

[BF64a] K. S. Banerjee and W. T. Federer. Estimates of effects for fractional replicates. Annals of Mathematical Statistics, 35(2):711–715, June 1964. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177703568.

Blackwell:1964:RCT

[BF64b] David Blackwell and David Freedman. A remark on the coin tossing game. Annals of Mathematical Statistics, 35(3):1345–1347, September 1964. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177703292.

Blackwell:1964:TFM

[BF64c] David Blackwell and David Freedman. The tail σ-field of a Markov chain and a theorem of Orey. Annals of Mathematical Statistics, 35 (3):1291–1295, September 1964. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177703284.

Banerjee:1966:ECF

[BF66] K. S. Banerjee and W. T. Federer. On estimation and construction in fractional replication. Annals of Mathematical Statistics, 37(4):1033– 1039, August 1966. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177699383.

Banerjee:1968:SAS

[BF68a] K. S. Banerjee and W. T. Federer. On the structure and analysis of singular fractional replicates. Annals of Mathematical Statistics, 39 (2):657–663, April 1968. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177698424.

Blackwell:1968:BM

[BF68b] David Blackwell and T. S. Ferguson. The big match. Annals of Mathematical Statistics, 39(1):159–163, February 1968. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177698513.

Blackwell:1968:LBM

[BF68c] David Blackwell and David Freedman. On the local behavior of Markov transition probabilities. Annals of Mathematical Statistics, 39 (6):2123–2127, December 1968. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177698042.

Bohrer:1972:SOS

[BF72] Robert Bohrer and George K. Francis. Sharp one-sided confidence bounds over positive regions. Annals of Mathematical Statistics, 43 (5):1541–1548, October 1972. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177692386.

Blackwell:1946:FSI

[BG46] D. Blackwell and M. A. Girshick. On functions of sequences of independent chance vectors with applications to the problem of the “Random Walk” in k dimensions. Annals of Mathematical Statistics, 17(3):310–317, September 1946. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177730943.

Blackwell:1947:LBV

[BG47] D. Blackwell and M. A. Girshick. A lower bound for the variance of some unbiased sequential estimates. Annals of Mathematical Statistics, 18(2):277–280, June 1947. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177730444.

Bahadur:1952:IDR

[BG52] Raghu Raj Bahadur and Leo A. Goodman. Impartial decision rules and sufficient statistics. Annals of Mathematical Statistics, 23(4):553– 562, December 1952. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177729334.

Blumenthal:1964:TST

[BG64] R. M. Blumenthal and R. K. Getoor. A theorem on stopping times. Annals of Mathematical Statistics, 35(3):1348–1350, September 1964. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177703293.

Bloch:1968:SER

[BG68] Daniel A. Bloch and Joseph L. Gastwirth. On a simple estimate of the reciprocal of the density function. Annals of Mathematical Statistics, 39(3):1083–1085, June 1968. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177698342.

Barlow:1969:SPR

[BG69a] Richard E. Barlow and Shanti S. Gupta. Selection procedures for restricted families of probability distributions. Annals of Mathematical Statistics, 40(3):905–917, June 1969. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177697596.

Basu:1969:IST

[BG69b] D. Basu and J. K. Ghosh. Invariant sets for translation-parameter families of measures. Annals of Mathematical Statistics, 40(1):162– 174, February 1969. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177697813.

Barlow:1969:DMM

[BGP69] R. E. Barlow, S. S. Gupta, and S. Panchapakesan. On the distribution of the maximum and minimum of ratios of order statistics. Annals of Mathematical Statistics, 40(3):918–934, June 1969. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177697597.

Blackwell:1957:DCS

[BH57a] David Blackwell and J. L. Hodges, Jr. Design for the control of selection bias. Annals of Mathematical Statistics, 28(2):449–460, June 1957. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177706973.

Box:1957:MFE

[BH57b] G. E. P. Box and J. S. Hunter. Multi-factor experimental designs for exploring response surfaces. Annals of Mathematical Statistics, 28(1): 195–241, March 1957. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177707047.

Blackwell:1959:PET

[BH59] David Blackwell and J. L. Hodges, Jr. The probability in the extreme tail of a convolution. Annals of Mathematical Statistics, 30(4):1113– 1120, December 1959. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177706094.

Birnbaum:1960:EPV

[BH60a] Allan Birnbaum and William C. Healy, Jr. Estimates with prescribed variance based on two-stage sampling. Annals of Mathematical Statistics, 31(3):662–676, September 1960. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177705793.

Birnbaum:1960:SSD

[BH60b] Z. W. Birnbaum and R. A. Hall. Small sample distributions for multisample statistics of the Smirnov type. Annals of Mathematical Statistics, 31(3):710–720, September 1960. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177705797.

Buhlmann:1963:PCR

[BH63] Hans Bühlmann and Peter J. Huber. Pairwise comparison and ranking in tournaments. Annals of Mathematical Statistics, 34(2):501– 510, June 1963. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177704161.

Bolger:1965:CSD

[BH65] E. M. Bolger and W. L. Harkness. Characterizations of some distributions by conditional moments. Annals of Mathematical Statistics, 36 (2):703–705, April 1965. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177700179.

Bickel:1967:ATG

[BH67] P. J. Bickel and J. L. Hodges, Jr. The asymptotic theory of Galton’s test and a related simple estimate of location. Annals of Mathematical Statistics, 38(1):73–89, February 1967. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177699059.

Bell:1969:BST

[BH69] C. B. Bell and H. Smith Haller. Bivariate symmetry tests: Parametric and nonparametric. Annals of Mathematical Statistics, 40(1):259– 269, February 1969. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177697821.

Bhapkar:1959:NMI

[Bha59] V. P. Bhapkar. A note on multiple independence under multivariate normal linear models. Annals of Mathematical Statistics, 30(4):1248– 1251, December 1959. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177706109.

Bhapkar:1960:CBC

[Bha60] V. P. Bhapkar. Confidence bounds connected with anova and manova for balanced and partially balanced incomplete block designs. Annals of Mathematical Statistics, 31(3):741–748, September 1960. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177705800.

Bhapkar:1961:NTP

[Bha61a] V. P. Bhapkar. A nonparametric test for the problem of several samples. Annals of Mathematical Statistics, 32(4):1108–1117, December 1961. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177704849.

Bhapkar:1961:SNM

[Bha61b] V. P. Bhapkar. Some nonparametric median procedures. Annals of Mathematical Statistics, 32(3):846–863, September 1961. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177704978.

Bhapkar:1961:STC

[Bha61c] V. P. Bhapkar. Some tests for categorical data. Annals of Mathematical Statistics, 32(1):72–83, March 1961. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177705140.

Bhat:1961:ADP

[Bha61d] B. R. Bhat. On the asymptotic distribution of the “psi-squared” goodness of fit criteria for Markov chains and Markov sequences.

Annals of Mathematical Statistics, 32(1):49–58, March 1961. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177705138.

Bhat:1961:SPR

[Bha61e] B. R. Bhat. Some properties of regular Markov chains. Annals of Mathematical Statistics, 32(1):59–71, March 1961. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177705139.

Bhattacharya:1962:SPL

[Bha62] P. K. Bhattacharya. Some properties of the least squares estimator in regression analysis when the predictor variables are stochastic. Annals of Mathematical Statistics, 33(4):1365–1374, December 1962. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177704369.

Bhattacharya:1963:ARA

[Bha63] P. K. Bhattacharya. On an analog of regression analysis. Annals of Mathematical Statistics, 34(4):1459–1473, December 1963. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177703878.

Bhat:1964:BSS

[Bha64] B. R. Bhat. Bayes solution of sequential decision problem for Markov dependent observations. Annals of Mathematical Statistics, 35(4): 1656–1662, December 1964. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177700388.

Bhattacharjee:1965:ENN

[Bha65] G. P. Bhattacharjee. Effect of non-normality on Stein’s two sample test. Annals of Mathematical Statistics, 36(2):651–663, April 1965. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177700172.

Bhattacharya:1966:EMM

[Bha66] P. K. Bhattacharya. Estimating the mean of a multivariate normal population with general quadratic loss function. Annals of Mathematical Statistics, 37(6):1819–1824, December 1966. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177699174.

Bhattacharya:1967:EES

[Bha67a] P. K. Bhattacharya. Efficient estimation of a shift parameter from grouped data. Annals of Mathematical Statistics, 38(6):1770–1787, December 1967. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177698611.

Bhattacharyya:1967:AEM

[Bha67b] G. K. Bhattacharyya. Asymptotic efficiency of multivariate normal score test. Annals of Mathematical Statistics, 38(6):1753–1758, December 1967. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177698609.

Bhat:1968:RJJ

[Bha68] B. R. Bhat. Review: J. J. Martin, Bayesian Decision Problems and Markov Chains. Annals of Mathematical Statistics, 39(6):2169–2170, December 1968. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177698057.

Bhattacharyya:1970:RSS

[Bha70] B. B. Bhattacharyya. Reverse submartingale and some functions of order statistics. Annals of Mathematical Statistics, 41(6):2155–2157, December 1970. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177696720.

Bhattacharya:1971:TGP

[Bha71a] P. K. Bhattacharya. Transformations of Gaussian processes to the Brownian motion. Annals of Mathematical Statistics, 42(6):2008– 2017, December 1971. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177693068.

Bhattacharya:1971:RWC

[Bha71b] R. N. Bhattacharya. Rates of weak convergence and asymptotic expansions for classical Central Limit Theorems. Annals of Mathematical Statistics, 42(1):241–259, February 1971. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177693509.

Bhattacharya:1972:ODS

[Bha72] P. K. Bhattacharya. Order of dependence in a stationary normally distributed two-way series. Annals of Mathematical Statistics, 43(6): 1792–1807, December 1972. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177690852.

Blum:1963:CSR

[BHP63] J. R. Blum, D. L. Hanson, and S. V. Parter. On a class of simple random walks. Annals of Mathematical Statistics, 34(2):413– 415, June 1963. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177704152.

Bhuchongkul:1964:CNT

[Bhu64] S. Bhuchongkul. A class of nonparametric tests for independence in bivariate populations. Annals of Mathematical Statistics, 35(1): 138–149, March 1964. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177703735.

Bickel:1964:SAE

[Bic64] Peter J. Bickel. On some alternative estimates for shift in the P - variate one sample problem. Annals of Mathematical Statistics, 35 (3):1079–1090, September 1964. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177703266.

Bickel:1965:CNC

[Bic65a] Peter J. Bickel. Correction notes: Corrections to “On Some Asymptotically Non-Parametric Competitors of Hotelling’s T^{2}”. Annals of Mathematical Statistics, 36(5):1583, October 1965. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177699921. See [Bic65b].

Bickel:1965:SAN

[Bic65b] Peter J. Bickel. On some asymptotically nonparametric competitors of Hotelling’s T^{2}. Annals of Mathematical Statistics, 36(1):160–173, February 1965. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/

1177700280. See corrections [Bic65a].

Bickel:1965:SRE

[Bic65c] Peter J. Bickel. On some robust estimates of location. Annals of Mathematical Statistics, 36(3):847–858, June 1965. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177700058.

Bickel:1969:DFV

[Bic69a] P. J. Bickel. A distribution free version of the Smirnov two sample test in the p-variate case. Annals of Mathematical Statistics, 40(1): 1–23, February 1969. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177697800.

Bickel:1969:RKP

[Bic69b] P. J. Bickel. A remark on the Kolmogorov–Petrovskii criterion. Annals of Mathematical Statistics, 40(3):1086–1090, June 1969. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177697614.

Bickel:1969:TMFb

[Bic69c] Peter J. Bickel. Tests for monotone failure rate. II. Annals of Mathematical Statistics, 40(4):1250–1260, August 1969. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177697500.

Billingsley:1956:ADT

[Bil56] Patrick Billingsley. Asymptotic distributions of two goodness of fit criteria. Annals of Mathematical Statistics, 27(4):1123–1129, December 1956. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177728078.

Billingsley:1961:CSM

[Bil61a] Patrick Billingsley. Corrections to “Statistical Methods in Markow Chains”. Annals of Mathematical Statistics, 32(4):1343, December 1961. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/

1177704874. See [Bil61c].

Billingsley:1961:CTN

[Bil61b] Patrick Billingsley. On the coding theorem for the noiseless channel. Annals of Mathematical Statistics, 32(2):594–601, June 1961.

CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177705065.

Billingsley:1961:SMM

[Bil61c] Patrick Billingsley. Statistical methods in Markov chains. Annals of Mathematical Statistics, 32(1):12–40, March 1961. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http:

//projecteuclid.org/euclid.aoms/1177705136. See corrections

[Bil61a].

Billingsley:1962:LTR

[Bil62] Patrick Billingsley. Limit theorems for randomly selected partial sums. Annals of Mathematical Statistics, 33(1):85–92, March 1962. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177704713.

Bingham:1972:AED

[Bin72] Christopher Bingham. An asymptotic expansion for the distribution of the eigenvalues of a 3 by 3 Wishart matrix. Annals of Mathematical Statistics, 43(5):1498–1506, October 1972. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177692382.

Birnbaum:1942:IMR

[Bir42] Z. W. Birnbaum. An inequality for Mill’s ratio. Annals of Mathematical Statistics, 13(2):245–246, June 1942. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177731611.

Birnbaum:1948:RVC

[Bir48] Z. W. Birnbaum. On random variables with comparable peakedness. Annals of Mathematical Statistics, 19(1):76–81, March 1948. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177730293.

Birnbaum:1949:RSM

[Bir49] Z. W. Birnbaum. Report on the Seattle meeting of the Institute. Annals of Mathematical Statistics, 20(1):151–152, March 1949. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177730109.

Birnbaum:1950:ELT

[Bir50] Z. W. Birnbaum. Effect of linear truncation on a multinormal population. Annals of Mathematical Statistics, 21(2):272–279, June 1950. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177729844.

Birnbaum:1953:DFT

[Bir53a] Z. W. Birnbaum. Distribution-free tests of fit for continuous distribution functions. Annals of Mathematical Statistics, 24(1):1–8, March 1953. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177729078.

Birnbaum:1953:POS

[Bir53b] Z. W. Birnbaum. On the power of a one-sided test of fit for continuous probability functions. Annals of Mathematical Statistics, 24(3):484– 489, September 1953. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177728989.

Birnbaum:1954:ATM

[Bir54] Allan Birnbaum. Admissible tests for the mean of a rectangular distribution. Annals of Mathematical Statistics, 25(1):157–161, March 1954. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177728856.

Birnbaum:1955:CCC

[Bir55] Allan Birnbaum. Characterizations of complete classes of tests of some multiparametric hypotheses, with applications to likelihood ratio tests. Annals of Mathematical Statistics, 26(1):21–36, March 1955. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177728590.

Birnbaum:1958:STV

[Bir58] Allan Birnbaum. Sequential tests for variance ratios and components of variance. Annals of Mathematical Statistics, 29(2):504– 514, June 1958. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177706625.

Birnbaum:1961:FSI

[Bir61a] Allan Birnbaum. On the foundations of statistical inference: Binary experiments. Annals of Mathematical Statistics, 32(2):414– 435, June 1961. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177705050.

Birnbaum:1961:UTE

[Bir61b] Allan Birnbaum. A unified theory of estimation, I. Annals of Mathematical Statistics, 32(1):112–135, March 1961. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177705145.

Birch:1962:AEF

[Bir62] John J. Birch. Approximations for the entropy for functions of Markov chains. Annals of Mathematical Statistics, 33(3):930–938, September 1962. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177704462.

Birch:1964:NPP

[Bir64] M. W. Birch. A new proof of the Pearson–Fisher theorem. Annals of Mathematical Statistics, 35(2):817–824, June 1964. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http:

//projecteuclid.org/euclid.aoms/1177703581. See acknowledg-

ment of priority [Bir65].

Birch:1965:APN

[Bir65] M. W. Birch. Acknowledgment of priority: “A New Proof of the Pearson–Fisher Theorem”. Annals of Mathematical Statistics, 36(1): 344, February 1965. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177700301. See [Bir64].

Bishop:1939:NCA

[Bis39] Morris C. Bishop. A note on computation for analysis of variance. Annals of Mathematical Statistics, 10(4):393–399, December 1939. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177732154.

Bhattacharyya:1968:ADE

[BJ68a] G. K. Bhattacharyya and Richard A. Johnson. Approach to degeneracy and the efficiency of some multivariate tests. Annals of Mathematical Statistics, 39(5):1654–1660, October 1968. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177698147.

Bhattacharyya:1968:NTS

[BJ68b] G. K. Bhattacharyya and Richard A. Johnson. Nonparametric tests for shift at an unknown time point. Annals of Mathematical Statistics, 39(5):1731–1743, October 1968. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177698156.

Bhattacharyya:1970:LRT

[BJ70] G. K. Bhattacharyya and Richard A. Johnson. A layer rank test for ordered bivariate alternatives. Annals of Mathematical Statistics, 41 (4):1296–1310, August 1970. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177696904.

Bradt:1956:SDM

[BJK56] R. N. Bradt, S. M. Johnson, and S. Karlin. On sequential designs for maximizing the sum of n observations. Annals of Mathematical Statistics, 27(4):1060–1074, December 1956. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177728073.

Bradt:1956:DCC

[BK56] Russell N. Bradt and Samuel Karlin. On the design and comparison of certain dichotomous experiments. Annals of Mathematical Statistics, 27(2):390–409, June 1956. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177728265.

Birnbaum:1957:BVM

[BK57a] Z. W. Birnbaum and Orval M. Klose. Bounds for the variance of the Mann–Whitney statistic. Annals of Mathematical Statistics, 28 (4):933–945, December 1957. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177706794.

Blackwell:1957:IPF

[BK57b] David Blackwell and Lambert Koopmans. On the identifiability problem for functions of finite Markov chains. Annals of Mathematical Statistics, 28(4):1011–1015, December 1957. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177706802.

Bochner:1958:LTP

[BK58a] Salomon Bochner and Tatsuo Kawata. A limit theorem for the periodogram. Annals of Mathematical Statistics, 29(4):1198–1208, December 1958. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177706451.

Bradley:1958:AIB

[BK58b] Ralph Allan Bradley and Clyde Young Kramer. Addenda to “Intra-Block Analysis for Factorials in Two-Associate Class Group Divisible Designs”. Annals of Mathematical Statistics, 29(3):933–935, September 1958. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177706556.

Bose:1960:GBS

[BK60] R. C. Bose and Roy R. Kuebler, Jr. A geometry of binary sequences associated with group alphabets in information theory. Annals of Mathematical Statistics, 31(1):113–139, March 1960. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177705992.

Baum:1963:IMA

[BK63] Leonard E. Baum and Melvin L. Katz. On the influence of moments on the asymptotic distribution of sums of random variables. Annals of Mathematical Statistics, 34(3):1042–1044, September 1963. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177704027.

Brock:1970:DCR

[BK70] Archie D. Brock and Richard G. Krutchkoff. Distributions of Z^{γ} and Z^{∗} for complex Z with results applied to the complex normal distribution. Annals of Mathematical Statistics, 41(5):1563–1567, October 1970. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177696800.

Blum:1961:DFT

[BKR61] J. R. Blum, J. Kiefer, and M. Rosenblatt. Distribution free tests of independence based on the sample distribution function. Annals of Mathematical Statistics, 32(2):485–498, June 1961. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177705055.

Borwanker:1971:BMT

[BKR71] J. Borwanker, G. Kallianpur, and B. L. S. Prakasa Rao. The Bernstein–von Mises theorem for Markov processes. Annals of Mathematical Statistics, 42(4):1241–1253, August 1971. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177693237.

Baum:1971:SLR

[BKS71] Leonard E. Baum, M. Katz, and H. H. Stratton. Strong laws for ruled sums. Annals of Mathematical Statistics, 42(2):625–629, April 1971. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177693412.

Bahadur:1955:TCS

[BL55] R. R. Bahadur and E. L. Lehmann. Two comments on “Sufficiency and Statistical Decision Functions”. Annals of Mathematical Statistics, 26(1):139–142, March 1955. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177728604. See [Bah54].

Bickel:1969:UEC

[BL69] P. J. Bickel and E. L. Lehmann. Unbiased estimation in convex families. Annals of Mathematical Statistics, 40(5):1523–1535, October 1969. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177697370.

Blackwell:1946:EW

[Bla46] David Blackwell. On an equation of Wald. Annals of Mathematical Statistics, 17(1):84–87, March 1946. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177731028.

Blackwell:1947:CEU

[Bla47] David Blackwell. Conditional expectation and unbiased sequential estimation. Annals of Mathematical Statistics, 18(1):105–110, March 1947. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177730497.

Blackwell:1951:TL

[Bla51a] David Blackwell. On a theorem of Lyapunov. Annals of Mathematical Statistics, 22(1):112–114, March 1951. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177729699.

Blackwell:1951:TPP

[Bla51b] David Blackwell. On the translation parameter problem for discrete variables. Annals of Mathematical Statistics, 22(3):393–399, September 1951. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177729585.

Blackwell:1953:ECE

[Bla53] David Blackwell. Equivalent comparisons of experiments. Annals of Mathematical Statistics, 24(2):265–272, June 1953. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177729032. See

remarks [LS08].

Blackwell:1954:OS

[Bla54] David Blackwell. On optimal systems. Annals of Mathematical Statistics, 25(2):394–397, June 1954. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177728798.

Blackman:1955:ADF

[Bla55a] Jerome Blackman. On the approximation of a distribution function by an empiric distribution. Annals of Mathematical Statistics, 26(2): 256–267, June 1955. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177728542.

Blackwell:1955:TMP

[Bla55b] David Blackwell. On transient Markov processes with a countable number of states and stationary transition probabilities. Annals of Mathematical Statistics, 26(4):654–658, December 1955. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177728425.

Blackman:1956:EKD

[Bla56] Jerome Blackman. An extension of the Kolmogorov distribution. Annals of Mathematical Statistics, 27(2):513–520, June 1956. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177728274. See correc-

tion [Bla58a].

Blackwell:1957:DVW

[Bla57a] David Blackwell. On discrete variables whose sum is absolutely continuous. Annals of Mathematical Statistics, 28(2):520–521, June 1957. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177706985.

Blasbalg:1957:TFR

[Bla57b] H. Blasbalg. Transformation of the fundamental relationships in sequential analysis. Annals of Mathematical Statistics, 28(4):1024– 1028, December 1957. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177706805.

Blackman:1958:CEK

[Bla58a] Jerome Blackman. Correction to “An Extension of the Kolmogorov Distribution”. Annals of Mathematical Statistics, 29(1):318–322, March 1958. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/

1177706737. See [Bla56].

Blackwell:1958:ACM

[Bla58b] David Blackwell. Another countable Markov process with only instantaneous states. Annals of Mathematical Statistics, 29(1):313– 316, March 1958. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177706735.

Blackwell:1959:ICM

[Bla59] David Blackwell. Infinite codes for memoryless channels. Annals of Mathematical Statistics, 30(4):1242–1244, December 1959. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177706107.

Blackwell:1962:DDP

[Bla62] David Blackwell. Discrete dynamic programming. Annals of Mathematical Statistics, 33(2):719–726, June 1962. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177704593.

Blackwell:1964:MSF

[Bla64] David Blackwell. Memoryless strategies in finite-stage dynamic programming. Annals of Mathematical Statistics, 35(2):863–865, June 1964. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177703586.

Blackwell:1965:DDP

[Bla65] David Blackwell. Discounted dynamic programming. Annals of Mathematical Statistics, 36(1):226–235, February 1965. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177700285.

Blackwell:1968:BSC

[Bla68] David Blackwell. A Borel set not containing a graph. Annals of Mathematical Statistics, 39(4):1345–1347, August 1968. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177698260.

Bleick:1940:LSA

[Ble40] W. E. Bleick. A least squares accumulation theorem. Annals of Mathematical Statistics, 11(2):225–226, June 1940. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177731919.

Bliss:1946:EDS

[Bli46] C. I. Bliss. An experimental design for slope-ratio assays. Annals of Mathematical Statistics, 17(2):232–237, June 1946. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177730984.

Blischke:1962:MEP

[Bli62] W. R. Blischke. Moment estimators for the parameters of a mixture of two binomial distributions. Annals of Mathematical Statistics, 33 (2):444–454, June 1962. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177704571.

Blom:1949:GWF

[Blo49] Gunnar Blom. A generalization of Wald’s fundamental identity. Annals of Mathematical Statistics, 20(3):439–444, September 1949. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177729996.

Blomqvist:1950:MDB

[Blo50] Nils Blomqvist. On a measure of dependence between two random variables. Annals of Mathematical Statistics, 21(4):593–600, December 1950. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177729754.

Blomqvist:1951:STB

[Blo51] Nils Blomqvist. Some tests based on dichotomization. Annals of Mathematical Statistics, 22(3):362–371, September 1951. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177729582.

Block:1957:EET

[Blo57] H. D. Block. Estimates of error for two modifications of the Robbins– Monro stochastic approximation process. Annals of Mathematical Statistics, 28(4):1003–1010, December 1957. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177706801.

Block:1970:EEL

[Blo70] H. W. Block. Error estimation for a limit theorem for dependent random variables. Annals of Mathematical Statistics, 41(4):1334– 1338, August 1970. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177696907.

Block:1971:ACS

[Blo71] H. W. Block. Accuracy of convergence of sums of dependent random variables with variances not necessarily finite. Annals of Mathematical Statistics, 42(6):2134–2138, December 1971. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177693080.

Breiman:1964:CEZ

[BLS64] Leo Breiman, Lucien Le Cam, and Lorraine Schwartz. Consistent estimates and zero-one sets. Annals of Mathematical Statistics, 35(1): 157–161, March 1964. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177703737.

Blum:1954:AMW

[Blu54a] Julius R. Blum. Approximation methods which converge with probability one. Annals of Mathematical Statistics, 25(2):382–386, June 1954. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177728794.

Blum:1954:MSA

[Blu54b] Julius R. Blum. Multidimensional stochastic approximation methods. Annals of Mathematical Statistics, 25(4):737–744, December 1954. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177728659.

Blum:1955:CED

[Blu55] J. R. Blum. On the convergence of empiric distribution functions. Annals of Mathematical Statistics, 26(3):527–529, September 1955. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177728499.

Blum:1960:WG

[Blu60] J. R. Blum. Werner Gautschi, 1927–1959. Annals of Mathematical Statistics, 31(3):557, September 1960. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177705782.

Blumenthal:1963:ANT

[Blu63] Saul Blumenthal. The asymptotic normality of two test statistics associated with the two-sample problem. Annals of Mathematical Statistics, 34(4):1513–1523, December 1963. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177703883.

Blumenthal:1966:CSSa

[Blu66a] Saul Blumenthal. Contributions to sample spacings theory, i: Limit distributions of sums of ratios of spacings. Annals of Mathematical Statistics, 37(4):904–924, August 1966. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177699372.

Blumenthal:1966:CSSb

[Blu66b] Saul Blumenthal. Contributions to sample spacings theory, II: Tests of the parametric goodness of fit and two-sample problems. Annals of Mathematical Statistics, 37(4):925–939, August 1966. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177699373.

Blumenthal:1967:RJN

[Blu67a] R. M. Blumenthal. Review: J. Neveu, Amiel Feinstein, Mathematical Foundations of the Calculus of Probability. Annals of Mathematical Statistics, 38(2):624–625, April 1967. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177698988.

Blumenthal:1967:LTF

[Blu67b] Saul Blumenthal. Limit theorems for functions of shortest two-sample spacings and a related test. Annals of Mathematical Statistics, 38(1): 108–116, February 1967. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177699061.

Blyth:1951:MSD

[Bly51] Colin R. Blyth. On minimax statistical decision procedures and their admissibility. Annals of Mathematical Statistics, 22(1):22–42, March 1951. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177729690.

Blyth:1958:NRE

[Bly58] Colin R. Blyth. Note on relative efficiency of tests. Annals of Mathematical Statistics, 29(3):898–903, September 1958. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177706547.

Blyth:1959:NEI

[Bly59] Colin R. Blyth. Note on estimating information. Annals of Mathematical Statistics, 30(1):71–79, March 1959. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177706360.

Blyth:1970:IDM

[Bly70] Colin R. Blyth. On the inference and decision models of statistics. Annals of Mathematical Statistics, 41(3):1034–1058, June 1970. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177696980.

Bush:1953:SMA

[BM53] Robert R. Bush and Frederick Mosteller. A stochastic model with applications to learning. Annals of Mathematical Statistics, 24(4):559– 585, December 1953. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177728914.

Birnbaum:1958:DFU

[BM58a] Z. W. Birnbaum and R. C. McCarty. A distribution-free upper confidence bound for Pr{Y < X}, based on independent samples of X and Y. Annals of Mathematical Statistics, 29(2):558–562, June 1958. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177706631.

Box:1958:NGR

[BM58b] G. E. P. Box and Mervin E. Muller. A note on the generation of random normal deviates. Annals of Mathematical Statistics, 29(2): 610–611, June 1958. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177706645;   http://www.jstor.org/stable/2237361.

Beale:1959:SMS

[BM59a] E. M. L. Beale and C. L. Mallows. Scale mixing of symmetric distributions with zero means. Annals of Mathematical Statistics, 30 (4):1145–1151, December 1959. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177706099.

Bose:1959:LAA

[BM59b] R. C. Bose and Dale M. Mesner. On linear associative algebras corresponding to association schemes of partially balanced designs. Annals of Mathematical Statistics, 30(1):21–38, March 1959. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177706356.

Birnbaum:1961:SMC

[BM61] Z. W. Birnbaum and Albert W. Marshall. Some multivariate Chebyshev inequalities with extensions to continuous parameter processes. Annals of Mathematical Statistics, 32(3):687–703, September 1961. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177704964.

Blumenson:1963:PGR

[BM63] L. E. Blumenson and K. S. Miller. Properties of generalized Rayleigh distributions. Annals of Mathematical Statistics, 34(3):903–910, September 1963. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177704013.

Barlow:1964:BDMa

[BM64a] Richard E. Barlow and Albert W. Marshall. Bounds for distributions with monotone hazard rate, I. Annals of Mathematical Statistics, 35 (3):1234–1257, September 1964. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177703281.

Barlow:1964:BDMb

[BM64b] Richard E. Barlow and Albert W. Marshall. Bounds for distributions with monotone hazard rate, II. Annals of Mathematical Statistics, 35 (3):1258–1274, September 1964. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177703282.

Barton:1965:SAR

[BM65] D. E. Barton and C. L. Mallows. Some aspects of the random sequence. Annals of Mathematical Statistics, 36(1):236–260, February 1965. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177700286.

Barlow:1963:PPD

[BMP63] Richard E. Barlow, Albert W. Marshall, and Frank Proschan. Properties of probability distributions with monotone hazard rate. Annals of Mathematical Statistics, 34(2):375–389, June 1963. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177704147.

Bell:1966:GCT

[BMT66] C. B. Bell, J. M. Moser, and Rory Thompson. Goodness criteria for two-sample distribution-free tests. Annals of Mathematical Statistics, 37(1):133–142, February 1966. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177699604.

Bronowski:1945:VMT

[BN45] J. Bronowski and J. Neyman. The variance of the measure of a two-dimensional random set. Annals of Mathematical Statistics, 16 (4):330–341, December 1945. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177731059.

Brainerd:1961:NSB

[BN61] B. Brainerd and T. V. Narayana. A note on simple binomial sampling plans. Annals of Mathematical Statistics, 32(3):906–908, September 1961. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177704987.

Barndorff-Nielsen:1963:LBE

[BN63] Ole Barndorff-Nielsen. On the limit behaviour of extreme order statistics. Annals of Mathematical Statistics, 34(3):992–1002, September 1963. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177704022.

Boas:1949:RPD

[Boa49] R. P. Boas, Jr. Representation of probability distributions by Charlier series. Annals of Mathematical Statistics, 20(3):376–392, September 1949. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177729990.

Boas:1967:LBI

[Boa67] R. P. Boas, Jr. Lipschitz behavior and integrability of characteristic functions. Annals of Mathematical Statistics, 38(1):32–36, February 1967. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177699056.

Boes:1966:EMD

[Boe66] D. C. Boes. On the estimation of mixing distributions. Annals of Mathematical Statistics, 37(1):177–188, February 1966. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177699607.

Boes:1969:CPC

[Boe69] Ardel J. Boes. Conditional probability on Σ-complete Boolean algebras. Annals of Mathematical Statistics, 40(3):970–978, June 1969.

CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177697601.

Bohrer:1968:OCS

[Boh68] Robert Bohrer. Operating characteristics of some sequential design rules. Annals of Mathematical Statistics, 39(4):1176–1185, August 1968. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177698241.

Boldyreff:1934:MFM

[Bol34] John W. Boldyreff. Mathematical foundation for a method of statistical analysis of household budgets. Annals of Mathematical Statistics, 5(3):216–226, September 1934. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177732677.

Bondar:1972:SDE

[Bon72] James V. Bondar. Structural distributions without exact transitivity. Annals of Mathematical Statistics, 43(1):326–339, February 1972. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177692726.

Book:1972:LDP

[Boo72] Stephen A. Book. Large deviation probabilities for weighted sums. Annals of Mathematical Statistics, 43(4):1221–1234, August 1972. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177692474.

Borgman:1967:RHF

[Bor67] Leon E. Borgman. Random hydrodynamic forces on objects. Annals of Mathematical Statistics, 38(1):37–51, February 1967. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177699057.

Bose:1949:NFI

[Bos49] R. C. Bose. A note on Fisher’s inequality for balanced incomplete block designs. Annals of Mathematical Statistics, 20(4):619–620, December 1949. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177729958.

Boswell:1966:ETT

[Bos66] M. T. Boswell. Estimating and testing trend in a stochastic process of Poisson type. Annals of Mathematical Statistics, 37(6):1564–1573, December 1966. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177699148.

Boudreau:1968:FFM

[Bou68] Paul E. Boudreau. Functions of finite Markov chains and exponential type processes. Annals of Mathematical Statistics, 39(3):1020– 1029, June 1968. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177698334.

Bowker:1944:NCP

[Bow44] Albert H. Bowker. Note on consistency of a proposed test for the problem of two samples. Annals of Mathematical Statistics, 15(1): 98–101, March 1944. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177731320.

Bowker:1946:CFT

[Bow46] Albert H. Bowker. Computation of factors for tolerance limits on a normal distribution when the sample is large. Annals of Mathematical Statistics, 17(2):238–240, June 1946. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177730985.

Bowker:1947:NM

[Bow47] Albert H. Bowker. On the norm of a matrix. Annals of Mathematical Statistics, 18(2):285–288, June 1947. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177730446.

Box:1954:STQa

[Box54a] G. E. P. Box. Some theorems on quadratic forms applied in the study of analysis of variance problems, I. Effect of inequality of variance in the one-way classification. Annals of Mathematical Statistics, 25(2): 290–302, June 1954. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177728786.

Box:1954:STQb

[Box54b] G. E. P. Box. Some theorems on quadratic forms applied in the study of analysis of variance problems, II. Effects of inequality of variance and of correlation between errors in the two-way classification. Annals of Mathematical Statistics, 25(3):484–498, September 1954. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177728717.

Boyd:1958:BMR

[Boy58] A. V. Boyd. Bounds for Mills’ ratio for the Type III population. Annals of Mathematical Statistics, 29(3):926–929, September 1958. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177706554.

Boylan:1971:EM

[Boy71] Edward S. Boylan. Equiconvergence of martingales. Annals of Mathematical Statistics, 42(2):552–559, April 1971. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177693405.

Boza:1971:AOT

[Boz71] Luis B. Boza. Asymptotically optimal tests for finite Markov chains. Annals of Mathematical Statistics, 42(6):1992–2007, December 1971. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177693067.

Birnbaum:1958:SDR

[BP58] Z. W. Birnbaum and Ronald Pyke. On some distributions related to the statistic D_{n}^{+}. Annals of Mathematical Statistics, 29(1):179– 187, March 1958. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/

1177706714. See alternative proof [Kui59].

Barlow:1964:CRP

[BP64] Richard E. Barlow and Frank Proschan. Comparison of replacement policies, and renewal theory implications. Annals of Mathematical Statistics, 35(2):577–589, June 1964. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177703557.

Bhuchongkul:1965:ECL

[BP65a] Subha Bhuchongkul and Madan L. Puri. On the estimation of contrasts in linear models. Annals of Mathematical Statistics, 36(1):

198–202, February 1965. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177700282.

Borges:1965:OPE

[BP65b] R. Borges and J. Pfanzagl. One-parameter exponential families generated by transformation groups. Annals of Mathematical Statistics, 36(1):261–271, February 1965. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177700287.

Barlow:1966:ILC

[BP66a] Richard E. Barlow and Frank Proschan. Inequalities for linear combinations of order statistics from restricted families. Annals of Mathematical Statistics, 37(6):1574–1592, December 1966. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177699149.

Barlow:1966:TCL

[BP66b] Richard E. Barlow and Frank Proschan. Tolerance and confidence limits for classes of distributions based on failure rate. Annals of Mathematical Statistics, 37(6):1593–1601, December 1966. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177699150. See correc-

tion [BP67a].

Baum:1966:SIP

[BP66c] Leonard E. Baum and Ted Petrie. Statistical inference for probabilistic functions of finite state Markov chains. Annals of Mathematical Statistics, 37(6):1554–1563, December 1966. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177699147.

Barlow:1967:CNC

[BP67a] Richard E. Barlow and Frank Proschan. Correction note: Correction to “Tolerance and Confidence Limits for Classes of Distributions Based on Failure Rate”. Annals of Mathematical Statistics, 38(3): 950, June 1967. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/

1177698895. See [BP66b].

Bartoo:1967:OAT

[BP67b] James B. Bartoo and Prem S. Puri. On optimal asymptotic tests of composite statistical hypotheses. Annals of Mathematical Statistics, 38(6):1845–1852, December 1967. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177698617.

Bildikar:1968:MET

[BP68] Sheela Bildikar and G. P. Patil. Multivariate exponential-type distributions. Annals of Mathematical Statistics, 39(4):1316–1326, August 1968. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177698257.

Barlow:1969:NTM

[BP69] R. E. Barlow and F. Proschan. A note on tests for monotone failure rate based on incomplete data. Annals of Mathematical Statistics, 40 (2):595–600, April 1969. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177697727.

Blum:1972:NZO

[BP72] Julius R. Blum and Pramod K. Pathak. A note on the Zero–One Law. Annals of Mathematical Statistics, 43(3):1008–1009, June 1972. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177692564.

Baum:1970:MTO

[BPSW70] Leonard E. Baum, Ted Petrie, George Soules, and Norman Weiss. A maximization technique occurring in the statistical analysis of probabilistic functions of Markov chains. Annals of Mathematical Statistics, 41(1):164–171, February 1970. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177697196.

Bahadur:1950:PGM

[BR50] Raghu Raj Bahadur and Herbert Robbins. The problem of the greater mean. Annals of Mathematical Statistics, 21(4):469–487; 22, 310 (1951), December 1950. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.

aoms/1177729746. See correction [BR51].

Bahadur:1951:CPG

[BR51] Raghu Raj Bahadur and Herbert Robbins. Correction to “The Problem of the Greater Mean”. Annals of Mathematical Statistics, 22 (2):310, June 1951. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177729656. See [BR50].

Birnbaum:1954:DFS

[BR54] Z. W. Birnbaum and H. Rubin. On distribution-free statistics. Annals of Mathematical Statistics, 25(3):593–598, September 1954. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177728728.

Burke:1958:MFM

[BR58] C. J. Burke and M. Rosenblatt. A Markovian function of a Markov chain. Annals of Mathematical Statistics, 29(4):1112–1122, December 1958. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177706444.

Bahadur:1960:DSM

[BR60] R. R. Bahadur and R. Ranga Rao. On deviations of the sample mean. Annals of Mathematical Statistics, 31(4):1015–1027, December 1960. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177705674.

Blum:1963:ME

[BR63] J. R. Blum and Judah Rosenblatt. On multistage estimation. Annals of Mathematical Statistics, 34(4):1452–1458, December 1963. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177703877.

Blum:1964:RSS

[BR64] J. R. Blum and Judah Rosenblatt. On random sampling from a stochastic process. Annals of Mathematical Statistics, 35(4):1713– 1717, December 1964. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177700393.

Blum:1967:PPI

[BR67] J. R. Blum and Judah Rosenblatt. On partial a priori information in statistical inference. Annals of Mathematical Statistics, 38(6):1671– 1678, December 1967. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177698602.

Blum:1969:FPE

[BR69] J. R. Blum and Judah Rosenblatt. On fixed precision estimation in time series. Annals of Mathematical Statistics, 40(3):1021–1032, June 1969. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177697606.

Bradley:1952:CNU

[Bra52] Ralph Allan Bradley. Corrections for nonnormality in the use of the two-sample t- and F -test at high significance levels. Annals of Mathematical Statistics, 23(1):103–113, March 1952. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177729488.

Bradu:1965:MEA

[Bra65] Dan Bradu. Main-effect analysis of the general non-orthogonal layout with any number of factors. Annals of Mathematical Statistics, 36(1): 88–97, February 1965. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177700273.

Breiman:1957:CTK

[Bre57a] Leo Breiman. A counterexample to a theorem of Kolmogorov. Annals of Mathematical Statistics, 28(3):811–814, September 1957. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177706900.

Breiman:1957:IET

[Bre57b] Leo Breiman. The individual ergodic theorem of information theory. Annals of Mathematical Statistics, 28(3):809–811, September 1957. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177706899. See

correction [Bre60a].

Breiman:1957:TMC

[Bre57c] Leo Breiman. On transient Markov chains with application to the uniqueness problem for Markov processes. Annals of Mathematical Statistics, 28(2):499–503, June 1957. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177706979.

Breiman:1958:ETC

[Bre58a] Leo Breiman. An elementary theorem concerning stationary ergodic processes. Annals of Mathematical Statistics, 29(2):592–595, June 1958. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177706638.

Breiman:1958:TAM

[Bre58b] Leo Breiman. Transient atomic Markov chains with a denumerable number of states. Annals of Mathematical Statistics, 29(1):212–218, March 1958. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177706718.

Breiman:1960:CNC

[Bre60a] Leo Breiman. Correction notes: Correction to “The Individual Ergodic Theorem of Information Theory”. Annals of Mathematical Statistics, 31(3):809–810, September 1960. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177705812. See [Bre57b].

Breiman:1960:SLL

[Bre60b] Leo Breiman. The strong law of large numbers for a class of Markov chains. Annals of Mathematical Statistics, 31(3):801–803, September 1960. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177705810.

Breiman:1963:PTT

[Bre63] Leo Breiman. The Poisson tendency in traffic distribution. Annals of Mathematical Statistics, 34(1):308–311, March 1963. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177704267.

Breiman:1968:DLI

[Bre68] Leo Breiman. A delicate law of the iterated logarithm for nondecreasing stable processes. Annals of Mathematical Statistics, 39 (6):1818–1824, December 1968. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid. org/euclid.aoms/1177698014. See addendum [Bre69] and erratum [Bre70].

Breiman:1969:AAD

[Bre69] Leo Breiman. Addendum: Addendum to “A Delicate Law of the Iterated Logarithm for Non-Decreasing Stable Processes”. Annals of Mathematical Statistics, 40(5):1855, October 1969. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177697403. See [Bre68].

Breiman:1970:CNC

[Bre70] Leo Breiman. Correction notes: Correction to “A Delicate Law of the Iterated Logarithm for Non-Decreasing Stable Processes”. Annals of Mathematical Statistics, 41(3):1126, June 1970. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177696996. See [Bre68].

Bridger:1938:NRF

[Bri38] Clyde A. Bridger. Note on regression functions in the case of three second order random variables. Annals of Mathematical Statistics, 9 (4):309–313, December 1938. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177732288.

Brillinger:1962:EBD

[Bri62] David R. Brillinger. Examples bearing on the definition of fiducial probability with a bibliography. Annals of Mathematical Statistics, 33 (4):1349–1355, December 1962. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177704367.

Brillinger:1963:NSC

[Bri63a] David R. Brillinger. Necessary and sufficient conditions for a statistical problem to be invariant under a Lie group. Annals of Mathematical Statistics, 34(2):492–500, June 1963. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177704160.

Brillinger:1963:NRU

[Bri63b] David R. Brillinger. A note on the re-use of samples. Annals of Mathematical Statistics, 34(1):341–343, March 1963. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177704276.

Brillinger:1965:IP

[Bri65] David R. Brillinger. An introduction to polyspectra. Annals of Mathematical Statistics, 36(5):1351–1374, October 1965. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177699896.

Brillinger:1972:NSS

[Bri72] David R. Brillinger. On the number of solutions of systems of random equations. Annals of Mathematical Statistics, 43(2):534–540, April 1972. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177692634.

Blackwell:1963:NEE

[BRN63] D. Blackwell and C. Ryll-Nardzewski. Non-existence of everywhere proper conditional distributions. Annals of Mathematical Statistics, 34(1):223–225, March 1963. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177704259.

Brown:1933:SCP

[Bro33] George Middleton Brown. On sampling from compound populations. Annals of Mathematical Statistics, 4(4):288–342, November 1933. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177732758.

Brown:1939:PTE

[Bro39] George W. Brown. On the power of the L_{1} test for equality of several variances. Annals of Mathematical Statistics, 10(2):119–128, June 1939. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177732210.

Brown:1940:NUP

[Bro40a] A. W. Brown. A note on the use of a Pearson Type III function in renewal theory. Annals of Mathematical Statistics, 11(4):448–453, December 1940. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177731831.

Brown:1940:RCC

[Bro40b] George W. Brown. Reduction of a certain class of composite statistical hypotheses. Annals of Mathematical Statistics, 11(3):254–270, September 1940. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177731866.

Brookner:1945:COA

[Bro45] Ralph J. Brookner. Choice of one among several statistical hypotheses. Annals of Mathematical Statistics, 16(3):221–242, September 1945. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177731087.

Brown:1947:DF

[Bro47a] George W. Brown. Discriminant functions. Annals of Mathematical Statistics, 18(4):514–528, December 1947. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177730342.

Brown:1947:SSE

[Bro47b] George W. Brown. On small-sample estimation. Annals of Mathematical Statistics, 18(4):582–585, December 1947. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177730349.

Bross:1959:NAS

[Bro59] Irwin D. J. Bross. Note on an application of the Schumann–Bradley table. Annals of Mathematical Statistics, 30(2):581–583, June 1959. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177706271.

Brown:1963:EC

[Bro63] Thomas A. Brown. Entropy and conjugacy. Annals of Mathematical Statistics, 34(1):226–232, March 1963. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177704260.

Brown:1964:SSC

[Bro64] L. Brown. Sufficient statistics in the case of independent random variables. Annals of Mathematical Statistics, 35(4):1456–1474, December 1964. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177700373.

Brown:1965:IMD

[Bro65] Barry W. Brown. On the iterative method of dynamic programming on a finite space discrete time Markov process. Annals of Mathematical Statistics, 36(4):1279–1285, August 1965. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177699999.

Brown:1966:AIE

[Bro66] Lawrence David Brown. On the admissibility of invariant estimators of one or more location parameters. Annals of Mathematical Statistics, 37(5):1087–1136, October 1966. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177699259.

Brown:1967:CLS

[Bro67] L. Brown. The conditional level of Student’s t test. Annals of Mathematical Statistics, 38(4):1068–1071, August 1967. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177698776.

Brown:1968:IUE

[Bro68] L. Brown. Inadmissibility of the usual estimators of scale parameters in problems with unknown location and scale parameters. Annals of Mathematical Statistics, 39(1):29–48, February 1968. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177698503.

Brown:1969:MSR

[Bro69] Bruce M. Brown. Moments of a stopping rule related to the Central Limit Theorem. Annals of Mathematical Statistics, 40(4):1236–1249, August 1969. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177697499.

Brown:1970:CFM

[Bro70a] B. M. Brown. Characteristic functions, moments, and the Central Limit Theorem. Annals of Mathematical Statistics, 41(2):658– 664, April 1970. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177697109.

Brown:1970:CDS

[Bro70b] Mark Brown. Convergence in distribution of stochastic integrals. Annals of Mathematical Statistics, 41(3):829–842, June 1970. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177696961.

Brown:1970:MGE

[Bro70c] Mark Brown. An M/G/∞ estimation problem. Annals of Mathematical Statistics, 41(2):651–654, April 1970. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177697107.

Brown:1970:PPP

[Bro70d] Mark Brown. A property of Poisson processes and its application to macroscopic equilibrium of particle systems. Annals of Mathematical Statistics, 41(6):1935–1941, December 1970. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http:

//projecteuclid.org/euclid.aoms/1177696694. See correction

[Bro71f].

Brockwell:1971:RCL

[Bro71a] Peter Brockwell. Review: Chin Long Chiang, Introduction to Stochastic Processes in Biostatistics. Annals of Mathematical Statistics, 42(3):1140–1141, June 1971. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177693351.

Brown:1971:MCL

[Bro71b] B. M. Brown. Martingale Central Limit Theorems. Annals of Mathematical Statistics, 42(1):59–66, February 1971. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177693494.

Brown:1971:NCM

[Bro71c] B. M. Brown. A note on convergence of moments. Annals of Mathematical Statistics, 42(2):777–779, April 1971. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177693430.

Brown:1971:AER

[Bro71d] L. D. Brown. Admissible estimators, recurrent diffusions, and insoluble boundary value problems. Annals of Mathematical Statistics, 42 (3):855–903, June 1971. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177693318.

Brown:1971:NLA

[Bro71e] L. D. Brown. Non-local asymptotic optimality of appropriate likelihood ratio tests. Annals of Mathematical Statistics, 42(4):1206– 1240, August 1971. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177693236.

Brown:1971:CNC

[Bro71f] Mark Brown. Correction notes: Correction to “A Property of Poisson Processes and Its Application to Macroscopic Equilibrium of Particle Systems”. Annals of Mathematical Statistics, 42(5):1777, October 1971. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/

1177693183. See [Bro70d].

Brown:1971:DPP

[Bro71g] Mark Brown. Discrimination of Poisson processes. Annals of Mathematical Statistics, 42(2):773–776, April 1971. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177693429.

Brunk:1955:MLE

[Bru55] H. D. Brunk. Maximum likelihood estimates of monotone parameters. Annals of Mathematical Statistics, 26(4):607–616, December 1955. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177728420.

Brunk:1958:EPR

[Bru58] H. D. Brunk. On the estimation of parameters restricted by inequalities. Annals of Mathematical Statistics, 29(2):437–454, June 1958. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177706621.

Brunk:1962:RDB

[Bru62] H. D. Brunk. On the range of the difference between hypothetical distribution function and Pyke’s modified empirical distribution function. Annals of Mathematical Statistics, 33(2):525–532, June 1962. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177704578.

Brunk:1965:CEG

[Bru65] H. D. Brunk. Conditional expectation given a σ-lattice and applications. Annals of Mathematical Statistics, 36(5):1339–1350, October 1965. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177699895.

Bruckner:1969:IHR

[Bru69] L. A. Bruckner. Interpolation of homogeneous random fields on discrete groups. Annals of Mathematical Statistics, 40(1):251–258, February 1969. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177697820.

Bruckner:1971:MAH

[Bru71] L. A. Bruckner. Moving averages of homogeneous random fields. Annals of Mathematical Statistics, 42(6):2147–2149, December 1971. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177693083.

Birnbaum:1947:GTI

[BRZ47] Z. W. Birnbaum, J. Raymond, and H. S. Zuckerman. A generalization of Tshebyshev’s inequality to two dimensions. Annals of Mathematical Statistics, 18(1):70–79, March 1947. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177730493.

Bechhofer:1954:SSMb

[BS54] Robert E. Bechhofer and Milton Sobel. A single-sample multiple decision procedure for ranking variances of normal populations. Annals of Mathematical Statistics, 25(2):273–289, June 1954. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177728785.

Bahadur:1956:NCS

[BS56] R. R. Bahadur and Leonard J. Savage. The nonexistence of certain statistical procedures in nonparametric problems. Annals of Mathematical Statistics, 27(4):1115–1122, December 1956. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177728077.

Bose:1964:BUT

[BS64] R. C. Bose and J. N. Srivastava. On a bound useful in the theory of factorial designs and error correcting codes. Annals of Mathematical Statistics, 35(1):408–414, March 1964. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177703764.

Berk:1968:IRO

[BS68] Robert H. Berk and I. Richard Savage. The information in a rankorder and the stopping time of some associated SPRT’s. Annals of Mathematical Statistics, 39(5):1661–1674, October 1968. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177698148.

Binmore:1969:NCF

[BS69] K. G. Binmore and H. H. Stratton. A note on characteristic functions. Annals of Mathematical Statistics, 40(1):303–307, February 1969. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177697826.

Boonyasombut:1970:AID

[BS70] Virool Boonyasombut and Jesse M. Shapiro. The accuracy of infinitely divisible approximations to sums of independent variables with application to stable laws. Annals of Mathematical Statistics, 41(1):237–250, February 1970. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177697205.

Bhattacharya:1972:SPR

[BS72] P. K. Bhattacharya and Robert P. Smith. Sequential probability ratio test for the mean value function of a Gaussian process. Annals of Mathematical Statistics, 43(6):1861–1873, December 1972. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177690857.

Bose:1953:CGD

[BSB53] R. C. Bose, S. S. Shrikhande, and K. N. Bhattacharya. On the construction of group divisible incomplete block designs. Annals of Mathematical Statistics, 24(2):167–195, June 1953. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177729027.

Brown:1946:SDS

[BT46] George W. Brown and John W. Tukey. Some distributions of sample means. Annals of Mathematical Statistics, 17(1):1–12, March 1946. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177731017;
http://projecteuclid.org/euclid.aop/1176993612.

Birnbaum:1951:OSC

[BT51] Z. W. Birnbaum and Fred H. Tingey. One-sided confidence contours for probability distribution functions. Annals of Mathematical Statistics, 22(4):592–596, December 1951. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177729550.

Box:1965:MPB

[BT65] G. E. P. Box and George C. Tiao. Multiparameter problems from a Bayesian point of view. Annals of Mathematical Statistics, 36 (5):1468–1482, October 1965. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177699906.

Bucy:1965:RS

[Buc65] R. S. Bucy. Recurrent sets. Annals of Mathematical Statistics, 36(2): 535–545, April 1965. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177700164.

Buehler:1959:SVC

[Bue59] Robert J. Buehler. Some validity criteria for statistical inferences. Annals of Mathematical Statistics, 30(4):845–863, December 1959. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177706070.

Buehler:1965:LPD

[Bue65] Robert J. Buehler. The limit of the n th power of a density. Annals of Mathematical Statistics, 36(6):1878–1882, December 1965. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177699823.

Buhler:1967:SBP

[Büh67a] Wolfgang J. Bühler. Slowly branching processes. Annals of Mathematical Statistics, 38(3):919–921, June 1967. CODEN AASTAD.

ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177698886.

Buhler:1967:TTW

[Büh67b] Wolfgang J. Bühler. The treatment of ties in the Wilcoxon test. Annals of Mathematical Statistics, 38(2):519–522, April 1967. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177698968.

Burgess:1932:SAM

[Bur32] Robert W. Burgess. A statistical approach to mathematical formulation of demand-supply-price relationships. Annals of Mathematical Statistics, 3(1):10–19, February 1932. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177732928.

Burr:1942:CFF

[Bur42] Irving W. Burr. Cumulative frequency functions. Annals of Mathematical Statistics, 13(2):215–232, June 1942. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177731607.

Burr:1955:CES

[Bur55] Irving W. Burr. Calculation of exact sampling distribution of ranges from a discrete population. Annals of Mathematical Statistics, 26 (3):530–532, September 1955. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177728500. See correction [Bur67].

Burkholder:1956:CSA

[Bur56] D. L. Burkholder. On a class of stochastic approximation processes. Annals of Mathematical Statistics, 27(4):1044–1059, December 1956. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177728072.

Burkholder:1959:EBA

[Bur59] D. L. Burkholder. On the existence of a best approximation of one distribution function by another of a given type. Annals of Mathematical Statistics, 30(3):738–742, September 1959. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177706202.

Burkholder:1960:EMC

[Bur60] D. L. Burkholder. Effect on the minimal complete class of tests of changes in the testing problem. Annals of Mathematical Statistics, 31 (2):325–331, June 1960. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177705896.

Burkholder:1961:SUC

[Bur61] D. L. Burkholder. Sufficiency in the undominated case. Annals of Mathematical Statistics, 32(4):1191–1200, December 1961. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177704859.

Burkholder:1962:OSS

[Bur62a] D. L. Burkholder. On the order structure of the set of sufficient subfields. Annals of Mathematical Statistics, 33(2):596–599, June 1962. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177704584.

Burkholder:1962:SCE

[Bur62b] D. L. Burkholder. Successive conditional expectations of an integrable function. Annals of Mathematical Statistics, 33(3):887–893, September 1962. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177704457.

Burr:1963:DTS

[Bur63] E. J. Burr. Distribution of the two-sample Cramér–von Mises criterion for small equal samples. Annals of Mathematical Statistics, 34 (1):95–101, March 1963. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177704245.

Burke:1964:DDT

[Bur64a] P. J. Burke. The dependence of delays in tandem queues. Annals of Mathematical Statistics, 35(2):874–875, June 1964. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177703590.

Burr:1964:SSD

[Bur64b] E. J. Burr. Small-sample distributions of the two-sample Cramér–von Mises’ W^{2} and Watson’s U^{2}. Annals of Mathematical Statistics, 35 (3):1091–1098, September 1964. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177703267.

Burke:1965:UET

[Bur65] George Burke. A uniform ergodic theorem. Annals of Mathematical Statistics, 36(6):1853–1858, December 1965. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177699815.

Burkholder:1966:MT

[Bur66] D. L. Burkholder. Martingale transforms. Annals of Mathematical Statistics, 37(6):1494–1504, December 1966. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177699141.

Burr:1967:CNC

[Bur67] Irving W. Burr. Correction note: Correction to “Calculation of Exact Sampling Distribution of Ranges from a Discrete Population”. Annals of Mathematical Statistics, 38(1):280, February 1967. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177699084. See [Bur55].

Burke:1968:OPS

[Bur68a] P. J. Burke. The output process of a stationary M/M/s queueing system. Annals of Mathematical Statistics, 39(4):1144–1152, August 1968. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177698238.

Burkholder:1968:ISS

[Bur68b] D. L. Burkholder. Independent sequences with the Stein property. Annals of Mathematical Statistics, 39(4):1282–1288, August 1968. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177698253.

Burdick:1972:DFT

[Bur72a] D. L. Burdick. Distribution free tests for symmetry based on the number of positive sums. Annals of Mathematical Statistics, 43(2): 428–438, April 1972. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177692623.

Burdick:1972:NSR

[Bur72b] David L. Burdick. A note on symmetric random variables. Annals of Mathematical Statistics, 43(6):2039–2040, December 1972. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177690880.

Bush:1952:GTD

[Bus52a] K. A. Bush. A generalization of a theorem due to MacNeish. Annals of Mathematical Statistics, 23(2):293–295, June 1952. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177729449.

Bush:1952:OAI

[Bus52b] K. A. Bush. Orthogonal arrays of index unity. Annals of Mathematical Statistics, 23(3):426–434, September 1952. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177729387.

Butler:1969:TSU

[But69] Calvin C. Butler. A test for symmetry using the sample distribution function. Annals of Mathematical Statistics, 40(6):2209–2210, December 1969. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177697302.

Berger:1949:DH

[BW49] Agnes Berger and Abraham Wald. On distinct hypotheses. Annals of Mathematical Statistics, 20(1):104–109, March 1949. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177730096.

Bankier:1957:CVA

[BW57a] J. D. Bankier and R. E. Walpole. Components of variance analysis for proportional frequencies. Annals of Mathematical Statistics, 28 (3):742–753, September 1957. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177706886.

Blum:1957:CCT

[BW57b] J. R. Blum and Lionel Weiss. Consistency of certain two-sample tests. Annals of Mathematical Statistics, 28(1):242–246, March 1957. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177707048.

Burkholder:1963:OPA

[BW63] D. L. Burkholder and R. A. Wijsman. Optimum properties and admissibility of sequential tests. Annals of Mathematical Statistics, 34 (1):1–17, March 1963. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177704238.

Bain:1964:NTE

[BW64a] Lee J. Bain and David L. Weeks. A note on the truncated exponential distribution. Annals of Mathematical Statistics, 35(3):1366–1367, September 1964. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177703298.

Breiman:1964:CPL

[BW64b] Leo Breiman and Zivia S. Wurtele. Convergence properties of a learning algorithm. Annals of Mathematical Statistics, 35(4):1819–1822, December 1964. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177700406.

Basu:1967:NNT

[BW67a] A. P. Basu and George Woodworth. A note on nonparametric tests for scale. Annals of Mathematical Statistics, 38(1):274–277, February 1967. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177699082.

Bloch:1967:BSM

[BW67b] Daniel A. Bloch and Geoffrey S. Watson. A Bayesian study of the multinomial distribution. Annals of Mathematical Statistics, 38 (5):1423–1435, October 1967. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177698697.

Bickel:1971:CCM

[BW71] P. J. Bickel and M. J. Wichura. Convergence criteria for multiparameter stochastic processes and some applications. Annals of Mathematical Statistics, 42(5):1656–1670, October 1971. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177693164.

Bickel:1965:RTP

[BY65] Peter J. Bickel and Joseph A. Yahav. Renewal theory in the plane. Annals of Mathematical Statistics, 36(3):946–955, June 1965. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177700067.

Bickel:1968:AOB

[BY68] Peter J. Bickel and Joseph A. Yahav. Asymptotically optimal Bayes and minimax procedures in sequential estimation. Annals of Mathematical Statistics, 39(2):442–456, April 1968. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177698408.

Bickel:1969:PRS

[BY69] Peter J. Bickel and Joseph A. Yahav. On an A.P.O. rule in sequential estimation with quadratic loss. Annals of Mathematical Statistics, 40 (2):417–426, April 1969. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177697706.

Byleen:1970:SMC

[Byl70] Karl Byleen. On Stanton and Mullin’s construction of room squares. Annals of Mathematical Statistics, 41(3):1122–1125, June 1970. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177696995.

Byrne:1935:TVD

[Byr35a] Lee Byrne. A theory of validation for derivative specifications and check lists. Annals of Mathematical Statistics, 6(3):146–157, September 1935. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177732592.

Byron:1935:PBP

[Byr35b] Frank H. Byron. The point binomial and probability paper. Annals of Mathematical Statistics, 6(1):21–26, March 1935. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177732638.

Birnbaum:1944:IDH

[BZ44] Z. W. Birnbaum and Herbert S. Zuckerman. An inequality due to H. Hornich. Annals of Mathematical Statistics, 15(3):328–329, September 1944. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177731241.

Birnbaum:1949:GDS

[BZ49] Z. W. Birnbaum and H. S. Zuckerman. A graphical determination of sample size for Wilks’ tolerance limits. Annals of Mathematical Statistics, 20(2):313–316, June 1949. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177730044.

Cacoullos:1961:CDD

[Cac61] T. Cacoullos. A combinatorial derivation of the distribution of the truncated Poisson sufficient statistic. Annals of Mathematical Statistics, 32(3):904–905, September 1961. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177704986.

Cacoullos:1966:CAP

[Cac66] T. Cacoullos. On a class of admissible partitions. Annals of Mathematical Statistics, 37(1):189–195, February 1966. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177699608.

Cacoullos:1967:CNC

[Cac67] Theophilos Cacoullos. Characterizations of normality by constant regression of linear statistics on another linear statistic. Annals of Mathematical Statistics, 38(6):1894–1898, December 1967. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177698623.

Cadwell:1953:DQR

[Cad53] J. H. Cadwell. The distribution of quasi-ranges in samples from a normal population. Annals of Mathematical Statistics, 24(4):603– 613, December 1953. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177728916.

Cadwell:1954:PIR

[Cad54] J. H. Cadwell. The probability integral of range for samples from a symmetrical unimodal population. Annals of Mathematical Statistics, 25(4):803–806, December 1954. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177728672.

Camp:1937:MOP

[Cam37a] Burton H. Camp. Methods of obtaining probability distributions. Annals of Mathematical Statistics, 8(2):90–102, June 1937. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177732429.

Camp:1937:NNE

[Cam37b] Chester C. Camp. Note on numerical evaluation of double series. Annals of Mathematical Statistics, 8(1):72–75, March 1937. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177732454.

Camp:1938:NSB

[Cam38a] B. H. Camp. Notice to subscribers of the Annals of Mathematical Statistics. Annals of Mathematical Statistics, 9(1):??, March 1938. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177732364.

Camp:1938:NDG

[Cam38b] Burton H. Camp. Notes on the distribution of the geometric mean. Annals of Mathematical Statistics, 9(3):221–226, September 1938. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177732312.

Camp:1942:SRA

[Cam42] Burton H. Camp. Some recent advances in mathematical statistics, I. Annals of Mathematical Statistics, 13(1):62–73, March 1942. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177731643.

Camp:1946:EDF

[Cam46] Burton H. Camp. The effect on a distribution function of small changes in the population function. Annals of Mathematical Statistics, 17(2):226–231, June 1946. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177730983.

Camp:1948:GDI

[Cam48] Burton H. Camp. Generalization to N dimensions of inequalities of the Tchebycheff type. Annals of Mathematical Statistics, 19(4):568– 574, December 1948. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177730152.

Camp:1951:APB

[Cam51] Burton H. Camp. Approximation to the point binomial. Annals of Mathematical Statistics, 22(1):130–131, March 1951. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177729705.

Cameron:1952:RST

[Cam52] J. M. Cameron. Results of some tests of randomness on pseudorandom on numbers (preliminary report). Annals of Mathematical Statistics, 23(1):138, March 1952. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http:/

/projecteuclid.org/euclid.aoms/1177729493;   http://www.
jstor.org/stable/2236409. Abstract only.

Cam:1964:SAS

[Cam64] L. Le. Cam. Sufficiency and approximate sufficiency. Annals of Mathematical Statistics, 35(4):1419–1455, December 1964. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177700372.

Campbell:1970:EGP

[Cam70] L. L. Campbell. Equivalence of Gauss’s principle and minimum discrimination information estimation of probabilities. Annals of Mathematical Statistics, 41(3):1011–1015, June 1970. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177696977.

Capon:1961:AEC

[Cap61] Jack Capon. Asymptotic efficiency of certain locally most powerful rank tests. Annals of Mathematical Statistics, 32(1):88–100, March 1961. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177705142.

Capon:1964:RND

[Cap64] Jack Capon. Radon–Nikodym derivatives of stationary Gaussian measures. Annals of Mathematical Statistics, 35(2):517–531, June 1964. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177703552.

Carver:1930:EFT

[Car30] H. C. Carver. Editorial: Fundamentals of the theory of sampling. Annals of Mathematical Statistics, 1(3):260–274, August 1930. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177733091.

Carlson:1932:SDM

[Car32a] John L. Carlson. A study of the distribution of means estimated from small samples by the method of maximum likelihood for Pearson’s Type II curve. Annals of Mathematical Statistics, 3(2):86– 107, May 1932. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177732908.

Carver:1932:ETR

[Car32b] Harry C. Carver. Editorial: Trapezoidal rule for computing seasonal indices. Annals of Mathematical Statistics, 3(4):361–367, November 1932. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177732841.

Carver:1934:ENT

[Car34a] H. C. Carver. Editorial: A new type of average for security prices. Annals of Mathematical Statistics, 5(1):73–80, March 1934. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177732735.

Carver:1934:EPC

[Car34b] H. C. Carver. Editorial: Punched card systems and statistics. Annals of Mathematical Statistics, 5(2):153–160, June 1934. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177732714.

Carlton:1946:EPR

[Car46] A. George Carlton. Estimating the parameters of a rectangular distribution. Annals of Mathematical Statistics, 17(3):355–358, September 1946. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177730949.

Carpenter:1950:NEC

[Car50] Osmer Carpenter. Note on the extension of Craig’s theorem to noncentral variates. Annals of Mathematical Statistics, 21(3):455–457, September 1950. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177729807.

Carlyle:1967:ISC

[Car67] J. W. Carlyle. Identification of state-calculable functions of finite Markov chains. Annals of Mathematical Statistics, 38(1):201–205, February 1967. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177699071.

Carney:1968:RGP

[Car68] Edward J. Carney. Relationship of generalized polykays to unrestricted sums for balanced complete finite populations. Annals of Mathematical Statistics, 39(2):643–656, April 1968. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177698423.

Carney:1970:MPU

[Car70] E. J. Carney. Multiplication of polykays using ordered partitions. Annals of Mathematical Statistics, 41(5):1749–1752, October 1970. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177696819.

Castellani:1950:MDL

[Cas50] Maria Castellani. On multinomial distributions with limited freedom: A stochastic Genesis of Pareto’s and Pearson’s curves. Annals of Mathematical Statistics, 21(2):289–293, June 1950. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177729847.

Causey:1968:SEM

[Cau68] B. D. Causey. Some examples of multi-dimensional incomplete block designs. Annals of Mathematical Statistics, 39(5):1577–1590, October 1968. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177698139.

Causey:1972:SRC

[Cau72] B. D. Causey. Sensitivity of raked contingency table totals to changes in problem conditions. Annals of Mathematical Statistics, 43(2):656– 658, April 1972. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177692647.

Cochran:1948:DFC

[CB48] W. G. Cochran and C. I. Bliss. Discriminant functions with covariance. Annals of Mathematical Statistics, 19(2):151–176, June 1948. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177730242.

Cernuschi:1946:CRE

[CC46] Felix Cernuschi and Louis Castagnetto. Chains of rare events. Annals of Mathematical Statistics, 17(1):53–61, March 1946. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177731022.

Cernuschi:1947:PSC

[CC47] Félix Cernuschi and Louis Castagnetto. Probability schemes with contagion in space and time. Annals of Mathematical Statistics, 18(1): 122–127, March 1947. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177730500.

Connor:1954:STP

[CC54] W. S. Connor and W. H. Clatworthy. Some theorems for partially balanced designs. Annals of Mathematical Statistics, 25(1):100–112, March 1954. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177728849.

Curnow:1962:NEC

[CD62] R. N. Curnow and C. W. Dunnett. The numerical evaluation of certain multivariate normal integrals. Annals of Mathematical Statistics, 33(2):571–579, June 1962. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177704581.

Cochran:1945:RMC

[CDDK45] W. G. Cochran, W. E. Deming, P. S. Dwyer, and T. Koopmans. Report of the Membership Committee of the Institute. Annals of Mathematical Statistics, 16(1):107–108, March 1945. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177731181.

Cover:1967:GPR

[CE67] Thomas M. Cover and Bradley Efron. Geometrical probability and random points on a hypersphere. Annals of Mathematical Statistics, 38(1):213–220, February 1967. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177699073.

Carlsson:1966:SPS

[CG66] Siv Carlsson and Ulf Grenander. Some properties of statistical reliability functions. Annals of Mathematical Statistics, 37(4):826–836, August 1966. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177699363.

Chernoff:1967:ADL

[CGJ67] Herman Chernoff, Joseph L. Gastwirth, and M. V. Johns, Jr. Asymptotic distribution of linear combinations of functions of order statistics with applications to estimation. Annals of Mathematical Statistics, 38(1):52–72, February 1967. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177699058.

Chung:1945:CFA

[CH45] Kai-Lai Chung and Lietz C. Hsu. A combinatorial formula and its application to the theory of probability of arbitrary events. Annals of Mathematical Statistics, 16(1):91–95, March 1945. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177731175.

Chu:1955:MSM

[CH55] John T. Chu and Harold Hotelling. The moments of the sample median. Annals of Mathematical Statistics, 26(4):593–606, December 1955. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177728419.

Chapman:1935:GPC

[Cha35] Dwight W. Chapman. The generalized problem of correct matchings. Annals of Mathematical Statistics, 6(2):85–95, June 1935. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177732613.

Chand:1950:DRC

[Cha50a] Uttam Chand. Distributions related to comparison of two means and two regression coefficients. Annals of Mathematical Statistics, 21 (4):507–522, December 1950. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177729748.

Chapman:1950:STS

[Cha50b] Douglas G. Chapman. Some two sample tests. Annals of Mathematical Statistics, 21(4):601–606, December 1950. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177729755.

Chand:1951:TCH

[Cha51] Uttam Chand. Test criteria for hypotheses of symmetry of a regression matrix. Annals of Mathematical Statistics, 22(4):513–522, December 1951. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177729541.

Chapman:1954:EBP

[Cha54] Douglas G. Chapman. The estimation of biological populations. Annals of Mathematical Statistics, 25(1):1–15, March 1954. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177728844.

Chapman:1956:EPT

[Cha56] Douglas G. Chapman. Estimating the parameters of a truncated gamma distribution. Annals of Mathematical Statistics, 27(2):498– 506, June 1956. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177728272.

Chapman:1958:CSS

[Cha58] Douglas G. Chapman. A comparative study of several one-sided goodness-of-fit tests. Annals of Mathematical Statistics, 29(3):655– 674, September 1958. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.

aoms/1177706529. See correction [Cha65].

Chakravarti:1961:SMC

[Cha61] I. M. Chakravarti. On some methods of construction of partially balanced arrays. Annals of Mathematical Statistics, 32(4):1181–1185, December 1961. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177704857.

Chanda:1962:BSC

[Cha62a] K. C. Chanda. On bounds of serial correlations. Annals of Mathematical Statistics, 33(4):1457–1460, December 1962. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177704378.

Chatterjee:1962:SIP

[Cha62b] Shoutir Kishore Chatterjee. Sequential inference procedures of Stein’s type for a class of multivariate regression problems. Annals of Mathematical Statistics, 33(3):1039–1064, September 1962. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177704471.

Chacko:1963:THA

[Cha63a] V. J. Chacko. Testing homogeneity against ordered alternatives. Annals of Mathematical Statistics, 34(3):945–956, September 1963. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177704017.

Chanda:1963:AEC

[Cha63b] K. C. Chanda. Asymptotic expansions for a class of distribution functions. Annals of Mathematical Statistics, 34(4):1302–1307, December 1963. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177703865.

Chanda:1963:ETS

[Cha63c] K. C. Chanda. On the efficiency of two-sample Mann–Whitney test for discrete populations. Annals of Mathematical Statistics, 34(2): 612–617, June 1963. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177704173.

Chazan:1964:NCS

[Cha64] Dan Chazan. A note on convergence of sub-martingales. Annals of Mathematical Statistics, 35(4):1811–1814, December 1964. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177700404.

Chapman:1965:CNC

[Cha65] D. G. Chapman. Correction notes: Correction to “A Comparative Study of Several One-Sided Goodness-of-Fit Tests”. Annals of Mathematical Statistics, 36(5):1583, October 1965. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177699922. See [Cha58].

Chatterjee:1966:BST

[Cha66] Shoutir Kishore Chatterjee. A bivariate sign test for location. Annals of Mathematical Statistics, 37(6):1771–1782, December 1966. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177699165.

Chan:1967:RLM

[Cha67] L. K. Chan. Remark on the linearized maximum likelihood estimate. Annals of Mathematical Statistics, 38(6):1876–1881, December 1967. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177698620.

Chatterji:1969:CT

[Cha69] S. D. Chatterji. An L^{p}-convergence theorem. Annals of Mathematical Statistics, 40(3):1068–1070, June 1969. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177697609.

Chang:1970:ARD

[Cha70a] Tseng C. Chang. On an asymptotic representation of the distribution of the characteristic roots of S_{1}S_{2}^{−1}. Annals of Mathematical Statistics, 41(2):440–445, April 1970. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177697083.

Chao:1970:ABB

[Cha70b] M. T. Chao. The asymptotic behavior of Bayes’ estimators. Annals of Mathematical Statistics, 41(2):601–608, April 1970. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177697100.

Chanda:1972:ETE

[Cha72] K. C. Chanda. On estimation of tail end probabilities of the sample mean for linear stochastic processes. Annals of Mathematical Statistics, 43(5):1680–1686, October 1972. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177692402.

Chernoff:1949:AST

[Che49] Herman Chernoff. Asymptotic Studentization in testing of hypotheses. Annals of Mathematical Statistics, 20(2):268–278, June 1949. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177730035.

Chernoff:1951:PST

[Che51] Herman Chernoff. A property of some type A regions. Annals of Mathematical Statistics, 22(3):472–474, September 1951. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177729598.

Chernoff:1952:MAE

[Che52] Herman Chernoff. A measure of asymptotic efficiency for tests of a hypothesis based on the sum of observations. Annals of Mathematical Statistics, 23(4):493–507, December 1952. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177729330.

Chernoff:1953:LOD

[Che53] Herman Chernoff. Locally optimal designs for estimating parameters. Annals of Mathematical Statistics, 24(4):586–602, December 1953. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177728915.

Chernoff:1954:DLR

[Che54] Herman Chernoff. On the distribution of the likelihood ratio. Annals of Mathematical Statistics, 25(3):573–578, September 1954. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177728725.

Chernoff:1956:LST

[Che56] Herman Chernoff. Large-sample theory: Parametric case. Annals of Mathematical Statistics, 27(1):1–22, March 1956. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177728347.

Chernoff:1959:SDE

[Che59] Herman Chernoff. Sequential design of experiments. Annals of Mathematical Statistics, 30(3):755–770, September 1959. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177706205.

Chernoff:1962:SMC

[Che62] Herman Chernoff. The scoring of multiple choice questionnaires. Annals of Mathematical Statistics, 33(2):375–393, June 1962. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177704565.

Chernoff:1965:STMa

[Che65a] Herman Chernoff. Sequential test for the mean of a normal distribution. III. (Small t). Annals of Mathematical Statistics, 36(1):28–54, February 1965. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177700269.

Chernoff:1965:STMb

[Che65b] Herman Chernoff. Sequential tests for the mean of a normal distribution IV (discrete case). Annals of Mathematical Statistics, 36(1): 55–68, February 1965. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177700270.

Chernoff:1966:RDV

[Che66] Herman Chernoff. Review: Dennis V. Lindley, Introduction to Probability and Statistics from a Bayesian Viewpoint, Vols. 1 and 2. Annals of Mathematical Statistics, 37(4):1056–1059, August 1966. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177699389.

Chernoff:1967:NRM

[Che67a] H. Chernoff. A note on risk and maximal regular generalized submartingales in stopping problems. Annals of Mathematical Statistics, 38(2):606–607, April 1967. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177698979.

Chew:1967:SSP

[Che67b] Milton C. Chew, Jr. A sequential search procedure. Annals of Mathematical Statistics, 38(2):494–502, April 1967. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177698965.

Cheng:1969:OPF

[Che69] M. C. Cheng. The orthant probabilities of four Gaussian variates. Annals of Mathematical Statistics, 40(1):152–161, February 1969. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177697812.

Chiang:1956:RBA

[Chi56] Chin Long Chiang. On regular best asymptotically normal estimates. Annals of Mathematical Statistics, 27(2):336–351, June 1956. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177728262.

Chow:1965:LCM

[Cho65] Y. S. Chow. Local convergence of martingales and the law of large numbers. Annals of Mathematical Statistics, 36(2):552–558, April 1965. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177700166.

Chow:1966:MSO

[Cho66a] Y. S. Chow. On the moments of some one-sided stopping rules. Annals of Mathematical Statistics, 37(2):382–387, April 1966. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177699519.

Chow:1966:SCT

[Cho66b] Y. S. Chow. Some convergence theorems for independent random variables. Annals of Mathematical Statistics, 37(6):1482–1493, December 1966. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177699140.

Chow:1967:SLL

[Cho67a] Y. S. Chow. On a strong law of large numbers for martingales. Annals of Mathematical Statistics, 38(2):610, April 1967. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177698981.

Chow:1967:EVS

[Cho67b] Y. S. Chow. On the expected value of a stopped submartingale. Annals of Mathematical Statistics, 38(2):608–609, April 1967. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177698980.

Chow:1968:CSS

[Cho68] Y. S. Chow. Convergence of sums of squares of martingale differences. Annals of Mathematical Statistics, 39(1):123–133, February 1968.

CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177698509.

Chow:1969:EGS

[Cho69a] T. R. Chow. Equivalence of Gaussian stationary processes. Annals of Mathematical Statistics, 40(1):197–202, February 1969. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177697816.

Chow:1969:MET

[Cho69b] Y. S. Chow. Martingale extensions of a theorem of Marcinkiewicz and Zygmund. Annals of Mathematical Statistics, 40(2):427–433, April 1969. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177697707.

Chow:1971:C

[Cho71] Y. S. Chow. On the L_{p}-convergence for n^{−1/p}S_{n}, 0 < p < 2. Annals of Mathematical Statistics, 42(1):393–394, February 1971. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177693530.

Chung:1941:POL

[Chu41] Kai Lai Chung. On the probability of the occurrence of at least m events among n arbitrary events. Annals of Mathematical Statistics, 12(3):328–338, September 1941. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177731715.

Chung:1942:MFE

[Chu42] Kai-Lai Chung. On mutually favorable events. Annals of Mathematical Statistics, 13(3):338–349, September 1942. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177731573.

Chung:1943:FRP

[Chu43a] Kai Lai Chung. Further results on probabilities of a finite number of events. Annals of Mathematical Statistics, 14(3):234–237, September 1943. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177731417.

Chung:1943:GPF

[Chu43b] Kai Lai Chung. Generalization of Poincaré’s formula in the theory of probability. Annals of Mathematical Statistics, 14(1):63–65, March 1943. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177731493.

Chung:1943:FSP

[Chu43c] Kai Lai Chung. On fundamental systems of probabilities of a finite number of events. Annals of Mathematical Statistics, 14(2): 123–133, June 1943. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177731453.

Chung:1946:ADS

[Chu46a] Kai-Lai Chung. The approximate distribution of Student’s statistic. Annals of Mathematical Statistics, 17(4):447–465, December 1946. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177730884.

Churchill:1946:IGO

[Chu46b] Edmund Churchill. Information given by odd moments. Annals of Mathematical Statistics, 17(2):244–246, June 1946. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177730987.

Chung:1948:LK

[Chu48] Kai-Lai Chung. On a lemma by Kolmogoroff. Annals of Mathematical Statistics, 19(1):88–91, March 1948. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177730296.

Chung:1954:SAM

[Chu54] K. L. Chung. On a stochastic approximation method. Annals of Mathematical Statistics, 25(3):463–483, September 1954. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177728716.

Chu:1955:DSM

[Chu55] John T. Chu. On the distribution of the sample median. Annals of Mathematical Statistics, 26(1):112–116, March 1955. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177728598.

Chu:1956:ENA

[Chu56] J. T. Chu. Errors in normal approximations to the t, τ, and similar types of distribution. Annals of Mathematical Statistics, 27(3):780– 789, September 1956. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177728184.

Chu:1957:SUQ

[Chu57] J. T. Chu. Some uses of quasi-ranges. Annals of Mathematical Statistics, 28(1):173–180, March 1957. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177707044.

Chung:1961:NET

[Chu61] K. L. Chung. A note on the ergodic theorem of information theory. Annals of Mathematical Statistics, 32(2):612–614, June 1961. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177705069.

Chapman:1956:ESS

[CJ56] D. G. Chapman and C. O. Junge, Jr. The estimation of the size of a stratified animal population. Annals of Mathematical Statistics, 27 (2):375–389, June 1956. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177728264.

Constantine:1958:GCC

[CJ58] A. G. Constantine and A. T. James. On the general canonical correlation distribution. Annals of Mathematical Statistics, 29(4):1146– 1166, December 1958. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177706447.

Chernoff:1954:UML

[CL54] Herman Chernoff and E. L. Lehmann. The use of maximum likelihood estimates in χ^{2} tests for goodness of fit. Annals of Mathematical Statistics, 25(3):579–586, September 1954. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177728726.

Chernoff:1956:UGP

[CL56] Herman Chernoff and Gerald J. Lieberman. The use of generalized probability paper for continuous distributions. Annals of Mathematical Statistics, 27(3):806–818, September 1956. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177728187.

Cramer:1965:MNC

[CL65] Harald Cramér and M. R. Leadbetter. The moments of the number of crossings of a level by a stationary normal process. Annals of Mathematical Statistics, 36(6):1656–1663, December 1965. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177699794.

Cuppens:1970:DDA

[CL70] R. Cuppens and E. Lukacs. On the domains of definition of analytic characteristic functions. Annals of Mathematical Statistics, 41(3): 1096–1101, June 1970. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.

aoms/1177696990. See correction [CL71].

Cuppens:1971:CNC

[CL71] R. Cuppens and E. Lukacs. Correction note: Correction to “On the Domains of Definition of Analytic Characteristic Functions”. Annals of Mathematical Statistics, 42(1):415, February 1971. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177693538. See [CL70].

Clark:1937:NBD

[Cla37] C. E. Clark. Note on the binomial distribution. Annals of Mathematical Statistics, 8(2):116–117, June 1937. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177732432.

Clarke:1956:WLP

[Cla56] A. B. Clarke. A waiting line process of Markov type. Annals of Mathematical Statistics, 27(2):452–459, June 1956. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177728268.

Clarke:1957:MLE

[Cla57] A. Bruce Clarke. Maximum likelihood estimates in a simple queue. Annals of Mathematical Statistics, 28(4):1036–1040, December 1957. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177706808.

Chakravarti:1962:ARE

[CLA62] I. M. Chakravarti, F. C. Leone, and J. D. Alanen. Asymptotic relative efficiency of Mood’s and Massey’s two sample tests against some parametric alternatives. Annals of Mathematical Statistics, 33(4): 1375–1383, December 1962. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177704370.

Clark:1970:RFB

[Cla70] J. M. C. Clark. The representation of functionals of Brownian motion by stochastic integrals. Annals of Mathematical Statistics, 41 (4):1282–1295, August 1970. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177696903. See correction [Cla71].

Clark:1971:CNC

[Cla71] J. M. C. Clark. Correction notes: Correction to “The Representation of Functionals of Brownian Motion by Stochastic Integrals”. Annals of Mathematical Statistics, 42(5):1778, October 1971. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177693186. See [Cla70].

Cleveland:1971:PWI

[Cle71] William S. Cleveland. Projection with the wrong inner product and its application to regression with correlated errors and linear filtering of time series. Annals of Mathematical Statistics, 42(2): 616–624, April 1971. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177693411.

Cohn:1960:MPA

[CMPT60] Richard Cohn, Frederick Mosteller, John W. Pratt, and Maurice Tatsuoka. Maximizing the probability that adjacent order statistics of samples from several populations form overlapping intervals. Annals of Mathematical Statistics, 31(4):1095–1104, December 1960. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177705681.

Cochran:1940:AVW

[Coc40a] W. G. Cochran. The analysis of variance when experimental errors follow the Poisson or binomial laws. Annals of Mathematical Statistics, 11(3):335–347, September 1940. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177731871.

Cochran:1940:NAF

[Coc40b] W. G. Cochran. Note on an approximate formula for the significance levels of Z. Annals of Mathematical Statistics, 11(1):93–95, March 1940. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177731945.

Cochran:1943:CDS

[Coc43] W. G. Cochran. The comparison of different scales of measurement for experimental results. Annals of Mathematical Statistics, 14(3):205– 216, September 1943. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177731414.

Cochran:1944:RPJ

[Coc44] W. G. Cochran. Review: E. S. Pearson, John Wishart, “Student’s” Collected Papers. Annals of Mathematical Statistics, 15(4):435–438, December 1944. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177731216.

Cochran:1946:RAS

[Coc46] W. G. Cochran. Relative accuracy of systematic and stratified random samples for a certain class of populations. Annals of Mathematical Statistics, 17(2):164–177, June 1946. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177730978.

Cochran:1947:ARP

[Coc47] W. G. Cochran. Annual report of the President of the Institute for 1946. Annals of Mathematical Statistics, 18(1):150–155, March 1947. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177730508.

Cochran:1952:TGF

[Coc52] William G. Cochran. The χ^{2} test of goodness of fit. Annals of Mathematical Statistics, 23(3):315–345, September 1952. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177729380.

Cocke:1972:CLT

[Coc72] W. J. Cocke. Central Limit Theorems for sums of dependent vector variables. Annals of Mathematical Statistics, 43(3):968–976, June 1972. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177692559.

Cogburn:1962:CPO

[Cog62] Robert Cogburn. Conditional probability operators. Annals of Mathematical Statistics, 33(2):634–658, June 1962. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177704587.

Cogburn:1967:SSS

[Cog67] Robert Cogburn. Stringent solutions to statistical decision problems. Annals of Mathematical Statistics, 38(2):447–463, April 1967. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177698960.

Cohen:1940:NCP

[Coh40] A. C. Cohen, Jr. The numerical computation of the product of conjugate imaginary gamma functions. Annals of Mathematical Statistics, 11(2):213–218, June 1940. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177731917.

Cohen:1950:EMV

[Coh50] A. C. Cohen, Jr. Estimating the mean and variance of normal populations from singly truncated and doubly truncated samples. Annals of Mathematical Statistics, 21(4):557–569, December 1950. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177729751.

Cohen:1951:EPT

[Coh51] A. C. Cohen, Jr. Estimation of parameters in truncated Pearson frequency distributions. Annals of Mathematical Statistics, 22(2): 256–265, June 1951. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177729645.

Cohen:1957:RSM

[Coh57] A. Clifford Cohen, Jr. Restriction and selection in multinormal distributions. Annals of Mathematical Statistics, 28(3):731–741, September 1957. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177706885.

Cohen:1958:MSS

[Coh58] Leonard Cohen. On mixed single sample experiments. Annals of Mathematical Statistics, 29(4):947–971, December 1958. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177706435.

Cohen:1965:ELC

[Coh65a] Arthur Cohen. Estimates of linear combinations of the parameters in the mean vector of a multivariate distribution. Annals of Mathematical Statistics, 36(1):78–87, February 1965. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177700272.

Cohen:1965:HPE

[Coh65b] Arthur Cohen. A hybrid problem on the exponential family. Annals of Mathematical Statistics, 36(4):1185–1206, August 1965. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177699991.

Cohen:1966:AAL

[Coh66] Arthur Cohen. All admissible linear estimates of the mean vector. Annals of Mathematical Statistics, 37(2):458–463, April 1966. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177699528.

Cohen:1968:NAP

[Coh68] Arthur Cohen. A note on the admissibility of pooling in the analysis of variance. Annals of Mathematical Statistics, 39(5):1744–1746, October 1968. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177698157.

Cohn:1970:TAF

[Coh70] Harry Cohn. On the tail Σ-algebra of the finite inhomogeneous Markov chains. Annals of Mathematical Statistics, 41(6):2175–2176, December 1970. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177696725.

Cohn:1972:LTS

[Coh72] Harry Cohn. Limit theorems for sums of random variables defined on finite inhomogeneous Markov chains. Annals of Mathematical Statistics, 43(4):1283–1292, August 1972. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177692479.

Coleman:1932:CLC

[Col32] J. B. Coleman. A coefficient of linear correlation based on the method of least squares and the line of best fit. Annals of Mathematical Statistics, 3(2):79–85, May 1932. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177732907.

Cole:1951:RBM

[Col51] Randal H. Cole. Relations between moments of order statistics. Annals of Mathematical Statistics, 22(2):308–310, June 1951. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177729655.

Comer:1964:SSA

[Com64] John P. Comer, Jr. Some stochastic approximation procedures for use in process control. Annals of Mathematical Statistics, 35(3):1136– 1146, September 1964. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177703271.

Connor:1952:SRA

[Con52a] W. S. Connor. Some relations among the blocks of symmetrical group divisible designs. Annals of Mathematical Statistics, 23(4):602–609, December 1952. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177729339.

Connor:1952:SBI

[Con52b] W. S. Connor, Jr. On the structure of balanced incomplete block designs. Annals of Mathematical Statistics, 23(1):57–71, March 1952. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177729485. See

correction [Con53].

Connor:1953:CSB

[Con53] W. S. Connor. Correction to “On the Structure of Balanced Incomplete Block Designs”. Annals of Mathematical Statistics, 24(1): 135, March 1953. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/

1177729096. See [Con52b].

Connor:1958:UTA

[Con58] W. S. Connor. The uniqueness of the triangular association scheme. Annals of Mathematical Statistics, 29(1):262–266, March 1958. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177706724.

Connor:1963:CDS

[Con63a] W. S. Connor. The conditional distribution of sets of tests on a system simulated from tests on its components. Annals of Mathematical Statistics, 34(4):1585–1587, December 1963. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177703891.

Constantine:1963:SNC

[Con63b] A. G. Constantine. Some non-central distribution problems in multivariate analysis. Annals of Mathematical Statistics, 34(4):1270–1285, December 1963. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177703863.

Conover:1965:SMO

[Con65a] W. J. Conover. A k-sample model in order statistics. Annals of Mathematical Statistics, 36(4):1223–1235, August 1965. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177699993. See note [Dav66].

Conover:1965:SSK

[Con65b] W. J. Conover. Several k-sample Kolmogorov–Smirnov tests. Annals of Mathematical Statistics, 36(3):1019–1026, June 1965. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177700073.

Constantine:1966:DHG

[Con66a] A. G. Constantine. The distribution of Hotelling’s generalised T_{0}^{2}. Annals of Mathematical Statistics, 37(1):215–225, February 1966. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177699611.

Consul:1966:EDL

[Con66b] P. C. Consul. On the exact distributions of the likelihood ratio criteria for testing linear hypotheses about regression coefficients. Annals of Mathematical Statistics, 37(5):1319–1330, October 1966. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177699276.

Conover:1967:DFT

[Con67a] W. J. Conover. The distribution functions of Tsao’s truncated Smirnov statistics. Annals of Mathematical Statistics, 38(4):1208– 1215, August 1967. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177698789.

Conover:1967:SEO

[Con67b] W. J. Conover. A k-sample extension of the one-sided two-sample Smirnov test statistic. Annals of Mathematical Statistics, 38(6):1726– 1730, December 1967. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177698606.

Consul:1967:EDL

[Con67c] P. C. Consul. On the exact distributions of likelihood ratio criteria for testing independence of sets of variates under the null hypothesis. Annals of Mathematical Statistics, 38(4):1160–1169, August 1967. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177698785.

Consul:1967:EDC

[Con67d] P. C. Consul. On the exact distributions of the criterion W for testing sphericity in a p-variate normal distribution. Annals of Mathematical Statistics, 38(4):1170–1174, August 1967. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177698786.

Consul:1969:EDV

[Con69] P. C. Consul. On the exact distributions of Votaw’s criteria for testing compound symmetry of a covariance matrix. Annals of Mathematical Statistics, 40(3):836–843, June 1969. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177697591.

Conner:1970:CAT

[Con70] William M. Conner. The capacity and ambiguity of a transducer. Annals of Mathematical Statistics, 41(6):2093–2104, December 1970. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177696707.

Copeland:1932:TPP

[Cop32] Arthur H. Copeland. The theory of probability from the point of view of admissible numbers. Annals of Mathematical Statistics, 3(3): 143–156, August 1932. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177732883.

Corsten:1960:PSR

[Cor60] L. C. A. Corsten. Proper spaces related to triangular partially balanced incomplete block designs. Annals of Mathematical Statistics, 31 (2):498–501, June 1960. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177705916.

Court:1944:RPN

[Cou44] Louis M. Court. A reciprocity principle for the Neyman–Pearson theory of testing statistical hypotheses. Annals of Mathematical Statistics, 15(3):326–327, September 1944. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177731240.

Cover:1966:PRG

[Cov66] Thomas M. Cover. The probability that a random game is unfair. Annals of Mathematical Statistics, 37(6):1796–1799, December 1966. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177699168.

Cover:1969:HTF

[Cov69] Thomas M. Cover. Hypothesis testing with finite statistics. Annals of Mathematical Statistics, 40(3):828–835, June 1969. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177697590.

Cox:1940:ECB

[Cox40] Gertrude M. Cox. Enumeration and construction of balanced incomplete block configurations. Annals of Mathematical Statistics, 11(1): 72–85, March 1940. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177731943.

Cox:1956:NWR

[Cox56] D. R. Cox. A note on weighted randomization. Annals of Mathematical Statistics, 27(4):1144–1151, December 1956. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177728081.

Cox:1958:SPC

[Cox58] D. R. Cox. Some problems connected with statistical inference. Annals of Mathematical Statistics, 29(2):357–372, June 1958. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177706618.

Cox:1962:RSW

[Cox62] D. R. Cox. Review: Samuel S. Wilks, Mathematical Statistics. Annals of Mathematical Statistics, 33(4):1473–1476, December 1962. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177704384.

Chapman:1951:MVE

[CR51] Douglas G. Chapman and Herbert Robbins. Minimum variance estimation without regularity assumptions. Annals of Mathematical Statistics, 22(4):581–586, December 1951. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177729548.

Chow:1963:RTR

[CR63] Y. S. Chow and Herbert Robbins. A renewal theorem for random variables which are dependent or non-identically distributed. Annals of Mathematical Statistics, 34(2):390–395, June 1963. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177704148.

Chernoff:1965:BSS

[CR65a] Herman Chernoff and S. N. Ray. A Bayes sequential sampling inspection plan. Annals of Mathematical Statistics, 36(5):1387–1407, October 1965. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177699898.

Chow:1965:ATF

[CR65b] Y. S. Chow and Herbert Robbins. On the asymptotic theory of fixed-width sequential confidence intervals for the mean. Annals of Mathematical Statistics, 36(2):457–462, April 1965. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177700156.

Craig:1931:NDM

[Cra31a] Cecil C. Craig. Note on the distribution of means of samples of N drawn from a Type A population. Annals of Mathematical Statistics, 2(1):99–101, February 1931. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177733045.

Craig:1931:PSI

[Cra31b] Cecil C. Craig. On a property of the semi-invariants of Thiele. Annals of Mathematical Statistics, 2(2):154–164, May 1931. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177732994.

Craig:1931:SCC

[Cra31c] Cecil C. Craig. Sampling in the case of correlated observations. Annals of Mathematical Statistics, 2(3):324–332, August 1931. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177732976.

Craig:1932:SDM

[Cra32] Allen T. Craig. The simultaneous distribution of mean and standard deviation in small samples. Annals of Mathematical Statistics, 3(2): 126–140, May 1932. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177732910.

Craig:1933:CBC

[Cra33a] Allen T. Craig. On the correlation between certain averages from small samples. Annals of Mathematical Statistics, 4(2):127–142, May 1933. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177732805.

Craig:1933:TIB

[Cra33b] Cecil C. Craig. On the Tchebychef inequality of Bernstein. Annals of Mathematical Statistics, 4(2):94–102, May 1933. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177732803.

Craig:1936:NEC

[Cra36a] Cecil C. Craig. A new exposition and chart for the Pearson system of frequency curves. Annals of Mathematical Statistics, 7(1): 16–28, March 1936. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177732542.

Craig:1936:FF

[Cra36b] Cecil C. Craig. On the frequency function of xy. Annals of Mathematical Statistics, 7(1):1–15, March 1936. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177732541.

Craig:1936:SCD

[Cra36c] Cecil C. Craig. Sheppard’s corrections for a discrete variable. Annals of Mathematical Statistics, 7(2):55–61, June 1936. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177732520.

Craig:1937:RAM

[Cra37] Allen T. Craig. Report of the annual meeting of the Institute of Mathematical Statistics. Annals of Mathematical Statistics, 8(1):76, March 1937. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177732455.

Craig:1938:ICE

[Cra38a] Allen T. Craig. On the independence of certain estimates of variance. Annals of Mathematical Statistics, 9(1):48–55, March 1938. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177732358.

Craig:1938:RAM

[Cra38b] Allen T. Craig. Report of the annual meeting of the Institute of Mathematical Statistics. Annals of Mathematical Statistics, 9(1):??, March 1938. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177732363.

Craig:1939:RAM

[Cra39a] A. T. Craig. Report of the annual meeting of the Institute. Annals of Mathematical Statistics, 10(1):81–83, March 1939. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177732250.

Craig:1939:MRM

[Cra39b] Allen T. Craig. On the mathematics of the representative method of sampling. Annals of Mathematical Statistics, 10(1):26–34, March 1939. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177732244.

Craig:1940:PSI

[Cra40] Cecil C. Craig. The product semi-invariants of the mean and a central moment in samples. Annals of Mathematical Statistics, 11(2):177– 185, June 1940. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177731911.

Craig:1941:NSC

[Cra41a] Cecil C. Craig. A note on Sheppard’s corrections. Annals of Mathematical Statistics, 12(3):339–345, September 1941. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177731716.

Craig:1941:NDN

[Cra41b] Cecil C. Craig. Note on the distribution of non-central t with an application. Annals of Mathematical Statistics, 12(2):224–228, June 1941. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177731752.

Craig:1942:RAM

[Cra42] Cecil C. Craig. Recent advances in mathematical statistics, II. Annals of Mathematical Statistics, 13(1):74–85, March 1942. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177731644.

Craig:1943:NBL

[Cra43a] Allen T. Craig. A note on the best linear estimate. Annals of Mathematical Statistics, 14(1):88–90, March 1943. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177731495.

Craig:1943:NIC

[Cra43b] Allen T. Craig. Note on the independence of certain quadratic forms. Annals of Mathematical Statistics, 14(2):195–197, June 1943. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177731461. See

note [Hot44a].

Craig:1943:ARP

[Cra43c] Cecil C. Craig. Annual report of the President of the Institute. Annals of Mathematical Statistics, 14(1):98–100, March 1943. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177731499.

Craig:1944:ARP

[Cra44a] C. C. Craig. Annual report of the President of the Institute. Annals of Mathematical Statistics, 15(1):111–112, March 1944. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177731323.

Crathorne:1944:HLR

[Cra44b] A. R. Crathorne. Henry Lewis Rietz — in memoriam. Annals of Mathematical Statistics, 15(1):102–108, March 1944. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177731321.

Craig:1947:BFN

[Cra47a] Allen T. Craig. Bilinear forms in normally correlated variables. Annals of Mathematical Statistics, 18(4):565–573, December 1947. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177730347.

Craig:1947:RNH

[Cra47b] C. C. Craig. Report on the New Haven meeting of the Institute. Annals of Mathematical Statistics, 18(4):616–618, December 1947. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177730359.

Cramer:1947:PPT

[Cra47c] Harald Cramér. Problems in probability theory. Annals of Mathematical Statistics, 18(2):165–193, June 1947. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177730437.

Cramer:1948:ABS

[Cra48] G. F. Cramer. An approximation to the binomial summation. Annals of Mathematical Statistics, 19(4):592–594, December 1948. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177730156.

Cramer:1953:RMW

[Cra53] Harald Cramér. Richard von Mises’ work in probability and statistics. Annals of Mathematical Statistics, 24(4):657–662, December 1953. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177728921.

Craig:1960:OSA

[Cra60] Allen T. Craig. Our Silver anniversary. Annals of Mathematical Statistics, 31(4):835–837, December 1960. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177705660.

Cramer:1962:KWM

[Cra62] Harald Cramér. A. I. Khinchin’s work in mathematical probability. Annals of Mathematical Statistics, 33(4):1227–1237, December 1962. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177704355.

Craswell:1965:DET

[Cra65] K. J. Craswell. Density estimation in a topological group. Annals of Mathematical Statistics, 36(3):1047–1048, June 1965. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177700078.

Crawford:1966:CGE

[Cra66] Gordon B. Crawford. Characterization of geometric and exponential distributions. Annals of Mathematical Statistics, 37(6):1790–1795, December 1966. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177699167.

Crawford:1970:HFH

[Cra70] Gordon B. Crawford. Harmonic functions and hitting distributions for Markov processes. Annals of Mathematical Statistics, 41(6):2123– 2125, December 1970. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177696710.

Crow:1958:PAC

[Cro58] Edwin L. Crow. A property of additively closed families of distributions. Annals of Mathematical Statistics, 29(3):892–897, September 1958. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177706546.

Crouse:1964:NMT

[Cro64] C. F. Crouse. Note on Mood’s test. Annals of Mathematical Statistics, 35(4):1825–1826, December 1964. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177700408.

Chow:1965:MRS

[CRT65] Y. S. Chow, Herbert Robbins, and Henry Teicher. Moments of randomly stopped sums. Annals of Mathematical Statistics, 36(3):789– 799, June 1965. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177700053.

Cryer:1972:MMR

[CRWC72] J. D. Cryer, Tim Robertson, F. T. Wright, and Robert J. Casady. Monotone median regression. Annals of Mathematical Statistics, 43 (5):1459–1469, October 1972. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177692378.

Cotterman:1939:TSM

[CS39] Chas. W. Cotterman and Laurence H. Snyder. Tests of simple Mendelian inheritance in randomly collected data of one and two generations. Journal of the American Statistical Association, 34 (207):511–523, September 1939. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic). URL http://www.jstor.org/ stable/2279484. See note [Man45a].

Chernoff:1952:GNP

[CS52] Herman Chernoff and Henry Scheffé. A generalization of the Neyman–Pearson fundamental lemma. Annals of Mathematical Statistics, 23(2):213–225, June 1952. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177729438.

Chernoff:1958:ANE

[CS58] Herman Chernoff and I. Richard Savage. Asymptotic normality and efficiency of certain nonparametric test statistics. Annals of Mathematical Statistics, 29(4):972–994, December 1958. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177706436.

Sun:1967:NUD

[cS67] Tze chien Sun. A note on the unimodality of distribution functions of class L. Annals of Mathematical Statistics, 38(4):1296–1299, August 1967. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177698804.

Chow:1968:M

[CS68] Y. S. Chow and W. J. Studden. On the monotonicity of E_{p} (S_{t}/t). Annals of Mathematical Statistics, 39(5):1755, October 1968. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177698161.

Chow:1969:MVU

[CS69] Y. S. Chow and W. J. Studden. Monotonicity of the variance under truncation and variations of Jensen’s inequality. Annals of Mathematical Statistics, 40(3):1106–1108, June 1969. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177697619.

Cohen:1970:ELM

[CS70] Arthur Cohen and Harold B. Sackrowitz. Estimation of the last mean of a monotone sequence. Annals of Mathematical Statistics, 41(6): 2021–2034, December 1970. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177696702.

Cohen:1971:UTH

[CS71] Arthur Cohen and William E. Strawderman. Unbiasedness of tests for homogeneity of variances. Annals of Mathematical Statistics, 42 (1):355–360, February 1971. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177693520.

Crow:1972:MLE

[CS72] L. H. Crow and I. N. Shimi. Maximum likelihood estimation of lifedistributions from renewal testing. Annals of Mathematical Statistics, 43(6):1827–1838, December 1972. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177690854.

Csorgo:1965:CNC

[Csö65a] Miklós Csörgő. Correction notes: Correction to “Some Rényi Type Limit Theorems for Empirical Distribution Functions”. Annals of Mathematical Statistics, 36(3):1069, June 1965. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177700085. See [Csö65b].

Csorgo:1965:SRT

[Csö65b] Miklós Csörgő. Some Rényi type limit theorems for empirical distribution functions. Annals of Mathematical Statistics, 36(1):322–326, February 1965. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/

1177700296. See correction [Csö65a].

Csorgo:1965:SST

[Csö65c] Miklós Csörgő. Some Smirnov type theorems of probability theory. Annals of Mathematical Statistics, 36(4):1113–1119, August 1965. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177699986.

Cornfield:1956:AVM

[CT56] Jerome Cornfield and John W. Tukey. Average values of mean squares in factorials. Annals of Mathematical Statistics, 27(4):907–949, December 1956. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177728067.

Chernoff:1958:CLT

[CT58] H. Chernoff and H. Teicher. A Central Limit Theorem for sums of interchangeable random variables. Annals of Mathematical Statistics, 29(1):118–130, March 1958. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177706709.

Chow:1966:SMS

[CT66] Y. S. Chow and H. Teicher. On second moments of stopping rules. Annals of Mathematical Statistics, 37(2):388–392, April 1966. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177699520.

Chow:1971:ACS

[CT71] Y. S. Chow and H. Teicher. Almost certain summability of independent, identically distributed random variables. Annals of Mathematical Statistics, 42(1):401–404, February 1971. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177693533.

Cuppens:1969:RBR

[Cup69a] R. Cuppens. Review: B. Ramanchandran, Advanced Theory of Characteristic Functions. Annals of Mathematical Statistics, 40(3):1139– 1141, June 1969. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177697630.

Cuppens:1969:FPP

[Cup69b] Roger Cuppens. On finite products of Poisson-type characteristic functions of several variables. Annals of Mathematical Statistics, 40 (2):434–444, April 1969. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177697708.

Curtiss:1941:DQT

[Cur41] J. H. Curtiss. On the distribution of the quotient of two chance variables. Annals of Mathematical Statistics, 12(4):409–421, December 1941. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177731679.

Curtiss:1942:NTM

[Cur42] J. H. Curtiss. A note on the theory of moment generating functions. Annals of Mathematical Statistics, 13(4):430–433, December 1942. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177731541.

Curtiss:1943:TUA

[Cur43] J. H. Curtiss. On transformations used in the analysis of variance. Annals of Mathematical Statistics, 14(2):107–122, June 1943. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177731452.

Curtiss:1946:NSS

[Cur46] J. H. Curtiss. A note on some single sampling plans requiring the inspection of a small number of items. Annals of Mathematical Statistics, 17(1):62–70, March 1946. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177731023.

Curtis:1968:PTS

[Cur68] E. Count Curtis. A potential theory for supermartingales. Annals of Mathematical Statistics, 39(3):802–814, June 1968. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177698312.

Clark:1958:DMO

[CW58] Charles E. Clark and G. Trevor Williams. Distributions of the members of an ordered sample. Annals of Mathematical Statistics, 29 (3):862–870, September 1958. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177706542.

Clark:1959:ADM

[CW59] Charles E. Clark and G. Trevor Williams. Addenda: Distributions of the members of an ordered sample. Annals of Mathematical Statistics, 30(2):610, June 1959. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177706280.

Chernoff:1964:ECM

[CZ64] H. Chernoff and S. Zacks. Estimating the current mean of a normal distribution which is subjected to changes in time. Annals of Mathematical Statistics, 35(3):999–1018, September 1964. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177700517.

daSilvaRodrigues:1945:ECM

[da 45] Milton da Silva Rodrigues. On an extension of the concept of moment with applications to measures of variability, general similarity, and overlapping. Annals of Mathematical Statistics, 16(1):74– 84, March 1945. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177731173.

David:1959:OSR

[DA59] H. A. David and Beverly E. Arens. Optimal spacing in regression analysis. Annals of Mathematical Statistics, 30(4):1072–1081, December 1959. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177706091.

Daly:1940:UCL

[Dal40] Joseph F. Daly. On the unbiased character of likelihood-ratio tests for independence in normal systems. Annals of Mathematical Statistics, 11(1):1–32, March 1940. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177731939.

Daly:1941:PE

[Dal41] Joseph F. Daly. A problem in estimation. Annals of Mathematical Statistics, 12(4):459–461, December 1941. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177731683.

Daly:1946:USR

[Dal46] Joseph F. Daly. On the use of the sample range in an analogue of Student’s t-test. Annals of Mathematical Statistics, 17(1):71–74, March 1946. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177731024.

Dalenius:1962:RAS

[Dal62] Tore Dalenius. Recent advances in sample survey theory and methods. Annals of Mathematical Statistics, 33(2):325–349, June 1962. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177704563.

Dalenius:1964:RWG

[Dal64a] Tore Dalenius. Review: William G. Cochran, Sampling Techniques. Annals of Mathematical Statistics, 35(3):1381–1382, September 1964. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177703301.

DallAglio:1964:PVR

[Dal64b] Giorgio Dall’Aglio. Present value of a renewal process. Annals of Mathematical Statistics, 35(3):1326–1331, September 1964. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177703289.

Daley:1968:CSO

[Dal68] D. J. Daley. The correlation structure of the output process of some single server queueing systems. Annals of Mathematical Statistics, 39(3):1007–1019, June 1968. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177698333.

Dalenius:1969:RMM

[Dal69a] Tore Dalenius. Review: M. N. Murthy, Sampling Theory and Methods. Annals of Mathematical Statistics, 40(3):1136–1138, June 1969. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177697629.

Daley:1969:QSB

[Dal69b] D. J. Daley. Quasi-stationary behaviour of a left-continuous random walk. Annals of Mathematical Statistics, 40(2):532–539, April 1969. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177697721.

Dantzig:1939:CDA

[Dan39] George B. Dantzig. On a class of distributions that approach the normal distribution function. Annals of Mathematical Statistics, 10 (3):247–253, September 1939. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177732183. See remark [Gre50a].

Dantzig:1940:NET

[Dan40] George B. Dantzig. On the non-existence of tests of “Student’s” hypothesis having power functions independent of σ. Annals of Mathematical Statistics, 11(2):186–192, June 1940. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177731912.

Daniels:1954:DFT

[Dan54a] H. E. Daniels. A distribution-free test for regression parameters. Annals of Mathematical Statistics, 25(3):499–513, September 1954. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177728718.

Daniels:1954:SAS

[Dan54b] H. E. Daniels. Saddlepoint approximations in statistics. Annals of Mathematical Statistics, 25(4):631–650, December 1954. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177728652.

Daniel:1971:ABH

[Dan71] James W. Daniel. Asymptotic behavior of high order means. Annals of Mathematical Statistics, 42(5):1761–1762, October 1971. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177693178.

Darling:1952:THE

[Dar52] D. A. Darling. On a test for homogeneity and extreme values. Annals of Mathematical Statistics, 23(3):450–456; correction, 24, 135– 136 (1953), September 1952. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/ euclid.aoms/1177729390. See correction [Dar53a].

Darling:1953:CTH

[Dar53a] D. A. Darling. Correction to “On a Test for Homogeneity and Extreme Values”. Annals of Mathematical Statistics, 24(1):135–136, March 1953. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/

1177729097. See [Dar52].

Darling:1953:CPR

[Dar53b] D. A. Darling. On a class of problems related to the random division of an interval. Annals of Mathematical Statistics, 24(2):239– 253, June 1953. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/

1177729030. See correction [Dar62].

Darling:1955:CST

[Dar55] D. A. Darling. The Cramér–Smirnov test in the parametric case. Annals of Mathematical Statistics, 26(1):1–20, March 1955. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177728589.

Darling:1957:KSC

[Dar57] D. A. Darling. The Kolmogorov–Smirnov, Cramér–von Mises tests. Annals of Mathematical Statistics, 28(4):823–838, December 1957. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177706788.

Darling:1962:CNC

[Dar62] D. A. Darling. Correction notes: Correction to “On a class of Problems Related to the Random Division of an Interval”. Annals of Mathematical Statistics, 33(2):812, June 1962. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177704603. See [Dar53b].

Darroch:1964:DNS

[Dar64a] J. N. Darroch. On the distribution of the number of successes in independent trials. Annals of Mathematical Statistics, 35(3):1317– 1321, September 1964. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177703287.

Darroch:1964:TLQ

[Dar64b] J. N. Darroch. On the traffic-light queue. Annals of Mathematical Statistics, 35(1):380–388, March 1964. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177703761.

Darling:1965:RFS

[Dar65a] D. A. Darling. Review: Frank Spitzer, Principles of Random Walk. Annals of Mathematical Statistics, 36(4):1319–1320, August 1965. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177700010.

Darling:1965:RHB

[Dar65b] D. A. Darling. Review: Harald Bergström, Limit Theorems for Convolutions. Annals of Mathematical Statistics, 36(3):1070–1072, June 1965. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177700088.

Darroch:1965:OPP

[Dar65c] J. N. Darroch. An optimal property of principal components. Annals of Mathematical Statistics, 36(5):1579–1582, October 1965. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177699920.

Darst:1971:UMP

[Dar71] R. B. Darst. On universal measurability and perfect probability. Annals of Mathematical Statistics, 42(1):352–354, February 1971. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177693519.

Davis:1933:PAM

[Dav33] H. T. Davis. Polynomial approximation by the method of least squares. Annals of Mathematical Statistics, 4(3):155–195, August 1933. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177732778.

David:1949:NRW

[Dav49] Herbert T. David. A note on random walk. Annals of Mathematical Statistics, 20(4):603–608, December 1949. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177729954.

Davis:1950:DBC

[Dav50] R. C. Davis. Derivation of a broad class of consistent estimates. Annals of Mathematical Statistics, 21(3):425–431, September 1950. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177729801.

Davis:1951:NUB

[Dav51a] R. C. Davis. Note on uniformly best unbiased estimates. Annals of Mathematical Statistics, 22(3):440–445, September 1951. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177729591.

Davis:1951:MVN

[Dav51b] R. C. Davis. On minimum variance in nonregular estimation. Annals of Mathematical Statistics, 22(1):43–57, March 1951. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177729691.

David:1958:TSK

[Dav58] Herbert T. David. A three-sample Kolmogorov–Smirnov test. Annals of Mathematical Statistics, 29(3):842–851, September 1958. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177706540.

David:1960:CPB

[Dav60] H. A. David. A conservative property of binomial tests. Annals of Mathematical Statistics, 31(4):1205–1207, December 1960. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177705692. See correc-

tion [Dav61].

David:1961:CCP

[Dav61] H. A. David. Corrections to “A Conservative Property of Binomial Tests”. Annals of Mathematical Statistics, 32(4):1343, December 1961. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/

1177704875. See [Dav60].

David:1962:SMA

[Dav62] Herbert T. David. The sample mean among the moderate order statistics. Annals of Mathematical Statistics, 33(3):1160–1166, September 1962. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177704478.

David:1963:RJA

[Dav63a] H. A. David. Review: J. Arthur Greenwood, H. O. Hartley, Guide to Tables in Mathematical Statistics. Annals of Mathematical Statistics, 34(1):353–354, March 1963. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177704282.

David:1963:SMA

[Dav63b] Herbert T. David. The sample mean among the extreme normal order statistics. Annals of Mathematical Statistics, 34(1):33–55, March 1963. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177704241.

David:1965:OSS

[Dav65] Herbert T. David. Order statistics and statistics of structure (d). Annals of Mathematical Statistics, 36(3):897–906, June 1965. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177700062.

David:1966:NSM

[Dav66] H. A. David. A note on “A k-Sample Model in Order Statistics” by W. J. Conover. Annals of Mathematical Statistics, 37(1):287–288, February 1966. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/

1177699622. See [Con65a].

Davis:1967:CER

[Dav67] A. W. Davis. A counter-example relating to certain multivariate generalizations of t and F. Annals of Mathematical Statistics, 38(2): 613–615, April 1967. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177698983.

Davis:1968:SLD

[Dav68a] A. W. Davis. A system of linear differential equations for the distribution of Hotelling’s generalized T_{0}^{2}. Annals of Mathematical Statistics, 39(3):815–832, June 1968. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177698313.

Davis:1968:CTC

[Dav68b] Burgess Davis. Comparison tests for the convergence of martingales. Annals of Mathematical Statistics, 39(6):2141–2144, December 1968. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177698046.

Davis:1968:CRP

[Dav68c] James Avery Davis. Convergence rates for probabilities of moderate deviations. Annals of Mathematical Statistics, 39(6):2016–2028, December 1968. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177698029.

Davis:1968:CRL

[Dav68d] James Avery Davis. Convergence rates for the law of the iterated logarithm. Annals of Mathematical Statistics, 39(5):1479–1485, October 1968. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177698127.

Davis:1969:CTM

[Dav69a] Burgess Davis. A comparison test for martingale inequalities. Annals of Mathematical Statistics, 40(2):505–508, April 1969. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177697717.

Davis:1969:DPS

[Dav69b] Burgess Davis. Divergence properties of some martingale transforms. Annals of Mathematical Statistics, 40(5):1852–1854, October 1969. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177697402.

Davis:1969:CUL

[Dav69c] James Avery Davis. A characterization of the upper and lower classes in terms of convergence rates. Annals of Mathematical Statistics, 40(3):1120–1123, June 1969. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177697623.

Davis:1970:NDS

[Dav70] A. W. Davis. On the null distribution of the sum of the roots of a multivariate beta distribution. Annals of Mathematical Statistics, 41 (5):1557–1562, October 1970. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177696799.

Davis:1972:MDL

[Dav72] A. W. Davis. On the marginal distributions of the latent roots of the multivariate beta matrix. Annals of Mathematical Statistics, 43 (5):1664–1670, October 1972. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177692399.

Dayhoff:1964:EPS

[Day64] Eugene Dayhoff. On the equivalence of polykays of the second degree and Σ’s. Annals of Mathematical Statistics, 35(4):1663–1672, December 1964. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/

1177700389. See erratum [Day65].

Dayhoff:1965:CNC

[Day65] Eugene Dayhoff. Correction notes: Correction to “On the Equivalence of Polykays of the Second Degree and Σ’s”. Annals of Mathematical Statistics, 36(3):1069, June 1965. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177700086. See [Day64].

Dayhoff:1966:CNC

[Day66a] Eugene Dayhoff. Correction notes: Correction to “Generalized Polykays, an Extension of Simple Polykays and Bipolykays”. Annals of Mathematical Statistics, 37(3):746, June 1966. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177699478. See [Day66b].

Dayhoff:1966:GPE

[Day66b] Eugene Dayhoff. Generalized polykays, an extension of simple polykays and bipolykays. Annals of Mathematical Statistics, 37(1): 226–241, February 1966. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.

aoms/1177699612. See correction [Day66a].

Draper:1971:EFL

[DB71] Norman R. Draper and William J. Beggs. Errors in the factor levels and experimental design. Annals of Mathematical Statistics, 42(1): 46–58, February 1971. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177693493.

Deberghes:1968:ENS

[DBTL68] Hélène Deberghes and Bùi-Trong-Liẽu. On the ergodicity for nonstationary multiple Markov processes. Annals of Mathematical Statistics, 39(5):1448–1462, October 1968. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177698124.

Danziger:1964:TDF

[DD64] L. Danziger and S. A. Davis. Tables of distribution-free tolerance limits. Annals of Mathematical Statistics, 35(3):1361–1365, September 1964. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177703297.

DeLury:1938:NC

[De 38] D. B. De Lury. Note on correlations. Annals of Mathematical Statistics, 9(2):149–151, June 1938. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177732333.

DeGroot:1959:USE

[DeG59] Morris H. DeGroot. Unbiased sequential estimation for binomial populations. Annals of Mathematical Statistics, 30(1):80–101, March 1959. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177706361.

DeGroot:1960:MST

[DeG60] M. H. DeGroot. Minimax sequential tests of some composite hypotheses. Annals of Mathematical Statistics, 31(4):1193–1200, December 1960. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177705690.

DeGroot:1961:ECC

[DeG61] M. H. DeGroot. The essential completeness of the class of generalized sequential probability ratio tests. Annals of Mathematical Statistics, 32(2):602–605, June 1961. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177705066.

DeGroot:1962:UIS

[DeG62] M. H. DeGroot. Uncertainty, information, and sequential experiments. Annals of Mathematical Statistics, 33(2):404–419, June 1962. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177704567.

DeHardt:1970:NCG

[DeH70] John DeHardt. A necessary condition for Glivenko–Cantelli convergence in E_{n}. Annals of Mathematical Statistics, 41(6):2177–2178, December 1970. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177696726.

Dehardt:1971:GGC

[Deh71] J. Dehardt. Generalizations of the Glivenko–Cantelli theorem. Annals of Mathematical Statistics, 42(6):2050–2055, December 1971. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177693073.

Deming:1940:DPH

[Dem40] W. Edwards Deming. Discussion of Professor Hotelling’s paper. Annals of Mathematical Statistics, 11(4):470–471, December 1940. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177731834.

Dempster:1958:HDT

[Dem58] A. P. Dempster. A high dimensional two sample significance test. Annals of Mathematical Statistics, 29(4):995–1010, December 1958. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177706437.

Dempster:1959:GS

[Dem59] A. P. Dempster. Generalized D_{n}^{+} statistics. Annals of Mathematical Statistics, 30(2):593–597, June 1959. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177706275.

Dempster:1960:RAD

[Dem60] A. P. Dempster. Random allocation designs i: On general classes of estimation methods. Annals of Mathematical Statistics, 31(4):885– 905, December 1960. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177705665.

Dempster:1961:RAD

[Dem61] A. P. Dempster. Random allocation designs II: Approximate theory for simple random allocation. Annals of Mathematical Statistics, 32 (2):387–405, June 1961. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177705048.

Dempster:1963:FEI

[Dem63a] A. P. Dempster. Further examples of inconsistencies in the fiducial argument. Annals of Mathematical Statistics, 34(3):884–891, September 1963. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177704011.

Dempster:1963:MTG

[Dem63b] A. P. Dempster. Multivariate theory for general stepwise methods. Annals of Mathematical Statistics, 34(3):873–883, September 1963. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177704010.

Dempster:1963:PCI

[Dem63c] A. P. Dempster. On a paradox concerning inference about a covariance matrix. Annals of Mathematical Statistics, 34(4):1414–1418, December 1963. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177703873.

Dempster:1964:TET

[Dem64] A. P. Dempster. Tests for the equality of two covariance matrices in relation to a best linear discriminator analysis. Annals of Mathematical Statistics, 35(1):190–199, March 1964. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177703741.

Deming:1965:PPS

[Dem65] W. Edwards Deming. Principles of professional statistical practice. Annals of Mathematical Statistics, 36(6):1883–1900, December 1965. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177699824.

Dempster:1966:NMR

[Dem66] A. P. Dempster. New methods for reasoning towards posterior distributions based on sample data. Annals of Mathematical Statistics, 37 (2):355–374, April 1966. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177699517.

Dempster:1967:ULP

[Dem67] A. P. Dempster. Upper and lower probabilities induced by a multivalued mapping. Annals of Mathematical Statistics, 38(2):325– 339, April 1967. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177698950.

Dempster:1968:ULP

[Dem68] A. P. Dempster. Upper and lower probabilities generated by a random closed interval. Annals of Mathematical Statistics, 39(3):957– 966, June 1968. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177698328.

Dempster:1969:ULP

[Dem69] A. P. Dempster. Upper and lower probability inferences for families of hypotheses with monotone density ratios. Annals of Mathematical Statistics, 40(3):953–969, June 1969. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177697600.

Dempster:1972:CRC

[Dem72] A. P. Dempster. A class of random convex polytopes. Annals of Mathematical Statistics, 43(1):260–272, February 1972. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177692719.

denBroeder:1955:PET

[den55] George Gerard den Broeder, Jr. On parameter estimation for truncated Pearson Type III distributions. Annals of Mathematical Statistics, 26(4):659–663, December 1955. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177728426.

Denny:1964:CSS

[Den64] J. L. Denny. On continuous sufficient statistics. Annals of Mathematical Statistics, 35(3):1229–1233, September 1964. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177703280.

Denny:1966:CPC

[Den66] J. L. Denny. Convergence properties of convergence with probability one. Annals of Mathematical Statistics, 37(6):1800–1802, December 1966. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177699169.

Denny:1969:NTD

[Den69] J. L. Denny. Note on a theorem of Dynkin on the dimension of sufficient statistics. Annals of Mathematical Statistics, 40(4):1474– 1476, August 1969. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177697518.

Denny:1970:CES

[Den70] J. L. Denny. Cauchy’s equation and sufficient statistics on arcwise connected spaces. Annals of Mathematical Statistics, 41(2): 401–411, April 1970. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177697079.

Denardo:1971:MRP

[Den71] Eric V. Denardo. Markov renewal programs with small interest rates. Annals of Mathematical Statistics, 42(2):477–496, April 1971. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177693399.

Denny:1972:SSD

[Den72] J. L. Denny. Sufficient statistics and discrete exponential families. Annals of Mathematical Statistics, 43(4):1320–1322, August 1972. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177692483.

Department:1930:MTS

[Dep30] Statistical Department. A mathematical theory of seasonals. Annals of Mathematical Statistics, 1(1):57–72, February 1930. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177733260.

Derksen:1939:SIS

[Der39] J. B. D. Derksen. On some infinite series introduced by Tschuprow. Annals of Mathematical Statistics, 10(4):380–383, December 1939. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177732150.

Derman:1956:ACL

[Der56a] Cyrus Derman. An application of Chung’s lemma to the Kiefer– Wolfowitz stochastic approximation procedure. Annals of Mathematical Statistics, 27(2):532–536, June 1956. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177728277.

Derman:1956:SA

[Der56b] Cyrus Derman. Stochastic approximation. Annals of Mathematical Statistics, 27(4):879–886, December 1956. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177728065.

Derman:1957:NPE

[Der57] Cyrus Derman. Non-parametric up-and-down experimentation. Annals of Mathematical Statistics, 28(3):795–798, September 1957. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177706895.

Derman:1961:RCT

[Der61] C. Derman. Remark concerning two-state semi-Markov processes. Annals of Mathematical Statistics, 32(2):615–616, June 1961. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177705070.

Derman:1964:RDR

[Der64a] C. Derman. Review: D. R. Cox, Renewal Theory. Annals of Mathematical Statistics, 35(1):448–449, March 1964. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177703776.

Derman:1964:SCP

[Der64b] Cyrus Derman. On sequential control processes. Annals of Mathematical Statistics, 35(1):341–349, March 1964. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177703757.

Derman:1966:DSM

[Der66a] Cyrus Derman. Denumerable state Markovian decision processesaverage cost criterion. Annals of Mathematical Statistics, 37(6):1545– 1553, December 1966. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177699146.

Derman:1966:RGL

[Der66b] Cyrus Derman. Review: G. De Leve, Generalized Markovian Decision Processes — Parts I and II. Annals of Mathematical Statistics, 37 (6):1854–1855, December 1966. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177699184.

Deshpande:1970:CMD

[Des70a] Jayant V. Deshpande. A class of multisample distribution-free tests. Annals of Mathematical Statistics, 41(1):227–236, February 1970. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177697204.

Desu:1970:SP

[Des70b] M. Mahamunulu Desu. A selection problem. Annals of Mathematical Statistics, 41(5):1596–1603, October 1970. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177696803.

Desu:1971:CED

[Des71] M. M. Desu. A characterization of the exponential distribution by order statistics. Annals of Mathematical Statistics, 42(2):837–838, April 1971. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177693444.

Devinatz:1959:TLR

[Dev59] A. Devinatz. On a theorem of Lévy-Raikov. Annals of Mathematical Statistics, 30(2):583–586, June 1959. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177706272.

Dubins:1965:SFB

[DF65] Lester E. Dubins and David A. Freedman. A sharper form of the Borel–Cantelli lemma and the strong law. Annals of Mathematical Statistics, 36(3):800–807, June 1965. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177700054.

Dubins:1966:IPC

[DF66a] Lester E. Dubins and David A. Freedman. Invariant probabilities for certain Markov processes. Annals of Mathematical Statistics, 37(4): 837–848, August 1966. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177699364.

Dubins:1966:EVS

[DF66b] Lester E. Dubins and David A. Freedman. On the expected value of a stopped martingale. Annals of Mathematical Statistics, 37(6):1505– 1509, December 1966. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177699142.

DeGroot:1971:M

[DFG71] Morris H. DeGroot, Paul I. Feder, and Prem K. Goel. Matchmaking. Annals of Mathematical Statistics, 42(2):578–593, April 1971.

CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177693408.

Dharmadhikari:1968:BMM

[DFJ68] S. W. Dharmadhikari, V. Fabian, and K. Jogdeo. Bounds on the moments of martingales. Annals of Mathematical Statistics, 39(5):1719– 1723, October 1968. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177698154.

Dawson:1957:EMP

[DG57a] Reed Dawson and I. J. Good. Exact Markov probabilities from oriented linear graphs. Annals of Mathematical Statistics, 28(4):946– 956, December 1957. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177706795.

Durand:1957:RUV

[DG57b] David Durand and J. Arthur Greenwood. Random unit vectors II: Usefulness of Gram–Charlier and related series in approximating distributions. Annals of Mathematical Statistics, 28(4):978–986, December 1957. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177706798.

Draper:1968:FSO

[DH68] Norman R. Draper and Agnes M. Herzberg. Further second order rotatable designs. Annals of Mathematical Statistics, 39(6):1995–2001, December 1968. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177698027.

Dupac:1969:ANS

[DH69] Václav Dupač and Jaroslav Hájek. Asymptotic normality of simple linear rank statistics under alternatives II. Annals of Mathematical Statistics, 40(6):1992–2017, December 1969. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177697281.

deHaan:1970:NPH

[dH70] Laurens de Haan. Note on a paper by H. G. Tucker. Annals of Mathematical Statistics, 41(2):729–732, April 1970. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177697124. See [Tuc68].

Dharmadhikari:1963:FFM

[Dha63a] S. W. Dharmadhikari. Functions of finite Markov chains. Annals of Mathematical Statistics, 34(3):1022–1032, September 1963. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177704025.

Dharmadhikari:1963:SCS

[Dha63b] S. W. Dharmadhikari. Sufficient conditions for a stationary process to be a function of a finite Markov chain. Annals of Mathematical Statistics, 34(3):1033–1041, September 1963. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177704026.

Dharmadhikari:1964:EPW

[Dha64] S. W. Dharmadhikari. Exchangeable processes which are functions of stationary Markov chains. Annals of Mathematical Statistics, 35(1): 429–430, March 1964. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177703768.

Dharmadhikari:1965:CCF

[Dha65] S. W. Dharmadhikari. A characterisation of a class of functions of finite Markov chains. Annals of Mathematical Statistics, 36(2):524– 528, April 1965. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177700162.

Dharmadhikari:1968:SSS

[Dha68] S. W. Dharmadhikari. Splitting a single state of a stationary process into Markovian states. Annals of Mathematical Statistics, 39(3):1069– 1077, June 1968. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177698339.

Dharmadhikari:1969:NEP

[Dha69] S. W. Dharmadhikari. A note on exchangeable processes with states of finite rank. Annals of Mathematical Statistics, 40(6):2207–2208, December 1969. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177697301.

deHaan:1972:RGS

[dHH72] Laurens de Haan and Arie Hordijk. The rate of growth of sample maxima. Annals of Mathematical Statistics, 43(4):1185–1196, August 1972. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177692470.

Diamond:1963:LPC

[Dia63] Earl L. Diamond. The limiting power of categorical data chi-square tests analogous to normal analysis of variance. Annals of Mathematical Statistics, 34(4):1432–1441, December 1963. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177703875.

Dickey:1967:EDR

[Dic67a] James M. Dickey. Expansions of t densities and related complete integrals. Annals of Mathematical Statistics, 38(2):503–510, April 1967. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177698966.

Dickey:1967:MGM

[Dic67b] James M. Dickey. Matricvariate generalizations of the multivariate t distribution and the inverted multivariate t distribution. Annals of Mathematical Statistics, 38(2):511–518, April 1967. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177698967.

Dickey:1968:SEM

[Dic68a] James M. Dickey. Smoothed estimates for multinomial cell probabilities. Annals of Mathematical Statistics, 39(2):561–566, April 1968. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177698417.

Dickey:1968:TMI

[Dic68b] James M. Dickey. Three multidimensional-integral identities with Bayesian applications. Annals of Mathematical Statistics, 39(5):1615– 1628, October 1968. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177698143.

Dickey:1969:SC

[Dic69] James M. Dickey. Smoothing by cheating. Annals of Mathematical Statistics, 40(4):1477–1482, August 1969. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177697519.

Dickey:1971:WLR

[Dic71] James M. Dickey. The weighted likelihood ratio, linear hypotheses on normal location parameters. Annals of Mathematical Statistics, 42(1):204–223, February 1971. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177693507.

Dieulefait:1942:NMS

[Die42] Carlos E. Dieulefait. Note on a method of sampling. Annals of Mathematical Statistics, 13(1):94–97, March 1942. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177731648.

Dixon:1940:CTH

[Dix40] W. J. Dixon. A criterion for testing the hypothesis that two samples are from the same population. Annals of Mathematical Statistics, 11 (2):199–204, June 1940. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177731914.

Dixon:1944:FCP

[Dix44] Wilfrid J. Dixon. Further contributions to the problem of serial correlation. Annals of Mathematical Statistics, 15(2):119–144, June 1944. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177731279;   http://www.jstor.org/stable/2236194.

Dixon:1948:TNP

[Dix48] W. J. Dixon. Table of normal probabilities for intervals of various lengths and locations. Annals of Mathematical Statistics, 19(3):424– 426, September 1948. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177730212.

Dixon:1950:AEV

[Dix50] W. J. Dixon. Analysis of extreme values. Annals of Mathematical Statistics, 21(4):488–506, December 1950. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177729747.

Dixon:1951:RIE

[Dix51] W. J. Dixon. Ratios involving extreme values. Annals of Mathematical Statistics, 22(1):68–78, March 1951. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177729693.

Dixon:1953:PFS

[Dix53] W. J. Dixon. Power functions of the sign test and power efficiency for normal alternatives. Annals of Mathematical Statistics, 24(3):467– 473, September 1953. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177728986.

Dixon:1954:PUN

[Dix54] W. J. Dixon. Power under normality of several nonparametric tests. Annals of Mathematical Statistics, 25(3):610–614, September 1954. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177728732.

Dixon:1957:EMS

[Dix57] W. J. Dixon. Estimates of the mean and standard deviation of a normal population. Annals of Mathematical Statistics, 28(3):806– 809, September 1957. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177706898.

Dixon:1960:SEC

[Dix60] W. J. Dixon. Simplified estimation from censored normal samples. Annals of Mathematical Statistics, 31(2):385–391, June 1960. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177705900.

David:1951:MIE

[DJ51] F. N. David and N. L. Johnson. A method of investigating the effect of nonnormality and heterogeneity of variance on tests of the general linear hypothesis. Annals of Mathematical Statistics, 22(3):382–392, September 1951. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177729584.

David:1952:EMI

[DJ52] F. N. David and N. L. Johnson. Extension of a method of investigating the properties of analysis of variance tests to the case of random and mixed models. Annals of Mathematical Statistics, 23(4):594– 601, December 1952. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177729338.

David:1968:RRB

[DJ68] H. A. David and P. C. Joshi. Recurrence relations between moments of order statistics for exchangeable variates. Annals of Mathematical Statistics, 39(1):272–274, February 1968. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177698532.

Dharmadhikari:1969:BMC

[DJ69] S. W. Dharmadhikari and Kumar Jogdeo. Bounds on moments of certain random variables. Annals of Mathematical Statistics, 40(4): 1506–1509, August 1969. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177697526.

Derman:1959:CSP

[DJL59] C. Derman, M. V. Johns, Jr., and G. J. Lieberman. Continuous sampling procedures without control. Annals of Mathematical Statistics, 30(4):1175–1191, December 1959. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177706103.

David:1956:WST

[DK56] Herbert T. David and William H. Kruskal. The WAGR sequential t-test reaches a decision with probability one. Annals of Mathematical Statistics, 27(3):797–805, September 1956. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http:

//projecteuclid.org/euclid.aoms/1177728186. See correction

[DK58].

David:1958:CWS

[DK58] Herbert T. David and William H. Kruskal. Correction to “The WAGR Sequential T -Test Reaches a Decision with Probability One”. Annals of Mathematical Statistics, 29(3):936, September 1958. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177706559. See

[DK56].

Dwass:1963:CLT

[DK63] Meyer Dwass and Samuel Karlin. Conditioned limit theorems. Annals of Mathematical Statistics, 34(4):1147–1167, December 1963. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177703854.

Deely:1968:CSE

[DK68a] J. J. Deely and R. L. Kruse. Construction of sequences estimating the mixing distribution. Annals of Mathematical Statistics, 39(1): 286–288, February 1968. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177698536.

Dempster:1968:DDC

[DK68b] A. P. Dempster and Robert M. Kleyle. Distributions determined by cutting a simplex with hyperplanes. Annals of Mathematical Statistics, 39(5):1473–1478, October 1968. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177698126.

Dupac:1972:RMP

[DK72] Václav Dupač and František Král. Robbins–Monro procedure with both variables subject to experimental error. Annals of Mathematical Statistics, 43(4):1089–1095, August 1972. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177692462.

Dvoretzky:1953:SDPb

[DKW53a] A. Dvoretzky, J. Kiefer, and J. Wolfowitz. Sequential decision problems for processes with continuous time parameter. problems of estimation. Annals of Mathematical Statistics, 24(3):403–415, September 1953. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177728980.

Dvoretzky:1953:SDPa

[DKW53b] A. Dvoretzky, J. Kiefer, and J. Wolfowitz. Sequential decision problems for processes with continuous time parameter. Testing hypotheses. Annals of Mathematical Statistics, 24(2):254–264, June 1953. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177729031.

Dvoretzky:1956:AMC

[DKW56] A. Dvoretzky, J. Kiefer, and J. Wolfowitz. Asymptotic minimax character of the sample distribution function and of the classical multinomial estimator. Annals of Mathematical Statistics, 27(3):642–669, September 1956. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177728174.

Dvoretzky:1959:CNC

[DKW59] A. Dvoretzky, J. Kiefer, and J. Wolfowitz. Correction notes: Corrections to “Sequential Decision Problems for Processes with Continuous Time Parameter-Testing Hypothesis”. Annals of Mathematical Statistics, 30(4):1265, December 1959. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177706115.

Dym:1966:MDB

[DL66] Harry Dym and Eugene M. Luks. On the mean duration of a ball and cell game; a first passage problem. Annals of Mathematical Statistics, 37(2):517–521, April 1966. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177699537.

Dickey:1970:CNC

[DL70a] James Dickey and B. P. Lientz. Correction notes: Correction to “The Weighted Likelihood Ratio, Sharp Hypotheses About Chances, the Order of a Markov Chain”. Annals of Mathematical Statistics, 41 (4):1374, August 1970. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177696919. See [DL70b].

Dickey:1970:WLR

[DL70b] James M. Dickey and B. P. Lientz. The weighted likelihood ratio, sharp hypotheses about chances, the order of a Markov chain. Annals of Mathematical Statistics, 41(1):214–226, February 1970. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177697203. See correc-

tions [DL70a, DL72].

Dickey:1972:CNC

[DL72] James Dickey and B. P. Lientz. Correction notes: Correction to “The Weighted Likelihood Ratio, Sharp Hypotheses About Chances, the Order of a Markov Chain”. Annals of Mathematical Statistics, 43 (4):1380, August 1972. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177692496. See [DL70b].

delaGarza:1954:SIP

[dlG54] A. de la Garza. Spacing of information in polynomial regression. Annals of Mathematical Statistics, 25(1):123–130, March 1954. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177728851.

Derman:1957:TML

[DLS57] C. Derman, S. Littauer, and H. Solomon. Tightened multi-level continuous sampling plans. Annals of Mathematical Statistics, 28(2): 395–404, June 1957. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177706967.

Darling:1972:OSP

[DLT72] D. A. Darling, T. Liggett, and H. M. Taylor. Optimal stopping for partial sums. Annals of Mathematical Statistics, 43(4):1363–1368, August 1972. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177692491.

Dwyer:1948:SMD

[DM48] Paul S. Dwyer and M. S. Macphail. Symbolic matrix derivatives. Annals of Mathematical Statistics, 19(4):517–534, December 1948. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177730148.

Draper:1967:CSD

[DM67] Norman R. Draper and Toby J. Mitchell. The construction of saturated 2_{R}^{k−p} designs. Annals of Mathematical Statistics, 38(4):1110– 1126, August 1967. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177698780.

Denardo:1968:OCD

[DM68a] E. V. Denardo and B. L. Miller. An optimality condition for discrete dynamic programming with no discounting. Annals of Mathematical Statistics, 39(4):1220–1227, August 1968. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177698247.

Draper:1968:CSR

[DM68b] Norman R. Draper and Toby J. Mitchell. Construction of the set of 256-run designs of resolution ≥ 5 and the set of even 512-run designs of resolution ≥ 6 with special reference to the unique Saturated designs. Annals of Mathematical Statistics, 39(1):246–255, February 1968. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177698525.

Draper:1970:CSR

[DM70a] Norman R. Draper and Toby J. Mitchell. Construction of a set of 512-run designs of resolution ≥ 5 and a set of even 1024-run designs of resolution ≥ 6. Annals of Mathematical Statistics, 41(3): 876–887, June 1970. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177696965.

Dym:1970:EIS

[DM70b] H. Dym and H. P. McKean. Extrapolation and interpolation of stationary Gaussian processes. Annals of Mathematical Statistics, 41 (6):1817–1844, December 1970. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177696684.

Das:1962:CRD

[DN62] M. N. Das and V. L. Narasimham. Construction of rotatable designs through balanced incomplete block designs. Annals of Mathematical Statistics, 33(4):1421–1439, December 1962. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177704374.

Dharmadhikari:1970:SRN

[DN70] S. W. Dharmadhikari and M. G. Nadkarni. Some regular and non-regular functions of finite Markov chains. Annals of Mathematical Statistics, 41(1):207–213, February 1970. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177697202.

deOliveira:1966:QLI

[dO66] J. Tiago de Oliveira. Quasi-linearly invariant prediction. Annals of Mathematical Statistics, 37(6):1684–1687, December 1966. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177699157.

Dodd:1930:ULF

[Dod30] Edward L. Dodd. The use of linear functions to detect hidden periods in data separated into small sets. Annals of Mathematical Statistics, 1(3):205–223, August 1930. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177733089.

Dodd:1936:SES

[Dod36] Stuart C. Dodd. The standard error of a “Social Force”. Annals of Mathematical Statistics, 7(4):202–209, December 1936. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177732481.

Dodd:1937:IEM

[Dod37] Edward L. Dodd. Internal and external means arising from the scaling of frequency functions. Annals of Mathematical Statistics, 8(1):12– 20, March 1937. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177732450.

Dodd:1938:IEM

[Dod38] Edward L. Dodd. Interior and exterior means obtained by the method of moments. Annals of Mathematical Statistics, 9(3):153– 157, September 1938. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177732306.

Dodd:1939:LCW

[Dod39] Edward L. Dodd. The length of the cycles which result from the graduation of chance elements. Annals of Mathematical Statistics, 10 (3):254–264, September 1939. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177732184.

Dodd:1940:SMC

[Dod40] Edward L. Dodd. The substitutive mean and certain subclasses of this general mean. Annals of Mathematical Statistics, 11(2):163– 176, June 1940. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177731910.

Dodd:1941:CEL

[Dod41a] Edward L. Dodd. The cyclic effects of linear graduations persisting in the differences of the graduated values. Annals of Mathematical Statistics, 12(2):127–136, June 1941. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177731744.

Dodd:1941:SGL

[Dod41b] Edward L. Dodd. Some generalizations of the logarithmic mean and of similar means of two variates which become indeterminate when the two variates are equal. Annals of Mathematical Statistics, 12 (4):422–428, December 1941. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177731680.

Dodge:1943:SIP

[Dod43] H. F. Dodge. A sampling inspection plan for continuous production. Annals of Mathematical Statistics, 14(3):264–279, September 1943. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177731420. See

note [Lie53a].

Doksum:1966:AMD

[Dok66] Kjell Doksum. Asymptotically minimax distribution-free procedures. Annals of Mathematical Statistics, 37(3):619–628, June 1966. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177699455.

Doksum:1967:AOS

[Dok67a] Kjell Doksum. Asymptotically optimal statistics in some models with increasing failure rate averages. Annals of Mathematical Statistics, 38 (6):1731–1739, December 1967. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177698607. See correction [Dok68].

Doksum:1967:RPS

[Dok67b] Kjell Doksum. Robust procedures for some linear models with one observation per cell. Annals of Mathematical Statistics, 38(3):878– 883, June 1967. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177698881.

Doksum:1968:CNC

[Dok68] Kjell Doksum. Correction notes: Correction to “Asymptotically Optimal Statistics in Some Models with Increasing Failure Rate Averages”. Annals of Mathematical Statistics, 39(2):684–685, April 1968. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177698431. See

[Dok67a].

Doksum:1969:MRI

[Dok69a] Kjell Doksum. Minimax results for IFRA scale alternatives. Annals of Mathematical Statistics, 40(5):1778–1783, October 1969. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177697390.

Doksum:1969:STP

[Dok69b] Kjell Doksum. Starshaped transformations and the power of rank tests. Annals of Mathematical Statistics, 40(4):1167–1176, August 1969. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177697493.

Doleans:1969:VQM

[Dol69] Catherine Doléans. Variation quadratique des martingales continues à droite. (French) [Quadratic variation of right-continuous martingales]. Annals of Mathematical Statistics, 40(1):284–289, February 1969. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177697823.

Donsker:1952:JED

[Don52] Monroe D. Donsker. Justification and extension of Doob’s heuristic approach to the Kolmogorov–Smirnov theorems. Annals of Mathematical Statistics, 23(2):277–281, June 1952. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177729445.

Doob:1935:LDC

[Doo35] J. L. Doob. The limiting distributions of certain statistics. Annals of Mathematical Statistics, 6(3):160–169, September 1935. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177732594.

Doob:1941:PM

[Doo41] J. L. Doob. Probability as measure. Annals of Mathematical Statistics, 12(2):206–214, June 1941. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177731749.

Doob:1944:EGP

[Doo44] J. L. Doob. The elementary Gaussian processes. Annals of Mathematical Statistics, 15(3):229–282, September 1944. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177731234.

Doob:1949:HAK

[Doo49] J. L. Doob. Heuristic approach to the Kolmogorov–Smirnov theorems. Annals of Mathematical Statistics, 20(3):393–403, September 1949. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177729991;   http://www.jstor.org/stable/2236535.

Doob:1951:RPI

[Doo51] J. L. Doob. Report of the President of the Institute for 1950. Annals of Mathematical Statistics, 22(1):159–160, March 1951. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177729712.

Doob:1968:REN

[Doo68] J. L. Doob. Review: Edward Nelson, Dynamical Theories of Brownian Motion. Annals of Mathematical Statistics, 39(2):686, April 1968. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177698432.

Dorfman:1943:DDM

[Dor43] Robert Dorfman. The detection of defective members of large populations. Annals of Mathematical Statistics, 14(4):436–440, December 1943. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177731363.

Doss:1969:CLE

[Dos69] D. C. Doss. Characterizations of the linear exponential family in a parameter by recurrence relations for functions of cumulants. Annals of Mathematical Statistics, 40(5):1721–1727, October 1969. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177697383.

Downton:1954:LSE

[Dow54] F. Downton. Least-squares estimates using ordered observations. Annals of Mathematical Statistics, 25(2):303–316, June 1954. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177728787.

DeGroot:1963:BEC

[DR63] M. H. DeGroot and M. M. Rao. Bayes estimation with convex loss. Annals of Mathematical Statistics, 34(3):839–846, September 1963. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177704006.

Darling:1967:FSF

[DR67] D. A. Darling and Herbert Robbins. Finding the size of a finite population. Annals of Mathematical Statistics, 38(5):1392–1398, October 1967. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177698695.

Darroch:1972:GIS

[DR72] J. N. Darroch and D. Ratcliff. Generalized iterative scaling for loglinear models. Annals of Mathematical Statistics, 43(5):1470–1480, October 1972. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177692379.

Draper:1960:SOR

[Dra60a] Norman R. Draper. Second order rotatable designs in four or more dimensions. Annals of Mathematical Statistics, 31(1):23–33, March 1960. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177705984.

Draper:1960:TORb

[Dra60b] Norman R. Draper. A third order rotatable design in four dimensions. Annals of Mathematical Statistics, 31(4):875–877, December 1960. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177705663.

Draper:1960:TORa

[Dra60c] Norman R. Draper. Third order rotatable designs in three dimensions. Annals of Mathematical Statistics, 31(4):865–874, December 1960. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177705662.

Draper:1961:TOR

[Dra61] Norman R. Draper. Third order rotatable designs in three dimensions: Some specific designs. Annals of Mathematical Statistics, 32 (3):910–913, September 1961. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177704989.

Dressel:1940:SST

[Dre40] Paul L. Dressel. Statistical seminvariants and their estimates with particular emphasis on their relation to algebraic invariants. Annals of Mathematical Statistics, 11(1):33–57, March 1940. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177731940.

Dressel:1941:SMO

[Dre41] Paul L. Dressel. A symmetric method of obtaining unbiased estimates and expected values. Annals of Mathematical Statistics, 12(1):84– 90, March 1941. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177731787.

Drion:1952:SDF

[Dri52] E. F. Drion. Some distribution-free tests for the difference between two empirical cumulative distribution functions. Annals of Mathematical Statistics, 23(4):563–574, December 1952. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177729335.

Drogin:1972:IPM

[Dro72] Richard Drogin. An invariance principle for martingales. Annals of Mathematical Statistics, 43(2):602–620, April 1972. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177692640.

Deming:1940:LSA

[DS40] W. Edwards Deming and Frederick F. Stephan. On a least squares adjustment of a sampled frequency table when the expected marginal totals are known. Annals of Mathematical Statistics, 11(4):427–444, December 1940. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177731829.

Darling:1953:FPP

[DS53] D. A. Darling and A. J. F. Siegert. The first passage problem for a continuous Markov process. Annals of Mathematical Statistics, 24 (4):624–639, December 1953. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177728918.

Derman:1959:DSA

[DS59] C. Derman and J. Sacks. On Dvoretzky’s stochastic approximation theorem. Annals of Mathematical Statistics, 30(2):601–606, June 1959. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177706277.

Darroch:1963:TMT

[DS63] J. N. Darroch and S. D. Silvey. On testing more than one hypothesis. Annals of Mathematical Statistics, 34(2):555–567, June 1963. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177704168.

Derman:1966:NMR

[DS66] Cyrus Derman and Ralph E. Strauch. A note on memoryless rules for controlling sequential control processes. Annals of Mathematical Statistics, 37(1):276–278, February 1966. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177699618.

DeGroot:1969:OTS

[DS69a] M. H. DeGroot and N. Starr. Optimal two-stage stratified sampling. Annals of Mathematical Statistics, 40(2):575–582, April 1969. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177697725.

Dubman:1969:EPT

[DS69b] M. Dubman and B. Sherman. Estimation of parameters in a transient Markov chain arising in a reliability growth model. Annals of Mathematical Statistics, 40(5):1542–1556, October 1969. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177697372.

Dwass:1957:IDR

[DT57] Meyer Dwass and Henry Teicher. On infinitely divisible random vectors. Annals of Mathematical Statistics, 28(2):461–470, June 1957. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177706974.

Dwyer:1964:CMP

[DT64] P. S. Dwyer and D. S. Tracy. A combinatorial method for products of two polykays with some general formulae. Annals of Mathematical Statistics, 35(3):1174–1185, September 1964. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177703275.

Dubins:1967:OSW

[DT67] Lester E. Dubins and Henry Teicher. Optimal stopping when the future is discounted. Annals of Mathematical Statistics, 38(2):601– 605, April 1967. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177698978.

Deo:1968:NWL

[DT68] C. M. Deo and D. R. Truax. A note on the weak law. Annals of Mathematical Statistics, 39(6):2159–2160, December 1968. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177698050.

Doksum:1971:PBA

[DT71] Kjell Doksum and Rory Thompson. Power bounds and asymptotic minimax results for one-sample rank tests. Annals of Mathematical Statistics, 42(1):12–34, February 1971. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177693491.

Dubins:1966:NUS

[Dub66] Lester E. Dubins. A note on upcrossings of semimartingales. Annals of Mathematical Statistics, 37(3):728, June 1966. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177699469.

Dubins:1968:TS

[Dub68a] Lester E. Dubins. On a theorem of Skorohod. Annals of Mathematical Statistics, 39(6):2094–2097, December 1968. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177698036.

Dubins:1968:SPS

[Dub68b] Lester E. Dubins. A simpler proof of Smith’s roulette theorem. Annals of Mathematical Statistics, 39(2):390–393, April 1968. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177698402.

Dubins:1972:SBT

[Dub72] Lester E. Dubins. Sharp bounds for the total variance of uniformly bounded semimartingales. Annals of Mathematical Statistics, 43 (5):1559–1565, October 1972. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177692388.

Dudley:1965:GPS

[Dud65] R. M. Dudley. Gaussian processes on several parameters. Annals of Mathematical Statistics, 36(3):771–788, June 1965. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177700052.

Dudley:1968:DPM

[Dud68] R. M. Dudley. Distances of probability measures and random variables. Annals of Mathematical Statistics, 39(5):1563–1572, October 1968. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177698137.

Dudewicz:1969:ASS

[Dud69a] Edward J. Dudewicz. An approximation to the sample size in selection problems. Annals of Mathematical Statistics, 40(2):492–497, April 1969. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177697715.

Dudley:1969:RKR

[Dud69b] R. M. Dudley. Review: K. R. Parthasarathy, Probability Measures on Metric Spaces. Annals of Mathematical Statistics, 40(1):328, February 1969. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177697834.

Dudley:1969:SMG

[Dud69c] R. M. Dudley. The speed of mean Glivenko–Cantelli convergence. Annals of Mathematical Statistics, 40(1):40–50, February 1969. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177697802. See note

[Gän70].

Dugue:1961:GD

[Dug61] D. Dugué. Georges Darmois, 1888–1960. Annals of Mathematical Statistics, 32(2):357–360, June 1961. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177705046.

Dunlap:1931:EDD

[Dun31] Hilda Frost Dunlap. An empirical determination of the distribution of means, standard deviations and correlation coefficients drawn from rectangular populations. Annals of Mathematical Statistics, 2(1):66– 81, February 1931. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177733043.

Dunn:1958:EMD

[Dun58] Olive Jean Dunn. Estimation of the means of dependent variables. Annals of Mathematical Statistics, 29(4):1095–1111, December 1958. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177706443.

Dunn:1959:EMD

[Dun59] Olive Jean Dunn. Estimation of the medians for dependent variables. Annals of Mathematical Statistics, 30(1):192–197, March 1959. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177706374.

Duncan:1961:BRC

[Dun61] David B. Duncan. Bayes rules for a common multiple comparisons problem and related Student-t problems. Annals of Mathematical Statistics, 32(4):1013–1033, December 1961. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177704842.

Dunn:1965:PMD

[Dun65] Olive Jean Dunn. A property of the multivariate t distribution. Annals of Mathematical Statistics, 36(2):712–714, April 1965. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177700183.

Duncan:1970:ACM

[Dun70] Tyrone E. Duncan. On the absolute continuity of measures. Annals of Mathematical Statistics, 41(1):30–38, February 1970. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177697185.

Dupac:1965:DSA

[Dup65] Václav Dupač. A dynamic stochastic approximation method. Annals of Mathematical Statistics, 36(6):1695–1702, December 1965. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177699797.

Dupac:1970:RMW

[Dup70] Vaclav Dupac. Review: M. T. Wasan, Stochastic Approximation. Annals of Mathematical Statistics, 41(3):1131–1132, June 1970. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177696999.

Durbin:1968:PSD

[Dur68] J. Durbin. The probability that the sample distribution function lies between two parallel straight lines. Annals of Mathematical Statistics, 39(2):398–411, April 1968. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177698404.

Durham:1971:PCG

[Dur71] S. D. Durham. A problem concerning generalized age-dependent branching processes with immigration. Annals of Mathematical Statistics, 42(3):1121–1123, June 1971. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177693344.

Deemer:1955:EPT

[DV55] Walter L. Deemer, Jr. and David F. Votaw, Jr. Estimation of parameters of truncated or censored exponential distributions. Annals of Mathematical Statistics, 26(3):498–504, September 1955. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177728494.

Derman:1967:SCS

[DV67] Cyrus Derman and Arthur F. Veinott, Jr. A solution to a countable system of equations arising in Markovian decision processes. Annals of Mathematical Statistics, 38(2):582–584, April 1967. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177698973.

Dvoretzky:1949:SSS

[Dvo49] Aryeh Dvoretzky. On the strong stability of a sequence of events. Annals of Mathematical Statistics, 20(2):296–299, June 1949. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177730039.

Dantzig:1951:FLN

[DW51] George B. Dantzig and Abraham Wald. On the fundamental lemma of Neyman and Pearson. Annals of Mathematical Statistics, 22(1): 87–93, March 1951. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177729695.

David:1965:CD

[DW65] H. A. David and F. W. Wolock. Cyclic designs. Annals of Mathematical Statistics, 36(5):1526–1534, October 1965. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177699912.

deWaal:1970:DCM

[dW70] D. J. de Waal. Distributions connected with a multivariate beta statistic. Annals of Mathematical Statistics, 41(3):1091–1095, June 1970. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/

1177696989. See correction [dW71].

deWaal:1971:CNC

[dW71] D. J. de Waal. Correction note: Correction to “Distributions Connected with a Multivariate Beta Statistic”. Annals of Mathematical Statistics, 42(6):2165–2166, December 1971. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177693088. See [dW70].

deWaal:1972:EVE

[dW72] D. J. de Waal. On the expected values of the elementary symmetric functions of a noncentral Wishart matrix. Annals of Mathematical Statistics, 43(1):344–347, February 1972. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177692728.

Dwass:1953:ANC

[Dwa53] Meyer Dwass. On the asymptotic normality of certain rank order statistics. Annals of Mathematical Statistics, 24(2):303–306, June 1953. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177729038.

Dwass:1955:NSC

[Dwa55a] Meyer Dwass. A note on simultaneous confidence intervals. Annals of Mathematical Statistics, 26(1):146–147, March 1955. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177728607.

Dwass:1955:ANS

[Dwa55b] Meyer Dwass. On the asymptotic normality of some statistics used in non-parametric tests. Annals of Mathematical Statistics, 26(2): 334–339, June 1955. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177728552.

Dwass:1956:LSP

[Dwa56] Meyer Dwass. The large-sample power of rank order tests in the two-sample problem. Annals of Mathematical Statistics, 27(2):352– 374, June 1956. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177728263.

Dwass:1957:MRT

[Dwa57a] Meyer Dwass. Modified randomization tests for nonparametric hypotheses. Annals of Mathematical Statistics, 28(1):181–187, March 1957. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177707045.

Dwass:1957:DRC

[Dwa57b] Meyer Dwass. On the distribution of ranks and of certain rank order statistics. Annals of Mathematical Statistics, 28(2):424–431, June 1957. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177706970.

Dwass:1958:SSR

[Dwa58] Meyer Dwass. On several statistics related to empirical distribution functions. Annals of Mathematical Statistics, 29(1):188–191, March 1958. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177706715.

Dwass:1959:DGS

[Dwa59] Meyer Dwass. The distribution of a generalized D_{n}^{+} statistic. Annals of Mathematical Statistics, 30(4):1024–1028, December 1959. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177706085.

Dwass:1962:FTC

[Dwa62] Meyer Dwass. A fluctuation theorem for cyclic random variables. Annals of Mathematical Statistics, 33(4):1450–1454, December 1962. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177704376.

Dwass:1964:EP

[Dwa64] Meyer Dwass. Extremal processes. Annals of Mathematical Statistics, 35(4):1718–1725, December 1964. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177700394.

Dwass:1967:RUP

[Dwa67a] Meyer Dwass. Review: N. U. Prabhu, Stochastic Processes. Annals of Mathematical Statistics, 38(2):626, April 1967. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177698989.

Dwass:1967:SRW

[Dwa67b] Meyer Dwass. Simple random walk and rank order statistics. Annals of Mathematical Statistics, 38(4):1042–1053, August 1967. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177698773.

Dwass:1972:UCT

[Dwa72] Meyer Dwass. Unbiased coin tossing with discrete random variables. Annals of Mathematical Statistics, 43(3):860–864, June 1972. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177692551.

Dvoretzky:1951:ERC

[DWW51] A. Dvoretzky, A. Wald, and J. Wolfowitz. Elimination of randomization in certain statistical decision procedures and zero-sum two-person games. Annals of Mathematical Statistics, 22(1):1–21, March 1951. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177729689.

Dwyer:1937:MRI

[Dwy37a] Paul S. Dwyer. Moments of any rational integral isobaric sample moment function. Annals of Mathematical Statistics, 8(1):21–65, March 1937. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177732451.

Dwyer:1937:SCG

[Dwy37b] Paul S. Dwyer. The simultaneous computation of groups of regression equations and associated multiple correlation coefficients. Annals of Mathematical Statistics, 8(4):224–231, December 1937. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177732391.

Dwyer:1938:CEPa

[Dwy38a] Paul S. Dwyer. Combined expansions of products of symmetric power sums and of sums of symmetric power products with application to sampling. Annals of Mathematical Statistics, 9(1):1–47, March 1938. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177732357.

Dwyer:1938:CMU

[Dwy38b] Paul S. Dwyer. The computation of moments with the use of cumulative totals. Annals of Mathematical Statistics, 9(4):288–304, December 1938. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177732286.

Dwyer:1938:CEPb

[Dwy38c] Paul S. Dwyer. On combined expansions of products of symmetric power sums and of sums of symmetric power products with applications to sampling (continued). Annals of Mathematical Statistics, 9 (2):97–132, June 1938. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177732331.

Dwyer:1940:CFT

[Dwy40a] P. S. Dwyer. Combinatorial formulas for the r th standard moment of the sample sum, of the sample mean, and of the normal curve. Annals of Mathematical Statistics, 11(3):353–355, September 1940. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177731873.

Dwyer:1940:CNT

[Dwy40b] P. S. Dwyer. The cumulative numbers and their polynomials. Annals of Mathematical Statistics, 11(1):66–71, March 1940. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177731942.

Dwyer:1941:SRL

[Dwy41a] P. S. Dwyer. The skewness of the residuals in linear regression theory. Annals of Mathematical Statistics, 12(1):104–110, March 1941. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177731790.

Dwyer:1941:DT

[Dwy41b] Paul S. Dwyer. The Doolittle technique. Annals of Mathematical Statistics, 12(4):449–458, December 1941. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177731682.

Dwyer:1942:GM

[Dwy42] P. S. Dwyer. Grouping methods. Annals of Mathematical Statistics, 13(2):138–155, June 1942. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177731602.

Dwyer:1944:RWM

[Dwy44a] P. S. Dwyer. Report on the Wellesley meeting of the Institute. Annals of Mathematical Statistics, 15(3):338–339, September 1944. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177731245.

Dwyer:1944:MPL

[Dwy44b] Paul S. Dwyer. A matrix presentation of least squares and correlation theory with matrix justification of improved methods of solution. Annals of Mathematical Statistics, 15(1):82–89, March 1944. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177731317.

Dwyer:1945:RRM

[Dwy45a] P. S. Dwyer. Report of the Rutgers meeting of the Institute. Annals of Mathematical Statistics, 16(3):317–318, September 1945. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177731099.

Dwyer:1945:ARS

[Dwy45b] Paul S. Dwyer. Annual report of the Secretary-Treasurer of the Institute. Annals of Mathematical Statistics, 16(1):105–106, March 1945. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177731180.

Dwyer:1946:RPM

[Dwy46a] P. S. Dwyer. Report of the Princeton meeting of the Institute. Annals of Mathematical Statistics, 17(4):516–517, December 1946. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177730894.

Dwyer:1946:RIM

[Dwy46b] P. S. Dwyer. Report on the Ithaca meeting of the Institute. Annals of Mathematical Statistics, 17(4):514–515, December 1946. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177730893.

Dwyer:1946:RCM

[Dwy46c] Paul S. Dwyer. Report on the Cleveland meeting of the Institute. Annals of Mathematical Statistics, 17(1):92–94, March 1946. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177731031.

Dwyer:1947:RAC

[Dwy47a] P. S. Dwyer. Report on the Atlantic City meeting of the Institute. Annals of Mathematical Statistics, 18(2):306–308, June 1947. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177730454.

Dwyer:1947:RBM

[Dwy47b] P. S. Dwyer. Report on the Boston meeting of the Institute. Annals of Mathematical Statistics, 18(1):149, March 1947. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177730507.

Dwyer:1947:RST

[Dwy47c] Paul S. Dwyer. Report of the Secretary-Treasurer of the Institute for 1946. Annals of Mathematical Statistics, 18(1):156–158, March 1947. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177730509.

Dwyer:1948:ANC

[Dwy48a] P. S. Dwyer. Adoption of the new constitution. Annals of Mathematical Statistics, 19(4):607, December 1948. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177730159.

Dwyer:1948:RST

[Dwy48b] Paul S. Dwyer. Report of the Secretary-Treasurer of the Institute for 1947. Annals of Mathematical Statistics, 19(1):141–143, March 1948. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177730308.

Dwyer:1948:RNY

[Dwy48c] Paul S. Dwyer. Report on the New York meeting of the Institute. Annals of Mathematical Statistics, 19(1):133–136, March 1948. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177730305.

Dwyer:1949:EOC

[Dwy49a] Paul S. Dwyer. Election of officers and council and revision of bylaws. Annals of Mathematical Statistics, 20(1):150–151, March 1949. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177730108.

Dwyer:1949:PCC

[Dwy49b] Paul S. Dwyer. Pearsonian correlation coefficients associated with least squares theory. Annals of Mathematical Statistics, 20(3):404– 416, September 1949. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177729992.

Dwyer:1949:RST

[Dwy49c] Paul S. Dwyer. Report of the Secretary-Treasurer of the Institute for 1948. Annals of Mathematical Statistics, 20(1):160–163, March 1949. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177730112.

Dwyer:1952:RPI

[Dwy52] Paul S. Dwyer. Report of the President of the Institute for 1951. Annals of Mathematical Statistics, 23(1):160–164, March 1952. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177729498.

Dwyer:1958:GGT

[Dwy58] Paul S. Dwyer. Generalizations of a Gaussian theorem. Annals of Mathematical Statistics, 29(1):106–117, March 1958. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177706708. See correc-

tion [Dwy60].

Dwyer:1960:CNC

[Dwy60] Paul S. Dwyer. Correction notes: Correction to “Generalizations of a Gaussian Theorem”. Annals of Mathematical Statistics, 31(1): 227, March 1960. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/

1177706001. See [Dwy58].

Dwyer:1964:PPD

[Dwy64] P. S. Dwyer. Properties of polykays of deviates. Annals of Mathematical Statistics, 35(3):1167–1173, September 1964. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177703274.

Dykstra:1970:CCE

[Dyk70a] Richard L. Dykstra. A characterization of a conditional expectation with respect to a σ-lattice. Annals of Mathematical Statistics, 41(2): 698–701, April 1970. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177697117.

Dykstra:1970:EPD

[Dyk70b] Richard L. Dykstra. Establishing the positive definiteness of the sample covariance matrix. Annals of Mathematical Statistics, 41(6):2153– 2154, December 1970. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177696719.

Dym:1966:NLT

[Dym66] Harry Dym. A note on limit theorems for the entropy of Markov chains. Annals of Mathematical Statistics, 37(2):522–524, April 1966. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177699538.

Dym:1970:RRM

[Dym70] Harry Dym. Review: R. M. Blumenthal, R. K. Getoor, Markov Processes and Potential Theory. Annals of Mathematical Statistics, 41 (2):749–750, April 1970. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177697131.

Darst:1967:PMS

[DZ67] R. B. Darst and R. E. Zink. A perfect measurable space that is not a Lusin space. Annals of Mathematical Statistics, 38(6):1918, December 1967. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177698628.

Eagleson:1964:PEB

[Eag64] G. K. Eagleson. Polynomial expansions of bivariate distributions. Annals of Mathematical Statistics, 35(3):1208–1215, September 1964. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177703278.

Eaton:1967:GVT

[Eat67a] Morris L. Eaton. The generalized variance: Testing and ranking problem. Annals of Mathematical Statistics, 38(3):941–943, June 1967. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177698892.

Eaton:1967:SOP

[Eat67b] Morris L. Eaton. Some optimum properties of ranking procedures. Annals of Mathematical Statistics, 38(1):124–137, February 1967.

CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177699063.

Eaton:1970:CCT

[Eat70a] Morris L. Eaton. A complete class theorem for multidimensional onesided alternatives. Annals of Mathematical Statistics, 41(6):1884– 1888, December 1970. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177696689.

Eaton:1970:GME

[Eat70b] Morris L. Eaton. Gauss–Markov estimation for multivariate linear models: A coordinate free approach. Annals of Mathematical Statistics, 41(2):528–538, April 1970. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177697093.

Eaton:1970:NSB

[Eat70c] Morris L. Eaton. A note on symmetric Bernoulli random variables. Annals of Mathematical Statistics, 41(4):1223–1226, August 1970. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177696897.

Eaves:1967:SFG

[Eav67] D. M. Eaves. Sample functions of Gaussian random homogeneous fields and either continuous or very irregular. Annals of Mathematical Statistics, 38(5):1579–1582, October 1967. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177698713.

Eaves:1969:CLP

[Eav69] D. M. Eaves. The central limit problem for generalized random fields. Annals of Mathematical Statistics, 40(1):203–215, February 1969. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177697817.

Epstein:1944:SSD

[EC44] Benjamin Epstein and C. West Churchman. On the statistics of sensitivity data. Annals of Mathematical Statistics, 15(1):90–96, March 1944. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177731318.

Eckler:1955:RS

[Eck55] Albert Ross Eckler. Rotation sampling. Annals of Mathematical Statistics, 26(4):664–685, December 1955. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177728427.

Efron:1965:IPP

[Efr65a] Bradley Efron. Increasing properties of Pólya frequency function. Annals of Mathematical Statistics, 36(1):272–279, February 1965. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177700288.

Efron:1965:NDP

[Efr65b] Bradley Efron. Note on decision procedures for finite decision problems under complete ignorance. Annals of Mathematical Statistics, 36 (2):691–697, April 1965. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177700177.

Efron:1967:PLR

[Efr67] Bradley Efron. The power of the likelihood ratio test. Annals of Mathematical Statistics, 38(3):802–806, June 1967. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177698874.

Efron:1979:RLB

[Efr79] B. Efron. The 1977 Rietz Lecture: Bootstrap methods: Another look at the jackknife. Annals of Statistics, 7(1):1–26, January 1979. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966

(electronic). URL http://www.jstor.org/stable/2958830. See

[Que49a, Que56, Tuk58a, Mil74].

Eisenberg:1959:CSP

[EG59] Edmund Eisenberg and David Gale. Consensus of subjective probabilities: The pari-mutuel method. Annals of Mathematical Statistics, 30(1):165–168, March 1959. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177706369.

EgydiodeOliveiraCarvalho:1959:DKS

[Egy59] Pedro Egydio de Oliveira Carvalho. On the distribution of the Kolmogorov–Smirnov D-statistic. Annals of Mathematical Statistics, 30(1):173–176, March 1959. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177706371.

Ehrenfeld:1955:EED

[Ehr55] Sylvain Ehrenfeld. On the efficiency of experimental designs. Annals of Mathematical Statistics, 26(2):247–255, June 1955. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177728541.

Ehrenfeld:1966:MEC

[Ehr66] Sylvain Ehrenfeld. On a minimal essentially complete class of experiments. Annals of Mathematical Statistics, 37(2):435–440, April 1966. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177699525.

Ehrenfeld:1970:SPS

[Ehr70] S. Ehrenfeld. On a scheduling problem in sequential analysis. Annals of Mathematical Statistics, 41(4):1206–1216, August 1970. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177696895.

Eicker:1963:ANC

[Eic63a] F. Eicker. Asymptotic normality and consistency of the least squares estimators for families of linear regressions. Annals of Mathematical Statistics, 34(2):447–456, June 1963. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177704156.

Eicker:1963:CLT

[Eic63b] F. Eicker. Central Limit Theorems for families of sequences of random variables. Annals of Mathematical Statistics, 34(2):439–446, June 1963. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177704155.

Eicker:1966:MCL

[Eic66] F. Eicker. A multivariate Central Limit Theorem for random linear vector forms. Annals of Mathematical Statistics, 37(6):1825–1828, December 1966. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177699175.

Eicker:1970:PSD

[Eic70] F. Eicker. On the probability that a sample distribution function lies below a line segment. Annals of Mathematical Statistics, 41(6):2075– 2092, December 1970. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177696706.

Eisenhart:1939:ICR

[Eis39a] C. Eisenhart. The interpretation of certain regression methods and their use in biological and industrial research. Annals of Mathematical Statistics, 10(2):162–186, June 1939. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177732214.

Eisenhart:1939:NPI

[Eis39b] C. Eisenhart. A note on a priori information. Annals of Mathematical Statistics, 10(4):390–393, December 1939. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177732153.

Eisenhart:1952:RWM

[Eis52] Churchill Eisenhart. Report of the Washington meeting of the Institute. Annals of Mathematical Statistics, 23(1):152–155, March 1952. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177729495.

Eisenberg:1969:REC

[Eis69] Bennett Eisenberg. The relation of the equivalence conditions for the Brownian motion to the equivalence conditions for certain stationary processes. Annals of Mathematical Statistics, 40(6):2143–2148, December 1969. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177697291.

Eisenberg:1970:TGP

[Eis70] Bennett Eisenberg. Translating Gaussian processes. Annals of Mathematical Statistics, 41(3):888–893, June 1970. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177696966.

Eisenberg:1972:NMT

[Eis72] Bennett Eisenberg. A note on metric transitivity for stationary Gaussian processes on groups. Annals of Mathematical Statistics, 43(2):

683–687, April 1972. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177692655.

Ekman:1959:AEC

[Ekm59a] Gunnar Ekman. Approximate expressions for the conditional mean and variance over small intervals of a continuous distribution. Annals of Mathematical Statistics, 30(4):1131–1134, December 1959. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177706096.

Ekman:1959:AUU

[Ekm59b] Gunnar Ekman. An approximation useful in univariate stratification. Annals of Mathematical Statistics, 30(1):219–229, March 1959. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177706377.

Elandt:1962:EAP

[Ela62] Regina C. Elandt. Exact and approximate power function of the nonparametric test of tendency. Annals of Mathematical Statistics, 33 (2):471–481, June 1962. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177704574.

Elfving:1952:OAL

[Elf52] G. Elfving. Optimum allocation in linear regression theory. Annals of Mathematical Statistics, 23(2):255–262, June 1952. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177729442.

Elias:1970:BAP

[Eli70] Peter Elias. Bounds and asymptotes for the performance of multivariate quantizers. Annals of Mathematical Statistics, 41(4):1249–1259, August 1970. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177696900.

Elias:1972:ECU

[Eli72] Peter Elias. The efficient construction of an unbiased random sequence. Annals of Mathematical Statistics, 43(3):865–870, June 1972. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177692552;   http://www.jstor.org/stable/2240383.

Ellison:1962:CPW

[Ell62] Bob E. Ellison. A classification problem in which information about alternative distributions is based on samples. Annals of Mathematical Statistics, 33(1):213–223, March 1962. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177704725.

Ellison:1964:TST

[Ell64] Bob E. Ellison. On two-sided tolerance intervals for a normal distribution. Annals of Mathematical Statistics, 35(2):762–772, June 1964. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177703575.

Ellison:1965:MNC

[Ell65] Bob E. Ellison. Multivariate-normal classification with covariances known. Annals of Mathematical Statistics, 36(6):1787–1793, December 1965. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177699807.

Elliott:1972:GSS

[Ell72] E. O. Elliott. A generalization of separable stochastic processes. Annals of Mathematical Statistics, 43(1):320–325, February 1972. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177692725.

Eaton:1970:AIU

[EM70] Morris L. Eaton and Carl N. Morris. The application of invariance to unbiased estimation. Annals of Mathematical Statistics, 41(5):1708– 1716, October 1970. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177696815.

Enlow:1934:QNC

[Enl34] E. R. Enlow. Quadrature of the normal curve. Annals of Mathematical Statistics, 5(2):136–145, June 1934. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177732712.

Enns:1970:DPM

[Enn70] E. G. Enns. The decision process for maximizing the probability of obtaining the random variable nearest to an arbitrary real number.

Annals of Mathematical Statistics, 41(5):1466–1471, October 1970. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177696792. See

correction [Enn73].

Enns:1973:CNC

[Enn73] E. G. Enns. Correction notes: Correction to “The Optimum Strategy for Maximising the Probability of Obtaining the Random Variable Nearest to an Arbitrary Real Number”. Annals of Statistics, 1(2): 400, March 1973. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.aos/

1176342382. See [Enn70].

Eaton:1972:RQB

[EO72] Morris L. Eaton and Richard A. Olshen. Random quotients and the Behrens–Fisher problem. Annals of Mathematical Statistics, 43(6): 1852–1860, December 1972. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177690856.

Esary:1972:RAS

[EP72] J. D. Esary and F. Proschan. Relationships among some concepts of bivariate dependence. Annals of Mathematical Statistics, 43(2): 651–655, April 1972. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177692646.

Epstein:1948:SAM

[Eps48] Benjamin Epstein. Some applications of the Mellin transform in statistics. Annals of Mathematical Statistics, 19(3):370–379, September 1948. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177730201.

Epstein:1949:DEV

[Eps49a] Benjamin Epstein. The distribution of extreme values in samples whose members are subject to a Markoff chain condition. Annals of Mathematical Statistics, 20(4):590–594, December 1949. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177729951. See correc-

tion [Eps51].

Epstein:1949:MEV

[Eps49b] Benjamin Epstein. A modified extreme value problem. Annals of Mathematical Statistics, 20(1):99–103, March 1949. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177730095.

Epstein:1951:CDE

[Eps51] Benjamin Epstein. Correction to “The Distribution of Extreme Values in Samples whose Members are Subject to a Markoff Chain Condition”. Annals of Mathematical Statistics, 22(1):133–134, March 1951. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177729707. See

[Eps49a].

Epstein:1954:TDN

[Eps54a] Benjamin Epstein. Tables for the distribution of the number of exceedances. Annals of Mathematical Statistics, 25(4):762–768, December 1954. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177728662.

Epstein:1954:TLT

[Eps54b] Benjamin Epstein. Truncated life tests in the exponential case. Annals of Mathematical Statistics, 25(3):555–564, September 1954. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177728723.

Esary:1967:ARV

[EPW67] J. D. Esary, F. Proschan, and D. W. Walkup. Association of random variables, with applications. Annals of Mathematical Statistics, 38 (5):1466–1474, October 1967. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177698701.

Eisenhart:1972:WJY

[ER72] Churchill Eisenhart and Joan R. Rosenblatt. W. J. Youden, 1900– 1971. Annals of Mathematical Statistics, 43(4):1035–1040, August 1972. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177692458.

Erdos:1949:THR

[Erd49] P. Erdős. On a theorem of Hsu and Robbins. Annals of Mathematical Statistics, 20(2):286–291, June 1949. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177730037. See remark [Erd50].

Erdos:1950:RMP

[Erd50] P. Erdős. Remark on my paper “On a Theorem of Hsu and Robbins”. Annals of Mathematical Statistics, 21(1):138, March 1950. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177729897. See

[Erd49].

Ericson:1966:MCS

[Eri66] Ragnar Ericson. On moments of cumulative sums. Annals of Mathematical Statistics, 37(6):1803–1805, December 1966. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177699170.

Erickson:1970:FMC

[Eri70a] R. V. Erickson. Functions of Markov chains. Annals of Mathematical Statistics, 41(3):843–850, June 1970. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177696962.

Ericson:1970:CUA

[Eri70b] W. A. Ericson. On a class of uniformly admissible estimators of a finite population total. Annals of Mathematical Statistics, 41(4): 1369–1372, August 1970. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177696916.

Erickson:1971:CCL

[Eri71a] R. V. Erickson. Constant coefficient linear differential equations driven by white noise. Annals of Mathematical Statistics, 42(2): 820–823, April 1971. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177693440.

Erickson:1971:EAM

[Eri71b] R. V. Erickson. On the existence of absolute moments for the extinction time of a Galton–Watson process. Annals of Mathematical Statistics, 42(3):1124–1128, June 1971. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177693345.

Erickson:1972:SMC

[Eri72] R. V. Erickson. Stationary measures for a class of storage models driven by a Markov chain. Annals of Mathematical Statistics, 43(3): 997–1007, June 1972. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177692563.

Epstein:1954:STR

[ES54] B. Epstein and M. Sobel. Some theorems relevant to life testing from an exponential distribution. Annals of Mathematical Statistics, 25 (2):373–381, June 1954. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177728793.

Epstein:1955:SLT

[ES55] Benjamin Epstein and Milton Sobel. Sequential life tests in the exponential case. Annals of Mathematical Statistics, 26(1):82–93, March 1955. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177728595.

Eicker:1972:MOP

[ESM72] Patrick J. Eicker, M. M. Siddiqui, and Paul W. Mielke, Jr. A matrix occupancy problem. Annals of Mathematical Statistics, 43(3):988– 996, June 1972. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177692562.

Epstein:1953:STB

[ET53] Benjamin Epstein and Chia Kuei Tsao. Some tests based on ordered observations from two exponential populations. Annals of Mathematical Statistics, 24(3):458–466, September 1953. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177728985.

Efron:1968:LDT

[ET68] Bradley Efron and Donald Truax. Large deviations theory in exponential families. Annals of Mathematical Statistics, 39(5):1402–1424, October 1968. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177698121.

Ephremides:1972:MCM

[ET72] A. Ephremides and J. B. Thomas. On the multiplicity of a class of multivariate random processes. Annals of Mathematical Statistics, 43 (6):2083–2089, December 1972. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177690889.

Evans:1940:NMB

[Eva40] W. D. Evans. Note on the moments of a binomially distributed variate. Annals of Mathematical Statistics, 11(1):106–107, March 1940. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177731949.

Evans:1963:MS

[Eva63] David H. Evans. Multiplex sampling. Annals of Mathematical Statistics, 34(4):1322–1346, December 1963. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177703868.

Ehrenfeld:1961:RFE

[EZ61] S. Ehrenfeld and S. Zacks. Randomization and factorial experiments. Annals of Mathematical Statistics, 32(1):270–297, March 1961. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177705159.

Ehrenfeld:1963:OSF

[EZ63] S. Ehrenfeld and S. Zacks. Optimal strategies in factorial experiments. Annals of Mathematical Statistics, 34(3):780–791, September 1963. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177704003.

Ehrenfeld:1967:THR

[EZ67] S. Ehrenfeld and S. Zacks. Testing hypotheses in randomized factorial experiments. Annals of Mathematical Statistics, 38(5):1494– 1507, October 1967. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177698704.

Ezekiel:1930:SVL

[Eze30] Mordecai Ezekiel. The sampling variability of linear and curvilinear regressions: A first approximation to the reliability of the results secured by the graphic “Successive Approximation” method. Annals of Mathematical Statistics, 1(4):275–315, 317–333, November 1930. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177733062.

Fabian:1962:MDM

[Fab62] Václav Fabian. On multiple decision methods for ranking population means. Annals of Mathematical Statistics, 33(1):248–254, March 1962. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177704728.

Fabius:1964:ABB

[Fab64] J. Fabius. Asymptotic behavior of Bayes’ estimates. Annals of Mathematical Statistics, 35(2):846–856, June 1964. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177703584.

Fabian:1967:SAM

[Fab67] Václav Fabian. Stochastic approximation of minima with improved asymptotic speed. Annals of Mathematical Statistics, 38(1):191–200, February 1967. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177699070.

Fabian:1968:ANS

[Fab68a] Václav Fabian. On asymptotic normality in stochastic approximation. Annals of Mathematical Statistics, 39(4):1327–1332, August 1968. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177698258.

Fabian:1968:CDS

[Fab68b] Václav Fabian. On the choice of design in stochastic approximation methods. Annals of Mathematical Statistics, 39(2):457–465, April 1968. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177698409.

Fabian:1969:RAA

[Fab69a] Vaclav Fabian. Review: Arthur E. Albert, Leland A. Gardner, Stochastic Approximation and Nonlinear Regression. Annals of Mathematical Statistics, 40(3):1134–1135, June 1969. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177697628.

Fabian:1969:SAS

[Fab69b] Václav Fabian. Stochastic approximation for smooth functions. Annals of Mathematical Statistics, 40(1):299–302, February 1969. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177697825.

Fabian:1972:NKT

[Fab72] Václav Fabian. Note on a “Kolmogorov-type” inequality. Annals of Mathematical Statistics, 43(2):663, April 1972. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177692649.

Farrell:1962:BLC

[Far62] R. H. Farrell. Bounded length confidence intervals for the zero of a regression function. Annals of Mathematical Statistics, 33(1):237– 247, March 1962. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177704727.

Farrell:1964:ABE

[Far64a] R. H. Farrell. Asymptotic behavior of expected sample size in certain one sided tests. Annals of Mathematical Statistics, 35(1):36– 72, March 1964. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177703731.

Farrell:1964:ELP

[Far64b] R. H. Farrell. Estimators of a location parameter in the absolutely continuous case. Annals of Mathematical Statistics, 35(3):949–998, September 1964. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177700516.

Farrell:1964:LTS

[Far64c] R. H. Farrell. Limit theorems for stopped random walks. Annals of Mathematical Statistics, 35(3):1332–1343, September 1964. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177703290.

Farrell:1966:BLCa

[Far66a] R. H. Farrell. Bounded length confidence intervals for the p-point of a distribution function, II. Annals of Mathematical Statistics, 37 (3):581–585, June 1966. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177699451.

Farrell:1966:BLCb

[Far66b] R. H. Farrell. Bounded length confidence intervals for the p-point of a distribution function, III. Annals of Mathematical Statistics, 37 (3):586–592, June 1966. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177699452.

Farrell:1966:LTSa

[Far66c] R. H. Farrell. Limit theorems for stopped random walks. II. Annals of Mathematical Statistics, 37(4):860–865, August 1966. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177699367.

Farrell:1966:LTSb

[Far66d] R. H. Farrell. Limit theorems for stopped random walks. III. Annals of Mathematical Statistics, 37(6):1510–1527, December 1966. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177699143.

Farrell:1967:LUC

[Far67] R. H. Farrell. On the lack of a uniformly consistent sequence of estimators of a density function in certain cases. Annals of Mathematical Statistics, 38(2):471–474, April 1967. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177698962.

Farrell:1968:NSC

[Far68a] R. H. Farrell. On a necessary and sufficient condition for admissibility of estimators when strictly convex loss is used. Annals of Mathematical Statistics, 39(1):23–28, February 1968. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177698502.

Farrell:1968:AWC

[Far68b] R. H. Farrell. On the admissibility at ∞, within the class of randomized designs, of balanced designs. Annals of Mathematical Statistics, 39(6):1978–1994, December 1968. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177698026.

Farrell:1968:TTG

[Far68c] R. H. Farrell. Towards a theory of generalized Bayes tests. Annals of Mathematical Statistics, 39(1):1–22, February 1968. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177698501.

Farrell:1969:ARS

[Far69a] R. H. Farrell. On the admissibility of a randomized symmetrical design for the problem of a one way classification. Annals of Mathematical Statistics, 40(2):356–365, April 1969. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177697699.

Farrell:1969:BCS

[Far69b] R. H. Farrell. On the Bayes character of a standard Model II analysis of variance test. Annals of Mathematical Statistics, 40(3):1094– 1097, June 1969. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177697616.

Farrell:1972:BOA

[Far72] R. H. Farrell. On the best obtainable asymptotic rates of convergence in estimation of a density function at a point. Annals of Mathematical Statistics, 43(1):170–180, February 1972. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177692711.

Fuchs:1965:PLM

[FD65] Carol E. Fuchs and H. T. David. Poisson limits of multivariate run distributions. Annals of Mathematical Statistics, 36(1):215–225, February 1965. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177700284.

Federer:1951:TPC

[Fed51] Walter T. Federer. Testing proportionality of covariance matrices. Annals of Mathematical Statistics, 22(1):102–106, March 1951. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177729697.

Feder:1968:DLL

[Fed68] Paul I. Feder. On the distribution of the log likelihood ratio test statistic when the true parameter is “near” the boundaries of the hypothesis regions. Annals of Mathematical Statistics, 39(6):2044– 2055, December 1968. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177698032.

Feldman:1932:DPC

[Fel32] H. M. Feldman. The distributions of the precision constant and its square in samples of η from a normal population. Annals of Mathematical Statistics, 3(1):20–31, February 1932. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177732929.

Feldman:1935:MEP

[Fel35] Hyman M. Feldman. Mathematical expectation of product moments of samples drawn from a set of infinite populations. Annals of Mathematical Statistics, 6(1):30–52, March 1935. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177732640.

Feller:1941:IER

[Fel41] Willy Feller. On the integral equation of renewal theory. Annals of Mathematical Statistics, 12(3):243–267, September 1941. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177731708.

Feller:1943:GCC

[Fel43] W. Feller. On a general class of “contagious” distributions. Annals of Mathematical Statistics, 14(4):389–400, December 1943. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177731359.

Feller:1945:NLL

[Fel45a] W. Feller. Note on the law of large numbers and “fair” games. Annals of Mathematical Statistics, 16(3):301–304, September 1945. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177731094.

Feller:1945:NAB

[Fel45b] W. Feller. On the normal approximation to the binomial distribution. Annals of Mathematical Statistics, 16(4):319–329, December 1945.

CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177731058.

Feller:1947:RHC

[Fel47] Will Feller. Review: Harald Cramér, Mathematical Methods of Statistics. Annals of Mathematical Statistics, 18(1):136–139, March 1947. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177730503.

Feller:1948:KSL

[Fel48a] W. Feller. On the Kolmogorov–Smirnov limit theorems for empirical distributions. Annals of Mathematical Statistics, 19(2):177– 189, June 1948. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177730243;   http://www.jstor.org/stable/2236265. See er-

rata [Fel50].

Feller:1948:RPI

[Fel48b] Will Feller. Report of the President of the Institute for 1947. Annals of Mathematical Statistics, 19(1):137–140, March 1948. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177730307.

Feller:1950:EKS

[Fel50] W. Feller. Errata: “On the Kolmogorov–Smirnov Limit Theorems for Empirical Distributions”. Annals of Mathematical Statistics, 21 (2):301–302, June 1950. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177729850. See [Fel48a].

Feller:1951:ADR

[Fel51] William Feller. The asymptotic distribution of the range of sums of independent random variables. Annals of Mathematical Statistics, 22 (3):427–432, September 1951. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177729589.

Feller:1959:NMP

[Fel59] William Feller. Non-Markovian processes with the semigroup property. Annals of Mathematical Statistics, 30(4):1252–1253, December 1959. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177706110.

Feldman:1962:CTA

[Fel62] Dorian Feldman. Contributions to the “Two-Armed Bandit” problem. Annals of Mathematical Statistics, 33(3):847–856, September 1962. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177704454.

Feldman:1965:IKI

[Fel65] J. Feldman. Integral kernels and invariant measures for Markoff transition functions. Annals of Mathematical Statistics, 36(2):517– 523, April 1965. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177700161.

Feldman:1968:CCC

[Fel68] J. Feldman. A clarification concerning certain equivalence classes of Gaussian processes on an interval. Annals of Mathematical Statistics, 39(3):1078–1079, June 1968. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177698340.

Feldman:1972:SPB

[Fel72] D. Feldman. Some properties of Bayesian orderings of experiments. Annals of Mathematical Statistics, 43(5):1428–1440, October 1972. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177692375.

Fend:1959:ACR

[Fen59] A. V. Fend. On the attainment of Cramér–Rao and Bhattacharyya bounds for the variance of an estimate. Annals of Mathematical Statistics, 30(2):381–388, June 1959. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177706258.

Fertig:1936:MTH

[Fer36] John W. Fertig. On a method of testing the hypothesis that an observed sample of n variables and of size N has been drawn from a specified population of the same number of variables. Annals of Mathematical Statistics, 7(3):113–121, September 1936. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177732501.

Ferguson:1958:MGB

[Fer58] Thomas S. Ferguson. A method of generating best asymptotically normal estimates with application to the estimation of bacterial densities. Annals of Mathematical Statistics, 29(4):1046–1062, December 1958. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177706440.

Ferguson:1962:LSP

[Fer62a] Thomas S. Ferguson. Location and scale parameters in exponential families of distributions. Annals of Mathematical Statistics, 33 (3):986–1001, September 1962. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177704466. See correction [Fer63].

Ferguson:1962:RSB

[Fer62b] Thomas S. Ferguson. A representation of the symmetric bivariate Cauchy distribution. Annals of Mathematical Statistics, 33(4):1256– 1266, December 1962. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177704357.

Ferguson:1963:CNC

[Fer63] Thomas S. Ferguson. Correction notes: Correction to “Location and Scale Parameters in Exponential Families of Distributions”. Annals of Mathematical Statistics, 34(4):1603, December 1963. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177703896. See [Fer62a].

Ferguson:1964:CED

[Fer64] Thomas S. Ferguson. A characterization of the exponential distribution. Annals of Mathematical Statistics, 35(3):1199–1207, September 1964. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177703277.

Ferguson:1970:CIG

[Fer70a] Thomas S. Ferguson. On a class of infinite games related to liar dice. Annals of Mathematical Statistics, 41(2):353–362, April 1970. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177697075.

Fernandez:1970:WCT

[Fer70b] Pedro J. Fernandez. A weak convergence theorem for random sums of independent random variables. Annals of Mathematical Statistics, 41 (2):710–712, April 1970. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177697120.

Fernandez:1971:WCT

[Fer71] Pedro J. Fernandez. A weak convergence theorem for random sums in a normed space. Annals of Mathematical Statistics, 42(5):1737– 1741, October 1971. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177693173.

Ferreira:1972:SPE

[Fer72a] Pedro E. Ferreira. On subjective probability and expected utilities. Annals of Mathematical Statistics, 43(3):928–933, June 1972. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177692556.

Fertig:1972:BPD

[Fer72b] K. W. Fertig. Bayesian prior distributions for systems with exponential failure-time data. Annals of Mathematical Statistics, 43(5):1441– 1448, October 1972. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177692376.

Fraser:1952:BBR

[FG52] D. A. S. Fraser and Irwin Guttman. Bhattacharyya bounds without regularity assumptions. Annals of Mathematical Statistics, 23(4):629– 632, December 1952. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177729344.

Fraser:1956:TR

[FG56] D. A. S. Fraser and Irwin Guttman. Tolerance regions. Annals of Mathematical Statistics, 27(1):162–179, March 1956. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177728355.

Freimer:1960:NDL

[FG60] Marshall Freimer and Bernard Gold. Note on the distribution of locally maximal elements in a random sample. Annals of Mathematical Statistics, 31(2):518, June 1960. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177705919.

Ferris:1946:OCC

[FGW46] Charles D. Ferris, Frank E. Grubbs, and Chalmers L. Weaver. Operating characteristics for the common statistical tests of significance. Annals of Mathematical Statistics, 17(2):178–197, June 1946. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177730979.

Fix:1955:SPW

[FH55] Evelyn Fix and J. L. Hodges, Jr. Significance probabilities of the Wilcoxon test. Annals of Mathematical Statistics, 26(2):301–312, June 1955. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177728547.

Hou:1969:WAT

[fH69] Tien fang Hou. Weak approachability in a two-person game. Annals of Mathematical Statistics, 40(3):789–813, June 1969. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177697588.

Hou:1971:ATP

[fH71] Tien fang Hou. Approachability in a two-person game. Annals of Mathematical Statistics, 42(2):735–744, April 1971. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177693422.

Fienberg:1968:GCT

[Fie68] Stephen E. Fienberg. The geometry of an r×c contingency table. Annals of Mathematical Statistics, 39(4):1186–1190, August 1968. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177698242.

Fienberg:1970:IPE

[Fie70] Stephen E. Fienberg. An iterative procedure for estimation in contingency tables. Annals of Mathematical Statistics, 41(3):907–917, June 1970. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/

1177696968. See correction [Fie71].

Fienberg:1971:CNC

[Fie71] Stephen E. Fienberg. Correction notes: Correction to “An Iterative Procedure for Estimation in Contingency Tables”. Annals of Mathematical Statistics, 42(5):1778, October 1971. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177693185.

Finney:1946:FDD

[Fin46] D. J. Finney. The frequency distribution of deviates from means and regression lines in samples from a multivariate normal population. Annals of Mathematical Statistics, 17(3):344–349, September 1946. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177730947.

Finch:1961:TBC

[Fin61] P. D. Finch. The transient behaviour of a coincidence variate in telephone traffic. Annals of Mathematical Statistics, 32(1):230–234, March 1961. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177705153.

Finch:1962:TBQ

[Fin62] P. D. Finch. On the transient behavior of a queueing system with bulk service and finite capacity. Annals of Mathematical Statistics, 33(3):973–985, September 1962. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177704465.

Fine:1966:HLS

[Fin66] Terrence Fine. On the Hodges and Lehmann shift estimator in the two sample problem. Annals of Mathematical Statistics, 37(6):1814– 1818, December 1966. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177699173.

Fine:1971:NEQ

[Fin71a] Terrence Fine. A note on the existence of quantitative probability. Annals of Mathematical Statistics, 42(4):1182–1186, August 1971. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177693234.

Finkelstein:1971:LIL

[Fin71b] Helen Finkelstein. The law of the iterated logarithm for empirical distributions. Annals of Mathematical Statistics, 42(2):607–615, April 1971. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177693410.

Fischer:1933:CSS

[Fis33a] Carl H. Fischer. On correlation surfaces of sums with a certain number of random elements in common. Annals of Mathematical Statistics, 4(2):103–126, May 1933. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177732804.

Fischer:1933:MPC

[Fis33b] Carl H. Fischer. On multiple and partial correlation coefficients of a certain sequence of sums. Annals of Mathematical Statistics, 4 (4):278–284, November 1933. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177732756.

Fisher:1939:NFI

[Fis39] R. A. Fisher. A note on fiducial inference. Annals of Mathematical Statistics, 10(4):383–388, December 1939. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177732151.

Fischer:1942:SDV

[Fis42] Carl H. Fischer. A sequence of discrete variables exhibiting correlation due to common elements. Annals of Mathematical Statistics, 13 (1):97–101, March 1942. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177731649.

Fischer:1950:MAM

[Fis50a] Carl H. Fischer. Minutes of the annual membership meeting, New York, December 28, 1949. Annals of Mathematical Statistics, 21(1): 155, March 1950. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177729901.

Fischer:1950:RST

[Fis50b] Carl H. Fischer. Report of the Secretary-Treasurer of the Institute for 1949. Annals of Mathematical Statistics, 21(1):160–163, March 1950. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177729903.

Fischer:1951:RST

[Fis51] Carl H. Fischer. Report of the Secretary-Treasurer of the Institute for 1950. Annals of Mathematical Statistics, 22(1):161–163, March 1951. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177729713.

Fischer:1952:MAM

[Fis52a] C. H. Fischer. Minutes of the annual membership meeting, Boston, December 28, 1951. Annals of Mathematical Statistics, 23(1):159– 160, March 1952. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177729497.

Fischer:1952:RST

[Fis52b] Carl H. Fischer. Report of the Secretary-Treasurer of the Institute for 1951. Annals of Mathematical Statistics, 23(1):164–166, March 1952. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177729499.

Fisz:1960:RMR

[Fis60] M. Fisz. On a result by M. Rosenblatt concerning the von Mises– Smirnov test. Annals of Mathematical Statistics, 31(2):427–429, June 1960. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177705905.

Fisz:1962:IDD

[Fis62] Marek Fisz. Infinitely divisible distributions: Recent results and applications. Annals of Mathematical Statistics, 33(1):68–84, March 1962. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177704712.

Fishburn:1967:BEU

[Fis67a] Peter C. Fishburn. Bounded expected utility. Annals of Mathematical Statistics, 38(4):1054–1060, August 1967. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177698774.

Fishburn:1967:PBD

[Fis67b] Peter C. Fishburn. Preference-based definitions of subjective probability. Annals of Mathematical Statistics, 38(6):1605–1617, December 1967. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177698594.

Fisher:1968:RDD

[Fis68] Lloyd Fisher. On recurrent denumerable decision processes. Annals of Mathematical Statistics, 39(2):424–432, April 1968. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177698406.

Fishburn:1969:GTS

[Fis69a] Peter C. Fishburn. A general theory of subjective probabilities and expected utilities. Annals of Mathematical Statistics, 40(4):1419– 1429, August 1969. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177697513.

Fishburn:1969:WQP

[Fis69b] Peter C. Fishburn. Weak qualitative probability on finite sets. Annals of Mathematical Statistics, 40(6):2118–2126, December 1969. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177697289.

Fisher:1969:LSC

[Fis69c] Lloyd Fisher. Limiting sets and convex hulls of samples from product measures. Annals of Mathematical Statistics, 40(5):1824–1832, October 1969. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177697395.

Fisk:1970:NCM

[Fis70] P. R. Fisk. A note on a characterization of the multivariate normal distribution. Annals of Mathematical Statistics, 41(2):486–494, April 1970. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177697088.

Fishburn:1971:OWE

[Fis71] Peter C. Fishburn. One-way expected utility with finite consequence spaces. Annals of Mathematical Statistics, 42(2):572–577, April 1971. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177693407.

Fishburn:1972:AAO

[Fis72a] Peter C. Fishburn. Alternative axiomatizations of one-way expected utility. Annals of Mathematical Statistics, 43(5):1648–1651, October 1972. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177692397.

Fishburn:1972:SEU

[Fis72b] Peter C. Fishburn. Subjective expected utility with mixture sets and Boolean algebras. Annals of Mathematical Statistics, 43(3):917– 927, June 1972. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177692555.

Frank:1951:ASB

[FK51] P. Frank and J. Kiefer. Almost subminimax and biased minimax procedures. Annals of Mathematical Statistics, 22(3):465–468, September 1951. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177729596.

Fleischer:1958:STS

[FK58] Isidore Fleischer and Anthony Kooharian. On the statistical treatment of stochastic processes. Annals of Mathematical Statistics, 29 (2):544–549, June 1958. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177706629.

Furstenberg:1960:PRM

[FK60] H. Furstenberg and H. Kesten. Products of random matrices. Annals of Mathematical Statistics, 31(2):457–469, June 1960. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177705909.

Ferguson:1972:RII

[FK72] Thomas S. Ferguson and Michael J. Klass. A representation of independent increment processes without Gaussian components. Annals of Mathematical Statistics, 43(5):1634–1643, October 1972. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177692395.

Flehinger:1962:GMR

[Fle62a] B. J. Flehinger. A general model for the reliability analysis of systems under various preventive maintenance policies. Annals of Mathematical Statistics, 33(1):137–156, March 1962. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177704719.

Flehinger:1962:MMA

[Fle62b] Betty J. Flehinger. A Markovian model for the analysis of the effects of marginal testing on system reliability. Annals of Mathematical Statistics, 33(2):754–766, June 1962. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177704595.

Feldt:1958:PFC

[FM58] Leonard S. Feldt and Moharram W. Mahmoud. Power function charts for specifying numbers of observations in analyses of variance of fixed effects. Annals of Mathematical Statistics, 29(3):871–877, September 1958. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177706543.

Foster:1961:QBD

[FN61] F. G. Foster and K. M. Nyunt. Queues with batch departures. I. Annals of Mathematical Statistics, 32(4):1324–1332, December 1961. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177704870.

Forsyth:1937:SSD

[For37] C. H. Forsyth. Some simple developments in the use of the coefficient of stability. Annals of Mathematical Statistics, 8(1):5–11, March 1937. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177732449.

Foster:1952:MCD

[Fos52] F. G. Foster. A Markov chain derivation of discrete distributions. Annals of Mathematical Statistics, 23(4):624–627, December 1952. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177729342.

Foster:1953:SMA

[Fos53] F. G. Foster. On the stochastic matrices associated with certain queuing processes. Annals of Mathematical Statistics, 24(3):355–360, September 1953. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177728976.

Foster:1971:LTB

[Fos71] James H. Foster. A limit theorem for a branching process with state-dependent immigration. Annals of Mathematical Statistics, 42 (5):1773–1776, October 1971. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177693182.

Fox:1956:CPT

[Fox56] Martin Fox. Charts of the power of the F -test. Annals of Mathematical Statistics, 27(2):484–497, June 1956. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177728271.

Fox:1970:EED

[Fox70] Richard J. Fox. Estimating the empiric distribution function of certain parameter sequences. Annals of Mathematical Statistics, 41(6): 1845–1852, December 1970. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177696685.

Fox:1971:ACL

[Fox71] Martin Fox. Admissibility of certain location invariant multiple decision procedures. Annals of Mathematical Statistics, 42(5):1553–1561, October 1971. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177693153.

Fertig:1937:TSV

[FP37] John W. Fertig and Elizabeth A. Proehl. A test of a sample variance based on both tail ends of the distribution. Annals of Mathematical Statistics, 8(4):193–205, December 1937. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177732389.

Freund:1956:SRR

[FP56] John E. Freund and Arthur N. Pozner. Some results on restricted occupancy theory. Annals of Mathematical Statistics, 27(2):537– 540, June 1956. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177728278.

Foster:1964:QBD

[FP64] F. G. Foster and A. G. A. D. Perera. Queues with batch departures. II. Annals of Mathematical Statistics, 35(3):1147–1156, September 1964. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177703272.

Freedman:1967:TPOb

[FP67] David A. Freedman and Roger Purves. Timid play is optimal, II. Annals of Mathematical Statistics, 38(4):1284–1285, August 1967. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177698800.

Fox:1969:IBI

[FP69a] Martin Fox and S. K. Perng. Inadmissibility of the best invariant test when the moment is infinite under one of the hypotheses. Annals of Mathematical Statistics, 40(4):1483–1485, August 1969. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177697520.

Freedman:1969:BMB

[FP69b] David A. Freedman and Roger A. Purves. Bayes’ method for bookies. Annals of Mathematical Statistics, 40(4):1177–1186, August 1969. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177697494.

Fraser:1971:RMD

[FP71] D. A. S. Fraser and R. L. Prentice. Randomized models and the dilution and bioassay problems. Annals of Mathematical Statistics, 42(1):141–146, February 1971. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177693501.

Fox:1964:AQE

[FR64] Martin Fox and Herman Rubin. Admissibility of quantile estimates of a single location parameter. Annals of Mathematical Statistics, 35 (3):1019–1030, September 1964. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177700518.

Fisher:1968:EDD

[FR68a] Lloyd Fisher and Sheldon M. Ross. An example in denumerable decision processes. Annals of Mathematical Statistics, 39(2):674– 675, April 1968. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177698426.

Fox:1968:FPM

[FR68b] Martin Fox and Herman Rubin. Functions of processes with Markovian states. Annals of Mathematical Statistics, 39(3):938– 946, June 1968. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177698326.

Fox:1969:FPM

[FR69] Martin Fox and Herman Rubin. Functions of processes with Markovian states. II. Annals of Mathematical Statistics, 40(3):865–869, June 1969. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177697594.

Fox:1970:FPM

[FR70] Martin Fox and Herman Rubin. Functions of processes with Markovian states. III. Annals of Mathematical Statistics, 41(2):472–479, April 1970. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177697086.

Fraser:1951:GHP

[Fra51a] D. A. S. Fraser. Generalized hit probabilities with a Gaussian target. Annals of Mathematical Statistics, 22(2):248–255, June 1951. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177729644.

Fraser:1951:SDS

[Fra51b] D. A. S. Fraser. Sequentially determined statistically equivalent blocks. Annals of Mathematical Statistics, 22(3):372–381, September 1951. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177729583.

Fraser:1952:CBS

[Fra52a] D. A. S. Fraser. Confidence bounds for a set of means. Annals of Mathematical Statistics, 23(4):575–585, December 1952. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177729336.

Fraser:1952:SSS

[Fra52b] D. A. S. Fraser. Sufficient statistics and selection depending on the parameter. Annals of Mathematical Statistics, 23(3):417–425, September 1952. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177729386.

Fraser:1953:GHP

[Fra53a] D. A. S. Fraser. Generalized hit probabilities with a Gaussian target, II. Annals of Mathematical Statistics, 24(2):288–294, June 1953. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177729035.

Fraser:1953:NTR

[Fra53b] D. A. S. Fraser. Nonparametric tolerance regions. Annals of Mathematical Statistics, 24(1):44–55, March 1953. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177729081.

Frasher:1953:BFP

[Fra53c] D. A. S. Frasher. The Behrens–Fisher problem for regression coefficients. Annals of Mathematical Statistics, 24(3):390–402, September 1953. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177728979.

Fraser:1956:SSN

[Fra56a] D. A. S. Fraser. Sufficient statistics with nuisance parameters. Annals of Mathematical Statistics, 27(3):838–842, September 1956. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177728191.

Fraser:1956:VFW

[Fra56b] D. A. S. Fraser. A vector form of the Wald–Wolfowitz–Hoeffding theorem. Annals of Mathematical Statistics, 27(2):540–543, June 1956. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177728279.

Fraser:1957:MPR

[Fra57a] D. A. S. Fraser. Most powerful rank-type tests. Annals of Mathematical Statistics, 28(4):1040–1043, December 1957. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177706809.

Fraser:1957:CII

[Fra57b] D. A. S. Fraser. On the combining of interblock and intrablock estimates. Annals of Mathematical Statistics, 28(3):814–816, September 1957. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177706901.

Fraser:1957:RAU

[Fra57c] D. A. S. Fraser. A regression analysis using the invariance method. Annals of Mathematical Statistics, 28(2):517–520, June 1957. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177706984.

Fraser:1961:FI

[Fra61] D. A. S. Fraser. On fiducial inference. Annals of Mathematical Statistics, 32(3):661–676, September 1961. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177704962.

Franklin:1963:CMC

[Fra63] Joel N. Franklin. The covariance matrix of a continuous autoregressive vector time-series. Annals of Mathematical Statistics, 34(4): 1259–1264, December 1963. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177703861.

Fraser:1965:IS

[Fra65] D. A. S. Fraser. On information in statistics. Annals of Mathematical Statistics, 36(3):890–896, June 1965. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177700061.

Fraser:1967:DTL

[Fra67a] D. A. S. Fraser. Data transformations and the linear model. Annals of Mathematical Statistics, 38(5):1456–1465, October 1967. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177698700.

Fraser:1967:SMI

[Fra67b] D. A. S. Fraser. Statistical models and invariance. Annals of Mathematical Statistics, 38(4):1061–1067, August 1967. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177698775.

Frank:1968:AEI

[Fra68a] Peter Frank. The asymptotic error of iterations. Annals of Mathematical Statistics, 39(1):266–269, February 1968. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177698530.

Fraser:1968:CWN

[Fra68b] D. A. S. Fraser. The conditional Wishart: Normal and nonnormal. Annals of Mathematical Statistics, 39(2):593–605, April 1968. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177698420.

Fraser:1972:BLS

[Fra72a] D. A. S. Fraser. Bayes, likelihood, or structural. Annals of Mathematical Statistics, 43(3):777–790, June 1972. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177692544.

Fraser:1972:DLT

[Fra72b] D. A. S. Fraser. The determination of likelihood and the transformed regression model. Annals of Mathematical Statistics, 43(3): 898–916, June 1972. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177692554.

Frechet:1947:DPD

[Fré47a] M. Fréchet. Definition of the probable deviation. Annals of Mathematical Statistics, 18(2):288–290, June 1947. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177730447.

Frechet:1947:GRB

[Fré47b] M. Fréchet. The general relation between the mean and the mode for a discontinuous variate. Annals of Mathematical Statistics, 18(2): 290–293, June 1947. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177730448.

Freeman:1957:SFM

[Fre57] G. H. Freeman. Some further methods of constructing regular group divisible incomplete block designs. Annals of Mathematical Statistics, 28(2):479–487, June 1957. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177706976.

Freeman:1958:FDT

[Fre58] G. H. Freeman. Families of designs for two successive experiments. Annals of Mathematical Statistics, 29(4):1063–1078, December 1958. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177706441.

Freeman:1961:SFD

[Fre61] G. H. Freeman. Some further designs of Type O:PP. Annals of Mathematical Statistics, 32(4):1186–1190, December 1961. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177704858.

Freedman:1962:IUM

[Fre62a] David A. Freedman. Invariants under mixing which generalize de Finetti’s theorem. Annals of Mathematical Statistics, 33(3):916–923, September 1962. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177704460.

Freedman:1962:MMP

[Fre62b] David A. Freedman. Mixtures of Markov processes. Annals of Mathematical Statistics, 33(1):114–118, March 1962. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177704716.

Freedman:1962:PPR

[Fre62c] David A. Freedman. Poisson processes with random arrival rate. Annals of Mathematical Statistics, 33(3):924–929, September 1962. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177704461.

Freedman:1963:IUM

[Fre63a] David A. Freedman. Invariants under mixing which generalize de Finetti’s theorem: Continuous time parameter. Annals of Mathematical Statistics, 34(4):1194–1216, December 1963. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177703856.

Freedman:1963:ABB

[Fre63b] David A. Freedman. On the asymptotic behavior of Bayes’ estimates in the discrete case. Annals of Mathematical Statistics, 34(4):1386– 1403, December 1963. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177703871.

Freedman:1965:BFU

[Fre65a] David A. Freedman. Bernard Friedman’s urn. Annals of Mathematical Statistics, 36(3):956–970, June 1965. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177700068.

Freedman:1965:ABB

[Fre65b] David A. Freedman. On the asymptotic behavior of Bayes estimates in the discrete case II. Annals of Mathematical Statistics, 36(2):454– 456, April 1965. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177700155.

Freedman:1966:ESF

[Fre66a] David A. Freedman. Equivalence and singularity for Friedman urns. Annals of Mathematical Statistics, 37(1):268–270, February 1966. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177699616.

Freedman:1966:NMS

[Fre66b] David A. Freedman. A note on mutual singularity of priors. Annals of Mathematical Statistics, 37(2):375–381, April 1966. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177699518.

Freedman:1966:TER

[Fre66c] David A. Freedman. On two equivalence relations between measures. Annals of Mathematical Statistics, 37(3):686–689, June 1966. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177699462.

Freeman:1966:SNO

[Fre66d] G. H. Freeman. Some non-orthogonal partitions of 4 × 4, 5 × 5 and 6 × 6 latin squares. Annals of Mathematical Statistics, 37(3):666– 681, June 1966. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177699460.

Freedman:1967:OS

[Fre67a] David A. Freedman. An oscillating semigroup. Annals of Mathematical Statistics, 38(3):924–926, June 1967. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177698888.

Freedman:1967:RSD

[Fre67b] David A. Freedman. A remark on sequential discrimination. Annals of Mathematical Statistics, 38(6):1666–1670, December 1967. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177698601. See correc-

tion [Fre68].

Freedman:1967:RLI

[Fre67c] David A. Freedman. A remark on the law of the iterated logarithm. Annals of Mathematical Statistics, 38(2):598–600, April 1967. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177698977.

Freedman:1967:SIP

[Fre67d] David A. Freedman. Some invariance principles for functionals of a Markov chain. Annals of Mathematical Statistics, 38(1):1–7, February 1967. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177699053.

Freedman:1967:TLP

[Fre67e] David A. Freedman. A theorem of Lévy and a peculiar semigroup. Annals of Mathematical Statistics, 38(5):1552–1557, October 1967. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177698709.

Freedman:1967:TPOa

[Fre67f] David A. Freedman. Timid play is optimal. Annals of Mathematical Statistics, 38(4):1281–1283, August 1967. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177698799.

Freedman:1968:CNC

[Fre68] David A. Freedman. Correction notes: Correction to “A Remark on Sequential Discrimination”. Annals of Mathematical Statistics, 39(6): 2161, December 1968. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177698051. See [Fre67b].

Friedman:1940:CAT

[Fri40] Milton Friedman. A comparison of alternative tests of significance for the problem of m rankings. Annals of Mathematical Statistics, 11 (1):86–92, March 1940. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177731944.

Friedman:1963:SA

[Fri63] Shmuel Friedman. On stochastic approximations. Annals of Mathematical Statistics, 34(1):343–346, March 1963. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177704277.

Fristedt:1969:SPK

[Fri69] Bert Fristedt. A short proof of a known limit theorem for sum of independent random variables with infinite expectations. Annals of Mathematical Statistics, 40(3):1114–1115, June 1969. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177697621.

Fristedt:1972:EDA

[Fri72] Bert Fristedt. Expansions for the density of the absolute value of a strictly stable vector. Annals of Mathematical Statistics, 43(2): 669–672, April 1972. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177692651.

Feldman:1971:BFI

[FS71] J. Feldman and M. Smorodinsky. Bernoulli flows with infinite entropy. Annals of Mathematical Statistics, 42(1):381–382, February 1971. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177693526.

Freeman:1950:TRA

[FT50] Murray F. Freeman and John W. Tukey. Transformations related to the angular and the square root. Annals of Mathematical Statistics, 21(4):607–611, December 1950. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/ euclid.aoms/1177729756. Reprinted in [Mal90, pp. 149–155].

Feldman:1966:ENU

[FT66] Dorian Feldman and Howard G. Tucker. Estimation of nonunique quantiles. Annals of Mathematical Statistics, 37(2):451–457, April 1966. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177699527.

Furukawa:1972:MDP

[Fur72] Nagata Furukawa. Markovian decision processes with compact action spaces. Annals of Mathematical Statistics, 43(5):1612–1622, October 1972. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177692393.

Fisher:1969:DPM

[FV69] Lloyd Fisher and John W. Van Ness. Distinguishability of probability measures. Annals of Mathematical Statistics, 40(2):381–392, April 1969. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177697702.

Fabius:1970:SRT

[FV70] J. Fabius and W. R. Van Zwet. Some remarks on the two-armed bandit. Annals of Mathematical Statistics, 41(6):1906–1916, December 1970. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177696692.

Fisher:1971:TD

[FV71] Lloyd Fisher and John W. Van Ness. The topology of distinguishability. Annals of Mathematical Statistics, 42(4):1262–1272, August 1971. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177693239.

Fraser:1951:NEI

[FW51] D. A. S. Fraser and R. Wormleighton. Nonparametric estimation. IV. Annals of Mathematical Statistics, 22(2):294–298, June 1951. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177729650.

Fisher:1969:EDB

[FW69] Lloyd Fisher and D. W. Walkup. An example of the difference between the Levy and Levy–Prokhorov metrics. Annals of Mathematical Statistics, 40(1):322–324, February 1969. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177697831.

Gabriel:1962:ADA

[Gab62] K. R. Gabriel. Ante-dependence analysis of an ordered set of variables. Annals of Mathematical Statistics, 33(1):201–212, March 1962. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177704724.

Gabriel:1969:STP

[Gab69] K. R. Gabriel. Simultaneous test procedures–some theory of multiple comparisons. Annals of Mathematical Statistics, 40(1):224–250, February 1969. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177697819.

Gaffey:1959:CEC

[Gaf59] William R. Gaffey. A consistent estimator of a component of a convolution. Annals of Mathematical Statistics, 30(1):198–205, March 1959. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177706375.

Galambos:1972:DMR

[Gal72] János Galambos. On the distribution of the maximum of random variables. Annals of Mathematical Statistics, 43(2):516–521, April 1972. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177692632.

Gupta:1964:MPF

[GAM64] S. Das Gupta, T. W. Anderson, and G. S. Mudholkar. Monotonicity of the power functions of some tests of the multivariate linear hypothesis. Annals of Mathematical Statistics, 35(1):200–205, March 1964. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177703742.

Gani:1961:FET

[Gan61] J. Gani. First emptiness of two dams in parallel. Annals of Mathematical Statistics, 32(1):219–229, March 1961. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177705152.

Ganser:1966:IDP

[Gan66] C. C. Ganser. Infinitely differentiable positive definite functions. Annals of Mathematical Statistics, 37(2):504–508, April 1966. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177699534.

Ganssler:1970:NRD

[Gän70] Peter Gänssler. Note on a result of Dudley on the speed of mean Glivenko–Cantelli convergence. Annals of Mathematical Statistics, 41(4):1339–1343, August 1970. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177696908. See [Dud69c].

Garver:1932:CLM

[Gar32] Raymond Garver. Concerning the limits of a measure of skewness. Annals of Mathematical Statistics, 3(4):358–360, November 1932. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177732840.

Gart:1959:ECR

[Gar59] John J. Gart. An extension of the Cramér-Rao inequality. Annals of Mathematical Statistics, 30(2):367–380, June 1959. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177706257.

Gardner:1969:DCM

[Gar69] L. A. Gardner, Jr. On detecting changes in the mean of normal variates. Annals of Mathematical Statistics, 40(1):116–126, February 1969. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177697808.

Gastwirth:1965:AMP

[Gas65] Joseph L. Gastwirth. Asymptotically most powerful rank tests for the two-sample problem with censored data. Annals of Mathematical Statistics, 36(4):1243–1247, August 1965. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177699995.

Gautschi:1957:SRS

[Gau57] Werner Gautschi. Some remarks on systematic sampling. Annals of Mathematical Statistics, 28(2):385–394, June 1957. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177706966.

Gautschi:1959:SRH

[Gau59] Werner Gautschi. Some remarks on Herbach’s paper, “Optimum Nature of the F -Test for Model II in the Balanced Case”. Annals of Mathematical Statistics, 30(4):960–963, December 1959. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177706078.

Gaver:1959:IMC

[Gav59] Donald P. Gaver, Jr. Imbedded Markov chain analysis of a waitingline process in continuous time. Annals of Mathematical Statistics, 30(3):698–720, September 1959. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177706200.

Graybill:1964:SSR

[GC64] Franklin A. Graybill and Terrence L. Connell. Sample size required for estimating the variance within d units of the true value. Annals of Mathematical Statistics, 35(1):438–440, March 1964. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177703771.

Gumbel:1965:NM

[GCM65] E. J. Gumbel, P. G. Carlson, and C. K. Mustafi. A note on midrange. Annals of Mathematical Statistics, 36(3):1052–1054, June 1965. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177700080.

Greenwood:1955:DLC

[GD55] J. Arthur Greenwood and David Durand. The distribution of length and components of the sum of n random unit vectors. Annals of Mathematical Statistics, 26(2):233–246, June 1955. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177728540.

Gebhard:1964:LTR

[Geb64a] Richard F. Gebhard. A limit theorem for random interval sampling of a stochastic process. Annals of Mathematical Statistics, 35(2):

866–868, June 1964. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177703587.

Gebhardt:1964:RSS

[Geb64b] Friedrich Gebhardt. On the risk of some strategies for outlying observations. Annals of Mathematical Statistics, 35(4):1524–1536, December 1964. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177700376.

Gebhardt:1966:ESR

[Geb66] Friedrich Gebhardt. On the effect of stragglers on the risk of some mean estimators in small samples. Annals of Mathematical Statistics, 37(2):441–450, April 1966. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177699526.

Gebhardt:1968:CNC

[Geb68] Friedrich Gebhardt. Correction notes: Correction to abstract “Some Numerical Results to R. Borges’ Approximation of the Binomial Distribution”. Annals of Mathematical Statistics, 39(6):2162, December 1968. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177698054.

Geertsema:1970:SCI

[Gee70] J. C. Geertsema. Sequential confidence intervals based on rank tests. Annals of Mathematical Statistics, 41(3):1016–1026, June 1970. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177696978.

Geiringer:1938:PTA

[Gei38] Hilda Geiringer. On the probability theory of arbitrarily linked events. Annals of Mathematical Statistics, 9(4):260–271, December 1938. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/

1177732283. See errata [Gei39].

Geiringer:1939:EPT

[Gei39] Hilda Geiringer. Errata: “On the Probability Theory of Arbitrarily Linked Events”. Annals of Mathematical Statistics, 10(2):202, June 1939. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/

1177732220. See [Gei38].

Geiringer:1940:GLL

[Gei40] Hilda Geiringer. A generalization of the law of large numbers. Annals of Mathematical Statistics, 11(4):393–401, December 1940. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177731826.

Geiringer:1942:NPA

[Gei42a] Hilda Geiringer. A note on the probability of arbitrary events. Annals of Mathematical Statistics, 13(2):238–245, June 1942. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177731610.

Geiringer:1942:OAV

[Gei42b] Hilda Geiringer. Observations on analysis of variance theory. Annals of Mathematical Statistics, 13(3):350–369, September 1942. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177731574.

Geiringer:1944:PTL

[Gei44] Hilda Geiringer. On the probability theory of linkage in Mendelian heredity. Annals of Mathematical Statistics, 15(1):25–57, March 1944. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177731313. See

remarks [Gei45a].

Geiringer:1945:FRL

[Gei45a] Hilda Geiringer. Further remarks on linkage theory in Mendelian heredity. Annals of Mathematical Statistics, 16(4):390–393, December 1945. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/

1177731066. See [Gei44].

Geiringer:1945:DDT

[Gei45b] Hilda Geiringer. On the definition of distance in the theory of the gene. Annals of Mathematical Statistics, 16(4):393–398, December 1945. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177731067.

Geisser:1956:MMS

[Gei56a] Seymour Geisser. The modified mean square successive difference and related statistics. Annals of Mathematical Statistics, 27(3):819– 824, September 1956. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177728188.

Geisser:1956:NND

[Gei56b] Seymour Geisser. A note on the normal distribution. Annals of Mathematical Statistics, 27(3):858–859, September 1956. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177728198.

Geisser:1957:DRC

[Gei57] Seymour Geisser. The distribution of the ratios of certain quadratic forms in time series. Annals of Mathematical Statistics, 28(3):724– 730, September 1957. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177706884.

Geisser:1965:BEM

[Gei65] Seymour Geisser. Bayesian estimation in multivariate analysis. Annals of Mathematical Statistics, 36(1):150–159, February 1965. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177700279.

Geisser:1967:EAL

[Gei67] Seymour Geisser. Estimation associated with linear discriminants. Annals of Mathematical Statistics, 38(3):807–817, June 1967. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177698875.

Geman:1972:VNZ

[Gem72] Donald Geman. On the variance of the number of zeros of a stationary Gaussian process. Annals of Mathematical Statistics, 43(3): 977–982, June 1972. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177692560.

Gerber:1971:DCL

[Ger71] Hans U. Gerber. The discounted Central Limit Theorem and its Berry–Esseen analogue. Annals of Mathematical Statistics, 42(1):

389–392, February 1971. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177693529.

Gessaman:1970:CNM

[Ges70] M. P. Gessaman. A consistent nonparametric multivariate density estimator based on statistically equivalent blocks. Annals of Mathematical Statistics, 41(4):1344–1346, August 1970. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177696909.

Getoor:1965:AFE

[Get65] R. K. Getoor. Additive functionals and excessive functions. Annals of Mathematical Statistics, 36(2):409–422, April 1965. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177700152.

Getoor:1966:RBD

[Get66] Ronald K. Getoor. Review: E. B. Dynkin, J. Fabins, V. Greenberg, A. Maitra, G. Majone, Markov Processes, Volumes I and II. Annals of Mathematical Statistics, 37(5):1408–1409, October 1966. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177699287.

Getoor:1967:RVS

[Get67a] R. K. Getoor. Review: A. V. Skorokhod, Studies in the Theory of Random Processes. Annals of Mathematical Statistics, 38(1):287, February 1967. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177699087.

Getoor:1967:SRC

[Get67b] R. K. Getoor. Some remarks on continuous additive functionals. Annals of Mathematical Statistics, 38(6):1655–1660, December 1967. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177698599.

Getoor:1971:REF

[Get71] R. K. Getoor. Regularity of excessive functions. II. Annals of Mathematical Statistics, 42(6):2056–2063, December 1971. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177693074.

Greenwood:1939:PAG

[GG39] Joseph A. Greenwood and T. N. E. Greville. On the probability of attaining a given standard deviation ratio in an infinite series of trials. Annals of Mathematical Statistics, 10(3):297–298, September 1939. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177732189.

Greenwood:1950:DMM

[GG50] Robert E. Greenwood and Mark O. Glasgow. Distribution of maximum and minimum frequencies in a sample drawn from a multinomial distribution. Annals of Mathematical Statistics, 21(3):416–424, September 1950. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177729800.

Gumbel:1951:TAD

[GG51] E. J. Gumbel and J. Arthur Greenwood. Table of the asymptotic distribution of the second extreme. Annals of Mathematical Statistics, 22(1):121–124, March 1951. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177729702.

Geisser:1958:EBR

[GG58] Seymour Geisser and Samuel W. Greenhouse. An extension of Box’s results on the use of the F distribution in multivariate analysis. Annals of Mathematical Statistics, 29(3):885–891, September 1958. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177706545.

Gardiner:1959:TOR

[GGH59] D. A. Gardiner, A. H. E. Grandage, and R. J. Hader. Third order rotatable designs for exploring response surfaces. Annals of Mathematical Statistics, 30(4):1082–1096, December 1959. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177706092.

Gumbel:1951:EDE

[GH51] E. J. Gumbel and L. H. Herbach. The exact distribution of the extremal quotient. Annals of Mathematical Statistics, 22(3):418–426, September 1951. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177729588.

Graybill:1961:TCE

[GH61] Franklin A. Graybill and Robert A. Hultquist. Theorems concerning Eisenhart’s model II. Annals of Mathematical Statistics, 32(1):261– 269, March 1961. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177705158.

Gross:1968:DGS

[GH68] Shulamith Gross and Paul W. Holland. The distribution of Galton’s statistic. Annals of Mathematical Statistics, 39(6):2114–2117, December 1968. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177698040.

Gilliland:1969:ECD

[GH69] Dennis C. Gilliland and James F. Hannan. On an extended compound decision problem. Annals of Mathematical Statistics, 40(5):1536– 1541, October 1969. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177697371.

Ghosh:1954:ADS

[Gho54] M. N. Ghosh. Asymptotic distribution of serial statistics and applications to problems of nonparametric tests of hypotheses. Annals of Mathematical Statistics, 25(2):218–251, June 1954. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177728783.

Ghosal:1960:EFD

[Gho60] A. Ghosal. Emptiness in the finite dam. Annals of Mathematical Statistics, 31(3):803–808, September 1960. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177705811.

Ghosh:1963:OST

[Gho63a] S. P. Ghosh. Optimum stratification with two characters. Annals of Mathematical Statistics, 34(3):866–872, September 1963. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177704009.

Ghosh:1963:PCS

[Gho63b] Sakti P. Ghosh. Post cluster sampling. Annals of Mathematical Statistics, 34(2):587–597, June 1963. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177704171.

Ghosh:1964:AST

[Gho64a] M. N. Ghosh. On the admissibility of some tests of MANOVA. Annals of Mathematical Statistics, 35(2):789–794, June 1964. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177703578.

Ghosh:1964:UAM

[Gho64b] M. N. Ghosh. Uniform approximation of minimax point estimates. Annals of Mathematical Statistics, 35(3):1031–1047, September 1964. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177703262.

Ghosh:1966:PS

[Gho66] Sakti P. Ghosh. Polychotomy sampling. Annals of Mathematical Statistics, 37(3):657–665, June 1966. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177699459.

Ghosh:1971:NPB

[Gho71a] J. K. Ghosh. A new proof of the Bahadur representation of quantiles and an application. Annals of Mathematical Statistics, 42(6):1957– 1961, December 1971. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177693063.

Ghosh:1971:WOR

[Gho71b] Malay Ghosh. On the Wald-optimality of rank-order tests for paired comparisons. Annals of Mathematical Statistics, 42(6):1970–1976, December 1971. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177693065.

Griego:1971:ASC

[GHRM71] Richard J. Griego, David Heath, and Alberto Ruiz-Moncayo. Almost sure convergence of uniform transport processes to Brownian motion. Annals of Mathematical Statistics, 42(3):1129–1131, June 1971. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177693346.

Ghurye:1954:RFS

[Ghu54] S. G. Ghurye. Random functions satisfying certain linear relations. Annals of Mathematical Statistics, 25(3):543–554, September 1954. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177728722.

Ghurye:1955:RFS

[Ghu55] S. G. Ghurye. Random functions satisfying certain linear relations, II. Annals of Mathematical Statistics, 26(1):105–111, March 1955. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177728597.

Ghurye:1958:NSS

[Ghu58] S. G. Ghurye. Note on sufficient statistics and two-stage procedures. Annals of Mathematical Statistics, 29(1):155–166, March 1958. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177706712.

Ghurye:1968:ISS

[Ghu68] S. G. Ghurye. Information and sufficient sub-fields. Annals of Mathematical Statistics, 39(6):2056–2066, December 1968. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177698033.

Gilbert:1959:RG

[Gil59a] E. N. Gilbert. Random graphs. Annals of Mathematical Statistics, 30 (4):1141–1144, December 1959. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177706098.

Gilbert:1959:IPF

[Gil59b] Edgar J. Gilbert. On the identifiability problem for functions of finite Markov chains. Annals of Mathematical Statistics, 30(3):688–697, September 1959. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177706199.

Gilbert:1962:RSS

[Gil62] E. N. Gilbert. Random subdivisions of space into crystals. Annals of Mathematical Statistics, 33(3):958–972, September 1962. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177704464.

Gilbert:1964:RHD

[Gil64a] E. N. Gilbert. Review: H. A. David, The Method of Paired Comparisons. Annals of Mathematical Statistics, 35(3):1386–1387, September 1964. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177703303.

Gilliland:1964:NMN

[Gil64b] Dennis C. Gilliland. A note on the maximization of a non-central chi-square probability. Annals of Mathematical Statistics, 35(1):441– 442, March 1964. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177703772.

Gilliland:1968:SCE

[Gil68] Dennis C. Gilliland. Sequential compound estimation. Annals of Mathematical Statistics, 39(6):1890–1904, December 1968. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177698019.

Gilliland:1969:ABR

[Gil69] Dennis C. Gilliland. Approximation to Bayes risk in sequences of non-finite games. Annals of Mathematical Statistics, 40(2):467–474, April 1969. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177697712.

Gilat:1971:NTS

[Gil71a] David Gilat. On the nonexistence of a three series condition for series of nonindependent random variables. Annals of Mathematical Statistics, 42(1):409, February 1971. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177693535.

Gilat:1971:SCU

[Gil71b] David Gilat. Some conditions under which two random variables are equal almost surely and a simple proof of a theorem of Chung and Fuchs. Annals of Mathematical Statistics, 42(5):1647–1655, October 1971. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177693163.

Gilat:1972:CDC

[Gil72] David Gilat. Convergence in distribution, convergence in probability and almost sure convergence of discrete martingales. Annals of Mathematical Statistics, 43(4):1374–1379, August 1972. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177692494.

Girshick:1939:STR

[Gir39] M. A. Girshick. On the sampling theory of roots of determinantal equations. Annals of Mathematical Statistics, 10(3):203–224, September 1939. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177732180.

Girshick:1942:CND

[Gir42a] M. A. Girshick. A correction: “Note on the Distribution of Roots of a Polynomial with Random Complex Coefficients”. Annals of Mathematical Statistics, 13(4):447, December 1942. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177731545. See [Gir42b].

Girshick:1942:NDR

[Gir42b] M. A. Girshick. Note on the distribution of roots of a polynomial with random complex coefficients. Annals of Mathematical Statistics, 13 (2):235–238, June 1942. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.

aoms/1177731609. See correction [Gir42a].

Girshick:1946:CTSa

[Gir46a] M. A. Girshick. Contributions to the theory of sequential analysis. I. Annals of Mathematical Statistics, 17(2):123–143, June 1946. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177730976.

Girshick:1946:CTSb

[Gir46b] M. A. Girshick. Contributions to the theory of sequential analysis, II, III. Annals of Mathematical Statistics, 17(3):282–298, September 1946. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177730941.

Girshick:1948:RAW

[Gir48] M. A. Girshick. Review: Abraham Wald, Sequential Analysis. Annals of Mathematical Statistics, 19(1):122–123, March 1948. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177730301.

Girshick:1958:EOP

[Gir58] M. A. Girshick. An extension of the optimum property of the sequential probability ratio test. Annals of Mathematical Statistics, 29(1): 288–290, March 1958. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177706728.

Giri:1964:CNC

[Gir64a] N. Giri. Correction notes: Correction to “On the Likelihood Ratio of a Normal Multivariate Testing Problem”. Annals of Mathematical Statistics, 35(3):1388, September 1964. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177703305. See [Gir64b].

Giri:1964:LRT

[Gir64b] N. Giri. On the likelihood ratio test of a normal multivariate testing problem. Annals of Mathematical Statistics, 35(1):181–189, March 1964. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/

1177703740. See correction [Gir64a].

Giri:1965:CAT

[Gir65a] N. Giri. On the complex analogues of T^{2}- and R^{2}-tests. Annals of Mathematical Statistics, 36(2):664–670, April 1965. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177700173.

Giri:1965:LRT

[Gir65b] N. Giri. On the likelihood ratio test of a normal multivariate testing problem II. Annals of Mathematical Statistics, 36(3):1061–1065, June 1965. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177700083.

Giri:1968:CNC

[Gir68a] N. Giri. Correction note: Correction to “On Tests of the Equality of Two Covariance Matrices”. Annals of Mathematical Statistics, 39 (5):1764, October 1968. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177698165. See [Gir68c].

Giri:1968:LAM

[Gir68b] N. Giri. Locally and asymptotically minimax tests of a multivariate problem. Annals of Mathematical Statistics, 39(1):171–178, February 1968. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177698515.

Giri:1968:TET

[Gir68c] N. Giri. On tests of the equality of two covariance matrices. Annals of Mathematical Statistics, 39(1):275–277, February 1968. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177698533. See correc-

tion [Gir68a].

Guthrie:1959:BAS

[GJ59] D. Guthrie, Jr. and M. V. Johns, Jr. Bayes acceptance sampling procedures for large lots. Annals of Mathematical Statistics, 30(4):896– 925, December 1959. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177706075.

Godambe:1965:ABE

[GJ65] V. P. Godambe and V. M. Joshi. Admissibility and Bayes estimation in sampling finite populations. I. Annals of Mathematical Statistics, 36(6):1707–1722, December 1965. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177699799.

Gumbel:1950:EQ

[GK50a] E. J. Gumbel and R. D. Keeney. The extremal quotient. Annals of Mathematical Statistics, 21(4):523–538, December 1950. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177729749.

Gumbel:1950:GRD

[GK50b] E. J. Gumbel and R. D. Keeney. The geometric range for distributions of Cauchy’s type. Annals of Mathematical Statistics, 21(1): 133–137, March 1950. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177729895.

Giri:1964:LAM

[GK64a] N. Giri and J. Kiefer. Local and asymptotic minimax properties of multivariate tests. Annals of Mathematical Statistics, 35(1):21– 35, March 1964. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177703730.

Giri:1964:MCT

[GK64b] N. Giri and J. Kiefer. Minimax character of the R^{2}-test in the simplest case. Annals of Mathematical Statistics, 35(4):1475–1490, December 1964. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177700374.

Girshick:1957:MSD

[GKR57] M. A. Girshick, S. Karlin, and H. L. Royden. Multistage statistical decision procedures. Annals of Mathematical Statistics, 28(1):111– 125, March 1957. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177707040.

Giri:1963:MCH

[GKS63] N. Giri, J. Kiefer, and C. Stein. Minimax character of Hotelling’s T^{2} test in the simplest case. Annals of Mathematical Statistics, 34 (4):1524–1535, December 1963. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177703884.

Goldberg:1946:AFP

[GL46] Henry Goldberg and Harriet Levine. Approximate formulas for the percentage points and normalization of t and χ^{2}. Annals of Mathematical Statistics, 17(2):216–225, June 1946. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177730982.

Glasgow:1959:NFM

[Gla59] M. O. Glasgow. Note on the factorial moments of the distribution of locally maximal elements in a random sample. Annals of Mathematical Statistics, 30(2):586–590, June 1959. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177706273.

Gleser:1964:MTE

[Gle64] Leon J. Gleser. On a measure of test efficiency proposed by R. R. Bahadur. Annals of Mathematical Statistics, 35(4):1537–1544, December 1964. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177700377.

Gleser:1965:ATF

[Gle65] Leon J. Gleser. On the asymptotic theory of fixed-size sequential confidence bounds for linear regression parameters. Annals of Mathematical Statistics, 36(2):463–467, April 1965. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http:

//projecteuclid.org/euclid.aoms/1177700157. See correction

[Gle66b].

Gleser:1966:NST

[Gle66a] Leon J. Gleser. A note on the sphericity test. Annals of Mathematical Statistics, 37(2):464–467, April 1966. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177699529. See correction [Gle68].

Gleser:1966:CNC

[Gle66b] Leon Jay Gleser. Correction note: Correction to “On the Asymptotic Theory of Fixed-Size Sequential Confidence Bounds for Linear Regression Parameters”. Annals of Mathematical Statistics, 37(4): 1053–1055, August 1966. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177699388. See [Gle65].

Gleser:1968:CNC

[Gle68] Leon J. Gleser. Correction notes: Correction to “A Note on the Sphericity Test”. Annals of Mathematical Statistics, 39(2):684, April 1968. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/

1177698429. See [Gle66a].

Gleser:1969:LDS

[Gle69] Leon Jay Gleser. On limiting distributions for sums of a random number of independent random vectors. Annals of Mathematical Statistics, 40(3):935–941, June 1969. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177697598.

Graybill:1957:IMQ

[GM57] Franklin A. Graybill and George Marsaglia. Idempotent matrices and quadratic forms in the general linear hypothesis. Annals of Mathematical Statistics, 28(3):678–686, September 1957. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177706879.

Geisser:1962:PIJ

[GM62] Seymour Geisser and Nathan Mantel. Pairwise independence of jointly dependent variables. Annals of Mathematical Statistics, 33(1): 290–291, March 1962. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177704732.

Graybill:1969:QFI

[GM69] Franklin A. Graybill and George Milliken. Quadratic forms and idempotent matrices with random elements. Annals of Mathematical Statistics, 40(4):1430–1438, August 1969. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177697514.

Gordon:1972:CMN

[GM72] F. S. Gordon and A. M. Mathai. Characterizations of the multivariate normal distribution using regression properties. Annals of Mathematical Statistics, 43(1):205–229, February 1972. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177692714.

Girshick:1946:UEC

[GMS46] M. A. Girshick, Frederick Mosteller, and L. J. Savage. Unbiased estimates for certain binomial sampling problems with applications. Annals of Mathematical Statistics, 17(1):13–23, March 1946. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177731018.

Gnanadesikan:1959:EMT

[Gna59] R. Gnanadesikan. Equality of more than two variances and of more than two dispersion matrices against certain alternatives. Annals of Mathematical Statistics, 30(1):177–184, March 1959. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177706372. See correc-

tion [Gna60].

Gnanadesikan:1960:CNC

[Gna60] R. Gnanadesikan. Correction notes: Correction to and comment on “Equality of More Than Two Variances and of More Than Two Dispersion Matrices Against Certain Alternatives”. Annals of Mathematical Statistics, 31(1):227–228, March 1960. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177706002. See [Gna59].

Gnanadesikan:1967:RTM

[Gna67] R. Gnanadesikan. Review: Tito A. Mijares, Percentage Points of the Sum V^{(}s)_{1} of s Roots (s = 1–50) A Unified Table for Tests of Significance in Various Univariate and Multivariate Hypotheses. Annals of Mathematical Statistics, 38(2):628–631, April 1967. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177698991.

Gnedenko:1964:RMR

[Gne64] Boris Gnedenko. Review: Murray Rosenblatt, Random Processes. Annals of Mathematical Statistics, 35(4):1832–1833, December 1964. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177700411.

Ghurye:1962:CMN

[GO62] S. G. Ghurye and Ingram Olkin. A characterization of the multivariate normal distribution. Annals of Mathematical Statistics, 33 (2):533–541, June 1962. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177704579.

Ghurye:1969:UES

[GO69] S. G. Ghurye and Ingram Olkin. Unbiased estimation of some multivariate probability densities and related functions. Annals of Mathematical Statistics, 40(4):1261–1271, August 1969. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177697501.

Godwin:1949:NKD

[God49a] H. J. Godwin. A note on Kac’s derivation of the distribution of the mean deviation. Annals of Mathematical Statistics, 20(1):127, March 1949. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/

1177730104. See [Kac48].

Godwin:1949:SLM

[God49b] H. J. Godwin. Some low moments of order statistics. Annals of Mathematical Statistics, 20(2):279–285, June 1949. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177730036.

Godambe:1960:OPR

[God60] V. P. Godambe. An optimum property of regular maximum likelihood estimation. Annals of Mathematical Statistics, 31(4):1208–1211, December 1960. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177705693.

Godambe:1961:AP

[God61a] V. P. Godambe. Acknowledgment of priority. Annals of Mathematical Statistics, 32(4):1343–1344, December 1961. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177704876.

Godambe:1961:TSP

[God61b] V. P. Godambe. On the two sample problem: A heuristic method for constructing tests. Annals of Mathematical Statistics, 32(4):1091– 1107, December 1961. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177704848.

Godambe:1969:ABE

[God69] V. P. Godambe. Admissibility and Bayes estimation in sampling finite populations–v. Annals of Mathematical Statistics, 40(2):672– 676, April 1969. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177697738.

Goffman:1953:NVM

[Gof53] Casper Goffman. Note on the variation of means. Annals of Mathematical Statistics, 24(2):307–309, June 1953. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177729039.

Gold:1963:TAT

[Gol63a] Ruth Z. Gold. Tests auxiliary to χ^{2} tests in a Markov chain. Annals of Mathematical Statistics, 34(1):56–74, March 1963. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177704242.

Golub:1963:CVM

[Gol63b] Gene H. Golub. Comparison of the variance of minimum variance and weighted least squares regression coefficients. Annals of Mathematical Statistics, 34(3):984–991, September 1963. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177704021.

Goldman:1967:SPP

[Gol67] Jay R. Goldman. Stochastic point processes: Limit theorems. Annals of Mathematical Statistics, 38(3):771–779, June 1967. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177698870.

Goldman:1971:FPI

[Gol71] Malcolm Goldman. On the first passage of the integrated Wiener process. Annals of Mathematical Statistics, 42(6):2150–2155, December 1971. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177693084.

Goodman:1949:ENC

[Goo49] Leo A. Goodman. On the estimation of the number of classes in a population. Annals of Mathematical Statistics, 20(4):572–579, December 1949. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177729949.

Goodman:1952:ASL

[Goo52] Leo A. Goodman. On the analysis of samples from k lists. Annals of Mathematical Statistics, 23(4):632–634, December 1952. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177729345.

Goodman:1953:SST

[Goo53a] Leo A. Goodman. Sequential sampling tagging for population size problems. Annals of Mathematical Statistics, 24(1):56–69, March 1953. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177729082.

Goodman:1953:SMI

[Goo53b] Leo A. Goodman. A simple method for improving some estimators. Annals of Mathematical Statistics, 24(1):114–117, March 1953. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177729089.

Good:1956:SIM

[Goo56] I. J. Good. The surprise index for the multivariate normal distribution. Annals of Mathematical Statistics, 27(4):1130–1135, December 1956. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/

1177728079. See correction [Goo57a].

Good:1957:CSI

[Goo57a] I. J. Good. Corrections to “The Surprise Index for the Multivariate Normal Distribution”. Annals of Mathematical Statistics, 28(4):1055, December 1957. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/

1177706815. See [Goo56].

Good:1957:STR

[Goo57b] I. J. Good. On the serial test for random sequences. Annals of Mathematical Statistics, 28(1):262–264, March 1957. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http:/

/projecteuclid.org/euclid.aoms/1177707053;   http://www.
jstor.org/stable/2237039.

Good:1957:SPM

[Goo57c] I. J. Good. Saddle-point methods for the multinomial distribution. Annals of Mathematical Statistics, 28(4):861–881, December 1957. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177706790. See

corrections [Goo61b].

Goodman:1958:ADP

[Goo58a] Leo A. Goodman. Asymptotic distributions of “psi–squared” goodness of fit criteria for m-th order Markov chains. Annals of Mathematical Statistics, 29(4):1123–1133, December 1958. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177706445.

Goodman:1958:EPA

[Goo58b] Leo A. Goodman. Exact probabilities and asymptotic relationships for some statistics from m-th order Markov chains. Annals of Mathematical Statistics, 29(2):476–490, June 1958. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177706623.

Goodman:1959:SST

[Goo59] Leo A. Goodman. On some statistical tests for M th order Markov chains. Annals of Mathematical Statistics, 30(1):154–164, March 1959. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177706368.

Good:1961:AFD

[Goo61a] I. J. Good. An asymptotic formula for the differences of the powers at zero. Annals of Mathematical Statistics, 32(1):249–256, March 1961. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177705156.

Good:1961:CNC

[Goo61b] I. J. Good. Corrections notes: Corrections to “Saddle-point Methods for the Multinomial Distribution”. Annals of Mathematical Statistics, 32(2):619, June 1961. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177705072. See [Goo57c].

Good:1961:FCM

[Goo61c] I. J. Good. The frequency count of a Markov chain and the transition to continuous time. Annals of Mathematical Statistics, 32(1): 41–48, March 1961. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177705137.

Good:1961:MSM

[Goo61d] I. J. Good. The multivariate saddlepoint method and chi-squared for the multinomial distribution. Annals of Mathematical Statistics, 32 (2):535–548, June 1961. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177705059.

Goodman:1961:SS

[Goo61e] Leo A. Goodman. Snowball sampling. Annals of Mathematical Statistics, 32(1):148–170, March 1961. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177705148.

Good:1963:MEH

[Goo63a] I. J. Good. Maximum entropy for hypothesis formulation, especially for multidimensional contingency tables. Annals of Mathematical Statistics, 34(3):911–934, September 1963. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177704014.

Goodman:1963:DDC

[Goo63b] N. R. Goodman. The distribution of the determinant of a complex Wishart distributed matrix. Annals of Mathematical Statistics, 34(1): 178–180, March 1963. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177704251.

Goodman:1963:SAB

[Goo63c] N. R. Goodman. Statistical analysis based on a certain multivariate complex Gaussian distribution (an introduction). Annals of Mathematical Statistics, 34(1):152–177, March 1963. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177704250.

Goodman:1964:IMC

[Goo64a] Leo A. Goodman. Interactions in multidimensional contingency tables. Annals of Mathematical Statistics, 35(2):632–646, June 1964. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177703561.

Goodman:1964:SCI

[Goo64b] Leo A. Goodman. Simultaneous confidence intervals for contrasts among multinomial populations. Annals of Mathematical Statistics, 35(2):716–725, June 1964. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177703569. See correction [Goo76].

Goodman:1968:HMM

[Goo68] Leo A. Goodman. How to minimize or maximize the probabilities of extinction in a Galton–Watson process and in some related multiplicative population processes. Annals of Mathematical Statistics, 39(5):1700–1710, October 1968. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177698152.

Goodman:1976:NCS

[Goo76] Leo A. Goodman. Notes: Correction to “Simultaneous Confidence Intervals for Contrasts Among Multinomial Populations”. Annals of Statistics, 4(4):821, July 1976. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.org/

euclid.aos/1176343558. See [Goo64b].

Gordon:1941:EQW

[Gor41a] Robert D. Gordon. The estimation of a quotient when the denominator is normally distributed. Annals of Mathematical Statistics, 12(1): 115–118, March 1941. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177731792.

Gordon:1941:VMR

[Gor41b] Robert D. Gordon. Values of Mills’ ratio of area to bounding ordinate and of the normal probability integral for large values of the argument. Annals of Mathematical Statistics, 12(3):364–366, September 1941. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177731721.

Gordon:1972:EMS

[Gor72] Louis Gordon. An equivalent to the martingale square function inequality. Annals of Mathematical Statistics, 43(6):1927–1934, December 1972. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177690863.

Gould:1963:NTB

[Gou63] H. W. Gould. Note on two binomial coefficient sums found by Riordan. Annals of Mathematical Statistics, 34(1):333–335, March 1963. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177704272.

Govindarajulu:1962:ELM

[Gov62] Zakkula Govindarajulu. Exact lower moments of order statistics in samples from the chi-distribution (1 d.f.). Annals of Mathematical Statistics, 33(4):1292–1305, December 1962. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177704362.

Govindarajulu:1963:MOS

[Gov63] Zakkula Govindarajulu. On moments of order statistics and quasiranges from normal populations. Annals of Mathematical Statistics, 34(2):633–651, June 1963. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177704176.

Govindarajulu:1966:CNG

[Gov66] Zakkula Govindarajulu. Characterization of normal and generalized truncated normal distributions using order statistics. Annals of Mathematical Statistics, 37(4):1011–1015, August 1966. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177699380.

Graybill:1958:SDT

[GP58] Franklin A. Graybill and William E. Pruitt. The staircase design: Theory. Annals of Mathematical Statistics, 29(2):523–533, June 1958. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177706627.

Gumbel:1967:PTE

[GP67] E. J. Gumbel and James Pickands III. Probability tables for the extremal quotient. Annals of Mathematical Statistics, 38(5):1541– 1551, October 1967. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177698708.

Ganssler:1971:CCE

[GP71] P. Gänssler and J. Pfanzagl. Convergence of conditional expectations. Annals of Mathematical Statistics, 42(1):315–324, February 1971. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177693514.

Gupta:1972:CSS

[GP72] Shanti S. Gupta and S. Panchapakesan. On a class of subset selection procedures. Annals of Mathematical Statistics, 43(3):814– 822, June 1972. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177692547.

Girshick:1952:BAQ

[GR52] M. A. Girshick and Herman Rubin. A Bayes approach to a quality control model. Annals of Mathematical Statistics, 23(1):114–125, March 1952. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177729489.

Grenander:1953:SSA

[GR53] Ulf Grenander and Murray Rosenblatt. Statistical spectral analysis of time series arising from stationary stochastic processes. Annals of Mathematical Statistics, 24(4):537–558, December 1953. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177728913.

Gastwirth:1969:RLE

[GR69] Joseph L. Gastwirth and Herman Rubin. On robust linear estimators. Annals of Mathematical Statistics, 40(1):24–39, February 1969. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177697801.

Getoor:1970:ALD

[GR70] R. K. Getoor and Murali Rao. Another look at Doob’s theorem. Annals of Mathematical Statistics, 41(2):503–506, April 1970. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177697090.

Graybill:1954:QEV

[Gra54] Franklin A. Graybill. On quadratic estimates of variance components. Annals of Mathematical Statistics, 25(2):367–372, June 1954. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177728792.

Graybill:1958:DSS

[Gra58] Franklin A. Graybill. Determining sample size for a specified width confidence interval. Annals of Mathematical Statistics, 29(1):282– 287, March 1958. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177706727.

Graham:1963:CTP

[Gra63] R. L. Graham. A combinatorial theorem for partial sums. Annals of Mathematical Statistics, 34(4):1600–1602, December 1963. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177703895.

Graybill:1967:RCR

[Gra67] Franklin A. Graybill. Review: C. R. Rao, Linear Statistical Inference and Its Applications. Annals of Mathematical Statistics, 38(1):281– 284, February 1967. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177699085.

Gray:1968:SSE

[Gra68] K. B. Gray, Jr. Sequential selection of experiments. Annals of Mathematical Statistics, 39(6):1953–1977, December 1968. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177698025.

Greenleaf:1932:CAM

[Gre32] Herrick E. H. Greenleaf. Curve approximation by means of functions analogous to the Hermite polynomials. Annals of Mathematical Statistics, 3(3):204–255, August 1932. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177732887.

Greenwood:1938:VGM

[Gre38] Joseph A. Greenwood. Variance of a general matching problem. Annals of Mathematical Statistics, 9(1):56–59, March 1938. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177732359.

Greville:1941:FDG

[Gre41] T. N. E. Greville. The frequency distribution of a general matching problem. Annals of Mathematical Statistics, 12(3):350–354, September 1941. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177731718.

Greville:1944:GWF

[Gre44a] T. N. E. Greville. A generalization of Waring’s formula. Annals of Mathematical Statistics, 15(2):218–219, June 1944. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177731288. See

remark [Gre47].

Greville:1944:MMO

[Gre44b] T. N. E. Greville. On multiple matching with one variable deck. Annals of Mathematical Statistics, 15(4):432–434, December 1944. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177731215.

Greville:1947:RNG

[Gre47] T. N. E. Greville. Remark on the note “A Generalization of Waring’s Formula”. Annals of Mathematical Statistics, 18(4):605–606, December 1947. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/

1177730356. See [Gre44a].

Greenwood:1949:NIL

[Gre49] Robert E. Greenwood. Numerical integration for linear sums of exponential functions. Annals of Mathematical Statistics, 20(4):608–611, December 1949. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177729955.

Greville:1950:RAC

[Gre50a] T. N. E. Greville. Remark on the article “On a Class of Distributions that Approach the Normal Distribution Function” by George B. Dantzig. Annals of Mathematical Statistics, 21(4):611–613, December 1950. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/

1177729757. See [Dan39].

Greville:1950:RWM

[Gre50b] T. N. E. Greville. Remark on W. M. Kincaid’s “Note on the Error in Interpolation of a Function of Two Independent Variables”. Annals of Mathematical Statistics, 21(1):137–138, March 1950. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177729896. See [Kin48a].

Green:1954:CIV

[Gre54a] J. R. Green. A confidence interval for variance components. Annals of Mathematical Statistics, 25(4):671–686, December 1954. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177728654.

Grenander:1954:ERC

[Gre54b] Ulf Grenander. On the estimation of regression coefficients in the case of an autocorrelated disturbance. Annals of Mathematical Statistics, 25(2):252–272, June 1954. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177728784.

Grenander:1964:RBV

[Gre64] Ulf Grenander. Review: B. V. Gnedenko, B. D. Seckler, The Theory of Probability. Annals of Mathematical Statistics, 35(2):920– 922, June 1964. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177703605.

Grenander:1965:LTS

[Gre65a] Ulf Grenander. A limit theorem for sums of minima of stochastic variables. Annals of Mathematical Statistics, 36(3):1041–1042, June 1965. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177700076.

Grenander:1965:SDE

[Gre65b] Ulf Grenander. Some direct estimates of the mode. Annals of Mathematical Statistics, 36(1):131–138, February 1965. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177700277.

Greenberg:1966:REI

[Gre66] Vida L. Greenberg. Robust estimation in incomplete blocks designs. Annals of Mathematical Statistics, 37(5):1331–1337, October 1966. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177699277.

Griffin:1931:FFD

[Gri31] Harold D. Griffin. Fundamental formulas for the Doolittle method, using zero-order correlation coefficients. Annals of Mathematical Statistics, 2(2):150–153, May 1931. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177732993.

Griffiths:1969:CCC

[Gri69] R. C. Griffiths. The canonical correlation coefficients of bivariate gamma distributions. Annals of Mathematical Statistics, 40(4):1401– 1408, August 1969. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177697511.

Girshick:1955:EBR

[GRS55] M. A. Girshick, H. Rubin, and R. Sitgreaves. Estimates of bounded relative error in particle counting. Annals of Mathematical Statistics, 26(2):276–285, June 1955. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177728544.

Gruzewska:1933:PWA

[Gru33] H. Milicer Gruzewska. The precision of the weighted average. Annals of Mathematical Statistics, 4(3):196–215, August 1933. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177732779.

Grummann:1936:SNE

[Gru36] H. R. Grummann. Some notes on exponential analysis. Annals of Mathematical Statistics, 7(3):133–144, September 1936. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177732504.

Grubbs:1944:DRS

[Gru44] Frank E. Grubbs. On the distribution of the radial standard deviation. Annals of Mathematical Statistics, 15(1):75–81, March 1944. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177731316.

Grubbs:1949:DSS

[Gru49] Frank E. Grubbs. On designing single sampling inspection plans. Annals of Mathematical Statistics, 20(2):242–256, June 1949. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177730033.

Grubbs:1950:SCT

[Gru50] Frank E. Grubbs. Sample criteria for testing outlying observations. Annals of Mathematical Statistics, 21(1):27–58, March 1950. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177729885.

Grad:1955:DQF

[GS55] Arthur Grad and Herbert Solomon. Distribution of quadratic forms and some applications. Annals of Mathematical Statistics, 26(3):464– 477, September 1955. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177728491.

Gupta:1957:SWA

[GS57] Shanti S. Gupta and Milton Sobel. On a statistic which arises in selection and ranking problems. Annals of Mathematical Statistics, 28(4):957–967, December 1957. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177706796.

Gupta:1958:SSW

[GS58a] Shanti S. Gupta and Milton Sobel. On selecting a subset which contains all populations better than a standard. Annals of Mathematical Statistics, 29(1):235–244, March 1958. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177706721.

Gupta:1958:DSB

[GS58b] Shanti S. Gupta and Milton Sobel. On the distribution of a statistic based on ordered uniform chance variables. Annals of Mathematical Statistics, 29(1):274–281, March 1958. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177706726.

Graybill:1960:UYM

[GS60] Franklin A. Graybill and V. Seshadri. On the unbiasedness of Yates’ method of estimation using interblock information. Annals of Mathematical Statistics, 31(3):786–787, September 1960. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177705805.

Gupta:1965:EMP

[GS65] Shanti S. Gupta and Bhupendra K. Shah. Exact moments and percentage points of the order statistics and the distribution of the range from the logistic distribution. Annals of Mathematical Statistics, 36 (3):907–920, June 1965. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177700063.

Ghosh:1966:UEL

[GS66] J. K. Ghosh and Rajinder Singh. Unbiased estimation of location and scale parameters. Annals of Mathematical Statistics, 37(6):1671– 1675, December 1966. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177699155.

Gundy:1967:SRC

[GS67] Richard F. Gundy and David Siegmund. On a stopping rule and the Central Limit Theorem. Annals of Mathematical Statistics, 38 (6):1915–1917, December 1967. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177698627.

Ghosh:1971:CRO

[GS71] Malay Ghosh and Pranab Kumar Sen. On a class of rank order tests for regression with partially informed stochastic predictors. Annals of Mathematical Statistics, 42(2):650–661, April 1971. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177693415.

Guenther:1964:RLC

[GT64a] William C. Guenther and Paul J. Terragno. A review of the literature on a class of coverage problems. Annals of Mathematical Statistics, 35(1):232–260, March 1964. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177703747.

Guttman:1964:BAS

[GT64b] Irwin Guttman and George C. Tiao. A Bayesian approach to some best population problems. Annals of Mathematical Statistics, 35(2): 825–835, June 1964. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177703582.

Guard:1949:RBM

[Gua49] Harris T. Guard. Report on the Boulder meeting of the Institute. Annals of Mathematical Statistics, 20(4):626–628, December 1949. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177729961.

Guest:1951:FPM

[Gue51] P. G. Guest. The fitting of polynomials by the method of weighted grouping. Annals of Mathematical Statistics, 22(4):537–548, December 1951. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177729543.

Guest:1958:SOP

[Gue58] P. G. Guest. The spacing of observations in polynomial regression. Annals of Mathematical Statistics, 29(1):294–299, March 1958. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177706730.

Guiasu:1971:ADS

[Gui71] Silviu Guiaşu. On the asymptotic distribution of the sequences of random variables with random indices. Annals of Mathematical Statistics, 42(6):2018–2028, December 1971. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177693069.

Gulati:1972:CTL

[Gul72] Bodh Raj Gulati. Construction of two-level Saturated symmetrical factorial designs of resolution VI. Annals of Mathematical Statistics, 43(5):1652–1663, October 1972. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177692398.

Gumbel:1941:RPF

[Gum41] E. J. Gumbel. The return period of flood flows. Annals of Mathematical Statistics, 12(2):163–190, June 1941. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177731747.

Gumbel:1943:SN

[Gum43a] E. J. Gumbel. On serial numbers. Annals of Mathematical Statistics, 14(2):163–178, June 1943. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177731458.

Gumbel:1943:RCC

[Gum43b] E. J. Gumbel. On the reliability of the classical chi-square test. Annals of Mathematical Statistics, 14(3):253–263, September 1943. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177731419.

Gumbel:1944:RM

[Gum44] E. J. Gumbel. Ranges and midranges. Annals of Mathematical Statistics, 15(4):414–422, December 1944. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177731211.

Gumbel:1946:IES

[Gum46] E. J. Gumbel. On the independence of the extremes in a sample. Annals of Mathematical Statistics, 17(1):78–81, March 1946. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177731026.

Gumbel:1947:DR

[Gum47] E. J. Gumbel. The distribution of the range. Annals of Mathematical Statistics, 18(3):384–412, September 1947. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177730387.

Gumbel:1954:MML

[Gum54] E. J. Gumbel. The maxima of the mean largest value and of the range. Annals of Mathematical Statistics, 25(1):76–84, March 1954. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177728847.

Gundy:1967:MVT

[Gun67] Richard F. Gundy. The martingale version of a theorem of Marcinkiewicz and Zygmund. Annals of Mathematical Statistics, 38 (3):725–734, June 1967. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177698865.

Gundy:1968:DBM

[Gun68] Richard F. Gundy. A decomposition for L^{1}-bounded martingales. Annals of Mathematical Statistics, 39(1):134–138, February 1968. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177698510.

Gupta:1961:PPM

[Gup61] Shanti S. Gupta. Percentage points and modes of order statistics from the normal distribution. Annals of Mathematical Statistics, 32 (3):888–893, September 1961. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177704982.

Gupta:1962:CNC

[Gup62] Somesh Das Gupta. Correction notes: Correction to abstract of “Optimal Classification Rules”: Abstracts of papers. Annals of Mathematical Statistics, 33(4):1466, December 1962. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177704382. See [Ano62b].

Gupta:1963:BMN

[Gup63a] Shanti S. Gupta. Bibliography on the multivariate normal integrals and related topics. Annals of Mathematical Statistics, 34(3):829– 838, September 1963. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177704005.

Gupta:1963:PIM

[Gup63b] Shanti S. Gupta. Probability integrals of multivariate normal and multivariate t. Annals of Mathematical Statistics, 34(3):792–828, September 1963. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177704004.

Gupta:1965:OCR

[Gup65] S. Das Gupta. Optimum classification rules for classification into two multivariate normal populations. Annals of Mathematical Statistics, 36(4):1174–1184, August 1965. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177699990. See correction [Gup70].

Gupta:1967:ANT

[Gup67] M. K. Gupta. An asymptotically nonparametric test of symmetry. Annals of Mathematical Statistics, 38(3):849–866, June 1967. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177698879.

Gupta:1969:PPF

[Gup69] Somesh Das Gupta. Properties of power functions of some tests concerning dispersion matrices of multivariate normal distributions. Annals of Mathematical Statistics, 40(2):697–701, April 1969. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177697745.

Gupta:1970:CNC

[Gup70] S. Das Gupta. Correction notes: Correction to “Optimum Classification Rules for Classification into Two Multivariate Normal Populations”. Annals of Mathematical Statistics, 41(1):326, February 1970. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177697214. See

[Gup65].

Gupta:1971:NDW

[Gup71] A. K. Gupta. Noncentral distribution of Wilks’ statistic in manova. Annals of Mathematical Statistics, 42(4):1254–1261, August 1971. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177693238.

Gupta:1972:RRB

[Gup72] Shanti S. Gupta. Review: R. E. Bechhofer, J. Kiefer, M. Sobel, Sequential Identification and Ranking Procedures (with Special Reference to Koopman–Darmois Populations). Annals of Mathematical Statistics, 43(2):697–700, April 1972. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177692660.

Gurland:1948:IFD

[Gur48] John Gurland. Inversion formulae for the distribution of ratios. Annals of Mathematical Statistics, 19(2):228–237, June 1948. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177730247.

Gurland:1953:DQF

[Gur53] John Gurland. Distribution of quadratic forms and ratios of quadratic forms. Annals of Mathematical Statistics, 24(3):416–427, September 1953. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177728981.

Gurland:1955:DDI

[Gur55a] John Gurland. Distribution of definite and of indefinite quadratic forms. Annals of Mathematical Statistics, 26(1):122–127, March 1955. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177728600. See

correction [Gur62].

Gurland:1955:DMA

[Gur55b] John Gurland. Distribution of the maximum of the arithmetic mean of correlated random variables. Annals of Mathematical Statistics, 26 (2):294–300, June 1955. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177728546.

Gurland:1959:CNC

[Gur59] John Gurland. Correction notes: Corrections to “Distribution of the Maximum of the Arithmetic Mean of Correlated Random Variables”. Annals of Mathematical Statistics, 30(4):1265–1266, December 1959. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177706116.

Gurland:1962:CNC

[Gur62] John Gurland. Correction notes: Correction to “Distribution of Definite and of Indefinite Quadratic Forms”. Annals of Mathematical Statistics, 33(2):813, June 1962. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177704605. See [Gur55a].

Guttman:1938:NDF

[Gut38] Louis Guttman. A note on the derivation of formulae for multiple and partial correlation. Annals of Mathematical Statistics, 9(4):305– 308, December 1938. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177732287.

Guttman:1946:AQP

[Gut46a] Louis Guttman. An approach for quantifying paired comparisons and rank order. Annals of Mathematical Statistics, 17(2):144–163, June 1946. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177730977.

Guttman:1946:EMC

[Gut46b] Louis Guttman. Enlargement methods for computing the inverse matrix. Annals of Mathematical Statistics, 17(3):336–343, September 1946. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177730946.

Guttman:1948:DFC

[Gut48a] Louis Guttman. A distribution-free confidence interval for the mean. Annals of Mathematical Statistics, 19(3):410–413, September 1948. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177730207.

Guttman:1948:IK

[Gut48b] Louis Guttman. An inequality for kurtosis. Annals of Mathematical Statistics, 19(2):277–278, June 1948. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177730255.

Guttman:1957:POT

[Gut57] Irwin Guttman. On the power of optimum tolerance regions when sampling from normal distributions. Annals of Mathematical Statistics, 28(3):773–778, September 1957. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177706890.

Guttman:1959:OTR

[Gut59] Irwin Guttman. Optimum tolerance regions and power when sampling from some non-normal universes. Annals of Mathematical Statistics, 30(4):926–938, December 1959. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177706076.

Guthery:1970:IAO

[Gut70a] Scott Bates Guthery. An inversion algorithm for one-dimensional F -expansions. Annals of Mathematical Statistics, 41(5):1472–1490, October 1970. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177696793.

Guttman:1970:CCT

[Gut70b] Irwin Guttman. Construction of β-content tolerance regions at confidence level γ for large samples from the k-variate normal distribution. Annals of Mathematical Statistics, 41(2):376–400, April 1970. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177697078.

Gumbel:1950:DNE

[GvS50] E. J. Gumbel and H. von Schelling. The distribution of the number of exceedances. Annals of Mathematical Statistics, 21(2):247– 262, June 1950. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177729842.

Ghurye:1959:CCM

[GW59a] S. G. Ghurye and David L. Wallace. A convolutive class of monotone likelihood ratio families. Annals of Mathematical Statistics, 30(4): 1158–1164, December 1959. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177706101.

Graybill:1959:CIB

[GW59b] Franklin A. Graybill and David L. Weeks. Combining inter-block and intra-block information in balanced incomplete blocks. Annals of Mathematical Statistics, 30(3):799–805, September 1959. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177706208.

Gupta:1965:SII

[GW65] Shanti S. Gupta and Mrudulla N. Waknis. A system of inequalities for the incomplete gamma function and the normal integral. Annals of Mathematical Statistics, 36(1):139–149, February 1965. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177700278.

Gastwirth:1968:EMO

[GW68] Joseph L. Gastwirth and Stephen S. Wolff. An elementary method for obtaining lower bounds on the asymptotic power of rank tests. Annals of Mathematical Statistics, 39(6):2128–2130, December 1968. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177698043.

Gnanadesikan:1970:UML

[GW70] R. Gnanadesikan and M. B. Wilk. Use of maximum likelihood for estimating error variance from a collection of analysis of variance mean squares. Annals of Mathematical Statistics, 41(1):292–304, February 1970. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177697208.

Gray:1972:JSE

[GWA72] H. L. Gray, T. A. Watkins, and J. E. Adams. On the jackknife statistic, its extensions, and its relation to e_{n}-transformations. Annals of Mathematical Statistics, 43(1):1–30, February 1972. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177692697.

Godwin:1961:CLT

[GZ61] H. J. Godwin and S. K. Zaremba. A Central Limit Theorem for partly dependent variables. Annals of Mathematical Statistics, 32(3):677– 686, September 1961. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177704963.

Goldman:1966:CST

[GZ66] Aaron S. Goldman and R. K. Zeigler. Comparisons of some two stage sampling methods. Annals of Mathematical Statistics, 37(4): 891–897, August 1966. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177699370.

Hultquist:1972:MEM

[HA72] Robert Hultquist and Erwin M. Atzinger. The mixed effects model and simultaneous diagonalization of symmetric matrices. Annals of Mathematical Statistics, 43(6):2024–2030, December 1972. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177690877.

Hachigian:1963:CMC

[Hac63] Jack Hachigian. Collapsed Markov chains and the Chapman– Kolmogorov equation. Annals of Mathematical Statistics, 34(1):233– 237, March 1963. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177704261.

Hadley:1963:RKJ

[Had63] George Hadley. Review: Kenneth J. Arrow, Samuel Karlin, Herbert Scarf, Studies in Applied Probability and Management Science. Annals of Mathematical Statistics, 34(2):674–677, June 1963. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177704184.

Haggstrom:1966:OSE

[Hag66] Gus W. Haggstrom. Optimal stopping and experimental design. Annals of Mathematical Statistics, 37(1):7–29, February 1966. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177699594.

Haggstrom:1967:OSP

[Hag67] Gus W. Haggstrom. Optimal sequential procedures when more than one stop is required. Annals of Mathematical Statistics, 38(6):1618– 1626, December 1967. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177698595.

Haight:1961:EUQ

[Hai61] Frank A. Haight. Expected utility for queues servicing Messages with exponentially decaying utility. Annals of Mathematical Statistics, 32 (2):587–593, June 1961. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177705064.

Hajek:1961:SEW

[Háj61] Jaroslav Hájek. Some extensions of the Wald–Wolfowitz–Noether theorem. Annals of Mathematical Statistics, 32(2):506–523, June 1961. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177705057.

Hajek:1962:AMP

[Háj62] Jaroslav Hájek. Asymptotically most powerful rank-order tests. Annals of Mathematical Statistics, 33(3):1124–1147, September 1962. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177704476.

Hajek:1964:ATR

[Háj64] Jaroslav Hájek. Asymptotic theory of rejective sampling with varying probabilities from a finite population. Annals of Mathematical Statistics, 35(4):1491–1523, December 1964. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177700375.

Hajian:1967:RPB

[Haj67] Arshag Hajian. Review: Patrick Billingsley, Ergodic Theory and Information. Annals of Mathematical Statistics, 38(1):285–286, February 1967. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177699086.

Hajek:1968:ANS

[Háj68] Jaroslav Hájek. Asymptotic normality of simple linear rank statistics under alternatives. Annals of Mathematical Statistics, 39(2):

325–346, April 1968. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177698394.

Halmos:1944:RA

[Hal44] Paul R. Halmos. Random alms. Annals of Mathematical Statistics, 15 (2):182–189, June 1944. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177731283.

Halmos:1946:TUE

[Hal46] Paul R. Halmos. The theory of unbiased estimation. Annals of Mathematical Statistics, 17(1):34–43, March 1946. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177731020.

Halperin:1951:NRT

[Hal51] Max Halperin. Normal regression theory in the presence of intraclass correlation. Annals of Mathematical Statistics, 22(4):573–580, December 1951. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177729547.

Halperin:1952:MLE

[Hal52] Max Halperin. Maximum likelihood estimation in truncated samples. Annals of Mathematical Statistics, 23(2):226–238, June 1952. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177729439.

Hall:1958:MEM

[Hal58] Wm. Jackson Hall. Most economical multiple-decision rules. Annals of Mathematical Statistics, 29(4):1079–1094, December 1958. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177706442.

Hall:1959:MEC

[Hal59] Wm. Jackson Hall. The most-economical character of some Bechhofer and sobel decision rules. Annals of Mathematical Statistics, 30 (4):964–969, December 1959. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177706079.

Halperin:1960:SAR

[Hal60] Max Halperin. Some asymptotic results for a coverage problem. Annals of Mathematical Statistics, 31(4):1063–1076, December 1960. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177705678.

Halperin:1961:CIC

[Hal61] Max Halperin. Confidence intervals from censored samples. Annals of Mathematical Statistics, 32(3):828–837, September 1961. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177704976.

Hammersley:1950:DDH

[Ham50] J. M. Hammersley. The distribution of distance in a hypersphere. Annals of Mathematical Statistics, 21(3):447–452, September 1950. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177729805.

Hamblen:1956:DRQ

[Ham56] John W. Hamblen. Distributions of roots of quadratic equations with random coefficients. Annals of Mathematical Statistics, 27(4):1136– 1143, December 1956. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177728080.

Hammersley:1957:PPL

[Ham57] J. M. Hammersley. Percolation processes: Lower bounds for the critical probability. Annals of Mathematical Statistics, 28(3):790–795, September 1957. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177706894.

Hampel:1971:GQD

[Ham71] Frank R. Hampel. A general qualitative definition of robustness. Annals of Mathematical Statistics, 42(6):1887–1896, December 1971. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177693054.

Hamilton:1972:SAA

[Ham72] Martin A. Hamilton. The stochastic approximation approach to a discrimination problem. Annals of Mathematical Statistics, 43(4): 1096–1109, August 1972. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177692463.

Hand:1943:RFM

[Han43a] Howard Hand. Report of the first meeting of the Pittsburgh chapter of the Institute. Annals of Mathematical Statistics, 14(3):303–304, September 1943. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177731425.

Hand:1943:RSM

[Han43b] Howard Hand. Report on the second meeting of the Pittsburgh chapter of the Institute. Annals of Mathematical Statistics, 14(4):445, December 1943. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177731367.

Hans:1959:MEC

[Han59] Otto Hanš. Measurability of extensions of continuous random transforms. Annals of Mathematical Statistics, 30(4):1152–1157, December 1959. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177706100.

Hanani:1961:ECB

[Han61] Haim Hanani. The existence and construction of balanced incomplete block designs. Annals of Mathematical Statistics, 32(2):361– 386, June 1961. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177705047.

Hanani:1965:BIB

[Han65] Haim Hanani. A balanced incomplete block design. Annals of Mathematical Statistics, 36(2):711, April 1965. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177700182.

Hannan:1966:RDR

[Han66a] E. J. Hannan. Review: D. R. Cox, P. A. W. Lewis, The Statistical Analysis of Series of Events. Annals of Mathematical Statistics, 37 (6):1852–1853, December 1966. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177699183.

Hanson:1966:RKK

[Han66b] D. L. Hanson. Review: Klaus Krickeberg, Probability Theory. Annals of Mathematical Statistics, 37(3):755–756, June 1966. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177699480.

Hanson:1967:SRR

[Han67] D. L. Hanson. Some results relating moment generating functions and convergence rates in the law of large numbers. Annals of Mathematical Statistics, 38(3):742–750, June 1967. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177698867.

Hanurav:1968:HAO

[Han68] T. V. Hanurav. Hyper-admissibility and optimum estimators for sampling finite populations. Annals of Mathematical Statistics, 39(2): 621–642, April 1968. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177698422.

Haq:1970:SAF

[Haq70] M. Safiul Haq. Structural analysis for the first order autoregressive stochastic models. Annals of Mathematical Statistics, 41(3): 970–978, June 1970. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177696973.

Hart:1942:SLR

[Har42] B. I. Hart. Significance levels for the ratio of the mean square successive difference to the variance. Annals of Mathematical Statistics, 13(4):445–447, December 1942. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177731544.

Harshbarger:1944:ACS

[Har44] Boyd Harshbarger. On the analysis of a certain six-by-six fourgroup lattice design. Annals of Mathematical Statistics, 15(3):307– 320, September 1944. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177731237.

Harshbarger:1945:ACS

[Har45] Boyd Harshbarger. On the analysis of a certain six-by-six four-group lattice design using the recovery of inter-block information. Annals of Mathematical Statistics, 16(4):387–390, December 1945. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177731065.

Harris:1947:NDU

[Har47] T. E. Harris. Note on differentiation under the expectation sign in the fundamental identity of sequential analysis. Annals of Mathematical Statistics, 18(2):294–295, June 1947. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177730449.

Harris:1948:BP

[Har48] T. E. Harris. Branching processes. Annals of Mathematical Statistics, 19(4):474–494, December 1948. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177730146.

Harter:1951:DWC

[Har51] Harman Leon Harter. On the distribution of Wald’s classification statistic. Annals of Mathematical Statistics, 22(1):58–67, March 1951. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177729692.

Harris:1952:LCD

[Har52] Lee B. Harris. On a limiting case for the distribution of exceedances, with an application to life-testing. Annals of Mathematical Statistics, 23(2):295–298, June 1952. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177729450.

Harris:1959:DBI

[Har59a] Bernard Harris. Determining bounds on integrals with applications to cataloging problems. Annals of Mathematical Statistics, 30(2): 521–548, June 1959. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177706266.

Harter:1959:USQ

[Har59b] H. Leon Harter. The use of sample quasi-ranges in estimating population standard deviation. Annals of Mathematical Statistics, 30 (4):980–999, December 1959. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177706081.

Harris:1960:PDR

[Har60a] Bernard Harris. Probability distributions related to random mappings. Annals of Mathematical Statistics, 31(4):1045–1062, December 1960. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177705677.

Harter:1960:CNC

[Har60b] H. Leon Harter. Correction notes: Correction to “The Use of Sample Quasi-Ranges in Estimating Population Standard Deviation”. Annals of Mathematical Statistics, 31(1):228, March 1960. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177706003.

Harter:1960:TRS

[Har60c] H. Leon Harter. Tables of range and Studentized range. Annals of Mathematical Statistics, 31(4):1122–1147, December 1960. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177705684.

Harter:1961:EPN

[Har61] H. Leon Harter. Estimating the parameters of negative exponential populations from one or two order statistics. Annals of Mathematical Statistics, 32(4):1078–1090, December 1961. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http:

//projecteuclid.org/euclid.aoms/1177704847. See correction

[Har64a].

Harris:1962:DBE

[Har62] Bernard Harris. Determining bounds on expected values of certain functions. Annals of Mathematical Statistics, 33(4):1454–1457, December 1962. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177704377.

Harter:1964:CEP

[Har64a] H. Leon Harter. Correction to “Estimating the Parameters of Negative Exponential Population from One or Two Order Statistics”. Annals of Mathematical Statistics, 35(2):925, June 1964. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177703610. See [Har61].

Hartigan:1964:IPD

[Har64b] J. Hartigan. Invariant prior distributions. Annals of Mathematical Statistics, 35(2):836–845, June 1964. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177703583.

Harkness:1965:PEH

[Har65a] W. L. Harkness. Properties of the extended hypergeometric distribution. Annals of Mathematical Statistics, 36(3):938–945, June 1965. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177700066.

Hartigan:1965:AUP

[Har65b] J. A. Hartigan. The asymptotically unbiased prior distribution. Annals of Mathematical Statistics, 36(4):1137–1152, August 1965. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177699988.

Harper:1966:FCI

[Har66a] Lawrence H. Harper. A family of combinatorial identities. Annals of Mathematical Statistics, 37(2):509–512, April 1966. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177699535.

Hartigan:1966:PCK

[Har66b] J. A. Hartigan. Probabilistic completion of a knockout tournament. Annals of Mathematical Statistics, 37(2):495–503, April 1966. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177699533.

Harper:1967:SBA

[Har67a] L. H. Harper. Stirling behavior is asymptotically normal. Annals of Mathematical Statistics, 38(2):410–414, April 1967. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177698956.

Harville:1967:EVC

[Har67b] David A. Harville. Estimability of variance components for the twoway classification with interaction. Annals of Mathematical Statistics, 38(5):1508–1519, October 1967. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177698705.

Hartigan:1968:IKT

[Har68] J. A. Hartigan. Inference from a knockout tournament. Annals of Mathematical Statistics, 39(2):583–592, April 1968. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177698419.

Harville:1969:EVC

[Har69a] David A. Harville. Expression of variance-component estimators as linear combinations of independent noncentral chi-square variates. Annals of Mathematical Statistics, 40(6):2189–2194, December 1969. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177697297.

Harville:1969:VVC

[Har69b] David A. Harville. Variances of variance-component estimators for the unbalanced 2-way cross classification with application to balanced incomplete block designs. Annals of Mathematical Statistics, 40(2): 408–416, April 1969. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177697705.

Harris:1970:AEV

[Har70a] Robert Harris. An application of extreme value theory to reliability theory. Annals of Mathematical Statistics, 41(5):1456–1465, October 1970. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177696791.

Harris:1970:MDI

[Har70b] Robert Harris. A multivariate definition for increasing hazard rate distribution functions. Annals of Mathematical Statistics, 41(2):713– 717, April 1970. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177697121.

Hartigan:1970:ECI

[Har70c] J. A. Hartigan. Exact confidence intervals in regression problems with independent symmetric errors. Annals of Mathematical Statistics, 41 (6):1992–1998, December 1970. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177696700.

Hartley:1971:RWJ

[Har71a] H. O. Hartley. Review: William J. Hemmerle, Statistical Computations on a Digital Computer. Annals of Mathematical Statistics, 42 (4):1468, August 1971. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177693266.

Harville:1971:DLC

[Har71b] David A. Harville. On the distribution of linear combinations of non-central chi-squares. Annals of Mathematical Statistics, 42(2): 809–811, April 1971. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177693437.

Harrison:1972:DDP

[Har72] J. Michael Harrison. Discrete dynamic programming with unbounded rewards. Annals of Mathematical Statistics, 43(2):636–644, April 1972. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177692643.

Hasel:1942:EVT

[Has42] A. A. Hasel. Estimation of volume in timber stands by strip sampling. Annals of Mathematical Statistics, 13(2):179–206, June 1942. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177731605.

Hasofer:1963:ICD

[Has63] A. M. Hasofer. On the integrability, continuity and differentiability of a family of functions introduced by L. Takacs. Annals of Mathematical Statistics, 34(3):1045–1049, September 1963. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177704028.

Hatke:1949:CCP

[Hat49] Mary Agnes Hatke. A certain cumulative probability function. Annals of Mathematical Statistics, 20(3):461–463, September 1949. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177730002.

Housner:1948:ELT

[HB48] G. W. Housner and J. F. Brennan. The estimation of linear trends. Annals of Mathematical Statistics, 19(3):380–388, September 1948. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177730202.

Hartley:1965:NLS

[HB65] H. O. Hartley and Aaron Booker. Nonlinear least squares estimation. Annals of Mathematical Statistics, 36(2):638–650, April 1965. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177700171.

Hora:1966:FTI

[HB66] R. B. Hora and R. J. Buehler. Fiducial theory and invariant estimation. Annals of Mathematical Statistics, 37(3):643–656, June 1966. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177699458.

Hora:1967:FTI

[HB67] R. B. Hora and R. J. Buehler. Fiducial theory and invariant prediction. Annals of Mathematical Statistics, 38(3):795–801, June 1967. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177698873.

Heyde:1970:DNC

[HB70] C. C. Heyde and B. M. Brown. On the departure from normality of a certain class of martingales. Annals of Mathematical Statistics, 41 (6):2161–2165, December 1970. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177696722.

Hendrickson:1971:PSP

[HB71] Arlo D. Hendrickson and Robert J. Buehler. Proper scores for probability forecasters. Annals of Mathematical Statistics, 42(6):1916– 1921, December 1971. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177693057.

Hogg:1958:DCV

[HC58] Robert V. Hogg and Allen T. Craig. On the decomposition of certain χ^{2} variables. Annals of Mathematical Statistics, 29(2):608– 610, June 1958. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177706644.

Hellman:1970:LFM

[HC70] Martin E. Hellman and Thomas M. Cover. Learning with finite memory. Annals of Mathematical Statistics, 41(3):765–782, June 1970. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177696958.

Hellman:1971:MSR

[HC71] Martin E. Hellman and Thomas M. Cover. On memory saved by randomization. Annals of Mathematical Statistics, 42(3):1075–1078, June 1971. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177693334.

Hartley:1954:UBM

[HD54] H. O. Hartley and H. A. David. Universal bounds for mean range and extreme observation. Annals of Mathematical Statistics, 25(1): 85–99, March 1954. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177728848.

Hill:1968:GAE

[HD68] G. W. Hill and A. W. Davis. Generalized asymptotic expansions of Cornish–Fisher type. Annals of Mathematical Statistics, 39(4):1264– 1273, August 1968. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177698251.

Healy:1956:TSP

[Hea56] W. C. Healy, Jr. Two-sample procedures in simultaneous estimation. Annals of Mathematical Statistics, 27(3):687–702, September 1956. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177728176.

Healy:1961:LVC

[Hea61a] W. C. Healy, Jr. Limits for a variance component with an exact confidence coefficient. Annals of Mathematical Statistics, 32(2):466– 476, June 1961. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177705053.

Heathcote:1961:QPM

[Hea61b] C. R. Heathcote. On the queueing process M/G/1. Annals of Mathematical Statistics, 32(3):770–773, September 1961. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177704972.

Heck:1960:CSU

[Hec60] D. L. Heck. Charts of some upper percentage points of the distribution of the largest characteristic root. Annals of Mathematical Statistics, 31(3):625–642, September 1960. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177705790.

Hedayat:1970:RVP

[Hed70] A. Hedayat. Review: S. Vajda, Patterns and Configurations in Finite Spaces. Number 22 of Griffins Statistical Monographs and Courses; S. Vajda, The Mathematics of Experimental Design. Number 23 of Griffins Statistical Monographs and Courses. Annals of Mathematical Statistics, 41(5):1780–1782, October 1970. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177696826.

Heller:1965:SPD

[Hel65] Alex Heller. On stochastic processes derived from Markov chains. Annals of Mathematical Statistics, 36(4):1286–1291, August 1965. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177700000.

Hellman:1970:ESU

[Hel70] Olavi Hellman. On the effect of a search upon the probability distribution of a target whose motion is a diffusion process. Annals of Mathematical Statistics, 41(5):1717–1724, October 1970. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177696816.

Hemelrijk:1952:NWT

[Hem52] J. Hemelrijk. Note on Wilcoxon’s two-sample test when ties are present. Annals of Mathematical Statistics, 23(1):133–135, March 1952. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177729491.

Hemelrijk:1960:SWD

[Hem60] J. Hemelrijk. The statistical work of David Van Dantzig (1900–1959). Annals of Mathematical Statistics, 31(2):269–275, June 1960. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177705893.

Hendricks:1931:URR

[Hen31] Walter A. Hendricks. The use of the relative residual in the application of the method of least squares. Annals of Mathematical Statistics, 2(4):458–478, November 1931. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177732952.

Henderson:1932:PO

[Hen32a] R. Henderson. A postulate for observations. Annals of Mathematical Statistics, 3(1):32–39, February 1932. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177732930.

Hendricks:1932:RRC

[Hen32b] Walter A. Hendricks. Relative residuals considered as weighted simple residuals in the application of the method of least squares. Annals of Mathematical Statistics, 3(3):157–162, August 1932. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177732884.

Hendricks:1934:SEA

[Hen34] Walter A. Hendricks. The standard error of any analytic function of a set of parameters evaluated by the method of least squares. Annals of Mathematical Statistics, 5(2):107–112, June 1934. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177732709.

Hendricks:1935:AVC

[Hen35a] Walter A. Hendricks. Analysis of variance considered as an application of simple error theory. Annals of Mathematical Statistics, 6 (3):117–126, September 1935. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177732586.

Hendricks:1935:PIL

[Hen35b] Walter A. Hendricks. A problem involving the Lexis theory of dispersion. Annals of Mathematical Statistics, 6(2):78–82, June 1935.

CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177732611.

Hendricks:1936:ASD

[Hen36] Walter A. Hendricks. An approximation to “Student’s” distribution. Annals of Mathematical Statistics, 7(4):210–221, December 1936. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177732482.

Hendricks:1972:HDP

[Hen72] W. J. Hendricks. Hausdorff dimension in a process with stable components — an interesting counterexample. Annals of Mathematical Statistics, 43(2):690–694, April 1972. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177692657.

Herbach:1948:BSF

[Her48] Leon H. Herbach. Bounds for some functions used in sequentially testing the mean of a Poisson distribution. Annals of Mathematical Statistics, 19(3):400–405, September 1948. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177730205.

Herbach:1959:PMI

[Her59] Leon H. Herbach. Properties of Model II-type analysis of variance tests, A: Optimum nature of the F -test for Model II in the balanced case. Annals of Mathematical Statistics, 30(4):939–959, December 1959. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177706077.

Herbst:1963:APV

[Her63] Laurence Herbst. Almost periodic variances. Annals of Mathematical Statistics, 34(4):1549–1557, December 1963. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177703887.

Herzberg:1964:TTO

[Her64] Agnes M. Herzberg. Two third order rotatable designs in four dimensions. Annals of Mathematical Statistics, 35(1):445–446, March 1964. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177703774.

Herzberg:1966:CRD

[Her66] Agnes M. Herzberg. Cylindrically rotatable designs. Annals of Mathematical Statistics, 37(1):242–247, February 1966. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177699613.

Herr:1967:AOT

[Her67a] David G. Herr. Asymptotically optimal tests for multivariate normal distributions. Annals of Mathematical Statistics, 38(6):1829–1844, December 1967. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177698616.

Herzberg:1967:CRD

[Her67b] Agnes M. Herzberg. Cylindrically rotatable designs of Types 1, 2 and 3. Annals of Mathematical Statistics, 38(1):167–176, February 1967. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177699067. See

correction [Her70].

Herzberg:1967:MCS

[Her67c] Agnes M. Herzberg. A method for the construction of second order rotatable designs in k dimensions. Annals of Mathematical Statistics, 38(1):177–180, February 1967. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177699068.

Hertz:1969:CRC

[Her69] Ellen S. Hertz. On convergence rates in the Central Limit Theorem. Annals of Mathematical Statistics, 40(2):475–479, April 1969. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177697713.

Herzberg:1970:CNC

[Her70] Agnes M. Herzberg. Correction notes: Correction to “Cylindrically Rotatable Designs of Types 1, 2 and 3”. Annals of Mathematical Statistics, 41(1):326–327, February 1970. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177697215. See [Her67b].

Hern:1972:EEW

[Her72] Thomas A. Hern. Error estimates for the weak convergence to certain infinitely divisible laws. Annals of Mathematical Statistics, 43 (5):1592–1602, October 1972. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177692391.

Hettmansperger:1968:TMW

[Het68] Thomas P. Hettmansperger. On the trimmed Mann–Whitney statistic. Annals of Mathematical Statistics, 39(5):1610–1614, October 1968. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177698142.

Heyde:1966:SRT

[Hey66] C. C. Heyde. Some renewal theorems with application to a first passage problem. Annals of Mathematical Statistics, 37(3):699–710, June 1966. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177699465.

Heyde:1967:LDP

[Hey67] C. C. Heyde. On large deviation problems for sums of random variables which are not attracted to the normal law. Annals of Mathematical Statistics, 38(5):1575–1578, October 1967. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177698712.

Heyde:1968:VRT

[Hey68] C. C. Heyde. Variations on a renewal theorem of Smith. Annals of Mathematical Statistics, 39(1):155–158, February 1968. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177698512.

Heyde:1969:FTP

[Hey69a] C. C. Heyde. On a fluctuation theorem for processes with independent increments II. Annals of Mathematical Statistics, 40(2):688– 691, April 1969. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177697742.

Heyde:1969:ERC

[Hey69b] C. C. Heyde. On extended rate of convergence results for the invariance principle. Annals of Mathematical Statistics, 40(6):2178–2179, December 1969. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177697294.

Heyde:1970:ERS

[Hey70] C. C. Heyde. Extension of a result of Seneta for the super-critical Galton–Watson process. Annals of Mathematical Statistics, 41(2): 739–742, April 1970. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177697127. See [Sen68e].

Hotelling:1938:TSS

[HF38] Harold Hotelling and Lester R. Frankel. The transformation of statistics to simplify their distribution. Annals of Mathematical Statistics, 9(2):87–96, June 1938. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177732330.

Hedayat:1970:EMG

[HF70] A. Hedayat and W. T. Federer. On the equivalence of Mann’s group automorphism method of constructing an O(n, n−1) set and Raktoe’s collineation method of constructing a balanced set of L-restrictional prime-powered lattice designs. Annals of Mathematical Statistics, 41 (5):1530–1540, October 1970. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177696797.

Hedayat:1971:EEO

[HF71] A. Hedayat and W. T. Federer. On embedding and enumeration of orthogonal latin squares. Annals of Mathematical Statistics, 42(2): 509–516, April 1971. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177693401.

Haynam:1966:EPM

[HG66] George E. Haynam and Zakkula Govindarajulu. Exact power of Mann–Whitney test for exponential and rectangular alternatives. Annals of Mathematical Statistics, 37(4):945–953, August 1966. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177699375.

Hansen:1943:TSF

[HH43] Morris H. Hansen and William N. Hurwitz. On the theory of sampling from finite populations. Annals of Mathematical Statistics, 14(4):333– 362, December 1943. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177731356.

Hansen:1949:DOP

[HH49] Morris H. Hansen and William N. Hurwitz. On the determination of optimum probabilities in sampling. Annals of Mathematical Statistics, 20(3):426–432, September 1949. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177729994.

Hannan:1963:NAD

[HH63] J. Hannan and W. Harkness. Normal approximation to the distribution of two independent binomials, conditional on fixed sum. Annals of Mathematical Statistics, 34(4):1593–1595, December 1963. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177703893.

Hannan:1972:LTQ

[HH72a] E. J. Hannan and C. C. Heyde. On limit theorems for quadratic functions of discrete time series. Annals of Mathematical Statistics, 43(6):2058–2066, December 1972. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177690885.

Hannan:1972:EPC

[HH72b] James Hannan and J. S. Huang. Equivariant procedures in the compound decision problem with finite state component problem. Annals of Mathematical Statistics, 43(1):102–112, February 1972. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177692706.

Hannan:1972:SSP

[HH72c] James Hannan and J. S. Huang. A stability of symmetrization of product measures with few distinct factors. Annals of Mathematical Statistics, 43(1):308–319, February 1972. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177692724.

Hildebrandt:1931:SPC

[Hil31] Emanuel Henry Hildebrandt. Systems of polynomials connected with the Charlier expansions and the Pearson differential and difference equations. Annals of Mathematical Statistics, 2(4):379–439, November 1931. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177732950.

Hill:1960:RBH

[Hil60] Bruce M. Hill. A relationship between Hodges’ bivariate sign test and a non-parametric test of Daniels. Annals of Mathematical Statistics, 31(4):1190–1192, December 1960. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177705689. See correction [Hil61].

Hill:1961:CNC

[Hil61] Bruce M. Hill. Correction notes: Corrections to “A Relationship Between Hodges’ Bivariate Sign Test and a Non-Parametric Test of Daniels”. Annals of Mathematical Statistics, 32(2):619, June 1961. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177705073. See

[Hil60].

Hill:1962:TLV

[Hil62] Bruce Marvin Hill. A test of linearity versus convexity of a median regression curve. Annals of Mathematical Statistics, 33(3):1096–1123, September 1962. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177704475.

Hildebrand:1969:DOT

[Hil69a] David K. Hildebrand. Domains of optimality of tests in simple random sampling. Annals of Mathematical Statistics, 40(1):308–312, February 1969. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177697827.

Hildreth:1969:ADM

[Hil69b] Clifford Hildreth. Asymptotic distribution of maximum likelihood estimators in a linear model with autoregressive disturbances. Annals of Mathematical Statistics, 40(2):583–594, April 1969. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177697726.

Hinkelmann:1964:EGD

[Hin64] Klaus Hinkelmann. Extended group divisible partially balanced incomplete block designs. Annals of Mathematical Statistics, 35(2): 681–695, June 1964. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177703565.

Hinich:1965:LSE

[Hin65] Melvin Hinich. Large-sample estimation of an unknown discrete waveform which is randomly repeating in Gaussian noise. Annals of Mathematical Statistics, 36(2):489–508, April 1965. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177700159.

Hurwitz:1944:SAC

[HK44] H. Hurwitz, Jr. and M. Kac. Statistical analysis of certain types of random functions. Annals of Mathematical Statistics, 15(2):173– 181, June 1944. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177731282.

Hanson:1964:TLC

[HK64a] D. L. Hanson and L. H. Koopmans. Tolerance limits for the class of distributions with increasing hazard rates. Annals of Mathematical Statistics, 35(4):1561–1570, December 1964. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177700380.

Harkness:1964:CPF

[HK64b] W. L. Harkness and L. Katz. Comparison of the power functions for the test of independence in 2 × 2 contingency tables. Annals of Mathematical Statistics, 35(3):1115–1127, September 1964. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177703269.

Hanson:1965:CRLb

[HK65a] D. L. Hanson and L. H. Koopmans. Convergence rates for the law of large numbers for linear combinations of exchangeable and ^{∗}-mixing stochastic processes. Annals of Mathematical Statistics, 36(6):1840– 1852, December 1965. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177699814.

Hanson:1965:CRLa

[HK65b] D. L. Hanson and L. H. Koopmans. On the convergence rate of the law of large numbers for linear combinations of independent random variables. Annals of Mathematical Statistics, 36(2):559–564, April 1965. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177700167.

Hall:1968:STG

[HK68] Irving J. Hall and Akio Kudô. On slippage tests —- I. A generalization of Neyman–Pearson’s lemma. Annals of Mathematical Statistics, 39 (5):1693–1699, October 1968. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177698151.

Hall:1968:STI

[HKcY68] Irving J. Hall, Akio Kudô, and Neng che Yeh. On slippage tests. II. Similar slippage tests. Annals of Mathematical Statistics, 39(6):2029– 2037, December 1968. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177698030.

Hodges:1950:SPM

[HL50] J. L. Hodges, Jr. and E. L. Lehmann. Some problems in minimax point estimation. Annals of Mathematical Statistics, 21(2): 182–197, June 1950. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177729838.

Hodges:1952:UPE

[HL52] J. L. Hodges, Jr. and E. L. Lehmann. The use of previous experience in reaching statistical decisions. Annals of Mathematical Statistics, 23(3):396–407, September 1952. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177729384.

Hodges:1954:MPC

[HL54] J. L. Hodges, Jr. and E. L. Lehmann. Matching in paired comparisons. Annals of Mathematical Statistics, 25(4):787–791, December 1954. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177728667.

Hodges:1956:ESN

[HL56] J. L. Hodges, Jr. and E. L. Lehmann. The efficiency of some nonparametric competitors of the t-test. Annals of Mathematical Statistics, 27 (2):324–335, June 1956. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177728261.

Hodges:1960:PAP

[HL60] J. L. Hodges, Jr. and Lucien Le Cam. The Poisson approximation to the Poisson binomial distribution. Annals of Mathematical Statistics, 31(3):737–740, September 1960. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177705799.

Hodges:1962:RMC

[HL62] J. L. Hodges, Jr. and E. L. Lehmann. Rank methods for combination of independent experiments in analysis of variance. Annals of Mathematical Statistics, 33(2):482–497, June 1962. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177704575.

Hodges:1963:ELB

[HL63] J. L. Hodges, Jr. and E. L. Lehmann. Estimates of location based on rank tests. Annals of Mathematical Statistics, 34(2):598–611, June 1963. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177704172.

Hoel:1964:OSW

[HL64] P. G. Hoel and A. Levine. Optimal spacing and weighting in polynomial prediction. Annals of Mathematical Statistics, 35(4):1553–1560, December 1964. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177700379.

Hodges:1968:CTP

[HL68] J. L. Hodges, Jr. and E. L. Lehmann. A compact table for power of the t-test. Annals of Mathematical Statistics, 39(5):1629–1637, October 1968. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177698144.

Hodges:1970:D

[HL70] J. L. Hodges, Jr. and E. L. Lehmann. Deficiency. Annals of Mathematical Statistics, 41(3):783–801, June 1970. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177696959.

Harter:1967:AVC

[HM67] H. Leon Harter and Albert H. Moore. Asymptotic variances and covariances of maximum-likelihood estimators, from censored samples, of the parameters of Weibull and gamma populations. Annals of Mathematical Statistics, 38(2):557–570, April 1967. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177698971.

Hoel:1968:EPS

[HM68] D. G. Hoel and M. Mazumdar. An extension of Paulson’s selection procedure. Annals of Mathematical Statistics, 39(6):2067–2074, December 1968. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177698034.

Hastings:1947:LMS

[HMTW47] Cecil Hastings, Jr., Frederick Mosteller, John W. Tukey, and Charles P. Winsor. Low moments for small samples: a comparative study of order statistics. Annals of Mathematical Statistics, 18 (3):413–426, September 1947. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177730388.

Hoadley:1967:PLD

[Hoa67] A. Bruce Hoadley. On the probability of large deviations of functions of several empirical CDF’s. Annals of Mathematical Statistics, 38(2): 360–381, April 1967. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177698952.

Hoadley:1971:APM

[Hoa71] Bruce Hoadley. Asymptotic properties of maximum likelihood estimators for the independent not identically distributed case. Annals of Mathematical Statistics, 42(6):1977–1991, December 1971. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177693066.

Hodges:1949:RBM

[Hod49] J. L. Hodges. Report on the Berkeley meeting of the Institute. Annals of Mathematical Statistics, 20(3):475–476, September 1949. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177730005.

Hodges:1950:RBM

[Hod50] J. L. Hodges. Report of the Berkeley meeting of the Institute. Annals of Mathematical Statistics, 21(4):622–624, December 1950. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177729762.

Hodges:1955:BST

[Hod55a] J. L. Hodges, Jr. A bivariate sign test. Annals of Mathematical Statistics, 26(3):523–527, September 1955. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177728498.

Hodges:1955:NBD

[Hod55b] J. L. Hodges, Jr. On the noncentral beta-distribution. Annals of Mathematical Statistics, 26(4):648–653, December 1955. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177728424.

Hoel:1937:STC

[Hoe37] Paul G. Hoel. A significance test for component analysis. Annals of Mathematical Statistics, 8(3):149–158, September 1937. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177732411.

Hoel:1938:CSD

[Hoe38] Paul G. Hoel. On the chi-square distribution for small samples. Annals of Mathematical Statistics, 9(3):158–165, September 1938. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177732307.

Hoel:1940:EIE

[Hoe40] Paul G. Hoel. The errors involved in evaluating correlation determinants. Annals of Mathematical Statistics, 11(1):58–65, March 1940. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177731941.

Hoel:1941:MSN

[Hoe41] Paul G. Hoel. On methods of solving normal equations. Annals of Mathematical Statistics, 12(3):354–359, September 1941. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177731719.

Hoel:1943:ASM

[Hoe43a] Paul G. Hoel. The accuracy of sampling methods in ecology. Annals of Mathematical Statistics, 14(3):289–300, September 1943. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177731422.

Hoel:1943:ID

[Hoe43b] Paul G. Hoel. On indices of dispersion. Annals of Mathematical Statistics, 14(2):155–162, June 1943. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177731457.

Hoel:1945:THP

[Hoe45] Paul G. Hoel. Testing the homogeneity of Poisson frequencies. Annals of Mathematical Statistics, 16(4):362–368, December 1945. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177731062.

Hoel:1946:EMM

[Hoe46] Paul G. Hoel. The efficiency of the mean moving range. Annals of Mathematical Statistics, 17(4):475–482, December 1946. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177730886.

Hoel:1947:DBB

[Hoe47] Paul G. Hoel. Discriminating between binomial distributions. Annals of Mathematical Statistics, 18(4):556–564, December 1947. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177730346.

Hoeffding:1948:CSA

[Hoe48a] Wassily Hoeffding. A class of statistics with asymptotically normal distribution. Annals of Mathematical Statistics, 19(3):293–325, September 1948. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177730196.

Hoeffding:1948:NPT

[Hoe48b] Wassily Hoeffding. A non-parametric test of independence. Annals of Mathematical Statistics, 19(4):546–557, December 1948. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177730150.

Hoel:1948:USR

[Hoe48c] Paul G. Hoel. On the uniqueness of similar regions. Annals of Mathematical Statistics, 19(1):66–71, March 1948. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177730291.

Hoeffding:1951:CCL

[Hoe51a] Wassily Hoeffding. A combinatorial Central Limit Theorem. Annals of Mathematical Statistics, 22(4):558–566, December 1951. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177729545.

Hoel:1951:CEE

[Hoe51b] Paul G. Hoel. Conditional expectation and the efficiency of estimates. Annals of Mathematical Statistics, 22(2):299–301, June 1951. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177729651.

Hoeffding:1952:LSP

[Hoe52] Wassily Hoeffding. The large-sample power of tests based on permutations of observations. Annals of Mathematical Statistics, 23(2): 169–192, June 1952. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177729436.

Hoeffding:1953:LBA

[Hoe53a] Wassily Hoeffding. A lower bound for the average sample number of a sequential test. Annals of Mathematical Statistics, 24(1):127– 130, March 1953. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177729092.

Hoeffding:1953:DEV

[Hoe53b] Wassily Hoeffding. On the distribution of the expected values of the order statistics. Annals of Mathematical Statistics, 24(1):93–100, March 1953. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177729086.

Hoel:1954:CBP

[Hoe54] Paul G. Hoel. Confidence bands for polynomial curves. Annals of Mathematical Statistics, 25(3):534–542, September 1954. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177728721.

Hoeffding:1955:EEV

[Hoe55a] Wassily Hoeffding. The extrema of the expected value of a function of independent random variables. Annals of Mathematical Statistics, 26 (2):268–275, June 1955. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177728543.

Hoel:1955:STG

[Hoe55b] Paul G. Hoel. On a sequential test for the general linear hypothesis. Annals of Mathematical Statistics, 26(1):136–139, March 1955. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177728603.

Hoeffding:1956:DNS

[Hoe56] Wassily Hoeffding. On the distribution of the number of successes in independent trials. Annals of Mathematical Statistics, 27(3):713– 721, September 1956. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177728178.

Hoel:1958:EPP

[Hoe58] Paul G. Hoel. Efficiency problems in polynomial estimation. Annals of Mathematical Statistics, 29(4):1134–1145, December 1958. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177706446.

Hoeffding:1960:LBE

[Hoe60] Wassily Hoeffding. Lower bounds for the expected sample size and the average risk of a sequential procedure. Annals of Mathematical Statistics, 31(2):352–368, June 1960. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177705898.

Hoel:1961:AEP

[Hoe61] Paul G. Hoel. Asymptotic efficiency in polynomial estimation. Annals of Mathematical Statistics, 32(4):1042–1047, December 1961. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177704844.

Hoeffding:1962:RSW

[Hoe62] Wassily Hoeffding. Review: Samuel S. Wilks, Mathematical Statistics. Annals of Mathematical Statistics, 33(4):1467–1473, December 1962. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177704383.

Hoeffding:1965:AOT

[Hoe65a] Wassily Hoeffding. Asymptotically optimal tests for multinomial distributions. Annals of Mathematical Statistics, 36(2):369–401, April 1965. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/

1177700150. See discussion [NCC65].

Hoel:1965:MDT

[Hoe65b] Paul G. Hoel. Minimax designs in two dimensional regression. Annals of Mathematical Statistics, 36(4):1097–1106, August 1965. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177699984.

Hoel:1965:ODP

[Hoe65c] Paul G. Hoel. Optimum designs for polynomial extrapolation. Annals of Mathematical Statistics, 36(5):1483–1493, October 1965. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177699907.

Hoel:1966:SSO

[Hoe66] Paul G. Hoel. A simple solution for optimal Chebyshev regression extrapolation. Annals of Mathematical Statistics, 37(3):720– 725, June 1966. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177699467.

Hoeffding:1968:RJH

[Hoe68] Wassily Hoeffding. Review: Jaroslav Hajek, Zbynek Sidak, Theory of Rank Tests. Annals of Mathematical Statistics, 39(5):1765–1768, October 1968. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177698166.

Hoem:1969:SDE

[Hoe69] Jan M. Hoem. The sampling distribution of an estimator arising in connection with the truncated exponential distribution. Annals of Mathematical Statistics, 40(2):702–703, April 1969. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177697746.

Hoel:1970:MOS

[Hoe70] David G. Hoel. On the monotonicity of the OC of an SPRT. Annals of Mathematical Statistics, 41(1):310–314, February 1970. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177697210.

Hoel:1971:MCS

[Hoe71] David G. Hoel. A method for the construction of sequential selection procedures. Annals of Mathematical Statistics, 42(2):630– 642, April 1971. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177693413.

Hoffman:1960:UTA

[Hof60] A. J. Hoffman. On the uniqueness of the triangular association scheme. Annals of Mathematical Statistics, 31(2):492–497, June 1960. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177705914.

Hoffman:1963:DSP

[Hof63] A. J. Hoffman. On the duals of symmetric partially-balanced incomplete block designs. Annals of Mathematical Statistics, 34(2): 528–531, June 1963. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177704164.

Hoffman:1964:LGC

[Hof64] A. J. Hoffman. On the line graph of the complete bipartite graph. Annals of Mathematical Statistics, 35(2):883–885, June 1964. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177703593.

Hogg:1951:RCA

[Hog51] Robert V. Hogg. On ratios of certain algebraic forms. Annals of Mathematical Statistics, 22(4):567–572, December 1951. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177729546.

Hogg:1956:DLR

[Hog56] Robert V. Hogg. On the distribution of the likelihood ratio. Annals of Mathematical Statistics, 27(2):529–532, June 1956. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177728276.

Hogg:1963:ICW

[Hog63] Robert V. Hogg. On the independence of certain Wishart variables. Annals of Mathematical Statistics, 34(3):935–939, September 1963. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177704015.

Hoglund:1972:ANS

[Hög72] Thomas Höglund. Asymptotic normality of sums of minima of random variables. Annals of Mathematical Statistics, 43(1):351–353, February 1972. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177692730.

Hollander:1966:ADF

[Hol66] Myles Hollander. An asymptotically distribution-free multiple comparison procedure — treatments vs. control. Annals of Mathematical Statistics, 37(3):735–738, June 1966. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177699472.

Hollander:1967:RTR

[Hol67] Myles Hollander. Rank tests for randomized blocks when the alternatives have an a priori ordering. Annals of Mathematical Statistics, 38 (3):867–877, June 1967. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177698880.

Holland:1968:SPA

[Hol68] Paul W. Holland. Some properties of an algebraic representation of stochastic processes. Annals of Mathematical Statistics, 39(1):164– 170, February 1968. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177698514.

Holewijn:1969:NWC

[Hol69] P. J. Holewijn. Note on Weyl’s criterion and the uniform distribution of independent random variables. Annals of Mathematical Statistics, 40(3):1124–1125, June 1969. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177697624.

Holm:1971:NAM

[Hol71a] Sture Holm. On nonparametric asymptotically minimax tests. Annals of Mathematical Statistics, 42(2):497–508, April 1971. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177693400.

Holst:1971:LTS

[Hol71b] Lars Holst. Limit theorems for some occupancy and sequential occupancy problems. Annals of Mathematical Statistics, 42(5):1671– 1680, October 1971. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177693165.

Holley:1972:MIP

[Hol72] Richard Holley. Markovian interaction processes with finite range interactions. Annals of Mathematical Statistics, 43(6):1961–1967, December 1972. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177690867.

Hooke:1956:SAB

[Hoo56a] Robert Hooke. Some applications of bipolykays to the estimation of variance components and their moments. Annals of Mathematical Statistics, 27(1):80–98, March 1956. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177728351.

Hooke:1956:SFT

[Hoo56b] Robert Hooke. Symmetric functions of a two-way array. Annals of Mathematical Statistics, 27(1):55–79, March 1956. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177728350.

Horst:1932:SMS

[Hor32] Paul Horst. A short method for solving for a coefficient of multiple correlation. Annals of Mathematical Statistics, 3(1):40–44, February 1932. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177732931.

Horst:1935:MDC

[Hor35] Paul Horst. A method for determining the coefficients of a characteristic equation. Annals of Mathematical Statistics, 6(2):83–84, June 1935. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177732612.

Horton:1948:MOR

[Hor48] H. Burke Horton. A method for obtaining random numbers. Annals of Mathematical Statistics, 19(1):81–85, March 1948. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177730294;   http:
//www.jstor.org/stable/2236060.

Horowitz:1971:NAS

[Hor71] Joseph Horowitz. A note on the arc-sine law and Markov random sets. Annals of Mathematical Statistics, 42(3):1068–1074, June 1971. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177693333.

Hotelling:1931:GSR

[Hot31] Harold Hotelling. The generalization of Student’s ratio. Annals of Mathematical Statistics, 2(3):360–378, August 1931. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177732979. See notes [Hsu38].

Hotelling:1940:SVU

[Hot40a] Harold Hotelling. The selection of variates for use in prediction with some comments on the general problem of nuisance parameters. Annals of Mathematical Statistics, 11(3):271–283, September 1940. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177731867.

Hotelling:1940:TS

[Hot40b] Harold Hotelling. The teaching of statistics. Annals of Mathematical Statistics, 11(4):457–470, December 1940. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177731833.

Hotelling:1941:EDM

[Hot41] Harold Hotelling. Experimental determination of the maximum of a function. Annals of Mathematical Statistics, 12(1):20–45, March 1941. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177731784.

Hotelling:1943:FPM

[Hot43a] Harold Hotelling. Further points on matrix calculation and simultaneous equations. Annals of Mathematical Statistics, 14(4):440–441, December 1943. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177731364.

Hotelling:1943:SNM

[Hot43b] Harold Hotelling. Some new methods in matrix calculation. Annals of Mathematical Statistics, 14(1):1–34, March 1943. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177731489. See note [Wau45].

Hotelling:1944:NMT

[Hot44a] Harold Hotelling. Note on a matric theorem of A. T. Craig. Annals of Mathematical Statistics, 15(4):427–429, December 1944. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177731213. See

[Cra43b].

Hotelling:1944:SIW

[Hot44b] Harold Hotelling. Some improvements in weighing and other experimental techniques. Annals of Mathematical Statistics, 15(3):297–306, September 1944. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177731236.

Houston:1966:SCL

[Hou66] Tom R. Houston. Sequential counterbalancing in Latin squares. Annals of Mathematical Statistics, 37(3):741–743, June 1966. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177699474.

Howell:1949:CCL

[How49] John M. Howell. Control chart for largest and smallest values. Annals of Mathematical Statistics, 20(2):305–309, June 1949. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177730041. See errata [How50].

Howell:1950:ECC

[How50] John M. Howell. Errata to “Control Chart for Largest and Smallest Values”. Annals of Mathematical Statistics, 21(4):615–616, December 1950. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/

1177729759. See [How49].

Hoyland:1965:RHL

[Høy65] Arnljot Høyland. Robustness of the Hodges–Lehmann estimates for shift. Annals of Mathematical Statistics, 36(1):174–197, February 1965. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177700281.

Hoyland:1968:RWE

[Høy68a] Arnljot Høyland. Robustness of the Wilcoxon estimate of location against a certain dependence. Annals of Mathematical Statistics, 39 (4):1196–1201, August 1968. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177698244.

Hoyle:1968:EVA

[Hoy68b] M. H. Hoyle. The estimation of variances after using a Gaussianating transformation. Annals of Mathematical Statistics, 39(4):1125–1143, August 1968. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177698237.

Hotelling:1936:RCT

[HP36] Harold Hotelling and Margaret Richards Pabst. Rank correlation and tests of significance involving no assumption of normality. Annals of Mathematical Statistics, 7(1):29–43, March 1936. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177732543.

Hoel:1949:SPO

[HP49] P. G. Hoel and R. P. Peterson. A solution to the problem of optimum classification. Annals of Mathematical Statistics, 20(3):433– 438, September 1949. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177729995.

Hobby:1963:CRF

[HP63a] Charles Hobby and Ronald Pyke. Combinatorial results in fluctuation theory. Annals of Mathematical Statistics, 34(4):1233–1242, December 1963. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/

1177703858. See correction [HP64].

Hobby:1963:CRM

[HP63b] Charles Hobby and Ronald Pyke. Combinatorial results in multidimensional fluctuation theory. Annals of Mathematical Statistics, 34 (2):402–404, June 1963. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177704150.

Hobby:1964:CCR

[HP64] Charles Hobby and Ronald Pyke. Correction to “Combinatorial Results in Fluctuation”. Annals of Mathematical Statistics, 35(2): 924, June 1964. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/

1177703609. See [HP63a].

Hollander:1972:TWN

[HP72] Myles Hollander and Frank Proschan. Testing whether new is better than used. Annals of Mathematical Statistics, 43(4):1136–1146, August 1972. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177692466.

Hogben:1964:ADV

[HPW64a] D. Hogben, R. S. Pinkham, and M. B. Wilk. An approximation to the distribution of Q (a variate related to the non-central t). Annals of Mathematical Statistics, 35(1):315–318, March 1964. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177703753.

Hogben:1964:MVR

[HPW64b] D. Hogben, R. S. Pinkham, and M. B. Wilk. The moments of a variate related to the non-central t. Annals of Mathematical Statistics, 35(1):298–314, March 1964. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177703752.

Hendricks:1936:SDC

[HR36] Walter A. Hendricks and Kate W. Robey. The sampling distribution of the coefficient of variation. Annals of Mathematical Statistics, 7 (3):129–132, September 1936. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177732503.

Hannan:1955:ASC

[HR55a] James F. Hannan and Herbert Robbins. Asymptotic solutions of the compound decision problem for two completely specified distributions. Annals of Mathematical Statistics, 26(1):37–51, March 1955. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177728591.

Hoeffding:1955:ET

[HR55b] Wassily Hoeffding and Joan Raup Rosenblatt. The efficiency of tests. Annals of Mathematical Statistics, 26(1):52–63, March 1955. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177728592.

Hartley:1962:SUP

[HR62] H. O. Hartley and J. N. K. Rao. Sampling with unequal probabilities and without replacement. Annals of Mathematical Statistics, 33(2): 350–374, June 1962. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177704564.

Hotelling:1932:LMS

[HS32] Harold Hotelling and Leonard M. Solomons. The limits of a measure of skewness. Annals of Mathematical Statistics, 3(2):141–142, May 1932. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177732911.

Halmos:1949:ARN

[HS49a] Paul R. Halmos and L. J. Savage. Application of the Radon–Nikodym theorem to the theory of sufficient statistics. Annals of Mathematical Statistics, 20(2):225–241, June 1949. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177730032.

Horton:1949:DMP

[HS49b] H. Burke Horton and R. Tynes Smith III. A direct method for producing random digits in any number system. Annals of Mathematical Statistics, 20(1):82–90, March 1949. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http:/

/projecteuclid.org/euclid.aoms/1177730092;   http://www.
jstor.org/stable/2236805.

Hoeffding:1955:BDF

[HS55] Wassily Hoeffding and S. S. Shrikhande. Bounds for the distribution function of a sum of independent, identically distributed random variables. Annals of Mathematical Statistics, 26(3):439–449, September 1955. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177728489.

Hoel:1961:CSM

[HS61] P. G. Hoel and E. M. Scheuer. Confidence sets for multivariate medians. Annals of Mathematical Statistics, 32(2):477–484, June 1961. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177705054.

Hedayat:1970:SOS

[HS70a] A. Hedayat and E. Seiden. F -square and orthogonal F -squares design: A generalization of latin square and orthogonal latin squares design. Annals of Mathematical Statistics, 41(6):2035–2044, December 1970. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177696703.

Hoeffding:1970:UCT

[HS70b] Wassily Hoeffding and Gordon Simons. Unbiased coin tossing with a biased coin. Annals of Mathematical Statistics, 41(2):341–352, April 1970. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177697074.

Horn:1970:EHS

[HS70c] Susan Horn and Siegfried Schach. An extension of the Hewitt–Savage zero-one law. Annals of Mathematical Statistics, 41(6):2130–2131, December 1970. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177696712.

Heath:1972:TFO

[HS72a] David C. Heath and William D. Sudderth. On a theorem of De Finetti, oddsmaking, and game theory. Annals of Mathematical Statistics, 43(6):2072–2077, December 1972. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177690887.

Hinich:1972:PED

[HS72b] Melvin J. Hinich and Paul Shaman. Parameter estimation for an R-dimensional plane wave observed with additive independent Gaussian errors. Annals of Mathematical Statistics, 43(1):153–169, February 1972. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177692710.

Hedayat:1972:SFD

[HSF72] A. Hedayat, E. Seiden, and W. T. Federer. Some families of designs for multistage experiments: Mutually balanced Youden designs when the number of treatments is prime power or twin primes. I. Annals of Mathematical Statistics, 43(5):1517–1527, October 1972. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177692384.

Hartley:1953:NIB

[HST53] H. O. Hartley, S. S. Shrikhande, and W. B. Taylor. A note on incomplete block designs with row balance. Annals of Mathematical Statistics, 24(1):123–126, March 1953. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177729091.

Hsu:1938:NHG

[Hsu38] P. L. Hsu. Notes on Hotelling’s generalized T. Annals of Mathematical Statistics, 9(4):231–243, December 1938. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177732280. See [Hot31].

Hsu:1940:SNB

[Hsu40] C. T. Hsu. On samples from a normal bivariate population. Annals of Mathematical Statistics, 11(4):410–426, December 1940. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177731828.

Hsu:1941:STB

[Hsu41] Chung Tsi Hsu. Samples from two bivariate normal populations. Annals of Mathematical Statistics, 12(3):279–292, September 1941. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177731710.

Hsu:1944:SCF

[Hsu44] L. C. Hsu. Some combinatorial formulas with applications to probable values of a polynomial-product and to differences of zero. Annals of Mathematical Statistics, 15(4):399–413, December 1944. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177731210.

Hsu:1945:SCF

[Hsu45a] L. C. Hsu. Some combinatorial formulas on mathematical expectation. Annals of Mathematical Statistics, 16(4):369–380, December 1945. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177731063.

Hsu:1945:ADM

[Hsu45b] P. L. Hsu. The approximate distributions of the mean and variance of a sample of independent variables. Annals of Mathematical Statistics, 16(1):1–29, March 1945. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177731168.

Hsu:1945:ADR

[Hsu45c] P. L. Hsu. On the approximate distribution of ratios. Annals of Mathematical Statistics, 16(2):204–210, June 1945. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177731122.

Hsu:1945:PFT

[Hsu45d] P. L. Hsu. On the power functions of the E^{2}-test and the T^{2}-test. Annals of Mathematical Statistics, 16(3):278–286, September 1945. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177731091.

Hsu:1946:ADC

[Hsu46] P. L. Hsu. On the asymptotic distributions of certain statistics used in testing the independence between successive observations from a normal population. Annals of Mathematical Statistics, 17(3):350– 354, September 1946. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177730948.

Hsu:1948:NAE

[Hsu48] L. C. Hsu. Note on an asymptotic expansion of the n th difference of zero. Annals of Mathematical Statistics, 19(2):273–277, June 1948. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177730254.

Huang:1972:NRC

[Hua72] J. S. Huang. A note on Robbins’ compound decision procedure. Annals of Mathematical Statistics, 43(1):348–350, February 1972. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177692729.

Huber:1963:PCR

[Hub63a] Peter J. Huber. Pairwise comparison and ranking: Optimum properties of the row sum procedure. Annals of Mathematical Statistics, 34 (2):511–520, June 1963. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177704162.

Huber:1963:RPT

[Hub63b] Peter J. Huber. A remark on a paper of Trawinski and David entitled: “Selection of the Best Treatment in a paired-Comparison Experiment”. Annals of Mathematical Statistics, 34(1):92–94, March 1963. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177704244. See

[TD63].

Huber:1964:REL

[Hub64] Peter J. Huber. Robust estimation of a location parameter. Annals of Mathematical Statistics, 35(1):73–101, March 1964. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177703732.

Huber:1965:RVP

[Hub65] Peter J. Huber. A robust version of the probability ratio test. Annals of Mathematical Statistics, 36(6):1753–1758, December 1965. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177699803.

Huber:1972:WLR

[Hub72] Peter J. Huber. The 1972 Wald lecture. Robust statistics: A review. Annals of Mathematical Statistics, 43(4):1041–1067, August 1972. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177692459.

Huff:1969:LSD

[Huf69] Barthel W. Huff. The loose subordination of differential processes to Brownian motion. Annals of Mathematical Statistics, 40(5):1603– 1609, October 1969. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177697375.

Huff:1971:SCI

[Huf71] Barthel W. Huff. Superpositions and the continuity of infinitely divisible distributions. Annals of Mathematical Statistics, 42(1):410–412, February 1971. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177693536.

Huntington:1939:FDP

[Hun39] E. V. Huntington. Frequency distribution of product and quotient. Annals of Mathematical Statistics, 10(2):195–198, June 1939. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177732218.

Huntsberger:1955:GPT

[Hun55] D. V. Huntsberger. A generalization of a preliminary testing procedure for pooling data. Annals of Mathematical Statistics, 26(4):734– 743, December 1955. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177728431.

Hurwicz:1948:RJN

[Hur48] Leonid Hurwicz. Review: John von Neumann, Oskar Morgenstern, The Theory of Games and Economic Behavior. Annals of Mathematical Statistics, 19(3):436–437, September 1948. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177730216.

Hannan:1965:RCC

[HV65] J. F. Hannan and J. R. Van Ryzin. Rate of convergence in the compound decision problem for two completely specified distributions.

Annals of Mathematical Statistics, 36(6):1743–1752, December 1965. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177699802.

Hart:1942:TPR

[HvN42] B. I. Hart and John von Neumann. Tabulation of the probabilities for the ratio of the mean square successive difference to the variance. Annals of Mathematical Statistics, 13(2):207–214, June 1942. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177731606. Reprinted

in [Tau62, Paper 36].

Hartman:1948:EDC

[HW48] Philip Hartman and Aurel Wintner. On the effect of decimal corrections on errors of observation. Annals of Mathematical Statistics, 19 (3):389–393, September 1948. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177730203.

Hoeffding:1958:DSD

[HW58] Wassily Hoeffding and J. Wolfowitz. Distinguishability of sets of distributions. (The case of independent and identically distributed chance variables). Annals of Mathematical Statistics, 29(3):700–718, September 1958. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177706531.

Hanson:1971:BTP

[HW71] D. L. Hanson and F. T. Wright. A bound on tail probabilities for quadratic forms in independent random variables. Annals of Mathematical Statistics, 42(3):1079–1083, June 1971. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177693335.

Hall:1965:RBS

[HWG65] W. J. Hall, R. A. Wijsman, and J. K. Ghosh. The relationship between sufficiency and invariance with applications in sequential analysis. Annals of Mathematical Statistics, 36(2):575–614, April 1965. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177700169. See

correction [HWG95].

Hall:1995:CRB

[HWG95] W. J. Hall, R. A. Wijsman, and J. K. Ghosh. Correction: “The Relationship Between Sufficiency and Invariance with Applications in Sequential Analysis”. Annals of Statistics, 23(2):705, April 1995. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176324543.

Hobby:1964:SST

[HY64] Charles Hobby and N. Donald Ylvisaker. Some structure theorems for stationary probability measures on finite state sequences. Annals of Mathematical Statistics, 35(2):550–556, June 1964. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177703554.

Ifram:1965:HFS

[Ifr65a] Adnan F. Ifram. Hypergeometric functions in sequential analysis. Annals of Mathematical Statistics, 36(6):1870–1872, December 1965. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177699819.

Ifram:1965:ABD

[Ifr65b] Adnan F. Ifram. On the asymptotic behavior of densities with applications to sequential analysis. Annals of Mathematical Statistics, 36 (2):615–637, April 1965. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177700170.

Ifram:1966:SSS

[Ifr66] Adnan F. Ifram. On the sample size and simplification of a class of sequential probability ratio tests. Annals of Mathematical Statistics, 37(2):425–434, April 1966. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177699524.

Iglehart:1964:MCP

[Igl64] Donald L. Iglehart. Multivariate competition processes. Annals of Mathematical Statistics, 35(1):350–361, March 1964. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177703758.

Iglehart:1965:LTQ

[Igl65] Donald L. Iglehart. Limit theorems for queues with traffic intensity one. Annals of Mathematical Statistics, 36(5):1437–1449, October 1965. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177699903.

Iglehart:1968:LTM

[Igl68] Donald L. Iglehart. Limit theorems for the multi-urn Ehrenfest model. Annals of Mathematical Statistics, 39(3):864–876, June 1968. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177698318.

Iglehart:1972:EVG

[Igl72] Donald L. Iglehart. Extreme values in the GI/G/1 queue. Annals of Mathematical Statistics, 43(2):627–635, April 1972. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177692642.

Imhof:1960:MMC

[Imh60] J. P. Imhof. A mixed model for the complete three-way layout with two random-effects factors. Annals of Mathematical Statistics, 31 (4):906–928, December 1960. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177705666.

Imhof:1962:THN

[Imh62] J. P. Imhof. Testing the hypothesis of no fixed main-effects in Scheffe’s mixed model. Annals of Mathematical Statistics, 33(3):1085–1095, September 1962. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177704474.

Imhof:1968:SIL

[Imh68] J. P. Imhof. Some invariant laws related to the arc sine law. Annals of Mathematical Statistics, 39(1):258–260, February 1968. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177698527.

Imhof:1971:SJL

[Imh71] J. P. Imhof. Some joint laws in fluctuation theory. Annals of Mathematical Statistics, 42(3):1099–1103, June 1971. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177693340.

Iosifescu:1965:SER

[Ios65] Marius Iosifescu. Sampling entropy for random homogeneous systems with complete connections. Annals of Mathematical Statistics, 36 (5):1433–1436, October 1965. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177699902. See correction [Ios69].

Iosifescu:1969:CNC

[Ios69] Marius Iosifescu. Correction note: Correction to “Sampling Entropy for Random Homogeneous Systems with Complete Connections”. Annals of Mathematical Statistics, 40(6):2215, December 1969. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177697304. See

[Ios65].

Iosifescu:1972:TRP

[Ios72] Marius Iosifescu. On two recent papers on ergodicity in nonhomogeneous Markov chains. Annals of Mathematical Statistics, 43(5):1732– 1736, October 1972. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177692411.

Isaacson:1951:TUT

[Isa51] Stanley L. Isaacson. On the theory of unbiased tests of simple statistical hypotheses specifying the values of two or more parameters. Annals of Mathematical Statistics, 22(2):217–234, June 1951. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177729642.

Isaac:1963:GVD

[Isa63] Richard Isaac. A general version of Doeblin’s condition. Annals of Mathematical Statistics, 34(2):668–671, June 1963. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177704180.

Isaac:1964:NSR

[Isa64a] Richard Isaac. Non-singular recurrent Markov processes have stationary measures. Annals of Mathematical Statistics, 35(2):869– 871, June 1964. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177703588.

Isaac:1964:UTS

[Isa64b] Richard Isaac. A uniqueness theorem for stationary measures of ergodic Markov processes. Annals of Mathematical Statistics, 35(4): 1781–1786, December 1964. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177700399.

Isaac:1967:SMP

[Isa67] Richard Isaac. On stationary Markov processes. Annals of Mathematical Statistics, 38(2):588–592, April 1967. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177698975.

Isaacson:1969:NTS

[Isa69a] Dean Isaacson. Note on the three series theorem. Annals of Mathematical Statistics, 40(5):1844, October 1969. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177697399.

Isaacson:1969:SID

[Isa69b] Dean Isaacson. Stochastic integrals and derivatives. Annals of Mathematical Statistics, 40(5):1610–1616, October 1969. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177697376.

Isaacson:1971:CMD

[Isa71] Dean Isaacson. Continuous martingales with discontinuous marginal distributions. Annals of Mathematical Statistics, 42(6):2139–2142, December 1971. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177693081.

Isaac:1972:LTM

[Isa72a] Richard Isaac. Limit theorems for Markov transition functions. Annals of Mathematical Statistics, 43(2):621–626, April 1972. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177692641.

Isaacson:1972:UIS

[Isa72b] Dean Isaacson. Uniform integrability of square integrable martingales. Annals of Mathematical Statistics, 43(2):688–689, April 1972. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177692656.

Isbell:1959:PR

[Isb59] J. R. Isbell. On a problem of Robbins. Annals of Mathematical Statistics, 30(2):606–610, June 1959. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177706278.

Iglehart:1968:WCS

[IT68] Donald L. Iglehart and Howard M. Taylor. Weak convergence of a sequence of quickest detection problems. Annals of Mathematical Statistics, 39(6):2149–2153, December 1968. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177698048.

Ito:1956:AFD

[Ito56] Koichi Ito. Asymptotic formulae for the distribution of Hotelling’s generalized T_{0}^{2} statistic. Annals of Mathematical Statistics, 27(4): 1091–1105, December 1956. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177728075.

Ito:1960:AFD

[Ito60] Koichi Ito. Asymptotic formulae for the distribution of Hotelling’s generalized T_{0}^{2} statistic. II. Annals of Mathematical Statistics, 31 (4):1148–1153, December 1960. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177705685.

Iglehart:1971:EFC

[IW71] Donald L. Iglehart and Ward Whitt. The equivalence of functional Central Limit Theorems for counting processes and associated partial sums. Annals of Mathematical Statistics, 42(4):1372–1378, August 1971. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177693249.

Iyer:1950:TPD

[Iye50] P. V. Krishna Iyer. The theory of probability distributions of points on a lattice. Annals of Mathematical Statistics, 21(2):198– 217, June 1950. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/

1177729839. See errata [Iye51].

Iyer:1951:ETP

[Iye51] P. V. Krishna Iyer. Errata: “The theory of probability distributions of points on a lattice”. Annals of Mathematical Statistics, 22(2): 310, June 1951. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/

1177729657. See [Iye50].

Iyer:1958:TFM

[Iye58] P. V. Krishna Iyer. A theorem on factorial moments and its applications. Annals of Mathematical Statistics, 29(1):254–261, March 1958. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177706723. See

correction [Iye61a].

Iyer:1961:CNCb

[Iye61a] P. V. Krishna Iyer. Correction notes: Corrections to “A Theorem on Factorial Moments and Its Applications”. Annals of Mathematical Statistics, 32(2):620, June 1961. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177705075. See [Iye58].

Iyer:1961:CNCa

[Iye61b] P. V. Krishna Iyer. Correction notes: Corrections to “The Theory of Probability Distributions of Points on a Lattice”. Annals of Mathematical Statistics, 32(2):619, June 1961. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177705074.

Jacobson:1969:MVR

[Jac69] Harold I. Jacobson. The maximum variance of restricted unimodal distributions. Annals of Mathematical Statistics, 40(5):1746–1752, October 1969. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177697386.

Jaeckel:1971:REL

[Jae71a] Louis A. Jaeckel. Robust estimates of location: Symmetry and asymmetric contamination. Annals of Mathematical Statistics, 42(3):1020– 1034, June 1971. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177693330.

Jaeckel:1971:SFE

[Jae71b] Louis A. Jaeckel. Some flexible estimates of location. Annals of Mathematical Statistics, 42(5):1540–1552, October 1971. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177693152.

Jaeckel:1972:ERC

[Jae72] Louis A. Jaeckel. Estimating regression coefficients by minimizing the dispersion of the residuals. Annals of Mathematical Statistics, 43 (5):1449–1458, October 1972. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177692377.

Jagerman:1963:AFS

[Jag63] David L. Jagerman. The autocorrelation function of a sequence uniformly distributed modulo 1. Annals of Mathematical Statistics, 34(4):1243–1252, December 1963. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http:/

/projecteuclid.org/euclid.aoms/1177703859;   http://www.
jstor.org/stable/2238334.

Jagers:1971:DAB

[Jag71] Peter Jagers. Diffusion approximations of branching processes. Annals of Mathematical Statistics, 42(6):2074–2078, December 1971. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177693076.

Jain:1966:NIM

[Jai66] N. C. Jain. A note on invariant measures. Annals of Mathematical Statistics, 37(3):729–732, June 1966. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177699470.

Jain:1969:SRL

[Jai69] Naresh C. Jain. The strong ratio limit property for some general Markov processes. Annals of Mathematical Statistics, 40(3):986– 992, June 1969. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177697603.

Jain:1970:RHT

[Jai70] Naresh C. Jain. Review: Howard Tucker, A Graduate Course in Probability. Annals of Mathematical Statistics, 41(2):748–749, April 1970.

CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177697130.

Jambunathan:1954:SPB

[Jam54a] M. V. Jambunathan. Some properties of Beta and Gamma distributions. Annals of Mathematical Statistics, 25(2):401–405, June 1954. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177728800.

James:1954:NMA

[Jam54b] A. T. James. Normal multivariate analysis and the orthogonal group. Annals of Mathematical Statistics, 25(1):40–75, March 1954. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177728846.

James:1957:RAE

[Jam57] A. T. James. The relationship algebra of an experimental design. Annals of Mathematical Statistics, 28(4):993–1002, December 1957. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177706800.

James:1960:DLR

[Jam60] Alan T. James. The distribution of the latent roots of the covariance matrix. Annals of Mathematical Statistics, 31(1):151–158, March 1960. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177705994.

James:1961:DNM

[Jam61] Alan T. James. The distribution of noncentral means with known covariance. Annals of Mathematical Statistics, 32(3):874–882, September 1961. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177704980.

James:1964:DMV

[Jam64] Alan T. James. Distributions of matrix variates and latent roots derived from normal samples. Annals of Mathematical Statistics, 35 (2):475–501, June 1964. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177703550.

James:1968:CZP

[Jam68] A. T. James. Calculation of zonal polynomial coefficients by use of the Laplace–Beltrami operator. Annals of Mathematical Statistics, 39(5):1711–1718, October 1968. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177698153.

Jamison:1970:RPS

[Jam70] Benton Jamison. Reciprocal processes: The stationary Gaussian case. Annals of Mathematical Statistics, 41(5):1624–1630, October 1970. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177696805.

Jaquette:1972:MDP

[Jaq72] Stratton C. Jaquette. Markov decision processes with a new optimality criterion: Small interest rates. Annals of Mathematical Statistics, 43(6):1894–1901, December 1972. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177690860.

Jackson:1961:ST

[JB61] J. Edward Jackson and Ralph A. Bradley. Sequential χ^{2}- and T^{2}-tests. Annals of Mathematical Statistics, 32(4):1063–1077, December 1961. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177704846.

Jeeves:1954:IEL

[Jee54] T. A. Jeeves. Identification and estimation of linear manifolds in n-dimensions. Annals of Mathematical Statistics, 25(4):714–723, December 1954. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177728657.

Jenkins:1932:SMT

[Jen32] Thomas N. Jenkins. A short method and tables for the calculation of the average and standard deviation of logarithmic distributions. Annals of Mathematical Statistics, 3(1):45–55, February 1932. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177732932.

Jennrich:1969:APN

[Jen69] Robert I. Jennrich. Asymptotic properties of non-linear least squares estimators. Annals of Mathematical Statistics, 40(2):633– 643, April 1969. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177697731.

Jensen:1970:CNC

[Jen70a] D. R. Jensen. Correction note: Correction to “The Joint Distribution of Traces of Wishart Matrices and Some Applications”. Annals of Mathematical Statistics, 41(6):2186, December 1970. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177696729. See [Jen70b].

Jensen:1970:JDT

[Jen70b] D. R. Jensen. The joint distribution of traces of Wishart matrices and some applications. Annals of Mathematical Statistics, 41(1):133–145, February 1970. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/

1177697194. See correction [Jen70a].

Johnson:1964:PMS

[JF64] Charles Henry Johnson and John Leroy Folks. A property of the method of steepest ascent. Annals of Mathematical Statistics, 35(1): 435–437, March 1964. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177703770.

Jirina:1972:CDR

[Jiř72] Miloslav Jiřina. Convergence in distribution of random measures. Annals of Mathematical Statistics, 43(5):1727–1731, October 1972. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177692410.

Joffe:1962:NDB

[JK62] A. Joffe and Jerome Klotz. Null distribution and Bahadur efficiency of the Hodges bivariate sign test. Annals of Mathematical Statistics, 33 (2):803–807, June 1962. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177704600.

Johansen:1966:SCT

[JK66] S. Johansen and J. Karush. On the semimartingale convergence theorem. Annals of Mathematical Statistics, 37(3):690–694, June 1966. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177699463.

Jain:1970:NUC

[JK70] Naresh C. Jain and G. Kallianpur. A note on uniform convergence of stochastic processes. Annals of Mathematical Statistics, 41(4):1360– 1362, August 1970. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177696914.

Johns:1963:ARL

[JM63] M. V. Johns, Jr. and R. G. Miller, Jr. Average renewal loss rates. Annals of Mathematical Statistics, 34(2):396–401, June 1963. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177704149.

Johnson:1972:LMP

[JM72] Richard A. Johnson and K. G. Mehrotra. Locally most powerful rank tests for the two-sample problem with censored data. Annals of Mathematical Statistics, 43(3):823–831, June 1972. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177692548.

Joffe:1963:CTM

[JN63] A. Joffe and P. E. Ney. Convergence theorems for multiple channel loss probabilities. Annals of Mathematical Statistics, 34(1):260– 273, March 1963. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177704263.

Jamison:1969:OST

[JO69] Benton Jamison and Steven Orey. An optional stopping theorem. Annals of Mathematical Statistics, 40(2):677–678, April 1969. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177697739.

Joffe:1967:GWB

[Jof67a] A. Joffe. On the Galton–Watson branching process with mean less than one. Annals of Mathematical Statistics, 38(1):264–266, February 1967. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177699079.

Joffe:1967:MCS

[Jof67b] A. D. Joffe. Minimum chi-squared estimation using independent statistics. Annals of Mathematical Statistics, 38(1):267–270, February 1967. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177699080.

Jogdeo:1966:RRS

[Jog66] Kumar Jogdeo. On randomized rank score procedures of Bell and doksum. Annals of Mathematical Statistics, 37(6):1697–1703, December 1966. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177699159.

Jogdeo:1967:MCB

[Jog67] Kumar Jogdeo. Monotone convergence of binomial probabilities with an application to maximum likelihood estimation. Annals of Mathematical Statistics, 38(5):1583–1586, October 1967. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177698714.

Jogdeo:1968:ANN

[Jog68a] Kumar Jogdeo. Asymptotic normality in nonparametric methods. Annals of Mathematical Statistics, 39(3):905–922, June 1968. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177698323.

Jogdeo:1968:CIC

[Jog68b] Kumar Jogdeo. Characterizations of independence in certain families of bivariate and multivariate distributions. Annals of Mathematical Statistics, 39(2):433–441, April 1968. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177698407.

Jogdeo:1970:SPI

[Jog70] Kumar Jogdeo. A simple proof of an inequality for multivariate normal probabilities of rectangles. Annals of Mathematical Statistics, 41(4):1357–1359, August 1970. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177696913.

Johnson:1940:PTL

[Joh40a] N. L. Johnson. Parabolic test for linkage. Annals of Mathematical Statistics, 11(3):227–253, September 1940. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177731865.

Johnson:1940:EPM

[Joh40b] Evan Johnson, Jr. Estimates of parameters by means of least squares. Annals of Mathematical Statistics, 11(4):453–456, December 1940. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177731832.

Johnson:1951:EPZ

[Joh51] N. L. Johnson. Estimators of the probability of the zero class in Poisson and certain related populations. Annals of Mathematical Statistics, 22(1):94–101, March 1951. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177729696.

Johnson:1953:SNA

[Joh53] N. L. Johnson. Some notes on the application of sequential methods in the analysis of variance. Annals of Mathematical Statistics, 24 (4):614–623, December 1953. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177728917.

Johnson:1954:SPC

[Joh54] N. L. Johnson. Sequential procedures in certain component of variance problems. Annals of Mathematical Statistics, 25(2):357–366, June 1954. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177728791.

John:1957:DTH

[Joh57a] Peter W. M. John. Divergent time homogeneous birth and death processes. Annals of Mathematical Statistics, 28(2):514–517, June 1957. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177706983.

Johns:1957:NPE

[Joh57b] M. V. Johns, Jr. Non-parametric empirical Bayes procedures. Annals of Mathematical Statistics, 28(3):649–669, September 1957. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177706877.

Johnson:1959:PWT

[Joh59] N. L. Johnson. A proof of Wald’s theorem on cumulative sums. Annals of Mathematical Statistics, 30(4):1245–1247, December 1959. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177706108. See

correction [Joh61b].

John:1961:ED

[Joh61a] S. John. Errors in discrimination. Annals of Mathematical Statistics, 32(4):1125–1144, December 1961. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177704851.

Johnson:1961:CPW

[Joh61b] N. L. Johnson. Correction to “A Proof of Wald’s Theorem on Cumulative Sums”. Annals of Mathematical Statistics, 32(4):1344, December 1961. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/

1177704877. See [Joh59].

John:1964:BDU

[Joh64a] Peter W. M. John. Balanced designs with unequal numbers of replicates. Annals of Mathematical Statistics, 35(2):897–899, June 1964. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177703597.

John:1964:PIA

[Joh64b] Peter W. M. John. Pseudo-inverses in the analysis of variance. Annals of Mathematical Statistics, 35(2):895–896, June 1964. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177703596.

Johnson:1965:ESQ

[Joh65] Bruce McK. Johnson. The ergodicity of series queues with general priorities. Annals of Mathematical Statistics, 36(6):1664–1676, December 1965. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177699795.

John:1966:IRD

[Joh66] Peter W. M. John. On identity relationships for 2_{n−r} designs having words of equal length. Annals of Mathematical Statistics, 37(6):1842– 1843, December 1966. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177699179.

John:1967:RGD

[Joh67a] J. A. John. Reduced group divisible paired comparison designs. Annals of Mathematical Statistics, 38(6):1887–1893, December 1967. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177698622.

John:1967:OBI

[Joh67b] Peter W. M. John. On obtaining balanced incomplete block designs from partially balanced association schemes. Annals of Mathematical Statistics, 38(2):618–619, April 1967. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177698985.

Johnson:1967:AEP

[Joh67c] R. A. Johnson. An asymptotic expansion for posterior distributions. Annals of Mathematical Statistics, 38(6):1899–1906, December 1967. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177698624.

Johnson:1967:AEA

[Joh67d] R. A. Johnson. Asymptotic expansions associated with the n th power of a density. Annals of Mathematical Statistics, 38(4):1266– 1272, August 1967. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177698796.

Johnson:1969:RFH

[Joh69] Norman L. Johnson. Review: Frank A. Haight, Handbook of the Poisson Distribution. Annals of Mathematical Statistics, 40(1):326– 327, February 1969. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177697833.

John:1970:FCA

[Joh70a] Peter W. M. John. The folded cubic association scheme. Annals of Mathematical Statistics, 41(6):1983–1991, December 1970. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177696699.

John:1970:NLB

[Joh70b] Peter W. M. John. The nonexistence of linked block designs with latin square association schemes. Annals of Mathematical Statistics, 41(3):1105–1107, June 1970. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177696992.

Johnson:1970:AEA

[Joh70c] Richard A. Johnson. Asymptotic expansions associated with posterior distributions. Annals of Mathematical Statistics, 41(3):851– 864, June 1970. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177696963.

Johnson:1971:AEC

[Joh71] Bruce McK. Johnson. On the admissible estimators for certain fixed sample binomial problems. Annals of Mathematical Statistics, 42 (5):1579–1587, October 1971. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177693156.

Johnson:1972:HTD

[Joh72] Dudley Paul Johnson. Hitting time distributions for general stochastic processes. Annals of Mathematical Statistics, 43(5):1716–1718, October 1972. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177692408.

Jones:1948:ELM

[Jon48] Howard L. Jones. Exact lower moments of order statistics in small samples from a normal distribution. Annals of Mathematical Statistics, 19(2):270–273, June 1948. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177730253.

Jones:1952:FGS

[Jon52] Howard L. Jones. Formulas for the group sequential sampling of attributes. Annals of Mathematical Statistics, 23(1):72–87, March 1952. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177729486.

Jones:1962:SAR

[Jon62] Richard H. Jones. Spectral analysis with regularly missed observations. Annals of Mathematical Statistics, 33(2):455–461, June 1962. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177704572.

Jones:1963:SPS

[Jon63] Richard H. Jones. Stochastic processes on a sphere. Annals of Mathematical Statistics, 34(1):213–218, March 1963. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177704257.

Jones:1969:PFE

[Jon69] Richard H. Jones. Phase free estimation of coherence. Annals of Mathematical Statistics, 40(2):540–548, April 1969. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177697722.

Jones:1970:RES

[Jon70] Richard H. Jones. Recursive estimation of a subset of regression coefficients. Annals of Mathematical Statistics, 41(2):688–691, April 1970. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177697115.

Jordan:1932:AGA

[Jor32] Charles Jordan. Approximation and graduation according to the principle of least squares by orthogonal polynomials. Annals of Mathematical Statistics, 3(4):257–357, November 1932. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http:

//projecteuclid.org/euclid.aoms/1177732839. See correction

[Jor59].

Jordan:1959:CNC

[Jor59] Charles Jordan. Correction notes: Corrections to “Approximation and Graduation According to the Principle of Least Squares by Orthogonal Polynomials”. Annals of Mathematical Statistics, 30(4): 1266, December 1959. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177706117. See [Jor32].

Joseph:1939:CE

[Jos39] J. A. Joseph. On the coefficients of the expansion of X^{(n)}. Annals of Mathematical Statistics, 10(3):293–296, September 1939. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177732188.

Joshi:1965:ABEa

[Jos65a] V. M. Joshi. Admissibility and Bayes estimation in sampling finite populations. II. Annals of Mathematical Statistics, 36(6):1723–1729, December 1965. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177699800.

Joshi:1965:ABEb

[Jos65b] V. M. Joshi. Admissibility and Bayes estimation in sampling finite populations. III. Annals of Mathematical Statistics, 36(6):1730–1742, December 1965. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177699801.

Joshi:1966:ABE

[Jos66a] V. M. Joshi. Admissibility and Bayes estimation in sampling finite populations-IV. Annals of Mathematical Statistics, 37(6):1658–1670, December 1966. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177699154.

Joshi:1966:ACI

[Jos66b] V. M. Joshi. Admissibility of confidence intervals. Annals of Mathematical Statistics, 37(3):629–638, June 1966. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177699456.

Joshi:1967:CIM

[Jos67a] V. M. Joshi. Confidence intervals for the mean of a finite population. Annals of Mathematical Statistics, 38(4):1180–1207, August 1967. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177698788.

Joshi:1967:IUC

[Jos67b] V. M. Joshi. Inadmissibility of the usual confidence sets for the mean of a multivariate normal population. Annals of Mathematical Statistics, 38(6):1868–1875, December 1967. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177698619.

Joshi:1968:ASM

[Jos68a] V. M. Joshi. Admissibility of the sample mean as estimate of the mean of a finite population. Annals of Mathematical Statistics, 39 (2):606–620, April 1968. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177698421.

Joshi:1968:NMD

[Jos68b] V. M. Joshi. Note on a minimax design for cluster sampling. Annals of Mathematical Statistics, 39(1):278–281, February 1968. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177698534.

Joshi:1969:AUC

[Jos69a] V. M. Joshi. Admissibility of the usual confidence sets for the mean of a univariate or bivariate normal population. Annals of Mathematical Statistics, 40(3):1042–1067, June 1969. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177697608.

Joshi:1969:TKR

[Jos69b] V. M. Joshi. On a theorem of Karlin regarding admissible estimates for exponential populations. Annals of Mathematical Statistics, 40 (1):216–223, February 1969. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177697818.

Joshi:1970:AIC

[Jos70] V. M. Joshi. Admissibility of invariant confidence procedures for estimating a location parameter. Annals of Mathematical Statistics, 41(5):1568–1581, October 1970. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/ euclid.aoms/1177696801. See correction [Jos71d].

Joshi:1971:RRM

[Jos71a] Prakash C. Joshi. Recurrence relations for the mixed moments of order statistics. Annals of Mathematical Statistics, 42(3):1096–1098, June 1971. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177693339.

Joshi:1971:AAE

[Jos71b] V. M. Joshi. Admissibility of arbitrary estimates as unbiased estimates of their expectations in a finite population. Annals of Mathematical Statistics, 42(2):839–841, April 1971. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177693445.

Joshi:1971:AUC

[Jos71c] V. M. Joshi. Admissibility of the usual confidence sets for a class of bivariate populations. Annals of Mathematical Statistics, 42(2): 662–679, April 1971. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177693416.

Joshi:1971:CNC

[Jos71d] V. M. Joshi. Correction note: Correction to “Admissibility of Invariant Confidence Procedures for a Location Parameter”. Annals of Mathematical Statistics, 42(4):1462, August 1971. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177693264. See [Jos70].

Joshi:1971:HEF

[Jos71e] V. M. Joshi. Hyperadmissibility of estimators for finite populations. Annals of Mathematical Statistics, 42(2):680–690, April 1971. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177693417.

Joshi:1971:NAS

[Jos71f] V. M. Joshi. A note on admissible sampling designs for a finite population. Annals of Mathematical Statistics, 42(4):1425–1428, August 1971. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177693253.

Joshi:1972:NHE

[Jos72] V. M. Joshi. A note on hyperadmissibility of estimators for finite populations. Annals of Mathematical Statistics, 43(4):1323–1328, August 1972. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177692484.

Jain:1972:LIL

[JP72] Naresh C. Jain and William E. Pruitt. The law of the iterated logarithm for the range of random walk. Annals of Mathematical Statistics, 43(5):1692–1697, October 1972. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177692404.

Johnson:1951:MPU

[JR51] N. L. Johnson and C. A. Rogers. The moment problem for unimodal distributions. Annals of Mathematical Statistics, 22(3):433– 439, September 1951. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177729590.

Johnson:1969:AMP

[JR69] Richard A. Johnson and George G. Roussas. Asymptotically most powerful tests in Markov processes. Annals of Mathematical Statistics, 40(4):1207–1215, August 1969. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177697497.

Johnson:1970:AOT

[JR70] Richard A. Johnson and George G. Roussas. Asymptotically optimal tests in Markov processes. Annals of Mathematical Statistics, 41(3): 918–938, June 1970. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177696969.

Jogdeo:1968:MCB

[JS68] Kumar Jogdeo and S. M. Samuels. Monotone convergence of binomial probabilities and a generalization of Ramanujan’s equation. Annals of Mathematical Statistics, 39(4):1191–1195, August 1968. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177698243.

Jacobs:1972:SOR

[JS72] David R. Jacobs, Jr. and Siegfried Schach. Stochastic order relationships between GI/G/k systems. Annals of Mathematical Statistics, 43(5):1623–1633, October 1972. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177692394.

Jureckova:1969:ALR

[Jur69] Jana Jurečková. Asymptotic linearity of a rank statistic in regression parameter. Annals of Mathematical Statistics, 40(6):1889–1900, December 1969. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177697273.

Jureckova:1971:NER

[Jur71] Jana Jurečková. Nonparametric estimate of regression coefficients. Annals of Mathematical Statistics, 42(4):1328–1338, August 1971. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177693245.

Johns:1971:CRE

[JV71] M. V. Johns, Jr. and J. Van Ryzin. Convergence rates for empirical Bayes two-action problems I. Discrete case. Annals of Mathematical Statistics, 42(5):1521–1539, October 1971. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177693151.

Johns:1972:CRE

[JV72] M. V. Johns, Jr. and J. Van Ryzin. Convergence rates for empirical Bayes two-action problems II. Continuous case. Annals of Mathematical Statistics, 43(3):934–947, June 1972. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177692557.

Kotz:1964:APD

[KA64a] Samuel Kotz and John W. Adams. Acknowledgment of priority: Distribution of sum of identically distributed exponentially correlated gamma-variables. Annals of Mathematical Statistics, 35(2): 925, June 1964. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177703611.

Kotz:1964:DSI

[KA64b] Samuel Kotz and John W. Adams. Distribution of sum of identically distributed exponentially correlated gamma-variables. Annals of Mathematical Statistics, 35(1):277–283, March 1964. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177703750.

Krishnaiah:1965:CNC

[KA65] P. R. Krishnaiah and J. V. Armitage. Correction notes: Correction to abstract “Tables for the Distribution of the Maximum of Correlated Chi-Square Variates with One Degree of Freedom”. Annals of Mathematical Statistics, 36(3):1069, June 1965. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177700087. See [Ano65a].

Kabe:1962:EDC

[Kab62] D. G. Kabe. On the exact distribution of a class of multivariate test criteria. Annals of Mathematical Statistics, 33(3):1197–1200, September 1962. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177704484.

Kabe:1963:SRD

[Kab63] D. G. Kabe. Some results on the distribution of two random matrices used in classification procedures. Annals of Mathematical Statistics, 34(1):181–185, March 1963. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/ euclid.aoms/1177704252. See corrections [Kab64].

Kabe:1964:CSR

[Kab64] D. G. Kabe. Corrections to: “Some Results on the Distribution of Two Random Matrices used in Classification Procedures”. Annals of Mathematical Statistics, 35(2):924, June 1964. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177703608. See [Kab63].

Kabe:1965:GSL

[Kab65] D. G. Kabe. Generalization of Sverdrup’s lemma and its applications to multivariate distribution theory. Annals of Mathematical Statistics, 36(2):671–676, April 1965. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177700174.

Kabe:1968:DDC

[Kab68] D. G. Kabe. On the distributions of direction and collinearity factors in discriminant analysis. Annals of Mathematical Statistics, 39(3): 855–858, June 1968. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177698316.

Kac:1945:RWP

[Kac45a] M. Kac. Random walk in the presence of absorbing barriers. Annals of Mathematical Statistics, 16(1):62–67, March 1945. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177731171.

Kac:1945:RIL

[Kac45b] M. Kac. A remark on independence of linear and quadratic forms involving independent Gaussian variables. Annals of Mathematical Statistics, 16(4):400–401, December 1945. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177731069.

Kac:1948:CFD

[Kac48] M. Kac. On the characteristic functions of the distributions of estimates of various deviations in samples from a normal population. Annals of Mathematical Statistics, 19(2):257–261, June 1948. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177730250. See

note [God49a].

Kadane:1966:SEC

[Kad66] Joseph B. Kadane. Some equivalence classes in paired comparisons. Annals of Mathematical Statistics, 37(2):488–494, April 1966. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177699532.

Kadiyala:1968:IRT

[Kad68] Koteswara Rao Kadiyala. An inequality for the ratio of two quadratic forms in normal variates. Annals of Mathematical Statistics, 39 (5):1762–1763, October 1968. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177698164.

Kadane:1971:MPO

[Kad71] Joseph B. Kadane. A moment problem for order statistics. Annals of Mathematical Statistics, 42(2):745–751, April 1971. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177693423. See correc-

tion [Kad73].

Kadane:1973:CNC

[Kad73] Joseph B. Kadane. Correction note: Correction to “A Moment Problem for Order Statistics”. Annals of Statistics, 1(4):797, July 1973. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176342478. See

[Kad71].

Kageyama:1971:IIB

[Kag71] S. Kageyama. An improved inequality for balanced incomplete block designs. Annals of Mathematical Statistics, 42(4):1448–1449, August 1971. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177693259.

Kageyama:1972:RAC

[Kag72] Sanpei Kageyama. On the reduction of associate classes for certain PBIB designs. Annals of Mathematical Statistics, 43(5):1528–1540, October 1972. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177692385.

Kailath:1967:MEW

[Kai67] Thomas Kailath. On measures equivalent to Wiener measure. Annals of Mathematical Statistics, 38(1):261–263, February 1967. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177699078.

Kailath:1971:SRN

[Kai71] Thomas Kailath. The structure of Radon–Nikodym derivatives with respect to Wiener and related measures. Annals of Mathematical Statistics, 42(3):1054–1067, June 1971. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177693332.

Kakumanu:1971:CDM

[Kak71] Prasadarao Kakumanu. Continuously discounted Markov decision model with countable state and action space. Annals of Mathematical Statistics, 42(3):919–926, June 1971. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177693321.

Kallianpur:1954:NRM

[Kal54] Gopinath Kallianpur. A note on the Robbins–Monro stochastic approximation method. Annals of Mathematical Statistics, 25(2):386– 388, June 1954. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/

1177728795. See [RM51].

Kallianpur:1959:NPP

[Kal59a] Gopinath Kallianpur. A note on perfect probability. Annals of Mathematical Statistics, 30(1):169–172, March 1959. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177706370.

Kallianpur:1959:POF

[Kal59b] Gopinath Kallianpur. A problem in optimum filtering with finite data. Annals of Mathematical Statistics, 30(3):659–669, September 1959. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177706196.

Kallianpur:1965:RUG

[Kal65] G. Kallianpur. Review: Ulf Grenander, Probabilities on Algebraic Structures. Annals of Mathematical Statistics, 36(1):345–347, February 1965. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177700302.

Kallianpur:1969:RLB

[Kal69] G. Kallianpur. Review: Leo Breiman, Probability. Annals of Mathematical Statistics, 40(3):1132–1133, June 1969. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177697627.

Kanter:1969:DTM

[Kan69] Marek Kanter. On distinguishing translates of measures. Annals of Mathematical Statistics, 40(5):1773–1777, October 1969. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177697389.

Kaplansky:1943:CND

[Kap43] Irving Kaplansky. A characterization of the normal distribution. Annals of Mathematical Statistics, 14(2):197–198, June 1943. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177731462.

Kaplansky:1945:ADR

[Kap45] Irving Kaplansky. The asymptotic distribution of runs of consecutive elements. Annals of Mathematical Statistics, 16(2):200–203, June 1945. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177731121.

Kapadia:1966:BSS

[Kap66] C. H. Kapadia. On the block structure of singular group divisible designs. Annals of Mathematical Statistics, 37(5):1398–1400, October 1966. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177699284.

Kappenman:1971:NMR

[Kap71] Russell F. Kappenman. A note on the multivariate t-ratio distribution. Annals of Mathematical Statistics, 42(1):349–351, February 1971. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177693518.

Karlin:1957:PTDa

[Kar57a] Samuel Karlin. Pólya type distributions, II. Annals of Mathematical Statistics, 28(2):281–308, June 1957. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177706960.

Karlin:1957:PTDb

[Kar57b] Samuel Karlin. Pólya-type distributions, III: Admissibility for multiaction problems. Annals of Mathematical Statistics, 28(4):839–860, December 1957. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177706789.

Karlin:1958:AEQ

[Kar58a] Samuel Karlin. Admissibility for estimation with quadratic loss. Annals of Mathematical Statistics, 29(2):406–436, June 1958. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177706620.

Karlin:1958:PTD

[Kar58b] Samuel Karlin. Pólya type distributions IV. Some principles of selecting a single procedure from a complete class. Annals of Mathematical Statistics, 29(1):1–21, March 1958. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177706703.

Karush:1961:CKE

[Kar61] Jack Karush. On the Chapman–Kolmogorov equation. Annals of Mathematical Statistics, 32(4):1333–1337, December 1961. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177704871.

Katz:1950:REB

[Kat50] Leo Katz. On the relative efficiencies of BAN estimates. Annals of Mathematical Statistics, 21(3):398–405, September 1950. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177729798.

Katz:1952:DNI

[Kat52] Leo Katz. The distribution of the number of isolates in a social group. Annals of Mathematical Statistics, 23(2):271–276, June 1952. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177729444.

Katz:1955:PIR

[Kat55] Leo Katz. Probability of indecomposability of a random mapping function. Annals of Mathematical Statistics, 26(3):512–517, September 1955. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177728496.

Katti:1960:MAD

[Kat60] S. K. Katti. Moments of the absolute difference and absolute deviation of discrete distributions. Annals of Mathematical Statistics, 31 (1):78–85, March 1960. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177705989.

Katti:1961:DLR

[Kat61a] S. K. Katti. Distribution of the likelihood ratio for testing multivariate linear hypotheses. Annals of Mathematical Statistics, 32(1):

333–335, March 1961. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177705162.

Katz:1961:AME

[Kat61b] Morris W. Katz. Admissible and minimax estimates of parameters in truncated spaces. Annals of Mathematical Statistics, 32(1):136– 142, March 1961. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177705146.

Katz:1963:NBE

[Kat63a] Melvin L. Katz. Note on the Berry–Esseen theorem. Annals of Mathematical Statistics, 34(3):1107–1108, September 1963. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177704037.

Katz:1963:PTD

[Kat63b] Melvin L. Katz. The probability in the tail of a distribution. Annals of Mathematical Statistics, 34(1):312–318, March 1963. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177704268.

Katz:1963:EOP

[Kat63c] Morris W. Katz. Estimating ordered probabilities. Annals of Mathematical Statistics, 34(3):967–972, September 1963. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177704019.

Katti:1967:IDI

[Kat67] S. K. Katti. Infinite divisibility of integer-valued random variables. Annals of Mathematical Statistics, 38(4):1306–1308, August 1967. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177698807.

Katz:1968:NWL

[Kat68] Melvin Katz. A note on the Weak Law of Large Numbers. Annals of Mathematical Statistics, 39(4):1348–1349, August 1968. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177698261.

Kavanagh:1941:NAO

[Kav41] Arthur J. Kavanagh. Note on the adjustment of observations. Annals of Mathematical Statistics, 12(1):111–114, March 1941. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177731791.

Kawada:1950:IQF

[Kaw50] Yukiyosi Kawada. Independence of quadratic forms in normally correlated variables. Annals of Mathematical Statistics, 21(4):614–615, December 1950. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177729758.

Kawata:1959:SCT

[Kaw59] T. Kawata. Some convergence theorems for stationary stochastic processes. Annals of Mathematical Statistics, 30(4):1192–1214, December 1959. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177706104.

Kramer:1957:IBA

[KB57] Clyde Young Kramer and Ralph Allan Bradley. Intra-block analysis for factorials in two-associate class group divisible designs. Annals of Mathematical Statistics, 28(2):349–361, June 1957. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177706963.

Kennedy:1971:MBP

[KB71] W. J. Kennedy and T. A. Bancroft. Model building for prediction in regression based upon repeated significance tests. Annals of Mathematical Statistics, 42(4):1273–1284, August 1971. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177693240.

Kulshreshtha:1972:BDU

[KDS72] A. C. Kulshreshtha, A. Dey, and G. M. Saha. Balanced designs with unequal replications and unequal block sizes. Annals of Mathematical Statistics, 43(4):1342–1345, August 1972. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177692487.

Keeping:1951:STE

[Kee51] E. S. Keeping. A significance test for exponential regression. Annals of Mathematical Statistics, 22(2):180–198, June 1951. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177729640.

Keilson:1962:QSS

[Kei62] Julian Keilson. Queues subject to service interruption. Annals of Mathematical Statistics, 33(4):1314–1322, December 1962. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177704364.

Keilson:1963:FPT

[Kei63] Julian Keilson. The first passage time density for homogeneous skipfree walks on the continuum. Annals of Mathematical Statistics, 34 (3):1003–1011, September 1963. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177704023.

Keilson:1966:LTP

[Kei66] Julian Keilson. A limit theorem for passage times in ergodic regenerative processes. Annals of Mathematical Statistics, 37(4):866–870, August 1966. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177699368.

Keilson:1969:MRF

[Kei69] J. Keilson. On the matrix renewal function for Markov renewal processes. Annals of Mathematical Statistics, 40(6):1901–1907, December 1969. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177697274.

Keilson:1972:TLC

[Kei72] J. Keilson. A threshold for log-concavity for probability generating functions and associated moment inequalities. Annals of Mathematical Statistics, 43(5):1702–1708, October 1972. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177692406.

Kelker:1971:IDV

[Kel71] Douglas Kelker. Infinite divisibility and variance mixtures of the normal distribution. Annals of Mathematical Statistics, 42(2):802– 808, April 1971. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177693436.

Kempthorne:1948:FAW

[Kem48] O. Kempthorne. The factorial approach to the weighing problem. Annals of Mathematical Statistics, 19(2):238–245, June 1948. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177730248.

Kempthorne:1953:CED

[Kem53] O. Kempthorne. A class of experimental designs using blocks of two plots. Annals of Mathematical Statistics, 24(1):76–84, March 1953. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177729084.

Kemperman:1956:GTL

[Kem56a] J. H. B. Kemperman. Generalized tolerance limits. Annals of Mathematical Statistics, 27(1):180–186, March 1956. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177728356.

Kempthorne:1956:EFI

[Kem56b] Oscar Kempthorne. The efficiency factor of an incomplete block design. Annals of Mathematical Statistics, 27(3):846–849, September 1956. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177728193.

Kemperman:1958:A

[Kem58] J. H. B. Kemperman. Appendix. Annals of Mathematical Statistics, 29(1):322–324, March 1958. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177706738.

Kemperman:1959:AES

[Kem59] J. H. B. Kemperman. Asymptotic expansions for the Smirnov test and for the range of cumulative sums. Annals of Mathematical Statistics, 30(2):448–462, June 1959. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177706262.

Kemperman:1963:WHT

[Kem63] J. H. B. Kemperman. A Wiener–Hopf type method for a general random walk with a two-sided boundary. Annals of Mathematical Statistics, 34(4):1168–1193, December 1963. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177703855.

Kempthorne:1964:RMC

[Kem64] Oscar Kempthorne. Review: M. C. Chakrabarti, Mathematics of Design and Analysis of Experiments. Annals of Mathematical Statistics, 35(2):911–912, June 1964. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177703601.

Kemp:1968:CJP

[Kem68a] L. F. Kemp, Jr. Construction of joint probability distributions. Annals of Mathematical Statistics, 39(4):1354–1357, August 1968. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177698263.

Kemperman:1968:GMP

[Kem68b] J. H. B. Kemperman. The general moment problem, a geometric approach. Annals of Mathematical Statistics, 39(1):93–122, February 1968. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177698508.

Kemperman:1969:ORT

[Kem69] J. H. B. Kemperman. On the optimum rate of transmitting information. Annals of Mathematical Statistics, 40(6):2156–2177, December 1969. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177697293.

Kemp:1970:GFC

[Kem70] Adrienne W. Kemp. General formulae for the central moments of certain serial correlation coefficient approximations. Annals of Mathematical Statistics, 41(4):1363–1368, August 1970. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177696915.

Kenney:1939:RST

[Ken39] J. F. Kenney. The regression systems of two sums having random elements in common. Annals of Mathematical Statistics, 10(1):70– 73, March 1939. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177732247.

Kendall:1940:CUP

[Ken40] M. G. Kendall. Conditions for uniqueness in the problem of moments. Annals of Mathematical Statistics, 11(4):402–409, December 1940. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177731827. See

corrections [Ken41].

Kendall:1941:CPU

[Ken41] M. G. Kendall. Corrections to a paper on the uniqueness problem of moments. Annals of Mathematical Statistics, 12(4):464–465, December 1941. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/

1177731685. See [Ken40].

Kendall:1948:GBD

[Ken48] David G. Kendall. On the generalized “birth-and-death” process. Annals of Mathematical Statistics, 19(1):1–15, March 1948. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177730285.

Kendall:1953:SPO

[Ken53] David G. Kendall. Stochastic processes occurring in the theory of queues and their analysis by the method of the imbedded Markov chain. Annals of Mathematical Statistics, 24(3):338–354, September 1953. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177728975.

Kendall:1964:RJR

[Ken64] David G. Kendall. Review: John Riordan, Stochastic Service Systems. Annals of Mathematical Statistics, 35(2):917–919, June 1964. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177703604.

Kennedy:1971:FCL

[Ken71] Douglas P. Kennedy. A functional Central Limit Theorem for kdimensional renewal theory. Annals of Mathematical Statistics, 42 (1):376–380, February 1971. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177693525.

Kennedy:1972:ERC

[Ken72] Douglas P. Kennedy. Estimates of the rates of convergence in limit theorems for the first passage times of random walks. Annals of Mathematical Statistics, 43(6):2090–2094, December 1972. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177690890.

Kerridge:1963:BFM

[Ker63] D. Kerridge. Bounds for the frequency of misleading Bayes inferences. Annals of Mathematical Statistics, 34(3):1109–1110, September 1963. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177704038.

Kesten:1958:ASA

[Kes58] Harry Kesten. Accelerated stochastic approximation. Annals of Mathematical Statistics, 29(1):41–59, March 1958. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177706705.

Kesten:1970:LPN

[Kes70] Harry Kesten. The limit points of a normalized random walk. Annals of Mathematical Statistics, 41(4):1173–1205, August 1970. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177696894.

Kesten:1972:RLS

[Kes72] Harry Kesten. The 1971 Rietz Lecture: Sums of independent random variables — without moment conditions. Annals of Mathematical Statistics, 43(3):701–732, June 1972. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177692541.

Keyfitz:1938:GTN

[Key38] Nathan Keyfitz. Graduation by a truncated normal. Annals of Mathematical Statistics, 9(1):66–67, March 1938. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177732362.

Kullback:1940:SSE

[KF40] S. Kullback and A. Frankel. A simple sampling experiment on confidence intervals. Annals of Mathematical Statistics, 11(2):209– 213, June 1940. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177731916.

Kingman:1970:NLC

[KG70] Albert Kingman and Franklin A. Graybill. A non-linear characterization of the normal distribution. Annals of Mathematical Statistics, 41 (6):1889–1895, December 1970. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177696690.

Killeen:1972:BTL

[KH72] Timothy J. Killeen and Thomas P. Hettmansperger. Bivariate tests for location and their Bahadur efficiencies. Annals of Mathematical Statistics, 43(5):1507–1516, October 1972. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177692383.

Khamis:1954:RMP

[Kha54] Salem H. Khamis. On the reduced moment problem. Annals of Mathematical Statistics, 25(1):113–122, March 1954. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177728850.

Khatri:1959:MIC

[Kha59] C. G. Khatri. On the mutual independence of certain statistics. Annals of Mathematical Statistics, 30(4):1258–1262, December 1959. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177706112. See

correction [Kha61].

Khatri:1961:CMI

[Kha61] C. G. Khatri. Correction to “On the Mutual Independence of Certain Statistics”. Annals of Mathematical Statistics, 32(4):1344, December 1961. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/

1177704878. See [Kha59].

Khatri:1962:CIG

[Kha62a] C. G. Khatri. A characterization of the inverse Gaussian distribution. Annals of Mathematical Statistics, 33(2):800–803, June 1962. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177704599.

Khatri:1962:CWI

[Kha62b] C. G. Khatri. Conditions for Wishartness and independence of second degree polynomials in a normal vector. Annals of Mathematical Statistics, 33(3):1002–1007, September 1962. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http:

//projecteuclid.org/euclid.aoms/1177704467. See corrections

[Kha73].

Khatri:1964:DGM

[Kha64a] C. G. Khatri. Distribution of the ‘generalised’ multiple correlation matrix in the dual case. Annals of Mathematical Statistics, 35(4): 1801–1806, December 1964. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177700402.

Khatri:1964:DLS

[Kha64b] C. G. Khatri. Distribution of the largest or the smallest characteristic root under null hypothesis concerning complex multivariate normal populations. Annals of Mathematical Statistics, 35(4):1807– 1810, December 1964. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177700403.

Khamis:1965:CNC

[Kha65a] Salem H. Khamis. Correction notes: Correction to “Some Basic Properties of the Incomplete Gamma Function Ratio”. Annals of Mathematical Statistics, 36(5):1584, October 1965. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177699923. See [Kha65b].

Khamis:1965:SBP

[Kha65b] Salem H. Khamis. Some basic properties of the incomplete gamma function ratio. Annals of Mathematical Statistics, 36(3):926–937; correction, 1584, June 1965. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.

aoms/1177700065. See correction [Kha65a].

Khatri:1965:CSA

[Kha65c] C. G. Khatri. Classical statistical analysis based on a certain multivariate complex Gaussian distribution. Annals of Mathematical Statistics, 36(1):98–114, February 1965. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177700274.

Khatri:1965:TRC

[Kha65d] C. G. Khatri. A test for reality of a covariance matrix in a certain complex Gaussian distribution. Annals of Mathematical Statistics, 36(1):115–119, February 1965. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177700275.

Khatri:1966:CDP

[Kha66] C. G. Khatri. On certain distribution problems based on positive definite quadratic functions in normal vectors. Annals of Mathematical Statistics, 37(2):468–479, April 1966. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177699530.

Khatri:1967:CIN

[Kha67a] C. G. Khatri. On certain inequalities for normal distributions and their applications to simultaneous confidence bounds. Annals of Mathematical Statistics, 38(6):1853–1867, December 1967. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177698618.

Khatri:1967:SDP

[Kha67b] C. G. Khatri. Some distribution problems connected with the characteristic roots of S_{1}S_{2}^{−1}. Annals of Mathematical Statistics, 38(3): 944–948, June 1967. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177698893.

Khan:1969:GMD

[Kha69] Rasul A. Khan. A general method of determining fixed-width confidence intervals. Annals of Mathematical Statistics, 40(2):704–709, April 1969. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177697747.

Khatri:1973:NCC

[Kha73] C. G. Khatri. Note: Corrections to “Conditions for Wishartness and Independence of Second Degree Polynomials in Normal Vector”. Annals of Statistics, 1(5):991, September 1973. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1176342522. See [Kha62b].

Killeen:1972:ETP

[KHS72] Timothy J. Killeen, Thomas P. Hettmansperger, and Gerald L. Sievers. An elementary theorem on the probability of large deviations. Annals of Mathematical Statistics, 43(1):181–192, February 1972. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177692712. See

correction [KHS74].

Killeen:1974:NCE

[KHS74] Timothy J. Killeen, Thomas P. Hettmansperger, and Gerald Sievers. Notes: Correction to “An Elementary Theorem on the Probability of Large Deviations”. Annals of Statistics, 2(6):1357, November 1974. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176342891. See

[KHS72].

Kiefer:1952:MVE

[Kie52a] J. Kiefer. On minimum variance estimators. Annals of Mathematical Statistics, 23(4):627–629, December 1952. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177729343.

Kiefer:1952:SME

[Kie52b] J. Kiefer. Sequential minimax estimation for the rectangular distribution with unknown range. Annals of Mathematical Statistics, 23 (4):586–593, December 1952. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177729337.

Kiefer:1953:CPW

[Kie53a] J. Kiefer. Correction of a proof: On Wald’s complete class theorems. Annals of Mathematical Statistics, 24(4):680, December 1953. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177728926. See

[Kie53b].

Kiefer:1953:WCC

[Kie53b] J. Kiefer. On Wald’s complete class theorems. Annals of Mathematical Statistics, 24(1):70–75, March 1953. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http:

//projecteuclid.org/euclid.aoms/1177729083. See correction

[Lie53a].

Kiefer:1957:IMS

[Kie57] J. Kiefer. Invariance, minimax sequential estimation, and continuous time processes. Annals of Mathematical Statistics, 28(3):573–601, September 1957. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177706874.

Kiefer:1958:NOR

[Kie58] J. Kiefer. On the nonrandomized optimality and randomized nonoptimality of symmetrical designs. Annals of Mathematical Statistics, 29(3):675–699, September 1958. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177706530.

Kiefer:1959:SAK

[Kie59] J. Kiefer. K-sample analogues of the Kolmogorov–Smirnov and Cramér–von Mises tests. Annals of Mathematical Statistics, 30(2): 420–447, June 1959. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177706261.

Kiefer:1961:ODR

[Kie61] J. Kiefer. Optimum designs in regression problems, II. Annals of Mathematical Statistics, 32(1):298–325, March 1961. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177705160.

Kiefer:1962:TMC

[Kie62] J. Kiefer. Two more criteria equivalent to D-optimality of designs. Annals of Mathematical Statistics, 33(2):792–796, June 1962. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177704597.

Kiefer:1964:RMG

[Kie64] J. Kiefer. Review: Maurice G. Kendall, Alan Stuart, The Advanced Theory of Statistics, Volume 2, “Inference and Relationship.”. Annals of Mathematical Statistics, 35(3):1371–1380, September 1964. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177703300.

Kiefer:1967:BRS

[Kie67] J. Kiefer. On Bahadur’s representation of sample quantiles. Annals of Mathematical Statistics, 38(5):1323–1342, October 1967. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177698690.

Kimball:1940:OPA

[Kim40] Bradford F. Kimball. Orthogonal polynomials applied to least square fitting of weighted observations. Annals of Mathematical Statistics, 11(3):348–352, September 1940. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177731872.

Kimball:1942:LTP

[Kim42] Bradford F. Kimball. Limited type of primary probability distribution applied to annual maximum flood flows. Annals of Mathematical Statistics, 13(3):318–325, September 1942. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177731570.

Kimball:1944:NAV

[Kim44] Bradford F. Kimball. Note on asymptotic value of probability distribution of sum of random variables which are greater than a set of arbitrarily chosen numbers. Annals of Mathematical Statistics, 15 (4):423–427, December 1944. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177731212.

Kimball:1946:SSE

[Kim46] Bradford F. Kimball. Sufficient statistical estimation functions for the parameters of the distribution of maximum values. Annals of Mathematical Statistics, 17(3):299–309, September 1946. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177730942.

Kimball:1947:SBT

[Kim47] Bradford F. Kimball. Some basic theorems for developing tests of fit for the case of the non-parametric probability distribution function, I. Annals of Mathematical Statistics, 18(4):540–548, December 1947. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177730344.

Kimball:1949:ASV

[Kim49] Bradford F. Kimball. An approximation to the sampling variance of an estimated maximum value of given frequency based on fit of doubly exponential distribution of maximum values. Annals of Mathematical Statistics, 20(1):110–113, March 1949. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177730097.

Kimball:1950:ADS

[Kim50] Bradford F. Kimball. On the asymptotic distribution of the sum of powers of unit frequency differences. Annals of Mathematical Statistics, 21(2):263–271, June 1950. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177729843.

Kimball:1951:DTS

[Kim51] A. W. Kimball. On dependent tests of significance in the analysis of variance. Annals of Mathematical Statistics, 22(4):600–602, December 1951. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177729552.

Kimball:1953:NCO

[Kim53] Bradford F. Kimball. Note on computation of orthogonal predictors. Annals of Mathematical Statistics, 24(2):299–303, June 1953. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177729037.

Kimball:1956:BCE

[Kim56] Bradford F. Kimball. The bias in certain estimates of the parameters of the extreme-value distribution. Annals of Mathematical Statistics, 27(3):758–767, September 1956. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177728182.

Kimura:1957:SPS

[Kim57] Motoo Kimura. Some problems of stochastic processes in genetics. Annals of Mathematical Statistics, 28(4):882–901, December 1957. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177706791.

Kimbleton:1968:CCS

[Kim68] Stephen R. Kimbleton. A characterization of certain sequences of norming constants. Annals of Mathematical Statistics, 39(2):394– 397, April 1968. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177698403.

Kimbleton:1969:SLT

[Kim69] Stephen R. Kimbleton. A stable limit theorem for Markov chains. Annals of Mathematical Statistics, 40(4):1467–1473, August 1969. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177697517.

King:1930:BMS

[Kin30] Wilford I. King. The Annals of Mathematical Statistics. Annals of Mathematical Statistics, 1(1):1–2, February 1930. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177733256.

Kincaid:1948:NEI

[Kin48a] W. M. Kincaid. Note on the error in interpolation of a function of two independent variables. Annals of Mathematical Statistics, 19(1): 85–88, March 1948. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.

aoms/1177730295. See remark [Gre50b].

Kincaid:1948:SEI

[Kin48b] W. M. Kincaid. Solution of equations by interpolation. Annals of Mathematical Statistics, 19(2):207–219, June 1948. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177730245.

King:1952:OCC

[Kin52] Edgar P. King. The operating characteristic of the control chart for sample means. Annals of Mathematical Statistics, 23(3):384–395, September 1952. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177729383.

Kingman:1961:TSQ

[Kin61a] J. F. C. Kingman. Two similar queues in parallel. Annals of Mathematical Statistics, 32(4):1314–1323, December 1961. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177704869.

Kinney:1961:CAP

[Kin61b] John R. Kinney. Correction and acknowledgment of priority to “First Passage Times for a Generalized Random Walk”. Annals of Mathematical Statistics, 32(4):1345, December 1961. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177704879. See [Kin61c].

Kinney:1961:FPT

[Kin61c] John R. Kinney. First passage times of a generalized random walk. Annals of Mathematical Statistics, 32(1):235–243, March 1961. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177705154. See correc-

tion and acknowledgment of priority [Kin61b].

Kinney:1962:TDT

[Kin62] John R. Kinney. A transient discrete time queue with finite storage. Annals of Mathematical Statistics, 33(1):130–136, March 1962. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177704718.

Kingman:1963:PCR

[Kin63a] J. F. C. Kingman. Poisson counts for random sequences of events. Annals of Mathematical Statistics, 34(4):1217–1232, December 1963. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177703857.

Kinney:1963:CHP

[Kin63b] J. R. Kinney. The convex Hull of plane Brownian motion. Annals of Mathematical Statistics, 34(1):327–329, March 1963. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177704270.

Kirkham:1935:MAA

[Kir35] W. J. Kirkham. Moments about the arithmetic mean of a binomial frequency distribution. Annals of Mathematical Statistics, 6(2):96– 101, June 1935. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177732614.

Kirkham:1937:NDM

[Kir37] William J. Kirkham. Note on the derivation of the multiple correlation coefficient. Annals of Mathematical Statistics, 8(1):68–71, March 1937. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177732453.

Kishen:1945:DEW

[Kis45] K. Kishen. On the design of experiments for weighing and making other types of measurements. Annals of Mathematical Statistics, 16 (3):294–300, September 1945. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177731093.

Kotz:1967:SRDa

[KJB67a] Samuel Kotz, N. L. Johnson, and D. W. Boyd. Series representations of distributions of quadratic forms in normal variables. I. Central case. Annals of Mathematical Statistics, 38(3):823–837, June 1967. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177698877.

Kotz:1967:SRDb

[KJB67b] Samuel Kotz, N. L. Johnson, and D. W. Boyd. Series representations of distributions of quadratic forms in normal variables II. Non-central case. Annals of Mathematical Statistics, 38(3):838–848, June 1967. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177698878.

Keilson:1960:TDQ

[KK60] J. Keilson and A. Kooharian. On time dependent queuing processes. Annals of Mathematical Statistics, 31(1):104–112, March 1960. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177705991.

Keilson:1962:GTD

[KK62] J. Keilson and A. Kooharian. On the general time dependent queue with a single server. Annals of Mathematical Statistics, 33(2):767– 791, June 1962. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177704596.

Kullback:1966:NMD

[KK66] S. Kullback and M. A. Khairat. A note on minimum discrimination information. Annals of Mathematical Statistics, 37(1):279–280, February 1966. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177699619.

Kac:1955:TNO

[KKW55] M. Kac, J. Kiefer, and J. Wolfowitz. On tests of normality and other tests of goodness of fit based on distance methods. Annals of Mathematical Statistics, 26(2):189–211, June 1955. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177728538.

Kullback:1951:IS

[KL51] S. Kullback and R. A. Leibler. On information and sufficiency. Annals of Mathematical Statistics, 22(1):79–86, March 1951. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177729694.

Kraft:1956:RRM

[KL56] C. Kraft and L. Le Cam. A remark on the roots of the maximum likelihood equation. Annals of Mathematical Statistics, 27(4):1174– 1177, December 1956. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177728087.

Kleyle:1971:ULP

[Kle71] Robert Kleyle. Upper and lower posterior probabilities for truncated means. Annals of Mathematical Statistics, 42(3):976–990, June 1971. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177693326.

Klimko:1967:GCT

[Kli67] Eugene M. Klimko. On the Glivenko–Cantelli theorem for infinite invariant measures. Annals of Mathematical Statistics, 38(4):1273– 1277, August 1967. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177698797.

Klimko:1969:UOE

[Kli69] E. M. Klimko. A uniform operator ergodic theorem. Annals of Mathematical Statistics, 40(3):1126–1129, June 1969. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177697625.

Klotz:1959:NDH

[Klo59] Jerome Klotz. Null distribution of the Hodges bivariate sign test. Annals of Mathematical Statistics, 30(4):1029–1033, December 1959.

CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177706086.

Klotz:1962:NTS

[Klo62] Jerome Klotz. Nonparametric tests for scale. Annals of Mathematical Statistics, 33(2):498–512, June 1962. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177704576.

Klotz:1963:SSP

[Klo63] Jerome Klotz. Small sample power and efficiency for the one sample Wilcoxon and normal scores tests. Annals of Mathematical Statistics, 34(2):624–632, June 1963. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177704175.

Klotz:1964:SSP

[Klo64] Jerome Klotz. Small sample power of the bivariate sign tests of Blumen and Hodges. Annals of Mathematical Statistics, 35(4):1576–1582, December 1964. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177700382.

Klotz:1965:AES

[Klo65] Jerome Klotz. Alternative efficiencies for signed rank tests. Annals of Mathematical Statistics, 36(6):1759–1766, December 1965. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177699804.

Klotz:1970:GDU

[Klo70] Jerome Klotz. The geometric density with unknown location parameter. Annals of Mathematical Statistics, 41(3):1078–1082, June 1970. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177696986.

Kesten:1959:PMD

[KM59] Harry Kesten and Norman Morse. A property of the multinomial distribution. Annals of Mathematical Statistics, 30(1):120–127, March 1959. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177706363.

Karlin:1967:PSM

[KM67] Samuel Karlin and James McGregor. Properties of the stationary measure of the critical case simple branching process. Annals of Mathematical Statistics, 38(4):977–991, August 1967. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177698766.

Kelker:1970:SSC

[KM70] Douglas Kelker and Ted K. Matthes. A sufficient statistics characterization of the normal distribution. Annals of Mathematical Statistics, 41(3):1086–1090, June 1970. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177696988.

Kalish:1971:ABS

[KM71] George Kalish and Piotr W. Mikulski. The asymptotic behavior of the Smirnov test compared to standard “Optimal Procedures”. Annals of Mathematical Statistics, 42(5):1742–1747, October 1971. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177693174.

Karlin:1959:NMS

[KMP59] S. Karlin, R. G. Miller, Jr., and N. U. Prabhu. Note on a moving single server problem. Annals of Mathematical Statistics, 30(1):243– 246, March 1959. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177706381.

Knight:1963:URP

[Kni63] William Knight. The use of the range in place of the standard deviation in Stein’s test. Annals of Mathematical Statistics, 34(1):346– 347, March 1963. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177704278.

Knight:1965:DAM

[Kni65a] Frank B. Knight. Distribution of the absolute maximum for certain Brownian motions. Annals of Mathematical Statistics, 36(1):311–312, February 1965. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177700293.

Knight:1965:LMI

[Kni65b] William Knight. A lemma for multiple inference. Annals of Mathematical Statistics, 36(6):1873–1874, December 1965. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177699820.

Knight:1965:MSE

[Kni65c] William Knight. A method of sequential estimation applicable to the hypergeometric, binomial, Poisson, and exponential distributions. Annals of Mathematical Statistics, 36(5):1494–1503, October 1965. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177699908.

Knight:1966:RPM

[Kni66] Frank Knight. Review: Paul A. Meyer, Probability and Potentials. Annals of Mathematical Statistics, 37(6):1856–1857, December 1966. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177699185.

Knight:1970:IDC

[Kni70] Frank B. Knight. An infinitesimal decomposition for a class of Markov processes. Annals of Mathematical Statistics, 41(5):1510– 1529, October 1970. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177696796.

Knight:1972:MPR

[Kni72] Frank B. Knight. On Markov processes with right-deterministic germ fields. Annals of Mathematical Statistics, 43(6):1968–1976, December 1972. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177690868.

Knott:1967:MCP

[Kno67] Martin Knott. Models for cataloguing problems. Annals of Mathematical Statistics, 38(4):1255–1260, August 1967. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177698794.

Kochen:1958:COS

[Koc58] Manfred Kochen. On the commutativity of operators in stochastic models for learning. Annals of Mathematical Statistics, 29(3):930– 933, September 1958. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177706555.

Koehn:1970:GCS

[Koe70] Uwe Koehn. Global cross sections and the densities of maximal invariants. Annals of Mathematical Statistics, 41(6):2045–2056, December 1970. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177696704.

Kolmogoroff:1941:CLU

[Kol41] A. Kolmogoroff. Confidence limits for an unknown distribution function. Annals of Mathematical Statistics, 12(4):461–463, December 1941. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177731684;   http://www.jstor.org/stable/2235958.

Konijn:1953:CCS

[Kon53] H. S. Konijn. On certain classes of statistical decision procedures. Annals of Mathematical Statistics, 24(3):440–448; corrections 25, 170– 171 (1954), September 1953. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177728983. See correction [Kon54].

Konijn:1954:CCC

[Kon54] H. S. Konijn. Correction to “On Certain Classes of Statistical Decision Procedures”. Annals of Mathematical Statistics, 25(1):170– 171, March 1954. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/

1177728860. See [Kon53].

Konijn:1956:PCT

[Kon56] H. S. Konijn. On the power of certain tests for independence in bivariate populations. Annals of Mathematical Statistics, 27(2):300– 323, June 1956. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/

1177728260. See correction [Kon58].

Konijn:1958:CPC

[Kon58] H. S. Konijn. Correction to “On the Power of Certain Tests for Independence in Bivariate Populations”. Annals of Mathematical Statistics, 29(3):935–936, September 1958. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177706558. See [Kon56].

Konijn:1966:SRO

[Kon66] H. S. Konijn. Statistical reproduction of orderings and translation subfamilies. Annals of Mathematical Statistics, 37(1):196–202, February 1966. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177699609.

Koopmans:1942:SCQ

[Koo42] Tjalling Koopmans. Serial correlation and quadratic forms in normal variables. Annals of Mathematical Statistics, 13(1):14–33, March 1942. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177731639.

Koopmans:1960:ARD

[Koo60] Lambert H. Koopmans. Asymptotic rate of discrimination for Markov processes. Annals of Mathematical Statistics, 31(4):982–994, December 1960. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177705671.

Koopmans:1961:EBS

[Koo61] L. H. Koopmans. An exponential bound on the strong law of large numbers for linear stochastic processes with absolutely convergent coefficients. Annals of Mathematical Statistics, 32(2):583– 586, June 1961. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177705063.

Koopmans:1964:CCW

[Koo64a] L. H. Koopmans. On the coefficient of coherence for weakly stationary stochastic processes. Annals of Mathematical Statistics, 35(2):532– 549, June 1964. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177703553.

Koopmans:1964:MAW

[Koo64b] L. H. Koopmans. On the multivariate analysis of weakly stationary stochastic processes. Annals of Mathematical Statistics, 35(4):1765– 1780, December 1964. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177700398.

Koopmans:1966:CRL

[Koo66] L. H. Koopmans. Convergence rates for the law of large numbers for linear combinations of Markov processes. Annals of Mathematical Statistics, 37(3):711–719, June 1966. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177699466.

Koop:1967:RIS

[Koo67] J. C. Koop. Replicated (or interpenetrating) samples of unequal sizes. Annals of Mathematical Statistics, 38(4):1142–1147, August 1967. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177698782.

Koopmans:1968:ERT

[Koo68] L. H. Koopmans. An extension of Rosen’s theorem to non-identically distributed random variables. Annals of Mathematical Statistics, 39 (3):897–904, June 1968. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177698322.

Koop:1971:SSS

[Koo71] J. C. Koop. On splitting a systematic sample for variance estimation. Annals of Mathematical Statistics, 42(3):1084–1087, June 1971. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177693336.

Kossack:1945:MC

[Kos45] Carl F. Kossack. On the mechanics of classification. Annals of Mathematical Statistics, 16(1):95–98, March 1945. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177731176.

Kotz:1961:EBP

[Kot61] Samuel Kotz. Exponential bounds on the probability of error for a discrete memoryless channel. Annals of Mathematical Statistics, 32 (2):577–582, June 1961. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177705062.

Kotlarski:1964:BRV

[Kot64] Ignacy Kotlarski. On bivariate random variables where the quotient of their coordinates follows some known distribution. Annals of Mathematical Statistics, 35(4):1673–1684, December 1964. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177700390.

Kotlarski:1965:GMT

[Kot65] Ignacy Kotlarski. On the generalized Mellin transform of a complex random variable and its applications. Annals of Mathematical Statistics, 36(5):1459–1467, October 1965. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177699905.

Kounias:1968:BPU

[Kou68] Eustratios G. Kounias. Bounds for the probability of a union, with applications. Annals of Mathematical Statistics, 39(6):2154–2158, December 1968. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177698049.

Koul:1969:ABW

[Kou69] Hira Lal Koul. Asymptotic behavior of Wilcoxon type confidence regions in multiple linear regression. Annals of Mathematical Statistics, 40(6):1950–1979, December 1969. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177697278.

Koul:1970:CAT

[Kou70a] Hira L. Koul. A class of ADF tests for subhypothesis in the multiple linear regression. Annals of Mathematical Statistics, 41(4):1273–1281, August 1970. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177696902.

Koul:1970:ANR

[Kou70b] Hira Lal Koul. Asymptotic normality of random rank statistics. Annals of Mathematical Statistics, 41(6):2144–2149, December 1970. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177696717.

Koul:1970:SCT

[Kou70c] Hira Lal Koul. Some convergence theorems for ranks and weighted empirical cumulatives. Annals of Mathematical Statistics, 41(5): 1768–1773, October 1970. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177696824.

Koul:1971:ABC

[Kou71] Hira Lal Koul. Asymptotic behavior of a class of confidence regions based on ranks in regression. Annals of Mathematical Statistics, 42 (2):466–476, April 1971. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177693398.

Koul:1972:SAR

[Kou72] Hira Lal Koul. Some asymptotic results on random rank statistics. Annals of Mathematical Statistics, 43(3):842–859, June 1972. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177692550.

Kraft:1972:ETD

[KOvE72] Charles H. Kraft, Ingram Olkin, and Constance van Eeden. Estimation and testing for differences in magnitude or displacement in the mean vectors of two multivariate normal populations. Annals of Mathematical Statistics, 43(2):455–467, April 1972. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177692626.

Kozakiewicz:1947:CSM

[Koz47] W. Kozakiewicz. On the convergence of sequences of moment generating functions. Annals of Mathematical Statistics, 18(1):61–69, March 1947. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177730492.

Kozakiewicz:1951:NSC

[Koz51] W. Kozakiewicz. On the necessary and sufficient conditions for the convergence of a sequence of moment generating functions. Annals of Mathematical Statistics, 22(3):478–480, September 1951. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177729601.

Kozelka:1956:AUP

[Koz56] Robert M. Kozelka. Approximate upper percentage points for extreme values in multinomial sampling. Annals of Mathematical Statistics, 27(2):507–512, June 1956. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177728273.

Krishnamoorthy:1951:MGT

[KP51] A. S. Krishnamoorthy and M. Parthasarathy. A multivariate gammatype distribution. Annals of Mathematical Statistics, 22(4):549–557, December 1951. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/

1177729544. See correction [KP60b].

Katz:1957:PDR

[KP57] Leo Katz and James H. Powell. Probability distributions of random variables associated with a structure of the sample space of sociometric investigations. Annals of Mathematical Statistics, 28(2): 442–448, June 1957. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177706972.

Karlin:1960:PTD

[KP60a] Samuel Karlin and Frank Proschan. Pólya type distributions of convolutions. Annals of Mathematical Statistics, 31(3):721–736, September 1960. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177705798.

Krishnamoorthy:1960:CNC

[KP60b] A. S. Krishnamoorthy and M. Parthasarathy. Correction notes: Correction to “A Multivariate Gamma-Type Distribution”. Annals of Mathematical Statistics, 31(1):229, March 1960. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177706004. See [KP51].

Khatri:1965:SRN

[KP65] C. G. Khatri and K. C. S. Pillai. Some results on the non-central multivariate beta distribution and moments of traces of two matrices. Annals of Mathematical Statistics, 36(5):1511–1520, October 1965. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177699910.

Khatri:1966:MTM

[KP66a] C. G. Khatri and K. C. S. Pillai. On the moments of the trace of a matrix and approximations to its non-central distribution. Annals of Mathematical Statistics, 37(5):1312–1318, October 1966. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177699275.

Kinney:1966:DPR

[KP66b] J. R. Kinney and T. S. Pitcher. Dimensional properties of a random distribution function on the square. Annals of Mathematical Statistics, 37(4):849–854, August 1966. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177699365.

Krishnaiah:1967:TEC

[KP67] P. R. Krishnaiah and P. K. Pathak. Tests for the equality of covariance matrices under the intraclass correlation model. Annals of Mathematical Statistics, 38(4):1286–1288, August 1967. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177698801. See correc-

tion [KP68c].

Khatri:1968:MES

[KP68a] C. G. Khatri and K. C. S. Pillai. On the moments of elementary symmetric functions of the roots of two matrices and approximations to a distribution. Annals of Mathematical Statistics, 39(4):1274–1281, August 1968. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177698252.

Khatri:1968:NCD

[KP68b] C. G. Khatri and K. C. S. Pillai. On the non-central distributions of two test criteria in multivariate analysis of variance. Annals of Mathematical Statistics, 39(1):215–226, February 1968. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177698521.

Krishnaiah:1968:CNC

[KP68c] P. R. Krishnaiah and P. K. Pathak. Correction notes: Corrections and comments on “Tests for the Equality of Covariance Matrices Under the Intraclass Correlation Model”. Annals of Mathematical Statistics, 39(4):1358, August 1968. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177698264. See [KP67].

Klotz:1969:MLE

[KP69] Jerome Klotz and Joseph Putter. Maximum likelihood estimation of multivariate covariance components for the balanced one-way layout. Annals of Mathematical Statistics, 40(3):1100–1105, June 1969.

CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177697618.

Krafft:1970:BPL

[KP70] Olaf Krafft and Detlef Plachky. Bounds for the power of likelihood ratio tests and their asymptotic properties. Annals of Mathematical Statistics, 41(5):1646–1654, October 1970. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177696808.

Kraft:1959:IPF

[KPS59] Charles H. Kraft, John W. Pratt, and A. Seidenberg. Intuitive probability on finite sets. Annals of Mathematical Statistics, 30(2):408– 419, June 1959. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177706260.

Koopmans:1972:EPB

[KQ72] L. H. Koopmans and C. Qualls. An exponential probability bound for the energy of a type of Gaussian process. Annals of Mathematical Statistics, 43(6):1953–1960, December 1972. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177690866.

Kaplansky:1945:MMR

[KR45] Irving Kaplansky and John Riordan. Multiple matching and runs by the symbolic method. Annals of Mathematical Statistics, 16(3):272– 277, September 1945. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177731090.

Koopmans:1950:ISC

[KR50] T. C. Koopmans and O. Reiersøl. The identification of structural characteristics. Annals of Mathematical Statistics, 21(2):165–181, June 1950. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177729837.

Karlin:1956:TDP

[KR56] Samuel Karlin and Herman Rubin. The theory of decision procedures for distributions with monotone likelihood ratio. Annals of Mathematical Statistics, 27(2):272–299, June 1956. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177728259.

Krishnaiah:1966:NRR

[KR66] P. R. Krishnaiah and M. Haseeb Rizvi. A note on recurrence relations between expected values of functions of order statistics. Annals of Mathematical Statistics, 37(3):733–734, June 1966. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177699471.

Kramer:1959:SMM

[Kra59] H. P. Kramer. Symmetrizable Markov matrices. Annals of Mathematical Statistics, 30(1):149–153, March 1959. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177706367.

Krauth:1971:LMP

[Kra71] J. Krauth. A locally most powerful tied rank test in a Wilcoxon situation. Annals of Mathematical Statistics, 42(6):1949–1956, December 1971. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177693062.

Krengel:1971:IET

[Kre71] Ulrich Krengel. On the individual ergodic theorem for subsequences. Annals of Mathematical Statistics, 42(3):1091–1095, June 1971. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177693338.

Krishnaiah:1968:STE

[Kri68] P. R. Krishnaiah. Simultaneous tests for the equality of covariance matrices against certain alternatives. Annals of Mathematical Statistics, 39(4):1303–1309, August 1968. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177698255.

Krishna:1970:PTL

[Kri70] J. Gopala Krishna. Probabilistic techniques leading to a Valirontype theorem in several complex variables. Annals of Mathematical Statistics, 41(6):2126–2129, December 1970. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177696711.

Krickeberg:1971:RCH

[Kri71a] K. Krickeberg. Review: C. A. Hayes, C. Y. Pauc, Derivation and Martingales. Annals of Mathematical Statistics, 42(5):1779, October 1971. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177693187.

Krishnaji:1971:NCP

[Kri71b] N. Krishnaji. Note on a characterizing property of the exponential distribution. Annals of Mathematical Statistics, 42(1):361–362, February 1971. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177693521.

Kruskal:1952:NTS

[Kru52] William H. Kruskal. A nonparametric test for the several sample problem. Annals of Mathematical Statistics, 23(4):525–540, December 1952. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177729332.

Kruskal:1954:MRT

[Kru54] William Kruskal. The monotonicity of the ratio of two noncentral t density functions. Annals of Mathematical Statistics, 25(1):162– 165, March 1954. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177728857.

Kruskal:1968:WGM

[Kru68] William Kruskal. When are Gauss–Markov and least squares estimators identical? a coordinate-free approach. Annals of Mathematical Statistics, 39(1):70–75, February 1968. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177698505.

Kendall:1939:PR

[KS39] M. G. Kendall and B. Babington Smith. The problem of m rankings. Annals of Mathematical Statistics, 10(3):275–287, September 1939. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177732186.

Kac:1947:ERS

[KS47] M. Kac and A. J. F. Siegert. An explicit representation of a stationary Gaussian process. Annals of Mathematical Statistics, 18(3):438–442, September 1947. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177730391.

Kemeny:1958:SMC

[KS58] John G. Kemeny and J. Laurie Snell. Semimartingales of Markov chains. Annals of Mathematical Statistics, 29(1):143–154, March 1958. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177706711.

Kac:1959:LEG

[KS59] Mark Kac and David Slepian. Large excursions of Gaussian processes. Annals of Mathematical Statistics, 30(4):1215–1228, December 1959. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177706105.

Kemeny:1961:MCP

[KS61] John G. Kemeny and J. Laurie Snell. On Markov chain potentials. Annals of Mathematical Statistics, 32(3):709–715, September 1961. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177704966.

Kamat:1962:APC

[KS62] A. R. Kamat and Y. S. Sathe. Asymptotic power of certain test criteria (based on first and second differences) for serial correlation between successive observations. Annals of Mathematical Statistics, 33(1):186–200, March 1962. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177704723.

Kiefer:1963:AOS

[KS63] J. Kiefer and J. Sacks. Asymptotically optimum sequential inference and design. Annals of Mathematical Statistics, 34(3):705–750, September 1963. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177704000.

Kiefer:1965:ABC

[KS65] J. Kiefer and R. Schwartz. Admissible Bayes character of T^{2}−, R^{2}−, and other fully invariant tests for classical multivariate normal problems. Annals of Mathematical Statistics, 36(3):747–770, June 1965. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177700051. See

correction [KS72b].

Karlin:1966:OED

[KS66a] Samuel Karlin and William J. Studden. Optimal experimental designs. Annals of Mathematical Statistics, 37(4):783–815, August 1966. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177699361.

Kesten:1966:ALT

[KS66b] H. Kesten and B. P. Stigum. Additional limit theorems for indecomposable multidimensional Galton–Watson processes. Annals of Mathematical Statistics, 37(6):1463–1481, December 1966. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177699139.

Kesten:1966:LTM

[KS66c] H. Kesten and B. P. Stigum. A limit theorem for multidimensional Galton–Watson processes. Annals of Mathematical Statistics, 37 (5):1211–1223, October 1966. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177699266.

Kallianpur:1968:ESS

[KS68] G. Kallianpur and C. Striebel. Estimation of stochastic systems: Arbitrary system process with additive white noise observation errors. Annals of Mathematical Statistics, 39(3):785–801, June 1968. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177698311.

Krengel:1969:NSI

[KS69] U. Krengel and L. Sucheston. Note on shift-invariant sets. Annals of Mathematical Statistics, 40(2):694–696, April 1969. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177697744.

Korwar:1970:ADU

[KS70] R. M. Korwar and R. J. Serfling. On averaging over distinct units in sampling with replacement. Annals of Mathematical Statistics, 41 (6):2132–2134, December 1970. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177696713.

Keilson:1972:FID

[KS72a] J. Keilson and F. W. Steutel. Families of infinitely divisible distributions closed under mixing and convolution. Annals of Mathematical Statistics, 43(1):242–250, February 1972. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177692717.

Kiefer:1972:CNC

[KS72b] J. Kiefer and R. Schwartz. Correction notes: Correction to “Admissible Bayes Character of T^{2}-, R^{2}-, and Other Fully Invariant Tests for Classical Multivariate Normal Problems”. Annals of Mathematical Statistics, 43(5):1742, October 1972. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177692413. See [KS65].

Korwar:1972:APA

[KS72c] R. M. Korwar and R. J. Serfling. Acknowledgment of priority to “On Averaging Over Distinct Units in Sampling with Replacement”. Annals of Mathematical Statistics, 43(5):1744, October 1972. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177692415.

Koul:1972:WCW

[KS72d] Hira Lal Koul and Robert G. Staudte, Jr. Weak convergence of weighted empirical cumulatives based on ranks. Annals of Mathematical Statistics, 43(3):832–841, June 1972. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177692549.

Kshirsagar:1958:NIB

[Ksh58] A. M. Kshirsagar. A note on incomplete block designs. Annals of Mathematical Statistics, 29(3):907–910, September 1958. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177706549.

Kshirsagar:1959:BDW

[Ksh59a] A. M. Kshirsagar. Bartlett decomposition and Wishart distribution. Annals of Mathematical Statistics, 30(1):239–241, March 1959. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177706379.

Kshirsagar:1959:DBA

[Ksh59b] A. M. Kshirsagar. Distribution of the “Blocks Adjusted for Treatments” sum of squares in incomplete block designs. Annals of Mathematical Statistics, 30(1):246–249, March 1959. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177706382.

Kshirsagar:1961:NCM

[Ksh61] A. M. Kshirsagar. The non-central multivariate beta distribution. Annals of Mathematical Statistics, 32(1):104–111, March 1961. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177705144.

Kounias:1969:IAS

[KsW69] Eustratios G. Kounias and Teng shan Weng. An inequality and almost sure convergence. Annals of Mathematical Statistics, 40(3): 1091–1093, June 1969. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177697615.

Karlin:1960:SP

[KT60a] Samuel Karlin and Donald Truax. Slippage problems. Annals of Mathematical Statistics, 31(2):296–324, June 1960. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177705895.

Katz:1960:EBF

[KT60b] Melvin Katz, Jr. and A. J. Thomasian. An exponential bound for functions of a Markov chain. Annals of Mathematical Statistics, 31 (2):470–474, June 1960. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177705910.

Katz:1961:BLL

[KT61] Melvin Katz, Jr. and A. J. Thomasian. A bound for the law of large numbers for discrete Markov processes. Annals of Mathematical Statistics, 32(1):336–337, March 1961. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177705163.

Kudo:1966:PWC

[Kud66] Hirokichi Kudō. On the property (w) of the class of statistical decision functions. Annals of Mathematical Statistics, 37(6):1631–1642, December 1966. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177699152.

Kuelbs:1968:IPL

[Kue68] J. Kuelbs. The invariance principle for a lattice of random variables. Annals of Mathematical Statistics, 39(2):382–389, April 1968. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177698401.

Kuiper:1959:APT

[Kui59] Nicolaas H. Kuiper. Alternative proof of a theorem of Birnbaum and Pyke. Annals of Mathematical Statistics, 30(1):251–252, March 1959. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177706384. See

[BP58].

Kullback:1934:ACF

[Kul34] Solomon Kullback. An application of characteristic functions to the distribution problem of statistics. Annals of Mathematical Statistics, 5(4):263–307, December 1934. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177732655.

Kullback:1935:NSC

[Kul35a] Solomon Kullback. A note on Sheppard’s corrections. Annals of Mathematical Statistics, 6(3):158–159, September 1935. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177732593.

Kullback:1935:NAV

[Kul35b] Solomon Kullback. A note on the analysis of variance. Annals of Mathematical Statistics, 6(2):76–77, June 1935. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177732610.

Kullback:1935:SMN

[Kul35c] Solomon Kullback. On samples from a multivariate normal population. Annals of Mathematical Statistics, 6(4):202–213, December 1935. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177732566.

Kullback:1936:DLD

[Kul36] Solomon Kullback. The distribution laws of the difference and quotient of variables independently distributed in Pearson Type III laws.

Annals of Mathematical Statistics, 7(1):51–53, March 1936. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177732546.

Kullback:1937:CDD

[Kul37] Solomon Kullback. On certain distributions derived from the multinomial distribution. Annals of Mathematical Statistics, 8(3):127–144, September 1937. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177732409.

Kullback:1939:NMP

[Kul39a] Solomon Kullback. Note on a matching problem. Annals of Mathematical Statistics, 10(1):77–80, March 1939. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177732249.

Kullback:1939:NNT

[Kul39b] Solomon Kullback. A note on Neyman’s theory of statistical estimation. Annals of Mathematical Statistics, 10(4):388–390, December 1939. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177732152.

Kullback:1946:RAM

[Kul46] S. Kullback. Report on the April meeting of the Washington chapter of the Institute. Annals of Mathematical Statistics, 17(2):254– 255, June 1946. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177730990.

Kullback:1947:CTB

[Kul47] S. Kullback. On the Charlier Type B series. Annals of Mathematical Statistics, 18(4):574–581, December 1947. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http:

//projecteuclid.org/euclid.aoms/1177730348. See correction

[Alb47].

Kullback:1948:CCT

[Kul48] S. Kullback. Correction to “On the Charlier Type B Series”. Annals of Mathematical Statistics, 19(3):427, September 1948. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177730214. See

[Kul47].

Kullback:1952:AIT

[Kul52] S. Kullback. An application of information theory to multivariate analysis. Annals of Mathematical Statistics, 23(1):88–102, March 1952. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177729487.

Kullback:1954:CII

[Kul54] S. Kullback. Certain inequalities in information theory and the Cramér–Rao inequality. Annals of Mathematical Statistics, 25(4): 745–751, December 1954. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177728660.

Kullback:1956:AIT

[Kul56a] S. Kullback. An application of information theory to multivariate analysis, II. Annals of Mathematical Statistics, 27(1):122–146, March 1956. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/

1177728353. See correction [Kul56b].

Kullback:1956:CAI

[Kul56b] S. Kullback. Correction to “An Application of Information Theory to Multivariate Analysis, II”. Annals of Mathematical Statistics, 27(3): 860, September 1956. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177728199. See [Kul56a].

Kulldorff:1963:EOT

[Kul63] Gunnar Kulldorff. Estimation of one or two parameters of the exponential distribution on the basis of suitably chosen order statistics. Annals of Mathematical Statistics, 34(4):1419–1431, December 1963. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177703874.

Kullback:1968:CNC

[Kul68a] S. Kullback. Correction notes: Correction to “Probability Densities with Given Marginals”. Annals of Mathematical Statistics, 39(6): 2161, December 1968. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177698052. See [Kul68b].

Kullback:1968:PDG

[Kul68b] S. Kullback. Probability densities with given marginals. Annals of Mathematical Statistics, 39(4):1236–1243, August 1968. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177698249. See correc-

tion [Kul68a].

Kullback:1969:BVG

[Kul69] S. Kullback. A bound for the variation of Gaussian densities. Annals of Mathematical Statistics, 40(6):2180–2182, December 1969. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177697295.

Kullback:1971:MHM

[Kul71] S. Kullback. Marginal homogeneity of multidimensional contingency tables. Annals of Mathematical Statistics, 42(2):594–606, April 1971. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177693409.

Kumar:1965:CNC

[Kum65] S. Kumar. Correction notes: Correction to abstract “A Group-Testing Problem”. Annals of Mathematical Statistics, 36(4):1318, August 1965. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/

1177700009. See [Ano65b].

Kunita:1967:NDS

[Kun67] Hiroshi Kunita. Note on Dynkin’s (α, ξ)-subprocess of standard Markov process. Annals of Mathematical Statistics, 38(6):1647–1654, December 1967. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177698598.

Kupperman:1958:CPH

[Kup58a] Morton Kupperman. Correction to “Probabilities of Hypotheses and Information-Statistics in Sampling From Exponential-Class Populations”. Annals of Mathematical Statistics, 29(4):1278, December 1958. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/

1177706462. See [Kup58b].

Kupperman:1958:PHI

[Kup58b] Morton Kupperman. Probabilities of hypotheses and informationstatistics in sampling from exponential-class populations. Annals of Mathematical Statistics, 29(2):571–575, June 1958. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http:

//projecteuclid.org/euclid.aoms/1177706633. See correction

[Kup58a].

Kurtz:1969:NSC

[Kur69] Thomas G. Kurtz. A note on sequences of continuous parameter Markov chains. Annals of Mathematical Statistics, 40(3):1078– 1082, June 1969. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177697612.

Kurtz:1970:AAD

[Kur70] Thomas G. Kurtz. Approximation of age dependent, multitype branching processes. Annals of Mathematical Statistics, 41(2):363– 368, April 1970. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177697076.

Kurtz:1971:CSM

[Kur71] Thomas G. Kurtz. Comparison of semi-Markov and Markov processes. Annals of Mathematical Statistics, 42(3):991–1002, June 1971. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177693327.

Kurtz:1972:ILL

[Kur72] Thomas G. Kurtz. Inequalities for the law of large numbers. Annals of Mathematical Statistics, 43(6):1874–1883, December 1972. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177690858.

Kushner:1966:NMS

[Kus66] Harold J. Kushner. A note on the maximum sample excursions of stochastic approximation processes. Annals of Mathematical Statistics, 37(2):513–516, April 1966. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177699536.

Kushner:1969:ASI

[Kus69] Harold J. Kushner. An application of the Sobolev imbedding theorems to criteria for the continuity of processes with a vector parameter. Annals of Mathematical Statistics, 40(2):517–526, April 1969. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177697719.

Kraft:1964:BBA

[KvE64a] Charles H. Kraft and Constance van Eeden. Bayesian bio-assay. Annals of Mathematical Statistics, 35(2):886–890, June 1964. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177703594.

Kraft:1964:FTD

[KvE64b] Charles H. Kraft and Constance van Eeden. A fluctuation theorem and a distribution-free test. Annals of Mathematical Statistics, 35 (3):1359–1360, September 1964. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177703296.

Kraft:1972:LRE

[KvE72] Charles H. Kraft and Constance van Eeden. Linearized rank estimates and signed-rank estimates for the general linear hypothesis. Annals of Mathematical Statistics, 43(1):42–57, February 1972. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177692699.

Kiefer:1952:SEM

[KW52] J. Kiefer and J. Wolfowitz. Stochastic estimation of the maximum of a regression function. Annals of Mathematical Statistics, 23(3):462– 466, September 1952. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177729392.

Kiefer:1956:CML

[KW56a] J. Kiefer and J. Wolfowitz. Consistency of the maximum likelihood estimator in the presence of infinitely many incidental parameters. Annals of Mathematical Statistics, 27(4):887–906, December 1956. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177728066.

Kiefer:1956:CGQ

[KW56b] J. Kiefer and J. Wolfowitz. On the characteristics of the general queueing process, with applications to random walk. Annals of Mathematical Statistics, 27(1):147–161, March 1956. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177728354.

Kiefer:1957:SPG

[KW57] J. Kiefer and Lionel Weiss. Some properties of generalized sequential probability ratio tests. Annals of Mathematical Statistics, 28(1):57– 74, March 1957. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177707037.

Kiefer:1959:AMC

[KW59a] J. Kiefer and J. Wolfowitz. Asymptotic minimax character of the sample distribution function for vector chance variables. Annals of Mathematical Statistics, 30(2):463–489, June 1959. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177706263.

Kiefer:1959:ODR

[KW59b] J. Kiefer and J. Wolfowitz. Optimum designs in regression problems. Annals of Mathematical Statistics, 30(2):271–294, June 1959. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177706252.

Kiefer:1965:THL

[KW65] J. Kiefer and J. Wolfowitz. On a theorem of Hoel and Levine on extrapolation designs. Annals of Mathematical Statistics, 36(6):1627– 1655, December 1965. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177699793.

Kimeldorf:1970:CBB

[KW70] George S. Kimeldorf and Grace Wahba. A correspondence between Bayesian estimation on stochastic processes and smoothing by splines. Annals of Mathematical Statistics, 41(2):495–502, April 1970. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177697089.

Kadota:1972:CTS

[KW72] T. T. Kadota and A. D. Wyner. Coding theorem for stationary, asymptotically memoryless, continuous-time channels. Annals of Mathematical Statistics, 43(5):1603–1611, October 1972. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177692392.

Kurkjian:1962:CFA

[KZ62] B. Kurkjian and M. Zelen. A calculus for factorial arrangements. Annals of Mathematical Statistics, 33(2):600–619, June 1962. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177704585.

Kander:1966:TPP

[KZ66] Z. Kander and S. Zacks. Test procedures for possible changes in parameters of statistical distributions occurring at unknown time points. Annals of Mathematical Statistics, 37(5):1196–1210, October 1966. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177699265.

Kailath:1971:ACR

[KZ71] Thomas Kailath and Moshe Zakai. Absolute continuity and Radon– Nikodym derivatives for certain measures relative to Wiener measure. Annals of Mathematical Statistics, 42(1):130–140, February 1971. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177693500.

Laderman:1939:DSR

[Lad39] Jack Laderman. The distribution of “Student’s” ratio for samples of two items drawn from non-normal universes. Annals of Mathematical Statistics, 10(4):376–379, December 1939. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177732149.

Laderman:1955:ABD

[Lad55] Jack Laderman. On the asymptotic behavior of decision procedures. Annals of Mathematical Statistics, 26(4):551–575, December 1955. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177728417.

Laha:1954:CGD

[Lah54] R. G. Laha. On a characterisation of the gamma distribution. Annals of Mathematical Statistics, 25(4):784–787, December 1954. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177728666.

Laha:1956:CSL

[Lah56a] R. G. Laha. On a characterization of the stable law with finite expectation. Annals of Mathematical Statistics, 27(1):187–195, March 1956. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177728357.

Laha:1956:SIT

[Lah56b] R. G. Laha. On the stochastic independence of two second-degree polynomial statistics in normally distributed variates. Annals of Mathematical Statistics, 27(3):790–796, September 1956. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177728185.

Laha:1957:CND

[Lah57a] R. G. Laha. On a characterization of the normal distribution from properties of suitable linear statistics. Annals of Mathematical Statistics, 28(1):126–139, March 1957. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177707041.

Laha:1957:SCP

[Lah57b] R. G. Laha. On some characterization problems connected with linear structural relations. Annals of Mathematical Statistics, 28(2):405– 414, June 1957. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177706968.

Laha:1958:FTT

[Lah58] R. G. Laha. On a factorization theorem in the theory of analytic characteristic functions. Annals of Mathematical Statistics, 29(3):922– 926, September 1958. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177706553.

Laha:1959:LCG

[Lah59] R. G. Laha. On the laws of Cauchy and Gauss. Annals of Mathematical Statistics, 30(4):1165–1174, December 1959. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177706102.

Lamperti:1960:FPM

[Lam60] John Lamperti. The first-passage moments and the invariant measure of a Markov chain. Annals of Mathematical Statistics, 31(2):515– 517, June 1960. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177705918.

Lamperti:1962:IPR

[Lam62] John Lamperti. An invariance principle in renewal theory. Annals of Mathematical Statistics, 33(2):685–696, June 1962. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177704590.

Lamperti:1963:CSP

[Lam63] John Lamperti. On a class of stochastic processes. Annals of Mathematical Statistics, 34(1):206–212, March 1963. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177704256.

Lamperti:1964:EOS

[Lam64] John Lamperti. On extreme order statistics. Annals of Mathematical Statistics, 35(4):1726–1737, December 1964. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177700395.

Lamperti:1965:LTG

[Lam65] John Lamperti. On limit theorems for Gaussian processes. Annals of Mathematical Statistics, 36(1):304–310, February 1965. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177700292.

Lamb:1971:NHF

[Lam71a] Charles W. Lamb. A note on harmonic functions and martingales. Annals of Mathematical Statistics, 42(6):2044–2049, December 1971. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177693072.

Lamb:1971:CCT

[Lam71b] Charles W. Lamb. On the construction of certain transition functions. Annals of Mathematical Statistics, 42(2):439–450, April 1971. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177693396.

LaMotte:1972:NCM

[LaM72] Lynn Roy LaMotte. Notes on the covariance matrix of a random, nested anova model. Annals of Mathematical Statistics, 43(2):659– 662, April 1972. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177692648.

Landau:1944:NVB

[Lan44] H. G. Landau. Note on the variance and best estimates. Annals of Mathematical Statistics, 15(2):219–221, June 1944. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177731289.

Lancaster:1958:SBD

[Lan58] H. O. Lancaster. The structure of bivariate distributions. Annals of Mathematical Statistics, 29(3):719–736, September 1958. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177706532. See correc-

tion [Lan64].

Lancaster:1963:CCD

[Lan63a] H. O. Lancaster. Correlation and complete dependence of random variables. Annals of Mathematical Statistics, 34(4):1315–1321, December 1963. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177703867.

Lancaster:1963:CCF

[Lan63b] H. O. Lancaster. Correlations and canonical forms of bivariate distributions. Annals of Mathematical Statistics, 34(2):532–538, June 1963. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177704165.

Lancaster:1964:CNC

[Lan64] H. O. Lancaster. Correction notes: Correction to “The Structure of Bivariate Distributions”. Annals of Mathematical Statistics, 35(3):

1388, September 1964. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177703304. See [Lan58].

Lancaster:1965:PSI

[Lan65] H. O. Lancaster. Pairwise statistical independence. Annals of Mathematical Statistics, 36(4):1313–1317, August 1965. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177700007.

Land:1971:CIL

[Lan71] Charles E. Land. Confidence intervals for linear functions of the normal mean and variance. Annals of Mathematical Statistics, 42 (4):1187–1205, August 1971. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177693235.

Landers:1972:ECM

[Lan72a] D. Landers. Existence and consistency of modified minimum contrast estimates. Annals of Mathematical Statistics, 43(1):74–83, February 1972. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177692702.

Lanke:1972:SDE

[Lan72b] Jan Lanke. Sampling distinguishable elements with replacement. Annals of Mathematical Statistics, 43(4):1329–1332, August 1972. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177692485.

Larguier:1935:CCU

[Lar35] Everett H. Larguier. On certain coefficients used in mathematical statistics. Annals of Mathematical Statistics, 6(4):220–226, December 1935. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177732568.

Larguier:1936:MEM

[Lar36] Everett H. Larguier. On a method for evaluating the moments of a Bernoulli distribution. Annals of Mathematical Statistics, 7(4):191– 195, December 1936. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177732479.

Larsen:1939:MAA

[Lar39] Harold D. Larsen. Moments about the arithmetic mean of a hypergeometric frequency distribution. Annals of Mathematical Statistics, 10 (2):198–201, June 1939. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177732219.

Laubscher:1960:NND

[Lau60] Nico F. Laubscher. Normalizing the noncentral t and F distributions. Annals of Mathematical Statistics, 31(4):1105–1112, December 1960. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177705682.

Laurent:1963:CDO

[Lau63] Andre G. Laurent. Conditional distribution of order statistics and distribution of the reduced i th order statistic of the exponential model. Annals of Mathematical Statistics, 34(2):652–657, June 1963. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177704177.

Lawther:1933:EPT

[Law33] H. P. Lawther. The extended probability theory for the continuous variable with particular application to the linear distribution. Annals of Mathematical Statistics, 4(4):241–262, November 1933. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177732754.

Lawrence:1955:RT

[Law55] Charles Lawrence. Report of the Treasurer for 1954. Annals of Mathematical Statistics, 26(2):361–364, June 1955. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177728557.

Lawley:1963:TSC

[Law63] D. N. Lawley. On testing a set of correlation coefficients for equality. Annals of Mathematical Statistics, 34(1):149–151, March 1963. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177704249.

Lawton:1965:SIC

[Law65] William H. Lawton. Some inequalities for central and non-central distributions. Annals of Mathematical Statistics, 36(5):1521–1525, October 1965. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177699911.

Lawton:1968:CRQ

[Law68] William H. Lawton. Concentration of random quotients. Annals of Mathematical Statistics, 39(2):466–480, April 1968. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177698410.

Lawless:1971:NCT

[Law71] J. F. Lawless. A note on certain types of BIBD’s balanced for residual effects. Annals of Mathematical Statistics, 42(4):1439–1441, August 1971. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177693256.

Layard:1972:LST

[Lay72] M. W. J. Layard. Large sample tests for the equality of two covariance matrices. Annals of Mathematical Statistics, 43(1):123–141, February 1972. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177692708.

Larson:1963:SMB

[LB63] Harold J. Larson and T. A. Bancroft. Sequential model building for prediction in regression analysis, I. Annals of Mathematical Statistics, 34(2):462–479, June 1963. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177704158.

Leadbetter:1965:MNC

[LC65] M. R. Leadbetter and J. D. Cryer. On the mean number of curve crossings by non-stationary normal processes. Annals of Mathematical Statistics, 36(2):509–516, April 1965. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177700160.

Lawing:1966:LRC

[LD66] William D. Lawing and H. T. David. Likelihood ratio computations of operating characteristics. Annals of Mathematical Statistics, 37 (6):1704–1716, December 1966. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177699160.

LeCam:1954:NTL

[Le 54] Lucien Le Cam. Note on a theorem of Lionel Weiss. Annals of Mathematical Statistics, 25(4):791–794, December 1954. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177728668. See [Wei53b].

LeCam:1955:EWT

[Le 55] L. Le Cam. An extension of Wald’s theory of statistical decision functions. Annals of Mathematical Statistics, 26(1):69–81, March 1955. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177728594.

LeCam:1970:AUP

[Le 70a] L. Le Cam. On the assumptions used to prove asymptotic normality of maximum likelihood estimates. Annals of Mathematical Statistics, 41(3):802–828, June 1970. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177696960.

LeCam:1970:WCP

[Le 70b] Lucien Le Cam. On the weak convergence of probability measures. Annals of Mathematical Statistics, 41(2):621–625, April 1970. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177697102.

Leadbetter:1966:CLC

[Lea66] M. R. Leadbetter. On crossings of levels and curves by a wide class of stochastic processes. Annals of Mathematical Statistics, 37(1):260– 267, February 1966. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177699615.

Lechner:1962:ODP

[Lec62] J. A. Lechner. Optimum decision procedures for a Poisson process parameter. Annals of Mathematical Statistics, 33(4):1384–1402, December 1962. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177704371.

Ledyard:1971:PMS

[Led71] J. O. Ledyard. A pseudo-metric space of probability measures and the existence of measurable utility. Annals of Mathematical Statistics, 42 (2):794–798, April 1971. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177693434.

Lee:1964:AIT

[Lee64] P. M. Lee. On the axioms of information theory. Annals of Mathematical Statistics, 35(1):415–418, March 1964. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177703765.

Lehmer:1944:ITP

[Leh44] Emma Lehmer. Inverse tables of probabilities of errors of the second kind. Annals of Mathematical Statistics, 15(4):388–398, December 1944. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177731209.

Lehmann:1947:FAT

[Leh47a] E. L. Lehmann. On families of admissible tests. Annals of Mathematical Statistics, 18(1):97–104, March 1947. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177730496.

Lehmann:1947:OTC

[Leh47b] E. L. Lehmann. On optimum tests of composite hypotheses with one constraint. Annals of Mathematical Statistics, 18(4):473–494, December 1947. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177730340.

Lehmann:1950:SPT

[Leh50] E. L. Lehmann. Some principles of the theory of testing hypotheses. Annals of Mathematical Statistics, 21(1):1–26, March 1950. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177729884.

Lehmann:1951:CUC

[Leh51a] E. L. Lehmann. Consistency and unbiasedness of certain nonparametric tests. Annals of Mathematical Statistics, 22(2):165–179, June 1951. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177729639.

Lehmann:1951:GCU

[Leh51b] E. L. Lehmann. A general concept of unbiasedness. Annals of Mathematical Statistics, 22(4):587–592, December 1951. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177729549.

Lehmann:1952:ELF

[Leh52a] E. L. Lehmann. On the existence of least favorable distributions. Annals of Mathematical Statistics, 23(3):408–416, September 1952. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177729385.

Lehmann:1952:TMH

[Leh52b] E. L. Lehmann. Testing multiparameter hypotheses. Annals of Mathematical Statistics, 23(4):541–552, December 1952. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177729333.

Lehmann:1953:PRT

[Leh53] E. L. Lehmann. The power of rank tests. Annals of Mathematical Statistics, 24(1):23–43, March 1953. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177729080.

Lehmann:1955:OFD

[Leh55] E. L. Lehmann. Ordered families of distributions. Annals of Mathematical Statistics, 26(3):399–419, September 1955. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177728487.

Lehmann:1957:TSMa

[Leh57a] E. L. Lehmann. A theory of some multiple decision problems, I. Annals of Mathematical Statistics, 28(1):1–25, March 1957. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177707034.

Lehmann:1957:TSMb

[Leh57b] E. L. Lehmann. A theory of some multiple decision problems. II. Annals of Mathematical Statistics, 28(3):547–572, September 1957. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177706873.

Lehmann:1958:SLP

[Leh58] E. L. Lehmann. Significance level and power. Annals of Mathematical Statistics, 29(4):1167–1176, December 1958. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177706448.

Lehmann:1959:OIT

[Leh59] E. L. Lehmann. Optimum invariant tests. Annals of Mathematical Statistics, 30(4):881–884, December 1959. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177706073.

Lehmann:1961:SMP

[Leh61] E. L. Lehmann. Some Model I problems of selection. Annals of Mathematical Statistics, 32(4):990–1012, December 1961. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177704841.

Lehmann:1963:ANI

[Leh63a] E. L. Lehmann. Asymptotically nonparametric inference: An alternative approach to linear models. Annals of Mathematical Statistics, 34 (4):1494–1506, December 1963. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177703881.

Lehmann:1963:NCI

[Leh63b] E. L. Lehmann. Nonparametric confidence intervals for a shift parameter. Annals of Mathematical Statistics, 34(4):1507–1512, December 1963. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177703882.

Lehmann:1963:REA

[Leh63c] E. L. Lehmann. Robust estimation in analysis of variance. Annals of Mathematical Statistics, 34(3):957–966, September 1963. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177704018.

Lehner:1963:ODR

[Leh63d] G. Lehner. One dimensional random walk with a partially reflecting barrier. Annals of Mathematical Statistics, 34(2):405–412, June 1963. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177704151.

Lehmann:1964:ANI

[Leh64] E. L. Lehmann. Asymptotically nonparametric inference in some linear models with one observation per cell. Annals of Mathematical Statistics, 35(2):726–734, June 1964. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177703570.

Lehmann:1966:TBG

[Leh66a] E. L. Lehmann. On a theorem of Bahadur and Goodman. Annals of Mathematical Statistics, 37(1):1–6, February 1966. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177699593.

Lehmann:1966:SCD

[Leh66b] E. L. Lehmann. Some concepts of dependence. Annals of Mathematical Statistics, 37(5):1137–1153, October 1966. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177699260.

Leipnik:1947:DSC

[Lei47] R. B. Leipnik. Distribution of the serial correlation coefficient in a circularly correlated universe. Annals of Mathematical Statistics, 18 (1):80–87, March 1947. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177730494.

Lengyel:1939:THT

[Len39] B. A. Lengyel. On testing the hypothesis that two samples have been drawn from a common normal population. Annals of Mathematical Statistics, 10(4):365–375, December 1939. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177732148.

Lentner:1969:GLS

[Len69] M. M. Lentner. Generalized least-squares estimation of a subvector of parameters in randomized fractional factorial experiments. Annals of Mathematical Statistics, 40(4):1344–1352, August 1969. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177697506.

LePage:1970:RKL

[LeP70] Raoul LePage. Review: Kai Lai Chung, A Course in Probability Theory. Annals of Mathematical Statistics, 41(2):746–748, April 1970.

CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177697129.

Lev:1949:PBC

[Lev49a] Joseph Lev. The point biserial coefficient of correlation. Annals of Mathematical Statistics, 20(1):125–126, March 1949. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177730103.

Levene:1949:MPA

[Lev49b] Howard Levene. On a matching problem arising in genetics. Annals of Mathematical Statistics, 20(1):91–94, March 1949. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177730093.

Levene:1952:PFT

[Lev52] Howard Levene. On the power function of tests of randomness based on runs up and down. Annals of Mathematical Statistics, 23(1):34– 56, March 1952. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177729484.

LeVeque:1966:RJK

[LeV66a] W. J. LeVeque. Review: J. Kubilius, Gretchen Burgie, Susan Schuur, Probabilistic Methods in the Theory of Numbers. Annals of Mathematical Statistics, 37(4):1063–1065, August 1966. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177699391.

Levine:1966:PMV

[Lev66b] A. Levine. A problem in minimax variance polynomial extrapolation. Annals of Mathematical Statistics, 37(4):898–903, August 1966. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177699371.

Levitan:1971:GDR

[Lev71] Michael L. Levitan. A generalized Doeblin ratio limit theorem. Annals of Mathematical Statistics, 42(3):904–911, June 1971. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177693319.

Lewis:1935:RSC

[Lew35] W. T. Lewis. A reconsideration of Sheppard’s corrections. Annals of Mathematical Statistics, 6(1):11–20, March 1935. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177732637.

Lewis:1961:DAD

[Lew61] Peter A. W. Lewis. Distribution of the Anderson–Darling statistic. Annals of Mathematical Statistics, 32(4):1118–1124, December 1961. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177704850.

Leysieffer:1967:FFM

[Ley67] Frederick W. Leysieffer. Functions of finite Markov chains. Annals of Mathematical Statistics, 38(1):206–212, February 1967. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177699072.

Leysieffer:1970:SMC

[Ley70] Frederick W. Leysieffer. A system of Markov chains with random life times. Annals of Mathematical Statistics, 41(2):576–584, April 1970. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177697097.

Littell:1970:RBB

[LF70] Ramon C. Littell and J. Leroy Folks. On the relation between Bahadur efficiency and power. Annals of Mathematical Statistics, 41 (6):2135–2136, December 1970. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177696714.

Lieberman:1953:NDC

[Lie53a] Gerald J. Lieberman. A note on Dodge’s continuous inspection plan. Annals of Mathematical Statistics, 24(3):480–484, September 1953. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177728988. See

[Dod43].

Lieblein:1953:EEV

[Lie53b] Julius Lieblein. On the exact evaluation of the variances and covariances of order statistics in samples from the extreme-value distribution. Annals of Mathematical Statistics, 24(2):282–287, June 1953.

CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177729034.

Lieblein:1955:MOS

[Lie55] Julius Lieblein. On moments of order statistics from the Weibull distribution. Annals of Mathematical Statistics, 26(2):330–333, June 1955. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177728551.

Liggett:1970:CSR

[Lig70] Thomas M. Liggett. Convergence of sums of random variables conditioned on a future change of sign. Annals of Mathematical Statistics, 41(6):1978–1982, December 1970. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177696698.

Liggett:1971:AOS

[Lig71] W. S. Liggett, Jr. On the asymptotic optimality of spectral analysis for testing hypotheses about time series. Annals of Mathematical Statistics, 42(4):1348–1358, August 1971. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177693247.

Link:1950:SDR

[Lin50] Richard F. Link. The sampling distribution of the ratio of two ranges from independent samples. Annals of Mathematical Statistics, 21(1): 112–116, March 1950. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.

aoms/1177729890. See correction [Lin52].

Link:1952:CSD

[Lin52] Richard F. Link. Correction to “The Sampling Distribution of the Ratio of Two Ranges from Independent Samples”. Annals of Mathematical Statistics, 23(2):298–299, June 1952. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177729451. See [Lin50].

Lindley:1956:MIP

[Lin56] D. V. Lindley. On a measure of the information provided by an experiment. Annals of Mathematical Statistics, 27(4):986–1005, December 1956. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177728069.

Linhart:1960:MPP

[Lin60] H. Linhart. A measure of predictive precision in regression analysis. Annals of Mathematical Statistics, 31(2):399–404, June 1960. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177705902.

Lindley:1964:BAC

[Lin64] Dennis V. Lindley. The Bayesian analysis of contingency tables. Annals of Mathematical Statistics, 35(4):1622–1643, December 1964. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177700386.

Lindley:1966:RRM

[Lin66] D. V. Lindley. Review: Richard von Mises, Hilda Geiringer, Mathematical Theory of Probability and Statistics. Annals of Mathematical Statistics, 37(3):747–754, June 1966. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177699479.

Lindgren:1970:SPN

[Lin70] George Lindgren. Some properties of a normal process near a local maximum. Annals of Mathematical Statistics, 41(6):1870–1883, December 1970. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177696688.

Lin:1972:SRL

[Lin72] Michael Lin. Strong ratio limit theorems for Markov processes. Annals of Mathematical Statistics, 43(2):569–579, April 1972. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177692637.

Lippman:1971:MAR

[Lip71] Steven A. Lippman. Maximal average-reward policies for semi-Markov decision processes with arbitrary state and action space. Annals of Mathematical Statistics, 42(5):1717–1726, October 1971. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177693170.

Littauer:1948:RNY

[Lit48] S. B. Littauer. Report on the New York meeting of the Institute. Annals of Mathematical Statistics, 19(2):290–291, June 1948. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177730259.

Littauer:1949:RNY

[Lit49] S. B. Littauer. Report on the New York meeting of the Institute. Annals of Mathematical Statistics, 20(2):325–326, June 1949. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177730048.

Littauer:1952:RBM

[Lit52] S. B. Littauer. Report of the Boston meeting of the Institute. Annals of Mathematical Statistics, 23(1):155–159, March 1952. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177729496.

Littell:1972:ECT

[Lit72] Ramon C. Littell. On the efficiency of a competitor of the twosample Kolmogorov–Smirnov and Kuiper tests. Annals of Mathematical Statistics, 43(6):1991–1992, December 1972. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177690871.

Lukacs:1954:PND

[LK54] Eugene Lukacs and Edgar P. King. A property of the normal distribution. Annals of Mathematical Statistics, 25(2):389–394, June 1954. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177728796.

Lowan:1939:DET

[LL39] Arnold N. Lowan and Jack Laderman. On the distribution of errors in N^{th} tabular differences. Annals of Mathematical Statistics, 10 (4):360–364, December 1939. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177732147.

Laha:1960:ISC

[LLN60] R. G. Laha, E. Lukacs, and M. Newman. On the independence of a sample central moment and the sample mean. Annals of Mathematical Statistics, 31(4):1028–1033, December 1960. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177705675.

Loeve:1950:FLT

[Loè50] M. Loève. Fundamental limit theorems of probability theory. Annals of Mathematical Statistics, 21(3):321–338, September 1950. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177729793.

Lohrding:1969:TET

[Loh69] R. K. Lohrding. A test of equality of two normal population means assuming homogeneous coefficients of variation. Annals of Mathematical Statistics, 40(4):1374–1385, August 1969. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177697509.

Lomnicki:1952:SEG

[Lom52] Z. A. Lomnicki. The standard error of Gini’s mean difference. Annals of Mathematical Statistics, 23(4):635–637, December 1952. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177729346.

Lonseth:1942:SLE

[Lon42] A. T. Lonseth. Systems of linear equations with coefficients subject to error. Annals of Mathematical Statistics, 13(3):332–337, September 1942. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177731572.

Lonseth:1944:RES

[Lon44] A. T. Lonseth. On relative errors in systems of linear equations. Annals of Mathematical Statistics, 15(3):323–325, September 1944. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177731239.

Lord:1954:DDH

[Lor54] R. D. Lord. The distribution of distance in a hypersphere. Annals of Mathematical Statistics, 25(4):794–798, December 1954. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177728669.

Lord:1959:JCT

[Lor59] Frederic M. Lord. The joint cumulants of true values and errors of measurement. Annals of Mathematical Statistics, 30(4):1000–1004, December 1959. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177706082.

Lorden:1967:IRA

[Lor67] Gary Lorden. Integrated risk of asymptotically Bayes sequential tests. Annals of Mathematical Statistics, 38(5):1399–1422, October 1967. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177698696.

Lorden:1970:EB

[Lor70] Gary Lorden. On excess over the boundary. Annals of Mathematical Statistics, 41(2):520–527, April 1970. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177697092.

Lorden:1971:PRC

[Lor71] G. Lorden. Procedures for reacting to a change in distribution. Annals of Mathematical Statistics, 42(6):1897–1908, December 1971. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177693055.

Lorden:1972:LRT

[Lor72] Gary Lorden. Likelihood ratio tests for sequential k-decision problems. Annals of Mathematical Statistics, 43(5):1412–1427, October 1972. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177692374.

Lotka:1939:CTS

[Lot39a] Alfred J. Lotka. A contribution to the theory of self-renewing aggregates, with special reference to industrial replacement. Annals of Mathematical Statistics, 10(1):1–25, March 1939. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177732243.

Lotka:1939:IEP

[Lot39b] Alfred J. Lotka. On an integral equation in population analysis. Annals of Mathematical Statistics, 10(2):144–161, June 1939. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177732213.

Lotka:1942:PEP

[Lot42] Alfred J. Lotka. The progeny of an entire population. Annals of Mathematical Statistics, 13(2):115–126, June 1942. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177731600.

Lotka:1948:ARS

[Lot48] Alfred J. Lotka. Application of recurrent series in renewal theory. Annals of Mathematical Statistics, 19(2):190–206, June 1948. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177730244.

Loynes:1962:SWT

[Loy62] R. M. Loynes. Stationary waiting-time distributions for single-server queues. Annals of Mathematical Statistics, 33(4):1323–1339, December 1962. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177704365.

Loynes:1965:EVU

[Loy65] R. M. Loynes. Extreme values in uniformly mixing stationary stochastic processes. Annals of Mathematical Statistics, 36(3):993– 999, June 1965. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177700071.

Loynes:1966:IM

[Loy66] R. M. Loynes. On idempotent matrices. Annals of Mathematical Statistics, 37(1):295–296, February 1966. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177699624.

Loynes:1969:CCS

[Loy69] R. M. Loynes. The consistency of certain sequential estimators. Annals of Mathematical Statistics, 40(2):568–574, April 1969. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177697724.

Li:1970:AED

[LPC70] Hung C. Li, K. C. S. Pillai, and Tseng C. Chang. Asymptotic expansions for distributions of the roots of two matrices from classical and complex Gaussian populations. Annals of Mathematical Statistics, 41 (5):1541–1556, October 1970. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177696798.

Loftsgaarden:1965:NEM

[LQ65] D. O. Loftsgaarden and C. P. Quesenberry. A nonparametric estimate of a multivariate density function. Annals of Mathematical Statistics, 36(3):1049–1051, June 1965. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177700079.

Lotkin:1953:SEA

[LR53] Mark Lotkin and Russell Remage. Scaling and error analysis for matrix inversion by partitioning. Annals of Mathematical Statistics, 24(3):428–439, September 1953. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177728982.

Landers:1972:SPT

[LR72a] D. Landers and L. Rogge. Short proofs of two convergence theorems for conditional expectations. Annals of Mathematical Statistics, 43 (4):1372–1373, August 1972. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177692493.

Landers:1972:MSF

[LR72b] Dieter Landers and Lothar Rogge. Minimal sufficient σ-fields and minimal sufficient statistics. Two counterexamples. Annals of Mathematical Statistics, 43(6):2045–2049, December 1972. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177690882.

Lind:1972:RQM

[LR72c] Bruce Lind and George Roussas. A remark on quadratic mean differentiability. Annals of Mathematical Statistics, 43(3):1030–1034, June 1972. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177692570.

Lehmann:1948:MPT

[LS48] E. L. Lehmann and C. Stein. Most powerful tests of composite hypotheses. I. Normal distributions. Annals of Mathematical Statistics, 19(4):495–516, December 1948. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177730147.

Lehmann:1949:TSN

[LS49] E. L. Lehmann and C. Stein. On the theory of some nonparametric hypotheses. Annals of Mathematical Statistics, 20(1): 28–45, March 1949. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177730089.

Lehmann:1950:CSC

[LS50] E. L. Lehmann and Charles Stein. Completeness in the sequential case. Annals of Mathematical Statistics, 21(3):376–385, September 1950. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177729796.

Lehmann:1953:ACI

[LS53] E. L. Lehmann and C. M. Stein. The admissibility of certain invariant statistical tests involving a translation parameter. Annals of Mathematical Statistics, 24(3):473–479, September 1953. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177728987.

Lieberman:1955:MLC

[LS55] Gerald J. Lieberman and Herbert Solomon. Multi-level continuous sampling plans. Annals of Mathematical Statistics, 26(4):686–704, December 1955. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177728428.

LeCam:1960:NSC

[LS60] Lucien Le Cam and Lorraine Schwartz. A necessary and sufficient condition for the existence of consistent estimates. Annals of Mathematical Statistics, 31(1):140–150, March 1960. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177705993.

Lehrer:2008:TRB

[LS08] Ehud Lehrer and Eran Shmaya. Two remarks on Blackwell’s theorem. Journal of Applied Probability, 45(2):580–586, June 2008. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL

http://www.jstor.org/stable/27595967. See [Bla53].

Luce:1967:SCE

[Luc67] R. Duncan Luce. Sufficient conditions for the existence of a finitely additive probability measure. Annals of Mathematical Statistics, 38 (3):780–786, June 1967. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177698871.

Luce:1968:NRQ

[Luc68] R. Duncan Luce. On the numerical representation of qualitative conditional probability. Annals of Mathematical Statistics, 39(2):481– 491, April 1968. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177698411.

Lukomski:1939:SPM

[Luk39] J. Lukomski. On some properties of multidimensional distributions. Annals of Mathematical Statistics, 10(3):236–246, September 1939. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177732182.

Lukacs:1942:CND

[Luk42] Eugene Lukacs. A characterization of the normal distribution. Annals of Mathematical Statistics, 13(1):91–93, March 1942. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177731647.

Lukacs:1952:SIS

[Luk52] Eugene Lukacs. The stochastic independence of symmetric and homogeneous linear and quadratic statistics. Annals of Mathematical Statistics, 23(3):442–449, September 1952. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177729389.

Lukacs:1955:CGD

[Luk55] Eugene Lukacs. A characterization of the gamma distribution. Annals of Mathematical Statistics, 26(2):319–324, June 1955. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177728549.

Lukacs:1957:RCC

[Luk57] Eugene Lukacs. Remarks concerning characteristic functions. Annals of Mathematical Statistics, 28(3):717–723, September 1957. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177706883.

Lukacs:1965:RYV

[Luk65] E. Lukacs. Review: Yu. V. Linnik, M. L. Gruel, Decomposition des Lois de Probabilities; Yu. V. Linnik, S. J. Taylor, Decomposition of Probability Distributions. Annals of Mathematical Statistics, 36(1): 347–350, February 1965. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177700303.

Luther:1965:DSM

[Lut65] Norman Y. Luther. Decomposition of symmetric matrices and distributions of quadratic forms. Annals of Mathematical Statistics, 36 (2):683–690, April 1965. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177700176.

Levene:1944:CMR

[LW44] H. Levene and J. Wolfowitz. The covariance matrix of runs up and down. Annals of Mathematical Statistics, 15(1):58–69, March 1944. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177731314;   http://www.jstor.org/stable/2236211.

Lomnicki:1957:EAT

[LZ57] Z. A. Lomnicki and S. K. Zaremba. On the estimation of autocorrelation in time series. Annals of Mathematical Statistics, 28(1): 140–158, March 1957. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177707042.

MacStewart:1941:NPS

[Mac41] W. Mac Stewart. A note on the power of the sign test. Annals of Mathematical Statistics, 12(2):236–239, June 1941. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177731755.

Maceda:1948:CGP

[Mac48] E. Cansado Maceda. On the compound and generalized Poisson distributions. Annals of Mathematical Statistics, 19(3):414–416, September 1948. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177730208.

MacKay:1959:AET

[Mac59] John H. MacKay. Asymptotically efficient tests based on the sums of observations. Annals of Mathematical Statistics, 30(3):806–813, September 1959. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177706209.

MacQueen:1961:PS

[Mac61] James B. MacQueen. A problem in survival. Annals of Mathematical Statistics, 32(2):605–610, June 1961. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177705067.

MacQueen:1967:IDM

[Mac67] J. MacQueen. An intrinsically determined Markov chain. Annals of Mathematical Statistics, 38(3):934–936, June 1967. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177698890.

MacNeill:1971:LPC

[Mac71] Ian B. MacNeill. Limit processes for co-spectral and quadrature spectral distribution functions. Annals of Mathematical Statistics, 42(1): 81–96, February 1971. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177693497.

Madow:1937:CTC

[Mad37] William G. Madow. Contributions to the theory of comparative statistical analysis. I. Fundamental theorems of comparative analysis. Annals of Mathematical Statistics, 8(3):159–176, September 1937. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177732412.

Madow:1940:DQF

[Mad40a] William G. Madow. The distribution of quadratic forms in noncentral normal random variables. Annals of Mathematical Statistics, 11(1):100–103, March 1940. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177731947.

Madow:1940:LDQ

[Mad40b] William G. Madow. Limiting distributions of quadratic and bilinear forms. Annals of Mathematical Statistics, 11(2):125–146, June 1940.

CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177731908.

Madow:1944:RWM

[Mad44] W. G. Madow. Report on the Washington meeting of the Institute. Annals of Mathematical Statistics, 15(2):226–227, June 1944. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177731291.

Madow:1945:NDS

[Mad45] William G. Madow. Note on the distribution of the serial correlation coefficient. Annals of Mathematical Statistics, 16(3):308–310, September 1945. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177731096.

Madow:1948:SDB

[Mad48a] William G. Madow. On a source of downward bias in the analysis of variance and covariance. Annals of Mathematical Statistics, 19 (3):351–359, September 1948. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177730199.

Madow:1948:LDE

[Mad48b] William G. Madow. On the limiting distributions of estimates based on samples from finite universes. Annals of Mathematical Statistics, 19(4):535–545, December 1948. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177730149.

Madow:1949:TSS

[Mad49] William G. Madow. On the theory of systematic sampling, II. Annals of Mathematical Statistics, 20(3):333–354, September 1949. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177729988.

Madow:1953:TSS

[Mad53] William G. Madow. On the theory of systematic sampling, III. Comparison of centered and random start systematic sampling. Annals of Mathematical Statistics, 24(1):101–106, March 1953. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177729087.

Madansky:1959:BEC

[Mad59] Albert Madansky. Bounds on the expectation of a convex function of a multivariate random variable. Annals of Mathematical Statistics, 30(3):743–746, September 1959. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177706203.

Madansky:1963:RMZ

[Mad63] Albert Madansky. Review: Marvin Zelen, Statistical Theory of Reliability. Annals of Mathematical Statistics, 34(4):1613–1614, December 1963. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177703900.

Madsen:1971:NSE

[Mad71] Richard W. Madsen. A note on some ergodic theorems of A. Paz. Annals of Mathematical Statistics, 42(1):405–408, February 1971. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177693534. See

[Paz70].

Mahamunulu:1963:SVE

[Mah63] D. M. Mahamunulu. Sampling variances of the estimates of variance components in the unbalanced 3-way nested classification. Annals of Mathematical Statistics, 34(2):521–527, June 1963. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177704163.

Mahamunulu:1967:SFS

[Mah67] D. M. Mahamunulu. Some fixed-sample ranking and selection problems. Annals of Mathematical Statistics, 38(4):1079–1091, August 1967. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177698778.

Maitra:1966:NUD

[Mai66] Ashok Maitra. A note on undiscounted dynamic programming. Annals of Mathematical Statistics, 37(4):1042–1044, August 1966. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177699385.

Maitra:1969:NPD

[Mai69] Ashok Maitra. A note on positive dynamic programming. Annals of Mathematical Statistics, 40(1):316–318, February 1969. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177697829.

Majumdar:1953:STC

[Maj53] Kulendra N. Majumdar. On some theorems in combinatorics relating to incomplete block designs. Annals of Mathematical Statistics, 24 (3):377–389, September 1953. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177728978.

Majindar:1962:IBM

[Maj62a] Kulendra N. Majindar. Improved bounds on a measure of skewness. Annals of Mathematical Statistics, 33(3):1192–1194, September 1962. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177704482.

Majindar:1962:PIB

[Maj62b] Kulendra N. Majindar. On the parameters and intersection of blocks of balanced incomplete block designs. Annals of Mathematical Statistics, 33(3):1200–1205, September 1962. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177704485.

Malmquist:1947:SPC

[Mal47] Sten Malmquist. A statistical problem connected with the counting of radioactive particles. Annals of Mathematical Statistics, 18(2): 255–264, June 1947. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177730441.

Malmquist:1954:CCC

[Mal54] Sten Malmquist. On certain confidence contours for distribution functions. Annals of Mathematical Statistics, 25(3):523–533, September 1954. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177728720.

Malik:1968:EDP

[Mal68a] Henrick John Malik. Exact distribution of the product of independent generalized gamma variables with the same shape parameter.

Annals of Mathematical Statistics, 39(5):1751–1752, October 1968. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177698159.

Mallows:1968:ICR

[Mal68b] C. L. Mallows. An inequality in constrained random variables. Annals of Mathematical Statistics, 39(3):1080–1082, June 1968. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177698341.

Malinvaud:1970:CNR

[Mal70] E. Malinvaud. The consistency of nonlinear regressions. Annals of Mathematical Statistics, 41(3):956–969, June 1970. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177696972.

Mallows:1972:NAJ

[Mal72] C. L. Mallows. A note on asymptotic joint normality. Annals of Mathematical Statistics, 43(2):508–515, April 1972. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177692631.

Mallows:1990:CWJ

[Mal90] Colin L. Mallows, editor. The collected works of John W. Tukey. Volume VI. More mathematical: 1938–1984. Wadsworth & Brooks/Cole Statistics/Probability Series. Wadsworth & Brooks/Cole Advanced Books & Software, Monterey, CA, USA, 1990. ISBN 0-534-05103-0. lxxii + 644 + i17 pp. With a preface by William S. Cleveland, and a biography by Frederick Mosteller.

Mann:1942:COL

[Man42] Henry B. Mann. The construction of orthogonal Latin squares. Annals of Mathematical Statistics, 13(4):418–423, December 1942. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177731539.

Mann:1943:CSO

[Man43] H. B. Mann. On the construction of sets of orthogonal latin squares. Annals of Mathematical Statistics, 14(4):401–414, December 1943. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177731360.

Mann:1945:NPC

[Man45a] H. B. Mann. Note on a paper by C. W. Cotterman and L. H. Snyder. Annals of Mathematical Statistics, 16(3):311–312, September 1945. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177731097. See

[CS39].

Mann:1945:PEO

[Man45b] Henry B. Mann. On a problem of estimation occurring in public opinion polls. Annals of Mathematical Statistics, 16(1):85–90, March 1945. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/

1177731174. See correction [Man46].

Mann:1945:TRB

[Man45c] Henry B. Mann. On a test for randomness based on signs of differences. Annals of Mathematical Statistics, 16(2):193–199, June 1945. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177731120.

Mann:1946:CPP

[Man46] H. B. Mann. Correction to the paper “On a Problem of Estimation Occurring in Public Opinion Polls”. Annals of Mathematical Statistics, 17(1):87–88, March 1946. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177731029. See [Man45b].

Mantel:1953:EBN

[Man53] Nathan Mantel. An extension of the Buffon needle problem. Annals of Mathematical Statistics, 24(4):674–677, December 1953. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177728924.

Mandel:1958:NCI

[Man58] John Mandel. A note on confidence intervals in regression problems. Annals of Mathematical Statistics, 29(3):903–907, September 1958. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177706548.

Mann:1960:ALH

[Man60] Henry B. Mann. The algebra of a linear hypothesis. Annals of Mathematical Statistics, 31(1):1–15, March 1960. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177705982.

Mandelbrot:1962:RSE

[Man62] Benoît Mandelbrot. The role of sufficiency and of estimation in thermodynamics. Annals of Mathematical Statistics, 33(3):1021–1038, September 1962. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177704470;   http://www.jstor.org/stable/2237878.

Mann:1969:NBI

[Man69a] Henry B. Mann. A note on balanced incomplete block designs. Annals of Mathematical Statistics, 40(2):679–680, April 1969. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177697740.

Mann:1969:OEL

[Man69b] Nancy R. Mann. Optimum estimators for linear functions of location and scale parameters. Annals of Mathematical Statistics, 40(6):2149– 2155, December 1969. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177697292.

Marks:1948:LBE

[Mar48] Eli S. Marks. A lower bound for the expected travel among m random points. Annals of Mathematical Statistics, 19(3):419–422, September 1948. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177730210.

Marshall:1958:SSD

[Mar58] A. W. Marshall. The small sample distribution of nω_{n}^{2}. Annals of Mathematical Statistics, 29(1):307–309, March 1958. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177706733.

Marshall:1960:OSA

[Mar60] Albert W. Marshall. A one-sided analog of Kolmogorov’s inequality. Annals of Mathematical Statistics, 31(2):483–487, June 1960. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177705912.

Marsaglia:1961:ERV

[Mar61a] G. Marsaglia. Expressing a random variable in terms of uniform random variables. Annals of Mathematical Statistics, 32(3):894–898, September 1961. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177704983;   http://www.jstor.org/stable/2237849.

Marsaglia:1961:GER

[Mar61b] G. Marsaglia. Generating exponential random variables. Annals of Mathematical Statistics, 32(3):899–900, September 1961. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177704984;   http:
//www.jstor.org/stable/2237850.

Mardia:1962:MPD

[Mar62] K. V. Mardia. Multivariate Pareto distributions. Annals of Mathematical Statistics, 33(3):1008–1015, September 1962. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http:

//projecteuclid.org/euclid.aoms/1177704468. See correction

[Mar63].

Mardia:1963:CNC

[Mar63] K. V. Mardia. Correction notes: Correction to “Multivariate Pareto Distributions”. Annals of Mathematical Statistics, 34(4):1603, December 1963. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/

1177703897. See [Mar62].

Mardia:1964:SRO

[Mar64] K. V. Mardia. Some results on the order statistics of the multivariate normal and Pareto type 1 populations. Annals of Mathematical Statistics, 35(4):1815–1818, December 1964. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177700405.

Marlow:1965:FD

[Mar65] W. H. Marlow. Factorial distributions. Annals of Mathematical Statistics, 36(3):1066–1068, June 1965. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177700084.

Martin:1969:UCF

[Mar69] N. F. G. Martin. Uniform convergence of families of martingales. Annals of Mathematical Statistics, 40(3):1071–1073, June 1969. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177697610.

Marcus:1970:BDM

[Mar70] Michael Marcus. A bound for the distribution of the maximum of continuous Gaussian processes. Annals of Mathematical Statistics, 41(1):305–309, February 1970. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177697209.

Marsaglia:1971:RVI

[Mar71a] George Marsaglia. Random variables with independent binary digits. Annals of Mathematical Statistics, 42(6):1922–1929, December 1971. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177693058;   http://www.jstor.org/stable/2240118.

Marshall:1971:RSK

[Mar71b] Albert W. Marshall. Review: Samuel Karlin, Total Positivity. Annals of Mathematical Statistics, 42(5):1779–1781, October 1971. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177693188.

Marcus:1972:UBA

[Mar72a] M. B. Marcus. Upper bounds for the asymptotic maxima of continuous Gaussian processes. Annals of Mathematical Statistics, 43(2): 522–533, April 1972. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177692633.

Marsaglia:1972:CPS

[Mar72b] George Marsaglia. Choosing a point from the surface of a sphere. Annals of Mathematical Statistics, 43(2):645–646, April 1972. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177692644;   http:
//www.jstor.org/stable/2240001.

Massey:1950:NPN

[Mas50a] F. J. Massey, Jr. A note on the power of a non-parametric test. Annals of Mathematical Statistics, 21(3):440–443, September 1950.

CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177729803. See

correction [Mas52a].

Massey:1950:NED

[Mas50b] Frank J. Massey, Jr. A note on the estimation of a distribution function by confidence limits. Annals of Mathematical Statistics, 21(1): 116–119, March 1950. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177729891;   http://www.jstor.org/stable/2236559.

Massey:1951:DMD

[Mas51a] Frank J. Massey, Jr. The distribution of the maximum deviation between two sample cumulative step functions. Annals of Mathematical Statistics, 22(1):125–128, March 1951. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177729703.

Massey:1951:NTS

[Mas51b] Frank J. Massey, Jr. A note on a two sample test. Annals of Mathematical Statistics, 22(2):304–306, June 1951. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177729653.

Massey:1952:CNP

[Mas52a] F. J. Massey, Jr. Correction to “A Note on the Power of a Nonparametric Test”. Annals of Mathematical Statistics, 23(4):637–638, December 1952. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/

1177729347. See [Mas50a].

Massey:1952:DTD

[Mas52b] Frank J. Massey, Jr. Distribution table for the deviation between two sample cumulatives. Annals of Mathematical Statistics, 23(3):435– 441, September 1952. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177729388.

Mason:1970:CSC

[Mas70a] J. David Mason. Convergence of sums to a convolution of stable laws. Annals of Mathematical Statistics, 41(3):1068–1070, June 1970. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177696983.

Mason:1970:CSL

[Mas70b] J. David Mason. Convolutions of stable laws as limit distributions of partial sums. Annals of Mathematical Statistics, 41(1):101–114, February 1970. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177697192.

Masani:1971:RYR

[Mas71a] P. Masani. Review: Yu. A. Rozanov, Stationary Random Processes. Annals of Mathematical Statistics, 42(4):1463–1467, August 1971. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177693265.

Mason:1971:LTS

[Mas71b] J. David Mason. Local theorems in strengthened form for lattice random variables. Annals of Mathematical Statistics, 42(2):784– 788, April 1971. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177693432.

Mason:1972:AAL

[Mas72] J. David Mason. Addendum: Addendum to “Local Theorems in Strengthened Form for Lattice Random Variables”. Annals of Mathematical Statistics, 43(2):696, April 1972. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177692659.

Mathisen:1943:MTH

[Mat43] Harold C. Mathisen. A method of testing the hypothesis that two samples are from the same population. Annals of Mathematical Statistics, 14(2):188–194, June 1943. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177731460.

Matern:1949:INN

[Mat49] Bertil Matérn. Independence of non-negative quadratic forms in normally correlated variables. Annals of Mathematical Statistics, 20(1): 119–120, March 1949. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177730100.

Matusita:1955:DRB

[Mat55] Kameo Matusita. Decision rules, based on the distance, for problems of fit, two samples, and estimation. Annals of Mathematical Statistics, 26(4):631–640, December 1955. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177728422.

Matthes:1963:OSP

[Mat63] Theodore K. Matthes. On the optimality of sequential probability ratio tests. Annals of Mathematical Statistics, 34(1):18–21, March 1963. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177704239.

Mauchly:1940:STS

[Mau40] John W. Mauchly. Significance test for sphericity of a normal nvariate distribution. Annals of Mathematical Statistics, 11(2):204– 209, June 1940. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177731915.

Mauldon:1959:GBD

[Mau59] J. G. Mauldon. A generalization of the beta-distribution. Annals of Mathematical Statistics, 30(2):509–520, June 1959. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177706265.

Mazumdar:1967:CCC

[Maz67] Sati Mazumdar. On the construction of cyclic collineations for obtaining a balanced set of L-restrictional prime-powered lattice designs. Annals of Mathematical Statistics, 38(4):1293–1295, August 1967. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177698803.

Mickey:1966:BDF

[MB66] M. Ray Mickey and Morton B. Brown. Bounds on the distribution functions of the Behrens–Fisher statistic. Annals of Mathematical Statistics, 37(3):639–642, June 1966. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177699457.

Mazuy:1965:STW

[MC65] Kay Knight Mazuy and W. S. Connor. Student’s t in a two-way classification with unequal variances. Annals of Mathematical Statistics, 36(4):1248–1255, August 1965. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177699996.

McCarthy:1939:ATR

[McC39] M. D. McCarthy. On the application of the Z-test to randomized blocks. Annals of Mathematical Statistics, 10(4):337–359, December 1939. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177732146.

McCarthy:1947:ASM

[McC47] Philip J. McCarthy. Approximate solutions for means and variances in a certain class of Box problems. Annals of Mathematical Statistics, 18(3):349–383, September 1947. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177730386.

McCord:1964:AME

[McC64] James R. McCord. On asymptotic moments of extreme statistics. Annals of Mathematical Statistics, 35(4):1738–1745, December 1964. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177700396.

McCall:1965:RRB

[McC65] J. J. McCall. Review: Richard E. Barlow, Frank Proschan, Larry C. Hunter, Mathematical Theory of Reliability. Annals of Mathematical Statistics, 36(6):1902–1904, December 1965. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177699826.

McCabe:1969:ABC

[McC69] B. J. McCabe. The asymptotic behavior of a certain Markov chain. Annals of Mathematical Statistics, 40(2):665–666, April 1969. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177697736.

McCabe:1972:SEP

[McC72] George P. McCabe, Jr. Sequential estimation of a Poisson integer mean. Annals of Mathematical Statistics, 43(3):803–813, June 1972. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177692546.

McFadden:1955:UMC

[McF55] J. A. McFadden. URN models of correlation and a comparison with the multivariate normal integral. Annals of Mathematical Statistics, 26(3):478–489, September 1955. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177728492.

McGilchrist:1968:EDE

[McG68] C. A. McGilchrist. Efficient difference equation estimators in exponential regression. Annals of Mathematical Statistics, 39(5):1638– 1645, October 1968. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177698145.

McMillan:1949:SML

[McM49] Brockway McMillan. Spread of minima of large samples. Annals of Mathematical Statistics, 20(3):444–447, September 1949. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177729997.

McMillan:1953:BTI

[McM53] Brockway McMillan. The basic theorems of information theory. Annals of Mathematical Statistics, 24(2):196–219, June 1953. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177729028.

McNolty:1962:CID

[McN62] Frank McNolty. A contour-integral derivation of the non-central chisquare distribution. Annals of Mathematical Statistics, 33(2):796– 800, June 1962. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177704598.

McNeil:1970:IFB

[McN70] Donald R. McNeil. Integral functionals of birth and death processes and related limiting distributions. Annals of Mathematical Statistics, 41(2):480–485, April 1970. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177697087.

McPherson:1941:MTP

[McP41] J. C. McPherson. On mechanical tabulation of polynomials. Annals of Mathematical Statistics, 12(3):317–327, September 1941. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177731714.

Morrison:1960:LDR

[MD60] Donald F. Morrison and H. A. David. The life distribution and reliability of a system with spare components. Annals of Mathematical Statistics, 31(4):1084–1094, December 1960. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177705680.

Murty:1968:DAE

[MD68] J. S. Murty and M. N. Das. Design and analysis of experiments with mixtures. Annals of Mathematical Statistics, 39(5):1517–1539, October 1968. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177698134.

Mensing:1971:APC

[MD71] Richard W. Mensing and H. T. David. Absorption probabilities for certain two-dimensional random walks. Annals of Mathematical Statistics, 42(1):304–314, February 1971. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177693513.

Meeden:1970:BTT

[Mee70] Glen Meeden. Best tests for testing hypotheses about a random parameter with unknown distribution. Annals of Mathematical Statistics, 41(2):585–591, April 1970. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177697098.

Meeden:1971:FBT

[Mee71] Glen Meeden. Finding best tests approximately for testing hypotheses about a random parameter. Annals of Mathematical Statistics, 42(4): 1452–1454, August 1971. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177693261.

Meeden:1972:BEM

[Mee72a] Glen Meeden. Bayes estimation of the mixing distribution, the discrete case. Annals of Mathematical Statistics, 43(6):1993–1999, December 1972. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177690872.

Meeden:1972:SAE

[Mee72b] Glen Meeden. Some admissible empirical Bayes procedures. Annals of Mathematical Statistics, 43(1):96–101, February 1972. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177692705.

Mehra:1964:RTP

[Meh64] K. L. Mehra. Rank tests for paired-comparison experiments involving several treatments. Annals of Mathematical Statistics, 35(1):122– 137, March 1964. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177703734.

Mehra:1972:BEJ

[Meh72] K. L. Mehra. On Bahadur efficiency of the joint-ranking procedure. Annals of Mathematical Statistics, 43(4):1155–1163, August 1972. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177692468.

Meilijson:1969:NSM

[Mei69] Isaac Meilijson. A note on sequential multiple decision procedures. Annals of Mathematical Statistics, 40(2):653–657, April 1969. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177697734.

Meilijson:1972:LPM

[Mei72] Isaac Meilijson. Limiting properties of the mean residual lifetime function. Annals of Mathematical Statistics, 43(1):354–357, February 1972. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177692731.

Menger:1952:FYA

[Men52] Karl Menger. The formative years of Abraham Wald and his work in geometry. Annals of Mathematical Statistics, 23(1):14–20, March 1952. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177729481.

Menon:1962:CCD

[Men62] M. V. Menon. A characterization of the Cauchy distribution. Annals of Mathematical Statistics, 33(4):1267–1271, December 1962. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177704358.

Menon:1966:ACP

[Men66] M. V. Menon. Another characteristic property of the Cauchy distribution. Annals of Mathematical Statistics, 37(1):289–294, February 1966. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177699623.

Merrell:1933:CRB

[Mer33] Margaret Merrell. On certain relationships between β_{1} and β_{2} for the point binomial. Annals of Mathematical Statistics, 4(3):216–228, August 1933. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177732780.

Mesner:1965:NPP

[Mes65] Dale M. Mesner. A note on the parameters of PBIB association schemes. Annals of Mathematical Statistics, 36(1):331–336, February 1965. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177700298.

Mesner:1967:NFP

[Mes67] Dale M. Mesner. A new family of partially balanced incomplete block designs with some latin square design properties. Annals of Mathematical Statistics, 38(2):571–581, April 1967. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177698972.

Meshalkin:1968:RSC

[Mes68] L. D. Meshalkin. On the robustness of some characterizations of the normal distribution. Annals of Mathematical Statistics, 39(5):1747– 1750, October 1968. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177698158.

Metivier:1969:EIM

[Met69] Michel Metivier. Existence of an invariant measure and an Ornstein’s ergodic theorem. Annals of Mathematical Statistics, 40(1): 79–96, February 1969. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177697806.

Meyer:1969:NMF

[Mey69] R. M. Meyer. Note on a ‘multivariate’ form of Bonferroni’s inequalities. Annals of Mathematical Statistics, 40(2):692–693, April 1969. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177697743.

Morse:1947:EDD

[MG47] Anthony P. Morse and Frank E. Grubbs. The estimation of dispersion from differences. Annals of Mathematical Statistics, 18(2): 194–214, June 1947. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177730438.

Maxfield:1955:NLH

[MG55] John E. Maxfield and Robert S. Gardner. Note on linear hypotheses with prescribed matrix of normal equations. Annals of Mathematical Statistics, 26(1):149–150, March 1955. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177728609.

Mehta:1969:SPA

[MG69] J. S. Mehta and John Gurland. Some properties and an application of a statistic arising in testing correlation. Annals of Mathematical Statistics, 40(5):1736–1745, October 1969. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177697385.

Murthy:1970:LDS

[MG70] V. K. Murthy and A. V. Gafarian. Limiting distributions of some variations of the chi-square statistic. Annals of Mathematical Statistics, 41(1):188–194, February 1970. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177697199.

McGregor:1962:LDA

[MH62] J. R. McGregor and Y. Y. Hui. Limiting distributions associated with certain stochastic learning models. Annals of Mathematical Statistics, 33(4):1281–1285, December 1962. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177704360.

Moore:1969:UCS

[MH69] D. S. Moore and E. G. Henrichon. Uniform consistency of some estimates of a density function. Annals of Mathematical Statistics, 40(4):1499–1502, August 1969. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177697524. See correction [MH70].

Moore:1970:CNC

[MH70] D. S. Moore and E. G. Henrichon. Correction notes: Correction to “Uniform Consistency of Some Estimates of a Density Function”. Annals of Mathematical Statistics, 41(3):1126–1127, June 1970. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177696997. See

[MH69].

Mickey:1959:SBD

[Mic59] Ray Mickey. Some bounds on the distribution functions of the largest and smallest roots of normal determinantal equations. Annals of Mathematical Statistics, 30(1):242–243, March 1959. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177706380.

Middleton:1966:RRD

[Mid66] David Middleton. Review: Ralph Deutsch, Estimation Theory. Annals of Mathematical Statistics, 37(5):1412–1413, October 1966. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177699289.

Mijares:1961:MES

[Mij61] Tito A. Mijares. The moments of elementary symmetric functions of the roots of a matrix in multivariate analysis. Annals of Mathematical Statistics, 32(4):1152–1160, December 1961. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177704853.

Mijares:1964:ESF

[Mij64] Tito A. Mijares. On elementary symmetric functions of the roots of a multivariate matrix: Distributions. Annals of Mathematical Statistics, 35(3):1186–1198, September 1964. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177703276.

Mijnheer:1972:LIL

[Mij72] J. L. Mijnheer. A law of the iterated logarithm for the asymmetric stable law with characteristic exponent one. Annals of Mathematical Statistics, 43(1):358–362, February 1972. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177692732.

Mikhail:1960:API

[Mik60a] Wadie F. Mikhail. Acknowledgment of priority: An inequality for balanced incomplete block designs. Annals of Mathematical Statistics, 31(4):1213, December 1960. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177705694.

Mikhail:1960:IBI

[Mik60b] Wadie F. Mikhail. An inequality for balanced incomplete block designs. Annals of Mathematical Statistics, 31(2):520–522; addendum, 1213, June 1960. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177705921.

Mikhail:1962:PTE

[Mik62] Wadie F. Mikhail. On a property of a test for the equality of two normal dispersion matrices against one-sided alternatives. Annals of Mathematical Statistics, 33(4):1463–1465, December 1962. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177704380.

Mikulski:1963:EON

[Mik63] Piotr Witold Mikulski. On the efficiency of optimal nonparametric procedures in the two sample case. Annals of Mathematical Statistics, 34(1):22–32, March 1963. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177704240.

Miller:1960:PQ

[Mil60] Rupert G. Miller, Jr. Priority queues. Annals of Mathematical Statistics, 31(1):86–103, March 1960. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177705990.

Miller:1961:CPT

[Mil61a] H. D. Miller. A convexity property in the theory of random variables defined on a finite Markov chain. Annals of Mathematical Statistics, 32(4):1260–1270, December 1961. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177704865.

Miller:1961:GWI

[Mil61b] H. D. Miller. A generalization of Wald’s identity with applications to random walks. Annals of Mathematical Statistics, 32(2): 549–560, June 1961. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177705060.

Miller:1964:RLT

[Mil64a] Rupert G. Miller. Review: Lajos Takacs, Introduction to the Theory of Queues. Annals of Mathematical Statistics, 35(1):450–453, March 1964. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177703777.

Miller:1964:TJ

[Mil64b] Rupert G. Miller, Jr. A trustworthy jackknife. Annals of Mathematical Statistics, 35(4):1594–1605, December 1964. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177700384.

Miller:1967:NSI

[Mil67a] H. D. Miller. A note on sums of independent random variables with infinite first moment. Annals of Mathematical Statistics, 38(3):751– 758, June 1967. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177698868.

Milton:1967:EPR

[Mil67b] Roy C. Milton. On exact probabilities of rank orders for two widely separated normal distributions. Annals of Mathematical Statistics, 38(5):1491–1493, October 1967. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177698703.

Milch:1968:MDL

[Mil68a] Paul R. Milch. A multi-dimensional linear growth birth and death process. Annals of Mathematical Statistics, 39(3):727–754, June 1968. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177698308.

Millar:1968:TSP

[Mil68b] P. Warwick Millar. Transforms of stochastic processes. Annals of Mathematical Statistics, 39(2):372–376, April 1968. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177698398.

Miller:1968:NSD

[Mil68c] K. S. Miller. A note on stochastic difference equations. Annals of Mathematical Statistics, 39(1):270–271, February 1968. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177698531.

Miller:1968:SMD

[Mil68d] K. S. Miller. Some multivariate t-distributions. Annals of Mathematical Statistics, 39(5):1605–1609, October 1968. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177698141.

Miller:1968:JV

[Mil68e] Rupert G. Miller, Jr. Jackknifing variances. Annals of Mathematical Statistics, 39(2):567–582, April 1968. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177698418.

Millar:1969:MII

[Mil69] P. Warwick Millar. Martingales with independent increments. Annals of Mathematical Statistics, 40(3):1033–1041, June 1969. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177697607.

Milliken:1971:NCE

[Mil71a] George A. Milliken. New criteria for estimability for linear models. Annals of Mathematical Statistics, 42(5):1588–1594, October 1971. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177693157.

Milne:1971:SPS

[Mil71b] R. K. Milne. Simple proofs of some theorems on point processes. Annals of Mathematical Statistics, 42(1):368–372, February 1971. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177693523.

Millar:1972:SRA

[Mil72a] P. W. Millar. Some remarks on asymmetric processes. Annals of Mathematical Statistics, 43(2):597–601, April 1972. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177692639.

Miller:1972:ELR

[Mil72b] Douglas R. Miller. Existence of limits in regenerative processes. Annals of Mathematical Statistics, 43(4):1275–1282, August 1972. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177692478.

Miller:1974:JR

[Mil74] Rupert G. Miller. The jackknife — a review. Biometrika, 61(1): 1–15, April 1974. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-

3510 (electronic). URL http://www.jstor.org/stable/2334280.

See [Que49a, Que56, Efr79].

Miner:1931:SEM

[Min31] John Rice Miner. The standard error of a multiple regression equation. Annals of Mathematical Statistics, 2(3):320–323, August 1931. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177732975.

Mineka:1968:MRR

[Min68] J. Mineka. A mixture of recurrent random walks need not be recurrent. Annals of Mathematical Statistics, 39(5):1753–1754, October 1968. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177698160.

Mineka:1972:LLT

[Min72] J. Mineka. Local limit theorems and recurrence conditions for sums of independent integer-valued random variables. Annals of Mathematical Statistics, 43(1):251–259, February 1972. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177692718.

Mitra:1958:LPF

[Mit58] Sujit Kumar Mitra. On the limiting power function of the frequency chi-square test. Annals of Mathematical Statistics, 29(4):1221–1233, December 1958. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177706453.

Madow:1944:TSS

[MM44] William G. Madow and Lillian H. Madow. On the theory of systematic sampling, I. Annals of Mathematical Statistics, 15(1):1–24, March 1944. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177731312.

Magness:1962:CLS

[MM62] T. A. Magness and J. B. McGuire. Comparison of least squares and minimum variance estimates of regression parameters. Annals of Mathematical Statistics, 33(2):462–470, June 1962. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177704573.

Magness:1963:APC

[MM63] T. A. Magness and J. B. McGuire. Acknowledgment of priority: “Comparison of Least Squares and Minimum Variance Estimates of Regression Parameters”. Annals of Mathematical Statistics, 34(1): 352, March 1963. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177704281.

Margolin:1969:ECL

[MM69] Barry H. Margolin and Frederick Mosteller. The expected coverage to the left of the i-th order statistic for arbitrary distributions. Annals of Mathematical Statistics, 40(2):644–647, April 1969. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177697732.

Marshall:1960:MCI

[MO60a] Albert W. Marshall and Ingram Olkin. Multivariate Chebyshev inequalities. Annals of Mathematical Statistics, 31(4):1001–1014, December 1960. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177705673.

Marshall:1960:OSI

[MO60b] Albert W. Marshall and Ingram Olkin. A one-sided inequality of the Chebyshev type. Annals of Mathematical Statistics, 31(2):488– 491, June 1960. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177705913.

Memon:1970:CSC

[MO70] Ahmed Zogo Memon and Masashi Okamoto. The classification statistic W^{∗} in covariate discriminant analysis. Annals of Mathematical Statistics, 41(5):1491–1499, October 1970. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177696794.

Mohanty:1971:SPS

[Moh71] S. G. Mohanty. A short proof of Steck’s result on two-sample Smirnov statistics. Annals of Mathematical Statistics, 42(1):413–414, February 1971. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177693537.

Molina:1931:BTE

[Mol31] Edward C. Molina. Bayes’ Theorem: An expository presentation. Annals of Mathematical Statistics, 2(1):23–37, February 1931. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177733040.

Molina:1946:SFC

[Mol46] Edward C. Molina. Some fundamental curves for the solution of sampling problems. Annals of Mathematical Statistics, 17(3):325– 335, September 1946. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177730945.

Montroll:1947:TMC

[Mon47] Elliott W. Montroll. On the theory of Markoff chains. Annals of Mathematical Statistics, 18(1):18–36, March 1947. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177730489.

Monroe:1972:ERC

[Mon72a] Itrel Monroe. On embedding right continuous martingales in Brownian motion. Annals of Mathematical Statistics, 43(4):1293–1311, August 1972. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177692480.

Monroe:1972:VPS

[Mon72b] Itrel Monroe. On the γ-variation of processes with stationary independent increments. Annals of Mathematical Statistics, 43(4):1213– 1220, August 1972. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177692473.

Mood:1939:NTM

[Moo39] A. M. Mood. Note on the L_{1} test for many samples. Annals of Mathematical Statistics, 10(2):187–190, June 1939. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177732215.

Mood:1940:DTR

[Moo40] A. M. Mood. The distribution theory of runs. Annals of Mathematical Statistics, 11(4):367–392, December 1940. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177731825.

Mood:1941:JDM

[Moo41] A. M. Mood. On the joint distribution of the medians in samples from a multivariate population. Annals of Mathematical Statistics, 12(3):268–278, September 1941. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177731709.

Mood:1943:DSI

[Moo43] Alexander M. Mood. On the dependence of sampling inspection plans upon population distributions. Annals of Mathematical Statistics, 14 (4):415–425, December 1943. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177731361.

Mood:1946:HWP

[Moo46] Alexander M. Mood. On Hotelling’s weighing problem. Annals of Mathematical Statistics, 17(4):432–446, December 1946. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177730883.

Mood:1948:RTL

[Moo48] A. M. Mood. Review: Truman Lee Kelley, Fundamentals of Statistics. Annals of Mathematical Statistics, 19(2):282–284, June 1948. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177730257.

Mood:1949:TIC

[Moo49] A. M. Mood. Tests of independence in contingency tables as unconditional tests. Annals of Mathematical Statistics, 20(1):114–116, March 1949. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177730098.

Mood:1951:DCR

[Moo51a] A. M. Mood. On the distribution of the characteristic roots of normal second-moment matrices. Annals of Mathematical Statistics, 22(2): 266–273, June 1951. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177729646.

Mood:1951:RSM

[Moo51b] A. M. Mood. Report of the Santa Monica meeting of the Institute. Annals of Mathematical Statistics, 22(3):492–494, September 1951. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177729605.

Mood:1954:AEC

[Moo54] A. M. Mood. On the asymptotic efficiency of certain nonparametric two-sample tests. Annals of Mathematical Statistics, 25(3):514–522, September 1954. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177728719.

Moon:1963:LGC

[Moo63a] J. W. Moon. On the line-graph of the complete bigraph. Annals of Mathematical Statistics, 34(2):664–667, June 1963. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177704179.

Moore:1963:DRS

[Moo63b] Calvin C. Moore. The degree of randomness in a stationary time series. Annals of Mathematical Statistics, 34(4):1253–1258, December 1963. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177703860.

Mood:1964:RRS

[Moo64] Alex M. Mood. Review: I. Richard Savage, Bibliography of Nonparametric Statistics. Annals of Mathematical Statistics, 35(2):915– 916, June 1964. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177703603.

Moon:1968:TCP

[Moo68a] J. W. Moon. A tree counting problem. Annals of Mathematical Statistics, 39(1):242–245, February 1968. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177698524.

Moore:1968:EPA

[Moo68b] D. S. Moore. An elementary proof of asymptotic normality of linear functions of order statistics. Annals of Mathematical Statistics, 39(1): 263–265, February 1968. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177698529.

Moore:1969:ANE

[Moo69] D. S. Moore. Asymptotically nearly efficient estimators of multivariate location parameters. Annals of Mathematical Statistics, 40 (5):1809–1823, October 1969. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177697394.

Moore:1971:CSS

[Moo71] D. S. Moore. A chi-square statistic with random cell boundaries. Annals of Mathematical Statistics, 42(1):147–156, February 1971. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177693502.

Moritz:1932:NTD

[Mor32] Robert E. Moritz. A new theory of depreciation of physical assets. Annals of Mathematical Statistics, 3(2):108–125, May 1932. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177732909.

Moriguti:1951:EPE

[Mor51] Sigeiti Moriguti. Extremal properties of extreme value distributions. Annals of Mathematical Statistics, 22(4):523–536, December 1951. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177729542.

Moriguti:1952:LBP

[Mor52] Sigeiti Moriguti. A lower bound for a probability moment of any absolutely continuous distribution with finite variance. Annals of Mathematical Statistics, 23(2):286–289, June 1952. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177729447.

Moriguti:1953:MSI

[Mor53] Sigeiti Moriguti. A modification of Schwarz’s inequality with applications to distributions. Annals of Mathematical Statistics, 24(1): 107–113, March 1953. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177729088.

Morrison:1962:DSS

[Mor62] Donald F. Morrison. On the distribution of sums of squares and cross products of normal variates in the presence of intra-class correlation. Annals of Mathematical Statistics, 33(4):1461–1463, December 1962. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177704379.

Moran:1966:RUP

[Mor66] P. A. P. Moran. Review: N. U. Prabhu, Queues and Inventories. Annals of Mathematical Statistics, 37(2):529–530, April 1966. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177699540.

Morimoto:1972:SSP

[Mor72] Haruki Morimoto. Statistical structure of the problem of sampling from finite populations. Annals of Mathematical Statistics, 43(2): 490–497, April 1972. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177692629.

Mosteller:1941:NAR

[Mos41] Frederick Mosteller. Note on an application of runs to quality control charts. Annals of Mathematical Statistics, 12(2):228–232, June 1941.

CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177731753.

Mosteller:1946:SUI

[Mos46] Frederick Mosteller. On some useful “inefficient” statistics. Annals of Mathematical Statistics, 17(4):377–408, December 1946. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177730881.

Mosteller:1948:SST

[Mos48a] Frederick Mosteller. A k-sample slippage test for an extreme population. Annals of Mathematical Statistics, 19(1):58–65, March 1948. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177730290.

Mosteller:1948:RGW

[Mos48b] Frederick Mosteller. Review: George W. Snedecor, Statistical Methods. Annals of Mathematical Statistics, 19(1):124–126, March 1948. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177730302.

Moshman:1951:ROR

[Mos51] Jack Moshman. Report of the Oak Ridge meeting of the Institute. Annals of Mathematical Statistics, 22(2):322–325, June 1951. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177729660.

Moshman:1952:TSM

[Mos52] Jack Moshman. Testing a straggler mean in a two-way classification using the range. Annals of Mathematical Statistics, 23(1):126– 132, March 1952. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177729490.

Moshman:1958:MSS

[Mos58] Jack Moshman. A method for selecting the size of the initial sample in Stein’s two sample procedure. Annals of Mathematical Statistics, 29 (4):1271–1275, December 1958. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177706460.

Moses:1962:UWT

[Mos62] Lincoln E. Moses. Use of Wilcoxon test theory in estimating the distribution of a ratio by Monte Carlo methods. Annals of Mathematical Statistics, 33(3):1194–1197, September 1962. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177704483.

Moses:1963:RTD

[Mos63] Lincoln E. Moses. Rank tests of dispersion. Annals of Mathematical Statistics, 34(3):973–983, September 1963. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177704020.

Moses:1965:RJW

[Mos65] Lincoln E. Moses. Review: John E. Walsh, Handbook of Nonparametric Statistics. Annals of Mathematical Statistics, 36(2):719–722, April 1965. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177700186.

Mote:1958:MPB

[Mot58] V. L. Mote. On a minimax property of a balanced incomplete block design. Annals of Mathematical Statistics, 29(3):910–914, September 1958. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177706550.

Mouzon:1930:EFD

[Mou30] Edwin D. Mouzon. Equimodal frequency distributions. Annals of Mathematical Statistics, 1(2):137–158, May 1930. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177733128.

Marshall:1965:MLE

[MP65] Albert W. Marshall and Frank Proschan. Maximum likelihood estimation for distributions with monotone failure rate. Annals of Mathematical Statistics, 36(1):69–77, February 1965. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177700271.

Mehra:1967:MSA

[MP67] K. L. Mehra and M. L. Puri. Multi-sample analogues of some one-sample tests. Annals of Mathematical Statistics, 38(2):523–549, April 1967. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177698969.

McEliece:1971:HSD

[MP71] Robert J. McEliece and Edward C. Posner. Hide and seek, data storage, and entropy. Annals of Mathematical Statistics, 42(5):1706– 1716, October 1971. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177693169.

McMillan:1957:MSS

[MR57] B. McMillan and J. Riordan. A moving single server problem. Annals of Mathematical Statistics, 28(2):471–478, June 1957. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177706975.

Mitra:1964:RSB

[MR64] S. K. Mitra and C. R. Rao. Review: A. E. Sarhan, B. G. Greenberg, Contributions to Order Statistics. Annals of Mathematical Statistics, 35(3):1383–1385, September 1964. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177703302.

MacQueen:1965:SAM

[MR65] J. MacQueen and R. M. Redheffer. Some applications of monotone operators in Markov processes. Annals of Mathematical Statistics, 36 (5):1421–1425, October 1965. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177699900.

Morton:1966:TKR

[MR66] Richard Morton and M. Raghavachari. On a theorem of Karlin regarding admissibility of linear estimates in exponential populations. Annals of Mathematical Statistics, 37(6):1809–1813, December 1966. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177699172.

Mudholkar:1967:SSM

[MR67] Govind S. Mudholkar and Poduri S. R. S. Rao. Some sharp multivariate Tchebycheff inequalities. Annals of Mathematical Statistics, 38 (2):393–400, April 1967. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177698954.

Mallows:1969:ICT

[MR69a] C. L. Mallows and Donald Richter. Inequalities of Chebyshev type involving conditional expectations. Annals of Mathematical Statistics, 40(6):1922–1932, December 1969. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177697276.

Mitra:1969:COV

[MR69b] Sujit Kumar Mitra and C. Radhakrishna Rao. Conditions for optimality and validity of simple least squares theory. Annals of Mathematical Statistics, 40(5):1617–1624, October 1969. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177697377.

Mathai:1971:EDW

[MR71] A. M. Mathai and P. N. Rathie. The exact distribution of Wilks’ criterion. Annals of Mathematical Statistics, 42(3):1010–1019, June 1971. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177693329.

Masterson:1963:QT

[MS63] Gregory E. Masterson and Seymour Sherman. On queues in tandem. Annals of Mathematical Statistics, 34(1):300–307, March 1963. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177704266.

Mott-Smith:1964:CNC

[MS64a] J. C. Mott-Smith. Correction note: Correction to “Two Estimates of the Binomial Distribution”. Annals of Mathematical Statistics, 35(4): 1827, December 1964. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177700409. See [MS64b].

Mott-Smith:1964:TEB

[MS64b] J. C. Mott-Smith. Two estimates of the binomial distribution. Annals of Mathematical Statistics, 35(2):809–816, June 1964. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177703580. See correc-

tion [MS64a].

Morse:1966:SI

[MS66] Norman Morse and Richard Sacksteder. Statistical isomorphism. Annals of Mathematical Statistics, 37(1):203–214, February 1966. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177699610.

Mehra:1967:AEC

[MS67] K. L. Mehra and J. Sarangi. Asymptotic efficiency of certain rank tests for comparative experiments. Annals of Mathematical Statistics, 38(1):90–107, February 1967. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177699060.

Maag:1968:TST

[MS68a] Urs R. Maag and M. A. Stephens. The V_{NM} two-sample test. Annals of Mathematical Statistics, 39(3):923–935, June 1968. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177698324.

Myhre:1968:CLR

[MS68b] Janet M. Myhre and Sam C. Saunders. On confidence limits for the reliability of systems. Annals of Mathematical Statistics, 39(5):1463– 1472, October 1968. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL ahttp://projecteuclid.org/euclid.
aoms/1177698125.

Mathai:1969:DPS

[MS69a] A. M. Mathai and R. K. Saxena. Distribution of a product and the structural set up of densities. Annals of Mathematical Statistics, 40 (4):1439–1448, August 1969. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177697515.

Mehra:1969:CCD

[MS69b] K. L. Mehra and P. K. Sen. On a class of conditionally distributionfree tests for interactions in factorial experiments. Annals of Mathematical Statistics, 40(2):658–664, April 1969. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177697735.

McCabe:1970:S

[MS70a] B. J. McCabe and L. A. Shepp. On the supremum of S_{n}/n. Annals of Mathematical Statistics, 41(6):2166–2168, December 1970. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177696723.

Mineka:1970:LLT

[MS70b] J. Mineka and S. Silverman. A local limit theorem and recurrence conditions for sums of independent non-lattice random variables. Annals of Mathematical Statistics, 41(2):592–600, April 1970. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177697099.

Myhre:1971:ACL

[MS71] J. M. Myhre and Sam C. Saunders. Approximate confidence limits for complex systems with exponential component lives. Annals of Mathematical Statistics, 42(1):342–348, February 1971. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177693517.

Martin:1972:MQM

[MS72a] R. D. Martin and S. C. Schwartz. On mixture, quasi-mixture and nearly normal random processes. Annals of Mathematical Statistics, 43(3):948–967, June 1972. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177692558.

Miller:1972:WCS

[MS72b] R. G. Miller, Jr. and Pranab Kumar Sen. Weak convergence of Ustatistics and von Mises’ differentiable statistical functions. Annals of Mathematical Statistics, 43(1):31–41, February 1972. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177692698.

Mosteller:1950:SLS

[MT50] Frederick Mosteller and John W. Tukey. Significance levels for a ksample slippage test. Annals of Mathematical Statistics, 21(1):120– 123, March 1950. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177729892.

Matthes:1965:OIR

[MT65] T. K. Matthes and D. R. Truax. Optimal invariant rank tests for the k-sample problem. Annals of Mathematical Statistics, 36(4):1207– 1222, August 1965. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177699992.

Matthes:1967:CCT

[MT67a] T. K. Matthes and D. R. Truax. Correction: Correction to tests of composite hypotheses for the multivariate exponential family. Annals of Mathematical Statistics, 38(6):1928, December 1967. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177698631. See

[MT67b].

Matthes:1967:TCH

[MT67b] T. K. Matthes and D. R. Truax. Tests of composite hypotheses for the multivariate exponential family. Annals of Mathematical Statistics, 38 (3):681–697, June 1967. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.

aoms/1177698862. See correction [MT67a].

Mudholkar:1965:CTM

[Mud65] Govind S. Mudholkar. A class of tests with monotone power functions for two problems in multivariate statistical analysis. Annals of Mathematical Statistics, 36(6):1794–1801, December 1965. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177699808.

Mudholkar:1966:CBA

[Mud66] Govind S. Mudholkar. On confidence bounds associated with multivariate analysis of variance and non-independence between two sets of variates. Annals of Mathematical Statistics, 37(6):1736–1746, December 1966. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177699162.

Muirhead:1970:ADS

[Mui70a] R. J. Muirhead. Asymptotic distributions of some multivariate tests. Annals of Mathematical Statistics, 41(3):1002–1010, June 1970. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177696976.

Muirhead:1970:SPD

[Mui70b] R. J. Muirhead. Systems of partial differential equations for hypergeometric functions of matrix argument. Annals of Mathematical Statistics, 41(3):991–1001, June 1970. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177696975.

Muirhead:1972:AND

[Mui72a] R. J. Muirhead. The asymptotic noncentral distribution of Hotelling’s generalized T_{0}^{2}. Annals of Mathematical Statistics, 43(5):1671–1677, October 1972. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177692400.

Muirhead:1972:TIB

[Mui72b] R. J. Muirhead. On the test of independence between two sets of variates. Annals of Mathematical Statistics, 43(5):1491–1497, October 1972. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177692381.

Muller:1956:SCM

[Mul56] Mervin E. Muller. Some continuous Monte Carlo methods for the Dirichlet problem. Annals of Mathematical Statistics, 27(3):569–589, September 1956. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177728169.

Muller:1966:RJM

[Mul66] Mervin E. Muller. Review: J. M. Hammersley, D. C. Handscomb, Monte Carlo Methods; Yu. A. Shreider, Methods of Statistical Testing/Monte Carlo Method. Annals of Mathematical Statistics, 37(2): 532–538, April 1966. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177699542.

Murphy:1948:NPT

[Mur48] R. B. Murphy. Non-parametric tolerance limits. Annals of Mathematical Statistics, 19(4):581–589, December 1948. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177730154.

Murty:1956:NBB

[Mur56] V. N. Murty. A note on Bhattacharyya bounds for the negative binomial distribution. Annals of Mathematical Statistics, 27(4):1182– 1183, December 1956. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177728089.

Murthy:1961:ES

[Mur61a] V. K. Murthy. Estimation of the spectrum. Annals of Mathematical Statistics, 32(3):730–738, September 1961. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177704968.

Murty:1961:IBI

[Mur61b] V. N. Murty. An inequality for balanced incomplete block designs. Annals of Mathematical Statistics, 32(3):908–909, September 1961. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177704988.

Murthy:1963:ECS

[Mur63] V. K. Murthy. Estimation of the cross-spectrum. Annals of Mathematical Statistics, 34(3):1012–1021, September 1963. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177704024.

Murthy:1965:EJR

[Mur65a] V. K. Murthy. Estimation of jumps, reliability and hazard rate. Annals of Mathematical Statistics, 36(3):1032–1040, June 1965. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177700075.

Murthy:1965:EPD

[Mur65b] V. K. Murthy. Estimation of probability density. Annals of Mathematical Statistics, 36(3):1027–1031, June 1965. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177700074.

Murty:1971:ODP

[Mur71a] V. N. Murty. Optimal designs with a polynomial spline regression with a single multiple knot at the center. Annals of Mathematical Statistics, 42(3):952–960, June 1971. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177693324.

Murty:1971:ODT

[Mur71b] V. N. Murty. Optimal designs with a Tchebycheffian spline regression function. Annals of Mathematical Statistics, 42(2):643–649, April 1971. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177693414.

Mureika:1972:MED

[Mur72] Roman A. Mureika. The maximization of entropy of discrete denumerably-valued random variables with known mean. Annals of Mathematical Statistics, 43(2):541–552, April 1972. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177692635.

Mustafi:1967:ICM

[Mus67] Chandan K. Mustafi. The inverse of a certain matrix, with an application. Annals of Mathematical Statistics, 38(4):1289–1292, August 1967. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177698802.

Mustafi:1968:IPA

[Mus68a] Chandan K. Mustafi. Inference problems about parameters which are subjected to changes over time. Annals of Mathematical Statistics, 39 (3):840–854, June 1968. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177698315.

Mustafi:1968:POA

[Mus68b] Chandan K. Mustafi. On the proportion of observations above sample means in a bivariate normal distribution. Annals of Mathematical Statistics, 39(4):1350–1353, August 1968. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177698262.

Mustafi:1971:CBD

[Mus71] Chandan K. Mustafi. On a class of bivariate distributions following a certain stochastic structure. Annals of Mathematical Statistics, 42 (4):1359–1371, August 1971. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177693248.

Molenaar:1967:HSF

[Mv67] W. Molenaar and E. A. van der Velde. How to survive a fixed number of fair bets. Annals of Mathematical Statistics, 38(4):1278–1280, August 1967. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177698798.

Miller:1969:DDP

[MV69] Bruce L. Miller and Arthur F. Veinott, Jr. Discrete dynamic programming with a small interest rate. Annals of Mathematical Statistics, 40 (2):366–370, April 1969. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177697700.

Molenaar:1966:MD

[MvZ66] W. Molenaar and W. R. van Zwet. On mixtures of distributions. Annals of Mathematical Statistics, 37(1):281–283, February 1966. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177699620.

Mann:1942:CNC

[MW42] H. B. Mann and A. Wald. On the choice of the number of class intervals in the application of the chi square test. Annals of Mathematical Statistics, 13(3):306–317, September 1942. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http:/

/projecteuclid.org/euclid.aoms/1177731569;   http://www.
jstor.org/stable/2235942.

Mann:1943:SLO

[MW43] H. B. Mann and A. Wald. On stochastic limit and order relationships. Annals of Mathematical Statistics, 14(3):217–226, September 1943. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177731415.

Mann:1947:TWO

[MW47] H. B. Mann and D. R. Whitney. On a test of whether one of two random variables is stochastically larger than the other. Annals of Mathematical Statistics, 18(1):50–60, March 1947. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177730491.

Mejzler:1969:SRV

[MW69] D. Mejzler and I. Weissman. On some results of N. V. Smirnov concerning limit distributions for variational series. Annals of Mathematical Statistics, 40(2):480–491, April 1969. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177697714.

Myers:1934:NKM

[Mye34] R. J. Myers. Note on Koshal’s method of improving the parameters of curves by the use of the method of maximum likelihood. Annals of Mathematical Statistics, 5(4):320–323, December 1934. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177732657.

McGregor:1965:LDR

[MZ65] J. R. McGregor and J. V. Zidek. Limiting distributions of response probabilities. Annals of Mathematical Statistics, 36(2):706– 707, April 1965. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177700180.

Martin:1966:CIU

[MZ66] Frank B. Martin and George Zyskind. On combinability of information from uncorrelated linear models by simple weighting. Annals of Mathematical Statistics, 37(5):1338–1347, October 1966. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177699278.

Nadler:1960:IBS

[Nad60] Jack Nadler. Inverse binomial sampling plans when an exponential distribution is sampled with censoring. Annals of Mathematical Statistics, 31(4):1201–1204, December 1960. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177705691.

Nadas:1969:ETC

[Nád69] Arthur Nádas. An extension of a theorem of Chow and Robbins on sequential confidence intervals for the mean. Annals of Mathematical Statistics, 40(2):667–671, April 1969. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177697737.

Nagabhushanam:1951:LTP

[Nag51] K. Nagabhushanam. Linear transformations and the product-moment matrix. Annals of Mathematical Statistics, 22(2):302–304, June 1951. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177729652.

Nair:1966:PLB

[Nai66] C. Ramankutty Nair. On partially linked block designs. Annals of Mathematical Statistics, 37(5):1401–1406, October 1966. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177699285.

Nair:1969:PBI

[Nai69] C. Ramankutty Nair. Partially balanced incomplete block designs with two-way classification of treatments. Annals of Mathematical Statistics, 40(1):175–187, February 1969. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177697814.

Nanda:1948:DRD

[Nan48a] D. N. Nanda. Distribution of a root of a determinantal equation. Annals of Mathematical Statistics, 19(1):47–57, March 1948. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177730289.

Nanda:1948:LDR

[Nan48b] D. N. Nanda. Limiting distribution of a root of a determinantal equation. Annals of Mathematical Statistics, 19(3):340–350, September 1948. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177730198.

Nanda:1950:DSR

[Nan50] D. N. Nanda. Distribution of the sum of roots of a determinantal equation under a certain condition. Annals of Mathematical Statistics, 21(3):432–439, September 1950. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177729802.

Narain:1950:CUC

[Nar50] R. D. Narain. On the completely unbiassed character of tests of independence in multivariate normal systems. Annals of Mathematical Statistics, 21(2):293–298, June 1950. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177729848.

Nash:1954:EBC

[Nas54] Stanley W. Nash. An extension of the Borel–Cantelli lemma. Annals of Mathematical Statistics, 25(1):165–167, March 1954. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177728858.

Neyman:1965:DHP

[NCC65] Jerzy Neyman, Herman Chernoff, and D. G. Chapman. Discussion of Hoeffding’s paper. Annals of Mathematical Statistics, 36(2):401– 408, April 1965. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/

1177700151. See [Hoe65a].

Neave:1970:IFA

[Nea70] Henry R. Neave. An improved formula for the asymptotic variance of spectrum estimates. Annals of Mathematical Statistics, 41(1):70–77, February 1970. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177697189.

Neave:1971:EES

[Nea71] Henry R. Neave. The exact error in spectrum estimates. Annals of Mathematical Statistics, 42(3):961–975, June 1971. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177693325.

Nelson:1966:MSS

[Nel66] Wayne Nelson. Minimax solution of statistical decision problems by iteration. Annals of Mathematical Statistics, 37(6):1643–1657, December 1966. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177699153.

Nelson:1970:COS

[Nel70] Paul I. Nelson. A class of orthogonal series related to martingales. Annals of Mathematical Statistics, 41(5):1684–1694, October 1970. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177696812.

Nemeth:1971:NFB

[Nem71] E. Nemeth. A new family of BIBD’s. Annals of Mathematical Statistics, 42(3):1118–1120, June 1971. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177693343.

Neuts:1964:GFM

[Neu64] Marcel F. Neuts. Generating functions for Markov renewal processes. Annals of Mathematical Statistics, 35(1):431–434, March 1964. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177703769.

Neuts:1967:GCB

[Neu67] Marcel F. Neuts. A general class of bulk queues with Poisson input. Annals of Mathematical Statistics, 38(3):759–770, June 1967. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177698869.

Neuhaus:1971:WCS

[Neu71] Georg Neuhaus. On weak convergence of stochastic processes with multidimensional time parameter. Annals of Mathematical Statistics, 42(4):1285–1295, August 1971. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177693241.

Neveu:1972:NTM

[Nev72] J. Neveu. Note on the tightness of the metric on the set of complete sub σ-algebras of a probability space. Annals of Mathematical Statistics, 43(4):1369–1371, August 1972. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177692492.

Newell:1960:QFC

[New60] G. F. Newell. Queues for a fixed-cycle traffic light. Annals of Mathematical Statistics, 31(3):589–597, September 1960. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177705787.

Newell:1964:AED

[New64] G. F. Newell. Asymptotic extremes for m-dependent random variables. Annals of Mathematical Statistics, 35(3):1322–1325, September 1964. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177703288.

Newman:1968:PDF

[New68] David S. Newman. On the probability distribution of a filtered random telegraph signal. Annals of Mathematical Statistics, 39(3):890– 896, June 1968. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177698321.

Neyman:1935:PCI

[Ney35] J. Neyman. On the problem of confidence intervals. Annals of Mathematical Statistics, 6(3):111–116, September 1935. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177732585.

Neyman:1938:TSH

[Ney38] J. Neyman. Tests of statistical hypotheses which are unbiased in the limit. Annals of Mathematical Statistics, 9(2):69–86, June 1938. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177732329.

Neyman:1939:NCC

[Ney39] J. Neyman. On a new class of “contagious” distributions, applicable in entomology and bacteriology. Annals of Mathematical Statistics, 10 (1):35–57, March 1939. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177732245.

Neyman:1941:SPA

[Ney41] J. Neyman. On a statistical problem arising in routine analyses and in sampling inspections of mass production. Annals of Mathematical Statistics, 12(1):46–76, March 1941. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177731785.

Neyman:1950:RPI

[Ney50] J. Neyman. Report of the President of the Institute for 1949. Annals of Mathematical Statistics, 21(1):155–160, March 1950. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177729902.

Neyman:1951:ECE

[Ney51] Jerzy Neyman. Existence of consistent estimates of the directional parameter in a linear structural relation between two variables. Annals of Mathematical Statistics, 22(4):497–512, December 1951. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177729540.

Ney:1962:RIF

[Ney62] P. E. Ney. A random interval filling problem. Annals of Mathematical Statistics, 33(2):702–718, June 1962. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177704592.

Ney:1963:LRC

[Ney63] P. E. Ney. The limit of a ratio of convolutions. Annals of Mathematical Statistics, 34(2):457–461, June 1963. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177704157.

Ney:1966:RWS

[Ney66] Peter Ney. Review: W. Smith, W. Wilkinson, Proceedings of Symposium on Congestion Theory. Annals of Mathematical Statistics, 37 (5):1410–1411, October 1966. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177699288.

Nicholson:1954:CFP

[Nic54] W. L. Nicholson. A computing formula for the power of the analysis of variance test. Annals of Mathematical Statistics, 25(3):607–610, September 1954. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177728731.

Nicholson:1956:NAH

[Nic56] W. L. Nicholson. On the normal approximation to the hypergeometric distribution. Annals of Mathematical Statistics, 27(2):471– 483, June 1956. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177728270.

Nicholson:1958:DRN

[Nic58] W. L. Nicholson. On the distribution of 2×2 random normal determinants. Annals of Mathematical Statistics, 29(2):575–580, June 1958. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177706634.

Nigam:1970:BDM

[Nig70] A. K. Nigam. Block designs for mixture experiments. Annals of Mathematical Statistics, 41(6):1861–1869, December 1970. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177696687. See correc-

tion [Nig76].

Nigam:1976:NCB

[Nig76] A. K. Nigam. Notes: Correction to “Block Designs for Mixture Experiments”. Annals of Statistics, 4(6):1294–1295, November 1976. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176343663. See [Nig70].

Niiniluoto:1972:NFT

[Nii72] Ilkka Niiniluoto. A note on fine and tight qualitative probabilities. Annals of Mathematical Statistics, 43(5):1581–1591, October 1972. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177692390.

Noack:1950:CRV

[Noa50] Albert Noack. A class of random variables with discrete distributions. Annals of Mathematical Statistics, 21(1):127–132, March 1950. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177729894.

Noether:1948:CLQ

[Noe48] Gottfried E. Noether. On confidence limits for quantiles. Annals of Mathematical Statistics, 19(3):416–419, September 1948. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177730209.

Noether:1949:TWW

[Noe49] Gottfried E. Noether. On a theorem by Wald and Wolfowitz. Annals of Mathematical Statistics, 20(3):455–458, September 1949. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177730000.

Noether:1950:APW

[Noe50] Gottfried Emanuel Noether. Asymptotic properties of the Wald– Wolfowitz test of randomness. Annals of Mathematical Statistics, 21 (2):231–246, June 1950. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177729841.

Noether:1951:CBC

[Noe51] Gottfried E. Noether. On a connection between confidence and tolerance intervals. Annals of Mathematical Statistics, 22(4):603–604, December 1951. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177729553.

Noether:1955:TP

[Noe55] Gottfried E. Noether. On a theorem of Pitman. Annals of Mathematical Statistics, 26(1):64–68, March 1955. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177728593.

Noether:1970:CLT

[Noe70] Gottfried E. Noether. A Central Limit Theorem with nonparametric applications. Annals of Mathematical Statistics, 41(5):1753–1755, October 1970. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177696820.

Noe:1972:CDT

[Noé72] Marc Noé. The calculation of distributions of two-sided Kolmogorov– Smirnov type statistics. Annals of Mathematical Statistics, 43(1): 58–64, February 1972. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177692700.

Norris:1935:IAA

[Nor35] Nilan Norris. Inequalities among averages. Annals of Mathematical Statistics, 6(1):27–29, March 1935. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177732639.

Norris:1937:CPG

[Nor37] Nilan Norris. Convexity properties of generalized mean value functions. Annals of Mathematical Statistics, 8(2):118–120, June 1937. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177732433.

Norris:1938:SEM

[Nor38] Nilan Norris. Some efficient measures of relative dispersion. Annals of Mathematical Statistics, 9(3):214–220, September 1938. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177732311.

Norris:1940:SEG

[Nor40] Nilan Norris. The standard errors of the geometric and harmonic means and their application to index numbers. Annals of Mathematical Statistics, 11(4):445–448, December 1940. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177731830.

Norman:1966:LDS

[Nor66] M. Frank Norman. Limiting distributions for some random walks arising in learning models. Annals of Mathematical Statistics, 37(2): 393–405, April 1966. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177699521.

Norvig:1967:CSP

[Nor67] Torsten Norvig. Consensus of subjective probabilities: A convergence theorem. Annals of Mathematical Statistics, 38(1):221–225, February 1967. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177699074.

Norman:1970:RMI

[Nor70] M. Frank Norman. Review: M. Iosifescu, R. Theodorescu, Random Processes and Learning. Annals of Mathematical Statistics, 41(4): 1381–1383, August 1970. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177696922.

Newton:1966:FAN

[NP66] D. Newton and W. Parry. On a factor automorphism of a normal dynamical system. Annals of Mathematical Statistics, 37(6):1528– 1533, December 1966. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177699144.

Newell:1962:ZCP

[NR62] G. F. Newell and M. Rosenblatt. Zero crossing probabilities for Gaussian stationary processes. Annals of Mathematical Statistics, 33(4): 1306–1313, December 1962. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177704363.

Na:1967:FPO

[NR67] Hwa Sung Na and Anatol Rapoport. A formula for the probability of obtaining a tree from a graph constructed randomly except for an “Exogamous Bias”. Annals of Mathematical Statistics, 38(1):226– 241, February 1967. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177699075.

Nadler:1971:SCP

[NR71] Jack Nadler and Naomi B. Robbins. Some characteristics of Page’s two-sided procedure for detecting a change in a location parameter. Annals of Mathematical Statistics, 42(2):538–551, April 1971. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177693404.

Neyman:1951:CME

[NS51] J. Neyman and Elizabeth L. Scott. On certain methods of estimating the linear structural relation. Annals of Mathematical Statistics, 22 (3):352–361, September 1951. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177729581. See correction [NS52].

Neyman:1952:CCM

[NS52] J. Neyman and Elizabeth L. Scott. Correction to “On Certain Methods of Estimating the Linear Structural Relation”. Annals of Mathematical Statistics, 23(1):135, March 1952. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177729492. See [NS51].

Neyman:1960:CBI

[NS60] Jerzy Neyman and Elizabeth L. Scott. Correction for bias introduced by a transformation of variables. Annals of Mathematical Statistics, 31(3):643–655, September 1960. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177705791.

Nagambal:1970:PTP

[NT70] P. N. Nagambal and D. S. Tracy. Products of two polykays when one has weight 5. Annals of Mathematical Statistics, 41(3):1114– 1121, June 1970. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177696994.

Nuesch:1966:PTL

[Nüe66] Peter E. Nüesch. On the problem of testing location in multivariate populations for restricted alternatives. Annals of Mathematical Statistics, 37(1):113–119, February 1966. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177699602.

Nelson:1960:EFS

[NV60] Edward Nelson and Dale Varberg. Expectations of functionals on a stochastic process. Annals of Mathematical Statistics, 31(3):574– 578, September 1960. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177705785.

Noe:1968:CDK

[NV68] Marc Noé and Georges Vandewiele. The calculation of distributions of Kolmogorov–Smirnov type statistics including a table of significance points for a particular case. Annals of Mathematical Statistics, 39(1): 233–241, February 1968. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177698523.

Oaten:1972:ABR

[Oat72] Allan Oaten. Approximation to Bayes risk in compound decision problems. Annals of Mathematical Statistics, 43(4):1164–1184, August 1972. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177692469.

OBrien:1972:NCM

[O’B72] George O’Brien. A note on comparisons of Markov processes. Annals of Mathematical Statistics, 43(1):365–368, February 1972. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177692734.

Oberg:1947:AFR

[Obe47] E. N. Oberg. Approximate formulas for the radii of circles which include a specified fraction of a normal bivariate distribution. Annals of Mathematical Statistics, 18(3):442–447, September 1947. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177730392.

Obenchain:1971:MPI

[Obe71] R. L. Obenchain. Multivariate procedures invariant under linear transformations. Annals of Mathematical Statistics, 42(5):1569–1578; correction, ibid. 43 (1972), 1742–1743, October 1971. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http:

//projecteuclid.org/euclid.aoms/1177693155. See correction

[Obe72a].

Obenchain:1972:CNC

[Obe72a] R. L. Obenchain. Correction notes: Correction to “Multivariate Procedures Invariant Under Linear Transformations”. Annals of Mathematical Statistics, 43(5):1742–1743, October 1972. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177692414. See [Obe71].

Obenchain:1972:ROP

[Obe72b] R. L. Obenchain. Regression optimality of principal components. Annals of Mathematical Statistics, 43(4):1317–1319, August 1972. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177692482.

Ogawara:1951:NTS

[Oga51] Masami Ogawara. A note on the test of serial correlation coefficients. Annals of Mathematical Statistics, 22(1):115–118, March 1951. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177729700.

Ogawa:1959:NCE

[Oga59] Junjiro Ogawa. A necessary condition for existence of regular and symmetrical experimental designs of triangular type, with partially balanced incomplete blocks. Annals of Mathematical Statistics, 30 (4):1063–1071, December 1959. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177706090.

Ogawa:1963:NDS

[Oga63] Junjiro Ogawa. On the null-distribution of the F -statistic in a randomized balanced incomplete block design under the Neyman model. Annals of Mathematical Statistics, 34(4):1558–1568, December 1963. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177703888.

Ogawa:1965:RAA

[OI65] Junjiro Ogawa and Goro Ishii. The relationship algebra and the analysis of variance of a partially balanced incomplete block design. Annals of Mathematical Statistics, 36(6):1815–1828, December 1965. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177699811.

Okamoto:1968:MEM

[OK68] Masashi Okamoto and Mitsuyo Kanazawa. Minimization of eigenvalues of a matrix and optimality of principal components. Annals of Mathematical Statistics, 39(3):859–863, June 1968. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177698317.

Okamoto:1963:AED

[Oka63] Masashi Okamoto. An asymptotic expansion for the distribution of the linear discriminant function. Annals of Mathematical Statistics, 34(4):1286–1301, December 1963. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177703864. See correction [Oka68].

Okamoto:1968:CNC

[Oka68] Masashi Okamoto. Correction notes: Correction to “An Asymptotic Expansion for the Distribution of the Linear Discriminant Function”. Annals of Mathematical Statistics, 39(4):1358–1359, August 1968. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177698265. See

[Oka63].

Olds:1938:DSS

[Old38] E. G. Olds. Distributions of sums of squares of rank differences for small numbers of individuals. Annals of Mathematical Statistics, 9 (2):133–148, June 1938. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177732332.

Olds:1940:MS

[Old40] E. G. Olds. On a method of sampling. Annals of Mathematical Statistics, 11(3):355–358, September 1940. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177731874.

Olds:1941:RCM

[Old41] Edwin G. Olds. Report of the Chicago meeting of the Institute. Annals of Mathematical Statistics, 12(4):468–469, December 1941. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177731687.

Olds:1942:ARS

[Old42a] Edwin G. Olds. Annual report of the Secretary-Treasurer of the Institute. Annals of Mathematical Statistics, 13(1):107–109, March 1942. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177731652.

Olds:1942:RDM

[Old42b] Edwin G. Olds. Report of the Dallas meeting of the Institute. Annals of Mathematical Statistics, 13(1):106–107, March 1942. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177731651.

Olds:1942:RNY

[Old42c] Edwin G. Olds. Report of the New York meeting of the Institute. Annals of Mathematical Statistics, 13(1):102–106, March 1942. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177731650.

Olds:1942:RPM

[Old42d] Edwin G. Olds. Report of the Poughkeepsie meeting of the Institute. Annals of Mathematical Statistics, 13(4):448–449, December 1942. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177731546.

Olds:1943:RWM

[Old43a] Edward G. Olds. Report of the Washington meeting of the Institute. Annals of Mathematical Statistics, 14(3):302–303, September 1943. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177731424.

Olds:1943:ARS

[Old43b] Edwin G. Olds. Annual report of the Secretary-Treasurer of the Institute. Annals of Mathematical Statistics, 14(1):100–102, March 1943. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177731500.

Olds:1943:RNB

[Old43c] Edwin G. Olds. Report on the New Brunswick meeting of the Institute. Annals of Mathematical Statistics, 14(4):444–445, December 1943. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177731366.

Olds:1944:ARS

[Old44] Edwin C. Olds. Annual report of the Secretary-Treasurer of the Institute. Annals of Mathematical Statistics, 15(1):113–114, March 1944. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177731324.

Olds:1949:SLS

[Old49] Edwin G. Olds. The 5% significance levels for sums of squares of rank differences and a correction. Annals of Mathematical Statistics, 20(1): 117–118, March 1949. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177730099.

Olds:1952:NCU

[Old52] Edwin G. Olds. A note on the convolution of uniform distributions. Annals of Mathematical Statistics, 23(2):282–285, June 1952. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177729446.

Olmstead:1940:NTO

[Olm40] P. S. Olmstead. Note on theoretical and observed distributions of repetitive occurrences. Annals of Mathematical Statistics, 11(3):363– 366, September 1940. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177731876.

Olmstead:1946:DSA

[Olm46] P. S. Olmstead. Distribution of sample arrangements for runs up and down. Annals of Mathematical Statistics, 17(1):24–33, March 1946. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177731019;   http://www.jstor.org/stable/2235901.

Olshen:1938:TPT

[Ols38] A. C. Olshen. Transformations of the Pearson Type III distribution. Annals of Mathematical Statistics, 9(3):176–200, September 1938.

CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177732309.

Olshevsky:1940:TPS

[Ols40] Louis Olshevsky. Two properties of sufficient statistics. Annals of Mathematical Statistics, 11(1):104–106, March 1940. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177731948.

Olshen:1967:SWT

[Ols67a] Richard A. Olshen. Sign and Wilcoxon tests for linearity. Annals of Mathematical Statistics, 38(6):1759–1769, December 1967. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177698610.

Olson:1967:TEO

[Ols67b] Milton Philip Olson. The theory of experiment: Operational definition of the probability space. Annals of Mathematical Statistics, 38 (2):401–409, April 1967. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177698955.

Olshen:1968:RFA

[Ols68] Richard A. Olshen. Representing finitely additive invariant probabilities. Annals of Mathematical Statistics, 39(6):2131–2135, December 1968. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177698044.

Olkin:1958:MTI

[OP58a] Ingram Olkin and John W. Pratt. A multivariate Tchebycheff inequality. Annals of Mathematical Statistics, 29(1):226–234, March 1958. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177706720.

Olkin:1958:UEC

[OP58b] Ingram Olkin and John W. Pratt. Unbiased estimation of certain correlation coefficients. Annals of Mathematical Statistics, 29(1):201– 211, March 1958. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177706717.

Olkin:1969:TEC

[OP69] I. Olkin and S. J. Press. Testing and estimation for a circular stationary model. Annals of Mathematical Statistics, 40(4):1358–1373, August 1969. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177697508.

OReilly:1972:USC

[OQ72] Federico J. O’Reilly and C. P. Quesenberry. Uniform strong consistency of Rao–Blackwell distribution function estimators. Annals of Mathematical Statistics, 43(5):1678–1679, October 1972. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177692401.

Olkin:1954:MDT

[OR54] I. Olkin and S. N. Roy. On multivariate distribution theory. Annals of Mathematical Statistics, 25(2):329–339, June 1954. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177728789.

Olkin:1962:CWD

[OR62] Ingram Olkin and Herman Rubin. A characterization of the Wishart distribution. Annals of Mathematical Statistics, 33(4):1272–1280, December 1962. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177704359.

Olkin:1964:MBD

[OR64] Ingram Olkin and Herman Rubin. Multivariate beta distributions and independence properties of the Wishart distribution. Annals of Mathematical Statistics, 35(1):261–269, March 1964. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177703748. See correc-

tion [OR66].

Olkin:1966:CNC

[OR66] Ingram Olkin and Herman Rubin. Correction note: Correction to “Multivariate Beta Distributions and Independence Properties of the Wishart Distribution”. Annals of Mathematical Statistics, 37(1):297, February 1966. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/

1177699625. See [OR64].

Ord:1968:DSD

[Ord68] J. K. Ord. The discrete Student’s t distribution. Annals of Mathematical Statistics, 39(5):1513–1516, October 1968. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177698133.

Ore:1964:RFD

[Ore64] Oystein Ore. Review: F. N. David, Games, Gods and Gambling. The Origins and History of Probability and Statistical Ideas from the Earliest Times to the Newtonian Era. Annals of Mathematical Statistics, 35(1):447, March 1964. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177703775.

Orey:1968:CPI

[Ore68a] Steven Orey. On continuity properties of infinitely divisible distribution functions. Annals of Mathematical Statistics, 39(3):936– 937, June 1968. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177698325.

Orey:1968:RJG

[Ore68b] Steven Orey. Review: J. G. Kemeny, L. J. Snell, A. W. Knapp, Denumerable Markov Chains. Annals of Mathematical Statistics, 39(6): 2168, December 1968. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177698056.

Owen:1962:MOS

[OS62] D. B. Owen and G. P. Steck. Moments of order statistics from the equicorrelated multivariate normal distribution. Annals of Mathematical Statistics, 33(4):1286–1291, December 1962. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177704361.

Olshen:1969:SFP

[OS69] R. A. Olshen and D. O. Siegmund. Some first passage problems for S_{n}/n^{\frac{1}{2}} . Annals of Mathematical Statistics, 40(2):648–652, April 1969. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177697733.

Olkin:1970:EWT

[OS70a] I. Olkin and S. S. Shrikhande. An extension of Wilks’ test for the equality of means. Annals of Mathematical Statistics, 41(2):683– 687, April 1970. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177697114.

Ornstein:1970:OTC

[OS70b] Donald Ornstein and Louis Sucheston. An operator theorem on L_{1} convergence to zero with applications to Markov kernels. Annals of Mathematical Statistics, 41(5):1631–1639, October 1970. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177696806.

OToole:1931:SFS

[O’T31] A. L. O’Toole. On symmetric functions and symmetric functions of symmetric functions. Annals of Mathematical Statistics, 2(2):102– 149, May 1931. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177732992.

OToole:1932:SFM

[O’T32] A. L. O’Toole. On symmetric functions of more than one variable and of frequency functions. Annals of Mathematical Statistics, 3(1): 56–63, February 1932. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177732933.

OToole:1933:MDC

[O’T33a] A. L. O’Toole. A method of determining the constants in the bimodal fourth degree exponential function. Annals of Mathematical Statistics, 4(2):79–93, May 1933. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177732802.

OToole:1933:DAC

[O’T33b] A. L. O’Toole. On the degree of approximation of certain quadrature formulas. Annals of Mathematical Statistics, 4(2):143–153, May 1933. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177732806.

OToole:1933:SCW

[O’T33c] A. L. O’Toole. On the system of curves for which the method of moments is the best method of fitting. Annals of Mathematical Statistics, 4(1):1–29, February 1933. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177732820.

OToole:1934:EBV

[O’T34] A. L. O’Toole. Editorial: On a best value of R in samples of R from a finite population of N. Annals of Mathematical Statistics, 5(2): 146–152, June 1934. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177732713.

Olmstead:1947:CTA

[OT47] Paul S. Olmstead and John W. Tukey. A corner test for association. Annals of Mathematical Statistics, 18(4):495–513, December 1947. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177730341.

Olkin:1961:MCM

[OT61] I. Olkin and R. F. Tate. Multivariate correlation models with mixed discrete and continuous variables. Annals of Mathematical Statistics, 32(2):448–465, June 1961. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177705052. See correction [OT65].

Olkin:1965:CNC

[OT65] I. Olkin and R. F. Tate. Correction note: Correction to “Multivariate Correlation Models with Mixed Discrete and Continuous Variables”. Annals of Mathematical Statistics, 36(1):343–344, February 1965. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177700300. See

[OT61].

Otter:1949:MP

[Ott49] Richard Otter. The multiplicative process. Annals of Mathematical Statistics, 20(2):206–224, June 1949. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177730031.

Owen:1953:DST

[Owe53] Donald B. Owen. A double sample test procedure. Annals of Mathematical Statistics, 24(3):449–457, September 1953. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177728984.

Owen:1956:TCB

[Owe56] Donald B. Owen. Tables for computing bivariate normal probabilities. Annals of Mathematical Statistics, 27(4):1075–1090, December 1956. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177728074.

Owen:1966:RBV

[Owe66] D. B. Owen. Review: L. N. Bolshev, V. N. Smirnov, Mathematical Statistical Tables. Annals of Mathematical Statistics, 37(3):757– 758, June 1966. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177699481.

Owen:1967:RPH

[Owe67a] D. B. Owen. Review: E. S. Pearson, H. O. Hartley, Biometrika Tables for Statisticians, Volume I. Annals of Mathematical Statistics, 38(2): 632–633, April 1967. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177698992.

Owen:1967:RBB

[Owe67b] D. B. Owen. Review: L. S. Bark, L. N. Bolshev, P. I. Kuznetsov, A. P. Cherenkov, Tables of the Rayleigh–Rice Distributions. Annals of Mathematical Statistics, 38(1):294–296, February 1967. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177699089.

Owen:1970:ODT

[Owe70] R. J. Owen. The optimum design of a two-factor experiment using prior information. Annals of Mathematical Statistics, 41(6):1917– 1934, December 1970. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177696693.

Pillai:1969:DRR

[PAAJ69] K. C. S. Pillai, S. Al-Ani, and G. M. Jouris. On the distributions of the ratios of the roots of a covariance matrix and Wilks’ criterion for tests of three hypotheses. Annals of Mathematical Statistics, 40 (6):2033–2040, December 1969. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177697283.

Pachares:1955:NDD

[Pac55] James Pachares. Note on the distribution of a definite quadratic form. Annals of Mathematical Statistics, 26(1):128–131, March 1955. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177728601.

Pachares:1961:TUT

[Pac61] James Pachares. Tables for unbiased tests on the variance of a normal population. Annals of Mathematical Statistics, 32(1):84–87, March 1961. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177705141.

Pakes:1971:SCC

[Pak71] A. G. Pakes. The serial correlation coefficients of waiting times in the stationary GI/M/1 queue. Annals of Mathematical Statistics, 42 (5):1727–1734, October 1971. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177693171.

Palmer:1930:EUS

[Pal30] Edgar Z. Palmer. Error and unreliability in seasonals. Annals of Mathematical Statistics, 1(4):345–351, November 1930. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177733064.

Panzone:1967:APC

[Pan67] Rafael Panzone. Alternative proofs for certain upcrossing inequalities. Annals of Mathematical Statistics, 38(3):735–741, June 1967. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177698866.

Parzen:1957:CLT

[Par57a] Emanuel Parzen. A Central Limit Theorem for multilinear stochastic processes. Annals of Mathematical Statistics, 28(1):252–256, March 1957. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177707050.

Parzen:1957:CESb

[Par57b] Emanuel Parzen. On choosing an estimate of the spectral density function of a stationary time series. Annals of Mathematical Statistics, 28(4):921–932, December 1957. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177706793.

Parzen:1957:CESa

[Par57c] Emanuel Parzen. On consistent estimates of the spectrum of a stationary time series. Annals of Mathematical Statistics, 28(2):329– 348, June 1957. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177706962.

Parzen:1958:CSP

[Par58] Emanuel Parzen. Conditions that a stochastic process be ergodic. Annals of Mathematical Statistics, 29(1):299–301, March 1958. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177706731.

Parthasarathy:1960:ESS

[Par60] K. R. Parthasarathy. On the estimation of the spectrum of a stationary stochastic process. Annals of Mathematical Statistics, 31(3):568– 573, September 1960. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177705784.

Parzen:1961:ATS

[Par61] Emanuel Parzen. An approach to time series analysis. Annals of Mathematical Statistics, 32(4):951–989, December 1961. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177704840.

Parzen:1962:EPD

[Par62] Emanuel Parzen. On estimation of a probability density function and mode. Annals of Mathematical Statistics, 33(3):1065–1076, September 1962. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177704472.

Park:1964:VNC

[Par64] John H. Park, Jr. Variations of the non-central t and beta distributions. Annals of Mathematical Statistics, 35(4):1583–1593, December 1964. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177700383.

Park:1970:MPG

[Par70] Won Joon Park. A multi-parameter Gaussian process. Annals of Mathematical Statistics, 41(5):1582–1595, October 1970. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177696802.

Park:1972:NCO

[Par72a] C. J. Park. A note on the classical occupancy problem. Annals of Mathematical Statistics, 43(5):1698–1701, October 1972. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177692405.

Park:1972:CEK

[Par72b] Won Joon Park. A consistent estimation of kernel functions in the multiple Wiener integrals. Annals of Mathematical Statistics, 43(3): 983–987, June 1972. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177692561.

Patel:1962:CFR

[Pat62] M. S. Patel. On constructing the fractional replicates of the 2^{m} designs with blocks. Annals of Mathematical Statistics, 33(4):1440– 1449, December 1962. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177704375.

Patil:1963:MVU

[Pat63] G. P. Patil. Minimum variance unbiased estimation and certain problems of additive number theory. Annals of Mathematical Statistics, 34 (3):1050–1056, September 1963. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177704029.

Pathak:1964:SSP

[Pat64a] P. K. Pathak. On sampling schemes providing unbiased ratio estimators. Annals of Mathematical Statistics, 35(1):222–231, March 1964. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177703746.

Pathak:1964:SST

[Pat64b] P. K. Pathak. Sufficiency in sampling theory. Annals of Mathematical Statistics, 35(2):795–808, June 1964. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177703579.

Paulson:1941:CLR

[Pau41] Edward Paulson. On certain likelihood-ratio tests associated with the exponential distribution. Annals of Mathematical Statistics, 12 (3):301–306, September 1941. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177731712.

Paulson:1942:ANA

[Pau42a] Edward Paulson. An approximate normalization of the analysis of variance distribution. Annals of Mathematical Statistics, 13(2):233– 235, June 1942. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177731608.

Paulson:1942:NES

[Pau42b] Edward Paulson. A note on the estimation of some mean values for a bivariate distribution. Annals of Mathematical Statistics, 13(4):440– 445, December 1942. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177731543.

Paulson:1943:NTL

[Pau43] Edward Paulson. A note on tolerance limits. Annals of Mathematical Statistics, 14(1):90–93, March 1943. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177731496.

Paulson:1947:NEW

[Pau47] Edward Paulson. A note on the efficiency of the Wald sequential test. Annals of Mathematical Statistics, 18(3):447–450, September 1947. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177730393.

Paulson:1949:MDP

[Pau49] Edward Paulson. A multiple decision procedure for certain problems in the analysis of variance. Annals of Mathematical Statistics, 20 (1):95–98, March 1949. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177730094.

Paull:1950:PTP

[Pau50] A. E. Paull. On a preliminary test for pooling mean squares in the analysis of variance. Annals of Mathematical Statistics, 21(4):539– 556, December 1950. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177729750.

Paulson:1952:CSE

[Pau52a] Edward Paulson. On the comparison of several experimental categories with a control. Annals of Mathematical Statistics, 23(2): 239–246, June 1952. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177729440.

Paulson:1952:OSS

[Pau52b] Edward Paulson. An optimum solution to the k-sample slippage problem for the normal distribution. Annals of Mathematical Statistics, 23(4):610–616, December 1952. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177729340.

Paulson:1962:SPC

[Pau62] Edward Paulson. A sequential procedure for comparing several experimental categories with a standard or control. Annals of Mathematical Statistics, 33(2):438–443, June 1962. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177704570.

Paulson:1963:SDP

[Pau63] Edward Paulson. A sequential decision procedure for choosing one of k hypotheses concerning the unknown mean of a normal distribution. Annals of Mathematical Statistics, 34(2):549–554, June 1963. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177704167.

Paulson:1964:SEC

[Pau64a] Edward Paulson. Sequential estimation and closed sequential decision procedures. Annals of Mathematical Statistics, 35(3):1048–1058, September 1964. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177703263.

Paulson:1964:SPS

[Pau64b] Edward Paulson. A sequential procedure for selecting the population with the largest mean from k normal populations. Annals of Mathematical Statistics, 35(1):174–180, March 1964. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177703739.

Paulson:1967:SPS

[Pau67] Edward Paulson. Sequential procedures for selecting the best one of several binomial populations. Annals of Mathematical Statistics, 38(1):117–123, February 1967. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177699062.

Paulson:1969:SIE

[Pau69] Edward Paulson. Sequential interval estimation for the means of normal populations. Annals of Mathematical Statistics, 40(2):509– 516, April 1969. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177697718.

Paulson:1970:NSP

[Pau70] Edward Paulson. A note on a sequential procedure for comparing two Koopman–Darmois populations. Annals of Mathematical Statistics, 41(5):1756–1759, October 1970. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177696821.

Paz:1970:ETI

[Paz70] A. Paz. Ergodic theorems for infinite probabilistic tables. Annals of Mathematical Statistics, 41(2):539–550, April 1970. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177697094. See note [Mad71].

Patil:1970:CPM

[PB70] G. P. Patil and M. T. Boswell. A characteristic property of the multivariate normal density function and some of its applications. Annals of Mathematical Statistics, 41(6):1970–1977, December 1970. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177696697.

Pierce:1971:NBF

[PB71] Donald A. Pierce and David A. Bogdanoff. Note on Bayes-fiducial intervals for problems of location and scale. Annals of Mathematical Statistics, 42(2):833–836, April 1971. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177693443.

Pearson:1962:STS

[Pea62] E. S. Pearson. Some thoughts on statistical inference. Annals of Mathematical Statistics, 33(2):394–403, June 1962. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177704566.

Pearson:1964:RDB

[Pea64] E. S. Pearson. Review: Donald B. Owen, Handbook of Statistical Tables. Annals of Mathematical Statistics, 35(2):903–906, June 1964. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177703599.

Peiser:1943:AFS

[Pei43] Alfred M. Peiser. Asymptotic formulas for significance levels of certain distributions. Annals of Mathematical Statistics, 14(1):56–62, March 1943. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/

1177731492. See correction [Pei49].

Peiser:1949:CAF

[Pei49] A. M. Peiser. Correction to “Asymptotic Formulas for Significance Levels of Certain Distributions”. Annals of Mathematical Statistics, 20(1):128–129, March 1949. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177730105. See [Pei43].

Perlman:1969:OST

[Per69a] Michael D. Perlman. One-sided testing problems in multivariate analysis. Annals of Mathematical Statistics, 40(2):549–567, April 1969. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177697723. See

correction [Per71].

Perng:1969:CAE

[Per69b] S. K. Perng. A comparison of the asymptotic expected sample sizes of two sequential procedures for ranking problem. Annals of Mathematical Statistics, 40(6):2198–2202, December 1969. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177697299.

Perng:1970:IVG

[Per70] S. K. Perng. Inadmissibility of various “good” statistical procedures which are translation invariant. Annals of Mathematical Statistics, 41(4):1311–1321, August 1970. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177696905.

Perlman:1971:CNC

[Per71] Michael D. Perlman. Correction notes: Correction to “One-Sided Testing Problems in Multivariate Analysis”. Annals of Mathematical Statistics, 42(5):1777, October 1971. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177693184. See [Per69a].

Peterson:1954:DUP

[Pet54] Raymond P. Peterson. Density unbiased point estimates. Annals of Mathematical Statistics, 25(2):398–401, June 1954. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177728799.

Petrie:1969:PFF

[Pet69] T. Petrie. Probabilistic functions of finite state Markov chains. Annals of Mathematical Statistics, 40(1):97–115, February 1969. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177697807.

Paik:1970:RPS

[PF70] U. B. Paik and W. T. Federer. A randomized procedure of Saturated main effect fractional replicates. Annals of Mathematical Statistics, 41 (2):369–375, April 1970. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177697077.

Paik:1972:RPS

[PF72] U. B. Paik and W. T. Federer. On a randomized procedure for saturated fractional replicates in a 2^{n}-factorial. Annals of Mathematical Statistics, 43(4):1346–1351, August 1972. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177692488.

Pfanzagl:1964:TSS

[Pfa64] J. Pfanzagl. On the topological structure of some ordered families of distributions. Annals of Mathematical Statistics, 35(3):1216–1228, September 1964. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177703279.

Pfanzagl:1967:CCE

[Pfa67a] J. Pfanzagl. Characterizations of conditional expectations. Annals of Mathematical Statistics, 38(2):415–421, April 1967. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177698957.

Pfanzagl:1967:TLM

[Pfa67b] J. Pfanzagl. A technical lemma for monotone likelihood ratio families. Annals of Mathematical Statistics, 38(2):611–612, April 1967. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177698982.

Pfanzagl:1969:CEL

[Pfa69a] J. Pfanzagl. Consistent estimation of a location parameter in the presence of an incidental scale parameter. Annals of Mathematical Statistics, 40(4):1353–1357, August 1969. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177697507.

Pfanzagl:1969:FRT

[Pfa69b] J. Pfanzagl. Further remarks on topology and convergence in some ordered families of distributions. Annals of Mathematical Statistics, 40(1):51–65, February 1969. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177697803.

Pfanzagl:1970:AEM

[Pfa70] J. Pfanzagl. On the asymptotic efficiency of median unbiased estimates. Annals of Mathematical Statistics, 41(5):1500–1509, October 1970. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177696795.

Pfanzagl:1971:CLD

[Pfa71] J. Pfanzagl. A counterexample to a lemma of L. D. Brown. Annals of Mathematical Statistics, 42(1):373–375, February 1971. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177693524.

Pfanzagl:1972:TGS

[Pfa72] J. Pfanzagl. Transformation groups and sufficient statistics. Annals of Mathematical Statistics, 43(2):553–568, April 1972. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177692636.

Pillai:1968:NCD

[PG68] K. C. Sreedharan Pillai and Arjun K. Gupta. On the non-central distribution of the second elementary symmetric function of the roots of a matrix. Annals of Mathematical Statistics, 39(3):833– 839, June 1968. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177698314.

Han:1969:DDF

[pH69] Chien pai Han. Distribution of discriminant function when covariance matrices are proportional. Annals of Mathematical Statistics, 40(3): 979–985, June 1969. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177697602.

Phatarfod:1963:AMS

[Pha63] R. M. Phatarfod. Application of methods in sequential analysis to dam theory. Annals of Mathematical Statistics, 34(4):1588–1592, December 1963. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177703892.

Philipp:1969:LIL

[Phi69a] Walter Philipp. The law of the iterated logarithm for mixing stochastic processes. Annals of Mathematical Statistics, 40(6):1985–1991, December 1969. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177697280.

Philipp:1969:RCL

[Phi69b] Walter Philipp. The remainder in the Central Limit Theorem for mixing stochastic processes. Annals of Mathematical Statistics, 40 (2):601–609, April 1969. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177697728.

Picard:1951:NMV

[Pic51] H. C. Picard. A note on the maximum value of kurtosis. Annals of Mathematical Statistics, 22(3):480–482, September 1951. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177729602.

Pickands:1967:SSM

[Pic67] James Pickands III. Sample sequences of maxima. Annals of Mathematical Statistics, 38(5):1570–1574, October 1967. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177698711.

Pickands:1968:MCS

[Pic68] James Pickands III. Moment convergence of sample extremes. Annals of Mathematical Statistics, 39(3):881–889, June 1968. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177698320.

Pickands:1969:EEP

[Pic69] James Pickands III. Efficient estimation of a probability density function. Annals of Mathematical Statistics, 40(3):854–864, June 1969. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177697593.

Pickands:1970:SER

[Pic70] James Pickands III. Spectral estimation with random truncation. Annals of Mathematical Statistics, 41(1):44–58, February 1970. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177697187.

Pierce:1940:SUF

[Pie40] Joseph A. Pierce. A study of a universe of n finite populations with application to moment-function adjustments for grouped data. Annals of Mathematical Statistics, 11(3):311–334, September 1940. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177731870.

Pierre:1969:CNC

[Pie69a] Percy A. Pierre. Correction note: Correction to “New Conditions for Central Limit Theorems”. Annals of Mathematical Statistics, 40 (5):1855, October 1969. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177697404. See [Pie69b].

Pierre:1969:NCC

[Pie69b] Percy A. Pierre. New conditions for Central Limit Theorems. Annals of Mathematical Statistics, 40(1):319–321, February 1969. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177697830. See correc-

tion [Pie69a].

Pierce:1970:DBA

[Pie70] David A. Pierce. A duality between autoregressive and moving average processes concerning their least squares parameter estimates. Annals of Mathematical Statistics, 41(2):422–426, April 1970. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177697081.

Pillai:1950:DMS

[Pil50] K. C. S. Pillai. On the distributions of midrange and semi-range in samples from a normal population. Annals of Mathematical Statistics, 21(1):100–105, March 1950. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177729888.

Pillai:1951:DAS

[Pil51] K. C. S. Pillai. On the distribution of an analogue of Student’s t. Annals of Mathematical Statistics, 22(3):469–472, September 1951. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177729597.

Pillai:1955:SNT

[Pil55] K. C. S. Pillai. Some new test criteria in multivariate analysis. Annals of Mathematical Statistics, 26(1):117–121, March 1955. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177728599.

Pillai:1956:SRU

[Pil56] K. C. S. Pillai. Some results useful in multivariate analysis. Annals of Mathematical Statistics, 27(4):1106–1114, December 1956. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177728076.

Pillai:1964:MES

[Pil64] K. C. Sreedharan Pillai. On the moments of elementary symmetric functions of the roots of two matrices. Annals of Mathematical Statistics, 35(4):1704–1712, December 1964. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177700392.

Pillai:1967:DLR

[Pil67] K. C. Sreedharan Pillai. On the distribution of the largest root of a matrix in multivariate analysis. Annals of Mathematical Statistics, 38 (2):616–617, April 1967. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177698984.

Pillai:1968:MGF

[Pil68] K. C. Sreedharan Pillai. On the moment generating function of Pillai’s V^{(s)} criterion. Annals of Mathematical Statistics, 39(3):877– 880, June 1968. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177698319.

Pillai:1971:SSL

[Pil71] R. N. Pillai. Semi stable laws as limit distributions. Annals of Mathematical Statistics, 42(2):780–783, April 1971. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177693431.

Pinney:1947:FCZ

[Pin47] Edmund Pinney. Fitting curves with zero or infinite end points. Annals of Mathematical Statistics, 18(1):127–131, March 1947. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177730501.

Pinkham:1961:DFS

[Pin61] Roger S. Pinkham. On the distribution of first significant digits. Annals of Mathematical Statistics, 32(4):1223–1230, December 1961. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990

(electronic). URL http://links.jstor.org/sici?sici=0003-
4851%28196112%2932%3A4%3C1223%3AOTDOFS%3E2.0.CO%3B2-T;

http://projecteuclid.org/euclid.aoms/1177704862;   http:
//www.jstor.org/stable/2237922.

Pitman:1956:DCF

[Pit56] E. J. G. Pitman. On the derivatives of a characteristic function at the origin. Annals of Mathematical Statistics, 27(4):1156–1160, December 1956. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177728083.

Pitcher:1957:SMA

[Pit57] T. S. Pitcher. Sets of measures not admitting necessity and sufficient statistics or subfields. Annals of Mathematical Statistics, 28(1):267– 268, March 1957. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177707055.

Pitcher:1963:SFP

[Pit63] T. S. Pitcher. On the sample functions of processes which can be added to a Gaussian process. Annals of Mathematical Statistics, 34(1):329–333, March 1963. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177704271.

Pitcher:1964:AIS

[Pit64] Tom S. Pitcher. On adding independent stochastic processes. Annals of Mathematical Statistics, 35(2):872–873, June 1964. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177703589.

Pitcher:1965:BLR

[Pit65] T. S. Pitcher. The behavior of likelihood ratios of stochastic processes related by groups of transformations. Annals of Mathematical Statistics, 36(2):529–534, April 1965. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177700163.

Pitcher:1968:ECM

[Pit68] T. S. Pitcher. The ε-entropy of certain measures on [0, 1]. Annals of Mathematical Statistics, 39(4):1310–1315, August 1968. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177698256.

Patil:1970:FRM

[PJ70] G. P. Patil and S. W. Joshi. Further results on minimum variance unbiased estimation and additive number theory. Annals of Mathematical Statistics, 41(2):567–575, April 1970. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177697096.

Pillai:1971:SDP

[PJ71] K. C. S. Pillai and G. M. Jouris. Some distribution problems in the multivariate complex Gaussian case. Annals of Mathematical Statistics, 42(2):517–525, April 1971. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177693402.

Peach:1946:NSI

[PL46] Paul Peach and S. B. Littauer. A note on sampling inspection. Annals of Mathematical Statistics, 17(1):81–84, March 1946. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177731027.

Patil:1970:DRM

[PL70] S. A. Patil and S. H. Liao. The distribution of the ratios of means to the square root of the sum of variances of a bivariate normal sample. Annals of Mathematical Statistics, 41(2):723–728, April 1970. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177697123.

Patil:1971:APD

[PL71] S. A. Patil and S. H. Liao. Acknowledgment of priority to: “The Distribution of the Ratio of Means to the Square Root of the Sum of Variances of a Bivariate Normal Sample”. Annals of Mathematical Statistics, 42(4):1461, August 1971. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177693263.

Plackett:1948:BMS

[Pla48] R. L. Plackett. Boundaries of minimum size in binomial sampling. Annals of Mathematical Statistics, 19(4):575–580, December 1948. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177730153.

Plackett:1958:LEC

[Pla58] R. L. Plackett. Linear estimation from censored data. Annals of Mathematical Statistics, 29(1):131–142, March 1958. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177706710.

Plachky:1971:TGS

[Pla71] Detlef Plachky. On a theorem of G. L. Sievers. Annals of Mathematical Statistics, 42(4):1442–1443, August 1971. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177693257.

Pillai:1959:MTM

[PM59] K. C. S. Pillai and Tito A. Mijares. On the moments of the trace of a matrix and approximations to its distribution. Annals of Mathematical Statistics, 30(4):1135–1140, December 1959. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177706097.

Pillai:1971:DST

[PN71] K. C. S. Pillai and B. N. Nagarsenker. On the distribution of the sphericity test criterion in classical and complex normal populations having unknown covariance matrices. Annals of Mathematical Statistics, 42(2):764–767, April 1971. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177693427.

Pollard:1934:RSM

[Pol34] Harry S. Pollard. On the relative stability of the median and arithmetic mean, with particular reference to certain frequency distributions which can be dissected into normal distributions. Annals of Mathematical Statistics, 5(3):227–262, September 1934. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177732678.

Polfeldt:1970:MVO

[Pol70a] Thomas Polfeldt. Minimum variance order when estimating the location of an irregularity in the density. Annals of Mathematical Statistics, 41(2):673–679, April 1970. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177697112.

Polfeldt:1970:OMV

[Pol70b] Thomas Polfeldt. The order of the minimum variance in a non-regular case. Annals of Mathematical Statistics, 41(2):667–672, April 1970. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177697111.

Pollak:1972:NID

[Pol72] Moshe Pollak. A note on infinitely divisible random vectors. Annals of Mathematical Statistics, 43(2):673–675, April 1972. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177692652.

Port:1964:EPH

[Por64a] Sidney C. Port. Escape probability for a half line. Annals of Mathematical Statistics, 35(3):1351–1355, September 1964. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177703294.

Port:1964:STF

[Por64b] Sidney C. Port. Some theorems on functionals of Markov chains. Annals of Mathematical Statistics, 35(3):1275–1290, September 1964. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177703283.

Port:1965:SPP

[Por65] Sidney C. Port. A simple probabilistic proof of the discrete generalized renewal theorem. Annals of Mathematical Statistics, 36(4): 1294–1297, August 1965. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177700002.

Port:1966:SDM

[Por66] S. C. Port. A system of denumerably many transient Markov chains. Annals of Mathematical Statistics, 37(2):406–411, April 1966. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177699522.

Port:1967:HPS

[Por67] Sidney C. Port. On hitting places for stable processes. Annals of Mathematical Statistics, 38(4):1021–1026, August 1967. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177698769.

Port:1968:RHP

[Por68] Sidney C. Port. A remark on hitting places for transient stable process. Annals of Mathematical Statistics, 39(2):365–371, April 1968. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177698397.

Port:1970:EDI

[Por70] Sidney C. Port. The exit distribution of an interval for completely asymmetric stable processes. Annals of Mathematical Statistics, 41 (1):39–43, February 1970. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177697186.

Portnoy:1971:FBE

[Por71] Stephen Portnoy. Formal Bayes estimation with application to a random effects model. Annals of Mathematical Statistics, 42(4):1379– 1402, August 1971. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177693250.

Potthoff:1963:UWS

[Pot63] Richard F. Potthoff. Use of the Wilcoxon statistic for a generalized Behrens–Fisher problem. Annals of Mathematical Statistics, 34(4): 1596–1599, December 1963. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177703894.

Powell:1930:SIM

[Pow30] Ralph W. Powell. Successive integration as a method for finding long period cycles. Annals of Mathematical Statistics, 1(2):123– 136, May 1930. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177733127.

Puri:1969:MDP

[PP69] Madan L. Puri and Prem S. Puri. Multiple decision procedures based on ranks for certain problems in analysis of variance. Annals of Mathematical Statistics, 40(2):619–632, April 1969. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177697730.

Pillai:1954:DRI

[PR54] K. C. S. Pillai and K. V. Ramachandran. On the distribution of the ratio of the ith observation in an ordered sample from a normal population to an independent estimate of the standard deviation. Annals of Mathematical Statistics, 25(3):565–572, September 1954. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177728724.

Pyke:1968:CMN

[PR68] Ronald Pyke and David Root. On convergence in r-mean of normalized partial sums. Annals of Mathematical Statistics, 39(2):379– 381, April 1968. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177698400.

Puri:1970:CBA

[PR70] Prem S. Puri and Herman Rubin. A characterization based on the absolute difference of two I.I.D. random variables. Annals of Mathematical Statistics, 41(6):2113–2122, December 1970. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177696709.

Pathak:1971:SRI

[PR71a] P. K. Pathak and Neil W. Rickert. Some results on invariant sets for translation parameter family of probability measures — I. Annals of Mathematical Statistics, 42(5):1562–1568, October 1971. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177693154.

Posner:1971:EED

[PR71b] Edward C. Posner and Eugene R. Rodemich. Epsilon entropy and data compression. Annals of Mathematical Statistics, 42(6):2079– 2125, December 1971. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177693077.

Prabhu:1958:SER

[Pra58a] N. U. Prabhu. Some exact results for the finite dam. Annals of Mathematical Statistics, 29(4):1234–1243, December 1958. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177706454.

Pratt:1958:AOS

[Pra58b] J. W. Pratt. Admissible one-sided tests for the mean of a rectangular distribution. Annals of Mathematical Statistics, 29(4):1268–1271, December 1958. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177706459.

Pratt:1959:GCP

[Pra59] John W. Pratt. On a general concept of “In Probability”. Annals of Mathematical Statistics, 30(2):549–558, June 1959. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177706267.

Prabhu:1960:AST

[Pra60a] N. U. Prabhu. Application of storage theory to queues with Poisson arrivals. Annals of Mathematical Statistics, 31(2):475–482, June 1960. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177705911.

Pratt:1960:ILI

[Pra60b] John W. Pratt. On interchanging limits and integrals. Annals of Mathematical Statistics, 31(1):74–77, March 1960. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177705988.

Prabhu:1961:RPC

[Pra61] N. U. Prabhu. On the ruin problem of collective risk theory. Annals of Mathematical Statistics, 32(3):757–764, September 1961. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177704970.

Pratt:1962:NUT

[Pra62] John W. Pratt. A note on unbiased tests. Annals of Mathematical Statistics, 33(1):292–294, March 1962. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177704733.

Pratt:1963:SCI

[Pra63] John W. Pratt. Shorter confidence intervals for the mean of a normal distribution with known variance. Annals of Mathematical Statistics, 34(2):574–586, June 1963. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177704170.

Prabhu:1966:RDR

[Pra66a] N. U. Prabhu. Review: D. R. Cox, H. D. Miller, The Theory of Stochastic Processes. Annals of Mathematical Statistics, 37(6):1848– 1851, December 1966. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177699182.

Pratt:1966:API

[Pra66b] John W. Pratt. Acknowledgment of priority: On interchanging limits and integrals. Annals of Mathematical Statistics, 37(5):1407, October 1966. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177699286.

Press:1966:LCN

[Pre66] Sheldon James Press. Linear combinations of non-central chisquare variates. Annals of Mathematical Statistics, 37(2):480–487, April 1966. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177699531.

Preece:1969:NCR

[Pre69a] D. A. Preece. Near-cyclic representations for some resolution VI fractional factorial plans. Annals of Mathematical Statistics, 40(5):1840– 1843, October 1969. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177697398.

Press:1969:SC

[Pre69b] S. James Press. On serial correlation. Annals of Mathematical Statistics, 40(1):188–196, February 1969. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177697815.

Preston:1972:CPS

[Pre72] Christopher Preston. Continuity properties of some Gaussian processes. Annals of Mathematical Statistics, 43(1):285–292, February 1972. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177692721.

Proschan:1965:PDC

[Pro65] Frank Proschan. Peakedness of distributions of convex combinations. Annals of Mathematical Statistics, 36(6):1703–1706, December 1965.

CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177699798.

Posner:1967:EES

[PRR67] Edward C. Posner, Eugene R. Rodemich, and Howard Rumsey, Jr. Epsilon entropy of stochastic processes. Annals of Mathematical Statistics, 38(4):1000–1020, August 1967. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177698768.

Posner:1969:EEG

[PRR69a] Edward C. Posner, Eugene R. Rodemich, and Howard Rumsey. Epsilon entropy of Gaussian processes. Annals of Mathematical Statistics, 40(4):1272–1296, August 1969. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177697502.

Posner:1969:PEG

[PRR69b] Edward C. Posner, Eugene R. Rodemich, and Howard Rumsey, Jr. Product entropy of Gaussian distributions. Annals of Mathematical Statistics, 40(3):870–904, June 1969. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177697595.

Pruitt:1964:ENN

[Pru64] William E. Pruitt. Eigenvalues of non-negative matrices. Annals of Mathematical Statistics, 35(4):1797–1800, December 1964. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177700401.

Pruitt:1971:RME

[Pru71] William E. Pruitt. Review: Martin Eisen, Introduction to Mathematical Probability Theory. Annals of Mathematical Statistics, 42 (6):2159–2161, December 1971. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177693086.

Pyke:1964:LTM

[PS64] Ronald Pyke and Ronald Schaufele. Limit theorems for Markov renewal processes. Annals of Mathematical Statistics, 35(4):1746–1764, December 1964. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177700397.

Pyke:1966:EUS

[PS66] Ronald Pyke and Ronald Schaufele. The existence and uniqueness of stationary measures for Markov renewal processes. Annals of Mathematical Statistics, 37(6):1439–1462, December 1966. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177699138.

Port:1967:LST

[PS67a] Sidney C. Port and Charles J. Stone. A large sample test for the independence of two renewal processes. Annals of Mathematical Statistics, 38(4):1037–1041, August 1967. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177698772.

Puri:1967:REI

[PS67b] Madan Lal Puri and Pranab Kumar Sen. On robust estimation in incomplete block designs. Annals of Mathematical Statistics, 38 (5):1587–1591, October 1967. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177698715.

Puri:1968:CST

[PS68a] Madan Lal Puri and Pranab Kumar Sen. On Chernoff–Savage tests for ordered alternatives in randomized blocks. Annals of Mathematical Statistics, 39(3):967–972, June 1968. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177698329.

Pyke:1968:WCC

[PS68b] Ronald Pyke and Galen R. Shorack. Weak convergence and a Chernoff–Savage theorem for random sample sizes. Annals of Mathematical Statistics, 39(5):1675–1685, October 1968. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177698149.

Pyke:1968:WCT

[PS68c] Ronald Pyke and Galen R. Shorack. Weak convergence of a twosample empirical process and a new approach to Chernoff–Savage theorems. Annals of Mathematical Statistics, 39(3):755–771, June 1968. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177698309.

Port:1969:ACP

[PS69a] S. C. Port and C. J. Stone. The asymmetric Cauchy processes on the line. Annals of Mathematical Statistics, 40(1):137–143, February 1969. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177697810.

Puri:1969:ACB

[PS69b] Madan Lal Puri and Pranab Kumar Sen. Analysis of covariance based on general rank scores. Annals of Mathematical Statistics, 40 (2):610–618, April 1969. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177697729.

Puri:1969:CRO

[PS69c] Madan Lal Puri and Pranab Kumar Sen. A class of rank order tests for a general linear hypothesis. Annals of Mathematical Statistics, 40(4):1325–1343, August 1969. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177697505.

Pyke:1969:NCS

[PS69d] Ronald Pyke and Galen R. Shorack. A note on Chernoff–Savage theorems. Annals of Mathematical Statistics, 40(3):1116–1119, June 1969. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177697622.

Pakshirajan:1970:LIL

[PS70] R. P. Pakshirajan and M. Sreehari. The law of the iterated logarithm for a Markov process. Annals of Mathematical Statistics, 41(3):945– 955, June 1970. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177696971.

Portnoy:1971:IBI

[PS71] Stephen Portnoy and Charles Stein. Inadmissibility of the best invariant test in three or more dimensions. Annals of Mathematical Statistics, 42(2):799–801, April 1971. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177693435.

Pop-Stojanovic:1972:RDW

[PS72] Zoran R. Pop-Stojanović. Riesz decomposition for weak Banachvalued quasi-martingales. Annals of Mathematical Statistics, 43(3): 1020–1026, June 1972. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177692568.

Puri:1964:AEC

[Pur64] Madan Lal Puri. Asymptotic efficiency of a class of c-sample tests. Annals of Mathematical Statistics, 35(1):102–121, March 1964. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177703733.

Putter:1955:TTS

[Put55] Joseph Putter. The treatment of ties in some nonparametric tests. Annals of Mathematical Statistics, 26(3):368–386, September 1955. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177728485.

Pinkham:1963:TAD

[PW63] R. S. Pinkham and M. B. Wilk. Tail areas of the t-distribution from a Mills’-ratio-like expansion. Annals of Mathematical Statistics, 34(1): 335–337, March 1963. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177704273.

Pitman:1967:CDF

[PW67] E. J. G. Pitman and E. J. Williams. Cauchy-distributed functions of Cauchy variates. Annals of Mathematical Statistics, 38(3):916– 918, June 1967. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177698885.

Pyke:1958:RPR

[Pyk58] Ronald Pyke. On renewal processes related to Type I and Type II counter models. Annals of Mathematical Statistics, 29(3):737–754, September 1958. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177706533.

Pyke:1959:SIP

[Pyk59] Ronald Pyke. The supremum and infimum of the Poisson process. Annals of Mathematical Statistics, 30(2):568–576, June 1959. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177706269.

Pyke:1960:CID

[Pyk60] Ronald Pyke. On centering infinitely divisible processes. Annals of Mathematical Statistics, 31(3):797–800, September 1960. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177705809.

Pyke:1961:MRPa

[Pyk61a] Ronald Pyke. Markov renewal processes: Definitions and preliminary properties. Annals of Mathematical Statistics, 32(4):1231–1242, December 1961. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177704863.

Pyke:1961:MRPb

[Pyk61b] Ronald Pyke. Markov renewal processes with finitely many states. Annals of Mathematical Statistics, 32(4):1243–1259, December 1961. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177704864.

Pyke:1963:RBD

[Pyk63] Ronald Pyke. Review: E. B. Dynkin, D. E. Brown, Theory of Markov Processes; E. B. Dynkin, Joseph Wloka, Die Grundlagen der Theorie der Markoffschen Prozesse. Annals of Mathematical Statistics, 34 (4):1604–1609, December 1963. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177703898.

Quesenberry:1968:NDU

[QG68] C. P. Quesenberry and M. P. Gessaman. Nonparametric discrimination using tolerance regions. Annals of Mathematical Statistics, 39 (2):664–673, April 1968. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177698425.

Quade:1965:APO

[Qua65] Dana Quade. On the asymptotic power of the one-sample Kolmogorov–Smirnov tests. Annals of Mathematical Statistics, 36(3): 1000–1018, June 1965. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177700072.

Quade:1966:AVS

[Qua66] Dana Quade. On analysis of variance for the K-sample problem. Annals of Mathematical Statistics, 37(6):1747–1758, December 1966. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177699163.

Qualls:1968:LDH

[Qua68] Clifford Qualls. On a limit distribution of high level crossings of a stationary Gaussian process. Annals of Mathematical Statistics, 39 (6):2108–2113, December 1968. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177698039.

Quensel:1936:MDR

[Que36] Carl-Erik Quensel. A method of determining the regression curve when the marginal distribution is of the normal logarithmic type. Annals of Mathematical Statistics, 7(4):196–201, December 1936. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177732480.

Quenouille:1949:ATC

[Que49a] M. H. Quenouille. Approximate tests of correlation in time-series. Journal of the Royal Statistical Society. Series B (Methodological), 11(1):68–84, ???? 1949. CODEN JSTBAJ. ISSN 0035-9246. URL

http://www.jstor.org/stable/2983696. See extension in [Que56,

Tuk58a], and reviews [Mil74, Efr79].

Quenouille:1949:JDS

[Que49b] M. H. Quenouille. The joint distribution of serial correlation coefficients. Annals of Mathematical Statistics, 20(4):561–571, December 1949. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177729948.

Quenouille:1949:PPS

[Que49c] M. H. Quenouille. Problems in plane sampling. Annals of Mathematical Statistics, 20(3):355–375, September 1949. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177729989.

Quenouille:1956:NBE

[Que56] M. H. Quenouille. Notes on bias in estimation. Biometrika, 43 (3–4):353–360, December 1956. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/ stable/2332914. This extends earlier work [Que49a, Tuk58a] on what is now known as the Quenouille–Tukey jacknife technique. See reviews [Mil74, Efr79].

Qualls:1971:ABG

[QW71] Clifford Qualls and Hisao Watanabe. An asymptotic 0–1 behavior of Gaussian processes. Annals of Mathematical Statistics, 42(6):2029– 2035, December 1971. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177693070.

Qualls:1972:APG

[QW72] Clifford Qualls and Hisao Watanabe. Asymptotic properties of Gaussian processes. Annals of Mathematical Statistics, 43(2):580–596, April 1972. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177692638.

Radner:1958:MEL

[Rad58] R. Radner. Minimax estimation for linear regressions. Annals of Mathematical Statistics, 29(4):1244–1250, December 1958. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177706455.

Radner:1962:TDP

[Rad62] R. Radner. Team decision problems. Annals of Mathematical Statistics, 33(3):857–881, September 1962. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177704455.

Raghavarao:1959:SOW

[Rag59] Damaraju Raghavarao. Some optimum weighing designs. Annals of Mathematical Statistics, 30(2):295–303, June 1959. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177706253.

Raghavarao:1960:GGD

[Rag60a] Damaraju Raghavarao. A generalization of group divisible designs. Annals of Mathematical Statistics, 31(3):756–771, September 1960. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177705802.

Raghavarao:1960:BSC

[Rag60b] Damaraju Raghavarao. On the block structure of certain PBIB designs with two associate classes having triangular and L_{2} association schemes. Annals of Mathematical Statistics, 31(3):787–791, September 1960. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177705806.

Raghavarao:1960:SAW

[Rag60c] Damaraju Raghavarao. Some aspects of weighing designs. Annals of Mathematical Statistics, 31(4):878–884, December 1960. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177705664.

Raghavarao:1962:SUB

[Rag62] Damaraju Raghavarao. Symmetrical unequal block arrangements with two unequal block sizes. Annals of Mathematical Statistics, 33 (2):620–633, June 1962. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177704586.

Raghavarao:1964:SWD

[Rag64] Damaraju Raghavarao. Singular weighing designs. Annals of Mathematical Statistics, 35(2):673–680, June 1964. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177703564.

Raghavachari:1965:ENS

[Rag65a] M. Raghavachari. On the efficiency of the normal scores test relative to the F -test. Annals of Mathematical Statistics, 36(4):1306–1307, August 1965. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177700005.

Raghavachari:1965:TSS

[Rag65b] M. Raghavachari. The two-sample scale problem when locations are unknown. Annals of Mathematical Statistics, 36(4):1236–1242, August 1965. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177699994.

Raghavarao:1966:DPB

[Rag66] Damaraju Raghavarao. Duals of partially balanced incomplete block designs and some nonexistence theorems. Annals of Mathematical Statistics, 37(4):1048–1052, August 1966. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177699387.

Raghavachari:1970:TBR

[Rag70a] M. Raghavachari. On a theorem of Bahadur on the rate of convergence of test statistics. Annals of Mathematical Statistics, 41(5):1695– 1699, October 1970. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177696813.

Raghavan:1970:CCF

[Rag70b] T. E. S. Raghavan. Convex cones and finite optimals. Annals of Mathematical Statistics, 41(2):702–705, April 1970. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177697118.

Raifaizen:1972:POD

[Rai72] C. H. Raifaizen. Pairs of one dimensional random walk paths. Annals of Mathematical Statistics, 43(6):2095–2098, December 1972. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177690891.

Raj:1953:MRT

[Raj53] Des Raj. On Mill’s ratio for the Type III population. Annals of Mathematical Statistics, 24(2):309–312, June 1953. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177729040.

Raj:1964:DSP

[Raj64] Des Raj. On double sampling for PPS estimation. Annals of Mathematical Statistics, 35(2):900–902, June 1964. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177703598.

Raj:1965:STO

[Raj65] Des Raj. On sampling over two occasions with probability proportionate to size. Annals of Mathematical Statistics, 36(1):327–330, February 1965. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177700297.

Raktoe:1967:ACC

[Rak67] B. L. Raktoe. Application of cyclic collineations to the construction of balanced L-restrictional prime powered lattice designs. Annals of Mathematical Statistics, 38(4):1127–1141, August 1967. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177698781.

Raktoe:1969:CED

[Rak69] B. L. Raktoe. Combining elements from distinct finite fields in mixed factorials. Annals of Mathematical Statistics, 40(2):498–504, April 1969. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177697716.

Raktoe:1970:GCE

[Rak70] B. L. Raktoe. Generalized combining of elements from finite fields. Annals of Mathematical Statistics, 41(5):1763–1767, October 1970. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177696823.

Ramachandran:1956:SAV

[Ram56a] K. V. Ramachandran. On the simultaneous analysis of variance test. Annals of Mathematical Statistics, 27(2):521–528, June 1956. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177728275.

Ramachandran:1956:TTE

[Ram56b] K. V. Ramachandran. On the Tukey test for the equality of means and the Hartley test for the equality of variances. Annals of Mathematical Statistics, 27(3):825–831, September 1956. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177728189.

Ramachandran:1962:OTE

[Ram62] B. Ramachandran. On the order and the type of entire characteristic functions. Annals of Mathematical Statistics, 33(4):1238–1255, December 1962. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177704356.

Ramachandramurty:1966:SNE

[Ram66a] P. V. Ramachandramurty. On some nonparametric estimates for shift in the Behrens–Fisher situation. Annals of Mathematical Statistics, 37(3):593–610, June 1966. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177699453.

Ramachandramurty:1966:PEO

[Ram66b] P. V. Ramachandramurty. On the Pitman efficiency of one-sided Kolmogorov and Smirnov tests for normal alternatives. Annals of Mathematical Statistics, 37(4):940–944, August 1966. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177699374.

Ramberg:1972:SSS

[Ram72] John S. Ramberg. Selection sample size approximations. Annals of Mathematical Statistics, 43(6):1977–1980, December 1972. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177690869.

Randles:1970:SRS

[Ran70] Ronald H. Randles. Some robust selection procedures. Annals of Mathematical Statistics, 41(5):1640–1645, October 1970. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177696807.

Rao:1954:GTD

[Rao54] C. Radhakrishna Rao. A general theory of discrimination when the information about alternative population distributions is based on samples. Annals of Mathematical Statistics, 25(4):651–670, December 1954. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177728653.

Rao:1958:CND

[Rao58a] J. N. K. Rao. A characterization of the normal distribution. Annals of Mathematical Statistics, 29(3):914–919, September 1958. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177706551.

Rao:1958:NBD

[Rao58b] V. R. Rao. A note on balanced designs. Annals of Mathematical Statistics, 29(1):290–294, March 1958. CODEN AASTAD.

ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177706729.

Rao:1959:APC

[Rao59] J. N. K. Rao. Acknowledgment of priority: A characterization of the normal distribution. Annals of Mathematical Statistics, 30(2): 610, June 1959. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177706279.

Rao:1960:DBI

[Rao60] P. V. Rao. The dual of a balanced incomplete block design. Annals of Mathematical Statistics, 31(3):779–785, September 1960. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177705804.

Rao:1961:CLD

[Rao61a] M. M. Rao. Consistency and limit distributions of estimators of parameters in explosive stochastic difference equations. Annals of Mathematical Statistics, 32(1):195–218, March 1961. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177705151.

Rao:1961:ACP

[Rao61b] P. V. Rao. Analysis of a class of PBIB designs with more than two associate classes. Annals of Mathematical Statistics, 32(3):800–808, September 1961. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177704974.

Rao:1962:RBW

[Rao62] R. Ranga Rao. Relations between weak and uniform convergence of measures with applications. Annals of Mathematical Statistics, 33 (2):659–680, June 1962. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177704588.

Rao:1965:NIB

[Rao65] M. Bhaskar Rao. A note on incomplete block designs with b = v. Annals of Mathematical Statistics, 36(6):1877, December 1965. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177699822.

Rao:1966:WDW

[Rao66] M. Bhaskar Rao. Weighing designs when n is odd. Annals of Mathematical Statistics, 37(5):1371–1381, October 1966. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177699281.

Rao:1967:CPS

[Rao67] T. Subba Rao. On the cross periodogram of a stationary Gaussian vector process. Annals of Mathematical Statistics, 38(2):593– 597, April 1967. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177698976.

Rao:1968:ELC

[Rao68a] B. L. S. Prakasa Rao. Estimation of the location of the cusp of a continuous density. Annals of Mathematical Statistics, 39(1):76–87, February 1968. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177698506.

Rao:1968:CSS

[Rao68b] B. L. S. Prakasa Rao. On a characterization of symmetric stable processes with finite mean. Annals of Mathematical Statistics, 39 (5):1498–1501, October 1968. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177698130.

Rao:1968:CSA

[Rao68c] T. Subba Rao. Cross spectral analysis of a Gaussian vector process in the presence of variance fluctuations. Annals of Mathematical Statistics, 39(5):1507–1512, October 1968. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177698132.

Rao:1969:DTV

[Rao69] C. Radhakrishna Rao. A decomposition theorem for vector variables with a linear structure. Annals of Mathematical Statistics, 40 (5):1845–1849, October 1969. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177697400.

Rao:1970:EDM

[Rao70a] B. L. S. Prakasa Rao. Estimation for distributions with monotone failure rate. Annals of Mathematical Statistics, 41(2):507–519, April 1970. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177697091.

Rao:1970:CWP

[Rao70b] B. L. S. Prakasa Rao. On a characterization of the Wiener process by constant regression. Annals of Mathematical Statistics, 41(1):321– 325, February 1970. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177697213.

Rao:1972:BES

[Rao72] J. S. Rao. Bahadur efficiencies of some tests for uniformity on the circle. Annals of Mathematical Statistics, 43(2):468–479, April 1972. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177692627.

Rasch:1948:FEW

[Ras48] G. Rasch. A functional equation for Wishart’s distribution. Annals of Mathematical Statistics, 19(2):262–266, June 1948. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177730251.

Ray:1957:PSP

[Ray57] W. D. Ray. A proof that the Sequential Probability Ratio Test (S.P.R.T.) of the general linear hypothesis terminates with probability unity. Annals of Mathematical Statistics, 28(2):521–523, June 1957. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177706986.

Ray:1965:BMS

[Ray65] S. N. Ray. Bounds on the maximum sample size of a Bayes sequential procedure. Annals of Mathematical Statistics, 36(3):859– 878, June 1965. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177700059.

Roy:1953:SCI

[RB53] S. N. Roy and R. C. Bose. Simultaneous confidence interval estimation. Annals of Mathematical Statistics, 24(4):513–536, December 1953. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177728912.

Roy:1958:TMI

[RB58] S. N. Roy and R. E. Bargmann. Tests of multiple independence and the associated confidence bounds. Annals of Mathematical Statistics, 29(2):491–503, June 1958. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177706624.

Roy:1960:EMF

[RC60a] J. Roy and I. M. Chakravarti. Estimating the mean of a finite population. Annals of Mathematical Statistics, 31(2):392–398, June 1960. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177705901.

Roy:1960:MMVa

[RC60b] S. N. Roy and Whitfield Cobb. Mixed model variance analysis with normal error and possibly non-normal other random effects: Part I: The univariate case. Annals of Mathematical Statistics, 31(4):939– 957, December 1960. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177705668.

Roy:1960:MMVb

[RC60c] S. N. Roy and Whitfield Cobb. Mixed model variance analysis with normal error and possibly non-normal other random effects: Part II: The multivariate case. Annals of Mathematical Statistics, 31(4):958– 968, December 1960. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177705669.

Ray-Chaudhuri:1962:AGQ

[RC62] D. K. Ray-Chaudhuri. Application of the geometry of quadrics for constructing PBIB designs. Annals of Mathematical Statistics, 33(3): 1175–1186, September 1962. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177704480.

Rajput:1972:GPG

[RC72] Balram S. Rajput and Stamatis Cambanis. Gaussian processes and Gaussian measures. Annals of Mathematical Statistics, 43(6):1944– 1952, December 1972. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177690865.

Read:1972:AIS

[Rea72] R. R. Read. The asymptotic inadmissibility of the sample distribution function. Annals of Mathematical Statistics, 43(1):89–95, February 1972. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177692704.

Redheffer:1951:NSI

[Red51] R. M. Redheffer. A note on the surprise index. Annals of Mathematical Statistics, 22(1):128–130, March 1951. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177729704.

Reiersol:1940:MRC

[Rei40] Olav Reiersôl. A method for recurrent computation of all the principal minors of a determinant, and its application in confluence analysis. Annals of Mathematical Statistics, 11(2):193–198, June 1940. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177731913.

Reich:1949:CCI

[Rei49] Edgar Reich. On the convergence of the classical iterative method of solving linear simultaneous equations. Annals of Mathematical Statistics, 20(3):448–451, September 1949. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177729998.

Reich:1957:WTW

[Rei57] Edgar Reich. Waiting times when queues are in tandem. Annals of Mathematical Statistics, 28(3):768–773, September 1957. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177706889.

Reich:1958:IET

[Rei58] Edgar Reich. On the integrodifferential equation of Takács. I. Annals of Mathematical Statistics, 29(2):563–570, June 1958. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177706632.

Reich:1959:IET

[Rei59] Edgar Reich. On the integrodifferential equation of Takács. II. Annals of Mathematical Statistics, 30(1):143–148, March 1959. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177706366.

Reinhardt:1961:ULF

[Rei61] H. E. Reinhardt. The use of least favorable distributions in testing composite hypotheses. Annals of Mathematical Statistics, 32(4):1034– 1041, December 1961. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177704843.

Reich:1963:NQT

[Rei63a] Edgar Reich. Note on queues in tandem. Annals of Mathematical Statistics, 34(1):338–341, March 1963. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177704275.

Reitsma:1963:ASD

[Rei63b] A. Reitsma. On approximations to sampling distributions of the mean for samples from non-normal populations. Annals of Mathematical Statistics, 34(4):1308–1314, December 1963. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177703866.

Reich:1964:RVB

[Rei64] Edgar Reich. Review: Vaclav E. Benes, General Stochastic Processes in the Theory of Queues. Annals of Mathematical Statistics, 35(2): 913–914, June 1964. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177703602.

Resnikoff:1962:TFC

[Res62] George J. Resnikoff. Tables to facilitate the computation of percentage points of the non-central t-distribution. Annals of Mathematical Statistics, 33(2):580–586, June 1962. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177704582.

Raktoe:1970:COS

[RF70a] B. L. Raktoe and W. T. Federer. Characterization of optimal Saturated main effect plans of the 2^{n} factorial. Annals of Mathematical Statistics, 41(1):203–206, February 1970. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177697201.

Raktoe:1970:TSP

[RF70b] B. L. Raktoe and W. T. Federer. A theorem on Saturated plans and their complements. Annals of Mathematical Statistics, 41(6):2184– 2185, December 1970. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177696728.

Raktoe:1971:NGS

[RF71] B. L. Raktoe and W. T. Federer. On the number of generators of Saturated main effect fractional replicates. Annals of Mathematical Statistics, 42(5):1758–1760, October 1971. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177693177.

Roy:1959:SCAa

[RG59a] S. N. Roy and R. Gnanadesikan. Some contributions to anova in one or more dimensions: I. Annals of Mathematical Statistics, 30(2): 304–317, June 1959. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177706254.

Roy:1959:SCAb

[RG59b] S. N. Roy and R. Gnanadesikan. Some contributions to anova in one or more dimensions: II. Annals of Mathematical Statistics, 30 (2):318–340, June 1959. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177706255.

Roy:1962:TSC

[RG62] S. N. Roy and R. Gnanadesikan. Two-sample comparisons of dispersion matrices for alternatives of intermediate specificity. Annals of Mathematical Statistics, 33(2):432–437, June 1962. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177704569.

Randles:1971:MSP

[RH71] Ronald H. Randles and Myles Hollander. Γ-minimax selection procedures in treatments versus control problems. Annals of Mathematical Statistics, 42(1):330–341, February 1971. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177693516.

Richards:1948:PCT

[Ric48] Paul I. Richards. Probability of coincidence for two periodically recurring events. Annals of Mathematical Statistics, 19(1):16–29, March 1948. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177730286.

Richter:1960:TSE

[Ric60] Donald Richter. Two-stage experiments for estimating a common mean. Annals of Mathematical Statistics, 31(4):1164–1173, December 1960. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177705687.

Rich:1966:SBA

[Ric66] R. N. Rich. A statistical basis for approximation and optimization. Annals of Mathematical Statistics, 37(1):59–65, February 1966. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177699598.

Rider:1931:SSC

[Rid31] Paul R. Rider. On small samples from certain non-normal universes. Annals of Mathematical Statistics, 2(1):48–65, February 1931. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177733042.

Rider:1940:RAM

[Rid40] Paul R. Rider. Report of the annual meeting of the Institute. Annals of Mathematical Statistics, 11(1):108–109, March 1940. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177731950.

Rider:1941:RCM

[Rid41] P. R. Rider. Report of the Chicago meeting of the Institute. Annals of Mathematical Statistics, 12(1):120–122, March 1941. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177731794.

Rider:1959:CNC

[Rid59a] Paul R. Rider. Correction notes: Corrections to “Quasi-Ranges in Samples from an Exponential Population”. Annals of Mathematical Statistics, 30(4):1266–1267, December 1959. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177706118. See [Rid59b].

Rider:1959:QRS

[Rid59b] Paul R. Rider. Quasi-ranges of samples from an exponential population. Annals of Mathematical Statistics, 30(1):252–254, March 1959. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177706385. See

correction [Rid59a].

Rider:1961:MMA

[Rid61] Paul R. Rider. The method of moments applied to a mixture of two exponential distributions. Annals of Mathematical Statistics, 32(1): 143–147, March 1961. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177705147.

Rietz:1931:CPF

[Rie31] H. L. Rietz. On certain properties of frequency distributions obtained by a linear fractional transformation of the variates of a given distribution. Annals of Mathematical Statistics, 2(1):38–47, February 1931. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177733041.

Rietz:1936:FDC

[Rie36] H. L. Rietz. On the frequency distribution of certain ratios. Annals of Mathematical Statistics, 7(3):145–153, September 1936. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177732505.

Rietz:1939:DSR

[Rie39] H. L. Rietz. On the distribution of the “student” ratio for small samples from certain non-normal populations. Annals of Mathematical Statistics, 10(3):265–274, September 1939. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177732185.

Riordan:1937:MRR

[Rio37] John Riordan. Moment recurrence relations for binomial, Poisson and hypergeometric frequency distributions. Annals of Mathematical Statistics, 8(2):103–111, June 1937. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177732430.

Riordan:1949:IFN

[Rio49] John Riordan. Inversion formulas in normal variable mapping. Annals of Mathematical Statistics, 20(3):417–425, September 1949. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177729993.

Riordan:1962:ELG

[Rio62a] John Riordan. Enumeration of linear graphs for mappings of finite sets. Annals of Mathematical Statistics, 33(1):178–185, March 1962. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177704722.

Riordan:1962:RFD

[Rio62b] John Riordan. Review: F. N. David, D. E. Barton, Combinatorial Chance. Annals of Mathematical Statistics, 33(4):1477–1479, December 1962. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177704385.

Riordan:1964:EER

[Rio64] John Riordan. The enumeration of election returns by number of lead positions. Annals of Mathematical Statistics, 35(1):369–379, March 1964. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177703760.

Riordan:1965:RPR

[Rio65] John Riordan. A recurrence for permutations without rising or falling successions. Annals of Mathematical Statistics, 36(2):708– 710, April 1965. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177700181.

Roy:1956:HNI

[RK56] S. N. Roy and Marvin A. Kastenbaum. On the hypothesis of no “interaction” in a multi-way contingency table. Annals of Mathematical Statistics, 27(3):749–757, September 1956. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177728181.

Ramachandran:1957:DPB

[RK57] K. V. Ramachandran and C. G. Khatri. On a decision procedure based on the Tukey statistic. Annals of Mathematical Statistics, 28 (3):802–806, September 1957. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177706897.

Raj:1958:SRS

[RK58] Des Raj and Salem H. Khamis. Some remarks on sampling with replacement. Annals of Mathematical Statistics, 29(2):550–557, June 1958. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177706630.

Roy:1957:PBL

[RL57] J. Roy and R. G. Laha. On partially balanced linked block designs. Annals of Mathematical Statistics, 28(2):488–493, June 1957. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177706977.

Robbins:1951:SAM

[RM51] Herbert Robbins and Sutton Monro. A stochastic approximation method. Annals of Mathematical Statistics, 22(3):400–407, September 1951. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/

1177729586. See note [Kal54].

Roy:1961:MCP

[RM61] S. N. Roy and W. F. Mikhail. On the monotonic character of the power functions of two multivariate tests. Annals of Mathematical Statistics, 32(4):1145–1151, December 1961. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177704852.

Ruiz-Moncayo:1968:OSF

[RM68] Alberto Ruiz-Moncayo. Optimal stopping for functions of Markov chains. Annals of Mathematical Statistics, 39(6):1905–1912, December 1968. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177698020.

Ruiz-Moncayo:1971:NFP

[RM71] A. Ruiz-Moncayo. A note on first passages for S_{n}/n^{\frac{1}{2}} . Annals of Mathematical Statistics, 42(1):398–400, February 1971. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177693532.

Rosenkrantz:1972:ASR

[RO72] Walter Rosenkrantz and Neville E. O’Reilly. Application of the Skorokhod representation theorem to rates of convergence for linear combinations of order statistics. Annals of Mathematical Statistics, 43(4): 1204–1212, August 1972. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177692472.

Robe:1930:MLR

[Rob30] Richard A. Robe. Modification of the link relative and interpolation methods of determining seasonal variation. Annals of Mathematical Statistics, 1(4):352–363, November 1930. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177733065.

Robinson:1933:ERT

[Rob33] Selby Robinson. An experiment regarding the χ^{2} test. Annals of Mathematical Statistics, 4(4):285–287, November 1933. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177732757.

Roberts:1935:SPI

[Rob35] John L. Roberts. Some practical interpolation formulas. Annals of Mathematical Statistics, 6(3):133–142, September 1935. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177732589. See errata

[Ano35c].

Roberts:1936:EPC

[Rob36] John L. Roberts. The elimination of perpetual calendars. Annals of Mathematical Statistics, 7(1):44–46, March 1936. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177732544.

Roberts:1937:ATO

[Rob37a] John L. Roberts. Applications of two osculatory formulas. Annals of Mathematical Statistics, 8(1):1–4, March 1937. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177732448.

Roberts:1937:CCB

[Rob37b] John L. Roberts. A coefficient of correlation between scholarship and salaries. Annals of Mathematical Statistics, 8(1):66–68, March 1937. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177732452.

Robbins:1944:EVT

[Rob44a] H. E. Robbins. On the expected values of two statistics. Annals of Mathematical Statistics, 15(3):321–323, September 1944. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177731238.

Robbins:1944:MRS

[Rob44b] H. E. Robbins. On the measure of a random set. Annals of Mathematical Statistics, 15(1):70–74, March 1944. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177731315.

Robbins:1944:DFT

[Rob44c] Herbert Robbins. On distribution-free tolerance limits in random sampling. Annals of Mathematical Statistics, 15(2):214–216, June 1944. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177731286.

Robbins:1945:MRS

[Rob45] H. E. Robbins. On the measure of a random set. II. Annals of Mathematical Statistics, 16(4):342–347, December 1945. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177731060.

Robbins:1947:APM

[Rob47] H. E. Robbins. Acknowledgment of priority: “On the Measure of a Random Set, [Parts I and II]”. Annals of Mathematical Statistics, 18(2):297, June 1947. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177730451.

Robbins:1948:CD

[Rob48a] Herbert Robbins. Convergence of distributions. Annals of Mathematical Statistics, 19(1):72–76, March 1948. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177730292.

Robbins:1948:DDQ

[Rob48b] Herbert Robbins. The distribution of a definite quadratic form. Annals of Mathematical Statistics, 19(2):266–270, June 1948. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177730252.

Robbins:1948:DSW

[Rob48c] Herbert Robbins. The distribution of Student’s t when the population means are unequal. Annals of Mathematical Statistics, 19(3):406– 410, September 1948. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177730206.

Robbins:1948:MD

[Rob48d] Herbert Robbins. Mixture of distributions. Annals of Mathematical Statistics, 19(3):360–369, September 1948. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177730200.

Robbins:1950:RCH

[Rob50] Herbert Robbins. Report of the Chapel Hill meeting of the Institute. Annals of Mathematical Statistics, 21(2):317–319, June 1950. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177729853.

Robbins:1954:RJD

[Rob54] Herbert Robbins. A remark on the joint distribution of cumulative sums. Annals of Mathematical Statistics, 25(3):614–616, September 1954. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177728733.

Robson:1958:AMI

[Rob58] D. S. Robson. Admissible and minimax integer-valued estimators of an integer-valued parameter. Annals of Mathematical Statistics, 29 (3):801–812, September 1958. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177706537.

Robbins:1961:RGP

[Rob61] Herbert Robbins. Recurrent games and the Petersburg Paradox. Annals of Mathematical Statistics, 32(1):187–194, March 1961. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177705150.

Roberts:1963:AOS

[Rob63] Charles DeWitt Roberts. An asymptotically optimal sequential design for comparing several experimental categories with a control. Annals of Mathematical Statistics, 34(4):1486–1493, December 1963. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177703880.

Robbins:1964:EBA

[Rob64a] Herbert Robbins. The empirical Bayes approach to statistical decision problems. Annals of Mathematical Statistics, 35(1):1–20, March 1964. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177703729.

Roberts:1964:AOF

[Rob64b] Charles DeWitt Roberts. An asymptotically optimal fixed sample size procedure for comparing several experimental categories with a control. Annals of Mathematical Statistics, 35(4):1571–1575, December 1964. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177700381.

Robertson:1965:PNL

[Rob65a] James B. Robertson. On a problem in non-linear prediction theory. Annals of Mathematical Statistics, 36(5):1554–1555, October 1965. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177699914.

Robertson:1965:NRC

[Rob65b] Tim Robertson. A note on the reciprocal of the conditional expectation of a positive random variable. Annals of Mathematical Statistics, 36(4):1302–1305, August 1965. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177700004.

Robertson:1966:RCE

[Rob66] Tim Robertson. A representation for conditional expectations given σ-lattices. Annals of Mathematical Statistics, 37(5):1279–1283, October 1966. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177699271.

Robertson:1967:EDW

[Rob67] Tim Robertson. On estimating a density which is measurable with respect to a σ-lattice. Annals of Mathematical Statistics, 38(2):482– 493, April 1967. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177698964.

Robbins:1968:ETP

[Rob68] Herbert E. Robbins. Estimating the total probability of the unobserved outcomes of an experiment. Annals of Mathematical Statistics, 39(1):256–257, February 1968. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177698526.

Robbins:1970:SMR

[Rob70a] Herbert Robbins. Statistical methods related to the law of the iterated logarithm. Annals of Mathematical Statistics, 41(5):1397–1409, October 1970. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177696786.

Robinson:1970:RAI

[Rob70b] J. Robinson. On relationship algebras of incomplete block designs. Annals of Mathematical Statistics, 41(2):648–650, April 1970. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177697106. See correc-

tion [Rob71b].

Robertson:1971:ACN

[Rob71a] James B. Robertson. An algorithm for computing the non-linear predictor. Annals of Mathematical Statistics, 42(1):75–80, February 1971. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177693496.

Robinson:1971:CNC

[Rob71b] J. Robinson. Correction note: Correction to “On Relationship Algebras of Incomplete Block Designs”. Annals of Mathematical Statistics, 42(2):842, April 1971. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177693446. See [Rob70b].

Robinson:1972:CCL

[Rob72a] J. Robinson. A converse to a combinatorial limit theorem. Annals of Mathematical Statistics, 43(6):2053–2057, December 1972. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177690884.

Robinson:1972:RMI

[Rob72b] J. Robinson. The randomization model for incomplete block designs. Annals of Mathematical Statistics, 43(2):480–489, April 1972. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177692628.

Rodemich:1966:SEU

[Rod66a] Eugene R. Rodemich. Spectral estimates using nonlinear functions. Annals of Mathematical Statistics, 37(5):1237–1256, October 1966. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177699268.

Rodine:1966:PPM

[Rod66b] Robert H. Rodine. Perfect probability measures and regular conditional probabilities. Annals of Mathematical Statistics, 37(5):1273– 1278, October 1966. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177699270.

Rodriguez:1971:POB

[Rod71] Dennis M. Rodriguez. Processes obtainable from Brownian motion by means of a random time change. Annals of Mathematical Statistics, 42(1):115–129, February 1971. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177693499.

Rogers:1959:NSI

[Rog59] Gerald S. Rogers. A note on the stochastic independence of functions of order statistics. Annals of Mathematical Statistics, 30(4):1263– 1264, December 1959. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177706113.

Rogers:1964:AGG

[Rog64] Gerald S. Rogers. An application of a generalized gamma distribution. Annals of Mathematical Statistics, 35(3):1368–1370, September 1964. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177703299.

Rolph:1968:BEM

[Rol68] John E. Rolph. Bayesian estimation of mixing distributions. Annals of Mathematical Statistics, 39(4):1289–1302, August 1968. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177698254.

Root:1969:ECS

[Roo69] D. H. Root. The existence of certain stopping times on Brownian motion. Annals of Mathematical Statistics, 40(2):715–718, April 1969. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177697749.

Root:1971:CRT

[Roo71] David Root. A counterexample in renewal theory. Annals of Mathematical Statistics, 42(5):1763–1766, October 1971. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177693179.

Root:1972:DAF

[Roo72] D. H. Root. Domains of attraction of first passage times. Annals of Mathematical Statistics, 43(2):681–682, April 1972. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177692654.

Rosenblatt:1952:LTA

[Ros52a] M. Rosenblatt. Limit theorems associated with variants of the von Mises statistic. Annals of Mathematical Statistics, 23(4):617–623, December 1952. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177729341.

Rosenblatt:1952:RMT

[Ros52b] Murray Rosenblatt. Remarks on a multivariate transformation. Annals of Mathematical Statistics, 23(3):470–472, September 1952. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177729394.

Rosenbaum:1953:TNT

[Ros53] S. Rosenbaum. Tables for a nonparametric test of dispersion. Annals of Mathematical Statistics, 24(4):663–668, December 1953. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177728922.

Rosenbaum:1954:TNT

[Ros54] S. Rosenbaum. Tables for a nonparametric test of location. Annals of Mathematical Statistics, 25(1):146–150, March 1954. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177728854.

Rosenblatt:1956:ERC

[Ros56a] Murray Rosenblatt. On the estimation of regression coefficients of a vector-valued time series with a stationary residual. Annals of Mathematical Statistics, 27(1):99–121, March 1956. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177728352.

Rosenblatt:1956:RSN

[Ros56b] Murray Rosenblatt. Remarks on some nonparametric estimates of a density function. Annals of Mathematical Statistics, 27(3):832– 837, September 1956. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177728190.

Rosenblatt:1962:NMG

[Ros62a] Judah Rosenblatt. Note on multivariate goodness-of-fit tests. Annals of Mathematical Statistics, 33(2):807–810, June 1962. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177704601.

Rosenblatt:1962:SMK

[Ros62b] Judah Rosenblatt. Some modified Kolmogorov–Smirnov tests of approximate hypotheses and their properties. Annals of Mathematical Statistics, 33(2):513–524, June 1962. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177704577.

Rosenblatt:1962:TAH

[Ros62c] Judah Rosenblatt. Testing approximate hypotheses in the composite case. Annals of Mathematical Statistics, 33(4):1356–1364, December 1962. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/

1177704368. See correction [Ros63a].

Rosenblatt:1963:CNC

[Ros63a] Judah Rosenblatt. Correction notes: Correction to “Testing Approximate Hypotheses in the Composite Case”. Annals of Mathematical Statistics, 34(2):673, June 1963. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177704182. See erratum [Ros62c].

Rosenblatt:1963:TCI

[Ros63b] Judah Rosenblatt. Tests and confidence intervals based on the metric d_{2}. Annals of Mathematical Statistics, 34(2):618–623, June 1963. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177704174.

Rosenblatt:1966:RRB

[Ros66a] M. Rosenblatt. Review: Richard Bellman, Stochastic Processes in Mathematical Physics and Engineering. Annals of Mathematical Statistics, 37(1):298–300, February 1966. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177699626.

Rosenkrantz:1966:LLT

[Ros66b] Walter A. Rosenkrantz. A local limit theorem for a certain class of random walks. Annals of Mathematical Statistics, 37(4):855–859, August 1966. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177699366.

Rosen:1967:IH

[Ros67] Bengt Rosén. On an inequality of Hoeffding. Annals of Mathematical Statistics, 38(2):382–392, April 1967. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177698953.

Ross:1968:ASM

[Ros68a] Sheldon M. Ross. Arbitrary state Markovian decision processes. Annals of Mathematical Statistics, 39(6):2118–2122, December 1968.

CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177698041.

Ross:1968:NDD

[Ros68b] Sheldon M. Ross. Non-discounted denumerable Markovian decision models. Annals of Mathematical Statistics, 39(2):412–423, April 1968. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177698405.

Rosen:1970:BCM

[Ros70a] Bengt Rosén. On bounds on the central moments of even order of a sum of independent random variables. Annals of Mathematical Statistics, 41(3):1074–1077, June 1970. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177696985.

Rosen:1970:CCW

[Ros70b] Bengt Rosén. On the coupon Collector’s waiting time. Annals of Mathematical Statistics, 41(6):1952–1969, December 1970. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177696696.

Rosenblatt:1971:CE

[Ros71a] Murray Rosenblatt. Curve estimates. Annals of Mathematical Statistics, 42(6):1815–1842, December 1971. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177693050.

Ross:1971:ILA

[Ros71b] Sheldon M. Ross. Infinitesimal look-ahead stopping rules. Annals of Mathematical Statistics, 42(1):297–303, February 1971. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177693512.

Ross:1971:NOR

[Ros71c] Sheldon M. Ross. On the nonexistence of ε-optimal randomized stationary policies in average cost Markov decision models. Annals of Mathematical Statistics, 42(5):1767–1768, October 1971. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177693180.

Rosen:1972:ATSa

[Ros72a] Bengt Rosen. Asymptotic theory for successive sampling with varying probabilities without replacement, I. Annals of Mathematical Statistics, 43(2):373–397, April 1972. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177692620.

Rosen:1972:ATSb

[Ros72b] Bengt Rosen. Asymptotic theory for successive sampling with varying probabilities without replacement, II. Annals of Mathematical Statistics, 43(3):748–776, June 1972. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177692543.

Rothman:1971:TCI

[Rot71] Edward D. Rothman. Tests of coordinate independence for a bivariate sample on a torus. Annals of Mathematical Statistics, 42(6):1962– 1969, December 1971. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177693064.

Roussas:1965:AIM

[Rou65] G. G. Roussas. Asymptotic inference in Markov processes. Annals of Mathematical Statistics, 36(3):978–992, June 1965. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177700070.

Roussas:1969:NET

[Rou69] George G. Roussas. Nonparametric estimation of the transition distribution function of a Markov process. Annals of Mathematical Statistics, 40(4):1386–1400, August 1969. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177697510.

Royer:1933:SMC

[Roy33] Elmer B. Royer. A simple method for calculating mean square contingency. Annals of Mathematical Statistics, 4(1):75–78, February 1933. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177732822.

Roy:1953:HMT

[Roy53a] S. N. Roy. On a heuristic method of test construction and its use in multivariate analysis. Annals of Mathematical Statistics, 24(2):

220–238, June 1953. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177729029.

Royden:1953:BDF

[Roy53b] H. L. Royden. Bounds on a distribution function when its first n moments are given. Annals of Mathematical Statistics, 24(3):361– 376, September 1953. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177728977.

Roy:1954:SFR

[Roy54] S. N. Roy. Some further results in simultaneous confidence interval estimation. Annals of Mathematical Statistics, 25(4):752–761, December 1954. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/

1177728661. See note [Roy56].

Roy:1956:NSF

[Roy56] S. N. Roy. A note on “Some Further Results in Simultaneous Confidence Interval Estimation”. Annals of Mathematical Statistics, 27 (3):856–858, September 1956. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177728197. See [Roy54].

Roy:1958:SPM

[Roy58] J. Roy. Step-down procedure in multivariate analysis. Annals of Mathematical Statistics, 29(4):1177–1187, December 1958. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177706449.

Royall:1970:FPS

[Roy70] Richard M. Royall. Finite population sampling—On labels in estimation. Annals of Mathematical Statistics, 41(5):1774–1779, October 1970. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177696825.

Robbins:1949:AMM

[RP49] Herbert Robbins and E. J. G. Pitman. Application of the method of mixtures to quadratic forms in normal variates. Annals of Mathematical Statistics, 20(4):552–560, December 1949. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177729947.

Root:1955:FSE

[RP55] W. L. Root and T. S. Pitcher. On the Fourier series expansion of random functions. Annals of Mathematical Statistics, 26(2):313– 318, June 1955. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177728548.

Roy:1958:CBV

[RP58] S. N. Roy and R. F. Potthoff. Confidence bounds on vector analogues of the “Ratio of Means” and the “Ratio of Variances” for two correlated normal variates and some associated tests. Annals of Mathematical Statistics, 29(3):829–841, September 1958. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177706539.

Ray:1963:CPO

[RP63] W. D. Ray and A. E. N. T. Pitman. Chebyshev polynomial and other new approximations to Mills’ ratio. Annals of Mathematical Statistics, 34(3):892–902, September 1963. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177704012.

Rumsey:1965:JDP

[RP65] Howard Rumsey, Jr. and Edward C. Posner. Joint distributions with prescribed moments. Annals of Mathematical Statistics, 36(1):286– 298, February 1965. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177700290.

Rayner:1967:NGI

[RP67] A. A. Rayner and R. M. Pringle. A note on generalized inverses in the linear hypothesis not of full rank. Annals of Mathematical Statistics, 38(1):271–273, February 1967. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177699081.

Roy:1959:NCP

[RR59] S. N. Roy and J. Roy. A note on a class of problems in ‘normal’ multivariate analysis of variance. Annals of Mathematical Statistics, 30 (2):577–581, June 1959. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177706270.

Raghavarao:1964:CD

[RR64a] D. Raghavarao and K. Chandrasekhara Rao. Cubic designs. Annals of Mathematical Statistics, 35(1):389–397, March 1964. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177703762.

Rosenblatt-Roth:1964:STC

[RR64b] M. Rosenblatt-Roth. Some theorems concerning the strong law of large numbers for non-homogeneous Markov chains. Annals of Mathematical Statistics, 35(2):566–576, June 1964. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177703556.

Rosenblatt-Roth:1965:QM

[RR65] M. Rosenblatt-Roth. Quantiles and medians. Annals of Mathematical Statistics, 36(3):921–925, June 1965. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177700064.

Rosenblatt-Roth:1966:SLT

[RR66] Millu Rosenblatt-Roth. Some limit theorems for non-homogeneous Markov chains. Annals of Mathematical Statistics, 37(5):1224–1236, October 1966. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177699267.

Rao:1972:CNC

[RR72a] K. P. S. Bhaskara Rao and M. Bhaskara Rao. Correction notes: Correction to “A Remark on Nonatomic Measures”. Annals of Mathematical Statistics, 43(4):1380, August 1972. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177692495. See [RR72b].

Rao:1972:RNM

[RR72b] K. P. S. Bhaskara Rao and M. Bhaskara Rao. A remark on nonatomic measures. Annals of Mathematical Statistics, 43(1):369– 370, February 1972. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.

aoms/1177692735. See correction [RR72a].

Rizvi:1967:NPS

[RS67] M. Haseeb Rizvi and Milton Sobel. Nonparametric procedures for selecting a subset containing the population with the largest α-quantile.

Annals of Mathematical Statistics, 38(6):1788–1803, December 1967. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177698612.

Robbins:1970:BCP

[RS70] Herbert Robbins and David Siegmund. Boundary crossing probabilities for the Wiener process and sample sums. Annals of Mathematical Statistics, 41(5):1410–1429, October 1970. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177696787.

Rodrigues:1972:RAR

[RS72a] Flavio Rodrigues and Gordon Simons. Review: Alfréd Rényi, Foundations of Probability. Annals of Mathematical Statistics, 43(5):1745– 1747, October 1972. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177692416.

Rolph:1972:CPS

[RS72b] John E. Rolph and Ralph E. Strauch. A countable policy set for sequential decision problems. Annals of Mathematical Statistics, 43 (6):2078–2082, December 1972. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177690888.

Rao:1960:CTR

[RSS60] U. V. R. Rao, I. R. Savage, and M. Sobel. Contributions to the theory of rank order statistics: The two-sample censored case. Annals of Mathematical Statistics, 31(2):415–426, June 1960. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177705904.

Robbins:1967:SAB

[RSS67] H. Robbins, G. Simons, and N. Starr. A sequential analogue of the Behrens–Fisher problem. Annals of Mathematical Statistics, 38 (5):1384–1391, October 1967. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177698694.

Robbins:1968:SPS

[RSS68] H. Robbins, M. Sobel, and N. Starr. A sequential procedure for selecting the largest of k means. Annals of Mathematical Statistics, 39(1):88–92, February 1968. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177698507.

Ruymgaart:1972:ANN

[RSvZ72] F. H. Ruymgaart, G. R. Shorack, and W. R. van Zwet. Asymptotic normality of nonparametric tests for independence. Annals of Mathematical Statistics, 43(4):1122–1135, August 1972. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177692465.

Rizvi:1968:NRP

[RSW68] M. Haseeb Rizvi, Milton Sobel, and George G. Woodworth. Nonparametric ranking procedures for comparison with a control. Annals of Mathematical Statistics, 39(6):2075–2093, December 1968. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177698035.

Rubin:1959:EPD

[RT59] Herman Rubin and Howard G. Tucker. Estimating the parameters of a differential process. Annals of Mathematical Statistics, 30(3):641– 658, September 1959. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177706195.

Rao:1969:RPE

[RT69] P. V. Rao and J. I. Thornby. A robust point estimator in a generalized regression model. Annals of Mathematical Statistics, 40(5):1784– 1790, October 1969. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177697391.

Rubin:1945:DSC

[Rub45] Herman Rubin. On the distribution of the serial correlation coefficient. Annals of Mathematical Statistics, 16(2):211–215, June 1945. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177731123.

Rubin:1956:UCR

[Rub56] Herman Rubin. Uniform convergence of random functions with applications to statistics. Annals of Mathematical Statistics, 27(1):200– 203, March 1956. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177728359.

Rubin:1958:ERL

[Rub58] Herman Rubin. Estimation of a regression line with both variables subject to error under an unusual identification condition. Annals of Mathematical Statistics, 29(2):606–608, June 1958. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177706643.

Ruben:1960:PCRa

[Rub60a] Harold Ruben. Probability content of regions under spherical normal distributions, I. Annals of Mathematical Statistics, 31(3):598–618, September 1960. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177705788.

Ruben:1960:PCRb

[Rub60b] Harold Ruben. Probability content of regions under spherical normal distributions, II: The distribution of the range in normal samples. Annals of Mathematical Statistics, 31(4):1113–1121, December 1960. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177705683. See

corrections [Rub61a].

Ruben:1961:CNC

[Rub61a] Harold Ruben. Correction notes: Corrections to “Probability Content of Regions Under Spherical Normal Distributions, II: The Distribution of the Range in Normal Samples”. Annals of Mathematical Statistics, 32(2):620, June 1961. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177705076. See [Rub60b].

Ruben:1961:PCR

[Rub61b] Harold Ruben. Probability content of regions under spherical normal distributions, III: The bivariate normal integral. Annals of Mathematical Statistics, 32(1):171–186, March 1961. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177705149.

Ruben:1962:PCR

[Rub62a] Harold Ruben. Probability content of regions under spherical normal distributions, IV: The distribution of homogeneous and nonhomogeneous quadratic functions of normal variables. Annals of Mathematical Statistics, 33(2):542–570, June 1962. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177704580.

Ruben:1962:SAE

[Rub62b] Harold Ruben. Some aspects of the emigration–immigration process. Annals of Mathematical Statistics, 33(1):119–129, March 1962. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177704717. See

[Rub63a].

Ruben:1963:EFI

[Rub63a] Harold Ruben. The estimation of a fundamental interaction parameter in an emigration–immigration process. Annals of Mathematical Statistics, 34(1):238–259, March 1963. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177704262. See [Rub62b].

Ruben:1963:NRD

[Rub63b] Harold Ruben. A new result on the distribution of quadratic forms. Annals of Mathematical Statistics, 34(4):1582–1584, December 1963. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177703890.

Ruben:1966:RKM

[Rub66] Harold Ruben. Review: Kenneth S. Miller, Multidimensional Gaussian Distributions. Annals of Mathematical Statistics, 37(1):301–307, February 1966. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177699627.

Rubin:1968:RDL

[Rub68] Herman Rubin. Review: D. N. Lawley, A. E. Maxwell, Factor Analysis as a Statistical Method. Annals of Mathematical Statistics, 39 (1):292, February 1968. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177698539.

Ruegg:1970:CCI

[Rue70] Alan Ruegg. A characterization of certain infinitely divisible laws. Annals of Mathematical Statistics, 41(4):1354–1356, August 1970. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177696912.

Ruegg:1971:NCI

[Rue71] Alan F. Ruegg. A necessary condition on the infinite divisibility of probability distributions. Annals of Mathematical Statistics, 42 (5):1681–1685, October 1971. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177693166.

Rustagi:1957:MMC

[Rus57] Jagdish Sharan Rustagi. On minimizing and maximizing a certain integral with statistical applications. Annals of Mathematical Statistics, 28(2):309–328, June 1957. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177706961.

Rutherford:1954:CD

[Rut54] R. S. G. Rutherford. On a contagious distribution. Annals of Mathematical Statistics, 25(4):703–713, December 1954. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177728656.

Rosenblatt:1965:EB

[RV65] M. Rosenblatt and J. W. Van Ness. Estimation of the bispectrum. Annals of Mathematical Statistics, 36(4):1120–1136, August 1965. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177699987.

Revesz:1972:AR

[RV72] P. Révész and I. Vincze. Alfréd Rényi, 1921–1970. Annals of Mathematical Statistics, 43(6):i–xvi, December 1972. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http:/

/projecteuclid.org/euclid.aoms/1177690849. With a bibliogra-

phy based on one compiled by P. Medgyessy.

Ross:1964:TSP

[RW64] John Ross and R. A. Weitzman. The Twenty-Seven Per Cent Rule. Annals of Mathematical Statistics, 35(1):214–221, March 1964. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177703745.

Robertson:1968:EMP

[RW68] Tim Robertson and Paul Waltman. On estimating monotone parameters. Annals of Mathematical Statistics, 39(3):1030–1039, June 1968.

CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177698335.

Rizvi:1970:SPB

[RW70] M. Haseeb Rizvi and George G. Woodworth. On selection procedures based on ranks: Counterexamples concerning least favorable configurations. Annals of Mathematical Statistics, 41(6):1942–1951, December 1970. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177696695.

Rothman:1972:CMT

[RW72] E. D. Rothman and Michael Woodroofe. A Cramér–von Mises type statistic for testing symmetry. Annals of Mathematical Statistics, 43 (6):2035–2038, December 1972. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177690879.

Solari:1957:MVM

[SA57] M. E. Solari and A. A. Anis. The mean and variance of the maximum of the adjusted partial sums of a finite number of independent normal variates. Annals of Mathematical Statistics, 28(3):706–716, September 1957. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177706882.

Saleh:1966:AOQ

[SA66] A. K. Md. Ehsanes Saleh and Mir M. Ali. Asymptotic optimum quantiles for the estimation of the parameters of the negative exponential distribution. Annals of Mathematical Statistics, 37(1):143– 151, February 1966. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177699605.

Srivastava:1970:SBP

[SA70] J. N. Srivastava and Donald A. Anderson. Some basic properties of multidimensional partially balanced designs. Annals of Mathematical Statistics, 41(5):1438–1445, October 1970. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177696789.

Srivastava:1971:FAS

[SA71] J. N. Srivastava and D. A. Anderson. Factorial association schemes with applications to the construction of multidimensional partially balanced designs. Annals of Mathematical Statistics, 42(4):1167– 1181, August 1971. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177693233.

Sacks:1958:ADS

[Sac58] Jerome Sacks. Asymptotic distribution of stochastic approximation procedures. Annals of Mathematical Statistics, 29(2):373–405, June 1958. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177706619.

Sacks:1960:ESQ

[Sac60] J. Sacks. Ergodocity [sic] of queues in series. Annals of Mathematical Statistics, 31(3):579–588, September 1960. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177705786.

Sacks:1963:GBS

[Sac63] Jerome Sacks. Generalized Bayes solutions in estimation problems. Annals of Mathematical Statistics, 34(3):751–768, September 1963. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177704001.

Sacks:1965:NST

[Sac65] J. Sacks. A note on the sequential t-test. Annals of Mathematical Statistics, 36(6):1867–1869, December 1965. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177699818.

Sacksteder:1967:NSE

[Sac67] Richard Sacksteder. A note on statistical equivalence. Annals of Mathematical Statistics, 38(3):787–794, June 1967. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177698872.

Sackrowitz:1970:EMP

[Sac70] Harold Sackrowitz. Estimation for monotone parameter sequences: The discrete case. Annals of Mathematical Statistics, 41(2):609– 620, April 1970. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177697101.

Sackrowitz:1972:NAT

[Sac72] Harold Sackrowitz. A note on approachability in a two-person game. Annals of Mathematical Statistics, 43(3):1017–1019, June 1972. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177692567.

Sade:1951:ONL

[Sad51] Albert Sade. An omission in Norton’s list of 7 × 7 squares. Annals of Mathematical Statistics, 22(2):306–307, June 1951. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177729654.

Sakrison:1964:CKW

[Sak64] David J. Sakrison. A continuous Kiefer–Wolfowitz procedure for random processes. Annals of Mathematical Statistics, 35(2):590– 599, June 1964. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/

1177703558. See correction [Sak66].

Sakrison:1966:CNC

[Sak66] David J. Sakrison. Correction notes: Correction to “A Continuous Kiefer–Wolfowitz Procedure for Random Processes”. Annals of Mathematical Statistics, 37(3):745, June 1966. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177699476. See [Sak64].

Salvosa:1930:TID

[Sal30a] L. R. Salvosa. Table III: Derivatives of the standardized Type III function. Annals of Mathematical Statistics, 1(3):127–187, August 1930. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177733088.

Salvosa:1930:TPT

[Sal30b] Luis R. Salvosa. Tables of Pearson’s Type III function. Annals of Mathematical Statistics, 1(2):191–198 + 1–125, May 1930. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177733130.

Saleh:1966:EPE

[Sal66] A. K. Md. Ehsanes Saleh. Estimation of the parameters of the exponential distribution based on optimum order statistics in censored samples. Annals of Mathematical Statistics, 37(6):1717–1735, December 1966. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177699161.

Salehi:1967:RWS

[Sal67] H. Salehi. Review: N. Wiener, A. Siegel, B. Franklin, W. T. Martin, Differential Space, Quantum Systems, and Prediction. Annals of Mathematical Statistics, 38(6):1929–1931, December 1967. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177698632.

Samuelson:1941:CRP

[Sam41] Paul A. Samuelson. Conditions that the roots of a polynomial be less than unity in absolute value. Annals of Mathematical Statistics, 12 (3):360–364, September 1941. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177731720.

Samuelson:1942:MDE

[Sam42] P. A. Samuelson. A method of determining explicitly the coefficients of the characteristic equation. Annals of Mathematical Statistics, 13 (4):424–429, December 1942. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177731540.

Samuelson:1943:FGG

[Sam43] Paul A. Samuelson. Fitting general Gram–Charlier series. Annals of Mathematical Statistics, 14(2):179–187, June 1943. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177731459.

Sampford:1953:SIM

[Sam53] M. R. Sampford. Some inequalities on Mill’s ratio and related functions. Annals of Mathematical Statistics, 24(1):130–132, March 1953. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177729093.

Samuel:1963:ASS

[Sam63a] Ester Samuel. Asymptotic solutions of the sequential compound decision problem. Annals of Mathematical Statistics, 34(3):1079–1094, September 1963. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177704032.

Samuel:1963:EBA

[Sam63b] Ester Samuel. An empirical Bayes approach to the testing of certain parametric hypotheses. Annals of Mathematical Statistics, 34 (4):1370–1385, December 1963. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177703870.

Samuel:1963:NSC

[Sam63c] Ester Samuel. Note on a sequential classification problem. Annals of Mathematical Statistics, 34(3):1095–1097, September 1963. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177704033.

Samuel:1964:CLC

[Sam64] Ester Samuel. Convergence of the losses of certain decision rules for the sequential compound decision problem. Annals of Mathematical Statistics, 35(4):1606–1621, December 1964. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177700385.

Samuel:1965:NEO

[Sam65a] Ester Samuel. Note on estimating ordered parameters. Annals of Mathematical Statistics, 36(2):698–702, April 1965. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177700178.

Samuel:1965:SCE

[Sam65b] Ester Samuel. Sequential compound estimators. Annals of Mathematical Statistics, 36(3):879–889, June 1965. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177700060.

Samuels:1965:NSI

[Sam65c] S. M. Samuels. On the number of successes in independent trials. Annals of Mathematical Statistics, 36(4):1272–1278, August 1965. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177699998.

Samuels:1966:CTI

[Sam66] S. M. Samuels. On a Chebyshev-type inequality for sums of independent random variables. Annals of Mathematical Statistics, 37(1): 248–259, February 1966. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177699614.

Samuel:1968:SML

[Sam68a] Ester Samuel. Sequential maximum likelihood estimation of the size of a population. Annals of Mathematical Statistics, 39(3):1057– 1068, June 1968. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177698338.

Samuels:1968:RRT

[Sam68b] S. M. Samuels. Randomized rules for the two-armed-bandit with finite memory. Annals of Mathematical Statistics, 39(6):2103–2107, December 1968. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177698038.

Samuels:1969:MIS

[Sam69] S. M. Samuels. The Markov inequality for sums of independent random variables. Annals of Mathematical Statistics, 40(6):1980–1984, December 1969. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177697279.

Santalo:1947:FTM

[San47] L. A. Santaló. On the first two moments of the measure of a random set. Annals of Mathematical Statistics, 18(1):37–49, March 1947. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177730490.

Sankaranarayanan:1961:LDT

[San61] G. Sankaranarayanan. Limit distributions in the theory of counters. Annals of Mathematical Statistics, 32(4):1271–1285, December 1961. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177704866. See

correction [San62].

Sankaranarayanan:1962:CNC

[San62] G. Sankaranarayanan. Correction notes: Correction to “Limit Distributions in the Theory of Counters”. Annals of Mathematical Statistics, 33(4):1466, December 1962. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177704381. See [San61].

Sandved:1968:ASE

[San68] Else Sandved. Ancillary statistics and estimation of the loss in estimation problems. Annals of Mathematical Statistics, 39(5):1756– 1758, October 1968. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177698162.

Sandved:1970:CCD

[San70] Else Sandved. A comment on the compound decision theory. Annals of Mathematical Statistics, 41(5):1760–1762, October 1970. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177696822.

Sanathanan:1972:ESM

[San72] Lalitha Sanathanan. Estimating the size of a multinomial population. Annals of Mathematical Statistics, 43(1):142–152, February 1972. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177692709.

Sard:1949:SAF

[Sar49] Arthur Sard. Smoothest approximation formulas. Annals of Mathematical Statistics, 20(4):612–615, December 1949. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177729956.

Sarhan:1954:EMS

[Sar54] A. E. Sarhan. Estimation of the mean and standard deviation by order statistics. Annals of Mathematical Statistics, 25(2):317–328, June 1954. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177728788.

Sarhan:1955:EMSa

[Sar55a] A. E. Sarhan. Estimation of the mean and standard deviation by order statistics. Part II. Annals of Mathematical Statistics, 26(3):505– 511, September 1955. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177728495.

Sarhan:1955:EMSb

[Sar55b] A. E. Sarhan. Estimation of the mean and standard deviation by order statistics. Part III. Annals of Mathematical Statistics, 26(4):576– 592, December 1955. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177728418.

Sarkadi:1957:DNE

[Sar57] K. Sarkadi. On the distribution of the number of exceedances. Annals of Mathematical Statistics, 28(4):1021–1023, December 1957. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177706804.

Sarkadi:1960:MDE

[Sar60] K. Sarkadi. On the median of the distribution of exceedances. Annals of Mathematical Statistics, 31(1):225–226, March 1960. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177706000.

Sarndal:1965:TRO

[Sär65] Carl-Erik Särndal. A theorem on rank orders for two censored samples. Annals of Mathematical Statistics, 36(1):316–321, February 1965. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177700295.

Satterthwaite:1941:CAC

[Sat41] Franklin E. Satterthwaite. A concise analysis of certain algebraic forms. Annals of Mathematical Statistics, 12(1):77–83, March 1941. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177731786.

Satterthwaite:1942:GPD

[Sat42a] F. E. Satterthwaite. Generalized Poisson distribution. Annals of Mathematical Statistics, 13(4):410–417, December 1942. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177731538.

Satterthwaite:1942:GAV

[Sat42b] Franklin E. Satterthwaite. A generalized analysis of variance. Annals of Mathematical Statistics, 13(1):34–41, March 1942. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177731640.

Satterthwaite:1942:LRC

[Sat42c] Franklin E. Satterthwaite. Linear restrictions on chi-square. Annals of Mathematical Statistics, 13(3):326–331, September 1942. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177731571.

Satterthwaite:1944:ECM

[Sat44] F. E. Satterthwaite. Error control in matrix calculation. Annals of Mathematical Statistics, 15(4):373–387, December 1944. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177731208.

Saunders:1960:STR

[Sau60] Sam C. Saunders. Sequential tolerance regions. Annals of Mathematical Statistics, 31(1):198–216, March 1960. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177705997.

Saunders:1963:SSC

[Sau63a] Sam C. Saunders. On the sample size and coverage for the Jiřina sequential procedure. Annals of Mathematical Statistics, 34(3):847– 856, September 1963. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177704007.

Saunders:1963:SAJ

[Sau63b] Sam C. Saunders. Some applications of the Jiřina sequential procedure to observations with trend. Annals of Mathematical Statistics, 34(3):857–865, September 1963. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177704008.

Savage:1947:UTU

[Sav47] L. J. Savage. A uniqueness theorem for unbiased sequential binomial estimation. Annals of Mathematical Statistics, 18(2):295– 297, June 1947. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177730450.

Savage:1956:CTR

[Sav56] I. Richard Savage. Contributions to the theory of rank order statistics-the two-sample case. Annals of Mathematical Statistics, 27(3):590–615, September 1956. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177728170.

Savage:1957:CTR

[Sav57a] I. Richard Savage. Contributions to the theory of rank order statistics — the “trend” case. Annals of Mathematical Statistics, 28(4):968– 977, December 1957. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177706797.

Savage:1957:WDP

[Sav57b] Leonard J. Savage. When different pairs of hypotheses have the same family of likelihood-ratio test regions. Annals of Mathematical Statistics, 28(4):1028–1032, December 1957. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177706806.

Savage:1959:CTR

[Sav59] I. Richard Savage. Contributions to the theory of rank order statistics — the one-sample case. Annals of Mathematical Statistics, 30(4): 1018–1023, December 1959. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177706084.

Savage:1960:CTR

[Sav60] I. Richard Savage. Contributions to the theory of rank order statistics: Computation rules for probabilities of rank orders. Annals of Mathematical Statistics, 31(2):519–520, June 1960. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177705920.

Savage:1971:RHD

[Sav71] Richard I. Savage. Review: H. A. David, Order Statistics. Annals of Mathematical Statistics, 42(3):1138, June 1971. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177693349.

Sawyer:1972:RCS

[Saw72] Stanley Sawyer. Rates of convergence for some functionals in probability. Annals of Mathematical Statistics, 43(1):273–284, February 1972. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177692720.

Saxena:1967:NC

[Sax67] Ashok K. Saxena. A note on classification. Annals of Mathematical Statistics, 38(5):1592–1593, October 1967. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177698716.

Schumann:1957:CSS

[SB57a] D. E. W. Schumann and R. A. Bradley. The comparison of the sensitivities of similar experiments: Theory. Annals of Mathematical Statistics, 28(4):902–920, December 1957. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177706792.

Szekeres:1957:BFF

[SB57b] G. Szekeres and F. E. Binet. On Borel fields over finite sets. Annals of Mathematical Statistics, 28(2):494–498, June 1957. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177706978.

Skibinsky:1963:ISS

[SC63] M. Skibinsky and L. Cote. On the inadmissibility of some standard estimates in the presence of prior information. Annals of Mathematical Statistics, 34(2):539–548, June 1963. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177704166.

Stout:1969:EVS

[SC69] William F. Stout and Y. S. Chow. On the expected value of a stopped stochastic sequence. Annals of Mathematical Statistics, 40(2):456– 461, April 1969. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177697710.

Srivastava:1971:CRI

[SC71a] J. N. Srivastava and D. V. Chopra. On the characteristic roots of the information matrix of 2^{m} balanced factorial designs of resolution v, with applications. Annals of Mathematical Statistics, 42(2):722– 734, April 1971. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177693421.

Strawderman:1971:AEM

[SC71b] William E. Strawderman and Arthur Cohen. Admissibility of estimators of the mean vector of a multivariate normal distribution with quadratic loss. Annals of Mathematical Statistics, 42(1):270– 296, February 1971. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177693511.

Sudbury:1972:DSR

[SC72] Aidan Sudbury and Peter Clifford. The division of a sequence of random variables to form two approximately equal sums. Annals of Mathematical Statistics, 43(1):236–241, February 1972. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177692716.

Scarf:1959:BSS

[Sca59] Herbert Scarf. Bayes solutions of the statistical inventory problem. Annals of Mathematical Statistics, 30(2):490–508, June 1959. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177706264.

Schmidt:1934:SAO

[Sch34] Robert Schmidt. Statistical analysis of one-dimensional distributions. Annals of Mathematical Statistics, 5(1):30–72, March 1934. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177732734.

Scheffe:1942:RVT

[Sch42a] Henry Scheffé. On the ratio of the variances of two normal populations. Annals of Mathematical Statistics, 13(4):371–388, December 1942. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177731535.

Scheffe:1942:TTC

[Sch42b] Henry Scheffé. On the theory of testing composite hypotheses with one constraint. Annals of Mathematical Statistics, 13(3):280–293, September 1942. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177731567.

Scheffe:1943:MPA

[Sch43a] Henry Scheffé. On a measure problem arising in the theory of nonparametric tests. Annals of Mathematical Statistics, 14(3):227–233, September 1943. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177731416.

Scheffe:1943:SBF

[Sch43b] Henry Scheffé. On solutions of the Behrens–Fisher problem, based on the t-distribution. Annals of Mathematical Statistics, 14(1):35– 44, March 1943. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177731490.

Scheffe:1943:SIN

[Sch43c] Henry Scheffé. Statistical inference in the non-parametric case. Annals of Mathematical Statistics, 14(4):305–332, December 1943. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177731355.

Scheffe:1944:NBF

[Sch44] Henry Scheffé. A note on the Behrens–Fisher problem. Annals of Mathematical Statistics, 15(4):430–432, December 1944. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177731214.

Scheffe:1947:UCT

[Sch47] Henry Scheffé. A useful convergence theorem for probability distributions. Annals of Mathematical Statistics, 18(3):434–438, September 1947. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177730390.

Scheffe:1956:AMA

[Sch56a] Henry Scheffé. Alternative models for the analysis of variance. Annals of Mathematical Statistics, 27(2):251–271, June 1956. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177728258.

Scheffe:1956:MMA

[Sch56b] Henry Scheffé. A “Mixed Model” for the analysis of variance. Annals of Mathematical Statistics, 27(1):23–36, March 1956. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177728348.

Schmid:1958:KSL

[Sch58] Paul Schmid. On the Kolmogorov and Smirnov limit theorems for discontinuous distribution functions. Annals of Mathematical Statistics, 29(4):1011–1027, December 1958. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http:/

/projecteuclid.org/euclid.aoms/1177706438;   http://www.
jstor.org/stable/2236943.

Schmetterer:1960:PJN

[Sch60a] L. Schmetterer. On a problem of J. Neyman and E. Scott. Annals of Mathematical Statistics, 31(3):656–661, September 1960. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177705792.

Schmetterer:1960:UE

[Sch60b] L. Schmetterer. On unbiased estimation. Annals of Mathematical Statistics, 31(4):1154–1163, December 1960. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177705686.

Schwarz:1960:CFD

[Sch60c] Gideon Schwarz. A class of factorial designs with unequal cellfrequencies. Annals of Mathematical Statistics, 31(3):749–755, September 1960. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177705801.

Scheuer:1961:CAM

[Sch61a] Ernest M. Scheuer. Correction to abstract “Moments of the Radial Error”: Abstracts of papers. Annals of Mathematical Statistics, 32 (3):926, September 1961. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.

aoms/1177704991. See erratum [Ano61b].

Schutzenberger:1961:SCR

[Sch61b] M. P. Schützenberger. On a special class of recurrent events. Annals of Mathematical Statistics, 32(4):1201–1213, December 1961. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177704860.

Schwarz:1962:ASB

[Sch62] Gideon Schwarz. Asymptotic shapes of Bayes sequential testing regions. Annals of Mathematical Statistics, 33(1):224–236, March 1962. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177704726.

Schmetterer:1964:RRB

[Sch64] Leopold K. Schmetterer. Review: A. Rényi, Wahrscheinlichkeitsrechnung, mit einem Anhang über Informationstheorie (German) [Probability Theory, with an Appendix on Information Theory]. Annals of Mathematical Statistics, 35(4):1828–1831, December 1964. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177700410.

Scheinok:1965:SAR

[Sch65] Perry A. Scheinok. Spectral analysis with randomly missed observations: The binomial case. Annals of Mathematical Statistics, 36 (3):971–977, June 1965. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177700069.

Schaufele:1967:PTP

[Sch67a] Ronald A. Schaufele. A potential theoretic proof of a theorem of Derman and Veinott. Annals of Mathematical Statistics, 38(2):585– 587, April 1967. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177698974.

Schuler:1967:MSP

[Sch67b] W. Schüler. Multistage sampling procedures based on prior distributions and costs. Annals of Mathematical Statistics, 38(2):464– 470, April 1967. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177698961.

Schwartz:1967:ATM

[Sch67c] Richard Schwartz. Admissible tests in multivariate analysis of variance. Annals of Mathematical Statistics, 38(3):698–710, June 1967. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177698863.

Schwartz:1967:LMT

[Sch67d] Richard E. Schwartz. Locally minimax tests. Annals of Mathematical Statistics, 38(2):340–359, April 1967. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177698951.

Schwartz:1967:EPD

[Sch67e] Stuart C. Schwartz. Estimation of probability density by an orthogonal series. Annals of Mathematical Statistics, 38(4):1261–1265, August 1967. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177698795.

Schaafsma:1968:CMS

[Sch68a] W. Schaafsma. A comparison of the most stringent and the most stringent somewhere most powerful test for certain problems with restricted alternative. Annals of Mathematical Statistics, 39(2):531– 546, April 1968. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177698415.

Schwarz:1968:ASS

[Sch68b] Gideon Schwarz. Asymptotic shapes for sequential testing of truncation parameters. Annals of Mathematical Statistics, 39(6):2038–2043, December 1968. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177698031.

Schaafsma:1969:MRU

[Sch69a] Willem Schaafsma. Minimax risk and unbiasedness for multiple decision problems of type I. Annals of Mathematical Statistics, 40 (5):1684–1720, October 1969. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177697382.

Schach:1969:ADS

[Sch69b] Siegfried Schach. The asymptotic distribution of some non linear functions of the two-sample rank vector. Annals of Mathematical Statistics, 40(3):1011–1020, June 1969. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177697605.

Schach:1969:CNT

[Sch69c] Siegfried Schach. On a class of nonparametric two-sample tests for circular distributions. Annals of Mathematical Statistics, 40(5):1791– 1800, October 1969. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177697392.

Schulzer:1969:CSP

[Sch69d] M. Schulzer. Contributions to the k-sample problem: A symmetric statistic. Annals of Mathematical Statistics, 40(6):1933–1949, December 1969. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177697277.

Schuster:1969:EPD

[Sch69e] Eugene F. Schuster. Estimation of a probability density function and its derivatives. Annals of Mathematical Statistics, 40(4):1187– 1195, August 1969. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177697495.

Schwartz:1969:IPB

[Sch69f] Richard Schwartz. Invariant proper Bayes tests for exponential families. Annals of Mathematical Statistics, 40(1):270–283, February 1969. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177697822.

Schwarz:1969:SOA

[Sch69g] Gideon Schwarz. A second-order approximation to optimal sampling regions. Annals of Mathematical Statistics, 40(1):313–315, February 1969. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177697828.

Schal:1970:MRP

[Sch70a] Manfred Schäl. Markov renewal processes with auxiliary paths. Annals of Mathematical Statistics, 41(5):1604–1623, October 1970. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177696804.

Schassberger:1970:WTQ

[Sch70b] R. Schassberger. On the waiting time in the queuing system GI/G/1. Annals of Mathematical Statistics, 41(1):182–187, February 1970.

CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177697198.

Scheffe:1970:MTV

[Sch70c] Henry Scheffé. Multiple testing versus multiple estimation. Improper confidence sets. Estimation of directions and ratios. Annals of Mathematical Statistics, 41(1):1–29, February 1970. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177697184.

Schilder:1970:SST

[Sch70d] Michael Schilder. Some structure theorems for the symmetric stable laws. Annals of Mathematical Statistics, 41(2):412–421, April 1970. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177697080.

Schuster:1970:NUC

[Sch70e] Eugene F. Schuster. Note on the uniform convergence of density estimates. Annals of Mathematical Statistics, 41(4):1347–1348, August 1970. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177696910.

Schach:1971:WCR

[Sch71a] Siegfried Schach. Weak convergence results for a class of multivariate Markov processes. Annals of Mathematical Statistics, 42(2):451– 465, April 1971. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177693397.

Schwarz:1971:SST

[Sch71b] Gideon Schwarz. A sequential Student test. Annals of Mathematical Statistics, 42(3):1003–1009, June 1971. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177693328.

Schuster:1972:JAD

[Sch72] Eugene F. Schuster. Joint asymptotic distribution of the estimated regression function at a finite number of distinct points. Annals of Mathematical Statistics, 43(1):84–88, February 1972. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177692703.

Scott:1950:NCE

[Sco50] Elizabeth L. Scott. Note on consistent estimates of the linear structural relation between two variables. Annals of Mathematical Statistics, 21(2):284–288, June 1950. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177729846.

Scott:1967:NCC

[Sco67] Alastair Scott. A note on conservative confidence regions for the mean of a multivariate normal. Annals of Mathematical Statistics, 38(1):278–280, February 1967. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177699083.

Scott:1968:CNC

[Sco68] Alastair Scott. Correction notes: Correction to “A Note on Conservative Confidence Intervals for the Mean of a Multivariate Normal”. Annals of Mathematical Statistics, 39(6):2161, December 1968. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177698053.

Scott:1971:LSP

[Sco71] Alastair Scott. Large-sample posterior distributions for finite populations. Annals of Mathematical Statistics, 42(3):1113–1117, June 1971. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177693342.

Sen:1968:PCP

[SD68] P. K. Sen and H. A. David. Paired comparisons for paired characteristics. Annals of Mathematical Statistics, 39(1):200–208, February 1968. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177698519.

Swed:1943:TTR

[SE43] Frieda S. Swed and C. Eisenhart. Tables for testing randomness of grouping in a sequence of alternatives. Annals of Mathematical Statistics, 14(1):66–87, March 1943. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http:/

/projecteuclid.org/euclid.aoms/1177731494;   http://www.
jstor.org/stable/2236004.

Seal:1955:CDP

[Sea55] K. C. Seal. On a class of decision procedures for ranking means of normal populations. Annals of Mathematical Statistics, 26(3):387– 398, September 1955. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177728486.

Seal:1956:MVA

[Sea56a] K. C. Seal. On minimum variance among certain linear functions of order statistics. Annals of Mathematical Statistics, 27(3):854–855, September 1956. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177728196.

Searle:1956:MMC

[Sea56b] S. R. Searle. Matrix methods in components of variance and covariance analysis. Annals of Mathematical Statistics, 27(3):737–748, September 1956. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177728180.

Searle:1958:SVE

[Sea58] S. R. Searle. Sampling variances of estimates of components of variance. Annals of Mathematical Statistics, 29(1):167–178, March 1958. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177706713.

Searle:1961:VCU

[Sea61] S. R. Searle. Variance components in the unbalanced 2-way nested classification. Annals of Mathematical Statistics, 32(4):1161–1166, December 1961. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177704854.

Seber:1964:LHL

[Seb64a] G. A. F. Seber. The linear hypothesis and large sample theory. Annals of Mathematical Statistics, 35(2):773–779, June 1964. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177703576.

Seber:1964:OAV

[Seb64b] G. A. F. Seber. Orthogonality in analysis of variance. Annals of Mathematical Statistics, 35(2):705–710, June 1964. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177703567.

Seelbinder:1953:STS

[See53] B. M. Seelbinder. On Stein’s two-stage sampling scheme. Annals of Mathematical Statistics, 24(4):640–649, December 1953. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177728919.

Seely:1970:LSUa

[See70a] Justus Seely. Linear spaces and unbiased estimation. Annals of Mathematical Statistics, 41(5):1725–1734, October 1970. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177696817.

Seely:1970:LSUb

[See70b] Justus Seely. Linear spaces and unbiased estimation — application to the mixed linear model. Annals of Mathematical Statistics, 41 (5):1735–1748, October 1970. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177696818.

Seely:1971:QSC

[See71] Justus Seely. Quadratic subspaces and completeness. Annals of Mathematical Statistics, 42(2):710–721, April 1971. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177693420.

Seeger:1972:NPS

[See72a] Paul Seeger. A note on partitioning a set of normal populations by their locations with respect to two controls. Annals of Mathematical Statistics, 43(6):2019–2023, December 1972. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177690876.

Seely:1972:CFM

[See72b] Justus Seely. Completeness for a family of multivariate normal distributions. Annals of Mathematical Statistics, 43(5):1644–1647, October 1972. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177692396.

Seiden:1954:PCO

[Sei54] Esther Seiden. On the problem of construction of orthogonal arrays. Annals of Mathematical Statistics, 25(1):151–156, March 1954. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177728855.

Seiden:1955:FRM

[Sei55a] Esther Seiden. Further remark on the maximum number of constraints of an orthogonal array. Annals of Mathematical Statistics, 26(4):759–763, December 1955. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177728434. See correction [Sei56].

Seiden:1955:MNC

[Sei55b] Esther Seiden. On the maximum number of constraints of an orthogonal array. Annals of Mathematical Statistics, 26(1):132–135; correction 27 (1956), 204, March 1955. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177728602.

Seiden:1956:CMN

[Sei56] Esther Seiden. Correction to “On the Maximum Number of Constraints of an Orthogonal Array”. Annals of Mathematical Statistics, 27(1):204, March 1956. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177728360. See [Sei55a].

Seiden:1961:GMC

[Sei61] Esther Seiden. On a geometrical method of construction of partially balanced designs with two associate classes. Annals of Mathematical Statistics, 32(4):1177–1180, December 1961. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177704856.

Seiden:1963:NCE

[Sei63] Esther Seiden. On necessary conditions for the existence of some symmetrical and unsymmetrical triangular PBIB designs and BIB designs. Annals of Mathematical Statistics, 34(1):348–351, March 1963. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177704279.

Seiden:1966:NCP

[Sei66] Esther Seiden. A note on construction of partially balanced incomplete block designs with parameters v = 28, n_{1} = 12, n_{2} = 15, and p_{11}^{2} = 4. Annals of Mathematical Statistics, 37(6):1783–1789, December 1966. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177699166.

Sekar:1951:TCC

[Sek51] C. Chandra Sekar. A theorem on the correlation coefficient for samples of three when the variables are independent. Annals of Mathematical Statistics, 22(1):132–133, March 1951. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177729706.

Sen:1955:SPS

[Sen55] A. R. Sen. On the selection of n primary sampling units from a stratum structure (n ≥ 2). Annals of Mathematical Statistics, 26 (4):744–751, December 1955. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177728432.

Sen:1966:DFM

[Sen66] Pranab Kumar Sen. On a distribution-free method of estimating asymptotic efficiency of a class of non-parametric tests. Annals of Mathematical Statistics, 37(6):1759–1770, December 1966. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177699164.

Sen:1967:AMP

[Sen67] Pranab Kumar Sen. Asymptotically most powerful rank order tests for grouped data. Annals of Mathematical Statistics, 38(4):1229– 1239, August 1967. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177698791.

Sen:1968:ANS

[Sen68a] Pranab Kumar Sen. Asymptotic normality of sample quantiles for mdependent processes. Annals of Mathematical Statistics, 39(5):1724– 1730, October 1968. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177698155.

Sen:1968:CAR

[Sen68b] Pranab Kumar Sen. On a class of aligned rank order tests in twoway layouts. Annals of Mathematical Statistics, 39(4):1115–1124, August 1968. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177698236.

Sen:1968:FRP

[Sen68c] Pranab Kumar Sen. On a further robustness property of the test and estimator based on Wilcoxon’s signed rank statistic. Annals of Mathematical Statistics, 39(1):282–285, February 1968. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177698535.

Sen:1968:RSN

[Sen68d] Pranab Kumar Sen. Robustness of some nonparametric procedures in linear models. Annals of Mathematical Statistics, 39(6):1913–1922, December 1968. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177698021.

Seneta:1968:RTC

[Sen68e] E. Seneta. On recent theorems concerning the supercritical Galton– Watson process. Annals of Mathematical Statistics, 39(6):2098–2102, December 1968. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/

1177698037. See extension [Hey70].

Sen:1969:CRO

[Sen69] Pranab Kumar Sen. On a class of rank order tests for the parallelism of several regression lines. Annals of Mathematical Statistics, 40(5):1668–1683, October 1969. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177697381.

Sen:1970:NOS

[Sen70a] Pranab Kumar Sen. A note on order statistics for heterogeneous distributions. Annals of Mathematical Statistics, 41(6):2137–2139, December 1970. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177696715.

Sen:1970:SCP

[Sen70b] Pranab Kumar Sen. On some convergence properties of one-sample rank order statistics. Annals of Mathematical Statistics, 41(6):2140– 2143, December 1970. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177696716.

Sen:1971:AEC

[Sen71a] Pranab Kumar Sen. Asymptotic efficiency of a class of aligned rank order tests for multiresponse experiments in some incomplete block designs. Annals of Mathematical Statistics, 42(3):1104–1112, June 1971. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177693341.

Sen:1971:HRT

[Sen71b] Pranab Kumar Sen. A Hajek–Rényi type inequality for stochastic vectors with applications to simultaneous confidence regions. Annals of Mathematical Statistics, 42(3):1132–1134, June 1971. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177693347.

Sen:1971:NWC

[Sen71c] Pranab Kumar Sen. A note on weak convergence of empirical processes for sequences of ϕ-mixing random variables. Annals of Mathematical Statistics, 42(6):2131–2133, December 1971. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177693079.

Sen:1972:SSA

[Sen72a] A. R. Sen. Successive sampling with p (p ≥ 1) auxiliary variables. Annals of Mathematical Statistics, 43(6):2031–2034, December 1972. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177690878.

Sen:1972:CAR

[Sen72b] Pranab Kumar Sen. On a class of aligned rank order tests for the identity of the intercepts of several regression lines. Annals of Mathematical Statistics, 43(6):2004–2012, December 1972. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177690874.

Sen:1972:WCE

[Sen72c] Pranab Kumar Sen. On weak convergence of extremal processes for random sample sizes. Annals of Mathematical Statistics, 43(4):1355– 1362, August 1972. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177692490.

Serfling:1968:CCL

[Ser68a] R. J. Serfling. Contributions to central limit theory for dependent variables. Annals of Mathematical Statistics, 39(4):1158–1175, August 1968. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177698240.

Serfling:1968:WTS

[Ser68b] R. J. Serfling. The Wilcoxon two-sample statistic on strongly mixing processes. Annals of Mathematical Statistics, 39(4):1202–1209, August 1968. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177698245.

Serfling:1970:CPU

[Ser70a] R. J. Serfling. Convergence properties of S_{n} under moment restrictions. Annals of Mathematical Statistics, 41(4):1235–1248, August 1970. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177696899.

Serfling:1970:MIM

[Ser70b] R. J. Serfling. Moment inequalities for the maximum cumulative sum. Annals of Mathematical Statistics, 41(4):1227–1234, August 1970. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177696898.

Serfozo:1971:RTT

[Ser71] Richard F. Serfozo. Random time transformations of semi-Markov processes. Annals of Mathematical Statistics, 42(1):176–188, February 1971. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177693505.

Seshadri:1966:CPM

[Ses66a] V. Seshadri. A characteristic property of the multivariate normal distribution. Annals of Mathematical Statistics, 37(6):1829–1831, December 1966. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177699176.

Seshadri:1966:CCE

[Ses66b] V. Seshadri. Comparison of combined estimators in balanced incomplete blocks. Annals of Mathematical Statistics, 37(6):1832–1835, December 1966. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177699177.

Seth:1949:VE

[Set49] G. R. Seth. On the variance of estimates. Annals of Mathematical Statistics, 20(1):1–27, March 1949. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177730088.

Seth:1950:DTC

[Set50] G. R. Seth. On the distribution of the two closest among a set of three observations. Annals of Mathematical Statistics, 21(2):298– 301, June 1950. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177729849.

Sethuraman:1964:PLD

[Set64] J. Sethuraman. On the probability of large deviations of families of sample means. Annals of Mathematical Statistics, 35(3):1304–1316, September 1964. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/

1177703286. See correction [Set70a].

Sethuraman:1965:PLD

[Set65] J. Sethuraman. On the probability of large deviations of the mean for random variables in D[0, 1]. Annals of Mathematical Statistics, 36(1):280–285, February 1965. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177700289.

Sethuraman:1970:CNC

[Set70a] J. Sethuraman. Correction notes: Correction to “On the Probability of Large Deviations of Families of Sample Means”. Annals of Mathematical Statistics, 41(4):1376–1380, August 1970. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177696921. See [Set64].

Sethuraman:1970:STR

[Set70b] J. Sethuraman. Stopping time of a rank-order sequential probability ratio test based on Lehmann alternatives II. Annals of Mathematical Statistics, 41(4):1322–1333, August 1970. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177696906.

Severo:1957:ABT

[Sev57] Norman C. Severo. Asymptotic behavior of tests on the mean of a logarithmico-normal distribution with known variance. Annals of Mathematical Statistics, 28(4):1044–1046, December 1957. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177706810.

Stock:1939:ASA

[SF39] J. Stevens Stock and Lester R. Frankel. The allocation of samplings among several strata. Annals of Mathematical Statistics, 10(3):288– 293, September 1939. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177732187.

Sugiura:1969:AEN

[SF69] Nariaki Sugiura and Yasunori Fujikoshi. Asymptotic expansions of the non-null distributions of the likelihood ratio criteria for multivariate linear hypothesis and independence. Annals of Mathematical Statistics, 40(3):942–952, June 1969. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177697599.

Sarhan:1956:ELS

[SG56] A. E. Sarhan and B. G. Greenberg. Estimation of location and scale parameters by order statistics from singly and doubly censored samples. Annals of Mathematical Statistics, 27(2):427–451, June 1956. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177728267. See

correction [SG69].

Sarhan:1958:ELS

[SG58] A. E. Sarhan and B. G. Greenberg. Estimation of location and scale parameters by order statistics from singly and doubly censored samples: Part II. Tables for the normal distribution for samples of size 11 ≤ n ≤ 15. Annals of Mathematical Statistics, 29(1):79–105, March 1958. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177706707.

Sarhan:1969:CNC

[SG69] A. E. Sarhan and B. G. Greenberg. Correction note: Correction to “Estimation of Location and Scale Parameters by Order Statistics from Singly and Doubly Censored Samples: Part I. The Normal Distribution up to Samples of Size 10”. Annals of Mathematical Statistics, 40(1):325, February 1969. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177697832. See [SG56].

Sen:1971:BLS

[SG71] Pranab Kumar Sen and Malay Ghosh. On bounded length sequential confidence intervals based on one-sample rank order statistics. Annals of Mathematical Statistics, 42(1):189–203, February 1971. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177693506.

Sarhan:1963:SEE

[SGO63] A. E. Sarhan, B. G. Greenberg, and Junjiro Ogawa. Simplified estimates for the exponential distribution. Annals of Mathematical Statistics, 34(1):102–116, March 1963. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177704246.

Shantaram:1972:CCL

[SH72] R. Shantaram and W. Harkness. On a certain class of limit distributions. Annals of Mathematical Statistics, 43(6):2067–2071, December 1972. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177690886.

Shapiro:1955:EEC

[Sha55] J. M. Shapiro. Error estimates for certain probability limit theorems. Annals of Mathematical Statistics, 26(4):617–630, December 1955. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177728421.

Shapiro:1957:SIT

[Sha57] J. M. Shapiro. Sums of independent truncated random variables. Annals of Mathematical Statistics, 28(3):754–761, September 1957.

CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177706887.

Shah:1958:BFE

[Sha58a] B. V. Shah. On balancing in factorial experiments. Annals of Mathematical Statistics, 29(3):766–779, September 1958. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177706535. See correc-

tion [Sha59a].

Shapiro:1958:SPI

[Sha58b] J. M. Shapiro. Sums of powers of independent random variables. Annals of Mathematical Statistics, 29(2):515–522, June 1958. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177706626.

Shah:1959:CNC

[Sha59a] B. V. Shah. Correction notes: Corrections to “On Balancing in Factorial Experiments”. Annals of Mathematical Statistics, 30(4):1267, December 1959. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/

1177706119. See [Sha58a].

Shah:1959:GPB

[Sha59b] B. V. Shah. A generalisation of partially balanced incomplete block designs. Annals of Mathematical Statistics, 30(4):1041–1050, December 1959. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177706088.

Shah:1959:GKP

[Sha59c] B. V. Shah. On a generalisation of the Kronecker product designs. Annals of Mathematical Statistics, 30(1):48–54, March 1959. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177706358.

Shah:1960:BFE

[Sha60a] B. V. Shah. Balanced factorial experiments. Annals of Mathematical Statistics, 31(2):502–514, June 1960. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177705917.

Shah:1960:MSM

[Sha60b] B. V. Shah. A matrix substitution method of constructing partially balanced designs. Annals of Mathematical Statistics, 31(1):34–42, March 1960. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177705985.

Shah:1960:OCI

[Sha60c] K. R. Shah. Optimality criteria for incomplete block designs. Annals of Mathematical Statistics, 31(3):791–794, September 1960. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177705807.

Shapiro:1960:SSP

[Sha60d] J. M. Shapiro. Sums of small powers of independent random variables. Annals of Mathematical Statistics, 31(1):222–224, March 1960. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177705999.

Shah:1963:DDI

[Sha63] B. K. Shah. Distribution of definite and of indefinite quadratic forms from a non-central normal distribution. Annals of Mathematical Statistics, 34(1):186–190, March 1963. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http:

//projecteuclid.org/euclid.aoms/1177704253. See correction

[Sha68].

Shah:1964:UIB

[Sha64a] K. R. Shah. Use of inter-block information to obtain uniformly better estimators. Annals of Mathematical Statistics, 35(3):1064–1078, September 1964. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177703265.

Shah:1964:UBN

[Sha64b] S. M. Shah. An upper bound for the number of disjoint blocks in certain PBIB designs. Annals of Mathematical Statistics, 35(1):398– 407, March 1964. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177703763.

Shah:1965:BNC

[Sha65a] S. M. Shah. Bounds for the number of common treatments between any two blocks of certain PBIB designs. Annals of Mathematical Statistics, 36(1):337–342, February 1965. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177700299.

Shanbhag:1965:NQS

[Sha65b] D. N. Shanbhag. A note on queueing systems with Erlangian service time distributions. Annals of Mathematical Statistics, 36(5):1574– 1578, October 1965. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177699919.

Shah:1966:BMO

[Sha66a] B. K. Shah. On the bivariate moments of order statistics from a logistic distribution. Annals of Mathematical Statistics, 37(4):1002– 1010, August 1966. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177699379.

Shah:1966:BSC

[Sha66b] S. M. Shah. On the block structure of certain partially balanced incomplete block designs. Annals of Mathematical Statistics, 37(4): 1016–1020, August 1966. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.

aoms/1177699381. See correction [Sha67].

Shah:1967:CNC

[Sha67] S. M. Shah. Correction note: Correction to “On the Block Structure of Certain Partially Balanced Incomplete Block Designs”. Annals of Mathematical Statistics, 38(2):624, April 1967. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177698987. See [Sha66b].

Shah:1968:CNC

[Sha68] B. K. Shah. Correction note: Correction to “Distribution of Definite and of Indefinite Quadratic Forms From a Non-central Normal Distribution”. Annals of Mathematical Statistics, 39(1):289, February 1968. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177698537. See

[Sha63].

Sharpe:1969:SPV

[Sha69a] Michael Sharpe. Sample path variations of homogeneous processes. Annals of Mathematical Statistics, 40(2):399–407, April 1969. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177697704.

Sharpe:1969:ZID

[Sha69b] Michael Sharpe. Zeroes of infinitely divisible densities. Annals of Mathematical Statistics, 40(4):1503–1505, August 1969. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177697525.

Shah:1970:DTP

[Sha70a] B. K. Shah. Distribution theory of a positive definite quadratic form with matrix argument. Annals of Mathematical Statistics, 41(2):692– 697, April 1970. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177697116.

Shah:1970:NML

[Sha70b] B. K. Shah. Note on moments of a logistic order statistics. Annals of Mathematical Statistics, 41(6):2150–2152, December 1970. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177696718.

Shah:1970:ARS

[Sha70c] K. R. Shah. Analysis of Room’s square design. Annals of Mathematical Statistics, 41(2):743–745, April 1970. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177697128.

Sharpe:1970:CVS

[Sha70d] Michael Sharpe. Centered variations of sample paths of homogeneous processes. Annals of Mathematical Statistics, 41(5):1430–1437, October 1970. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177696788.

Shah:1971:UTE

[Sha71a] K. R. Shah. Use of truncated estimator of variance ratio in recovery of inter-block information. Annals of Mathematical Statistics, 42(2): 816–819, April 1971. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177693439.

Shane:1971:IOS

[Sha71b] Harold D. Shane. On an inequality for order statistics. Annals of Mathematical Statistics, 42(5):1748–1751, October 1971. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177693175.

Srivastava:1964:ADD

[SHB64] O. P. Srivastava, W. L. Harkness, and J. B. Bartoo. Asymptotic distribution of distances between order statistics from bivariate populations. Annals of Mathematical Statistics, 35(2):748–754, June 1964. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177703573.

Shen:1936:FTI

[She36] Ching-Lai Shen. Fundamentals of the theory of inverse sampling. Annals of Mathematical Statistics, 7(2):62–112, June 1936. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177732521.

Shewhart:1945:ARP

[She45] Walter A. Shewhart. Annual report of the President of the Institute. Annals of Mathematical Statistics, 16(1):103–104, March 1945. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177731179.

Sherman:1950:RVR

[She50] B. Sherman. A random variable related to the spacing of sample values. Annals of Mathematical Statistics, 21(3):339–361, September 1950. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177729794.

Sherman:1955:TCS

[She55] S. Sherman. A theorem on convex sets with applications. Annals of Mathematical Statistics, 26(4):763–767, December 1955. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177728435.

Sherman:1957:PS

[She57] B. Sherman. Percentiles of the ω_{n} statistic. Annals of Mathematical Statistics, 28(1):259–261, March 1957. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177707052.

Sherman:1958:LDB

[She58] B. Sherman. The limiting distribution of Brownian motion in a bounded region with instantaneous return. Annals of Mathematical Statistics, 29(1):267–273, March 1958. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177706725.

Shepp:1964:LLC

[She64a] L. A. Shepp. A limit law concerning moving averages. Annals of Mathematical Statistics, 35(1):424–428, March 1964. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177703767.

Shepp:1964:LLT

[She64b] L. A. Shepp. A local limit theorem. Annals of Mathematical Statistics, 35(1):419–423, March 1964. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177703766.

Shepp:1965:DSR

[She65] L. A. Shepp. Distinguishing a sequence of random variables from a translate of itself. Annals of Mathematical Statistics, 36(4):1107– 1112, August 1965. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177699985.

Shepp:1966:RND

[She66] L. A. Shepp. Radon-Nikodým derivatives of Gaussian measures. Annals of Mathematical Statistics, 37(2):321–354, April 1966. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177699516.

Shepp:1967:FPP

[She67] L. A. Shepp. A first passage problem for the Wiener process. Annals of Mathematical Statistics, 38(6):1912–1914, December 1967. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177698626.

Shepp:1969:ESS

[She69] L. A. Shepp. Explicit solutions to some problems of optimal stopping. Annals of Mathematical Statistics, 40(3):993–1010, June 1969. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177697604.

Shepp:1971:FPT

[She71] L. A. Shepp. First passage time for a particular Gaussian process. Annals of Mathematical Statistics, 42(3):946–951, June 1971. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177693323.

Shih:1971:CMP

[Shi71] C. T. Shih. Construction of Markov processes from hitting distributions II. Annals of Mathematical Statistics, 42(1):97–114, February 1971. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177693498.

Shohat:1930:SIM

[Sho30a] J. Shohat. Stieltjes integrals in mathematical statistics. Annals of Mathematical Statistics, 1(1):73–94, February 1930. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177733261.

Shook:1930:SEMa

[Sho30b] B. L. Shook. Synopsis of elementary mathematical statistics. Annals of Mathematical Statistics, 1(1):14–41, February 1930. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177733258.

Shook:1930:SEMb

[Sho30c] B. L. Shook. Synopsis of elementary mathematical statistics. Annals of Mathematical Statistics, 1(3):224–259, August 1930. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177733090.

Shorack:1966:RGD

[Sho66] Roger A. Shorack. Recursive generation of the distribution of the Mann–Whitney–Wilcoxon U-statistic under generalized Lehmann alternatives. Annals of Mathematical Statistics, 37(1):284–286, February 1966. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177699621.

Shorack:1967:TAO

[Sho67] Galen R. Shorack. Testing against ordered alternatives in Model I analysis of variance; normal theory and nonparametric. Annals of Mathematical Statistics, 38(6):1740–1752, December 1967. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177698608.

Shorack:1969:ANL

[Sho69] Galen R. Shorack. Asymptotic normality of linear combinations of functions of order statistics. Annals of Mathematical Statistics, 40 (6):2041–2050, December 1969. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177697284.

Shorack:1972:CQS

[Sho72a] Galen R. Shorack. Convergence of quantile and spacings processes with applications. Annals of Mathematical Statistics, 43(5):1400– 1411, October 1972. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177692373.

Shorack:1972:FOS

[Sho72b] Galen R. Shorack. Functions of order statistics. Annals of Mathematical Statistics, 43(2):412–427, April 1972. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177692622.

Shrikhande:1950:ICS

[Shr50] S. S. Shrikhande. The impossibility of certain symmetrical balanced incomplete block designs. Annals of Mathematical Statistics, 21(1): 106–111, March 1950. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177729889.

Shrikhande:1951:DTW

[Shr51] S. S. Shrikhande. Designs for two-way elimination of heterogeneity. Annals of Mathematical Statistics, 22(2):235–247, June 1951. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177729643.

Shrikhande:1959:CTA

[Shr59a] S. S. Shrikhande. On a characterization of the triangular association scheme. Annals of Mathematical Statistics, 30(1):39–47, March 1959. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177706357.

Shrikhande:1959:UAS

[Shr59b] S. S. Shrikhande. The uniqueness of the L_{2} association scheme. Annals of Mathematical Statistics, 30(3):781–798, September 1959. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177706207.

Shrikhande:1965:CPB

[Shr65] S. S. Shrikhande. On a class of partially balanced incomplete block designs. Annals of Mathematical Statistics, 36(6):1807–1814, December 1965. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177699810.

Shubert:1969:BMD

[Shu69] Bruno O. Shubert. Bayesian model of decision-making as a result of learning from experience. Annals of Mathematical Statistics, 40 (6):2127–2142, December 1969. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177697290.

Shuster:1972:GWV

[Shu72] Jonathan Shuster. A generalization of the weak version of the Glivenko–Cantelli theorem. Annals of Mathematical Statistics, 43(3): 1010–1011, June 1972. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177692565.

Siddiqui:1958:CLS

[Sid58a] M. M. Siddiqui. Covariances of least-squares estimates when residuals are correlated. Annals of Mathematical Statistics, 29(4):1251–1256, December 1958. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177706456.

Siddiqui:1958:DSC

[Sid58b] M. M. Siddiqui. Distribution of a serial correlation coefficient near the ends of the range. Annals of Mathematical Statistics, 29(3):852– 861, September 1958. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177706541.

Siddiqui:1958:ISC

[Sid58c] M. M. Siddiqui. On the inversion of the sample covariance matrix in a stationary autoregressive process. Annals of Mathematical Statistics, 29(2):585–588, June 1958. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177706636.

Siddiqui:1960:TRC

[Sid60] M. M. Siddiqui. Tests for regression coefficients when errors are correlated. Annals of Mathematical Statistics, 31(4):929–938, December 1960. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177705667.

Siddiqui:1962:AMS

[Sid62a] M. M. Siddiqui. Approximations to the moments of the sample median. Annals of Mathematical Statistics, 33(1):157–168, March 1962. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177704720.

Siddiqui:1962:GAR

[Sid62b] M. M. Siddiqui. Games associated with a renewal process. Annals of Mathematical Statistics, 33(2):697–701, June 1962. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177704591.

Siddiqui:1963:OEP

[Sid63] M. M. Siddiqui. Optimum estimators of the parameters of negative exponential distributions from one or two order statistics. Annals of Mathematical Statistics, 34(1):117–121, March 1963. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177704247.

Siddiqui:1965:ADQ

[Sid65] M. M. Siddiqui. Approximations to the distribution of quadratic forms. Annals of Mathematical Statistics, 36(2):677–682, April 1965. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177700175.

Siddiqui:1967:BD

[Sid67a] M. M. Siddiqui. A bivariate t distribution. Annals of Mathematical Statistics, 38(1):162–166, February 1967. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http:

//projecteuclid.org/euclid.aoms/1177699066. See correction

[Sid67b].

Siddiqui:1967:CNC

[Sid67b] M. M. Siddiqui. Correction note: Correction to “A Bivariate t Distribution”. Annals of Mathematical Statistics, 38(5):1594, October 1967. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/

1177698717. See [Sid67a].

Sidak:1968:MNP

[Šid68] Zbyněk Šidák. On multivariate normal probabilities of rectangles: Their dependence on correlations. Annals of Mathematical Statistics, 39(5):1425–1434, October 1968. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177698122.

Sidak:1971:PRM

[Šid71] Zbyněk Šidák. On probabilities of rectangles in multivariate Student distributions: Their dependence on correlations. Annals of Mathematical Statistics, 42(1):169–175, February 1971. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177693504.

Siegmund:1967:SOS

[Sie67a] David O. Siegmund. Some one-sided stopping rules. Annals of Mathematical Statistics, 38(6):1641–1646, December 1967. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177698597.

Siegmund:1967:SPT

[Sie67b] David Oliver Siegmund. Some problems in the theory of optimal stopping rules. Annals of Mathematical Statistics, 38(6):1627–1640, December 1967. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177698596.

Siegmund:1968:ANO

[Sie68] D. Siegmund. On the asymptotic normality of one-sided stopping rules. Annals of Mathematical Statistics, 39(5):1493–1497, October 1968. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177698129.

Siegmund:1969:MMN

[Sie69a] David Siegmund. On moments of the maximum of normed partial sums. Annals of Mathematical Statistics, 40(2):527–531, April 1969. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177697720.

Siegmund:1969:VOS

[Sie69b] David Siegmund. The variance of one-sided stopping rules. Annals of Mathematical Statistics, 40(3):1074–1077, June 1969. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177697611.

Sievers:1969:PLD

[Sie69c] Gerald L. Sievers. On the probability of large deviations and exact slopes. Annals of Mathematical Statistics, 40(6):1908–1921, December 1969. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177697275.

Silber:1948:MSV

[Sil48] Jack Silber. Multiple sampling for variables. Annals of Mathematical Statistics, 19(2):246–256, June 1948. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177730249.

Silvey:1959:LMT

[Sil59] S. D. Silvey. The Lagrangian multiplier test. Annals of Mathematical Statistics, 30(2):389–407, June 1959. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177706259.

Silvey:1964:MA

[Sil64] S. D. Silvey. On a measure of association. Annals of Mathematical Statistics, 35(3):1157–1166, September 1964. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177703273.

Simon:1943:STH

[Sim43] Herbert A. Simon. Symmetric tests of the hypothesis that the mean of one normal population exceeds that of another. Annals of Mathematical Statistics, 14(2):149–154, June 1943. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177731456.

Simpson:1951:NEB

[Sim51] Paul B. Simpson. Note on the estimation of a bivariate distribution function. Annals of Mathematical Statistics, 22(3):476–478, September 1951. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177729600.

Simons:1967:CSP

[Sim67a] Gordon Simons. A class of sequential procedures for choosing one of K hypotheses concerning the unknown drift parameter of the Wiener process. Annals of Mathematical Statistics, 38(5):1376–1383, October 1967. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177698693.

Simons:1967:LBA

[Sim67b] Gordon Simons. Lower bounds for average sample number of sequential multihypothesis tests. Annals of Mathematical Statistics, 38 (5):1343–1364, October 1967. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177698691.

Simons:1967:STH

[Sim67c] Gordon Simons. A sequential three hypothesis test for determining the mean of a normal population with known variance. Annals of Mathematical Statistics, 38(5):1365–1375, October 1967. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177698692.

Simons:1968:CKV

[Sim68] Gordon Simons. On the cost of not knowing the variance when making a fixed-width confidence interval for the mean. Annals of Mathematical Statistics, 39(6):1946–1952, December 1968. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177698024.

Simons:1970:MDT

[Sim70] Gordon Simons. A martingale decomposition theorem. Annals of Mathematical Statistics, 41(3):1102–1104, June 1970. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177696991.

Simons:1971:IPL

[Sim71a] Gordon Simons. Identifying probability limits. Annals of Mathematical Statistics, 42(4):1429–1433, August 1971. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177693254.

Sims:1971:DLE

[Sim71b] Christopher A. Sims. Distributed lag estimation when the parameter space is explicitly infinite-dimensional. Annals of Mathematical Statistics, 42(5):1622–1636, October 1971. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177693161.

Singleton:1940:MMS

[Sin40] Robert R. Singleton. A method for minimizing the sum of absolute values of deviations. Annals of Mathematical Statistics, 11(3):301– 310, September 1940. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177731869.

Singh:1963:EBL

[Sin63] Rajinder Singh. Existence of bounded length confidence intervals. Annals of Mathematical Statistics, 34(4):1474–1485, December 1963. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177703879.

Sinkhorn:1964:RBA

[Sin64] Richard Sinkhorn. A relationship between arbitrary positive matrices and doubly stochastic matrices. Annals of Mathematical Statistics, 35 (2):876–879, June 1964. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177703591.

Singh:1966:CNC

[Sin66] S. N. Singh. Correction note: Correction to abstract “A Test of Inflation in Poisson Distribution”. Annals of Mathematical Statistics, 37(6):1847, December 1966. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177699181. See [Ano66d].

Singh:1967:RES

[Sin67] M. P. Singh. The relative efficiency of some two-phase sampling schemes. Annals of Mathematical Statistics, 38(3):937–940, June 1967. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177698891.

Singer:1970:SAR

[Sin70] Burton Singer. Some asymptotic results in a model of population growth. I. A class of birth and death processes. Annals of Mathematical Statistics, 41(1):115–132, February 1970. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177697193.

Singer:1971:SAR

[Sin71] Burton Singer. Some asymptotic results in a model of population growth II. Positive recurrent chains. Annals of Mathematical Statistics, 42(4):1296–1315, August 1971. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177693242.

Siotani:1971:AEN

[Sio71] Minoru Siotani. An asymptotic expansion of the non-null distribution of Hotelling’s generalized T_{0}^{2}-statistic. Annals of Mathematical Statistics, 42(2):560–571, April 1971. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177693406.

Sitgreaves:1952:DTR

[Sit52] Rosedith Sitgreaves. On the distribution of two random matrices used in classification procedures. Annals of Mathematical Statistics, 23(2): 263–270, June 1952. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177729443.

Simons:1971:CBP

[SJ71] Gordon Simons and N. L. Johnson. On the convergence of binomial to Poisson distributions. Annals of Mathematical Statistics, 42(5):1735– 1736, October 1971. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177693172.

Shah:1961:DDQ

[SK61] B. K. Shah and C. G. Khatri. Distribution of a definite quadratic form for non-central normal variates. Annals of Mathematical Statistics, 32(3):883–887, September 1961. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177704981. See correction [SK63].

Shah:1963:CNC

[SK63] B. K. Shah and C. G. Khatri. Correction notes: Correction to “Distribution of a Definite Quadratic Form for Non-central Normal Variates”. Annals of Mathematical Statistics, 34(2):673, June 1963. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177704183.

Skibinsky:1960:SPC

[Ski60] Morris Skibinsky. Some properties of a class of Bayes two-stage tests. Annals of Mathematical Statistics, 31(2):332–351, June 1960. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177705897.

Skibinsky:1967:ASS

[Ski67] Morris Skibinsky. Adequate subfields and sufficiency. Annals of Mathematical Statistics, 38(1):155–161, February 1967. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177699065.

Skibinsky:1968:MER

[Ski68] Morris Skibinsky. Minimax estimation of a random probability whose first N moments are known. Annals of Mathematical Statistics, 39 (2):492–501, April 1968. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177698412.

Skibinsky:1969:SSP

[Ski69] Morris Skibinsky. Some striking properties of binomial and beta moments. Annals of Mathematical Statistics, 40(5):1753–1764, October 1969. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177697387.

Lee:1971:DCC

[sL71] Yoong sin Lee. Distribution of the canonical correlations and asymptotic expansions for distributions of certain independence test statistics. Annals of Mathematical Statistics, 42(2):526–537, April 1971. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177693403.

Slepian:1961:FPT

[Sle61] D. Slepian. First passage time for a particular Gaussian process. Annals of Mathematical Statistics, 32(2):610–612, June 1961. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177705068.

Sherman:1950:AIM

[SM50] Jack Sherman and Winifred J. Morrison. Adjustment of an inverse matrix corresponding to a change in one element of a given matrix. Annals of Mathematical Statistics, 21(1):124–127, March 1950. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177729893.

Stanton:1966:IMB

[SM66] R. G. Stanton and R. C. Mullin. Inductive methods for balanced incomplete block designs. Annals of Mathematical Statistics, 37 (5):1348–1354, October 1966. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177699279.

Stanton:1968:CRS

[SM68] R. G. Stanton and R. C. Mullin. Construction of Room squares. Annals of Mathematical Statistics, 39(5):1540–1548, October 1968. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177698135.

Smith:1939:TAD

[Smi39] C. D. Smith. On Tchebycheff approximation for decreasing functions. Annals of Mathematical Statistics, 10(2):190–192, June 1939. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177732216.

Smith:1947:ELF

[Smi47] John H. Smith. Estimation of linear functions of cell proportions. Annals of Mathematical Statistics, 18(2):231–254, June 1947. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177730440.

Smirnov:1948:TEG

[Smi48] N. Smirnov. Table for estimating the goodness of fit of empirical distributions. Annals of Mathematical Statistics, 19(2):279–281, June 1948. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177730256;   http://www.jstor.org/stable/2236278.

Smith:1957:NTS

[Smi57] Walter L. Smith. A note on truncation and sufficient statistics. Annals of Mathematical Statistics, 28(1):247–252, March 1957. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177707049.

Smiley:1964:STF

[Smi64] M. F. Smiley. The spectral theorem for finite matrices and Cochran’s theorem. Annals of Mathematical Statistics, 35(1):443– 444, March 1964. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177703773.

Smith:1965:REP

[Smi65] Walter L. Smith. Review: Emanuel Parzen, Stochastic Processes. Annals of Mathematical Statistics, 36(5):1585–1591, October 1965. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177699925.

Smith:1968:NCA

[Smi68a] Walter L. Smith. Necessary conditions for almost sure extinction of a branching process with random environment. Annals of Mathematical Statistics, 39(6):2136–2140, December 1968. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177698045.

Smith:1968:IDL

[Smi68b] Walter L. Smith. On infinitely divisible laws and a renewal theorem for non-negative random variables. Annals of Mathematical Statistics, 39(1):139–154, February 1968. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177698511.

Sclove:1972:NOP

[SMR72] Stanley L. Sclove, Carl Morris, and R. Radhakrishnan. Nonoptimality of preliminary-test estimators for the mean of a multivariate normal distribution. Annals of Mathematical Statistics, 43 (5):1481–1490, October 1972. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177692380.

Sugiura:1968:UST

[SN68] Nariaki Sugiura and Hisao Nagao. Unbiasedness of some test criteria for the equality of one or two covariance matrices. Annals of Mathematical Statistics, 39(5):1686–1692, October 1968. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177698150.

Sugiura:1969:BTL

[SN69] Nariaki Sugiura and Hisao Nagao. On Bartlett’s test and Lehmann’s test for homogeneity of variances. Annals of Mathematical Statistics, 40(6):2018–2032, December 1969. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177697282.

Sugiura:1970:APB

[SN70] Nariaki Sugiura and Hisao Nagao. Acknowledgment of priority to “On Bartlett’s Test and Lehmann’s Test for Homogeneity of Variances”. Annals of Mathematical Statistics, 41(4):1373, August 1970. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177696917.

Snyders:1972:EFO

[Sny72] Jakov Snyders. Error formulae for optimal linear filtering, prediction and interpolation of stationary time series. Annals of Mathematical Statistics, 43(6):1935–1943, December 1972. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177690864.

Severo:1956:CTM

[SO56] Norman C. Severo and Edwin G. Olds. A comparison of tests on the mean of a logarithmico-normal distribution with known variance. Annals of Mathematical Statistics, 27(3):670–686, September 1956. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177728175.

Srivastava:1968:PSS

[SO68] M. S. Srivastava and J. Ogilvie. The performance of some sequential procedures for a ranking problem. Annals of Mathematical Statistics, 39(3):1040–1047, June 1968. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177698336.

Sobel:1953:ECC

[Sob53] Milton Sobel. An essentially complete class of decision functions for certain standard sequential problems. Annals of Mathematical Statistics, 24(3):319–337, September 1953. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177728974.

Sobel:1967:NPS

[Sob67] Milton Sobel. Nonparametric procedures for selecting the t population with the largest α-quantiles. Annals of Mathematical Statistics, 38(6):1804–1816, December 1967. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177698613.

Sobel:1971:NSG

[Sob71] Matthew J. Sobel. Noncooperative stochastic games. Annals of Mathematical Statistics, 42(6):1930–1935, December 1971. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177693059.

Solomon:1953:DMR

[Sol53] Herbert Solomon. Distribution of the measure of a random twodimensional set. Annals of Mathematical Statistics, 24(4):650–656, December 1953. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177728920.

Somerville:1958:TON

[Som58] Paul N. Somerville. Tables for obtaining non-parametric tolerance limits. Annals of Mathematical Statistics, 29(2):599–601, June 1958. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177706640.

St-Pierre:1958:DLC

[SP58] Jacques St-Pierre. Distribution of linear contrasts of order statistics. Annals of Mathematical Statistics, 29(4):1264–1268, December 1958.

CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177706458.

Smith:1965:RIT

[SP65] Carter Vincent Smith and Ronald Pyke. The Robbins–Isbell twoarmed-bandit problem with finite memory. Annals of Mathematical Statistics, 36(5):1375–1386, October 1965. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177699897.

Sen:1967:TRO

[SP67] Pranab Kumar Sen and Madan Lal Puri. On the theory of rank order tests for location in the multivariate one sample problem. Annals of Mathematical Statistics, 38(4):1216–1228, August 1967. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177698790. See correc-

tion [SP68].

Sen:1968:CNC

[SP68] Pranab Kumar Sen and Madan Lal Puri. Correction notes: Correction to “On the Theory of Rank Order Tests for Location in the Multivariate One Sample Problem”. Annals of Mathematical Statistics, 39(2):684, April 1968. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177698430. See [SP67].

Shane:1969:ROT

[SP69] Harold D. Shane and Madan L. Puri. Rank order tests for multivariate paired comparisons. Annals of Mathematical Statistics, 40 (6):2101–2117, December 1969. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177697288.

Sen:1970:ATL

[SP70] Pranab Kumar Sen and Madan Lal Puri. Asymptotic theory of likelihood ratio and rank order tests in some multivariate linear models. Annals of Mathematical Statistics, 41(1):87–100, February 1970. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177697191.

Spitzer:1966:RKI

[Spi66] F. Spitzer. Review: K. Ito, H. P. Mc Kean, Diffusion Processes and their Sample Paths. Annals of Mathematical Statistics, 37(4):1060– 1062, August 1966. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177699390.

Spitzer:1967:RWF

[Spi67] Frank Spitzer. Review: William Feller, An Introduction to Probability Theory and Its Applications. Annals of Mathematical Statistics, 38 (3):951–952, June 1967. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177698896.

Spitzer:1972:VCF

[Spi72] Frank Spitzer. A variational characterization of finite Markov chains. Annals of Mathematical Statistics, 43(1):303–307, February 1972. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177692723.

Spjotvoll:1967:OIT

[Spj67] Emil Spjøtvoll. Optimum invariant tests in unbalanced variance components models. Annals of Mathematical Statistics, 38(2):422– 428, April 1967. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177698958.

Spjotvoll:1968:MPT

[Spj68a] Emil Spjøtvoll. Most powerful tests for some non-exponential families. Annals of Mathematical Statistics, 39(3):772–784, June 1968. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177698310.

Spjotvoll:1968:NRE

[Spj68b] Emil Spjøtvoll. A note on robust estimation in analysis of variance. Annals of Mathematical Statistics, 39(5):1486–1492, October 1968. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177698128.

Spjotvoll:1972:MCR

[Spj72a] Emil Spjøtvoll. Multiple comparison of regression functions. Annals of Mathematical Statistics, 43(4):1076–1088, August 1972. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177692461.

Spjotvoll:1972:OSM

[Spj72b] Emil Spjøtvoll. On the optimality of some multiple comparison procedures. Annals of Mathematical Statistics, 43(2):398–411, April 1972. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177692621.

Springer:1953:JSD

[Spr53] Melvin D. Springer. Joint sampling distribution of the mean and standard deviation for probability density functions of doubly infinite range. Annals of Mathematical Statistics, 24(1):118–122, March 1953. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177729090.

Sprott:1955:BIB

[Spr55] D. A. Sprott. Balanced incomplete block designs and tactical configurations. Annals of Mathematical Statistics, 26(4):752–758, December 1955. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177728433.

Sprott:1956:NCI

[Spr56] D. A. Sprott. A note on combined interblock and intrablock estimation in incomplete block designs. Annals of Mathematical Statistics, 27(3):633–641, September 1956. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177728173. See correction [Spr57].

Sprott:1957:CNC

[Spr57] D. A. Sprott. Correction to “A Note on Combined Interblock and Intrablock Estimation in Incomplete Block Designs”. Annals of Mathematical Statistics, 28(1):269, March 1957. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177707056. See [Spr56].

Sprott:1959:SSG

[Spr59] D. A. Sprott. A series of symmetrical group divisible incomplete block designs. Annals of Mathematical Statistics, 30(1):249–251, March 1959. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177706383.

Sprott:1961:EAS

[Spr61] D. A. Sprott. An example of an ancillary statistic and the combination of two samples by Bayes’ Theorem. Annals of Mathematical Statistics, 32(2):616–618, June 1961. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177705071.

St-Pierre:1956:NDD

[SPZ56] J. St-Pierre and A. Zinger. The null distribution of the difference between the two largest sample values. Annals of Mathematical Statistics, 27(3):849–851, September 1956. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177728194.

Seheult:1971:UED

[SQ71] A. H. Seheult and C. P. Quesenberry. On unbiased estimation of density functions. Annals of Mathematical Statistics, 42(4):1434–1438, August 1971. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177693255.

Schwartz:1968:DAA

[SR68] S. C. Schwartz and W. L. Root. On dominating an average associated with dependent Gaussian vectors. Annals of Mathematical Statistics, 39(6):1844–1848, December 1968. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177698016.

Srikantan:1962:RRB

[Sri62] K. S. Srikantan. Recurrence relations between the PDF’s of order statistics, and some applications. Annals of Mathematical Statistics, 33(1):169–177, March 1962. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177704721.

Srivastava:1965:CNC

[Sri65a] M. S. Srivastava. Correction notes: Correction to abstract “On a Multivariate Slippage Problem”. Annals of Mathematical Statistics, 36 (5):1584, October 1965. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177699924. See [Ano64d].

Srivastava:1965:CWD

[Sri65b] M. S. Srivastava. On the complex Wishart distribution. Annals of Mathematical Statistics, 36(1):313–315, February 1965. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177700294.

Srivastava:1965:STI

[Sri65c] M. S. Srivastava. Some tests for the intraclass correlation model. Annals of Mathematical Statistics, 36(6):1802–1806, December 1965. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177699809.

Srivastava:1967:CDB

[Sri67] M. S. Srivastava. Comparing distances between multivariate populations — the problem of minimum distance. Annals of Mathematical Statistics, 38(2):550–556, April 1967. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177698970.

Srivastava:1968:GCD

[Sri68a] J. N. Srivastava. On a general class of designs for multiresponse experiments. Annals of Mathematical Statistics, 39(6):1825–1843, December 1968. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177698015.

Srivastava:1968:CNC

[Sri68b] M. S. Srivastava. Correction notes: Correction to “On the Distribution of a Multiple Correlation Matrix: Non-Central Multivariate Beta-Distributions”. Annals of Mathematical Statistics, 39(4):1359, August 1968. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/

1177698266. See [Sri68c].

Srivastava:1968:DMC

[Sri68c] M. S. Srivastava. On the distribution of a multiple correlation matrix: Non-central multivariate beta distributions. Annals of Mathematical Statistics, 39(1):227–232, February 1968. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http:

//projecteuclid.org/euclid.aoms/1177698522. See correction

[Sri68b].

Srivastava:1968:EPS

[Sri68d] R. C. Srivastava. Estimation of the parameter in the stochastic model for phage attachment to bacteria. Annals of Mathematical Statistics, 39(1):183–192, February 1968. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177698517.

Srivastava:1969:CNC

[Sri69] Surendra K. Srivastava. Correction notes: Correction to abstract “A Generalized Estimator for the Mean of a Finite Population Using Multi-Auxiliary Information”. Annals of Mathematical Statistics, 40 (4):1510, August 1969. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177697528. See [Ano68f].

Srivastava:1971:FWC

[Sri71] M. S. Srivastava. On fixed-width confidence bounds for regression parameters. Annals of Mathematical Statistics, 42(4):1403–1411, August 1971. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177693251.

Savage:1966:STR

[SS66a] I. Richard Savage and J. Sethuraman. Stopping time of a rank-order sequential probability ratio test based on Lehmann alternatives. Annals of Mathematical Statistics, 37(5):1154–1160, October 1966. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177699261. See correc-

tion [SS67].

Schaafsma:1966:MSS

[SS66b] W. Schaafsma and L. J. Smid. Most stringent somewhere most powerful tests against alternatives restricted by a number of linear inequalities. Annals of Mathematical Statistics, 37(5):1161–1172, October 1966. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177699262.

Savage:1967:CNC

[SS67] I. Richard Savage and J. Sethuraman. Correction note: Correction to “Stopping Time of a Rank-Order Sequential Probability Ratio Test Based on Lehmann Alternatives”. Annals of Mathematical Statistics, 38(4):1309, August 1967. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177698808. See [SS66a].

Savage:1969:BPR

[SS69] I. R. Savage and I. N. Shimi. A branching process without rebranching. Annals of Mathematical Statistics, 40(5):1850–1851, October 1969. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177697401.

Siegmund:1968:EOS

[SSF68] David Siegmund, Gordon Simons, and Paul Feder. Existence of optimal stopping rules for rewards related to S_{n}/n. Annals of Mathematical Statistics, 39(4):1228–1235, August 1968. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177698248.

Savage:1966:FSO

[SSW66] I. Richard Savage, Milton Sobel, and George Woodworth. Fine structure of the ordering of probabilities of rank orders in the two sample case. Annals of Mathematical Statistics, 37(1):98–112, February 1966. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177699601.

Scheffe:1944:FSS

[ST44] H. Scheffé and J. W. Tukey. A formula for sample sizes for population tolerance limits. Annals of Mathematical Statistics, 15(2): 217, June 1944. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177731287;   http://www.jstor.org/stable/2236202.

Scheffe:1945:NPE

[ST45] H. Scheffé and J. W. Tukey. Non-parametric estimation. I. Validation of order statistics. Annals of Mathematical Statistics, 16(2): 187–192, June 1945. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177731119;   http://www.jstor.org/stable/2235830.

Stratton:1964:LDB

[ST64] Howard H. Stratton, Jr. and Howard G. Tucker. Limit distributions of a branching stochastic process. Annals of Mathematical Statistics, 35 (2):557–565, June 1964. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177703555.

Singh:1968:TTP

[ST68] Jagbir Singh and W. A. Thompson, Jr. A treatment of ties in paired comparisons. Annals of Mathematical Statistics, 39(6):2002–2015, December 1968. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177698028.

Starkey:1938:TSD

[Sta38] Daisy M. Starkey. A test of the significance of the difference between means of samples from two normal populations without assuming equal variances. Annals of Mathematical Statistics, 9(3):201– 213, September 1938. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177732310.

Starkey:1939:DMC

[Sta39] D. M. Starkey. The distribution of the multiple correlation coefficient in periodogram analysis. Annals of Mathematical Statistics, 10 (4):327–336, December 1939. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177732145.

Standish:1956:DDC

[Sta56] Charles Standish. N-dimensional distributions containing a normal component. Annals of Mathematical Statistics, 27(4):1161–1165, December 1956. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177728084.

Stanton:1957:NB

[Sta57] Ralph G. Stanton. A note on BIBDS. Annals of Mathematical Statistics, 28(4):1054–1055, December 1957. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177706814.

Stanton:1959:AP

[Sta59] Ralph G. Stanton. Acknowledgment of priority. Annals of Mathematical Statistics, 30(1):254, March 1959. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177706386.

Stacy:1962:GGD

[Sta62] E. W. Stacy. A generalization of the gamma distribution. Annals of Mathematical Statistics, 33(3):1187–1192, September 1962. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177704481.

Stapleton:1963:CUD

[Sta63] James H. Stapleton. A characterization of the uniform distribution on a compact topological group. Annals of Mathematical Statistics, 34(1):319–326, March 1963. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177704269.

Starr:1965:OLT

[Sta65] Norton Starr. On an operator limit theorem of Rota. Annals of Mathematical Statistics, 36(6):1864–1866, December 1965. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177699817.

Starr:1966:AES

[Sta66a] Norman Starr. On the asymptotic efficiency of a sequential procedure for estimating the mean. Annals of Mathematical Statistics, 37 (5):1173–1185, October 1966. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177699263.

Starr:1966:PSP

[Sta66b] Norman Starr. The performance of a sequential procedure for the fixed-width interval estimation of the mean. Annals of Mathematical Statistics, 37(1):36–50, February 1966. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177699596.

Staudte:1971:CIL

[Sta71] Robert G. Staudte, Jr. A characterization of invariant loss functions. Annals of Mathematical Statistics, 42(4):1322–1327, August 1971. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177693244.

Stanley:1972:BCP

[Sta72a] Richard M. Stanley. Boundary crossing probabilities for the Kolmogorov–Smirnov statistics. Annals of Mathematical Statistics, 43(2):664–668, April 1972. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177692650.

Starr:1972:HWW

[Sta72b] Norman Starr. How to win a war if you must: Optimal stopping based on success runs. Annals of Mathematical Statistics, 43(6):1884– 1893, December 1972. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177690859.

Stephan:1942:IMA

[Ste42] Frederick F. Stephan. An iterative method of adjusting sample frequency tables when expected marginal totals are known. Annals of Mathematical Statistics, 13(2):166–178, June 1942. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177731604.

Stein:1945:TST

[Ste45a] Charles Stein. A two-sample test for a linear hypothesis whose power is independent of the variance. Annals of Mathematical Statistics, 16 (3):243–258, September 1945. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177731088.

Stephan:1945:EVV

[Ste45b] Frederick F. Stephan. The expected value and variance of the reciprocal and other negative powers of a positive Bernoullian variate. Annals of Mathematical Statistics, 16(1):50–61, March 1945. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177731170.

Stein:1946:NCS

[Ste46] Charles Stein. A note on cumulative sums. Annals of Mathematical Statistics, 17(4):498–499, December 1946. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177730890.

Stein:1950:UEM

[Ste50] Charles Stein. Unbiased estimates with minimum variance. Annals of Mathematical Statistics, 21(3):406–415, September 1950. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177729799.

Steel:1951:MGV

[Ste51a] Robert G. D. Steel. Minimum generalized variance for a set of linear functions. Annals of Mathematical Statistics, 22(3):456–460, September 1951. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177729594.

Stein:1951:PST

[Ste51b] C. M. Stein. A property of some tests of composite hypotheses. Annals of Mathematical Statistics, 22(3):475–476, September 1951. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177729599.

Stein:1955:NSC

[Ste55] Charles Stein. A necessary and sufficient condition for admissibility. Annals of Mathematical Statistics, 26(3):518–522, September 1955. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177728497.

Stein:1956:AHT

[Ste56] Charles Stein. The admissibility of Hotelling’s T^{2}-test. Annals of Mathematical Statistics, 27(3):616–623, September 1956. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177728171.

Steinhaus:1957:PE

[Ste57a] H. Steinhaus. The problem of estimation. Annals of Mathematical Statistics, 28(3):633–648, September 1957. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177706876.

Sterrett:1957:DDM

[Ste57b] Andrew Sterrett. On the detection of defective members of large populations. Annals of Mathematical Statistics, 28(4):1033–1036, December 1957. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177706807.

Steck:1958:TCT

[Ste58a] George P. Steck. A table for computing trivariate normal probabilities. Annals of Mathematical Statistics, 29(3):780–800, September 1958. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/

1177706536. See correction [Ste59a].

Steck:1958:UPE

[Ste58b] George P. Steck. A uniqueness property not enjoyed by the normal distribution. Annals of Mathematical Statistics, 29(2):604–606, June 1958. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177706642.

Steck:1959:CNC

[Ste59a] George P. Steck. Correction notes: Corrections to “A Table for Computing Trivariate Normal Probabilities”. Annals of Mathematical Statistics, 30(4):1267, December 1959. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177706120. See [Ste58a].

Stein:1959:APE

[Ste59b] Charles Stein. The admissibility of Pitman’s estimator of a single location parameter. Annals of Mathematical Statistics, 30(4):970– 979, December 1959. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177706080.

Stein:1959:EWD

[Ste59c] Charles Stein. An example of wide discrepancy between fiducial and confidence intervals. Annals of Mathematical Statistics, 30(4):877– 880, December 1959. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177706072.

Steck:1960:APU

[Ste60] G. P. Steck. Acknowledgement of priority: A uniqueness property not enjoyed by the normal distribution. Annals of Mathematical Statistics, 31(2):523, June 1960. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177705922.

Steutel:1967:NID

[Ste67] F. W. Steutel. Note on the infinite divisibility of exponential mixtures. Annals of Mathematical Statistics, 38(4):1303–1305, August 1967. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177698806.

Steutel:1968:CID

[Ste68] F. W. Steutel. A class of infinitely divisible mixtures. Annals of Mathematical Statistics, 39(4):1153–1157, August 1968. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177698239.

Steck:1969:STS

[Ste69a] G. P. Steck. The Smirnov two sample tests as rank tests. Annals of Mathematical Statistics, 40(4):1449–1466, August 1969. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177697516.

Steiger:1969:BPK

[Ste69b] W. L. Steiger. A best possible Kolmogoroff-type inequality for martingales and a characteristic property. Annals of Mathematical Statistics, 40(3):764–769, June 1969. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177697586. See erratum [Ste73].

Stephens:1969:RRB

[Ste69c] M. A. Stephens. Results from the relation between two statistics of the Kolmogorov–Smirnov type. Annals of Mathematical Statistics, 40(5):1833–1837, October 1969. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177697396.

Steutel:1969:IDR

[Ste69d] F. W. Steutel. Infinitely divisible renewal distributions. Annals of Mathematical Statistics, 40(3):1109–1113, June 1969. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177697620.

Steutel:1969:NCM

[Ste69e] F. W. Steutel. Note on completely monotone densities. Annals of Mathematical Statistics, 40(3):1130–1131, June 1969. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177697626.

Stefansky:1970:NSC

[Ste70] Wilhelmine Stefansky. A necessary and sufficient condition that for regular multiple decision problems of Type I every unbiased procedure has minimax risk. Annals of Mathematical Statistics, 41(2):

736–738, April 1970. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177697126.

Steck:1971:RPU

[Ste71a] G. P. Steck. Rectangle probabilities for uniform order statistics and the probability that the empirical distribution function lies between two distribution functions. Annals of Mathematical Statistics, 42(1): 1–11, February 1971. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177693490.

Stefansky:1971:ROM

[Ste71b] Wilhelmine Stefansky. Rejecting outliers by maximum normed residual. Annals of Mathematical Statistics, 42(1):35–45, February 1971. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177693492.

Steutel:1971:ZID

[Ste71c] F. W. Steutel. On the zeros of infinitely divisible densities. Annals of Mathematical Statistics, 42(2):812–815, April 1971. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177693438.

Steiger:1973:CNC

[Ste73] W. L. Steiger. Correction notes: Correction to “A Best Possible Kolmogoroff-Type Inequality for Martingales and a Characteristic Property”. Annals of Statistics, 1(2):400, March 1973. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1176342383. See [Ste69b].

Stigum:1963:DSP

[Sti63] Bernt P. Stigum. Dynamic stochastic processes. Annals of Mathematical Statistics, 34(1):274–283, March 1963. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177704264.

Stigum:1966:TGW

[Sti66] Bernt P. Stigum. A theorem on the Galton–Watson process. Annals of Mathematical Statistics, 37(3):695–698, June 1966. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177699464.

Stigler:1969:LFO

[Sti69] Stephen Mack Stigler. Linear functions of order statistics. Annals of Mathematical Statistics, 40(3):770–788, June 1969. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177697587.

Stone:1959:AMI

[Sto59] M. Stone. Application of a measure of information to the design and comparison of regression experiments. Annals of Mathematical Statistics, 30(1):55–70, March 1959. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177706359.

Stone:1961:NEC

[Sto61a] M. Stone. Non-equivalent comparisons of experiments and their use for experiments involving location parameters. Annals of Mathematical Statistics, 32(1):326–332, March 1961. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177705161.

Stone:1961:OP

[Sto61b] M. Stone. The opinion pool. Annals of Mathematical Statistics, 32 (4):1339–1342, December 1961. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177704873.

Stone:1963:PD

[Sto63] M. Stone. The posterior t distribution. Annals of Mathematical Statistics, 34(2):568–573, June 1963. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177704169.

Stone:1965:LLT

[Sto65a] Charles Stone. A local limit theorem for nonlattice multi-dimensional distribution functions. Annals of Mathematical Statistics, 36(2):546– 551, April 1965. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177700165.

Stone:1965:MGF

[Sto65b] Charles Stone. On moment generating functions and renewal theory. Annals of Mathematical Statistics, 36(4):1298–1301, August 1965.

CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177700003.

Stone:1965:RHM

[Sto65c] M. Stone. Right Haar measure for convergence in probability to quasi posterior distributions. Annals of Mathematical Statistics, 36 (2):440–453, April 1965. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177700154.

Stone:1966:GRR

[Sto66a] Charles Stone. The growth of a recurrent random walk. Annals of Mathematical Statistics, 37(4):1040–1041, August 1966. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177699384.

Stone:1966:ACC

[Sto66b] Charles Stone. On absolutely continuous components and renewal theory. Annals of Mathematical Statistics, 37(1):271–275, February 1966. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177699617.

Stone:1967:GBD

[Sto67] M. Stone. Generalized Bayes decision functions, admissibility and the exponential family. Annals of Mathematical Statistics, 38(3):818– 822, June 1967. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177698876.

Stone:1968:TD

[Sto68a] Charles Stone. On a theorem by Dobrushin. Annals of Mathematical Statistics, 39(5):1391–1401, October 1968. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177698120.

Stone:1968:DMS

[Sto68b] Lawrence D. Stone. On the distribution of the maximum of a semi-Markov process. Annals of Mathematical Statistics, 39(3):947–956, June 1968. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177698327.

Stout:1968:SRC

[Sto68c] William F. Stout. Some results on the complete and almost sure convergence of linear combinations of independent random variables and martingale differences. Annals of Mathematical Statistics, 39 (5):1549–1562, October 1968. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177698136.

Stone:1969:GRW

[Sto69a] Charles J. Stone. The growth of a random walk. Annals of Mathematical Statistics, 40(6):2203–2206, December 1969. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177697300.

Stone:1969:DSF

[Sto69b] Lawrence D. Stone. On the distribution of the supremum functional for semi-Markov processes with continuous state spaces. Annals of Mathematical Statistics, 40(3):844–853, June 1969. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177697592.

Stone:1970:NSC

[Sto70a] M. Stone. Necessary and sufficient condition for convergence in probability to invariant posterior distributions. Annals of Mathematical Statistics, 41(4):1349–1353, August 1970. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177696911.

Stout:1970:HWL

[Sto70b] William F. Stout. The Hartman–Wintner law of the iterated logarithm for martingales. Annals of Mathematical Statistics, 41(6):2158– 2160, December 1970. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177696721.

Stone:1972:UBR

[Sto72] Charles J. Stone. An upper bound for the renewal function. Annals of Mathematical Statistics, 43(6):2050–2052, December 1972. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177690883.

Strecker:1935:ECP

[Str35] Grace Strecker. On evaluating a coefficient of partial correlation. Annals of Mathematical Statistics, 6(3):143–145, September 1935. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177732591.

Striebel:1958:EET

[Str58] Charlotte T. Striebel. On the efficiency of estimates of trend in the Ornstein Uhlenbeck process. Annals of Mathematical Statistics, 29(1):192–200, March 1958. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177706716.

Striebel:1959:DSP

[Str59] Charlotte T. Striebel. Densities for stochastic processes. Annals of Mathematical Statistics, 30(2):559–567, June 1959. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177706268.

Striebel:1961:EER

[Str61] Charlotte T. Striebel. Efficient estimation of a regression parameter for certain second order processes. Annals of Mathematical Statistics, 32(4):1299–1313, December 1961. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177704868.

Strassen:1965:EPM

[Str65a] V. Strassen. The existence of probability measures with given marginals. Annals of Mathematical Statistics, 36(2):423–439, April 1965. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177700153.

Strauch:1965:CER

[Str65b] R. E. Strauch. Conditional expectations of random variables without expectations. Annals of Mathematical Statistics, 36(5):1556–1559, October 1965. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177699915.

Strauch:1966:NDP

[Str66a] Ralph E. Strauch. Negative dynamic programming. Annals of Mathematical Statistics, 37(4):871–890, August 1966. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177699369.

Striebel:1966:RNJ

[Str66b] Charlotte Striebel. Review: Norman T. J. Bailey, The Elements of Stochastic Processes with Applications to the Natural Sciences. Annals of Mathematical Statistics, 37(2):531, April 1966. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177699541.

Strawderman:1971:PBM

[Str71a] William E. Strawderman. Proper Bayes minimax estimators of the multivariate normal mean. Annals of Mathematical Statistics, 42(1): 385–388, February 1971. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177693528.

Stroud:1971:OLS

[Str71b] T. W. F. Stroud. On obtaining large-sample tests from asymptotically normal estimators. Annals of Mathematical Statistics, 42(4):1412– 1424, August 1971. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.

aoms/1177693252. See correction [Str72c].

Stratton:1972:FPE

[Str72a] Howard H. Stratton. Feller’s parametric equations for laws of the iterated logarithm. Annals of Mathematical Statistics, 43(6):2104– 2109, December 1972. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177690893.

Stratton:1972:MOR

[Str72b] Howard H. Stratton. Moments of oscillations and ruled sums. Annals of Mathematical Statistics, 43(3):1012–1016, June 1972. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177692566.

Stroud:1972:CNC

[Str72c] T. W. F. Stroud. Correction note: Correction to “On Obtaining Large-Sample Tests from Asymptotically Normal Estimators”. Annals of Mathematical Statistics, 43(2):695, April 1972. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177692658. See [Str71b].

Stroud:1972:FAW

[Str72d] T. W. F. Stroud. Fixed alternatives and Wald’s formulation of the noncentral asymptotic behavior of the likelihood ratio statistic. Annals of Mathematical Statistics, 43(2):447–454, April 1972. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177692625.

Studden:1967:SSP

[Stu67] W. J. Studden. On selecting a subset of k populations containing the best. Annals of Mathematical Statistics, 38(4):1072–1078, August 1967. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177698777.

Studden:1968:ODT

[Stu68] W. J. Studden. Optimal designs on Tchebycheff points. Annals of Mathematical Statistics, 39(5):1435–1447, October 1968. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177698123.

Studden:1971:ETO

[Stu71a] W. J. Studden. Elfving’s theorem and optimal designs for quadratic loss. Annals of Mathematical Statistics, 42(5):1613–1621, October 1971. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177693160.

Studden:1971:ODM

[Stu71b] W. J. Studden. Optimal designs for multivariate polynomial extrapolation. Annals of Mathematical Statistics, 42(2):828–832, April 1971. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177693442.

Sobel:1972:BTI

[SU72] Milton Sobel and V. R. R. Uppuluri. On Bonferroni-type inequalities of the same degree for the probability of unions and intersections. Annals of Mathematical Statistics, 43(5):1549–1558, October 1972. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177692387.

Subrahmanya:1966:NBE

[Sub66] M. T. Subrahmanya. A note on a biased estimator in sampling with probability proportional to size with replacement. Annals of Mathematical Statistics, 37(4):1045–1047, August 1966. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177699386.

Sudderth:1969:NTS

[Sud69a] William D. Sudderth. A note on thrifty strategies and martingales in a finitely additive setting. Annals of Mathematical Statistics, 40 (6):2211–2214, December 1969. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177697303.

Sudderth:1969:MNG

[Sud69b] William D. Sudderth. On measurable, nonleavable gambling houses with a goal. Annals of Mathematical Statistics, 40(1):66–70, February 1969. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177697804.

Sudderth:1971:FER

[Sud71a] William D. Sudderth. A “Fatou Equation” for randomly stopped variables. Annals of Mathematical Statistics, 42(6):2143–2146, December 1971. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177693082.

Sudderth:1971:GTO

[Sud71b] William D. Sudderth. A gambling theorem and optimal stopping theory. Annals of Mathematical Statistics, 42(5):1697–1705, October 1971. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177693168.

Sudderth:1971:MGP

[Sud71c] William D. Sudderth. On measurable gambling problems. Annals of Mathematical Statistics, 42(1):260–269, February 1971. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177693510.

Sudderth:1972:DSC

[Sud72] William D. Sudderth. On the Dubins and Savage characterization of optimal strategies. Annals of Mathematical Statistics, 43(2):498– 507, April 1972. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177692630.

Sugiyama:1965:DLV

[Sug65] T. Sugiyama. On the distribution of the latent vectors for principal component analysis. Annals of Mathematical Statistics, 36(6):1875– 1876, December 1965. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177699821.

Sugiyama:1966:DLL

[Sug66] T. Sugiyama. On the distribution of the largest latent root and the corresponding latent vector for principal component analysis. Annals of Mathematical Statistics, 37(4):995–1001, August 1966. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177699378.

Sugiyama:1967:DLLb

[Sug67a] T. Sugiyama. Distribution of the largest latent root and the smallest latent root of the generalized B statistic and F statistic in multivariate analysis. Annals of Mathematical Statistics, 38(4):1152–1159, August 1967. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177698784.

Sugiyama:1967:DLLa

[Sug67b] T. Sugiyama. On the distribution of the largest latent root of the covariance matrix. Annals of Mathematical Statistics, 38(4):1148– 1151, August 1967. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177698783.

Sugiura:1969:AED

[Sug69] Nariaki Sugiura. Asymptotic expansions of the distributions of the likelihood ratio criteria for covariance matrix. Annals of Mathematical Statistics, 40(6):2051–2063, December 1969. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177697285.

Sugiyama:1970:JDE

[Sug70] T. Sugiyama. Joint distribution of the extreme roots of a covariance matrix. Annals of Mathematical Statistics, 41(2):655–657, April 1970. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177697108.

Sugiura:1971:NSF

[Sug71] Nariaki Sugiura. Note on some formulas for weighted sums of zonal polynomials. Annals of Mathematical Statistics, 42(2):768– 772, April 1971. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177693428.

Sugiura:1972:LBI

[Sug72] Nariaki Sugiura. Locally best invariant test for sphericity and the limiting distributions. Annals of Mathematical Statistics, 43(4):1312– 1316, August 1972. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177692481.

Sukhatme:1957:CTS

[Suk57] Balkrishna V. Sukhatme. On certain two-sample nonparametric tests for variances. Annals of Mathematical Statistics, 28(1):188– 194, March 1957. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177707046.

Sukhatme:1958:THT

[Suk58] Balkrishna V. Sukhatme. Testing the hypothesis that two populations differ only in location. Annals of Mathematical Statistics, 29(1):60– 78, March 1958. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177706706.

Sukhatme:1972:FDP

[Suk72] Shashikala Sukhatme. Fredholm determinant of a positive definite kernel of a special type and its application. Annals of Mathematical Statistics, 43(6):1914–1926, December 1972. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177690862.

Sundrum:1954:LTS

[Sun54] R. M. Sundrum. On Lehmann’s two-sample test. Annals of Mathematical Statistics, 25(1):139–145, March 1954. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177728853.

Surendran:1968:AMK

[Sur68a] P. U. Surendran. Association matrices and the Kronecker product of designs. Annals of Mathematical Statistics, 39(2):676–680, April 1968. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177698427.

Surendran:1968:CTB

[Sur68b] P. U. Surendran. Common treatments between blocks of certain partially balanced incomplete block designs. Annals of Mathematical Statistics, 39(3):999–1006, June 1968. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177698332.

Suryanarayana:1971:PBW

[Sur71] K. V. Suryanarayana. Partially balanced weighing designs with two association classes. Annals of Mathematical Statistics, 42(4):1316– 1321, August 1971. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177693243.

Sathe:1969:MVU

[SV69a] Y. S. Sathe and S. D. Varde. On minimum variance unbiased estimation of reliability. Annals of Mathematical Statistics, 40(2):710– 714, April 1969. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177697748.

Studden:1969:ADP

[SV69b] W. J. Studden and D. J. Van Arman. Admissible designs for polynomial spline regression. Annals of Mathematical Statistics, 40(5):1557– 1569, October 1969. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177697373.

Stein:1947:SCI

[SW47] Charles Stein and Abraham Wald. Sequential confidence intervals for the mean of a normal distribution with known variance. Annals of Mathematical Statistics, 18(3):427–433, September 1947. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177730389.

Sobel:1949:SDP

[SW49] Milton Sobel and Abraham Wald. A sequential decision procedure for choosing one of three hypotheses concerning the unknown mean of a normal distribution. Annals of Mathematical Statistics, 20(4):502– 522, December 1949. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177729944.

Smith:1969:BPR

[SW69] Walter L. Smith and William E. Wilkinson. On branching processes in random environments. Annals of Mathematical Statistics, 40(3): 814–827, June 1969. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177697589.

Sobel:1972:PWR

[SW72a] Milton Sobel and George H. Weiss. Play-the-winner rule and inverse sampling for selecting the best of k ≥ 3 binomial populations. Annals of Mathematical Statistics, 43(6):1808–1826, December 1972. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177690853.

Starr:1972:FRS

[SW72b] Norman Starr and Michael Woodroofe. Further remarks on sequential estimation: The exponential case. Annals of Mathematical Statistics, 43(4):1147–1154, August 1972. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177692467.

Sweat:1970:SSD

[Swe70] Calvin W. Sweat. Sequential search with discounted income, the discount a function of the cell searched. Annals of Mathematical Statistics, 41(5):1446–1455, October 1970. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177696790.

Switzer:1964:SPB

[Swi64] Paul Switzer. Significance probability bounds for rank orderings. Annals of Mathematical Statistics, 35(2):891–894, June 1964. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177703595.

Switzer:1965:RSP

[Swi65] Paul Switzer. A random set process in the plane with a Markovian property. Annals of Mathematical Statistics, 36(6):1859–1863, December 1965. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177699816.

Switzer:1967:RPS

[Swi67] Paul Switzer. Reconstructing patterns from sample data. Annals of Mathematical Statistics, 38(1):138–154, February 1967. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177699064.

Sacks:1966:DRP

[SY66] Jerome Sacks and Donald Ylvisaker. Designs for regression problems with correlated errors. Annals of Mathematical Statistics, 37(1):66– 89, February 1966. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177699599.

Sacks:1968:DRP

[SY68] Jerome Sacks and Donald Ylvisaker. Designs for regression problems with correlated errors: Many parameters. Annals of Mathematical Statistics, 39(1):49–69, February 1968. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http:

//projecteuclid.org/euclid.aoms/1177698504. See comments

[Wah71a].

Sacks:1970:DRP

[SY70a] Jerome Sacks and Donald Ylvisaker. Designs for regression problems with correlated errors III. Annals of Mathematical Statistics, 41(6): 2057–2074, December 1970. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177696705.

Sawaragi:1970:DTM

[SY70b] Yoshikazu Sawaragi and Tsuneo Yoshikawa. Discrete-time Markovian decision processes with incomplete state observation. Annals of Mathematical Statistics, 41(1):78–86, February 1970. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177697190.

Sacks:1972:NHR

[SY72] Jerome Sacks and Donald Ylvisaker. A note on Huber’s robust estimation of a location parameter. Annals of Mathematical Statistics, 43(4):1068–1075, August 1972. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177692460.

Seiden:1966:OA

[SZ66] Esther Seiden and Rita Zemach. On orthogonal arrays. Annals of Mathematical Statistics, 37(5):1355–1370, October 1966. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177699280.

Seely:1971:LSM

[SZ71] Justus Seely and George Zyskind. Linear spaces and minimum variance unbiased estimation. Annals of Mathematical Statistics, 42(2): 691–703, April 1971. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177693418.

Szasz:1972:LTD

[Szá72] D. Szász. Limit theorems for the distributions of the sums of a random number of random variables. Annals of Mathematical Statistics, 43 (6):1902–1913, December 1972. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177690861.

Tainiter:1966:SHT

[Tai66] M. Tainiter. Sequential hypothesis tests for r-dependent marginally stationary processes. Annals of Mathematical Statistics, 37(1):90–97, February 1966. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177699600.

Takacs:1958:PPT

[Tak58] L. Takács. On a probability problem in the theory of counters. Annals of Mathematical Statistics, 29(4):1257–1263, December 1958. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177706457.

Takacs:1959:LDN

[Tak59] L. Takács. On the limiting distribution of the number of coincidences concerning telephone traffic. Annals of Mathematical Statistics, 30(1):

134–142, March 1959. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177706365.

Takacs:1961:CJ

[Tak61a] L. Takács. Charles Jordan, 1871–1959. Annals of Mathematical Statistics, 32(1):1–11, March 1961. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177705135.

Takacs:1961:CPC

[Tak61b] Lajos Takács. On a coincidence problem concerning particle counters. Annals of Mathematical Statistics, 32(3):739–756, September 1961. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177704969.

Takacs:1961:TBS

[Tak61c] Lajos Takács. The transient behavior of a single server queuing process with recurrent input and gamma service time. Annals of Mathematical Statistics, 32(4):1286–1298, December 1961. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177704867.

Takacs:1962:TDS

[Tak62] Lajos Takács. The time dependence of a single-server queue with Poisson input and general service times. Annals of Mathematical Statistics, 33(4):1340–1348, December 1962. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177704366.

Takacs:1964:ABT

[Tak64] Lajos Takács. An application of a ballot theorem in order statistics. Annals of Mathematical Statistics, 35(3):1356–1358, September 1964. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177703295.

Takacs:1969:EF

[Tak69a] Lajos Takács. On Erlang’s formula. Annals of Mathematical Statistics, 40(1):71–78, February 1969. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177697805.

Takeuchi:1969:NTL

[Tak69b] Kei Takeuchi. A note on the test for the location parameter of an exponential distribution. Annals of Mathematical Statistics, 40 (5):1838–1839, October 1969. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177697397.

Takacs:1971:CTE

[Tak71a] Lajos Takács. On the comparison of two empirical distribution functions. Annals of Mathematical Statistics, 42(4):1157–1166, August 1971. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177693232.

Takacs:1971:RJW

[Tak71b] Lajos Takacs. Review: J. W. Cohen, The Single Server Queue. Annals of Mathematical Statistics, 42(6):2162–2164, December 1971. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177693087.

Tallis:1963:ERT

[Tal63] G. M. Tallis. Elliptical and radial truncation in normal populations. Annals of Mathematical Statistics, 34(3):940–944, September 1963. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177704016.

Tamura:1963:MCR

[Tam63] Ryoji Tamura. On a modification of certain rank tests. Annals of Mathematical Statistics, 34(3):1101–1103, September 1963. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177704035.

Tamura:1966:MNS

[Tam66] Ryoji Tamura. Multivariate nonparametric several-sample tests. Annals of Mathematical Statistics, 37(3):611–618, June 1966. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177699454.

Tamhankar:1967:CN

[Tam67] M. V. Tamhankar. A characterization of normality. Annals of Mathematical Statistics, 38(6):1924–1927, December 1967. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177698630.

Tamura:1969:SMC

[Tam69] Ryoji Tamura. Some multivariate comparison procedures based on ranks. Annals of Mathematical Statistics, 40(4):1486–1491, August 1969. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177697521.

Tamura:1971:STB

[Tam71] Ryoji Tamura. On a c-sample test based on trimmed samples. Annals of Mathematical Statistics, 42(4):1455–1460, August 1971. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177693262.

Tanis:1964:LFO

[Tan64] Elliot A. Tanis. Linear forms in the order statistics from an exponential distribution. Annals of Mathematical Statistics, 35(1):270– 276, March 1964. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177703749.

Tanaka:1970:ETH

[Tan70] Masao Tanaka. Estimation and testing hypotheses for one, two, or several samples from general multivariate distributions. Annals of Mathematical Statistics, 41(6):1999–2020, December 1970. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177696701.

Tate:1953:DIN

[Tat53] Robert F. Tate. On a double inequality of the normal distribution. Annals of Mathematical Statistics, 24(1):132–134, March 1953. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177729094.

Tate:1954:CBD

[Tat54] Robert F. Tate. Correlation between a discrete and a continuous variable. Point-biserial correlation. Annals of Mathematical Statistics, 25(3):603–607, September 1954. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177728730.

Tate:1959:UEF

[Tat59] R. F. Tate. Unbiased estimation: Functions of location and scale parameters. Annals of Mathematical Statistics, 30(2):341–366, June 1959. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177706256.

Taub:1962:JNC

[Tau62] A. H. Taub, editor. John von Neumann: Collected Works. Volume IV: Continuous Geometry and Other Topics. Pergamon Press, New York, NY, USA, 1962. x + 516 pp. LCCN ????

Taylor:1953:DFR

[Tay53] William F. Taylor. Distance functions and regular best asymptotically normal estimates. Annals of Mathematical Statistics, 24(1): 85–92, March 1953. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177729085.

Taylor:1965:MSR

[Tay65] Howard M. Taylor, III. Markovian sequential replacement processes. Annals of Mathematical Statistics, 36(6):1677–1694, December 1965. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177699796.

Taylor:1968:OSM

[Tay68] Howard M. Taylor. Optimal stopping in a Markov process. Annals of Mathematical Statistics, 39(4):1333–1344, August 1968. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177698259.

Taylor:1970:MSP

[Tay70] Howard M. Taylor. A Markov stopping problem for which no entry time is ε-optimal. Annals of Mathematical Statistics, 41(6):2105– 2112, December 1970. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177696708.

Taylor:1972:BSP

[Tay72a] Howard M. Taylor. Bounds for stopped partial sums. Annals of Mathematical Statistics, 43(3):733–747, June 1972. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177692542.

Taylor:1972:WLL

[Tay72b] Robert Lee Taylor. Weak laws of large numbers in normed linear spaces. Annals of Mathematical Statistics, 43(4):1267–1274, August 1972. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177692477.

Trawinski:1964:MLE

[TB64] Irene Monahan Trawinski and R. E. Bargmann. Maximum likelihood estimation with incomplete multivariate data. Annals of Mathematical Statistics, 35(2):647–657, June 1964. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177703562.

Thompson:1971:EOS

[TBO71] M. E. Thompson, A. K. Basu, and W. L. Owen. On the existence of the optimal stopping rule in the S_{n}/n problem when the second moment is infinite. Annals of Mathematical Statistics, 42(6):1936– 1942, December 1971. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177693060.

Tarter:1965:PMO

[TC65] Michael E. Tarter and Virginia A. Clark. Properties of the median and other order statistics of logistic variates. Annals of Mathematical Statistics, 36(6):1779–1786, December 1965. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http:

//projecteuclid.org/euclid.aoms/1177699806. See correction

[TC80].

Tarter:1980:COS

[TC80] Michael E. Tarter and Virginia A. Clark. Correction to “Order Statistics of Logistic Variates”. Annals of Statistics, 8(4):935, July 1980. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176345086. See

[TC65].

Moy:1959:SRT

[tCM59] Shu teh Chen Moy. Successive recurrence times in a stationary process. Annals of Mathematical Statistics, 30(4):1254–1257, December 1959. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177706111.

Moy:1960:ESP

[tCM60] Shu teh Chen Moy. Equalities for stationary processes similar to an equality of Wald. Annals of Mathematical Statistics, 31(4):995– 1000, December 1960. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177705672.

Moy:1967:ELT

[tCM67] Shu teh C. Moy. Extensions of a limit theorem of Everett, Ulam and Harris on multitype branching processes to a branching process with countably many types. Annals of Mathematical Statistics, 38(4): 992–999, August 1967. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177698767.

Trawinski:1963:SBT

[TD63] B. J. Trawinski and H. A. David. Selection of the best treatment in a paired-comparison experiment. Annals of Mathematical Statistics, 34 (1):75–91, March 1963. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.

aoms/1177704243. See remark [Hub63b].

Thomas:1967:GVDa

[TD67] David R. Thomas and H. T. David. Game value distributions. I. Annals of Mathematical Statistics, 38(1):242–250, February 1967. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177699076.

Tjoe-Tie:1963:MTC

[Teh63] Tjoe-Tie Teh. Minimax theorems on conditionally compact sets. Annals of Mathematical Statistics, 34(4):1536–1540, December 1963. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177703885.

Teicher:1954:CD

[Tei54a] Henry Teicher. On the convolution of distributions. Annals of Mathematical Statistics, 25(4):775–778, December 1954. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177728664.

Teicher:1954:FD

[Tei54b] Henry Teicher. On the factorization of distributions. Annals of Mathematical Statistics, 25(4):769–774, December 1954. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177728663.

Teicher:1955:IPP

[Tei55a] Henry Teicher. An inequality on Poisson probabilities. Annals of Mathematical Statistics, 26(1):147–149, March 1955. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177728608.

Teichroew:1955:EPF

[Tei55b] D. Teichroew. Empirical power functions for nonparametric twosample tests for small samples. Annals of Mathematical Statistics, 26 (2):340–344, June 1955. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177728553.

Teichroew:1956:TEV

[Tei56] D. Teichroew. Tables of expected values of order statistics and products of order statistics for samples of size twenty and less from the normal distribution. Annals of Mathematical Statistics, 27(2):410– 426, June 1956. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/

1177728266. See correction [Tei61c].

Teichroew:1957:MND

[Tei57] D. Teichroew. The mixture of normal distributions with different variances. Annals of Mathematical Statistics, 28(2):510–512, June 1957. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177706981.

Teicher:1960:MD

[Tei60] Henry Teicher. On the mixture of distributions. Annals of Mathematical Statistics, 31(1):55–73, March 1960. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177705987.

Teicher:1961:IM

[Tei61a] Henry Teicher. Identifiability of mixtures. Annals of Mathematical Statistics, 32(1):244–248, March 1961. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177705155.

Teicher:1961:MLC

[Tei61b] Henry Teicher. Maximum likelihood characterization of distributions. Annals of Mathematical Statistics, 32(4):1214–1222, December 1961.

CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177704861.

Teichroew:1961:CTE

[Tei61c] D. Teichroew. Correction to “Tables of Expected Values of Order Statistics and Products of Order Statistics for Samples of Size Twenty and Less from the Normal Distribution”. Annals of Mathematical Statistics, 32(4):1345, December 1961. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177704880. See [Tei56].

Teicher:1963:IFM

[Tei63] Henry Teicher. Identifiability of finite mixtures. Annals of Mathematical Statistics, 34(4):1265–1269, December 1963. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177703862.

Teicher:1965:RSR

[Tei65] Henry Teicher. On random sums of random vectors. Annals of Mathematical Statistics, 36(5):1450–1458, October 1965. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177699904.

Teicher:1967:IMP

[Tei67] Henry Teicher. Identifiability of mixtures of product measures. Annals of Mathematical Statistics, 38(4):1300–1302, August 1967. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177698805.

Teicher:1971:CDE

[Tei71] Henry Teicher. Completion of a dominated ergodic theorem. Annals of Mathematical Statistics, 42(6):2156–2158, December 1971. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177693085.

Terry:1952:SRO

[Ter52] Milton E. Terry. Some rank order tests which are most powerful against specific parametric alternatives. Annals of Mathematical Statistics, 23(3):346–366, September 1952. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177729381.

Teugels:1968:RTW

[Teu68] Jozef L. Teugels. Renewal theorems when the first or the second moment is infinite. Annals of Mathematical Statistics, 39(4):1210– 1219, August 1968. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177698246.

Teugels:1972:NPS

[Teu72] Jozef L. Teugels. A note on Poisson-subordination. Annals of Mathematical Statistics, 43(2):676–680, April 1972. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177692653.

Tate:1958:MVU

[TG58] R. F. Tate and R. L. Goen. Minimum variance unbiased estimation for the truncated Poisson distribution. Annals of Mathematical Statistics, 29(3):755–765, September 1958. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177706534.

Tharthare:1963:RAD

[Tha63] Suresh K. Tharthare. Right angular designs. Annals of Mathematical Statistics, 34(3):1057–1067, September 1963. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177704030.

Tharthare:1965:GRA

[Tha65] Suresh K. Tharthare. Generalized right angular designs. Annals of Mathematical Statistics, 36(5):1535–1553, October 1965. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177699913.

Thedeen:1964:NPT

[The64] Torbjörn Thedéen. A note on the Poisson tendency in traffic distribution. Annals of Mathematical Statistics, 35(4):1823–1824, December 1964. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177700407.

Thiele:1931:C

[Thi31a] T. N. Thiele. Contents. Annals of Mathematical Statistics, 2(2): ??, May 1931. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177732996.

Thiele:1931:EN

[Thi31b] T. N. Thiele. Editor’s note. Annals of Mathematical Statistics, 2(2): ??, May 1931. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177732995.

Thiele:1931:LC

[Thi31c] T. N. Thiele. I. The law of causality. Annals of Mathematical Statistics, 2(2):165–169, May 1931. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177732997.

Thiele:1931:ILE

[Thi31d] T. N. Thiele. II. Laws of errors. Annals of Mathematical Statistics, 2 (2):169–172, May 1931. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177732998.

Thiele:1931:ITA

[Thi31e] T. N. Thiele. III. Tabular arrangements. Annals of Mathematical Statistics, 2(2):172–174, May 1931. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177732999.

Thiele:1931:ICE

[Thi31f] T. N. Thiele. IV. Curves of errors. Annals of Mathematical Statistics, 2(2):174–179, May 1931. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177733000.

Thiele:1931:IFF

[Thi31g] T. N. Thiele. IX. Free functions. Annals of Mathematical Statistics, 2 (2):217–228, May 1931. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177733005.

Thiele:1931:VFL

[Thi31h] T. N. Thiele. V. Functional laws of errors. Annals of Mathematical Statistics, 2(2):179–186, May 1931. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177733001.

Thiele:1931:VLEa

[Thi31i] T. N. Thiele. VI. Laws of errors: Expressed by symmetrical functions. Annals of Mathematical Statistics, 2(2):186–194, May 1931. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177733002.

Thiele:1931:VRB

[Thi31j] T. N. Thiele. VII. Relations between functional laws of errors and half-invariants. Annals of Mathematical Statistics, 2(2):194– 199, May 1931. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177733003.

Thiele:1931:VLEb

[Thi31k] T. N. Thiele. VIII. Laws of errors of function of observations. Annals of Mathematical Statistics, 2(2):199–217, May 1931. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177733004.

Thiele:1931:XA

[Thi31l] T. N. Thiele. X. Adjustment. Annals of Mathematical Statistics, 2 (2):228–235, May 1931. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177733006.

Thiele:1931:XAC

[Thi31m] T. N. Thiele. XI. Adjustment by correlates. Annals of Mathematical Statistics, 2(2):235–241, May 1931. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177733007.

Thiele:1931:XAE

[Thi31n] T. N. Thiele. XII. Adjustment by elements. Annals of Mathematical Statistics, 2(2):241–258, May 1931. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177733008.

Thiele:1931:XSA

[Thi31o] T. N. Thiele. XIII. Special auxiliary methods. Annals of Mathematical Statistics, 2(2):258–280, May 1931. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177733009.

Thiele:1931:XTP

[Thi31p] T. N. Thiele. XIV. The theory of probability. Annals of Mathematical Statistics, 2(2):280–283, May 1931. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177733010.

Thiele:1931:XFT

[Thi31q] T. N. Thiele. XV. The formal theory of probability. Annals of Mathematical Statistics, 2(2):283–293, May 1931. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177733011.

Thiele:1931:XDP

[Thi31r] T. N. Thiele. XVI. The determination of probabilities a priori and a posteriori. Annals of Mathematical Statistics, 2(2):293–301, May 1931. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177733012.

Thiele:1931:XME

[Thi31s] T. N. Thiele. XVII. Mathematical expectation and its mean error. Annals of Mathematical Statistics, 2(2):301–308, May 1931. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177733013.

Thompson:1935:CRO

[Tho35] William R. Thompson. On a criterion for the rejection of observations and the distribution of the ratio of deviation to sample standard deviation. Annals of Mathematical Statistics, 6(4):214–219, December 1935. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177732567.

Thompson:1936:CRM

[Tho36] William R. Thompson. On confidence ranges for the median and other expectation distributions for populations of unknown distribution form. Annals of Mathematical Statistics, 7(3):122–128, September 1936. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177732502.

Thompson:1938:BAN

[Tho38] William R. Thompson. Biological applications of normal range and associated significance tests in ignorance of original distribution forms. Annals of Mathematical Statistics, 9(4):281–287, December 1938. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177732285.

Thompson:1955:RVM

[Tho55a] W. A. Thompson, Jr. On the ratio of variances in the mixed incomplete block model. Annals of Mathematical Statistics, 26(4):721–733, December 1955. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177728430.

Thompson:1955:RVV

[Tho55b] W. A. Thompson, Jr. The ratio of variances in a variance components model. Annals of Mathematical Statistics, 26(2):325–329, June 1955. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177728550.

Thompson:1956:NBI

[Tho56] W. A. Thompson, Jr. A note on the balanced incomplete block designs. Annals of Mathematical Statistics, 27(3):842–846, September 1956. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177728192.

Thomasian:1957:MNS

[Tho57] A. J. Thomasian. Metrics and norms on spaces of random variables. Annals of Mathematical Statistics, 28(2):512–514, June 1957. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177706982.

Thompson:1958:NPB

[Tho58] W. A. Thompson, Jr. A note on P.B.I.B. design matrices. Annals of Mathematical Statistics, 29(3):919–922, September 1958. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177706552.

Thomasian:1960:EPA

[Tho60] A. J. Thomasian. An elementary proof of the AEP of information theory. Annals of Mathematical Statistics, 31(2):452–456, June 1960.

CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177705908.

Thompson:1962:PNE

[Tho62] W. A. Thompson, Jr. The problem of negative estimates of variance components. Annals of Mathematical Statistics, 33(1):273–289, March 1962. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177704731.

Thomasian:1963:FCI

[Tho63] A. J. Thomasian. A finite criterion for indecomposable channels. Annals of Mathematical Statistics, 34(1):337–338, March 1963. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177704274.

Thompson:1964:PSS

[Tho64] J. W. Thompson. A property of some symmetric two-stage sequential procedures. Annals of Mathematical Statistics, 35(2):755– 761, June 1964. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177703574.

Thomas:1967:GVDb

[Tho67] David R. Thomas. Game value distributions. II. Annals of Mathematical Statistics, 38(1):251–260, February 1967. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177699077.

Thomas:1972:LDR

[Tho72] D. L. Thomas. On limiting distributions of a random number of dependent random variables. Annals of Mathematical Statistics, 43 (5):1719–1726, October 1972. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177692409.

Tintner:1939:TST

[Tin39] G. Tintner. On tests of significance in time series. Annals of Mathematical Statistics, 10(2):139–143, June 1939. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177732212.

Tintner:1945:NRM

[Tin45] Gerhard Tintner. A note on rank, multicollinearity and multiple regression. Annals of Mathematical Statistics, 16(3):304–308, September 1945. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177731095.

Tintner:1952:AWC

[Tin52] G. Tintner. Abraham Wald’s contributions to econometrics. Annals of Mathematical Statistics, 23(1):21–28, March 1952. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177729482.

Tarter:1970:MDE

[TK70] Michael Tarter and Richard Kronmal. On multivariate density estimates based on orthogonal expansions. Annals of Mathematical Statistics, 41(2):718–722, April 1970. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177697122.

Thompson:1972:SOS

[TO72] M. E. Thompson and W. L. Owen. The structure of the optimal stopping rule in the S_{n}/n problem. Annals of Mathematical Statistics, 43(4):1110–1121, August 1972. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177692464.

Todorovic:1970:SPI

[Tod70] P. Todorovic. On some problems involving random number of random variables. Annals of Mathematical Statistics, 41(3):1059–1063, June 1970. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177696981.

Tomkins:1971:ILT

[Tom71] R. James Tomkins. An iterated logarithm theorem for some weighted averages of independent random variables. Annals of Mathematical Statistics, 42(2):760–763, April 1971. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177693426.

Tomkins:1972:SPI

[Tom72] R. J. Tomkins. Some probability inequalities related to the law of large numbers. Annals of Mathematical Statistics, 43(1):230–235, February 1972. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177692715.

Tong:1969:PSN

[Ton69] Yung Liang Tong. On partitioning a set of normal populations by their locations with respect to a control. Annals of Mathematical Statistics, 40(4):1300–1324, August 1969. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177697504.

Toops:1934:SFS

[Too34] Herbert A. Toops. On the systematic fitting of straight line trends by stencil and calculating machine. Annals of Mathematical Statistics, 5 (1):21–29, March 1934. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177732733.

Topsoe:1967:PWC

[Top67] Flemming Topsøe. Preservation of weak convergence under mappings. Annals of Mathematical Statistics, 38(6):1661–1665, December 1967. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177698600.

Toranzos:1952:ABS

[Tor52] Fausto I. Toranzos. An asymmetric bell-shaped frequency curve. Annals of Mathematical Statistics, 23(3):467–469, September 1952. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177729393.

Torgersen:1971:CTI

[Tor71] Erik N. Torgersen. A counterexample on translation invariant estimators. Annals of Mathematical Statistics, 42(4):1450–1451, August 1971. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177693260.

Torgersen:1972:CTE

[Tor72] Erik Nikolai Torgersen. Comparison of translation experiments. Annals of Mathematical Statistics, 43(5):1383–1399, October 1972. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177692372.

Thompson:1964:RPC

[TR64] W. A. Thompson, Jr. and Russell Remage, Jr. Rankings from paired comparisons. Annals of Mathematical Statistics, 35(2):739– 747, June 1964. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177703572.

Tracy:1968:SRC

[Tra68] Derrick S. Tracy. Some rules for a combinatorial method for multiple products of generalized k-statistics. Annals of Mathematical Statistics, 39(3):983–998, June 1968. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177698331.

Tracy:1969:SMP

[Tra69] Derrick S. Tracy. Some multiple products of polykays. Annals of Mathematical Statistics, 40(4):1297–1299, August 1969. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177697503.

Trench:1965:ESC

[Tre65] William F. Trench. On the extrapolation of a special class of stationary time series. Annals of Mathematical Statistics, 36(5):1426–1432, October 1965. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177699901.

Troskie:1971:DST

[Tro71] C. G. Troskie. The distributions of some test criteria in multivariate analysis. Annals of Mathematical Statistics, 42(5):1752–1757, October 1971. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177693176.

Truesdell:1947:NPC

[Tru47] C. Truesdell. A note on the Poisson–Charlier functions. Annals of Mathematical Statistics, 18(3):450–454, September 1947. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177730394.

Truax:1953:OST

[Tru53] Donald R. Truax. An optimum slippage test for the variances of k normal distributions. Annals of Mathematical Statistics, 24(4):669– 674, December 1953. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177728923.

Truelove:1970:BSF

[Tru70] Alan J. Truelove. Betting systems in favorable games. Annals of Mathematical Statistics, 41(2):551–566, April 1970. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177697095.

Trybula:1958:SPS

[Try58] Stanisław Trybuła. Some problems of simultaneous minimax estimation. Annals of Mathematical Statistics, 29(1):245–253, March 1958. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177706722.

Tsao:1954:EMD

[Tsa54a] Chia Kuei Tsao. An extension of Massey’s distribution of the maximum deviation between two-sample cumulative step functions. Annals of Mathematical Statistics, 25(3):587–592, September 1954. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177728727.

Tsao:1954:SSP

[Tsa54b] Chia Kuei Tsao. A simple sequential procedure for testing statistical hypotheses. Annals of Mathematical Statistics, 25(4):687–702, December 1954. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177728655.

Tsao:1955:RST

[Tsa55] Chia Kuei Tsao. Rank sum tests of fit. Annals of Mathematical Statistics, 26(1):94–104, March 1955. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177728596.

Tsao:1956:DSR

[Tsa56] Chia Kuei Tsao. Distribution of the sum in random samples from a discrete population. Annals of Mathematical Statistics, 27(3):703– 712, September 1956. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177728177.

Tsao:1957:APR

[Tsa57] Chia Kuei Tsao. Approximations to the power of rank tests. Annals of Mathematical Statistics, 28(1):159–172, March 1957. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177707043.

Tsao:1968:ADS

[Tsa68] Chia Kuei Tsao. Admissibility and distribution of some probabilistic functions of discrete finite state Markov chains. Annals of Mathematical Statistics, 39(5):1646–1653, October 1968. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177698146.

Tsutakawa:1967:APB

[Tsu67] R. K. Tsutakawa. Asymptotic properties of the block up-and-down method in bio-assay. Annals of Mathematical Statistics, 38(6):1822– 1828, December 1967. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177698615.

Tsutakawa:1968:ELD

[Tsu68] R. K. Tsutakawa. An example of large discrepancy between measures of asymptotic efficiency of tests. Annals of Mathematical Statistics, 39(1):179–182, February 1968. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177698516.

Tuckerman:1941:MSF

[Tuc41] L. B. Tuckerman. On the mathematically significant figures in the solution of simultaneous linear equations. Annals of Mathematical Statistics, 12(3):307–316, September 1941. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177731713.

Tucker:1959:CNC

[Tuc59a] Howard G. Tucker. Correction notes: Corrections to “A Generalization of the Glivenko–Cantelli Theorem”. Annals of Mathematical Statistics, 30(4):1267–1268, December 1959. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177706121. See [Tuc59b].

Tucker:1959:GGC

[Tuc59b] Howard G. Tucker. A generalization of the Glivenko–Cantelli theorem. Annals of Mathematical Statistics, 30(3):828–830, September 1959. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/

1177706212. See corrections [Tuc59a].

Tucker:1964:CSI

[Tuc64] Howard G. Tucker. On continuous singular infinitely divisible distribution functions. Annals of Mathematical Statistics, 35(1):330– 335, March 1964. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177703755.

Tucker:1968:CDA

[Tuc68] Howard G. Tucker. Convolutions of distributions attracted to stable laws. Annals of Mathematical Statistics, 39(5):1381–1390, October 1968. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/

1177698119. See note [dH70].

Tukey:1946:IDM

[Tuk46] John W. Tukey. An inequality for deviations from medians. Annals of Mathematical Statistics, 17(1):75–78, March 1946. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177731025.

Tukey:1947:NPE

[Tuk47a] John W. Tukey. Non-parametric estimation. II. Statistically equivalent blocks and tolerance regions–the continuous case. Annals of Mathematical Statistics, 18(4):529–539, December 1947. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177730343.

Tukey:1947:RSR

[Tuk47b] John W. Tukey. Review: Statistical Research Group [Columbia University, for the Office of Scientific Research and Development], Sequential Analysis of Statistical Data: Applications. Annals of Mathematical Statistics, 18(1):142–144, March 1947. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177730505.

Tukey:1948:AW

[Tuk48a] John W. Tukey. Approximate weights. Annals of Mathematical Statistics, 19(1):91–92, March 1948. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177730297.

Tukey:1948:NEI

[Tuk48b] John W. Tukey. Nonparametric estimation. III. Statistically equivalent blocks and multivariate tolerance regions–the discontinuous case. Annals of Mathematical Statistics, 19(1):30–39, March 1948. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177730287.

Tukey:1949:MRGb

[Tuk49a] John W. Tukey. Moments of random group size distributions. Annals of Mathematical Statistics, 20(4):523–539, December 1949. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177729945.

Tukey:1949:STS

[Tuk49b] John W. Tukey. Sufficiency, truncation and selection. Annals of Mathematical Statistics, 20(2):309–311, June 1949. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177730042.

Tukey:1956:KML

[Tuk56a] John W. Tukey. Keeping moment-like sampling computations simple. Annals of Mathematical Statistics, 27(1):37–54, March 1956. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177728349.

Tukey:1956:VVC

[Tuk56b] John W. Tukey. Variances of variance components. I. Balanced designs. Annals of Mathematical Statistics, 27(3):722–736, September 1956. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177728179.

Tukey:1957:CAT

[Tuk57a] John W. Tukey. On the comparative anatomy of transformations. Annals of Mathematical Statistics, 28(3):602–632, September 1957. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177706875.

Tukey:1957:SEF

[Tuk57b] John W. Tukey. Some examples with fiducial relevance. Annals of Mathematical Statistics, 28(3):687–695, September 1957. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177706880.

Tukey:1957:SRP

[Tuk57c] John W. Tukey. Sums of random partitions of ranks. Annals of Mathematical Statistics, 28(4):987–992, December 1957. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177706799.

Tukey:1957:VVCa

[Tuk57d] John W. Tukey. Variances of variance components. II. The unbalanced single classification. Annals of Mathematical Statistics, 28(1): 43–56, March 1957. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177707036.

Tukey:1957:VVCb

[Tuk57e] John W. Tukey. Variances of variance components. III. Third moments in a balanced single classification. Annals of Mathematical Statistics, 28(2):378–384, June 1957. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177706965.

Tukey:1958:BCQ

[Tuk58a] John W. Tukey. Bias and confidence in not-quite large samples (abstract). Annals of Mathematical Statistics, 29(2):614, June 1958. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177706647. This

is the other half of the Quenouille–Tukey jackknife technique; see [Que49a, Que56, Mil74, Efr79].

Tukey:1958:PBM

[Tuk58b] John W. Tukey. A problem of Berkson, and minimum variance orderly estimators. Annals of Mathematical Statistics, 29(2):588– 592, June 1958. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177706637.

Tukey:1958:SIT

[Tuk58c] John W. Tukey. A smooth invertibility theorem. Annals of Mathematical Statistics, 29(2):581–584, June 1958. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177706635.

Tukey:1962:CNC

[Tuk62a] John W. Tukey. Correction notes: Correction to “The Future of Data Analysis”. Annals of Mathematical Statistics, 33(2):812, June 1962. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177704604. See

[Tuk62b].

Tukey:1962:FDA

[Tuk62b] John W. Tukey. The future of data analysis. Annals of Mathematical Statistics, 33(1):1–67, March 1962. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177704604;   http:
//projecteuclid.org/euclid.aoms/1177704711;   http://www.
jstor.org/stable/2237638. See correction [Tuk62a].

Tulcea:1965:LPV

[Tul65] A. Ionescu Tulcea. On the lifting property. V. Annals of Mathematical Statistics, 36(3):819–828, June 1965. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177700056.

Treloar:1934:ASC

[TW34] Alan E. Treloar and Marian A. Wilder. The adequacy of “Student’s” criterion of deviations in small sample means. Annals of Mathematical Statistics, 5(4):324–341, December 1934. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177732658.

Tukey:1946:ADP

[TW46] John W. Tukey and S. S. Wilks. Approximation of the distribution of the product of beta variables by a single beta variable. Annals of Mathematical Statistics, 17(3):318–324, September 1946. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177730944.

Tretter:1979:CFI

[TW79] Marietta J. Tretter and G. William Walster. Continued fractions for the incomplete beta function: Additions and corrections. Annals of Statistics, 7(2):462–465, March 1979. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176344629. See [Aro41a].

Tweedie:1957:SPIa

[Twe57a] M. C. K. Tweedie. Statistical properties of inverse Gaussian distributions. I. Annals of Mathematical Statistics, 28(2):362–377, June 1957. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177706964.

Tweedie:1957:SPIb

[Twe57b] M. C. K. Tweedie. Statistical properties of inverse Gaussian distributions. II. Annals of Mathematical Statistics, 28(3):696–705, September 1957. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177706881.

Tweedie:1967:MSE

[Twe67] M. C. K. Tweedie. A mean-square-error characterization of binomialtype distributions. Annals of Mathematical Statistics, 38(2):620– 623, April 1967. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177698986.

ud-Din:1954:EST

[uD54] M. Zia ud Din. Expression of the k-statistics k_{9} and K_{10} in terms of power sums and sample moments. Annals of Mathematical Statistics, 25(4):800–803, December 1954. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177728671.

Ullman:1944:PCI

[Ull44] Joseph Ullman. The probability of convergence of an iterative process of inverting a matrix. Annals of Mathematical Statistics, 15(2):205– 213, June 1944. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177731285.

Usiskin:1964:MMP

[Usi64] Zalman Usiskin. Max-min probabilities in the voting paradox. Annals of Mathematical Statistics, 35(2):857–862, June 1964. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177703585.

Uzawa:1960:LMP

[Uza60] Hirofumi Uzawa. Locally most powerful rank tests for two-sample problems. Annals of Mathematical Statistics, 31(3):685–702, September 1960. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177705795.

Villars:1943:SST

[VA43] D. S. Villars and T. W. Anderson. Some significance tests for normal bivariate distributions. Annals of Mathematical Statistics, 14(2):141– 148, June 1943. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177731455.

Vajda:1945:CIR

[Vaj45] S. Vajda. On the constituent items of the reduction and the remainder in the method of least squares. Annals of Mathematical Statistics, 16 (4):381–386, December 1945. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177731064.

Vajda:1951:RBV

[Vaj51] S. Vajda. Relations between variously defined effects and interactions in analysis of variance. Annals of Mathematical Statistics, 22(2): 283–288, June 1951. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177729648.

Valand:1968:IIE

[Val68] R. S. Valand. Invariant interval estimation of a location parameter. Annals of Mathematical Statistics, 39(1):193–199, February 1968. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177698518.

VanVoorhis:1949:RCM

[Van49] W. R. Van Voorhis. Report on the Cleveland meeting of the Institute. Annals of Mathematical Statistics, 20(1):152–156, March 1949.

CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177730110.

vanderVaart:1961:CCD

[van61a] H. R. van der Vaart. On certain characteristics of the distribution of the latent roots of a symmetric random matrix under general conditions. Annals of Mathematical Statistics, 32(3):864–873, September 1961. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177704979.

vanderVaart:1961:SEI

[van61b] H. R. van der Vaart. Some extensions of the idea of bias. Annals of Mathematical Statistics, 32(2):436–447, June 1961. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177705051.

vanEeden:1963:RBP

[Van63] Constance Van Eeden. The relation between Pitman’s asymptotic relative efficiency of two tests and the correlation coefficient between their test statistics. Annals of Mathematical Statistics, 34(4):1442– 1451, December 1963. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177703876.

vanderVaart:1965:NWI

[van65] H. Robert van der Vaart. A note on Wilks’ internal scatter. Annals of Mathematical Statistics, 36(4):1308–1312, August 1965. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177700006.

VanNess:1966:ANB

[Van66a] John W. Van Ness. Asymptotic normality of bispectral estimates. Annals of Mathematical Statistics, 37(5):1257–1272, October 1966. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177699269.

VanRyzin:1966:RPF

[Van66b] J. Van Ryzin. Repetitive play in finite statistical games with unknown distributions. Annals of Mathematical Statistics, 37(4):976– 994, August 1966. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177699377.

VanRyzin:1966:SCD

[Van66c] J. Van Ryzin. The sequential compound decision problem with m×n finite loss matrix. Annals of Mathematical Statistics, 37(4):954–975, August 1966. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177699376.

VanRyzin:1966:CDP

[Van66d] J. R. Van Ryzin. The compound decision problem with m × n finite loss matrix. Annals of Mathematical Statistics, 37(2):412–424, April 1966. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177699523.

vanZwet:1967:IEV

[Van67] W. R. Van Zwet. An inequality for expected values of sample quantiles. Annals of Mathematical Statistics, 38(6):1817–1821, December 1967. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177698614.

VanRyzin:1969:SCD

[Van69] J. Van Ryzin. On strong consistency of density estimates. Annals of Mathematical Statistics, 40(5):1765–1772, October 1969. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177697388.

vanEeden:1970:ERE

[Van70] Constance Van Eeden. Efficiency-robust estimation of location. Annals of Mathematical Statistics, 41(1):172–181, February 1970. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177697197.

Vartak:1955:AKP

[Var55] Manohar Narhar Vartak. On an application of Kronecker product of matrices to statistical designs. Annals of Mathematical Statistics, 26 (3):420–438, September 1955. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177728488.

Varadarajan:1958:PMS

[Var58] V. S. Varadarajan. On a problem in measure-spaces. Annals of Mathematical Statistics, 29(4):1275–1278, December 1958. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177706461.

Vartak:1959:NEC

[Var59] Manohar Narhar Vartak. The non-existence of certain PBIB designs. Annals of Mathematical Statistics, 30(4):1051–1062, December 1959. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177706089.

Vartak:1960:RAB

[Var60] Manohar Narhar Vartak. Relations among the blocks of the Kronecker product of designs. Annals of Mathematical Statistics, 31 (3):772–778, September 1960. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177705803.

Varberg:1966:CQF

[Var66] Dale E. Varberg. Convergence of quadratic forms in independent random variables. Annals of Mathematical Statistics, 37(3):567– 576, June 1966. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177699449.

Varberg:1967:EGM

[Var67] Dale E. Varberg. Equivalent Gaussian measures with a particularly simple Radon–Nikodym derivative. Annals of Mathematical Statistics, 38(4):1027–1030, August 1967. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177698770.

Varberg:1968:ASC

[Var68] Dale E. Varberg. Almost sure convergence of quadratic forms in independent random variables. Annals of Mathematical Statistics, 39 (5):1502–1506, October 1968. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177698131.

Vatnsdal:1946:MVR

[Vat46] J. R. Vatnsdal. Minimal variance and its relation to efficient moment tests. Annals of Mathematical Statistics, 17(2):198–207, June 1946. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177730980.

vonBortkiewicz:1931:RBS

[vB31] L. von Bortkiewicz. The relations between stability and homogeneity. Annals of Mathematical Statistics, 2(1):1–22, February 1931. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177733039.

vonBahr:1965:CMC

[vB65] Bengt von Bahr. On the convergence of moments in the Central Limit Theorem. Annals of Mathematical Statistics, 36(3):808–818, June 1965. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177700055.

vonBahr:1965:IAM

[vBE65] Bengt von Bahr and Carl-Gustav Esseen. Inequalities for the r th absolute moment of a sum of random variables, 1 ≤ r ≤ 2. Annals of Mathematical Statistics, 36(1):299–303, February 1965. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177700291.

vanderWaerden:1960:SIM

[vdW60] B. L. van der Waerden. Sampling inspection as a minimum loss problem. Annals of Mathematical Statistics, 31(2):369–384, June 1960. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177705899.

vanEeden:1972:ASR

[vE72] Constance van Eeden. An analogue, for signed rank statistics, of Jureçková’s asymptotic linearity theorem for rank statistics. Annals of Mathematical Statistics, 43(3):791–802, June 1972. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177692545.

Veinott:1966:FOP

[Vei66] Arthur F. Veinott, Jr. On finding optimal policies in discrete dynamic programming with no discounting. Annals of Mathematical Statistics, 37(5):1284–1294, October 1966. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177699272.

Veinott:1969:DDP

[Vei69] Arthur F. Veinott, Jr. Discrete dynamic programming with sensitive discount optimality criteria. Annals of Mathematical Statistics, 40 (5):1635–1660, October 1969. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177697379.

Venter:1966:DSA

[Ven66] J. H. Venter. On Dvoretzky stochastic approximation theorems. Annals of Mathematical Statistics, 37(6):1534–1544, December 1966. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177699145.

Venter:1967:ERM

[Ven67a] J. H. Venter. An extension of the Robbins–Monro procedure. Annals of Mathematical Statistics, 38(1):181–190, February 1967. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177699069.

Venter:1967:CKW

[Ven67b] J. H. Venter. On convergence of the Kiefer–Wolfowitz approximation procedure. Annals of Mathematical Statistics, 38(4):1031–1036, August 1967. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177698771.

Venter:1967:EM

[Ven67c] J. H. Venter. On estimation of the mode. Annals of Mathematical Statistics, 38(5):1446–1455, October 1967. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177698699.

Villars:1947:STE

[Vil47] D. S. Villars. A significance test and estimation in the case of exponential regression. Annals of Mathematical Statistics, 18(4):596–600, December 1947. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177730353.

Villegas:1961:MLE

[Vil61] C. Villegas. Maximum likelihood estimation of a linear functional relationship. Annals of Mathematical Statistics, 32(4):1048–1062, December 1961. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177704845.

Villegas:1964:CRL

[Vil64a] C. Villegas. Confidence region for a linear relation. Annals of Mathematical Statistics, 35(2):780–788, June 1964. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177703577.

Villegas:1964:QPA

[Vil64b] C. Villegas. On qualitative probability σ-algebras. Annals of Mathematical Statistics, 35(4):1787–1796, December 1964. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177700400.

Villegas:1966:AEL

[Vil66] C. Villegas. On the asymptotic efficiency of least squares estimators. Annals of Mathematical Statistics, 37(6):1676–1683, December 1966. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177699156.

Villegas:1969:PDC

[Vil69a] C. Villegas. On the a priori distribution of the covariance matrix. Annals of Mathematical Statistics, 40(3):1098–1099, June 1969. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177697617.

Villegas:1969:LSE

[Vil69b] C. Villegas. On the least squares estimation of non-linear relations. Annals of Mathematical Statistics, 40(2):462–466, April 1969. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177697711.

Villegas:1972:BIL

[Vil72] C. Villegas. Bayesian inference in linear relations. Annals of Mathematical Statistics, 43(6):1767–1791, December 1972. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177690851.

Vere-Jones:1966:NTK

[VJ66] D. Vere-Jones. Note on a theorem of Kingman and a theorem of Chung. Annals of Mathematical Statistics, 37(6):1844–1846, December 1966. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177699180.

vonMises:1938:MBP

[vM38] R. von Mises. A modification of Bayes’ Problem. Annals of Mathematical Statistics, 9(4):256–259, December 1938. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177732282.

vonMises:1939:LDF

[vM39] R. von Mises. The limits of a distribution function if two expected values are given. Annals of Mathematical Statistics, 10(2):99–104, June 1939. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177732208.

vonMises:1941:FPS

[vM41] R. von Mises. On the foundations of probability and statistics. Annals of Mathematical Statistics, 12(2):191–205, June 1941. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177731748.

vonMises:1942:CUB

[vM42] R. von Mises. On the correct use of Bayes’ Formula. Annals of Mathematical Statistics, 13(2):156–165, June 1942. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177731603.

vonMises:1943:PTH

[vM43] R. von Mises. On the problem of testing hypotheses. Annals of Mathematical Statistics, 14(3):238–252, September 1943. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177731418.

vonMises:1945:COD

[vM45] R. von Mises. On the classification of observation data into distinct groups. Annals of Mathematical Statistics, 16(1):68–73, March 1945. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177731172.

vonMises:1947:ADD

[vM47] R. von Mises. On the asymptotic distribution of differentiable statistical functions. Annals of Mathematical Statistics, 18(3):309–348, September 1947. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177730385.

vonMises:1941:DPP

[vMD41] R. von Mises and J. L. Doob. Discussion of papers on probability theory. Annals of Mathematical Statistics, 12(2):215–217, June 1941. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177731750.

vonNeumann:1941:DRM

[vN41] John von Neumann. Distribution of the ratio of the mean square successive difference to the variance. Annals of Mathematical Statistics, 12(4):367–395, December 1941. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/ euclid.aoms/1177731677. Reprinted in [Tau62, Paper 34].

vonNeumann:1942:FRC

[vN42] John von Neumann. A further remark concerning the distribution of the ratio of the mean square successive difference to the variance. Annals of Mathematical Statistics, 13(1):86–88, March 1942. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177731645. Reprinted

in [Tau62, Paper 35].

Vandewiele:1967:ICT

[VN67] Georges Vandewiele and Marc Noé. An inequality concerning tests of fit of the Kolmogorov–Smirnov type. Annals of Mathematical Statistics, 38(4):1240–1244, August 1967. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177698792.

vonNeumann:1941:MSS

[vNKBH41] John von Neumann, R. H. Kent, H. R. Bellinson, and B. I. Hart. The mean square successive difference. Annals of Mathematical Statistics, 12(2):153–162, June 1941. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid. org/euclid.aoms/1177731746. Reprinted in [Tau62, Paper 33].

Vogel:1960:AMT

[Vog60a] Walter Vogel. An asymptotic minimax theorem for the two armed bandit problem. Annals of Mathematical Statistics, 31(2):444–451, June 1960. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177705907.

Vogel:1960:SDT

[Vog60b] Walter Vogel. A sequential design for the two armed bandit. Annals of Mathematical Statistics, 31(2):430–443, June 1960. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177705906.

Vorlickova:1972:APR

[Vor72] Dana Vorlíčková. Asymptotic properties of rank tests of symmetry under discrete distributions. Annals of Mathematical Statistics, 43 (6):2013–2018, December 1972. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177690875.

Votaw:1946:PDM

[Vot46] David F. Votaw, Jr. The probability distribution of the measure of a random linear set. Annals of Mathematical Statistics, 17(2): 240–244, June 1946. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177730986.

Votaw:1948:TCS

[Vot48] David F. Votaw, Jr. Testing compound symmetry in a normal multivariate distribution. Annals of Mathematical Statistics, 19(4):447– 473, December 1948. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177730145.

Votaw:1960:SP

[Vot60] D. F. Votaw, Jr. Statistical programming. Annals of Mathematical Statistics, 31(4):1077–1083, December 1960. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177705679.

Varadarajan:1958:TMS

[VR58] V. S. Varadarajan and R. Ranga Rao. On a theorem in metric spaces. Annals of Mathematical Statistics, 29(2):612–613, June 1958. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177706646.

vonSchelling:1949:FPS

[vS49] Hermann von Schelling. A formula for the partial sums of some hypergeometric series. Annals of Mathematical Statistics, 20(1):120– 122, March 1949. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177730101.

vonSchelling:1950:SFP

[vS50] Hermann von Schelling. A second formula for the partial sum of hypergeometric series having unity as the fourth argument. Annals of Mathematical Statistics, 21(3):458–460, September 1950. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177729808.

vonSchelling:1951:DON

[vS51] Hermann von Schelling. Distribution of the ordinal number of simultaneous events which last during a finite time. Annals of Mathematical Statistics, 22(3):452–455, September 1951. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177729593.

vanZwet:1967:CIT

[vZO67] W. R. van Zwet and J. Oosterhoff. On the combination of independent test statistics. Annals of Mathematical Statistics, 38(3): 659–680, June 1967. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177698861.

Wells:1962:DPT

[WAC62] W. T. Wells, R. L. Anderson, and John W. Cell. The distribution of the product of two central or non-central chi-square variates. Annals of Mathematical Statistics, 33(3):1016–1020, September 1962. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177704469.

Wagner:1969:RCL

[Wag69] Terry J. Wagner. On the rate of convergence for the law of large numbers. Annals of Mathematical Statistics, 40(6):2195–2197, December 1969. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177697298.

Wahba:1968:DSS

[Wah68] Grace Wahba. On the distribution of some statistics useful in the analysis of jointly stationary time series. Annals of Mathematical Statistics, 39(6):1849–1862, December 1968. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177698017.

Wahba:1971:RDP

[Wah71a] Grace Wahba. On the regression design problem of Sacks and Ylvisaker. Annals of Mathematical Statistics, 42(3):1035–1053, June 1971. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/

1177693331. See [SY68].

Wahba:1971:PAD

[Wah71b] Grace Wahba. A polynomial algorithm for density estimation. Annals of Mathematical Statistics, 42(6):1870–1886, December 1971. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177693053.

Wald:1938:GIM

[Wal38] A. Wald. Generalization of the inequality of Markoff. Annals of Mathematical Statistics, 9(4):244–255, December 1938. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177732281.

Wald:1939:CTS

[Wal39] Abraham Wald. Contributions to the theory of statistical estimation and testing hypotheses. Annals of Mathematical Statistics, 10(4):299– 326, December 1939. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177732144.

Wald:1940:FSL

[Wal40a] Abraham Wald. The fitting of straight lines if both variables are subject to error. Annals of Mathematical Statistics, 11(3):284–300, September 1940. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177731868.

Wald:1940:NAV

[Wal40b] Abraham Wald. A note on the analysis of variance with unequal class frequencies. Annals of Mathematical Statistics, 11(1):96–100, March 1940. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177731946.

Wald:1941:AMP

[Wal41a] Abraham Wald. Asymptotically most powerful tests of statistical hypotheses. Annals of Mathematical Statistics, 12(1):1–19, March 1941. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177731783.

Wald:1941:AVC

[Wal41b] Abraham Wald. On the analysis of variance in case of multiple classifications with unequal class frequencies. Annals of Mathematical Statistics, 12(3):346–350, September 1941. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177731717.

Wald:1941:SEA

[Wal41c] Abraham Wald. Some examples of asymptotically most powerful tests. Annals of Mathematical Statistics, 12(4):396–408, December 1941. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177731678.

Wald:1942:ASC

[Wal42a] Abraham Wald. Asymptotically shortest confidence intervals. Annals of Mathematical Statistics, 13(2):127–137, June 1942. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177731601.

Wald:1942:PFA

[Wal42b] Abraham Wald. On the power function of the analysis of variance test. Annals of Mathematical Statistics, 13(4):434–439, December 1942. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177731542.

Wald:1942:STL

[Wal42c] Abraham Wald. Setting of tolerance limits when the sample is large. Annals of Mathematical Statistics, 13(4):389–399, December 1942. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177731536.

Wald:1943:EWM

[Wal43a] Abraham Wald. An extension of Wilks’ method for setting tolerance limits. Annals of Mathematical Statistics, 14(1):45–55, March 1943.

CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177731491.

Wald:1943:EDS

[Wal43b] Abraham Wald. On the efficient design of statistical investigations. Annals of Mathematical Statistics, 14(2):134–140, June 1943. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177731454.

Wald:1944:NL

[Wal44a] Abraham Wald. Note on a lemma. Annals of Mathematical Statistics, 15(3):330–333, September 1944. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177731242.

Wald:1944:SPA

[Wal44b] Abraham Wald. On a statistical problem arising in the classification of an individual into one of two groups. Annals of Mathematical Statistics, 15(2):145–162, June 1944. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177731280.

Wald:1944:CSR

[Wal44c] Abraham Wald. On cumulative sums of random variables. Annals of Mathematical Statistics, 15(3):283–296, September 1944. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177731235.

Wald:1945:STS

[Wal45a] A. Wald. Sequential tests of statistical hypotheses. Annals of Mathematical Statistics, 16(2):117–186, June 1945. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177731118.

Wald:1945:SGT

[Wal45b] Abraham Wald. Some generalizations of the theory of cumulative sums of random variables. Annals of Mathematical Statistics, 16 (3):287–293, September 1945. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177731092.

Wald:1946:DUE

[Wal46a] Abraham Wald. Differentiation under the expectation sign in the fundamental identity of sequential analysis. Annals of Mathematical Statistics, 17(4):493–497, December 1946. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177730889.

Wald:1946:SIS

[Wal46b] Abraham Wald. Some improvements in setting limits for the expected number of observations required by a sequential probability ratio test. Annals of Mathematical Statistics, 17(4):466–474, December 1946. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177730885.

Walsh:1946:PFS

[Wal46c] John E. Walsh. On the power function of the sign test for slippage of means. Annals of Mathematical Statistics, 17(3):358–362, September 1946. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177730950.

Walsh:1946:SOS

[Wal46d] John E. Walsh. Some order statistic distributions for samples of size four. Annals of Mathematical Statistics, 17(2):246–248, June 1946. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177730988.

Walsh:1946:SST

[Wal46e] John E. Walsh. Some significance tests based on order statistics. Annals of Mathematical Statistics, 17(1):44–52, March 1946. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177731021.

Wald:1947:ECC

[Wal47a] Abraham Wald. An essentially complete class of admissible decision functions. Annals of Mathematical Statistics, 18(4):549–555, December 1947. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177730345.

Wald:1947:NRA

[Wal47b] Abraham Wald. A note on regression analysis. Annals of Mathematical Statistics, 18(4):586–589, December 1947. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177730350.

Walsh:1947:CEI

[Wal47c] John E. Walsh. Concerning the effect of intraclass correlation on certain significance tests. Annals of Mathematical Statistics, 18(1): 88–96, March 1947. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177730495.

Walsh:1947:ETP

[Wal47d] John E. Walsh. An extension to two populations of an analogue of Student’s t-test using the sample range. Annals of Mathematical Statistics, 18(2):280–285, June 1947. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177730445.

Walsh:1947:PET

[Wal47e] John E. Walsh. On the power efficiency of a t-test formed by pairing sample values. Annals of Mathematical Statistics, 18(4):601–604, December 1947. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177730354.

Wald:1948:APM

[Wal48a] Abraham Wald. Asymptotic properties of the maximum likelihood estimate of an unknown parameter of a discrete stochastic process. Annals of Mathematical Statistics, 19(1):40–46, March 1948. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177730288.

Wald:1948:EPW

[Wal48b] Abraham Wald. Estimation of a parameter when the number of unknown parameters increases indefinitely with the number of observations. Annals of Mathematical Statistics, 19(2):220–227, June 1948. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177730246.

Walsh:1948:UNC

[Wal48c] John E. Walsh. On the use of the non-central t-distribution for comparing percentage points of normal populations. Annals of Mathematical Statistics, 19(1):93–94, March 1948. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177730298.

Wald:1949:NCM

[Wal49a] Abraham Wald. Note on the consistency of the maximum likelihood estimate. Annals of Mathematical Statistics, 20(4):595–601, December 1949. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177729952.

Wald:1949:RPI

[Wal49b] Abraham Wald. Report of the President of the Institute for 1948. Annals of Mathematical Statistics, 20(1):156–160, March 1949. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177730111.

Wald:1949:SDF

[Wal49c] Abraham Wald. Statistical decision functions. Annals of Mathematical Statistics, 20(2):165–205, June 1949. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177730030.

Walsh:1949:CCR

[Wal49d] John E. Walsh. Concerning compound randomization in the binary system. Annals of Mathematical Statistics, 20(4):580–589, December 1949. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177729950;   http://www.jstor.org/stable/2236313.

Walsh:1949:PFB

[Wal49e] John E. Walsh. On the power function of the “best” t-test solution of the Behrens–Fisher problem. Annals of Mathematical Statistics, 20(4):616–618, December 1949. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177729957.

Walsh:1949:RMT

[Wal49f] John E. Walsh. On the range-midrange test and some tests with bounded significance levels. Annals of Mathematical Statistics, 20 (2):257–267, June 1949. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177730034.

Walsh:1949:SST

[Wal49g] John E. Walsh. Some significance tests for the median which are valid under very general conditions. Annals of Mathematical Statistics, 20 (1):64–81, March 1949. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177730091.

Walsh:1950:SET

[Wal50a] John E. Walsh. Some estimates and tests based on the r smallest values in a sample. Annals of Mathematical Statistics, 21(3):386– 397, September 1950. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177729797.

Walsh:1950:SNT

[Wal50b] John E. Walsh. Some nonparametric tests of whether the largest observations of a set are too large or too small. Annals of Mathematical Statistics, 21(4):583–592, December 1950. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http:

//projecteuclid.org/euclid.aoms/1177729753. See correction

[Wal53].

Walsh:1951:SBS

[Wal51] John E. Walsh. Some bounded significance level properties of the equal-tail sign test. Annals of Mathematical Statistics, 22(3):408– 417, September 1951. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177729587.

Walsh:1953:CSN

[Wal53] John E. Walsh. Correction to “Some Nonparametric Tests of Whether the Largest Observations of a set are too Large or too Small”. Annals of Mathematical Statistics, 24(1):134–135, March 1953. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177729095. See

[Wal50b].

Walsh:1957:EFP

[Wal57] John E. Walsh. Estimating future from past in life testing. Annals of Mathematical Statistics, 28(2):432–441, June 1957. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177706971.

Wallace:1958:AAD

[Wal58a] David L. Wallace. Asymptotic approximations to distributions. Annals of Mathematical Statistics, 29(3):635–654, September 1958. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177706528.

Wallace:1958:IRC

[Wal58b] David L. Wallace. Intersection region confidence procedures with an application to the location of the maximum in quadratic regression. Annals of Mathematical Statistics, 29(2):455–475, June 1958. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177706622.

Walsh:1958:NES

[Wal58c] John E. Walsh. Nonparametric estimation of sample percentage point standard deviation. Annals of Mathematical Statistics, 29(2):601– 604, June 1958. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177706641.

Wallace:1959:BNA

[Wal59a] David L. Wallace. Bounds on normal approximations to Student’s and the chi-square distributions. Annals of Mathematical Statistics, 30(4):1121–1130, December 1959. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177706095. See correction [Bre57b].

Wallace:1959:CCL

[Wal59b] David L. Wallace. Conditional confidence level properties. Annals of Mathematical Statistics, 30(4):864–876, December 1959. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177706071.

Walsh:1959:ENT

[Wal59c] John E. Walsh. Exact nonparametric tests for randomized blocks. Annals of Mathematical Statistics, 30(4):1034–1040, December 1959. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177706087.

Wallace:1960:CNC

[Wal60] David L. Wallace. Correction notes: Correction to “Bounds on Normal Approximations to Student’s and the Chi-Square Distributions”. Annals of Mathematical Statistics, 31(3):810, September 1960. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177705813. See

[Wal59a].

Walsh:1962:DFT

[Wal62] John E. Walsh. Distribution-free tolerance intervals for continuous symmetrical populations. Annals of Mathematical Statistics, 33(3): 1167–1174, September 1962. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177704479.

Walk:1970:CPM

[Wal70a] Harro Walk. Convergence properties of martingale transforms. Annals of Mathematical Statistics, 41(2):706–709, April 1970. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177697119.

Walsh:1970:SRF

[Wal70b] John B. Walsh. Some remarks on the Feller property. Annals of Mathematical Statistics, 41(5):1672–1683, October 1970. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177696811.

Wang:1968:RWD

[Wan68] Peter C. C. Wang. Review: Winifred D. Ashton, Theory of Road Traffic Flow. Annals of Mathematical Statistics, 39(1):293, February 1968. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177698540.

Warren:1960:APL

[War60] Richard Warren. An associated polynomial for least squares approximations. Annals of Mathematical Statistics, 31(4):969–981, December 1960. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177705670.

Wasow:1951:DRW

[Was51] Wolfgang Wasow. On the duration of random walks. Annals of Mathematical Statistics, 22(2):199–216, June 1951. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177729641.

Watson:1954:EVS

[Wat54] G. S. Watson. Extreme values in samples from m-dependent stationary stochastic processes. Annals of Mathematical Statistics, 25 (4):798–800, December 1954. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177728670.

Watterson:1959:LEC

[Wat59] G. A. Watterson. Linear estimation in censored samples from multivariate normal populations. Annals of Mathematical Statistics, 30 (3):814–824, September 1959. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177706210.

Watterson:1961:MCA

[Wat61] G. A. Watterson. Markov chains with absorbing states: A genetic example. Annals of Mathematical Statistics, 32(3):716–729, September 1961. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177704967.

Watterson:1962:STA

[Wat62] G. A. Watterson. Some theoretical aspects of diffusion theory in population genetics. Annals of Mathematical Statistics, 33(3):939– 957, September 1962. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.

aoms/1177704463. See correction [Wat63].

Watterson:1963:CST

[Wat63] G. A. Watterson. Correction to “Some Theoretical Aspects of Diffusion Theory in Population Genetics”. Annals of Mathematical Statistics, 34(1):352, March 1963. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177704280. See [Wat62].

Watson:1966:RFM

[Wat66] G. S. Watson. Review: Frederick Mosteller, David L. Wallace, Inference and Disputed Authorship: The Federalist. Annals of Mathematical Statistics, 37(1):308–312, February 1966. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177699628.

Watson:1967:LLS

[Wat67] Geoffrey S. Watson. Linear least squares regression. Annals of Mathematical Statistics, 38(6):1679–1699, December 1967. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177698603.

Watson:1969:DEO

[Wat69] Geoffrey S. Watson. Density estimation by orthogonal series. Annals of Mathematical Statistics, 40(4):1496–1498, August 1969. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177697523.

Waterman:1971:NRE

[Wat71a] Michael S. Waterman. A note on the reparametrization of an exponential family. Annals of Mathematical Statistics, 42(2):752–754, April 1971. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177693424.

Watson:1971:RMG

[Wat71b] G. S. Watson. Review: M. G. Kendall, A. Stuart, The Advanced Theory of Statistics. Annals of Mathematical Statistics, 42(3):1138– 1140, June 1971. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177693350.

Waugh:1945:NCH

[Wau45] F. V. Waugh. A note concerning Hotelling’s method of inverting a partitioned matrix. Annals of Mathematical Statistics, 16(2):216– 217, June 1945. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/

1177731124. See [Hot43b].

Wald:1941:DWS

[WB41] A. Wald and R. J. Brookner. On the distribution of Wilks’ statistic for testing the independence of several groups of variates. Annals of Mathematical Statistics, 12(2):137–152, June 1941. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177731745.

Williamson:1964:TLS

[WB64] E. Williamson and M. H. Bretherton. Tables of the logarithmic series distribution. Annals of Mathematical Statistics, 35(1):284–297, March 1964. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177703751.

Wilks:1939:OPC

[WD39] S. S. Wilks and J. F. Daly. An optimum property of confidence regions associated with the likelihood function. Annals of Mathematical Statistics, 10(3):225–235, September 1939. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177732181.

Waugh:1945:CCI

[WD45] Frederick V. Waugh and Paul S. Dwyer. Compact computation of the inverse of a matrix. Annals of Mathematical Statistics, 16(3):259– 271, September 1945. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177731089.

Watson:1951:ETS

[WD51] G. S. Watson and J. Durbin. Exact tests of serial correlation using noncircular statistics. Annals of Mathematical Statistics, 22(3):446– 451, September 1951. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177729592.

Weesakul:1961:RWB

[Wee61] B. Weesakul. The random walk between a reflecting and an absorbing barrier. Annals of Mathematical Statistics, 32(3):765–769, September 1961. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177704971.

Wegner:1956:PST

[Weg56] L. H. Wegner. Properties of some two-sample tests based on a particular measure of discrepancy. Annals of Mathematical Statistics, 27 (4):1006–1016, December 1956. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177728070.

Wegman:1969:NEU

[Weg69] Edward J. Wegman. A note on estimating a unimodal density. Annals of Mathematical Statistics, 40(5):1661–1667, October 1969. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177697380.

Wegman:1970:MLEa

[Weg70a] Edward J. Wegman. Maximum likelihood estimation of a unimodal density function. Annals of Mathematical Statistics, 41(2):457–471, April 1970. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177697085.

Wegman:1970:MLEb

[Weg70b] Edward J. Wegman. Maximum likelihood estimation of a unimodal density, II. Annals of Mathematical Statistics, 41(6):2169–2174, December 1970. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177696724.

Wegman:1971:NEM

[Weg71] Edward J. Wegman. A note on the estimation of the mode. Annals of Mathematical Statistics, 42(6):1909–1915, December 1971. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177693056.

Weida:1934:MC

[Wei34] Frank M. Weida. On measures of contingency. Annals of Mathematical Statistics, 5(4):308–319, December 1934. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177732656.

Weida:1935:CDF

[Wei35] Frank M. Weida. On certain distribution functions when the law of the universe is Poisson’s first law of error. Annals of Mathematical Statistics, 6(2):102–110, June 1935. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177732615. See erratum [Ano35b].

Weiler:1952:MES

[Wei52] H. Weiler. On the most economical sample size for controlling the mean of a population. Annals of Mathematical Statistics, 23(2):247– 254, June 1952. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177729441.

Weiss:1953:HOC

[Wei53a] Lionel Weiss. A higher order complete class theorem. Annals of Mathematical Statistics, 24(4):677–680, December 1953. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177728925.

Weiss:1953:TOS

[Wei53b] Lionel Weiss. Testing one simple hypothesis against another. Annals of Mathematical Statistics, 24(2):273–281, June 1953. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177729033. See note [Le 54].

Weil:1954:DRE

[Wei54a] Herschel Weil. The distribution of radial error. Annals of Mathematical Statistics, 25(1):168–170, March 1954. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177728859.

Weiss:1954:SPC

[Wei54b] Lionel Weiss. Sequential procedures that control the individual probabilities of coming to the various decisions. Annals of Mathematical Statistics, 25(4):779–784, December 1954. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177728665.

Weiss:1955:NCS

[Wei55a] Lionel Weiss. A note on confidence sets for random variables. Annals of Mathematical Statistics, 26(1):142–144, March 1955. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177728605.

Weiss:1955:CIG

[Wei55b] Lionel Weiss. On confidence intervals of given length for the mean of a normal distribution with unknown variance. Annals of Mathematical Statistics, 26(2):348–352, June 1955. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177728555.

Weiss:1955:SCF

[Wei55c] Lionel Weiss. The stochastic convergence of a function of sample successive differences. Annals of Mathematical Statistics, 26(3):532– 536, September 1955. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177728501.

Weingarten:1956:PLD

[Wei56a] Harry Weingarten. On the probability of large deviations for sums of bounded chance variables. Annals of Mathematical Statistics, 27 (4):1170–1174, December 1956. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177728086.

Weiss:1956:CCS

[Wei56b] Lionel Weiss. A certain class of solutions to a moment problem. Annals of Mathematical Statistics, 27(3):851–853, September 1956. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177728195.

Weiss:1956:CCT

[Wei56c] Lionel Weiss. A certain class of tests of fit. Annals of Mathematical Statistics, 27(4):1165–1170, December 1956. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177728085.

Weiss:1956:UWS

[Wei56d] Lionel Weiss. On the uniqueness of Wald sequential tests. Annals of Mathematical Statistics, 27(4):1178–1181, December 1956. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177728088.

Weingarten:1957:TTC

[Wei57a] Harry Weingarten. Tables for type a critical regions. Annals of Mathematical Statistics, 28(4):1052–1053, December 1957. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177706813.

Weiss:1957:APC

[Wei57b] Lionel Weiss. The asymptotic power of certain tests of fit based on sample spacings. Annals of Mathematical Statistics, 28(3):783–786, September 1957. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177706892.

Weiss:1957:CCF

[Wei57c] Lionel Weiss. The convergence of certain functions of sample spacings. Annals of Mathematical Statistics, 28(3):778–782, September 1957. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177706891.

Weiss:1957:TFP

[Wei57d] Lionel Weiss. Tests of fit in the presence of nuisance location and scale parameters. Annals of Mathematical Statistics, 28(4):1016–1020, December 1957. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177706803.

Weiss:1958:LDS

[Wei58a] Irving Weiss. Limiting distributions in some occupancy problems. Annals of Mathematical Statistics, 29(3):878–884, September 1958. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177706544.

Weiss:1958:LDH

[Wei58b] Lionel Weiss. Limiting distributions of homogeneous functions of sample spacings. Annals of Mathematical Statistics, 29(1):310–312, March 1958. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177706734.

Weiss:1958:SGM

[Wei58c] Lionel Weiss. Spacings generated by mixed samples. Annals of Mathematical Statistics, 29(1):316–318, March 1958. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177706736.

Weiss:1958:TFM

[Wei58d] Lionel Weiss. A test of fit for multivariate distributions. Annals of Mathematical Statistics, 29(2):595–599, June 1958. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177706639.

Weiss:1959:LJD

[Wei59] Lionel Weiss. The limiting joint distribution of the largest and smallest sample spacings. Annals of Mathematical Statistics, 30(2):590– 593, June 1959. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177706274.

Weiss:1960:TST

[Wei60] Lionel Weiss. Two-sample tests for multivariate distributions. Annals of Mathematical Statistics, 31(1):159–164, March 1960. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177705995.

Weiss:1961:TFB

[Wei61] Lionel Weiss. Tests of fit based on the number of observations falling in the shortest sample spacings determined by earlier observations. Annals of Mathematical Statistics, 32(3):838–845, September 1961. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177704977.

Weiner:1965:APA

[Wei65a] H. J. Weiner. Asymptotic properties of an age dependent branching process. Annals of Mathematical Statistics, 36(5):1565–1568, October 1965. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177699917.

Weiner:1965:IEA

[Wei65b] H. J. Weiner. An integral equation in age dependent branching processes. Annals of Mathematical Statistics, 36(5):1569–1573, October 1965. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177699918.

Weiner:1966:MCM

[Wei66] Howard J. Weiner. Monotone convergence of moments in age dependent branching processes. Annals of Mathematical Statistics, 37 (6):1806–1808, December 1966. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177699171.

Weisberg:1971:DLC

[Wei71] Herbert Weisberg. The distribution of linear combinations of order statistics from the uniform distribution. Annals of Mathematical Statistics, 42(2):704–709, April 1971. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177693419.

Weiner:1972:CAD

[Wei72a] Howard J. Weiner. A critical age-dependent branching process with immigration. Annals of Mathematical Statistics, 43(6):2099–2103, December 1972. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177690892.

Weiss:1972:OTS

[Wei72b] N. A. Weiss. The occupation time of a set by countably many recurrent random walks. Annals of Mathematical Statistics, 43(1):293– 302, February 1972. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177692722.

Welch:1939:CLS

[Wel39] B. L. Welch. On confidence limits and sufficiency, with particular reference to parameters of location. Annals of Mathematical Statistics, 10(1):58–69, March 1939. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177732246.

Welch:1947:SSV

[Wel47a] B. L. Welch. On the Studentization of several variances. Annals of Mathematical Statistics, 18(1):118–122, March 1947. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177730499.

Welker:1947:DM

[Wel47b] E. L. Welker. The distribution of the mean. Annals of Mathematical Statistics, 18(1):111–117, March 1947. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177730498.

Welch:1964:RML

[Wel64a] B. L. Welch. Review: M. O. Locks, M. J. Alexander, B. J. Byars, New Tables of the Non-central t Distribution; D. B. Owen, Factors for One-Sided Tolerance Limits and for Variables Sampling Plans. Annals of Mathematical Statistics, 35(2):907–910, June 1964. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177703600.

Welch:1964:PER

[Wel64b] Peter D. Welch. On pre-emptive resume priority queues. Annals of Mathematical Statistics, 35(2):600–612, June 1964. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177703559.

Welsh:1965:RTH

[Wel65] D. J. A. Welsh. Review: Theodore E. Harris, The Theory of Branching Processes. Annals of Mathematical Statistics, 36(2):723–725, April 1965. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177700187.

Welsch:1971:WCT

[Wel71] Roy E. Welsch. A weak convergence theorem for order statistics from strong-mixing processes. Annals of Mathematical Statistics, 42 (5):1637–1646, October 1971. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177693162.

Welsch:1972:LLE

[Wel72] Roy E. Welsch. Limit laws for extreme order statistics from strongmixing processes. Annals of Mathematical Statistics, 43(2):439–446, April 1972. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177692624.

Wendel:1957:GCM

[Wen57] J. G. Wendel. Groups and conditional Monte Carlo. Annals of Mathematical Statistics, 28(4):1048–1052, December 1957. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177706812.

Wendel:1960:OSP

[Wen60] J. G. Wendel. Order statistics of partial sums. Annals of Mathematical Statistics, 31(4):1034–1044, December 1960. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177705676.

Wendel:1961:NAC

[Wen61] J. G. Wendel. The non-absolute convergence of Gil–Pelaez’ inversion integral. Annals of Mathematical Statistics, 32(1):338–339, March 1961. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177705164.

Wendel:1968:RJK

[Wen68] J. G. Wendel. Review: J. Keilson, Green’s Function Methods in Probability Theory. Annals of Mathematical Statistics, 39(5):1769, October 1968. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177698167.

Wertheimer:1932:GEF

[Wer32] Albert Wertheimer. A generalized error function. Annals of Mathematical Statistics, 3(1):64–77, February 1932. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177732934.

Wertheimer:1937:NZP

[Wer37] Albert Wertheimer. Note on Zoch’s paper on the postulate of the arithmetic mean. Annals of Mathematical Statistics, 8(2):112–115, June 1937. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/

1177732431. See [Zoc35a].

Wertheimer:1939:NCI

[Wer39] Albert Wertheimer. A note on confidence intervals and inverse probability. Annals of Mathematical Statistics, 10(1):74–76, March 1939. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177732248.

Wesler:1959:CPI

[Wes59a] Oscar Wesler. A classification problem involving multinomials. Annals of Mathematical Statistics, 30(1):128–133, March 1959. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177706364.

Wesler:1959:ITM

[Wes59b] Oscar Wesler. Invariance theory and a modified minimax principle. Annals of Mathematical Statistics, 30(1):1–20, March 1959. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177706355.

Wine:1957:EDP

[WF57a] R. L. Wine and John E. Freund. On the enumeration of decision patterns involving n means. Annals of Mathematical Statistics, 28(1): 256–259, March 1957. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177707051.

Wold:1957:SER

[WF57b] H. Wold and P. Faxér. On the specification error in regression analysis. Annals of Mathematical Statistics, 28(1):265–267, March 1957.

CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177707054.

Wilk:1964:GMI

[WG64] M. B. Wilk and R. Gnanadesikan. Graphical methods for internal comparisons in multiresponse experiments. Annals of Mathematical Statistics, 35(2):613–631, June 1964. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177703560.

Wilk:1966:SPE

[WGL66] M. B. Wilk, R. Gnanadesikan, and Elizabeth Lauh. Scale parameter estimation from the order statistics of unequal gamma components. Annals of Mathematical Statistics, 37(1):152–176, February 1966. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177699606.

White:1965:CCP

[WH65] David White and Robert A. Hultquist. Construction of confounding plans for mixed factorial designs. Annals of Mathematical Statistics, 36(4):1256–1271, August 1965. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177699997.

Wherry:1931:NFP

[Whe31] R. J. Wherry. A new formula for predicting the shrinkage of the coefficient of multiple correlation. Annals of Mathematical Statistics, 2(4):440–457, November 1931. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177732951.

Wherry:1935:SBS

[Whe35] Robert J. Wherry. The shrinkage of the Brown–Spearman prophecy formula. Annals of Mathematical Statistics, 6(4):183–189, December 1935. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177732563.

Wheeler:1956:VPD

[Whe56] Ruric E. Wheeler. A variable probability distribution function. Annals of Mathematical Statistics, 27(1):196–199, March 1956. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177728358.

Whitney:1951:BES

[Whi51] D. R. Whitney. A bivariate extension of the U statistic. Annals of Mathematical Statistics, 22(2):274–282, June 1951. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177729647.

White:1957:AMS

[Whi57a] John S. White. Approximate moments for the serial correlation coefficient. Annals of Mathematical Statistics, 28(3):798–802, September 1957. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177706896.

White:1957:TSC

[Whi57b] John S. White. A t-test for the serial correlation coefficient. Annals of Mathematical Statistics, 28(4):1046–1048, December 1957. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177706811.

White:1958:APT

[Whi58a] John S. White. Acknowledgment of priority: A t-test for the serial correlation coefficient. Annals of Mathematical Statistics, 29(3):935, September 1958. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177706557.

White:1958:LDS

[Whi58b] John S. White. The limiting distribution of the serial correlation coefficient in the explosive case. Annals of Mathematical Statistics, 29 (4):1188–1197, December 1958. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177706450.

White:1959:LDS

[Whi59] John S. White. The limiting distribution of the serial correlation coefficient in the explosive case II. Annals of Mathematical Statistics, 30(3):831–834, September 1959. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177706213.

White:1961:NO

[Whi61a] John S. White. The nonparametric ordering: (1001) → (0110). Annals of Mathematical Statistics, 32(1):101–103, March 1961. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177705143.

Whittle:1961:SER

[Whi61b] P. Whittle. Some exact results for one-sided distribution tests of the Kolmogorov–Smirnov type. Annals of Mathematical Statistics, 32 (2):499–505, June 1961. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177705056.

White:1965:MDM

[Whi65] Leon S. White. Markovian decision models for the evaluation of a large class of continuous sampling inspection plans. Annals of Mathematical Statistics, 36(5):1408–1420, October 1965. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177699899.

Whitt:1970:WCP

[Whi70] Ward Whitt. Weak convergence of probability measures on the function space C[0, ∞). Annals of Mathematical Statistics, 41(3):939– 944, June 1970. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177696970.

Wicksell:1930:RR

[Wic30] S. D. Wicksell. Remarks on regression. Annals of Mathematical Statistics, 1(1):3–13, February 1930. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177733257.

Wichura:1969:IAW

[Wic69] Michael J. Wichura. Inequalities with applications to the weak convergence of random processes with multi-dimensional time parameters. Annals of Mathematical Statistics, 40(2):681–687, April 1969. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177697741.

Wichura:1970:CAU

[Wic70] Michael J. Wichura. On the construction of almost uniformly convergent random variables with given weakly convergent image laws. Annals of Mathematical Statistics, 41(1):284–291, February 1970. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177697207.

Wichura:1971:NWC

[Wic71] M. J. Wichura. A note on the weak convergence of stochastic processes. Annals of Mathematical Statistics, 42(5):1769–1772, October 1971. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177693181.

Wijsman:1957:ROT

[Wij57] Robert A. Wijsman. Random orthogonal transformations and their use in some classical distribution problems in multivariate analysis. Annals of Mathematical Statistics, 28(2):415–423, June 1957. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177706969. See correc-

tion [Wij59c].

Wijsman:1958:ISS

[Wij58] Robert A. Wijsman. Incomplete sufficient statistics and similar tests. Annals of Mathematical Statistics, 29(4):1028–1045, December 1958. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177706439.

Wijsman:1959:ACR

[Wij59a] Robert A. Wijsman. Applications of a certain representation of the Wishart matrix. Annals of Mathematical Statistics, 30(2):597– 601, June 1959. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177706276.

Wijsman:1959:CNC

[Wij59b] Robert A. Wijsman. Correction notes: Corrections to “On the Theory of BAN Estimates”. Annals of Mathematical Statistics, 30(4): 1268–1270, December 1959. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177706122. See [Wij59d].

Wijsman:1959:CRO

[Wij59c] Robert A. Wijsman. Correction to “Random Orthogonal Transformations and Their Use in Some Classical Distribution Problems in Multivariate Analysis”. Annals of Mathematical Statistics, 30(1): 254, March 1959. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/

1177706387. See [Wij57].

Wijsman:1959:TBE

[Wij59d] Robert A. Wijsman. On the theory of BAN estimates. Annals of Mathematical Statistics, 30(1):185–191, March 1959. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177706373. See correc-

tion [Wij59b].

Wijsman:1960:MPS

[Wij60] Robert A. Wijsman. A monotonicity property of the sequential probability ratio test. Annals of Mathematical Statistics, 31(3):677–684, September 1960. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/

1177705794. See correction [Wij75].

Wijsman:1963:EUM

[Wij63] R. A. Wijsman. Existence, uniqueness and monotonicity of sequential probability ratio tests. Annals of Mathematical Statistics, 34(4):1541– 1548, December 1963. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177703886.

Wijsman:1967:GPT

[Wij67] R. A. Wijsman. General proof of termination with probability one of invariant sequential probability ratio tests based on multivariate normal observations. Annals of Mathematical Statistics, 38(1):8–24, February 1967. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177699054.

Wijsman:1968:BSS

[Wij68a] R. A. Wijsman. Bounds on the sample size distribution for a class of invariant sequential probability ratio tests. Annals of Mathematical Statistics, 39(3):1048–1056, June 1968. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177698337.

Wijsman:1968:RTF

[Wij68b] Robert A. Wijsman. Review: Thomas S. Ferguson, Mathematical Statistics, A Decision Theoretic Approach. Annals of Mathematical Statistics, 39(6):2163–2167, December 1968. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177698055.

Wijsman:1970:CBR

[Wij70] Robert A. Wijsman. Continuity of the Bayes risk. Annals of Mathematical Statistics, 41(3):1083–1085, June 1970. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177696987.

Wijsman:1971:EBS

[Wij71] R. A. Wijsman. Exponentially bounded stopping time of sequential probability ratio tests for composite hypotheses. Annals of Mathematical Statistics, 42(6):1859–1869, December 1971. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177693052.

Wijsman:1972:TOD

[Wij72] R. A. Wijsman. A theorem on obstructive distributions. Annals of Mathematical Statistics, 43(5):1709–1715, October 1972. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177692407.

Wijsman:1975:NCM

[Wij75] Robert A. Wijsman. Notes: Correction to “A Monotonicity Property of the Sequential Probability Ratio Test”. Annals of Statistics, 3(3): 796, May 1975. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.aos/

1176343148. See [Wij60].

Will:1930:FCO

[Wil30] Harry S. Will. On fitting curves to observational series by the method of differences. Annals of Mathematical Statistics, 1(2):159– 190, May 1930. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177733129.

Wilks:1932:MDE

[Wil32a] S. S. Wilks. Moments and distributions of estimates of population parameters from fragmentary samples. Annals of Mathematical Statistics, 3(3):163–195, August 1932. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177732885.

Wilks:1932:SDM

[Wil32b] S. S. Wilks. On the sampling distribution of the multiple correlation coefficient. Annals of Mathematical Statistics, 3(3):196–203, August 1932. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177732886.

Wilks:1935:LTI

[Wil35] S. S. Wilks. The likelihood test of independence in contingency tables. Annals of Mathematical Statistics, 6(4):190–196, December 1935. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177732564.

Will:1936:GSP

[Wil36] Harry Sylvester Will. On a general solution for the parameters of any function with application to the theory of organic growth. Annals of Mathematical Statistics, 7(4):165–190, December 1936. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177732478.

Wilks:1938:FDF

[Wil38a] S. S. Wilks. Fiducial distributions in fiducial inference. Annals of Mathematical Statistics, 9(4):272–280, December 1938. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177732284.

Wilks:1938:LSD

[Wil38b] S. S. Wilks. The large-sample distribution of the likelihood ratio for testing composite hypotheses. Annals of Mathematical Statistics, 9 (1):60–62, March 1938. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177732360.

Wilks:1938:SAC

[Wil38c] S. S. Wilks. Shortest average confidence intervals from large samples. Annals of Mathematical Statistics, 9(3):166–175, September 1938. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177732308.

Wilks:1941:DSS

[Wil41a] S. S. Wilks. Determination of sample sizes for setting tolerance limits. Annals of Mathematical Statistics, 12(1):91–96, March 1941. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177731788.

Williams:1941:MRM

[Wil41b] J. D. Williams. Moments of the ratio of the mean square successive difference to the mean square difference in samples from a normal universe. Annals of Mathematical Statistics, 12(2):239–241, June 1941. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177731756.

Wilks:1942:SPS

[Wil42] S. S. Wilks. Statistical prediction with special reference to the problem of tolerance limits. Annals of Mathematical Statistics, 13(4):400– 409, December 1942. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177731537.

Wilkins:1944:NSK

[Wil44] J. Ernest Wilkins, Jr. A note on skewness and kurtosis. Annals of Mathematical Statistics, 15(3):333–335, September 1944. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177731243.

Wilks:1946:SCT

[Wil46a] S. S. Wilks. Sample criteria for testing equality of means, equality of variances, and equality of covariances in a normal multivariate distribution. Annals of Mathematical Statistics, 17(3):257–281, September 1946. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177730940.

Williams:1946:API

[Wil46b] J. D. Williams. An approximation to the probability integral. Annals of Mathematical Statistics, 17(3):363–365, September 1946. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177730951.

Wilks:1947:ARE

[Wil47] S. S. Wilks. Annual report of the Editor for 1946. Annals of Mathematical Statistics, 18(1):159, March 1947. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177730510.

Wilks:1948:RE

[Wil48] S. S. Wilks. Report of the Editor for 1947. Annals of Mathematical Statistics, 19(1):144, March 1948. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177730309.

Wilks:1949:RE

[Wil49] S. S. Wilks. Report of the Editor for 1948. Annals of Mathematical Statistics, 20(1):163–164, March 1949. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177730113.

Wilks:1950:REB

[Wil50] S. S. Wilks. Report of the Editor of the Annals for 1949. Annals of Mathematical Statistics, 21(1):163–164, March 1950. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177729904.

Williams:1965:ETT

[Wil65] Britain J. Williams. The effect of truncation on tests of hypotheses for normal populations. Annals of Mathematical Statistics, 36(5):1504– 1510, October 1965. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177699909.

Wilson:1967:AST

[Wil67] A. L. Wilson. An approach to simultaneous tolerance intervals in regression. Annals of Mathematical Statistics, 38(5):1536–1540, October 1967. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177698707.

Williams:1969:CDF

[Wil69] E. J. Williams. Cauchy-distributed functions and a characterization of the Cauchy distribution. Annals of Mathematical Statistics, 40(3): 1083–1085, June 1969. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177697613.

Williamson:1970:FW

[Wil70] John A. Williamson. Fluctuations when E(|X_{1}|) = ∞. Annals of Mathematical Statistics, 41(3):865–875, June 1970. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177696964.

Wintner:1947:SAC

[Win47] Aurel Wintner. On the shape of the angular case of Cauchy’s distribution curves. Annals of Mathematical Statistics, 18(4):589–593, December 1947. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177730351.

Wind:1972:SJE

[Win72] Serge L. Wind. Stein–James estimators of a multivariate location parameter. Annals of Mathematical Statistics, 43(1):340–343, February 1972. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177692727.

Wisniewski:1937:PLS

[Wis37] Jan K. Wisniewski. A problem in least squares. Annals of Mathematical Statistics, 8(3):145–148, September 1937. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177732410.

Wishart:1956:QSS

[Wis56] David M. G. Wishart. A queueing system with χ^{2} service-time distribution. Annals of Mathematical Statistics, 27(3):768–779, September 1956. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177728183.

Wishart:1957:AFC

[Wis57] John Wishart. An approximate formula for the cumulative zdistribution. Annals of Mathematical Statistics, 28(2):504–510, June 1957. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177706980.

Witting:1960:GPE

[Wit60] H. Witting. A generalized Pitman efficiency for nonparametric tests. Annals of Mathematical Statistics, 31(2):405–414, June 1960. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177705903.

Wilk:1956:SAA

[WK56] M. B. Wilk and O. Kempthorne. Some aspects of the analysis of factorial experiments in a completely randomized design. Annals of Mathematical Statistics, 27(4):950–985, December 1956. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177728068.

Woodall:1962:EOC

[WK62] Rosalie C. Woodall and Badrig M. Kurkjian. Exact operating characteristic for truncated sequential life tests in the exponential case. Annals of Mathematical Statistics, 33(4):1403–1412, December 1962. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177704372.

Kim:1968:GIT

[wK68] Choo whan Kim. A generalization of Itô’s theorem concerning the pointwise ergodic theorem. Annals of Mathematical Statistics, 39(6): 2145–2148, December 1968. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177698047.

Woinsky:1969:SNT

[WK69] Melvin N. Woinsky and Ludwik Kurz. Sequential nonparametric twoway classification with a prescribed maximum asymptotic error probability. Annals of Mathematical Statistics, 40(2):445–455, April 1969. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177697709.

Wani:1970:NCI

[WK70] J. K. Wani and D. G. Kabe. Note on a characterization of the inverse Gaussian distribution. Annals of Mathematical Statistics, 41(3):1071– 1073, June 1970. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177696984.

Warde:1971:IDD

[WK71] W. D. Warde and S. K. Katti. Infinite divisibility of discrete distributions, II. Annals of Mathematical Statistics, 42(3):1088–1090, June 1971. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177693337.

Watson:1963:EPD

[WL63] G. S. Watson and M. R. Leadbetter. On the estimation of the probability density, I. Annals of Mathematical Statistics, 34(2):480– 491, June 1963. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177704159.

Woinsky:1972:CNT

[Woi72] Melvin N. Woinsky. A composite nonparametric test for a scale slippage alternative. Annals of Mathematical Statistics, 43(1):65–73, February 1972. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177692701.

Wolfowitz:1942:APF

[Wol42] J. Wolfowitz. Additive partition functions and a class of statistical hypotheses. Annals of Mathematical Statistics, 13(3):247–279, September 1942. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177731566.

Wolfowitz:1943:TRS

[Wol43] J. Wolfowitz. On the theory of runs with some applications to quality control. Annals of Mathematical Statistics, 14(3):280–288, September 1943. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177731421.

Wolfowitz:1944:ADR

[Wol44a] J. Wolfowitz. Asymptotic distribution of runs up and down. Annals of Mathematical Statistics, 15(2):163–172, June 1944. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177731281;   http:
//www.jstor.org/stable/2236196.

Wolfowitz:1944:NRC

[Wol44b] J. Wolfowitz. Note on runs of consecutive elements. Annals of Mathematical Statistics, 15(1):97–98, March 1944. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177731319.

Wolfowitz:1946:CLF

[Wol46a] J. Wolfowitz. Confidence limits for the fraction of a normal population which lies between two given limits. Annals of Mathematical Statistics, 17(4):483–488, December 1946. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177730887.

Wolfowitz:1946:SBE

[Wol46b] J. Wolfowitz. On sequential binomial estimation. Annals of Mathematical Statistics, 17(4):489–493, December 1946. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177730888.

Wolfowitz:1947:CSB

[Wol47a] J. Wolfowitz. Consistency of sequential binomial estimates. Annals of Mathematical Statistics, 18(1):131–135, March 1947. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177730502.

Wolfowitz:1947:ESE

[Wol47b] J. Wolfowitz. The efficiency of sequential estimates and Wald’s equation for sequential processes. Annals of Mathematical Statistics, 18 (2):215–230, June 1947. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177730439.

Wold:1948:PST

[Wol48] Herman O. A. Wold. On prediction in stationary time series. Annals of Mathematical Statistics, 19(4):558–567, December 1948. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177730151.

Wolfowitz:1949:WPC

[Wol49a] J. Wolfowitz. On Wald’s proof of the consistency of the maximum likelihood estimate. Annals of Mathematical Statistics, 20(4):601– 602, December 1949. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177729953.

Wolfowitz:1949:PCT

[Wol49b] J. Wolfowitz. The power of the classical tests associated with the normal distribution. Annals of Mathematical Statistics, 20(4):540– 551, December 1949. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177729946.

Wolfowitz:1950:MEM

[Wol50] J. Wolfowitz. Minimax estimates of the mean of a normal distribution with known variance. Annals of Mathematical Statistics, 21(2):218– 230, June 1950. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177729840.

Wolfowitz:1951:CCD

[Wol51] J. Wolfowitz. On ε-complete classes of decision functions. Annals of Mathematical Statistics, 22(3):461–465, September 1951. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177729595.

Wolfowitz:1952:AW

[Wol52a] J. Wolfowitz. Abraham Wald, 1902–1950. Annals of Mathematical Statistics, 23(1):1–13, March 1952. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177729480.

Wolfowitz:1952:SAM

[Wol52b] J. Wolfowitz. On the stochastic approximation method of Robbins and Monro. Annals of Mathematical Statistics, 23(3):457–461, September 1952. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177729391.

Wolfowitz:1954:EMD

[Wol54a] J. Wolfowitz. Estimation by the minimum distance method in nonparametric stochastic difference equations. Annals of Mathematical Statistics, 25(2):203–217, June 1954. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177728782.

Wolfowitz:1954:GTG

[Wol54b] J. Wolfowitz. Generalization of the theorem of Glivenko–Cantelli. Annals of Mathematical Statistics, 25(1):131–138, March 1954. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177728852.

Wolfowitz:1956:SAM

[Wol56] J. Wolfowitz. On stochastic approximation methods. Annals of Mathematical Statistics, 27(4):1151–1156, December 1956. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177728082.

Wolfowitz:1957:MDM

[Wol57] J. Wolfowitz. The minimum distance method. Annals of Mathematical Statistics, 28(1):75–88, March 1957. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177707038.

Wolfowitz:1958:ITM

[Wol58] J. Wolfowitz. Information theory for mathematicians. Annals of Mathematical Statistics, 29(2):351–356, June 1958. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177706617.

Wold:1961:OA

[Wol61] Herman Wold. Oskar Anderson, 1887–1960. Annals of Mathematical Statistics, 32(3):651–660, September 1961. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177704961.

Wolfowitz:1966:ROC

[Wol66] J. Wolfowitz. Remark on the optimum character of the sequential probability ratio test. Annals of Mathematical Statistics, 37(3):726– 727, June 1966. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177699468.

Wold:1967:RCW

[Wol67a] H. O. Wold. Review: C. W. J. Granger, M. Hatanaka, Spectral Analysis of Economic Time Series. Annals of Mathematical Statistics, 38(1):288–293, February 1967. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177699088.

Wolfowitz:1967:RRA

[Wol67b] J. Wolfowitz. Review: Robert Ash, Information Theory. Annals of Mathematical Statistics, 38(2):627, April 1967. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177698990.

Wolfe:1971:IMF

[Wol71a] Stephen James Wolfe. Inequalities for modes of L functions. Annals of Mathematical Statistics, 42(6):2126–2130, December 1971. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177693078.

Wolfe:1971:MID

[Wol71b] Stephen James Wolfe. On moments of infinitely divisible distribution functions. Annals of Mathematical Statistics, 42(6):2036–2043, December 1971. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177693071.

Wolfe:1971:CPF

[Wol71c] Stephen James Wolfe. On the continuity properties of L functions. Annals of Mathematical Statistics, 42(6):2064–2073, December 1971. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177693075.

Wolfe:1971:UF

[Wol71d] Stephen James Wolfe. On the unimodality of L functions. Annals of Mathematical Statistics, 42(3):912–918, June 1971. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177693320.

Wolfe:1972:NCC

[Wol72] Stephen James Wolfe. A note on the complete convergence of stable distribution functions. Annals of Mathematical Statistics, 43(1):363– 364, February 1972. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177692733.

Wong:1935:AOP

[Won35] Y. K. Wong. An application of orthogonalization process to the theory of least squares. Annals of Mathematical Statistics, 6(2):53– 75, June 1935. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177732609.

Wong:1936:SEL

[Won36] Y. K. Wong. On standard error for the line of mutual regression. Annals of Mathematical Statistics, 7(1):47–50, March 1936. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177732545.

Wong:1968:AOP

[Won68] Seok Pin Wong. Asymptotically optimum properties of certain sequential tests. Annals of Mathematical Statistics, 39(4):1244–1263, August 1968. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177698250.

Wong:1969:HGM

[Won69] E. Wong. Homogeneous Gauss–Markov random fields. Annals of Mathematical Statistics, 40(5):1625–1634, October 1969. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177697378.

Woodbury:1940:RCW

[Woo40] Max A. Woodbury. Rank correlation when there are equal variates. Annals of Mathematical Statistics, 11(3):358–362, September 1940. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177731875.

Woodbury:1949:PD

[Woo49] Max A. Woodbury. On a probability distribution. Annals of Mathematical Statistics, 20(2):311–313, June 1949. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177730043.

Woodroofe:1966:SPN

[Woo66] Michael Woodroofe. Statistical properties of the number of positive sums. Annals of Mathematical Statistics, 37(5):1295–1304, October 1966. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177699273.

Woodroofe:1967:MDSa

[Woo67] Michael Woodroofe. On the maximum deviation of the sample density. Annals of Mathematical Statistics, 38(2):475–481, April 1967. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177698963.

Wood:1970:NRP

[Woo70a] Richard G. Wood. A note on random polynomials. Annals of Mathematical Statistics, 41(2):680–682, April 1970. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177697113.

Woodroofe:1970:CDS

[Woo70b] Michael Woodroofe. On choosing a delta-sequence. Annals of Mathematical Statistics, 41(5):1665–1671, October 1970. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177696810.

Woodroofe:1970:FT

[Woo70c] Michael Woodroofe. On the first time |S_{n}| ≥ cn^{\frac{1}{2}} . Annals of Mathematical Statistics, 41(6):2179–2183, December 1970. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177696727.

Woodworth:1970:LDB

[Woo70d] George G. Woodworth. Large deviations and Bahadur efficiency of linear rank statistics. Annals of Mathematical Statistics, 41(1):251– 283, February 1970. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177697206.

Woodroofe:1972:MLE

[Woo72] Michael Woodroofe. Maximum likelihood estimation of a translation parameter of a truncated distribution. Annals of Mathematical Statistics, 43(1):113–122, February 1972. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.

org/euclid.aoms/1177692707.

Wormleighton:1959:STP

[Wor59] R. Wormleighton. Some tests of permutation symmetry. Annals of Mathematical Statistics, 30(4):1005–1017, December 1959. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177706083.

Wormleighton:1960:UGS

[Wor60] R. Wormleighton. A useful generalization of the Stein twosample procedure. Annals of Mathematical Statistics, 31(1):217–221, March 1960. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177705998.

Wright:1934:MPC

[Wri34] Sewall Wright. The method of path coefficients. Annals of Mathematical Statistics, 5(3):161–215, September 1934. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177732676.

Wright:1972:RCW

[Wri72] F. T. Wright. Rates of convergence for weighted sums of random variables. Annals of Mathematical Statistics, 43(5):1687–1691, October 1972. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177692403.

Walker:1934:ACA

[WS34] Helen M. Walker and Vera Sanford. The accuracy of computation with approximate numbers. Annals of Mathematical Statistics, 5(1): 1–12, March 1934. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177732731.

Woodroofe:1967:MDSb

[WV67] Michael B. Woodroofe and John W. Van Ness. The maximum deviation of sample spectral densities. Annals of Mathematical Statistics, 38(5):1558–1569, October 1967. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177698710.

Wald:1939:CLC

[WW39] A. Wald and J. Wolfowitz. Confidence limits for continuous distribution functions. Annals of Mathematical Statistics, 10(2):105– 118, June 1939. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177732209.

Wald:1940:TWT

[WW40] A. Wald and J. Wolfowitz. On a test whether two samples are from the same population. Annals of Mathematical Statistics, 11(2):147– 162, June 1940. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177731909.

Wald:1941:NCL

[WW41] A. Wald and J. Wolfowitz. Note on confidence limits for continuous distribution functions. Annals of Mathematical Statistics, 12(1):118– 119, March 1941. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177731793.

Wald:1943:ETR

[WW43] A. Wald and J. Wolfowitz. An exact test for randomness in the non-parametric case, based on serial correlation. Annals of Mathematical Statistics, 14(4):378–388, December 1943. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http:/

/projecteuclid.org/euclid.aoms/1177731358;   http://www.
jstor.org/stable/2235925.

Wald:1944:STB

[WW44] A. Wald and J. Wolfowitz. Statistical tests based on permutations of the observations. Annals of Mathematical Statistics, 15(4):358– 372, December 1944. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177731207.

Wald:1945:SIP

[WW45] A. Wald and J. Wolfowitz. Sampling inspection plans for continuous production which insure a prescribed limit on the outgoing quality. Annals of Mathematical Statistics, 16(1):30–49, March 1945. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177731169.

Wald:1946:TLN

[WW46] A. Wald and J. Wolfowitz. Tolerance limits for a normal distribution. Annals of Mathematical Statistics, 17(2):208–215, June 1946. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177730981.

Wald:1948:OCS

[WW48] A. Wald and J. Wolfowitz. Optimum character of the sequential probability ratio test. Annals of Mathematical Statistics, 19(3):326– 339, September 1948. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177730197.

Wald:1950:BSS

[WW50] A. Wald and J. Wolfowitz. Bayes solutions of sequential decision problems. Annals of Mathematical Statistics, 21(1):82–99, March 1950. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177729887.

Wyner:1969:ADC

[Wyn69] A. D. Wyner. On the asymptotic distribution of a certain functional of the Wiener process. Annals of Mathematical Statistics, 40(4):1409– 1418, August 1969. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177697512.

Wynn:1970:SGO

[Wyn70] Henry P. Wynn. The sequential generation of D-optimum experimental designs. Annals of Mathematical Statistics, 41(5):1655–1664, October 1970. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177696809.

Wong:1965:COI

[WZ65] Eugene Wong and Moshe Zakai. On the convergence of ordinary integrals to stochastic integrals. Annals of Mathematical Statistics, 36(5):1560–1564, October 1965. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177699916.

Yackel:1968:RTC

[Yac68] James Yackel. A random time change relating semi-Markov and Markov processes. Annals of Mathematical Statistics, 39(2):358– 364, April 1968. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177698396.

Yahav:1966:OS

[Yah66] Joseph A. Yahav. On optimal stopping. Annals of Mathematical Statistics, 37(1):30–35, February 1966. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177699595.

Yakowitz:1969:CEI

[Yak69] S. Yakowitz. A consistent estimator for the identification of finite mixtures. Annals of Mathematical Statistics, 40(5):1728–1735, October 1969. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177697384.

Yalavigi:1968:SBI

[Yal68] C. C. Yalavigi. A series of balanced incomplete block designs. Annals of Mathematical Statistics, 39(2):681–683, April 1968. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177698428.

Yalavigi:1971:STD

[Yal71] C. C. Yalavigi. A series of tournamental designs. Annals of Mathematical Statistics, 42(1):395–397, February 1971. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177693531.

Yang:1968:CSM

[Yan68] Grace Lo Yang. Contagion in stochastic models for epidemics. Annals of Mathematical Statistics, 39(6):1863–1889, December 1968. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177698018.

Yang:1971:IMS

[Yan71] Y. S. Yang. Invariant measures on some Markov processes. Annals of Mathematical Statistics, 42(5):1686–1696, October 1971. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177693167.

Yarbrough:1971:SDL

[Yar71] Charles Yarbrough. Sequential discrimination with likelihood ratios. Annals of Mathematical Statistics, 42(4):1339–1347, August 1971. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177693246.

Yarnold:1972:AAP

[Yar72] James K. Yarnold. Asymptotic approximations for the probability that a sum of lattice random vectors lies in a convex set. Annals of Mathematical Statistics, 43(5):1566–1580, October 1972. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177692389.

Yen:1964:TSN

[Yen64] Elizabeth H. Yen. On two-stage non-parametric estimation. Annals of Mathematical Statistics, 35(3):1099–1114, September 1964. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177703268.

Ylvisaker:1961:GTB

[Ylv61] N. Donald Ylvisaker. A generalization of a theorem of Balakrishnan. Annals of Mathematical Statistics, 32(4):1337–1339, December 1961.

CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177704872.

Ylvisaker:1962:LER

[Ylv62] N. Donald Ylvisaker. On linear estimation for regression problems on time series. Annals of Mathematical Statistics, 33(3):1077–1084, September 1962. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177704473.

Ylvisaker:1964:LBM

[Ylv64] N. Donald Ylvisaker. Lower bounds for minimum covariance matrices in time series regression problems. Annals of Mathematical Statistics, 35(1):362–368, March 1964. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177703759.

Ylvisaker:1965:ENZ

[Ylv65] N. Donald Ylvisaker. The expected number of zeros of a stationary Gaussian process. Annals of Mathematical Statistics, 36(3):1043– 1046, June 1965. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177700077.

Ylvisaker:1966:TCL

[Ylv66] N. Donald Ylvisaker. On a theorem of Cramér and Leadbetter. Annals of Mathematical Statistics, 37(3):682–685, June 1966. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177699461.

Ylvisaker:1968:NAT

[Ylv68a] Donald Ylvisaker. A note on the absence of tangencies in Gaussian sample paths. Annals of Mathematical Statistics, 39(1):261–262, February 1968. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177698528.

Ylvisaker:1968:RHC

[Ylv68b] Donald Ylvisaker. Review: Harald Cramér, M. R. Leadbetter, Stationary and Related Stochastic Processes/Sampling Function Properties and Their Applications. Annals of Mathematical Statistics, 39(1): 290–291, February 1968. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177698538.

Yosida:1949:BMS

[Yos49] Kôsaku Yosida. Brownian motion on the surface of the 3-sphere. Annals of Mathematical Statistics, 20(2):292–296, June 1949. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177730038.

Young:1941:ROS

[You41] L. C. Young. On randomness in ordered sequences. Annals of Mathematical Statistics, 12(3):293–300, September 1941. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http:/

/projecteuclid.org/euclid.aoms/1177731711;   http://www.
jstor.org/stable/2235858.

Yakowitz:1968:IFM

[YS68] Sidney J. Yakowitz and John D. Spragins. On the identifiability of finite mixtures. Annals of Mathematical Statistics, 39(1):209–214, February 1968. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177698520.

Yu:1971:PEK

[Yu71] C. S. Yu. Pitman efficiencies of Kolmogorov–Smirnov tests. Annals of Mathematical Statistics, 42(5):1595–1605, October 1971. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177693158.

Yuan:1933:LFD

[Yua33] Pae-Tsi Yuan. On the logarithmic frequency distribution and the semi-logarithmic correlation surface. Annals of Mathematical Statistics, 4(1):30–74, February 1933. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177732821.

Zacks:1963:CCL

[Zac63] S. Zacks. On a complete class of linear unbiased estimators for randomized factorial experiments. Annals of Mathematical Statistics, 34 (3):769–779, September 1963. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177704002.

Zacks:1964:GLS

[Zac64] S. Zacks. Generalized least squares estimators for randomized fractional replication designs. Annals of Mathematical Statistics, 35(2):

696–704, June 1964. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177703566.

Zacks:1966:RFW

[Zac66a] S. Zacks. Randomized fractional weighing designs. Annals of Mathematical Statistics, 37(5):1382–1395, October 1966. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177699282.

Zacks:1966:SEM

[Zac66b] S. Zacks. Sequential estimation of the mean of a log-normal distribution having a prescribed proportional closeness. Annals of Mathematical Statistics, 37(6):1688–1696, December 1966. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177699158.

Zacks:1967:NEF

[Zac67] S. Zacks. On the non-existence of a fixed sample estimator of the mean of a log-normal distribution having a prescribed proportional closeness. Annals of Mathematical Statistics, 38(3):949, June 1967. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177698894.

Zacks:1968:BSD

[Zac68] S. Zacks. Bayes sequential design of fractional factorial experiments for the estimation of a subgroup of pre-assigned parameters. Annals of Mathematical Statistics, 39(3):973–982, June 1968. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177698330.

Zacks:1970:BFE

[Zac70a] S. Zacks. Bayes and fiducial equivariant estimators of the common mean of two normal distributions. Annals of Mathematical Statistics, 41(1):59–69, February 1970. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177697188.

Zacks:1970:RSJ

[Zac70b] S. Zacks. Review: Sidney J. Yakowitz, Mathematics of Adaptive Control Processes. Annals of Mathematical Statistics, 41(3):1128– 1131, June 1970. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177696998.

Zaremba:1965:NWM

[Zar65] S. K. Zaremba. Note on the Wilcoxon–Mann–Whitney statistic. Annals of Mathematical Statistics, 36(3):1058–1060, June 1965. CO-DEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177700082.

Ziaud-Din:1938:DOD

[ZD38] M. Ziaud-Din. On differential operator developed by o’toole. Annals of Mathematical Statistics, 9(1):63–65, March 1938. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177732361.

Ziaud-Din:1959:EST

[ZD59] M. Ziaud-Din. The expression of k-statistic k_{11} in terms of power sums and sample moments. Annals of Mathematical Statistics, 30 (3):825–828, September 1959. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177706211.

Zehna:1966:IML

[Zeh66] Peter W. Zehna. Invariance of maximum likelihood estimators. Annals of Mathematical Statistics, 37(3):744, June 1966. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177699475.

Zeigler:1950:NAS

[Zei50] R. K. Zeigler. A note on the asymptotic simultaneous distribution of the sample median and the mean deviation from the sample median. Annals of Mathematical Statistics, 21(3):452–455, September 1950. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177729806.

Zelen:1954:NPB

[Zel54] Marvin Zelen. A note on partially balanced designs. Annals of Mathematical Statistics, 25(3):599–602, September 1954. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177728729.

Zelen:1958:UGD

[Zel58] Marvin Zelen. The use of group divisible designs for confounded asymmetrical factorial arrangements. Annals of Mathematical Statistics, 29(1):22–40, March 1958. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177706704.

Zelen:1964:ACF

[ZF64] M. Zelen and W. Federer. Applications of the calculus for factorial arrangements II: Two way elimination of heterogeneity. Annals of Mathematical Statistics, 35(2):658–672, June 1964. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177703563.

Zidek:1969:IBI

[Zid69] J. V. Zidek. Inadmissibility of the best invariant estimator of extreme quantiles of the normal law under squared error loss. Annals of Mathematical Statistics, 40(5):1801–1808, October 1969. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177697393.

Zidek:1970:SCA

[Zid70] J. V. Zidek. Sufficient conditions for the admissibility under squared errors loss of formal Bayes estimators. Annals of Mathematical Statistics, 41(2):446–456, April 1970. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177697084.

Zidek:1971:ICE

[Zid71] James V. Zidek. Inadmissibility of a class of estimators of a normal quantile. Annals of Mathematical Statistics, 42(4):1444–1447, August 1971. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177693258.

Zimmerman:1972:SSF

[Zim72] Grenith J. Zimmerman. Some sample function properties of the twoparameter Gaussian process. Annals of Mathematical Statistics, 43 (4):1235–1246, August 1972. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/

euclid.aoms/1177692475.

Zelen:1959:WCT

[ZJ59] M. Zelen and L. S. Joel. The weighted compounding of two independent significance tests. Annals of Mathematical Statistics, 30(4):885– 895, December 1959. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.
aoms/1177706074.

Zoch:1934:ICC

[Zoc34] Richmond T. Zoch. Invariants and covariants of certain frequency curves. Annals of Mathematical Statistics, 5(2):124–135, June 1934. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177732711.

Zoch:1935:PAM

[Zoc35a] Richmond T. Zoch. On the postulate of the arithmetic mean. Annals of Mathematical Statistics, 6(4):171–182, December 1935. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177732562. See note

[Wer37].

Zoch:1935:SIF

[Zoc35b] Richmond T. Zoch. Some interesting features of frequency curves. Annals of Mathematical Statistics, 6(1):1–10, March 1935. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177732636.

Zoch:1937:RMW

[Zoc37] Richmond T. Zoch. Reply to Mr. Wertheimer’s paper. Annals of Mathematical Statistics, 8(3):177–178, September 1937. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177732413.

Zelen:1960:GBN

[ZS60] Marvin Zelen and Norman C. Severo. Graphs for bivariate normal probabilities. Annals of Mathematical Statistics, 31(3):619–624, September 1960. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.aoms/
1177705789.

Zygmund:1947:RCF

[Zyg47] A. Zygmund. A remark on characteristic functions. Annals of Mathematical Statistics, 18(2):272–276, June 1947. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177730443.

Zyskind:1963:SCR

[Zys63] George Zyskind. Some consequences of randomization in a generalization of the balanced incomplete block design. Annals of Mathematical Statistics, 34(4):1569–1581, December 1963. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177703889.

Zyskind:1967:CFN

[Zys67] George Zyskind. On canonical forms, non-negative covariance matrices and best and simple least squares linear estimators in linear models. Annals of Mathematical Statistics, 38(4):1092–1109, August 1967. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177698779.